A Complete Bibliography of Publications in Marine Policy (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

10 April 2024 Version 1.41

Title word cross-reference

30 × 30 [LMC^{+}23b]. _{2} [EV22, GZW^{+}20, MVSP^{+}21, OOGK^{+}24, SSK23]. Q [POM24].

[MWM^{+}23b].

İskenderun [CA21].

0 [Tan23a].

107 [GZW^{+}20]. 11 [GA20]. 115 [EHM^{+}22, Mot22]. 124 [SDG^{+}22]. 126 [TGM22]. 14 [Gul20]. 14.7.1 [Liu23]. 143 [SCSR^{+}23]. 151 [RP23a]. 160 [Smi23b]. 19 [AEL^{+}23, ASGZ22, ABME22, BRM^{+}21a, BSI^{+}21, BASM22, CDNS23, CCKK22, CRCP^{+}23, De 21, DME23, Ero22, FTM^{+}22, FGPPGG22, GPTASM23, GABSD21, GNMGWM21, GMCMJA^{+}21, GS21, HDCH21, HHB^{+}21, HBP^{+}21, Isr22, KFP20, KAC22, Kob22, LWSS21, LSB^{+}21, LC20b, LEERRM^{+}21, LKE^{+}22, MCBN22, MÇ23, MAK22, MMK^{+}22, NK21, NGRM^{+}22, NVP^{+}23, NAS^{+}22, PGK22, PAK22, PBD^{+}23, RDD^{+}21, SVV24, SSB^{+}22, SCH^{+}24, SdCMN^{+}22, SCLM^{+}21, SSP^{+}21a, SNK^{+}20, SRH^{+}22, SSP^{+}21b, Tan22, TAK^{+}22, VMM21, VDP^{+}22, WHL22, WPWM^{+}22, XSCL21, YCBW21, ZTZ^{+}21, ZTW23, ZTM^{+}23, ZT23a, ZCS^{+}23]. 1960s [GMVM23].

20-year [VBd^{+}21]. 2001 [LC23]. 2005 [MK24]. 2006 [CK23]. 2018 [ZPX^{+}22]. 2020 [Ano20a, Ano20b, Ano20n, Ano20o, Ano20q, Ano20p, Ano20r, Ano20s, Ano20t, Ano20u, Ano20v, Mar21a]. 2021 [Ano21a, Ano21b, Ano21c, Ano21q, Ano21r, Ano21t, Ano21s, Ano21u, Ano21v, Ano21w, Ano21x, Ano21y, SWGZ22]. 2022 [ALLS23, Ano22a, Ano22b, Ano22c, Ano22p, Ano22q, Ano22s, Ano22r, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x]. 2023 [Ano23a, Ano23b, Ano23p, Ano23s, Ano23r, Ano23t, Ano23u, Ano23v, Ano23w, Ano23x, BÖ24]. 2024 [Ano24a, Ano24g, Ano24h]. 21st [PZPB21, SF22, CTH^{+}23]. 2nd [Tan23b].

3 [Cha20b, Lv20]. 305 [Smi22]. 34 [SFGEPC22].

4 [Tan23b]. 450 [Smi20]. 48.3 [CHC^{+}21]. 49.99 [Rub24].

5.1 [ALLS23].

6 [Niu22]. 65-day [BSI^{+}21].

7 [Ngu23].

82 [Ege23].

9 [Mar21a]. 97 [Mal21]. 978 [Cha20b, Lv20, Ngu23, Niu22, Tan23a, Tan23b]. 978-1-108-84226-6 [Niu22]. 978-1-138-38814-7 [Ngu23]. 978-1-316-51742-0 [Tan23a]. 978-1-78643-673-3 [Cha20b, Lv20]. 978-1-78990-907-4 [Tan23b]. 978-3-031-10126-7 [Rub24]. 978-3-031-20739-6 [Sta23a]. 978-3-031-20740-2 [Sta23a]. 978-3-031-20742-6 [Sta23a]. 978-90-823810-4-7 [HH21].

A.-K [Sta23a]. Aarhus [RPY^{+}24]. abalone [HCSBM^{+}23]. Abandoned [GRS^{+}23, GHW^{+}22, GAD^{+}23, HGG^{+}22a, Lov23, Sam22]. abatement [NNR^{+}21, SBW^{+}21]. abating [Osm23]. Aberdeen [BHCL23]. ability [FWL^{+}22, GCV^{+}21, Mag23]. ABMTs [TCZ21]. aboard [MLE^{+}24]. abundance [DT23, GSAF21, MJ20, SGPF21]. abuses [WS23]. ACAP [BKW^{+}24]. Accelerating [BDM21, Cha22]. acceptability [EV22, HHEJB20, SMC^{+}21]. acceptable [HWT23]. acceptance [KNY20, LYQM21]. Access [BBS^{+}21, GBLBV20, MKS21, Art20, DTBP21, FHBS23, KL21b, LP20, NM23, OA20, PFG^{+}21, PGK22, SPFA^{+}20, SCF^{+}22, AGK22]. accessible [LHV^{+}20]. Accidents [Cha20b, FFE20, FFE^{+}24, KLK23, RE23, YHU22]. accomplish [ALLS23]. Accord [HvdBH23]. according [dSJNS20, KS22]. accountability [BDC^{+}21]. accounting [CVH^{+}20, DJC^{+}23, GPA^{+}22, GGP^{+}22, MMNC22, PMN^{+}22]. accounts [CNG^{+}23, GMG^{+}23, NMCL24]. accumulated [vdBSvdV^{+}23]. accuracy [DQJ23, dSOW^{+}21]. accurate [ZLD^{+}21]. achievable [HRC^{+}24]. achieve [CMOB^{+}20, TFH^{+}23]. achievement [PWL^{+}23a]. Achievements [GJS22, CCX23, Gul20, Tan24, Fis22]. Achieving [AAA^{+}24, BPA^{+}22, MMSB^{+}20, PDB^{+}23, Puh23, GKP^{+}22, HvPPV20b, Kob22, OKNB22, OGS24, DUYEA21]. Achilles [BD23]. acidification [AHN^{+}23, GBC^{+}20]. acknowledge [VBd^{+}21]. acoupa [dSJNS20, dSJNS20]. acoustic [KDM^{+}20, MGBS24, WCF^{+}21]. across [BDC23, DDAB^{+}23, DPG^{+}22, FRVB22, GSAF21, HHEJB20, KBD^{+}20, LJ20, Lot22, MWM^{+}23a, OPL20, SHPW21, SHP^{+}21, SOLN20, TOHK20, XWH^{+}22, YT20]. Act [BGK^{+}22, WMKG21, BMR23, Squ21]. Actiniaria [YNZ22]. Action [FS22, MPSF20, AHB^{+}23, BGK^{+}22, BCB20, KK22, LA23, MWM^{+}23a, Moh23, OGGC20, PN20, PKB^{+}23, WW21a, WWD23, Zhu22, vdBSvdV^{+}23]. actionable [SB20]. actions [BDM21, Li23a, MB22, PCG^{+}23, WYH21]. Activating [Men23]. active [CCK^{+}21, DXS^{+}21, GWAPRPRU22, JG20, SKL^{+}20]. activities [ABI24, CWR20, DNR21, Fei21, GCL21, Gug22, HHBC20, Kim23, LCL20, LWX^{+}20, OAAMS24, Puh23, QWBY21, SOT^{+}22, SV22, SSMG21, SHAC23, TLT^{+}22, YG20, YSEK22]. activity [CLS^{+}23, KDM^{+}20, NAN^{+}24, Okş23, SGB21, SVP^{+}23, War21, Yu22]. activity-based [Okş23]. actor [QRK^{+}23]. actor-driven [QRK^{+}23]. actors [BPA^{+}22, BE20b, CDNS23, CC23, SFB24, YWYL22, KWMM20a, MAMCHP^{+}20]. Adaptation [CDNS23, BOOO23, BG20b, Bir20, BB20, BvdH23, CBAPM20, DP23, GOH^{+}20, JTR^{+}20, MKM23, O’D21, OCD^{+}20, OGGC20, PCCOS24, RKR20, RHL20, SMLRB^{+}21, SSF23, VHSV20]. adaptations [BRM^{+}21a, IRK^{+}21]. Adapting [LMSLMQ21, SOO20, SGA^{+}21]. Adaption [RHO24]. Adaptive [Dun21a, Kim23, dV21, BBJ^{+}22, Cra21, FSF^{+}21, HCWC21, TFH^{+}23, WGP^{+}21]. added [WCC^{+}20]. Adding [KY20, LTW23]. address [LTJ^{+}23, McC24, PFWM23, SHZ^{+}21, Tel23b]. Addressing [GTMP20, JSW22b, LJ23, OPL20, SGB21, CWKW20, FWGM24, SRY24, VMV^{+}24, Wen20, GSM21, PMM^{+}21]. Adjacency [MS20]. adjacent [BDMW22, QTPL21, Tod22b]. Adjustment [SWXM21, BBK21a, CQ20]. administrative [AKFY22, Oan22, Ros22, Sak21]. ado [OSV21]. adopted [BRP^{+}23, MB22]. adopting [CGGA20, KSS23]. Adoption [dlPdlLLC^{+}22, MB22, NCD23, TM23b, WMRB22]. Adriatic [SPN^{+}21]. advance [KFP20, KSB^{+}23a, LGW^{+}21]. Advances [SS23, TLT^{+}22, KSB^{+}23a, dS21]. Advantages [Jen23]. adverse [BMR20]. advice [AMR21b, CPR21, BKW^{+}24]. advisory [GLT23, LNRM23]. advocacy [Shi23b]. Aegean [SDL^{+}22]. aerial [SSOP^{+}21]. affairs [LTSL22, Wil21b]. affect [ÁCT20, HvPPV20a, JD20, NAN^{+}24, RJ21, WYD^{+}22]. affected [LWSS21, MNGM24, RCT20, SVV24]. affecting [LSY20, ZZW^{+}21, ZZD^{+}23]. Africa [AAA^{+}24, AWBC20, AGA^{+}21, BRM^{+}21a, BBHN20, CES20, DAN^{+}23, Ege23, FTGD^{+}20, FMLMM23, GSG^{+}24, MF22, NRF21, NAS^{+}22, OYECdOP24, SRvNN21, SON21, SS21b, SSP^{+}21a, SN23, UOO20, WLCG23]. African [BCBS22, BGMCG24, OYKBA22]. after [BMGDS^{+}20, NAP^{+}21, PBG^{+}22, SMC^{+}21, SFB24, TMDN23, ZTM^{+}23]. aftermath [BVvKL21]. Afuera [dPJ21]. against [ABME22, BSI^{+}21, Gar22, GY23, Gul23, HJPL23, KAN^{+}23, Ma20, Oan22, POO23, PGSS22]. age [HVN21]. agencies [Han21]. agency [BEM^{+}20]. agenda [GEB^{+}21, LMC^{+}23b, TSH^{+}22a, ZLY23]. agent [MF23]. agents [GC22]. agglomeration [ZW21]. Aggregating [CMT^{+}23, MSZ^{+}23, DPB^{+}23, EMH^{+}23, WLRH^{+}20, ZMM^{+}23]. Agoa [MLJ^{+}22]. Agreement [ALLS23, SLP^{+}22, Bal22, CMSN^{+}22, Cla20, GNJR24, HGL^{+}20, Kim20, Mar21a, MH23, Shi20a, BJH21, EB23, Has23, Kim21, Moh23, Tan24, VSV^{+}23, VYB20, Wan23a]. agreements [Aqu23, PK20b]. agricultural [ZCG^{+}23]. Agriculture [Gug22]. agrifood [MdCLL^{+}23]. ahead [GJS22, NCE21]. AHOY [KDS22]. AHP [AHBAKH24, CA21, DG20, LZL23]. AHP-entropy-based [DG20]. Aichi [GA20]. aids [Kar22c]. aim [CLD^{+}24]. aiming [NW20]. Air [VGK23, BG20a, CAC23, HKMS21, KSS23, Kon20, SOS20, VMV^{+}24]. AIS [ATF^{+}21, CLS^{+}23, MLJ^{+}22, SSOP^{+}21]. al [LLM^{+}23, MJMAA22, MJ23, SCD21, TGM22]. Alaska [Bir20, Bra22, DCC^{+}20, GBSC22, SMF^{+}23, SCH^{+}24, Szy20, SSF23]. albacore [Abo23]. albatross [DPG^{+}22]. Albert [GAA^{+}20]. ALDFG [GRS^{+}23]. Aleutians [Ree20]. algae [GBLBV20]. algal [EHLT21, SLMG^{+}21]. algorithm [KP24]. alien [DUYEA21, SSDQ20]. alienation [TNY23]. align [CPR21]. Aligning [SPHO23]. Aline [Tan23b]. Aliwal [FMLMM23]. alleviate [YSWW21]. alleviation [OWS20]. Allocation [FDF^{+}21, HGSD23, HM22b, MKS21, Ren21b, TKP23]. allocations [AGN^{+}20]. Allowable [DSJ^{+}23, ZLD^{+}21, Sta23b]. along [Alf20, CDRS22, FFE^{+}24, GSA23, HLLK22, KSV^{+}23, MDLA21, OWCFDC21, RPK23, ZWG^{+}21, ZNA^{+}23]. Alopias [MH20b]. already [Men23]. also [LZZ22]. Alternative [GM21, WL20, Gor21, Han21, HRS^{+}24, MF23, MCTN23, Pha20, WMS21, dSOW^{+}21]. alternatives [PCD^{+}23, dS22a]. Amanda [Rub24]. Amapá [JAdS^{+}20]. amateur [DD22]. Amazon [AML^{+}21, JAdS^{+}20]. Ambarlı [Okş23]. ambiguity [BvdH23, KvTNL20, Li23b]. ambition [Obu20]. ambitious [RSS^{+}20]. Ambon [LLLF22]. Ameghinomya [BRR^{+}22]. Amendment [ML22]. amendments [VDP^{+}22]. America [AE23, KABT21, MBC^{+}24]. American [MM20, PBH^{+}23, SK24, dOPM^{+}20]. americanus [MM20]. Americas [OMC^{+}22]. amid [HCW^{+}23]. amidst [SNK^{+}20]. among [BPA^{+}22, CLW^{+}21, DDD22, GWCD22, IOHC22, Lev20, MRKM23, MJ23, MMK^{+}22, ONN^{+}20, SKS21, TNY23, TSWC23, VM21, WSD21]. amount [HGG^{+}22a]. Ampat [SWI^{+}22, WMRB22]. Amphiprion [YNZ22]. Amsterdam [De 20, DPG^{+}22]. Anadromous [SCBdA20]. analyse [FFD^{+}23]. analyses [GÁPSAP^{+}23]. Analysis [CJ21, FFE^{+}24, HZ23, HCSBM^{+}23, JL21c, Kao23, MCRPC20, PKC23, PCD^{+}23, RJAT23, Şan20, SFGEPC22, SV22, AHBAKH24, ASKI20, AOO^{+}23b, AH23, AGA^{+}21, AONF21, ANR^{+}22, BG20a, BDLO23, BISSBdlMM22, BDCJ22, BGK^{+}22, BMC20, BBBS22, CSLGFV20, CMC^{+}23, CQ20, CDHY22, CLL^{+}22b, CCCC21, COSS21, CLB^{+}23, DGG22, DC20, DAN^{+}23, EAGV20, FY24, FFE20, FGPPGG22, FS21, For22, GMM^{+}20, GPM23, GBC21a, GBC21b, GY23, GMRSP23, GOH^{+}20, HW22, HSGNMGC21, HNMSG21, HSHS24, JY20, JLY23, JWX^{+}20, JHL21, Jon21, JKKT20, KV20, KJE21a, KJE21b, KLK23, KF23, KABT21, LHW21, LLW22, LXW^{+}21, LLL^{+}24, LTJ^{+}21, LJRHH21, Ma20, Mar21a, MBORGN24, MF20, MGF20, MCKZ21, Mur21b, ORSB22, PCSS^{+}21, QTD21, RZ21a, RS24, RHL20, SRvNN21, SC21, STIS20, Sha20, SGK21, Shi20a, SKL^{+}20, SJ21, SWH^{+}23, SPT^{+}20, SCY24]. analysis [SLS^{+}20, STF21, TJ21a, TMM^{+}23, TMS^{+}22, TPG^{+}23, VMC^{+}21, VVR^{+}22, WFR23, WZZC21, WSD21, WW22, XSCL21, YHU22, YT20, Yu22, YWYL22, YYD23, ZCCQ22a, ZT21, ZZD^{+}23, ZWC^{+}23, ZLH21, dKMK21, dPJ21, DFC^{+}20, WYD^{+}22]. analytic [HF20]. analytical [TJ21b]. analyzed [FLC^{+}23]. Analyzing [EMS^{+}23, RBA23, YU20, ABRA20, Chr20]. Anambas [AGSI22]. anchor [AKL^{+}22]. anchored [DPB^{+}23]. anchoring [GBFMT23]. Anchors [MUR^{+}21a]. Anchoveta [CDG^{+}21]. anchovy [CC21, DGG22, SWI^{+}22, UISM^{+}23]. Andalusia [CRCP^{+}23]. Andaman [TJS22]. Andreas [Smi23b]. Andres [PCCOS24]. anemone [YNZ22]. Angeles [WEF^{+}21]. Angler [BST23, CD24, GCM^{+}23, CLL20, GS21]. anglers [SSA^{+}23]. angling [ES22, GS21, QCM^{+}20, WDC^{+}20]. animal [HACCM22, SRLLG23]. Anne [CJ21]. Annex [VSV^{+}23]. annual [DDAB^{+}23, ZTA23]. anomalies [BK23]. answer [HMG^{+}21a]. answers [WLCC23]. Antarctic [GB21, AM22, ADB^{+}22, BL23a, CMC22, DKW20, Gar20, Hon21, TLYC20, TH22, Tra23b]. Antarctica [TBH^{+}21, BBK^{+}21b, KGW20, SB20]. Anthropocene [AHB^{+}23, BAF^{+}23, Dun21b]. Anthropogenic [YSEK22, FMGV20, MNGM24, MG22, YG20]. Anthropopause [YNZ22]. Anticipated [SR22, TKP23]. Antigua [SWF^{+}24]. antiqua [BRR^{+}22]. Anzac [Dun24]. AOPS [Laj21]. Aotearoa [LLT^{+}20, MAR^{+}20b, RRM^{+}24]. apparel [VVR^{+}22]. appendix [Sel20, HdM23]. applicants [BI21]. Application [FFE^{+}24, GCG^{+}21, POM24, VRBC^{+}21, BMB^{+}24, FAC^{+}23, HJLN24, KGL22, MXZ23, MFRQ23, Oan22, TOAPJ21, Zhu23, AE21, CBAPM20, SKL^{+}20, SRSGMGG21b, XWL^{+}20]. applications [AdSMS21]. applied [ABVPM20, dlFGFFR^{+}20, SRSGMGG21a]. Applying [JL21d, QLU^{+}23, VBH^{+}23, WWC23, QUK21, DAA21, WHPF20]. appraisal [CK23, COM20, HF20, TK23]. approach [AHBAKH24, AMK^{+}21, AVF^{+}23a, Alf20, ASM21a, ACG21, BRR^{+}22, CBO^{+}21, CMR21, CPBO20, CTH^{+}21, CAM^{+}23, CAC23, CMWZ24, CGGA20, CTH^{+}23, CSH23, CHB23, EMN^{+}24, FGK^{+}20, FMCT^{+}21, FFD^{+}23, GSGOP^{+}21, dlFGFFR^{+}20, dHJTGO21, GMG^{+}23, GJY23, GR20b, GSB20, GML^{+}24, HvT22, Han20, HGL^{+}20, IO23, LMC^{+}23a, LZL23, LKE^{+}22, LFVCF24, MGS21, MSBA22, NGP^{+}21, NSPRSS20, OKR^{+}20, Osm23, OAM20, PVH^{+}21, PBT^{+}20, PMAVM20, PS21, RBB^{+}24, RKS^{+}23, SWI^{+}22, SEG^{+}22, SOLN21, Sta23b, THP22a, VGS^{+}20, VRBC^{+}21, VBC^{+}20, WH20, ZMR20, ZCG^{+}23, ZGY21, JL21d]. approaches [BB20, BEH^{+}21, DSL^{+}21, GCS21, KAN^{+}23, Kul23, LRD^{+}22, MM23, NCD23, QLU^{+}23, QRK^{+}23, SG20c, TFMT^{+}23, VBH^{+}23, WCF^{+}21, ZWJvL20, dSOW^{+}21]. appropriate [Kai21]. approval [Kak20]. April [Ano21a, Ano22a, Ano23a, Ano24a]. Aquaculture [MF20, OOGA22, TVN23, Åm21, AHBAKH24, AOO^{+}23a, AO20, AELR^{+}20, BVvKL21, BE20a, BE20b, BDLO23, BBGHR23, BABC20, BBO21, BLS21, BTE21, CFM^{+}21, CMR21, CGGA20, CVA22a, CVA22b, CWMM^{+}21, Ero22, EMS^{+}23, FBC^{+}23, HJLN24, HS20, Han21, HHB^{+}21, HCSBM^{+}23, Her22, HSHS24, IKB21, JFB^{+}24, JY20, JKV^{+}22, KCCC^{+}21, Kob22, KBD^{+}20, LGF24, LFPBD^{+}20, LLL23, LFVCF24, MGS21, MCPS20, MVSP^{+}21, MM22, MMNC22, MBC^{+}24, MFF^{+}23, NSZ^{+}21, NVP^{+}23, OOSG21, POM24, PŚ22, RMF^{+}21, RFG24, RBNE20, Sah22, SZ20, Squ20, TCS^{+}20, TFF22, VMD^{+}24, VSPMS^{+}22, WFG22a, WMKG21, YHU22, YK20, YTG^{+}23]. aquarium [AP23, SC21]. aquatic [BRM^{+}21a, MCBN22, SN23, SSP^{+}21b, VGS^{+}20]. Arab [Gul20, Squ20]. Arabian [DMS^{+}24, SONB^{+}22]. Arabian/Persian [SONB^{+}22]. Archipelagic [VKS23]. Archipelago [HF20, MIH^{+}24, ASM21a, CWB^{+}20, TOAPJ21]. archipelagos [GMGLV21]. Arctic [VYB20, Bal22, Boy21, BVT^{+}21, BTV^{+}22, CDRS22, De 22, FFE20, Geh23, Gun21, GGL22, HLLK22, JHL21, KKA^{+}21, Kra22, KY20, Laj21, Osm23, Pri22, QLZ^{+}24, RCT20, Tod22a, VYB20, WA24, WCT^{+}20, WSDB20, YK21, ZXGL21]. Arctocephalus [PBH^{+}23]. Area [Cho22, FMLMM23, GBC21b, HSSM24, LTJ^{+}21, ML22, RCM^{+}22, RJAT23, TJ21a, TCT21, VMC^{+}21, WJLSH21, ACA^{+}21, AHAB^{+}22, ASM^{+}21b, BCJ21, CAGBM22, CJ22, DLJ^{+}20, DKW20, DB20, DSC^{+}23, EAA^{+}22, Fei21, GBC21a, HLW^{+}23, HPPH22, HLM^{+}20, HCT^{+}21, Jae21, JPM^{+}23, JL21c, Kim21, KTC^{+}22, LZRC23, MM21a, MCRPC20, PBH^{+}23, RP23a, Rus20, SAO^{+}21, SMLRB^{+}21, SPHO23, SJ21, SDG^{+}21, SDG^{+}22, SB20, TLYC20, Tes20, Tra23a, VWFW22, WD20, WMRB22, YT20, YCL22b, YC21b, dS22a, FMGV20, VBd^{+}21, GABA^{+}22, RP23b, SOT^{+}22, VCC^{+}21, War21, dS21]. area-based [EAA^{+}22, KTC^{+}22]. Areas [Has23, HSGNMGC21, KJE21b, MFG^{+}22, SFGEPC22, ACA^{+}21, BDM20, BD23, CBO^{+}21, CMGF20, CLRW22, Cha21, CMWZ24, DDAB^{+}23, Col20, CWR20, CMOJ^{+}24, De 20, Dua22, EV22, FDGM24, FTGD^{+}20, FMS^{+}22, GFD21, GNJR24, Gar20, GCMT20, GBLBV20, GCL21, GBFMT23, GBC21b, GWAPT21, GMVM23, HJ21, HLW^{+}23, HM22a, JWX^{+}20, JL21b, KFM^{+}20, KDM^{+}20, KVL^{+}23, KOCL23, cL22, MSBA22, MCR22, MPSF20, MMO^{+}24, NNM^{+}21, ONA21, OKH^{+}24, OKR^{+}20, OCCSC23, PBG23, Pha20, dlÁPAFGO20, PL23, Pri22, PKMT23, RGK^{+}20, RPY^{+}24, RKS^{+}23, SON21, Sol23a, SO20, THT^{+}23, TBH^{+}21, Tod20, TEA^{+}22, TJ21b, VHD20, VBC^{+}20, WZZC21, WSD21, WGJ^{+}21, YG20, YCL22a, YSWW21, YD22, ZJ22, ZPX^{+}22, Zha22, AIG^{+}20, ÁFFSC20, AVF23b, BvBN^{+}20, BEH^{+}21, CCS23, NJBM^{+}23, ORSB22, QTBL23, RA20, RCM^{+}22, Teo23, TCD^{+}20]. Argentina [SV22, TDC21]. Argentine [SV22]. arguments [BE20b, Sel20]. arid [AS23b]. armed [CLS20]. arrangement [Li23b]. arrangements [AGSI22, Cla20, FFD^{+}23, HIP^{+}21, Kra22, MMSB^{+}20]. Arte [ACV^{+}21]. Article [WL20, Fla21, Guo20, ALLS23, Shi23a, WL20]. Artificial [OVSVPP^{+}22, TDP20]. Artisanal [DRCdMCJ21, GRPRVG23, SBFHD^{+}23, AACdlT23, ACV^{+}21, BCBS22, CDNS23, hBDF22, FLB^{+}21, GSAF21, GWAPRPRU22, JGWM^{+}23, JTR^{+}20, KWS20, KLSG21, MS21, MH20b, OCD^{+}23, PRM^{+}20, PCCOS24, RPK23, RMBR22, SPG20]. ASC [GVL22]. Ascension [ASM^{+}21b]. ASEAN [SF22]. Asia [War23, BBC24, BRM^{+}21a, DAN^{+}23, FOKM23, JF24, MGDJ22, PGP22, PK20b, QK24, WM22a, YS21, AFHM23]. Asian [HH21, PKB^{+}23, GLAVL21, Ton22b]. asks [Has23]. aspect [Cha22]. aspects [BN20, CMC22, MLVT21, MSSS22]. aspirations [GAD^{+}24]. assemblage [OKFN24]. assess [CSVC23, DDD22, EAGV20, GSAF21, GKP^{+}22, MLJ^{+}22, MLVT21, PCD^{+}23, QLU^{+}23, TLT^{+}22]. assessed [ASV^{+}21]. Assessing [AR23a, BJGG^{+}24, BUR^{+}22, BHKK23, CMR21, CACJR^{+}21, CA20, CWB^{+}20, CAM^{+}23, Cho22, DME23, EAAG21, FGPPGG23, FMGV20, For22, GAD^{+}20, HGH^{+}20, MBL24, NH20, NGT^{+}23, ONA21, OKH^{+}24, Pra20, QTBL23, RPK23, SSP^{+}21b, WRL23, WABR^{+}22, dOPM^{+}20, FAGS21, MGBS24, WNMS^{+}24, AAKI21, ABM21, DKKS21, LTRA20, OPL20, Szy20, WGP^{+}21]. Assessment [AGM^{+}22, GAA^{+}20, HM22a, HF20, Koj23, LDL^{+}23, LFPBD^{+}20, MFC^{+}24, AANM^{+}22, AP23, BB21, Bov22, BSK^{+}20, CAC23, CTCD21, Cha22, CJL^{+}23, CL21, DBS22a, DGIS^{+}22, Fuj21, GSG^{+}24, GJY23, GMCMJA^{+}21, HJLN24, HYN^{+}23, HBR21, HBR^{+}23, HMY^{+}23, HSHS24, KHH^{+}23, KMJ24, LLCH21, Lov23, MPD^{+}23, MVR21b, MPL^{+}23, MN23, Moh23, MIH^{+}24, Oan22, OKR^{+}20, OMC^{+}22, PSMR20, PWL^{+}23b, QUK21, RXN23, SAO^{+}21, SC21, SNK^{+}20, WHL22, YTBM^{+}22, ZTZ^{+}21, ZCA^{+}23, PS21]. assessments [BBB23, CTH^{+}21, DQJ23, FOH^{+}20, SSOP^{+}21, SDG^{+}21, SDG^{+}22, TFMT^{+}23, CDC21b]. Asset [HBCA^{+}24]. assets [BS22, FSKJ23]. assignments [DSC^{+}23]. assistance [Sar23]. associated [DWC^{+}20, PMAVM20, QWBY21, SKS21, SKW^{+}22, ZCA^{+}23]. Association [LA23, CLL^{+}22b, SL23]. Associations [HVN21]. assumptions [LLM^{+}23]. Assuring [BBO21]. Asturias [dlFGFFR^{+}20, SPG20]. asymmetric [ZLGX20]. at-sea [EHM^{+}20, EHM^{+}22]. Atacama [CBO^{+}21]. Atlantic [BUM^{+}24, CMI^{+}22, FMGV20, MFT^{+}22, Mar20a, MF20, MAG^{+}22, RHST21, dOSTB^{+}20, WCF^{+}21, ÁFFSC20, Bjø23, CPV^{+}21, CCGM21, CNB^{+}23, CSGC22, DSL20, EB24, FRVB22, FTSG21, GSG20, Gra21, GCM^{+}23, GCC^{+}21b, HBHJ20, KLSG21, LCL20, MSGAR20, MMGAA^{+}21, MM20, NOL^{+}23, PCG^{+}23, PBG23, SGR21, SGL^{+}23, Sha22, SNS23, SBT^{+}23, SGZH^{+}21, SGH22, SBW^{+}21, TFF22, TCD^{+}20, VBH^{+}23, YK20]. atlanticus [GCM^{+}23]. Atoll [YCL22b]. Attack [PB20, HPG20, Los23, Los24]. attempts [OYKBA22]. attitude [GBB^{+}21, KCCC^{+}21, KHMM20, LM24, RPK23]. Attitudes [PC20, AAKI21, CVA22a, DDD22, HWM22, dOJCSLdSB21, ONN^{+}20, SAB^{+}20, WJB^{+}20, YTÇ22]. attitudes/behaviours [DDD22]. attitudinal [NGP^{+}21]. attribute [TDC21]. attributes [CSVC23, OMC^{+}22, QWBY21]. auction [AMIM21, KV20]. audiences [WWD23]. augment [SCCW21]. August [Ano20a, Ano21b, Ano22b, Ano23b]. Australasia [BBBS22]. Australia [BLL22, BR22, BBHN20, BBGG22, CCK^{+}21, DPP^{+}20, Dun24, FFD^{+}23, HPG20, HWM22, HMG^{+}21a, HGH^{+}20, Jon21, MCW^{+}22a, RHC20, RDD^{+}21, SGLF^{+}23, SHAC23, DR22, WWG^{+}23, YT20]. Australia-wide [HMG^{+}21a]. Australian [KB20, LHW21, MCW^{+}22b, MKS21, OAM20, OAM21]. Australianised [Dun24]. australis [PBH^{+}23]. Authority [Sin21, BMM^{+}20, JPM^{+}23, AR22, BHKJT23, BRM21b, GSM21]. automated [KP24, SKL^{+}20]. Autonomous [DBZ24, KZR22, Kle21, Ste23, dKMK21]. autonomy [Sar23]. autoregressive [SGK21]. availability [MAA^{+}24, MMNC22]. Available [RJAT23, MH24]. Aveiro [BASM22]. Aviv [PC20]. avoid [McC20]. avoidance [KP24]. awareness [DG24b, DDD22, GCV^{+}21, HT23, HWMO20, Lev20, OCD^{+}20, Pra24]. Away [JFB^{+}24, dO23]. Azerbaijan [AB20]. Azores [ASM21a, CVCV^{+}22, GMGLV22, MCRPC20].

Back [Bjø23]. Badouzi [CCCC21]. Bags [NW20]. Bahamian [Doe21]. Bahia [MN21, SBFHD^{+}23]. bait [LMC^{+}23a]. baited [HMG^{+}21a]. balance [AED20, Pre21]. Balancing [EV22]. Balearic [GBFMT23]. ballast [KLS22]. Baltic [PKH22, EVPMKP22, Kiz24, LEVET21, MCKZ21, PS21, RKS^{+}23, SLK^{+}23]. ban [BSI^{+}21, CHEC21, CD23, CLB^{+}20, GTA^{+}20, KGP^{+}20, Lev20, Pen22, PBT^{+}20, PBG^{+}22, vLM22]. Banda [EABA^{+}22]. bane [MCBN22]. Bangladesh [BUR^{+}22, BSI^{+}21, DAAP21, HHB^{+}21, HBP^{+}21, HSN^{+}21, INW20, IKB21, IRK^{+}21, PPK23, RKR20, SRR^{+}21, SSP^{+}21b]. Bangladeshi [ASKI20]. Bank [GÁPSAP^{+}23, WCF^{+}21, CLS^{+}23, Ken20]. bans [CLB^{+}20, WLW^{+}22]. Barbuda [JMB^{+}20, SWF^{+}24]. Barclay [AFHM23]. barcoding [BdLAdC^{+}20]. Barents [BFMS21, DD22]. barnacle [GRA^{+}22]. Barrier [DR22, Wat22, YHU22]. Barriers [KMS20, Lot22, MKM23, AHM22, BDHI21, CWMM^{+}21, FSF^{+}21, FHBS23, KSS23, LHTR23, NCD23, OBNK24, SDI^{+}20, SBP24, XZ23]. base [GSM21]. Based [HJPL23, Öİ24, AKFY22, ASM21a, ACG21, ALF^{+}22, AGA^{+}21, BMMF22, BV21, CPV^{+}21, CLS^{+}23, CC21, COJE24, CMR21, CMT^{+}23, CdLN^{+}21, CSLGFV20, CS21, CLL^{+}22b, CJL^{+}23, CCCC21, CES20, Col20, CSVC23, CSL21, DRY22, DBH^{+}21, DG20, EAGV20, EAA^{+}22, FHBS23, GFD21, GPG^{+}21, GLAVL21, GSB20, Gug24, HFD^{+}20, HLW^{+}20, HGG^{+}22b, HVN21, HGG22c, HC21, HSHS24, HLO^{+}21, IRK^{+}21, JKV^{+}22, KS23, Kar22a, KP24, KTC^{+}22, LLT^{+}20, LLW22, LYQM21, LJ22, LXW^{+}21, Liu23, LLS^{+}20, LFVCF24, MUR^{+}21a, Moh23, NSPRSS20, Okş23, OCCSC23, OVSVPP^{+}22, PBT^{+}20, PGSS22, PGA^{+}20, PS21, QYD20, QTD21, QLU^{+}23, QRK^{+}23, RZ21a, RR20, RXN23, RXX24, RSFvH^{+}22, RMU^{+}22, RDM^{+}20, RKS^{+}23, RDD^{+}21, SVF20, SVP^{+}23, STKC22, TM23a, THP22a, TCT21, VRBC^{+}21, VBH^{+}23, WLQD21, WHPF20, WLX21, WYD^{+}22, WHL22, XHZ^{+}21, YWY20, YYD23, ZZH21, ZCCQ22a, ZY21]. based [ZCS^{+}23, Zhu22, ZSWJ23, dOPM^{+}20, vTHS21, QS21, FMGV20]. baseline [DLJ^{+}20, GRS^{+}23, MBC^{+}24, TPG^{+}23, Wat22]. baselines [EMH^{+}23, VKS23]. basepoints [VKS23]. bases [Kuł22, SV22]. Basin [JZW21, Kas21, SV22]. basis [CC23, TML^{+}21, YLSC23]. Basque [RHO24]. bass [CAM^{+}23]. bathers [AMLC^{+}23]. batoids [PMFR^{+}22]. Batten [SMF^{+}23]. Bay [AAKI21, Bra22, DSQ^{+}22, HM21, Jon21, KJE21b, MF20, MMNC22, SGPF21, SHAC23, YTG^{+}23, ADR^{+}21, Gir22, IBC^{+}20, PGA^{+}20, PMT^{+}20, SL20, SJ21, THGR^{+}20, UISM^{+}23, WLX21, dZ20]. Bayesian [PBT^{+}20]. BBNJ [Bal22, Cla20, De 22, Dua22, GNJR24, HHD20, HGL^{+}20, Kra22, LV23, Lot23, MH23, MS20, MFH^{+}20, Shi20a, Tan24, Wan23a]. be [BdLAdC^{+}20, Lev20, Pha20, SWF^{+}24, Sin21, Sol23b, Ste23]. Beach [BISSBdlMM22, CDP22, SAB^{+}20, Cat20, CLC21, KL21b, Lev20, PBG^{+}22, YSEK22]. beach-activities [YSEK22]. Beach-user [SAB^{+}20]. beaches [Che21a, WFR23]. beacons [BMRM23]. bear [Geh23]. beauty [CDP22, WFR23]. because [SCF^{+}22]. becomes [AH21]. Becoming [AO20, Shi20b]. beds [KWY^{+}22]. been [Gal22, HWMO20, ZD20b]. begins [TM23b]. behavior [DNR21, Fen23, HZ23, dSJNS20, LLJZ22, MLM23, SKS21, WKO^{+}22, WFG^{+}22b, XXLW24, DLYM22]. Behavioral [BC21, Muk21, GPM20, NGP^{+}21]. behaviour [CBR21, LYQM21, SFGEPC22, YBW20]. behaviours [DDD22, PC20]. behind [YL21]. Beibu [TLT^{+}22, WLW^{+}22]. being [AVF^{+}23a, GCF^{+}23, MNGM24, MTC^{+}23, RGH^{+}22, SCF^{+}22]. beliefs [Fro21]. Belize [AGK22, Mur21b]. Belizean [GABA^{+}23]. Below [NSN22, Gul20, MH20a]. Belt [HH21, TBPC20]. Benchmarking [WD23]. Beneficial [PGK22, GOH^{+}20, SOK^{+}22]. benefit [AH23, BTE21, LWX^{+}20, MB21, SC21]. Benefits [BBB^{+}20, KTC^{+}22, PSTW21, BBS^{+}21, BvBN^{+}20, BS22, DD22, JKKT20, KJS^{+}22, MRKM23, PFG^{+}21, QUK21, SBDC24, Squ20, TSA^{+}22, TFH^{+}23, TCS^{+}20, WWG^{+}23, WLCC23, WBPG20]. Bengal [CDP22]. Benin [GSG^{+}24]. benthic [DGIS^{+}22, DS22b, FARBV^{+}24, GSAF21, NJBM^{+}23]. benthopelagic [CDC21a]. Bering [Tod22b]. Bermuda [WLRH^{+}20]. Best [BKW^{+}24, Squ21, ABM21, GG21, MF22, MWM^{+}23a, MH24]. better [CAHR^{+}24, FMLMM23, GMRSP23, OBNK24, SOO20]. Between [hBDF22, ABM21, BCN23, BDDE20, BMB^{+}21, CSL21, DC23, DAAP21, GPA^{+}22, GLAVL21, dHJTGO21, GCL21, GVH^{+}23, GGL22, HTP^{+}23, Hor23, HWY22, IGO21, LJ23, LHPJ23, LZRC23, MZ24, Mar21a, PDB^{+}23, PSPR20, PL20b, PKMT23, RTMCVMDE22, RHST21, SL23, SLB21, SSA^{+}23, Shi23a, SS23, SSMG21, TH22, Tra23a, TJ21b, VGK23, Zha22, ZY23, vdGD21]. Beyond [CMD^{+}20, FDF^{+}21, Has23, MTN23, SDG^{+}21, SDG^{+}22, Teo23, TM23b, TCD^{+}20, VM21, WBPG20, dZ20, AS23a, BLL22, BBO21, CMOB^{+}20, Cla20, CWR20, DMNT20, EAA^{+}22, FARBV^{+}24, FMS^{+}22, GNJR24, Gar20, HJ21, HM22a, HGL^{+}20, MDJ^{+}22, MH24, MCR22, NRF21, OKR^{+}20, PSTW21, Pri22, QK24, TTA^{+}20, TD21, TMDN23, VHD20, VBC^{+}20, WPWM^{+}22, WGJ^{+}21, YG20, ZJ22, Sam22]. Bhitarkanika [KHMM20]. bias [PBG23, RZ21b, TDP20, WFR23]. biased [KSB^{+}23b, Ren21a]. Bibliography [Smi22, Smi23b]. bibliometric [BMR23, CTCD21, RXN23, TMM^{+}23, WHL22, WNDZ23]. bibliometric-based [RXN23, WHL22]. bicinctus [YNZ22]. bid [SKE22]. Bigi [DB20]. Bill [Dbo22, NW20]. billfish [KWS20]. BIMCO [Bov22]. binding [CWR20, HvT22, LR20, Pri22]. Bio [DAAP21]. Bio-economic [DAAP21]. Biodegradability [DFF23]. Biodegradable [ZMM^{+}23, EMH^{+}23, MSZ^{+}23, MZT^{+}23, SGL20]. Biodiversity [Has23, BLL22, DMNT20, FWGM24, FFGJ22, GNJR24, GABA^{+}22, HJLN24, HGL^{+}20, MDJ^{+}22, MH24, PKB^{+}21, PT23, Pri22, RSS^{+}20, TD21, TMDN23, TV21, DR22, MTN23, TM23b]. bioeconomic [BCBS22, Mot20, Mot22, PFWM23]. Biological [MG22, SLP^{+}22, FMS^{+}22, HM22a, LHV^{+}20, SMF^{+}23, TDC21]. Biology [dSJNS20, GSA23]. bioluminescent [SWW^{+}21]. biomass [MDG^{+}23, PMAVM20]. Biosecurity [MBC^{+}24, KCCC^{+}21, SSDQ20, VWJMAJ20]. Biosphere [MMSB^{+}20]. biota [CDC21a, JHHG23]. Bird [SEG^{+}22]. birds [BEM^{+}20, Ber23]. Biscay [ADR^{+}21, IBC^{+}20, PGA^{+}20, PMT^{+}20, UISM^{+}23]. bite [SAB^{+}20]. bivalve [BHB^{+}24, NNM^{+}21, PL23]. Bjørn [Fla21]. Black [SSZ^{+}22, GEB^{+}21, MJMAA21, QLZ^{+}24, SCD21, TGM22, BMM^{+}24, DGG22, DUYEA21, ES22, MÇ23, Şan20]. blanket [CLB^{+}20]. bleaching [DDAB^{+}23]. blind [CN23, NBB^{+}23]. Blockchain [NCD23, How20, LFVCF24, MBL24, ZSLY20]. blockchain-based [LFVCF24]. bloom [XLLS21]. blooms [EHLT21, SLMG^{+}21]. Blue [AC23, BBWC21, BE20c, CS21, HSGS^{+}23, SLW^{+}23, Tho22, TFMT^{+}23, Als20, AHB^{+}23, ANR^{+}22, DKKS21, DČFW23, DSL20, FFD^{+}23, FAB^{+}23, GdACT20, HJ21, Has20, HBCA^{+}24, HSN^{+}21, KNY22, KWY^{+}22, Lou22, MMB^{+}23, MFC^{+}24, MCKZ21, OCZ^{+}22, PSD23, Puh23, QZZX20, Qi22, Qi24, QUK21, QK24, RRM^{+}24, SDI^{+}20, SHPW21, SSA^{+}23, SWHT22, SR22, SOK^{+}22, TFH^{+}23, TT22, TAP^{+}23, WXLD21, WLQD21, AC23, CFM^{+}21, GFL^{+}21, HvT22, IMA22, LMMT23, MFC^{+}24, SCJ22, SWW^{+}21, WMKG21, ZFGL20]. bluefin [Bjø23, GSG20, HWT23]. bluefish [ES22]. board [ASKI20, MRD^{+}20, OVA22, SSDQ20, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano21p, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano22o, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano24b, Ano24c, Ano24d, Ano24e]. boat [CRJ22, Kle21, MAR^{+}20b, RDD^{+}21, SCGC20]. boat-based [RDD^{+}21]. boater [DWC^{+}20]. boating [DWC^{+}20, PHF20, VMC^{+}21]. boats [ÁCT20, WDC^{+}20, BEM^{+}20]. bodies [BLL22, SHFP21]. Body [GNJR24]. Bohai [LYHL23, LSH21]. bolts [SG20a]. bonds [Tho22]. Bonferroni [CMR21]. Bonn [VSV^{+}23]. Book [AFHM23, Cha20b, De 20, HH21, Lv20, Ngu23, Niu22, Rub24, Smi20, Smi22, Smi23a, Smi23b, Sos21a, Sta23a, Tan23a, Tan23b]. Boom [BDD^{+}21]. Boon [MCBN22]. border [De 21, GSGOP^{+}21, IO23, MMK^{+}22, Tse21]. border-crossing [Tse21]. Bosco [Smi22]. Bosomtwi [GAA^{+}20]. Bosomtwi-Sam [GAA^{+}20]. both [VHD20, WWD23]. Bothnian [EVPMKP22]. Bottom [FARBV^{+}24, BMGDS^{+}20, DAA21, DDF^{+}20, DGIS^{+}22, LCAX21, MVR21b, NSK20, RQB^{+}20, SDG^{+}23, SLB21, SC22, ZV20]. bottom-up [SLB21]. boundaries [BDC23, SDG^{+}21, SDG^{+}22]. Boundary [Smi23b, CG22, hC24, Fra23, LLT^{+}20, Øst20, PKH22, PK20b, RB22, SHPW21]. boundary-crossing [LLT^{+}20]. box [GEB^{+}21, Tod20]. brave [KFP20]. Brazil [LFLF22, dOSTB^{+}20, AFMdSdTdCC22, AML^{+}21, BNF21, BdLAdC^{+}20, CMI^{+}22, dFVCGdS21, DMd^{+}20, FSF^{+}21, GCMT20, HSM^{+}23, JAdS^{+}20, dSJNS20, KHdSM^{+}22, MIMS^{+}20, MN21, MMS24, ORSB22, PMC24, PBH^{+}23, RCM^{+}22, SdACdSLdAS24, SA21, SBFHD^{+}23, SSB^{+}22, dS21, dSdLSGB22]. Brazilian [VBd^{+}21, GCF^{+}23, GdACT20, MM21a, Mar20a, NH20, NAP^{+}21, NBB^{+}23, SSZ^{+}24, SS23, dS21, dS22a]. break [MMS24, WCF^{+}21]. Breaking [AC23]. breeding [OKH^{+}24, WCT^{+}20]. Brexit [BFMS21, BM21, DGS^{+}24, EB24, FDF^{+}21, Fox22, KFP20]. bridge [FS21, Osm23]. bridging [WMRB22]. Brief [CMWZ24]. brink [Bjø23]. Bristol [Bra22]. British [DBS^{+}22b, SSB^{+}23, WJLSH21]. broad [HMG^{+}21a]. broad-scale [HMG^{+}21a]. Broadening [SS21a]. broader [CMS^{+}21]. Brochmann [Fla21]. Brooks [Sos21a]. brown [GABSD21]. Brushett [Rub24]. BRUVs [HMG^{+}21a]. buds [PBG^{+}22]. build [Jon21, MZ24, PALB20, SCLM^{+}21]. Building [DB20, How20, RJAT23, XT23, AKFY22, AAA^{+}21, ACG21, FMS^{+}22, TYS24, TZC23, VHD20, WNBRSG22]. Bulgaria [SSZ^{+}22]. bullying [MLE^{+}24, ÖB22]. bunkering [HPS23, MBL24]. Busch [Niu22]. Business [LEVET21, BDR20, MLVT21]. business-oriented [MLVT21]. bust [BDD^{+}21]. buy [CDB20]. buy-in [CDB20]. by-catch [MSGAR20]. Bycatch [GCB^{+}23, RQB^{+}20, WLW^{+}22, BKW^{+}24, BEM^{+}20, CPBO20, CLB^{+}23, DER^{+}21, DBB^{+}22, HBHJ20, JSW22b, MM23, MWB^{+}21, SBW^{+}21]. Bycatch-neutral [GCB^{+}23]. bycatches [BUM^{+}24].

CA [AS23b, KL21b, NHD23]. CA-Markov [AS23b]. cabled [AIG^{+}20]. Cabo [HSGNMGC21, HSGNMGC21]. cabotage [CC22]. Cacão [BdLAdC^{+}20]. Cadiz [GMPT^{+}20, GTA^{+}20]. cage [HACCM22, SMHPMO^{+}21]. Cages [DMS^{+}24]. calculation [CA21, VGK23]. Caledonia [LDAC20, DPP^{+}20]. Caledonian [GA20]. calendars [HDCH21]. California [ABDRV^{+}23, ALLR^{+}24, BCMCM21, OKS^{+}22, PLOG24, QCM^{+}20, RC22, VWFW22, WEF^{+}21]. call [AVF^{+}23a, CCB^{+}21, FFYE21, VJL^{+}21]. Calling [AGSI22]. Cambodia [LHTR23]. Cambridge [Niu22, Smi20, Tan23a]. cameras [AIG^{+}20]. Cameron [Ngu23]. Cameroon [NAS^{+}22]. campaigns [WWD23]. Campeche [GÁPSAP^{+}23]. Can [AACdlT23, AAW^{+}23, Fen23, Gir22, NWZ^{+}21, PMN^{+}22, RM21b, Sol23b, Ste23, TLYC20, WS21, WKH22, YSWW21, BEM^{+}20, BdLAdC^{+}20, CBR21, CMS^{+}21, MWB^{+}21, RLS20, RMBR22, SWF^{+}24, SWHT22, TSA^{+}22, TFH^{+}23, YFD^{+}21, CDC21b, HHEJB20, Pha20]. Canada [BBS^{+}21, BMRM23, BMB^{+}21, CA20, CMNH21, CGP23, JHHG23, JSW22b, KDH^{+}21, Laj21, LTRA20, NGRM^{+}22, OCT^{+}22, Sha22, SSB^{+}23, WA24, WJLSH21, WH20, ZSWJ23]. Canadian [AMR21a, AMR21b, CMS20, CLD^{+}24, DKKS21, Gor21, HBHJ20, WMKG21]. Canary [GMCMJA^{+}21, MBORGN24]. canned [Kaw22]. canning [dHJTGO21, RAFLRP21]. canyon [SCA^{+}22]. Capability [GDLC20]. capacities [BBS^{+}21]. Capacity [LLLF22, LLCH21, OCD^{+}23, VHD20, ACG21, BXV20, FAC^{+}23, FMS^{+}22, GBFMT23, HCWC21, JDF^{+}23, Kim23, OEV23, PHF20, SMHPMO^{+}21, WGP^{+}21]. Cape [LA23, SAO^{+}21, Aqu23]. Capital [JL21d, RAFLRP21, RC22, SZ20, Sol23b, SOLN20, TV21, XXLW24, CLZ23]. captain [Ste23]. captains [CSGC22]. capture [CCKK22, GSA23, TMS^{+}22]. captured [TSS^{+}24]. Capturing [IO23, MHEW20, HKKY22]. Carbon [Okş23, WW22, ANR^{+}22, CDHY22, DGG22, DXS^{+}21, DG24b, Dom23, FAB^{+}23, GZW^{+}20, HZ23, HSGS^{+}23, KNY22, KWY^{+}22, NHD23, QLZ^{+}24, QTBL23, QUK21, QK24, RLRS21, SLL^{+}24, SGA^{+}21, SLW^{+}23, SOK^{+}22, TBV^{+}23, TMTC20, TAP^{+}23, WXLD21, WLQD21, WL22b, WLXY22, ZFGL20]. carbonate [PCG^{+}23]. carding [KL21a]. care [VKN20]. Caribbean [BJH21, Has23, Lov23, Mah22, Moh23, SWF^{+}24, WLRH^{+}20]. CARICOM [Has20]. Carolina [ALML23]. carrying [FAC^{+}23, GBFMT23, PHF20, SMHPMO^{+}21]. Case [CJ23, HdM23, LA23, MLJ^{+}22, PKC23, RJAT23, SWI^{+}22, SN23, APV^{+}20, ASV^{+}21, AOO^{+}23b, AACdlT23, ACV^{+}21, Aqu23, AS23a, AGSI22, AR23a, ASGZ22, BSBD21, BE20a, BDLO23, BMGDS^{+}20, BJGG^{+}24, BDDE20, BUM^{+}24, BMB^{+}21, CVCV^{+}22, CC21, CWB^{+}20, CMC^{+}23, CDHY22, CTIH21, CLC21, CCCC21, CCKK22, CES20, CLB^{+}20, CVA22b, dFVCGdS21, Cro21, COJZ^{+}21, DRCdMCJ21, DSQ^{+}22, DUYEA21, DFF23, DG20, Dun21b, EAGV20, EVPMKP22, FTGD^{+}20, FY24, FGPPPV21, Fro21, FMWW21, FMS^{+}22, GPA^{+}22, Geh23, GPM20, GPM23, GCC^{+}21b, GAA^{+}20, Han21, HLW^{+}23, HRR^{+}20, HHB^{+}21, Hor23, HCT^{+}21, IBC^{+}20, JY20, JTR^{+}20, JAH^{+}24, JLY23, JAdS^{+}20, JG23, JMB^{+}20, JL21c, Kim23, Koe22, Kon20, KABT21, LrS22, cL22, LWG23, LTJ^{+}23, LJ22, LJL22, LLSY22, LJ21, LSY20, MAA^{+}24, MVSP^{+}21, MSBA22, MMSB^{+}20, MGF20, MBL24, NHD23, NNRS24, NEH^{+}23]. case [OCD^{+}20, Okş23, OVSVPP^{+}22, OAM21, PL24a, PCA21, PRM^{+}20, PKR^{+}22, PBG23, PMR23, PC20, Qi24, RBA23, RMF^{+}21, RBB^{+}24, RDM^{+}20, RNGRC22, Rus20, SPG20, SGR21, SVV24, SSZ^{+}24, SMYH22, SCGC20, SCCW21, SOS20, SGZH^{+}21, SCLM^{+}21, SZBA20, SSZ^{+}22, SRSGMGG21b, SLK^{+}23, TM23a, Tei23, TBH^{+}21, THGR^{+}20, TSH22b, TLT^{+}22, Tru23, TDP20, TYS22, UISM^{+}23, VGK23, WD23, WJB^{+}20, WW21a, WSD21, WRL23, WFG^{+}22b, WKBS20, WDC^{+}20, WLW^{+}22, WWC23, YFD^{+}21, YS21, YCL22b, YD22, YC21b, ZJ20, ZCCQ22a, ZCCQ22b, AANM^{+}22, CMOJ^{+}24, GLADGGVL23, IKB21, LDAC20, ZWJvL20]. case-study [Cro21]. cases [dlÁPAFGO20, STG^{+}20]. cash [ABRA20, BK23, NRYG20]. Caspian [AB20, Pie21]. Cassandra [Sos21a]. Cassurubá [MN21]. Castellón [DL21]. castle [FTG^{+}23]. casualty [FFE^{+}24]. Catalonia [GPM23]. catalyze [FLB^{+}21]. catalyzing [MGDJ22]. Catch [DSJ^{+}23, HBJ21, ZLD^{+}21, BXV20, GÁPSAP^{+}23, GMM^{+}20, MLMD21, MMM^{+}22, Mat20, MAK22, MAG^{+}22, MSGAR20, STIS20, SCSR^{+}22, SCSR^{+}23, SVV24, TSLK23, TYS22, YCL22a]. catches [CPR21, CHB^{+}20, IBC^{+}20, KMG^{+}20, MH20a, SFGEPC22, SIW^{+}21, Sta23b, VAMR^{+}20]. Catching [Mar21b]. categorising [SRS23]. category [Kai21]. Caught [SH22, YS21, MM22]. causality [dHJTGO21]. causative [LM24]. cause [De 21, RE21]. caused [KLK23, MRKM23, SL20, WABR^{+}22, dOPM^{+}20]. Causes [GAD^{+}23, DT23]. causing [MÇ23]. Cavallo [HNMSG21]. Cayman [BDMW22, Dun21a]. CCAMLR [CHC^{+}21, SB20, TBH^{+}21, Wen20]. cell [LJ22]. cell-based [LJ22]. Cellular [Tel23a]. Celtic [FGPPGG22, CRP^{+}20, FGPPGG23]. census [APF22]. centered [SDG^{+}21, SDG^{+}22]. Central [Abo23, MAG^{+}22, TCS^{+}20, CS21, FMCT^{+}21, RNGRC22, AS23a, CMC^{+}23, COJZ^{+}21, DMd^{+}20, VYB20]. central-southern [FMCT^{+}21]. Centralization [Sak21]. centre [LMMT23]. centuries [Ken20, TUBA22]. Century [CTH^{+}23, PZPB21, SF22]. CEO [VJO23]. Cerastoderma [BASM22]. Certifiability [NRF21]. certification [AO20, BI21, BGMCG24, HPG20, LJ21, NK21, NRF21, PCK22, RFG24, Sah22, WS21]. certifications [JAK^{+}23]. certified [AO20, LLM^{+}23, SPG20, TOAPJ21, TOL^{+}22, WEF^{+}21]. cessation [LBL21]. Cetacean [OWCFDC21, DER^{+}21, RDM^{+}20, SPJB22]. cetacean-based [RDM^{+}20]. Cetaceans [WtDD^{+}21, BBBS22, DBB^{+}22, KDH^{+}21, MLJ^{+}22, MIMS^{+}20]. Cg [Yu22]. Cg-18 [Yu22]. chain [ALM^{+}21b, Bot20, CLB^{+}23, CSCGC20, dFVCGdS21, GSG^{+}24, GNMGWM21, How20, HP20, JAdS^{+}20, JKKT20, KWMM20a, KWMM20b, LCL20, LLL^{+}24, MGS21, MMM^{+}22, MMW^{+}23, NAS^{+}22, PCSS^{+}21, Seu22, SdCMN^{+}22, THP^{+}22b, TGHF^{+}22, TPG^{+}23, VKN20, WCC^{+}20, WEF^{+}21, YTG^{+}23, ZW21, ZCSK23, BMR20, dS22a]. chains [AANM^{+}22, BRM^{+}21a, DLJ^{+}20, FFYE21, HRC^{+}24, RMU^{+}22, TOHK20, VVR^{+}22, WKBS20, WMS21]. challenge [AML^{+}21, Kaw22, LC20a, SH23, WEF^{+}21, YK21, YYL20, AMK^{+}21]. challenged [TJ21a]. Challenges [SONB^{+}22, DBZ24, GG21, GMGLV21, JF24, KWF^{+}21, KHdSM^{+}22, KOCL23, Liu23, RWVH21, RCM^{+}22, Sha22, SCGC20, SF22, Tel23a, WD20, WdSRY23, AM22, BH23a, BI21, CC22, CWKW20, CGGA20, Cra21, DQJ23, DG23, EMN^{+}24, Fis22, FMLMM23, GSR21, GN22, GE23, GJS22, HIP^{+}21, HFD^{+}20, HLM^{+}20, INW20, Kob22, Koj23, Kul23, KSV^{+}23, LMSLMQ21, LJ23, LFVCF24, LL24, Mag23, MIH^{+}24, NSZ^{+}21, NJBM^{+}23, OPL20, OYECdOP24, PVH^{+}21, QK24, RFPS21, RSFvH^{+}22, SAA^{+}21, SSZ^{+}24, SEG^{+}22, Tod22a, WvdMG^{+}21, WWG^{+}23, WRV21, Wen20, Wil21b, WLXY22, Wu22, YL21, ZZH21, ZSZ20, ZSLY20, CLS20, CWR20, JL21c, SIW^{+}21, WMKG21]. Challenging [Cro21, KLSG21, MFT^{+}22]. Cham [AFHM23, KJE21b, Rub24, Sta23a]. Chamelea [BSBD21]. Change [Kar23, LDC^{+}20, Niu22, Smi20, vHSSK20, AMLC^{+}23, BG20b, BSM23, BvdH23, CWB^{+}20, CWKW20, CWKE20, De 21, DP23, EAGV20, FWGM24, GCV^{+}21, GB21, HCWC21, HBR^{+}23, Jon21, Ken20, KHH^{+}23, Kon20, KDVR20, LMSLMQ21, Mal21, MGDJ22, Mat20, McC24, MHN^{+}20, O’D21, OGGC20, PBS^{+}23, PK23, QF20, RB20, RHO24, RDLA^{+}20, RHL20, SOO20, SMLRB^{+}21, SBD^{+}23, SGL^{+}23, SRH23, SOK^{+}22, SN23, SHFP21, TKP23, TMTC20, VKS23, WWD23, WtDD^{+}21, WGP^{+}21, YG20, dSdLSGB22]. Changes [GS21, AFMdSdTdCC22, BASM22, CBR21, CLL20, CAM^{+}23, FRVB22, GTMP20, HCW^{+}23, KPN20, KMJ24, MdCLL^{+}23, MK24, NHA23, OCZ^{+}22, SSK23, TLT^{+}22, TGHF^{+}22, VGC^{+}22, XWH^{+}22, ZZH21]. Changing [CMOB^{+}20, DKKS21, Fen23, MSO23, GMG^{+}23, HSH^{+}21, HACCM22, Tra23a, Wen20, YMG23]. Channel [DFF23]. channels [ZZW^{+}21]. Chanos [SGPF21]. Chaos [ABME22]. Characterising [KDVR20]. Characteristics [YYD23, EdLP^{+}21, FLC^{+}23, JKV^{+}22, WL22b, ZSZ20, ZWC^{+}23, CMS20, ZLZ21]. Characterization [SVP^{+}23, VMC^{+}21]. charter [Bov22, CRJ22, WDC^{+}20]. Charting [MBDG23]. charts [Alm21a]. check [dOPM^{+}20]. Checking [BASM22]. Cheilinus [HdM23]. Cheltenham [Smi23b]. chemical [ALML23, Lev20, MLM23]. chicken [HM22b]. child [TOL^{+}22]. Children [KL21b]. Chile [DSQ^{+}22, PMR23, AELR^{+}20, AS23a, BVvKL21, BMH^{+}22, BRR^{+}22, CBO^{+}21, FMCT^{+}21, GRPRVG23, GCV^{+}20, MS21, RMBR22]. Chilean [AS23a, BBGHR23, CC21, GSAF21, LKE^{+}22, RQB^{+}20, SCLM^{+}21, WFG^{+}22b]. Chiloé [DSQ^{+}22, BMH^{+}22, PMR23]. China [CHB^{+}20, GC22, Guo20, HJPL23, HH21, HGLR22, Hsi20, PL20a, PWL^{+}23b, SW22, SHZC21, Zha21, ZFGL20, ZLGX20, Zou23, AWX21, BHdMCM^{+}21, CC22, CC23, CLL22a, CLZ23, CQ20, CDHY22, CMWZ24, CSHO^{+}20, DLX21, DSJ^{+}23, DG20, Dua22, FXC^{+}22, FZYF23, FWL^{+}22, GY23, GW21, Guo20, GGL22, HJPL23, HPH22, Hon21, HLM^{+}20, HSHS24, JL21a, JLY23, JWX^{+}20, JF20, JL21b, JZW21, JF22, JK22, KGP^{+}22, LDC^{+}20, LXSX21, LW22, LLHG23, Li23b, LMC^{+}23b, LC23, LYQM21, LXH^{+}22, LYL^{+}22, LDL^{+}23, Lin24, LWX^{+}20, Liu20, Liu21, LXW^{+}21, LJL22, LHPJ23, MCL22, MZ24, MX21, NWZ^{+}21, PZZ^{+}21, Qi24, QYD20, QTD21, QWBY21, RZ21b, RJ21, Ren21a, Ren21b, RXN23, RXX24, Sha20, SC22, SLY21, Shi23b, SGWL21, SF22, SFB24, SGRW22, SWXM21, SWW^{+}21, STKC22, SCY24, TLYC20, Tan22, TZZ^{+}21, TKC23, TBPC20, WDG^{+}20, WSD21, WCZS21, WLX21, WW22, WdSRY23, WLW^{+}22, WL22b, XZ23, XLLS21, XHY22, XT23, YWY20]. China [YSWW21, YW20, YYL20, YH20, YD22, YWYL22, YYD23, YLSC23, ZV20, ZL20, ZTZ^{+}21, ZMZ23, ZJ20, ZPX^{+}22, Zha22, ZCCQ22a, ZCCQ22b, ZZW^{+}21, ZT21, ZLZ21, ZZD^{+}23, ZWC^{+}23, ZCSK23, ZCS^{+}23, Zho23, ZLH21, ZLD^{+}21, Zou23, WRL23]. Chinese [Alf20, CWKW20, Cha22, DLYM22, LC20a, LCAX21, LL24, Mar21b, MLAS23, MLM23, SGK21, SHZ^{+}21, SFN^{+}21, YTG^{+}23, ZWJvL20, ZZH21, ZTW23, ZSZ20, ZWXL21]. choice [CDD^{+}23, CLL20, Faj22, KNY22, SNP21, SL20, TOAPJ21, THT^{+}23, Tra23a, WLX21, WYD^{+}22]. chokes [CBR21]. chondrichthyans [MT20]. Chorrera [GBC21b]. chronic [PGK22]. chum [KS23]. ciguatera [Nal24]. circulation [Cha22]. Circumpolar [BB20]. CITES’ [Mag23, HdM23, Kai21, Sel20, SMW^{+}21]. cities [GPM20, LHPJ23, RZ21b, RJ21, RXX24]. Citizen [SCA^{+}20, GS21, GVH^{+}23, HLCM21, HMG^{+}21b, TSLK23, WFG^{+}22b]. City [PCSS^{+}21, TM23a, DRCdMCJ21, SdCMN^{+}22, TCT21]. civilization [LYL^{+}22, RXX24, TZC23]. claim [SCF^{+}22, SMCW22]. claims [CG22]. clam [BSBD21, BRR^{+}22, LCMBGP22, SVP^{+}23]. clarifications [dS22a]. Clarifying [TAP^{+}23]. Clark [De 20]. Class [Dun24]. classification [HLW^{+}23, HLO^{+}21, PL23]. clauses [Bov22]. clean [BST23, CCX23, TSLK23, THS^{+}24]. clean-up [TSLK23]. cleaning [CMI^{+}22]. Clew [MMNC22]. client [RMU^{+}22]. Climate [GB21, Kar23, MHN^{+}20, Niu22, OOGK^{+}24, Pre21, SN23, VKS23, AH23, BG20b, Bir20, BB20, BvdH23, CWB^{+}20, CAVJ^{+}20, CWKW20, CWKE20, DKKS21, FWGM24, GLT23, HCWC21, HSH^{+}21, HBR^{+}23, HGLR22, HSM^{+}23, HSGS^{+}23, Jon21, KS23, KHH^{+}23, Kon20, KDVR20, LMSLMQ21, LGW^{+}21, Mal21, MKM23, MWM^{+}23a, MN23, McC24, Moh23, O’D21, OCD^{+}20, OGGC20, PK23, PCCOS24, RB20, RHO24, RDLA^{+}20, RHL20, SOO20, SMLRB^{+}21, SGL^{+}23, SOK^{+}22, SHFP21, SSF23, Tel23b, TFH^{+}23, TLT^{+}22, TKP23, WWD23, Wen20, WtDD^{+}21, YG20, dSdLSGB22, Smi20]. climate-adaptive [TFH^{+}23]. climate-ready [LGW^{+}21, MWM^{+}23a]. climate-responsiveness [DKKS21]. climate-shifted [Tel23b]. Climate-smart [OOGK^{+}24]. climates [DKKS21]. climatic [IRK^{+}21, SDL^{+}22]. clock [AR22]. close [HSSM24]. closed [MLMD21, MMM^{+}22, MMC^{+}22, OA20]. closer [LC23]. Closing [GR20b, MCPS20]. closure [Pra20]. closures [Ber23, De 21]. Clownfish [YNZ22]. club [GR20b]. clubs [Sak21]. clusters [SPT^{+}20, ZYTT21a, ZYTT21b]. CO [GZW^{+}20, BLY^{+}24, HYN^{+}23, MSBA22, SB20, CPV^{+}21, CdLN^{+}21, CSL21, DL21, DBS^{+}22b, Dun21b, EMS^{+}23, Gar20, GRPRVG23, GWAPRPRU22, GEB^{+}21, HW22, INW20, Kim23, KJS^{+}22, LJ22, LA23, MRD^{+}20, MAR^{+}20b, PL20b, PWL^{+}23b, SCJ22, SSZ^{+}24, SPHO23, WJB^{+}20, WBPG20, MVSP^{+}21, OOGK^{+}24, SSK23]. co-benefits [WBPG20]. co-development [CPV^{+}21, EMS^{+}23]. co-evolution [LJ22]. co-exist [Gar20]. co-existence [WJB^{+}20]. co-existing [GRPRVG23]. co-governance [MAR^{+}20b]. co-led [DBS^{+}22b]. Co-location [BLY^{+}24]. co-managed [CdLN^{+}21, Dun21b]. Co-management [HYN^{+}23, MSBA22, CSL21, DL21, GWAPRPRU22, HW22, INW20, Kim23, KJS^{+}22, LA23, MRD^{+}20, PL20b, SSZ^{+}24, SPHO23]. co-operation [PWL^{+}23b]. co-produced [GEB^{+}21]. Co-production [SB20, SCJ22]. Coast [BBS^{+}21, KABT21, SBFHD^{+}23, TCS^{+}20, AB20, ADR^{+}21, BOOO23, CHEC21, CACJR^{+}21, DXS^{+}21, GSA23, GBFMT23, HDAPG^{+}24, JAdS^{+}20, dSJNS20, KWS20, KSV^{+}23, LCL20, MDLA21, Mba22, OWS20, PMC24, RPK23, SRR^{+}21, WDC^{+}20, DBS^{+}22b, DP23, GC22, Lin24, MBB^{+}22, RTMCVMDE22]. Coastal [Bir20, BB20, CAGBM22, COM20, EVPMKP22, HM21, KWG23, O’D21, PFM20, RS24, SH22, YMG23, ALML23, AH23, Alf20, AC21, BOOO23, BG20b, BSM23, Bea22, BMR23, BASM22, BBGG22, CBO^{+}21, CDP22, CVH^{+}20, CQ20, CBD^{+}23, DC20, DJC^{+}23, DHC22, DSQ^{+}22, DB20, EdLP^{+}21, EMS^{+}23, FMGV20, FRVB22, FSF^{+}21, FAGS21, GCF^{+}23, GSA23, GPM20, GOH^{+}20, Gro20, HPC^{+}23, HC22, Hon20, HHEJB20, HBP^{+}21, HP23, IRK^{+}21, IGS^{+}23, JL21b, KS22, KKO^{+}22, KL21b, LSB^{+}21, LDAC20, cL22, LBWS23, LW22, LLCH21, LTJ^{+}23, LLSY22, Liu22, LLJZ22, LHPJ23, LRC24, MFG^{+}22, MKAF21, MKSG20, MS20, NNRS24, NAP^{+}21, NWZ^{+}21, NEH^{+}23, NBB^{+}23, OOGS20a, OOGS20b, OA20, PBS^{+}23, dlÁPAFGO20, PMR23, RCT20, Ree20, RZ21b, RJ21, RXX24, RGH^{+}22, RHL20, SA21, SVHV20, Sha20, SGK21, SBJF^{+}21, SCCW21, SBT^{+}23, SGZH^{+}21]. coastal [SGH22, SF22, SPA^{+}21, SNK^{+}20, SLK^{+}23, THT^{+}23, TLT^{+}22, TVN23, UOO20, VTRS20, DR22, VHSV20, WKO^{+}22, WJB^{+}20, WSD21, WLX21, WYD^{+}22, XWH^{+}22, YCL22a, YSWW21, Zha22, ZCG^{+}23, dSOW^{+}21, dSdLSGB22, BMR23, NAP^{+}21, TCT21]. coastline [OWCFDC21]. coasts [PBG^{+}22]. cockle [BASM22]. Coco [GASLCP20]. cod [BDR20, BFMS21, EAGV20, SGL^{+}23, SGZH^{+}21, SGH22]. Coda [HNMSG21]. Code [FFE20, KOBS21, Sin21]. cogens [Kuł22]. coherence [BTE21, PDB^{+}23]. coin [dOPM^{+}20]. Cold [MSGAR20]. Cold-water [MSGAR20]. Collaboration [HWY22, WP23, CdLN^{+}21, CS21, KMT^{+}21, MIMS^{+}20, MHEW20, TV21]. collaborative [CMS20, DSL^{+}21, DBS^{+}22b, HC22, LJ23, MMSB^{+}20, DB20]. collapse [Ber22]. collapses [VGC^{+}22]. collecting [GRS^{+}23]. collection [BMC20, CLL22a]. collective [BCB20, PN20, SVF20]. collegial [GR20a]. collision [KP24]. collisions [HSH^{+}21, SNP21]. Colombia [GMVM23, LZRC23, PCCOS24, PMFR^{+}22]. Colombian [MVR21a, MVR21b]. colonial [Bra22, LTRA20]. colonial-nesting [LTRA20]. colonization [Qaf24]. colour [SKL^{+}20]. COLREGs [KP24]. Columbia [SSB^{+}23, WJLSH21, DBS^{+}22b]. comanagement [Col20]. combat [BJH21, Men23]. combating [Wan23b]. Combatting [VVR^{+}22, HM22c]. combined [EMN^{+}24, KDS22, OCCSC23]. Combining [GMRSP23, MFG^{+}23, SVKP21, PBT^{+}20]. combustion [GZW^{+}20]. comes [FR23]. Commentary [LLM^{+}23, SGRW22]. comments [ML22]. Commercial [Shi22, AP23, BC20b, BDD^{+}21, CSGC22, DSB21, dlFGFFR^{+}20, GSG20, GWCD22, HKMS21, MRKM23, MCW^{+}22b, MGF20, MIH^{+}24, MHEW20, OCD^{+}20, SS22, SHP^{+}21, SBW^{+}21, SKW^{+}22, SPMN23, SO20, VÖB20a, VÖB20b, VSD^{+}23, VWJMAJ20, YTBM^{+}22]. commercial-recreational [GSG20]. commercially [RZ21a]. Commission [Kar23, SOT^{+}22, Abo23, GB21, Kuł22, SBS21]. Commitments [HBL^{+}21, Cha24, LBL21]. Committee [CR23]. commodities [KgJqY^{+}24]. ’Common [CAHR^{+}24, BBB23, DMG^{+}23, PV23, AELR^{+}20, BDDE20, Che21b, OKR^{+}20, SPG20, STS^{+}23, dS21, HBCA^{+}24, SPFA^{+}20]. commons [HBCA^{+}24, SII22, VHS^{+}20, Zha21]. communality [NGP^{+}21]. Communicating [AMK^{+}21, PBG23, TDC21]. communication [BDLO23, BMMF22, MBB^{+}22, MT20, Pau21, TFF22]. communitie [YMG23]. communities [BB20, DC20, DP23, Fro21, GCF^{+}23, GPTASM23, HHEJB20, HBP^{+}21, HC21, HM21, IRK^{+}21, IGS^{+}23, IAB^{+}20, JTR^{+}20, KHMM20, LSB^{+}21, Liu22, MKM23, MRD^{+}20, MTC^{+}23, NRYG20, NEH^{+}23, Pha20, RM21b, RGH^{+}22, SOP^{+}22, SSB^{+}22, SSP^{+}21a, SNK^{+}20, THT^{+}23, UOO20, VHS^{+}20, WKO^{+}22, WJLSH21, WNMS^{+}24, WGP^{+}21, dlPdlLLC^{+}22, CFM^{+}21]. Community [AAKI21, GLAVL21, HWM22, JY20, SAO^{+}21, BMM^{+}20, CFM^{+}21, Col20, CBD^{+}23, DBH^{+}21, DCC^{+}20, EMN^{+}24, GFD21, GE23, HFD^{+}20, HCWC21, LJRHH21, LLS^{+}20, MCBN22, MFM21, MM21b, NHD23, Ree20, RC22, SSZ^{+}24, SGA^{+}21, SCH^{+}24, SVKP21, SSF23, Wan23a, XT23, ZCZ20, Has23, MZ24]. Community-based [GLAVL21, Col20, DBH^{+}21, GFD21, HFD^{+}20, LLS^{+}20]. community-driven [BMM^{+}20]. Community-level [JY20]. companies [BHEE22, BK23, DSWB20, KLS22, LCL20, RAFLRP21, WW21a, ZWJvL20]. Company [HJLN24, FGPPPV21, HT23, LrS22, Ree20]. Comparative [AP23, RHL20, CSHO^{+}20, KGP^{+}22, LHW21, MBORGN24, MB21, NSZ^{+}21, NNRS24, OC20, QTD21, VVR^{+}22, ZTM^{+}23, dKMK21]. compared [GCC^{+}21b, SC21]. compares [HWU^{+}23]. Comparing [Cha24, dlFGFFR^{+}20, HW22, SW22, WJLSH21, ADR^{+}21, BBC24, PSPR20]. Comparison [KB20, KP24, ZWJvL20, GLAVL21, GMM^{+}20, PWL^{+}23a, TOHK20, Zha22]. comparisons [GVH^{+}23, LCAX21]. compatible [GLS^{+}21, JZW21]. Compensation [YTÇ22, HRR^{+}20, JF22, LYHL23, RXN23]. competing [BC21, Tea20]. competition [WP23]. competitive [ZT21]. competitiveness [BHEE22, DRCdMCJ21, KKDK20, ZCCQ22b]. complementary [OOGK^{+}24]. complete [Sin21]. Complex [MN21, DJC^{+}23, KLK23, KY20, LV23, MEG^{+}21, Tea20, TJ21b, VYB20, WCF^{+}21, XHZ^{+}21, YK21]. complexities [O’D21]. complexity [SRS23]. Compliance [BD23, BC20b, FMCD20, AL20, Dam24, DNR21, EHM^{+}20, EHM^{+}22, GDLC20, GRPE^{+}21, GHK20, Lin23, LFVCF24, MJ23, MJ20, MAK22, PMK^{+}21, PMAVM20, STS^{+}23, SOLN21, BC21]. compliant [WH20]. component [EAA^{+}21]. components [FZYF23, KBD^{+}20]. composition [BMC20, GÁPSAP^{+}23, PBG^{+}22, TSLK23]. comprehensive [AHBAKH24, AC23, GSG^{+}24, JCMD20, LWX^{+}20, Ren21b]. compromised [Mur21b]. Compulsory [Set23]. computable [WHPF20]. computer [VAMR^{+}20]. Con [KJE21a]. concentration [BADO20, FGPPPV21, SS22]. concept [RKS^{+}23, Wan23a, COJZ^{+}21]. concepts [JL21d]. conceptual [LLL23, Qi22, SWHT22, ZYTT21b]. Conceptualising [BE20c]. Conceptualizing [Day24]. concerning [BISSBdlMM22]. Concerns [CJ22, DGG22, AKFY22, DBB^{+}22, MFG^{+}23, MTC^{+}23, PZZ^{+}21]. concessions [VSD^{+}23]. conch [MVR21b]. conclusions [MGBS24]. Condition [MAG^{+}22, MKB21, VHSV20, vHSSK20]. conditioned [YMG23]. conditions [BLY^{+}24, BTE21, BvdH23, CDG^{+}21, Col20, GAMR20, HSH^{+}21, SOS^{+}21b, VM21, ZBF^{+}22]. conducted [KOBS21]. confer [FTM^{+}22]. Conference [DRY22, DČFW23, DC23, LLHG23, TZC23, Ton22a, Ton22b, TC23, War23, ABVPM20, DMNT20, MDJ^{+}22, MdCLL^{+}23, SCA^{+}22, TMDN23, Cha24, CLS20, CS20, HGG22c, LSH21, QYX21, RFPS21, SHZC21, YC21a, ZCZ20]. configurations [SCSR^{+}22, SCSR^{+}23]. Conflict [EV22, CLS20, CVA22b, Fen23, KHMM20, KL21b, MHN^{+}20, PMC24, STG^{+}20, TNY23, VGAH^{+}21, VLS^{+}23]. Conflicting [SPHO23, PFWM23, Gal21]. conflicts [CMGF20, KWS20, KOCL23, Li23b, MN21, ONA21, PMC24, dSOW^{+}21, GCL21]. Conformity [WKO^{+}22]. confronting [GE23]. Congo [Kas21]. Connecting [MWM^{+}23b, ADB^{+}22, YS21]. connections [KKA^{+}21]. connectivity [Jen21, KTC^{+}22, Lot22]. consciousness [MCL22]. Consensus [TYS24, EV22, GV22, MDJ^{+}22, TMDN23]. Consequence [BVT^{+}21]. Consequences [BNF21, BFMS21, HvT22, JPM^{+}23, Kiz24, Kuł22, dS21]. consequentiality [XLLS21]. Conservation [FMWW21, GB21, RDT^{+}23, SLP^{+}22, YC21b, AAA^{+}24, AAKI21, AML^{+}21, AVF23b, AAFK^{+}21, AGSI22, AdSMS21, BC20a, BMB^{+}24, BUR^{+}22, BGK^{+}22, BBHN20, BMB^{+}21, CMS20, CHB^{+}20, CMWN^{+}23, COSS21, CTC^{+}23, DH22, DBH^{+}22, DPG^{+}22, EAA^{+}21, EAA^{+}22, EABA^{+}22, FWGM24, FMS^{+}22, GNJR24, Geh23, GBB^{+}21, GMGLV21, GMGLV22, GEB^{+}21, Har22, HMG^{+}21b, HDAPG^{+}24, HGG22c, How20, HCT^{+}21, Kai21, Kar22c, KSB^{+}23b, KHMM20, KGL22, KVL^{+}23, LDL21, LMC^{+}23b, Liu20, LM24, MFT^{+}22, MBDG23, MJ20, Mba22, McC20, MWM^{+}23b, NGP^{+}21, OT23, OWCFDC21, OVSVPP^{+}22, PCG^{+}23, PBG23, PT23, PC20, Pri22, RPK23, RSS^{+}20, SEG^{+}22, SMW^{+}21, SWHT22, SGA^{+}21, SBDC24, SLW^{+}23, SVKP21, SLH22, Teo23, Tho22, TEA^{+}22, TMTC20, TMS^{+}22, VJL^{+}21, VPB^{+}24, WKO^{+}22, WLCG23, Mur21b, TBPC20]. Conservationist [Mar20a]. conserve [Ber23]. Consideration [FAC^{+}23, MMNC22, OCT^{+}22]. Considerations [FRVB22, KFS^{+}22, Sar23, CMD^{+}20, HKKY22, JACD20, LSD^{+}22, OI23, THGR^{+}20, VMC^{+}21, WCC^{+}20]. considered [Pha20]. Considering [VHS^{+}20, ONA21, OKH^{+}24, SKE22]. consistency [ABM21, CA20, MPD^{+}23]. consistent [CMR21]. consolidation [SS22]. constabulary [Day24]. Constantinople [Qaf24]. constraining [SGL^{+}21]. constraints [BTV^{+}22, WH20, Niu22]. construct [XXMZ23]. Constructing [MCL22, CKY20]. construction [Bau20, CLZ23, Los23, SHZC21, YLSC23]. Constructive [GR20a, CMSN^{+}22]. consultants [OOGS20a]. consultation [PVH^{+}21]. Consumer [WMS21, CMR21, CLCV21, DLYM22, HT23, MLM23, WM21]. consumerism [HWMO20]. consumers [AAA^{+}24, HWMO20, Nal24]. consumption [BABC20, BBHN20, HKMS21, HDAPG^{+}24, HM21, LLW22, MFF^{+}23, NGA23, YFD^{+}21, ZY21]. consumption-based [LLW22]. contagious [Bov22]. container [BHEE22, COJZ^{+}21, JHL21, KKDK20, KMM21, KABT21, NCD23, Okş23, SSK23, WW21a, ZLWY22]. containing [Lev20]. containment [SOK^{+}22]. contaminated [LWZ23, MXZ23, WL22a]. contamination [WM21]. contemporary [HLW^{+}20]. contested [PFM20]. context [CK23, FGPPPV21, GTTM21, HZ23, JF20, LDAC20, RRR^{+}23, Wat22]. contexts [MAK22]. continental [CHEC21, KS22, MFRQ23, MB22, dS21, dS22a]. continents [CMGF20]. contingent [CKY20, LrS22, ZMR20]. continue [dMLB^{+}20]. continues [RE23]. continuous [VRBC^{+}21]. contract [LC20a]. Contractor [Fei21]. contribute [BBHN20, PMN^{+}22]. contributing [LGF24, RE23]. Contribution [LRN^{+}22, NRYG20, RGH^{+}22, GCF^{+}23, dlFGFFR^{+}20, Gul23, KL22, WDC^{+}20, ZCCQ22a, ZT21, dSdLSGB22]. contributions [BBC24, Ege23, Moh23, PJN24, QUK21, VCC^{+}21, ZT23b]. Control [DNR21, Squ21, CC21, CWR20, GML^{+}24, HSdT21, JWX^{+}20, KLS22, LMT21, MX21, Mot20, Mot22, SGB21, Şan20, Tes20, WZZC21, YZG23, Kar22a, XWL^{+}20]. controlled [HACCM22, KZR22, Stø21]. controlling [Zhu23]. controls [BSSL^{+}20]. Controversial [CNB^{+}23, OOGS20a]. Controversy [KL21b, HM21]. convenience [Fel22, PSMR20]. Convention [CK23, GB21, SOT^{+}22, WL20, KS22, KGW20, LC23, RCH22, CCX23, Gul23, MG22, RPY^{+}24]. conventional [BMB^{+}21]. Convergent [JL21d]. cooperate [LWZ23, WL22a]. Cooperation [EB23, SWGZ22, VGAH^{+}21, FOKM23, GSGOP^{+}21, GBCNCL^{+}21, GY23, Hsi20, JF24, POO23, RCT20, SRY24, TCZ21, Tel23b, Ton22b, WRL23, XHZ^{+}21, YC21a, ZFGL20, ZLGX20]. cooperative [BLM24, MHEW20, TSS^{+}24, WJLSH21, ZHL^{+}23, BDD^{+}21, PSTW21]. cooperatives [Sol23b]. coordinating [SZBA20, YK21]. coordination [HCT^{+}21, LSH21, LHPJ23, MMSB^{+}20, TCZ21]. cope [GTMP20, RKR20]. Coping [GRA^{+}22, ZTZ^{+}21]. Coral [CLRW22, RBEA^{+}22, SVF20, AHN^{+}23, BDMW22, BS22, CWB^{+}20, DDAB^{+}23, Dun21a, Dun21b, GA20, GN22, HVS22, HF20, LRD^{+}22, Lev20, McC20, McC21, MDG^{+}23, NTH^{+}21, SKE22, SPCCvT21, YFD^{+}21]. corals [AKFY22, MSGAR20]. cordatus [dFVCGdS21]. core [EdLP^{+}21, HPS23]. Cork [OGGC20]. coronavirus [Gor21, IKB21, SFB24]. corporate [RFG24, UOO20, WS23]. Corrected [LLM^{+}23]. correlated [JY20]. Corridors [Wan23c]. Corrigendum [EHM^{+}22, GZW^{+}20, Mal21, Mot22, RP23a, SCSR^{+}23, SDG^{+}22]. corruption [Gar24]. Coryphaena [MAG^{+}22]. Cost [MPL^{+}23, AH23, CdLN^{+}21, JV20, KABT21, OAM21, SC21, SBW^{+}21, WBH^{+}23, ZMR20, ZNA^{+}23]. cost-benefit [AH23]. cost-effective [CdLN^{+}21]. Costa [PSPR20, RJAT23, Teo23]. costly [BXV20, COJS^{+}22]. costs [BBB^{+}20, BvBN^{+}20, CCGM21, CvHF^{+}24, CBAPM20, DD22, NSPRSS20, RLRS21, VGK23]. cotton [PBG^{+}22]. could [GLS^{+}21, SLMG^{+}21]. Council [BDC23, BI21, JAK^{+}23, LNRM23, SK20a, LJ21, MPH^{+}23, NRF21, RCT20, RROP23, Sta23b, TOL^{+}22]. count [KPN20]. counter [Sel20]. countermeasures [LYHL23, SOK^{+}22, YWY20]. Countries [OAAMS24, AB20, AGA^{+}21, BHdMCM^{+}21, GLAVL21, IGO21, IKB21, Kak20, LCC20, MAA^{+}24, NRF21, PMK^{+}21, SF22, TTA^{+}20, dKMK21, MZ24]. country [Gar22, Her22, HLO^{+}21, MCKZ21, Shi20b, TTA^{+}20]. country-level [MCKZ21]. County [LA23]. coupling [LHPJ23, WSD21]. Cournot [Han20]. course [BH23a, BB21]. court [NT20, Pen22]. courts [Pat22]. coverage [SMW^{+}21]. COVID [AEL^{+}23, ASGZ22, BRM^{+}21a, BSI^{+}21, BASM22, CDNS23, CCKK22, CRCP^{+}23, De 21, DME23, Ero22, FTM^{+}22, FGPPGG22, GPTASM23, GABSD21, GNMGWM21, HDCH21, HHB^{+}21, HBP^{+}21, Isr22, KFP20, KAC22, Kob22, LWSS21, LSB^{+}21, LBL21, LC20b, LEERRM^{+}21, LKE^{+}22, MCBN22, MAK22, NK21, NGRM^{+}22, NAS^{+}22, PGK22, PAK22, PBD^{+}23, RDD^{+}21, SVV24, SSB^{+}22, SCH^{+}24, SdCMN^{+}22, SCLM^{+}21, SSP^{+}21a, SNK^{+}20, SSP^{+}21b, Tan22, TAK^{+}22, VMM21, WHL22, WPWM^{+}22, XSCL21, YCBW21, ZTZ^{+}21, ZTW23, ZTM^{+}23, ZT23a, ZCS^{+}23, ABME22, GMCMJA^{+}21, GS21, MÇ23, MMK^{+}22, NVP^{+}23, SRH^{+}22, VDP^{+}22]. COVID-19 [AEL^{+}23, ASGZ22, BRM^{+}21a, BSI^{+}21, BASM22, CDNS23, CCKK22, CRCP^{+}23, De 21, DME23, Ero22, FTM^{+}22, FGPPGG22, GPTASM23, GNMGWM21, HDCH21, HHB^{+}21, HBP^{+}21, Isr22, KFP20, KAC22, Kob22, LSB^{+}21, LC20b, LEERRM^{+}21, LKE^{+}22, MCBN22, MAK22, NK21, NGRM^{+}22, NAS^{+}22, PGK22, PAK22, PBD^{+}23, RDD^{+}21, SVV24, SSB^{+}22, SCH^{+}24, SdCMN^{+}22, SSP^{+}21a, SNK^{+}20, SSP^{+}21b, Tan22, TAK^{+}22, VMM21, WHL22, WPWM^{+}22, XSCL21, YCBW21, ZTZ^{+}21, ZTW23, ZTM^{+}23, ZT23a, ZCS^{+}23, ABME22, GMCMJA^{+}21, GS21, MÇ23, MMK^{+}22, NVP^{+}23, SRH^{+}22, VDP^{+}22]. COVID-19-affected [LWSS21]. COVID’19 [Kiz24]. crab [ADR^{+}23, dFVCGdS21, DD22, Gor21, KKA^{+}21, KV20, LMC^{+}23a, LJRHH21, RKR20, VKN20, YSEK22]. crabs [MMB^{+}23, OCZ^{+}22]. Crafting [GdACT20, Wil20]. crass [NW20]. Creating [RB20, CMWN^{+}23]. creation [GV22, SOK^{+}22]. credit [AACdlT23]. crediting [KWY^{+}22]. crew [CSGC22, De 21, HCW^{+}23, YS23]. crews [Pau21]. crime [Álv23, BE20c, Gar24, PGP22, SKL23, Urb22]. Criminal [Tei23, Álv23, Pat22, Ros22]. criminalise [Faj22]. crises [BMH^{+}22, HSM^{+}23, LCL20]. crisis [AVF23b, BVvKL21, De 21, FGL^{+}21, Hon20, McC20, Tan22, WKBS20]. criteria [ASV^{+}21, ABM21, CPBO20, EAGV20, GCG^{+}21, HW22, OCCSC23, SLK^{+}23, TMS^{+}22]. Critical [BL23b, CLP21, DL22, LTJ^{+}21, MFG^{+}22, NAP^{+}21, Åm21, CLD^{+}24, EAA^{+}21, PMN^{+}22, SRvNN21, SMDM20, SM23, YT20, ZSLY20, ZLY23]. critically [GPH^{+}21]. critique [CMSN^{+}22]. Croatian [PBD^{+}23]. crocodile [KHMM20]. Croix [GMRSP23]. Cros [HNMSG21]. Cross [Tel23b, GSGOP^{+}21, IO23, MMK^{+}22, PWL^{+}23a, SL23]. cross-border [GSGOP^{+}21]. cross-national [SL23]. cross-regional [PWL^{+}23a]. crossing [LLT^{+}20, Tse21]. Crowdsourced [SR22]. Cruise [ZCSK23, Kiz24, LMT21, LZZ22, LLHG23, LC20b, Set23, SCY24, WCC^{+}20, YCBW21, ZCS^{+}23]. CSR [UOO20]. Cu [KJE21b]. Cuba [AVPAFM^{+}22]. cucumber [GÁPSAP^{+}23, HFD^{+}20, Pra20]. cultivation [BvdBvdFK20, Chr20, GBLBV20, YZH22]. Cultural [GCL21, LC23, MWWS21, WM22a, War23, BEH^{+}21, DBH^{+}22, FFYE21, GW21, KOCL23, LHW21, LKB^{+}23, Lin24, ML22, PT23, QK24, SSA^{+}23, SPHO23, SRS23, ZCG^{+}23, HVS22]. culture [Lin23, MCL22, MWM^{+}23b, SL23]. cumulative [OCT^{+}22, SBT^{+}23, HCWC21]. cure [LC20a]. curfews [MAK22]. Curieuse [COSS21]. curled [BBB23]. Current [BLY^{+}24, CLS20, CLZ23, JTR^{+}20, Kul23, LYHL23, LSD^{+}22, OKS^{+}22, QLZ^{+}24, SKW^{+}22, TYS24, VSV^{+}23, VCC^{+}21, WNDZ23, WM22b, ZMM^{+}23, BB21, BHKJT23, BRM21b, CCB^{+}21, DL21, DFC^{+}20, Kai21, MX21, NGT^{+}23, PCG^{+}23, TAK^{+}22, WSDB20, GBC^{+}20, Kas21]. customary [Kar22c, Mba22, SRH^{+}22, ZY23]. cuttlefish [GPM23]. CVM [LrS22]. Cyber [Lv20, SM23]. Cyber-seaworthiness [SM23]. cybernetics [SOS20]. cybersecurity [Kar22b]. cycle [RS24, YS21]. cyclone [MDLA21]. cyclones [SNK^{+}20]. cylinder [GMM^{+}20]. Cynoscion [dSJNS20].

dagaa [IOHC22]. daily [MLM23]. Dakar [CDNS23]. damage [Fei21, JF22, SL20, Tel23b, ZMR20]. dams [DDGdS^{+}21]. dangerous [RE21]. Danish [HVN21, NHA23, NSK20, SII22]. Danish-flagged [HVN21]. Danube [Tea20]. Dao [KJE21a]. Darwin [TOAPJ21]. dashed [Gul23]. Data [VBC^{+}20, Alm21a, ASM^{+}21b, ATF^{+}21, AGA^{+}21, BMC20, CRP^{+}20, CLS^{+}23, CPBO20, CLL22a, CDB20, CNG^{+}23, DBS22a, DGMF^{+}20, FTW^{+}23, Fuj21, GRS^{+}23, GR20a, GCS21, GAD^{+}20, HMG^{+}21a, KMG^{+}20, KPN20, LWSS21, LHV^{+}20, LLLF22, MLJ^{+}22, MFG^{+}23, MMNC22, MAG^{+}22, MSGAR20, PBS^{+}23, PWL^{+}23b, RZ21a, Şan20, SRR^{+}21, SCD21, STIS20, SSOP^{+}21, SWH^{+}23, SSMG21, TLT^{+}22, TK23, WCF^{+}21, XSCL21, XWH^{+}22, YTG^{+}23, ZCG^{+}23, vSHHT20, PCD^{+}23]. Data-driven [VBC^{+}20, GR20a]. data-limited [CDB20, GCS21, PWL^{+}23b]. data-poor [Alm21a, PBS^{+}23, YTG^{+}23]. data-scarse [CPBO20]. date [GJS22]. Davao [MLMD21, MMC^{+}22]. David [Smi22, Smi23a]. day [BSI^{+}21, Fra23]. de-privatization [CAHR^{+}24]. DEA [PCD^{+}23, FMGV20]. DEA-based [FMGV20]. deadline [Sin21]. Dealing [MSGAR20]. death [CJ23]. debate [BE20b, MT20, SS21a]. debates [Wan23b]. debris [SONB^{+}22, DT23, HLCM21, HWY22, SGLF^{+}23, Wu22, XXMZ23]. Decade [MZ24, Pol23, DG24b, VMC^{+}21]. decades [Ehl21b, GV22, KKO^{+}22, LJ20, ZV20]. Decarbonisation [Ada22, vLM22, BH23a]. decarbonization [Dom23]. deceived [LBL21]. December [Ano20b, Ano21c, Ano22c, Ano23c]. Decent [LFLF22, LSD^{+}22, YTÇ22]. decentralisation [JL21c]. decentralization [Sak21]. decentralized [MMSB^{+}20]. decides [Pre21]. Decision [GSG20, AMR21b, BL23a, BBBS22, CDB20, DJC^{+}23, DBBL21, FRM22, GCP^{+}20, GSM21, HDH21, HDCH21, HW22, JKV^{+}22, KVL^{+}23, MXZ23, MWM^{+}23b, MKB21, MMGAA^{+}21, PHF20, RB20, TDC21, Wat22]. decision-makers [RB20]. Decision-making [GSG20, AMR21b, BL23a, BBBS22, DJC^{+}23, FRM22, GCP^{+}20, GSM21, MWM^{+}23b, MKB21, MMGAA^{+}21, TDC21]. decision-support [CDB20]. decisions [BC20b, LMC^{+}23a, MKSG20, Sta23b, SN23, ZCG^{+}23]. deck [KZR22]. decline [Alm21a, ALLR^{+}24, Fla21, IAB^{+}20, OOSG21]. declines [SKS21]. declining [BB21, SSH^{+}23]. decommissioning [TN21]. Decomposing [Seu22]. decomposition [LLW22, ZZD^{+}23]. decoupling [WW22]. decreasing [ZD20a]. deemed [TW22]. Deep [CHEC21, Kak20, Wil21a, AGM^{+}22, AFHHD22, AR22, BDCJ22, BHKJT23, BRM21b, CDD^{+}23, CGHV21, CDC21a, CDC21b, CL21, Cra21, DFC^{+}20, Fei21, GG21, HSSM24, HSK^{+}23, HSHS24, Jae21, JACD20, MM21b, MSGAR20, OKR^{+}20, OVA22, PL24b, RFPS21, Squ21, SCA^{+}22, TML^{+}21, VJL^{+}21, WLCC23, Wil20, Wil22, vdGD21, KGW20]. deep-ocean [AFHHD22]. Deep-sea [CHEC21, Kak20, AR22, BHKJT23, CDD^{+}23, CDC21b, HSSM24, HSHS24, JACD20, MSGAR20, OKR^{+}20, PL24b, VJL^{+}21, WLCC23, vdGD21, KGW20]. deep-seabed [AGM^{+}22, HSK^{+}23]. deep-water [BDCJ22]. deepest [VJL^{+}21]. defects [CLL^{+}22b]. defences [COM20]. defend [HPG20]. deficiency [MÇ23]. deficit [OCD^{+}23]. deficits [CGP23]. defined [MKSG20, VvSvH^{+}22]. Defining [JG20, CVA22b, QCM^{+}20]. definition [GA20, Urb22]. Definitions [SRH23]. degazettement [ACA^{+}21]. degrade [PK23]. Déjà [SBT^{+}23]. del [GASLCP20, MMK^{+}23, MNGM24]. Delayed [AMR21b]. deliberate [VJL^{+}21]. Deliberating [HVS22]. Deliberative [BBGG22, SL20]. delimitation [Mar21a]. Delimiting [PMR23]. Delineating [AR23b]. delivering [Tel23a]. delivery [RFG24]. Delphi [HRS^{+}24, LZL23, MBDG23]. Delphi-AHP [LZL23]. Delta [BSBD21, dFVCGdS21, NHD23, Tea20, QYX21, UOO20, WHPF20]. demand [BHdMCM^{+}21, CLW^{+}21, DHC22, HCSBM^{+}23, Kaw22, Tan22, TOAPJ21, TCC^{+}22]. DEMATEL [DL22]. demersal [GMM^{+}20, LSGG^{+}20, NHA23, OCCSC23, SDL^{+}22]. democracy [AELR^{+}20, AE23, BBGG22]. demography [GS21]. demonstration [AC21, CKY22, RXX24]. Denmark [NAN^{+}24]. density [PHF20]. dependence [TUBA22]. dependency [LP20]. dependent [Ber22, IRK^{+}21, MAK22, WL22b]. deposition [RB22]. deposits [JG20, VCC^{+}20]. depredation [CRJ22]. deregulation [GWCD22]. Derelict [ABDRV^{+}23, DT23, ADR^{+}23, DSB21, Lov23]. derived [Öİ24, SOLN21]. desalinated [LYQM21]. description [MMS24]. design [Dun24, Fen23, MSZ^{+}23]. designation [CMWZ24, Cho22, KOCL23, Tes20, WD20]. Designing [DN20, OOGA22, RBB^{+}24, ABM21, JZW21]. designs [KvTNL20, SNP21]. desirability [PSMR20]. despite [HSM^{+}23, SSH^{+}23]. destinations [LLJZ22]. destructive [SBFHD^{+}23]. detail [MDJ^{+}22]. detect [GAD^{+}20]. detention [CLL^{+}22b]. detentions [MÇ23]. Determinants [SZ20, WW21a, CACJR^{+}21, HPS23, Kao23, KPN20, KJS^{+}22, LCC20, ZZD^{+}23]. Determination [Kar22a, BRR^{+}22]. determined [GÁPSAP^{+}23, MGS21, Moh23]. Determining [ADR^{+}23, Öİ24]. developed [HHH^{+}23, TTA^{+}20]. Developing [CC23, CMT^{+}23, CTIH21, IGS^{+}23, Moh23, SGL^{+}21, CdLN^{+}21, Gun21, GML^{+}24, IKB21, KMS20, MAA^{+}24, MCW^{+}22b, NSN22, NW20, NRF21, ONA21, RLRS21, SAA^{+}21, TTA^{+}20, VSV^{+}23, WH20, BGLTC21]. Development [ASGS23, BCN23, DQJ23, MZ24, RXN23, Run24, WM22b, vSHHT20, Alf20, AVPAFM^{+}22, Aqu23, BLS21, BTE21, CPV^{+}21, CA20, CLZ23, CLW^{+}21, CJL^{+}23, COSS21, DKW20, DČFW23, Ege23, EMN^{+}24, EVPMKP22, EMS^{+}23, FXC^{+}22, GLAVL21, GMG^{+}23, GW21, HS20, HFD^{+}20, Han21, HPH22, HSK^{+}23, HSN^{+}21, HC21, HWY22, IMA22, JC23, JL21a, JLY23, JD20, JV20, KWF^{+}21, Kim20, KWMM20a, KGP^{+}22, LHTR23, LHV^{+}20, LZZ22, LLHG23, Li23a, LLLF22, LJ22, LJL22, LLL^{+}24, ML21, MF22, MG22, MPH^{+}23, MMNC22, MBC^{+}24, MPSF20, Mur21b, NSN22, NSPRSS20, PMN^{+}22, PDJ22, QZZ20, QYX21, RA20, RJ21, Ren21a, RXX24, Sak21, SdACdSLdAS24, SVHV20, SWHT22, SGWL21, SHZ^{+}21, SSZ^{+}22, THP22a, Ton22a, Ton22b, UOO20, VÖB20a, VÖB20b, WLY21, WYD^{+}22, WNDZ23, WA24, WdSRY23, YK20, YSWW21, YZH22, YC21b, ZWG^{+}21, Zha22, ZCCQ22b, ZZW^{+}21, ZLZ21, ZWC^{+}23]. development [Zho23, ZT23b, LMSLMQ21]. Developments [Qaf24, AE23, AR22, Fra23, GAD^{+}24, Guo20, MX21]. deviation [RE21]. device [RQB^{+}20, WLRH^{+}20]. devices [DPB^{+}23, EMH^{+}23, MKM^{+}20, MM23, ZMM^{+}23, CMT^{+}23, MSZ^{+}23]. diagnose [STKC22]. diagnosis [CES20]. dichotomy [KLSG21]. did [Kim20]. die-off [SOP^{+}22]. Diego [PB20]. difference [RM21b, WZZC21]. difference-in-differences [WZZC21]. Differences [BMB^{+}21, MÇ23, SCH^{+}24, BDDE20, JLY23, PSPR20, WZZC21]. Different [HC21, Gue23, HRS^{+}24, Her22, MMW^{+}23, PMK^{+}21, RHST21, VPRR^{+}21]. Differential [CRCP^{+}23]. differentiated [Che21b]. difficulties [LL24]. diffusion [ZLPT23]. Digital [LR20, HPH22, Jen21, LEERRM^{+}21]. Dikkowita [JAH^{+}24]. dilemma [RSAP22, SVF20, Zha21]. Dilemmas [GY23]. dimension [KBD^{+}20, MCW^{+}22a, RSAP22, Cha20b]. dimensional [LRD^{+}22]. dimensions [BDM20, Har22, MK24]. Diomede [Tod22b]. dioxide [GZW^{+}20]. diplomacy [Geh23, Hon21, TLYC20]. Dipturus [GPH^{+}21]. Dire [Tod22a]. direct [AED20, BBC24, NTH^{+}21, SRSGMGG21a]. Direction [MDJ^{+}22]. directional [OAM20]. directionality [BBGHR23]. directions [Fuj21, NJBM^{+}23, PSD23, RXN23, WNDZ23]. Directive [GKP^{+}22, PDB^{+}23, BN20, Puh23, ANS^{+}20, MPD^{+}23, MFRQ23, Mor20, PKB^{+}21, VRBC^{+}21, VBH^{+}23]. Direnzo [Lv20]. disadvantages [Jen23]. disappearance [CJ23]. Disaster [IGS^{+}23, SWF^{+}24, DMd^{+}20, KWG23, LDAC20, OGS24, SOP^{+}22, dOSTB^{+}20, YSWW21]. disasters [FAGS21, RMBR22]. Discard [GMPT^{+}20, PBT^{+}20, ABVPM20, GTA^{+}20]. discarded [GRS^{+}23, GHW^{+}22, GAD^{+}23, HGG^{+}22a, Lov23, NSK20]. discarding [VJL^{+}21]. discards [BMGDS^{+}20, CBR21, GMM^{+}20, SG20c]. Discharge [PL24a, LWZ23, XTY22]. disclosure [ZLWY22]. discolorations [SCA^{+}20]. discounts [PYVF24]. discourse [SG21, SH22]. discourses [Los24, SPCCvT21, Tea20, VMD^{+}24]. discovered [WWG^{+}23]. discrete [CDD^{+}23, TOAPJ21]. discretionary [OAM21]. discussion [CMWZ24, Dua22, JL21c]. discussions [HGSD23]. disease [LMT21, MAK22]. diseases [Bov22]. Disentangling [HTP^{+}23, Mil22, SBS21]. disincentives [BC21]. disparities [GWCD22]. Displacement [SKS21, Tru23]. dispossession [Bra22]. dispute [HJPL23, PL24a]. disputed [Cha21, SCGC20]. Disputes [Smi23b, CG22, HGLR22, NT20, Øst20, PK20b, Shi20a, Zou23]. Disrupting [JD20]. disruptions [IGS^{+}23]. disservices [RFG24]. distance [FMCD20, OAM20, vRB24]. distant [HCS21, Shi23b, YS23, YL21, Zho23]. distinct [Yu22]. distraction [WWD23]. distress [AONF21]. Distribution [AKFY22, AMIM21, AGN^{+}20, AS23b, Che21a, GASLCP20, HRC^{+}24, QF20, SSV^{+}23, TZM^{+}22, TV21]. distributional [Muk21]. distributions [CD24]. District [MKAF21, TCT21, YC21b, CDP22]. disturbances [Dam24]. disturbing [WWCS21]. diver [ALLR^{+}24, LM24]. divergent [JL21d]. divers [AHN^{+}23, HMG^{+}21b]. Diverse [BLS21, DSL^{+}21, DCC^{+}20, SHP^{+}21]. diversification [SBD^{+}23]. diversify [CWMM^{+}21]. diversity [FMS^{+}22, HM22a, HMG^{+}21b, LGF24, RLS20, WvdMG^{+}21, MG22]. divide [LEERRM^{+}21]. Diving [AAFK^{+}21, CACJR^{+}21, GMGLV21, GMGLV22, HACCM22, MMO^{+}24, SMHPMO^{+}21]. DNA [BdLAdC^{+}20, CS21, GCC^{+}21a]. Do [JL21b, THT^{+}23, ALML23, BEM^{+}20, MS21, Øst20, RJ21, WLQD21, WYD^{+}22, BHEE22, CPR21]. dockside [KMG^{+}20]. Documenting [DLJ^{+}20, LA23]. documents [AKFY22]. Does [ÁCT20, ADB^{+}22, CLL20, HPH22, HvPPV20a, LWG23, LJL22, MAA^{+}24, NAN^{+}24, UOO20, GSA23, PBG^{+}22, QZZ20, SMW^{+}21, ZLZ21, RHC20, TJZ^{+}21]. doesn’t [SCF^{+}22]. Doing [MFG^{+}22, GB21]. dolphin [STS^{+}23]. dolphinfish [MAG^{+}22]. Dome [RJAT23, Teo23]. domestic [Cha22, LMC^{+}23b, NNR^{+}21]. dominance [PBG^{+}22]. Dominica [SWF^{+}24]. Dominican [PSPR20]. Donax [SVP^{+}23]. Donggang [LSY20]. Dongsha [YCL22b]. don’t [BEM^{+}20]. double [Gug24]. double-edged [Gug24]. doubled [SLL^{+}24]. doubloons [LHW21]. down [AFMdSdTdCC22, RC22, SLB21, SMF^{+}23]. downgrading [ACA^{+}21]. downs [CSQ^{+}24]. downscaling [MMNC22]. downsizing [ACA^{+}21]. DPSIR [CBO^{+}21, QUK21, QLU^{+}23]. draft [CMSN^{+}22, SK20b, ML22]. dragon [AWX21]. Dragons [LHW21]. drawing [QZZX20]. dredged [SOK^{+}22]. dressing [Lot23]. drifting [CMT^{+}23, EMH^{+}23, MZT^{+}23, ZMM^{+}23]. drilling [JF22]. driven [BMM^{+}20, For22, GR20a, LTJ^{+}23, QRK^{+}23, VBC^{+}20]. driver [JDF^{+}23]. Drivers [Ken20, VLS^{+}23, WLXY22, CCCC21, GHW^{+}22, JAdS^{+}20, KL21b, LXH^{+}22, MGS21, SDI^{+}20, SCBB^{+}22, SRH23, THS^{+}24]. drives [SRLLG23]. Driving [DLX21, FZYF23, TJ21a, WFG22a, ZWXL21]. drone [SAB^{+}20]. drones [Kle21]. drop [GCC^{+}21a]. Drumhiller [Lv20]. dual [Cha22]. due [Gor21, GWAPRPRU22, MS20, SKS21, SLMG^{+}21]. dumping [CMWZ24]. Dungeness [ADR^{+}23]. duplication [CCB^{+}21]. Duration [LCC20]. during [ASGZ22, BvdBvdFK20, Ber22, Cha21, hC24, EHLT21, FTM^{+}22, GS21, Isr22, JG23, MÇ23, NGRM^{+}22, SPJB22, SOS^{+}21b, TAK^{+}22, VDP^{+}22, WPWM^{+}22, XWH^{+}22, ZTW23, ZTM^{+}23]. Dutch [CWKE20, HvdBH23, SZBA20, SVKP21, vHSSK20]. duty [Dev22, LWZ23, NOB22, PMFR^{+}22, WL22a]. dykelands [ZCG^{+}23]. Dyking [dZ20]. Dynamic [HSH^{+}21, MWB^{+}21, CMD^{+}20, FDGM24, FWL^{+}22, Ren21b, YK21]. dynamics [AH21, BDM21, CDHY22, GFD21, JAdS^{+}20, Kas21, MMK^{+}22, PGSFRMHH24, SN23, YMG23].

e-logbooks [CLS^{+}23]. Early [SK24, SNK^{+}20, WCZS21, YC21b, INW20]. Earth [VJL^{+}21]. Earthscan [Smi23a]. East [Gra21, GC22, PGP22, PK20b, ADR^{+}21, ÁFFSC20, FFE^{+}24, HGLR22, SRR^{+}21, TLYC20, VBH^{+}23]. Eastern [CDG^{+}21, YMG23, CLRW22, CDP22, PDJ22, WWG^{+}23, DMS^{+}24, LTW23, LTRA20, Lov23, Mah22]. EBM [GCP^{+}20]. eBook [Sta23a]. Ebro [BSBD21]. ECAN [MFG^{+}22]. eco [Abo23, AH21, Jon21, JAK^{+}23, RFG24, WWC23, DČFW23]. eco-certification [RFG24]. eco-label [AH21]. eco-labelling [JAK^{+}23, WWC23]. eco-scalar [Abo23]. ecolabels [CGGA20, GCV^{+}21]. Ecological [AIG^{+}20, ABM21, DG20, EMN^{+}24, FAC^{+}23, FBC^{+}23, GCS21, LDL^{+}23, TZC23, AE21, Alm21a, AELR^{+}20, BC20a, BMK^{+}22, CAVJ^{+}20, CVA22b, DFC^{+}20, DBS22a, DRY22, DL22, Dun21b, FAGS21, GABA^{+}23, GTMP20, HRR^{+}20, HMG^{+}21a, JTR^{+}20, JF22, JMB^{+}20, KKA^{+}21, Kar22c, LMC^{+}23a, LYHL23, LYL^{+}22, MMB^{+}23, NGT^{+}23, OCCSC23, PBS^{+}23, PWL^{+}23a, QS21, RXN23, RXX24, RDLA^{+}20, SOO20, SRR^{+}21, SBD^{+}23, SCCW21, SVP^{+}23, STF21, VGC^{+}22, WLX21, WYD^{+}22, XXLW24, NAP^{+}21]. ecologicalization [XHY22]. ecologically [KVL^{+}23]. ecology [dSJNS20, WCZS21, WCF^{+}21]. ecology-economy [WCZS21]. Economic [BDR20, Cha20a, DMS^{+}24, FY24, GMGLV22, JGWM^{+}23, Kiz24, LrS22, PV23, QTPL21, SGWL21, SOLN20, SPA^{+}21, TM23a, AB20, AE21, ADR^{+}21, AVPAFM^{+}22, AVF23b, BBC24, BDCJ22, CMC22, CDHY22, CLP21, DGG22, DRCdMCJ21, DAAP21, DME23, Ero22, EB24, EAA^{+}21, FGPPGG22, FGPPGG23, FFYE21, FMLMM23, Fro21, GLADGGVL23, GASLCP20, GWCD22, HFD^{+}20, HBP^{+}21, JL21a, JK22, JKKT20, KKA^{+}21, LP20, LXSX21, LW22, LXW^{+}21, LFPBD^{+}20, LRC24, Ma20, OKFN24, PKR^{+}22, PALB20, PDJ22, Puh23, Qi22, Qi24, RGK^{+}20, SBD^{+}23, SCBB^{+}22, Sha20, SBW^{+}21, Squ20, SRSGMGG21a, SRSGMGG21b, TMS^{+}22, TOHK20, VSD^{+}23, VPRR^{+}21, WDG^{+}20, WCC^{+}20, WW22, WDC^{+}20, YSWW21, ZFGL20, ZT21, ZNA^{+}23, ZCA^{+}23, dOPM^{+}20, AAFK^{+}21, GWAPRPRU22, NAP^{+}21]. economics [BHCL23, DMS^{+}24, Gro20, TW22]. economies [AED20, DCC^{+}20, dlFGFFR^{+}20, LRN^{+}22, RLRS21]. Economy [TFMT^{+}23, AAA^{+}21, Als20, BG20a, BBWC21, CS21, DKKS21, FFD^{+}23, GdACT20, Has20, HSN^{+}21, LXSX21, LWG23, LHPJ23, Lou22, Mal21, MFC^{+}24, PSD23, QYX21, RJ21, RXX24, RRM^{+}24, SdACdSLdAS24, SHPW21, SGK21, SR22, SPH^{+}20, Tho22, TT22, WLY21, WCZS21, WNDZ23, ZZD^{+}23, ZLH21, AC23, CFM^{+}21, LMMT23, SGWL21, WMKG21]. Ecosystem [CVH^{+}20, FRVB22, HGG^{+}22b, KS23, KN21, OI23, RRR^{+}23, WHPF20, CDD^{+}23, CMI^{+}22, CTH^{+}21, CWKE20, CSL21, CNG^{+}23, DKW20, De 22, DJC^{+}23, DA23, EAGV20, GPG^{+}21, GCC^{+}21a, HRR^{+}20, HF20, HPPH22, HPC^{+}23, LLT^{+}20, LW22, Liu23, LRC24, MUR^{+}21a, MKAF21, MFC^{+}24, MWWS21, MKB21, MZT^{+}23, OVSVPP^{+}22, POM24, PGA^{+}20, PS21, QK24, QRK^{+}23, RFG24, RR20, RSFvH^{+}22, RKS^{+}23, RGH^{+}22, SAO^{+}21, SWI^{+}22, SL20, THGR^{+}20, Tra23a, WD23, ZCG^{+}23, dSOW^{+}21, dO23, vTHS21, JL21d, OKS^{+}22, SLP^{+}22]. Ecosystem-based [HGG^{+}22b, KS23, WHPF20, CSL21, EAGV20, GPG^{+}21, LLT^{+}20, Liu23, MUR^{+}21a, OVSVPP^{+}22, PGA^{+}20, PS21, QRK^{+}23, RR20, RKS^{+}23]. ecosystems [BMR20, BMMF22, COM20, DMG^{+}23, GCG^{+}21, HGG^{+}22b, Lev20, Lot22, OKR^{+}20, QUK21, QLU^{+}23, QK24, SOK^{+}22, dO23]. ecotourism [CLW^{+}21]. Ecuador [COJE24, DRCdMCJ21, MTC^{+}23]. Ecuadorian [MTC^{+}23]. Ed [Rub24, Smi20]. edged [Gug24]. Edition [Tan23b]. Editorial [Als20, BG20b, AEJ24, ZT23b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano21p, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano22o, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano24b, Ano24c, Ano24d, Ano24e]. eDNA [CCS23]. Eds [Cha20b, De 20, Lv20, Sos21a, Tan23a, Tan23b]. educating [RB20]. education [FLB^{+}21, cL22, OBNK24, WSD21]. edule [BASM22]. Edward [Cha20b, Smi23b, Sos21a, Tan23b]. eel [KYMW21, SK24, SFN^{+}21]. EEZ [ABI24, GAD^{+}24, TYS22]. Effect [CQ20, CRCP^{+}23, DA23, GAMR20, HSHS24, PKR^{+}22, PAK22, QTPL21, RAB22, SKL^{+}20, XLLS21, XSCL21, YTÇ22, ZYTT21a]. Effective [YCL22b, ABM21, AGM^{+}22, BvdH23, CdLN^{+}21, Dau23, EAA^{+}22, HHH^{+}23, JAH^{+}24, KMS20, MBDG23, OCD^{+}23, PKMT23, TJ21a, VWFW22, WW21a, WMKG21, XXMZ23, ZD20b, ZBF^{+}22]. Effectiveness [OKNB22, SGLF^{+}23, AHAB^{+}22, ADR^{+}23, CDC21b, CES20, Day24, Fel22, VBd^{+}21, GKP^{+}22, KF23, RP23a, RP23b, SSDQ20, SFN^{+}21, SCY24, TAK^{+}22, XWL^{+}20, YU20, ZPX^{+}22]. Effects [CCX23, MJ20, OWS20, QS21, SHAC23, WM21, BST23, CDNS23, CDC21a, CDB20, hBDF22, DME23, FGPPGG23, Fla21, GGL22, GLT23, HRR^{+}20, JK22, KBD^{+}20, LKE^{+}22, LM24, MVSP^{+}21, NUD^{+}23, OCT^{+}22, PGK22, Sha20, SSB^{+}22, SO20, TDC21]. efficacy [ALML23, HSHS24, PK23]. efficiencies [Cro21]. Efficiency [COJS^{+}22, GABSD21, ÁCT20, BDD^{+}21, DG23, GAMR20, JL21a, LWG23, MWB^{+}21, QS21, Ren21b, SLY21, VPRR^{+}21, XHY22, ZW21, ZLH21, OAM20]. efficient [DG24a, DG24b]. effort [BCJ21, HBJ21, MAG^{+}22, MM20, Obu20]. efforts [ABDRV^{+}23, FMS^{+}22, LC23, WW22]. egg [HM22b]. EGT [Alf20]. Egypt [AHBAKH24, SK20b, TCC^{+}22]. Eilat [YNZ22]. eke [MAR^{+}20b]. elasmobranch [HHBC20, KGL22, BdLAdC^{+}20]. elasmobranchs [GAD^{+}20, WLCG23]. electric [Pen22, RPK23]. Electronic [BDC^{+}21, GDLC20, MFG^{+}23]. element [VYB20, BAP21]. elementary [cL22]. eleven [CQ20]. Elgar [Cha20b, Smi23b, Sos21a, Tan23b]. Eliciting [CLB^{+}23]. eliminating [FR23]. Elise [Niu22]. Elmina [GE23]. Elsevier [De 20]. embayment [MCW^{+}22b]. Embracing [BvdH23, NMCL24]. Emergence [Sah22, SVF20]. emergency [BMRM23, LC20b]. Emerging [CLRW22, MWM^{+}23a, RSS^{+}20, GPA^{+}22, HSGS^{+}23, LV23, TFMT^{+}23, YZH22, Tan23a]. Emirates [Gul20, Squ20]. Emission [WLXY22, DXS^{+}21, HRS^{+}24, JWX^{+}20, Okş23, Tes20, VGK23, WZZC21, WW22]. Emissions [LLW22, BAD21, CMC^{+}23, Che21b, CDHY22, GZW^{+}20, HKMS21, HSSM24, Kon20, MVSP^{+}21, NNR^{+}21, OOGK^{+}24, QLZ^{+}24, SSK23, WL22b, ZLPT23, Zhu22, Zhu23]. EMODnet [SSMG21]. emotion [BDLO23]. emotional [TNY23]. Emperor [BMR20]. Emphasis [Los23, AVPAFM^{+}22]. Empirical [CC21, KF23, AED20, CQ20, DMS^{+}24, Han20, cL22, LJRHH21, NCD23, PSMR20, PPK23, RZ21a, XSCL21, ZYTT21b, ZCSK23, ZCS^{+}23, Pra20, RZ21b]. employed [KPN20]. employment [LC20a, WFG22c]. Empowering [LLS^{+}20, SSB^{+}22, AHB^{+}23]. empowerment [FS22, dOJCSLdSB21]. Empty [HBL^{+}21]. enable [PGSFRMHH24]. enabled [MWWS21]. enabler [WOG20]. enablers [MKM23]. Enabling [Col20, LMMT23, ZBF^{+}22, BLY^{+}24, SGL^{+}21, CMOB^{+}20]. enact [Gal21]. Encounter [SHAC23]. encountered [JPM^{+}23]. Encouraging [HM22a]. end [BBBS22, WEF^{+}21]. end-of-life [BBBS22]. endangered [BUM^{+}24, CMI^{+}22, Dam24, DPG^{+}22, GPH^{+}21, OVSVPP^{+}22]. endemic [YSEK22]. ending [BBHN20, FTSG21]. endowment [RZ21b]. ends [MZ22]. endurance [KL21b]. energy [BST23, CA20, CKY22, DG23, DG24a, DG24b, GSR21, HHEJB20, KGP^{+}22, KMM21, MBORGN24, MVSP^{+}21, MCRR23, MCTN23, OC20, ONA21, OKH^{+}24, SZBA20, WNMS^{+}24, vRB24, COJS^{+}22]. energy-efficient [DG24b]. energy-intensive [KMM21]. Enforcement [KAN^{+}23, Sha22, AGK22, ASM^{+}21b, Cha21, DG24b, FSKJ23, Hsi20, Lin24, MWB^{+}21, Tse21, VSV^{+}23]. engage [GSA23, MM23]. engaged [WWD23]. Engagement [MBB^{+}22, ZK21, BOOO23, BDC23, DG24a, EMN^{+}24, Has20, HMG^{+}21b, LHTR23, RCM^{+}22, TMM^{+}23, WWD23]. Engaging [GBSC22, VPB^{+}24]. England [BD20, BHB^{+}24, CSGC22, DGS^{+}24, HHEJB20, SSDQ20, WDC^{+}20]. English [ASKI20, CWMM^{+}21, NJBM^{+}23]. ENGOs [WYH21]. enhance [DMG^{+}23, GSM21, MAA^{+}24, TSA^{+}22]. Enhanced [EAA^{+}21, DSWB20, PSTW21]. Enhancing [BMMF22, MMGAA^{+}21, PKH22, VSPMS^{+}22, BDR20, GGP^{+}22, MF22, SGH22]. enough [GB21, MH24, dlPdlLLC^{+}22]. Enrollment [JAK^{+}23]. ensemble [GCS21]. ensure [PK20a]. entangle [MTC^{+}23]. entanglements [AVF^{+}23a, MM20]. entangling [EMH^{+}23]. enterprise [LZZ22, Wil21b]. entrant [WFG22c]. entrepreneurialism [Zho23]. entropy [DG20]. entry [BDHI21]. Envelopment [PCD^{+}23]. environment [AS23b, CA21, CCK^{+}21, CSHO^{+}20, DBBL21, GPM20, Kuł22, KHLS23, LP21, MCBN22, MSSS22, MDG^{+}23, Tra23a, Dbo22]. Environmental [BAF^{+}23, CR23, CG22, CGHV21, CDC21b, DXS^{+}21, GKP^{+}22, GVR20, HM22a, JACD20, KMM21, MN21, PYVF24, RCM^{+}22, SGK21, Tan23b, AAKI21, AGM^{+}22, AR22, AKL^{+}22, AH21, BSY^{+}20, BDD^{+}21, BRM21b, CMS20, CS20, CQ20, DMd^{+}20, Dbo22, DUYEA21, Doe21, FY24, Fei21, VBd^{+}21, GLADGGVL23, GG21, GR20a, GCC^{+}21a, GCL21, Gul23, Guo20, HLW^{+}23, HSM^{+}23, HWY22, Jae21, JKV^{+}22, JCMD20, JL21b, Koj23, KDVR20, LSH21, LC22, LFVCF24, MLVT21, MPL^{+}23, MAK22, NNRS24, NWZ^{+}21, OKR^{+}20, OOGA22, PKH22, PC20, Puh23, RJ21, SC21, Squ20, SH22, SRLLG23, TLT^{+}22, TOHK20, TML^{+}21, VGS^{+}20, VCC^{+}20, WWCS21, WDG^{+}20, WCC^{+}20, War21, WFG^{+}22b, YS21, YZG23, ZW21, MFG^{+}22, PS21, dOSTB^{+}20, TML^{+}21]. environmentally [FBC^{+}23, LLJZ22]. environments [AFHHD22, SBJF^{+}21, SBT^{+}23, YMG23]. Envisioning [PCA^{+}23, VGC^{+}22]. epidemic [LLHG23, LC20b, XSCL21]. Epinephelus [HDAPG^{+}24]. Epistemic [BC20a]. equality [BGLTC21, OKNB22, OBNK24, PGSFRMHH24]. equatorial [NOL^{+}23]. equilibrium [WLQD21, WHPF20]. equipment [HPH22, Okş23]. Equitable [WLCC23, AVF^{+}23a, BTE21, Dom23, GABA^{+}23, Gug24, JDF^{+}23]. equity [FHBS23, How20, MM22, OGS24]. era [KAC22, LLHG23, SS22]. erosion [CLC21, FSF^{+}21, NNRS24]. error [CBR21]. ESC [FFYE21]. escalating [Kar22b]. escalation [KL21b]. escape [SKL^{+}20]. escaped [CAHR^{+}24]. Escapement [MKM^{+}20, ADR^{+}23, SKL^{+}20]. escapes [CAHR^{+}24, YHU22]. especially [ZJ22]. essential [LHV^{+}20]. established [QTPL21]. Establishing [AC23, Wat22, BN20, DTBP21, WWG^{+}23]. establishment [CLL22a, SAO^{+}21]. estimate [Kaw22, SRSGMGG21a]. estimates [RWVH21]. Estimating [BBC24, CDD^{+}23, SRSGMGG21b, TSA^{+}22, WCT^{+}20, XSCL21, ZMR20, JV20]. Estimation [SKE22]. Estonia [LKB^{+}23]. Estonian [SK20a]. estuarine [AS23b]. estuary [DPP^{+}20, SCBdA20, SCA^{+}22]. ethical [GCL21]. ethnobiological [NEH^{+}23]. ethnography [GMRSP23]. ethos [Dun24]. EU [CHEC21, CCGM21, DL21, DN20, DMG^{+}23, EB23, Faj22, GMPT^{+}20, GCC^{+}21b, IGO21, KL21a, KA21, LFVCF24, MPD^{+}23, MMGAA^{+}21, OSV21, PKB^{+}21, PMK^{+}21, Pen22, dlÁPAFGO20, Puh23, SOLN20, Sta23b, SFN^{+}21, THP^{+}22b]. EU-level [CHEC21]. EU-Marine [MPD^{+}23]. EU28 [CMR21]. Eubalaena [MM20]. Eulachon [BMM^{+}20]. EUR [Rub24]. Europe [Smi23a, ALF^{+}22, BG20b, BB21, BRP^{+}23, dlFGFFR^{+}20, GGL22, LJ20, LJ23, OMC^{+}22, PY20, RHL20, SVV24]. European [Ada22, Álv23, ACG21, Aqu23, ALF^{+}22, BDCJ22, BUM^{+}24, CVCV^{+}22, CAM^{+}23, CGGA20, DER^{+}21, FGPPGG23, FMGV20, FGC21, GBC^{+}20, GLAVL21, GSGOP^{+}21, GRA^{+}22, KvTNL20, KGP^{+}20, LFPBD^{+}20, MFRQ23, MGBS24, NVP^{+}23, PMK^{+}21, PL23, PV23, PŚ22, RPY^{+}24, SPFA^{+}20, SGR21, SSMG21, SOLN21, TDP20, VDP^{+}22, VSDR^{+}22, ZWJvL20]. evacuations [BVT^{+}21, HVN21]. evaluate [DQJ23, HLCM21, SA21, SNP21, SG20b]. Evaluating [DDAB^{+}23, CSL21, Dam24, GKP^{+}22, HBHJ20, KCCC^{+}21, MVR21a, MLM23, RA20, RP23a, RP23b, RHC20, SII22, SHP^{+}21, Squ20, BTV^{+}22, PKB^{+}21, SLH22, SRS23, HM21, SSDQ20]. Evaluation [AVF23b, COJE24, FWL^{+}22, IBC^{+}20, JLY23, LWX^{+}20, LXW^{+}21, MBORGN24, ZCCQ22b, ASGS23, DSL20, DG20, EHM^{+}20, EHM^{+}22, KMS20, KOBS21, LYL^{+}22, LFLF22, MZ22, MM23, Mor20, NTH^{+}21, OWS20, Ren21b, SGWL21, TM23a, Tan24, TSS^{+}24, ZPX^{+}22, ZLD^{+}21, vdBSvdV^{+}23, YU20]. evasion [WOG20]. eve [AR22]. events [SMKS22, SPA^{+}21]. Every [SCBB^{+}22]. everyday [dZ20]. Everything [FR23, SCF^{+}22]. Evidence [BK23, CVA22a, LTW23, NBB^{+}23, PGP22, SPG20, Tan22, TSS^{+}24, ZT23a, DPG^{+}22, DNR21, GAD^{+}23, NRYG20, QLU^{+}23, RJ21, TSH^{+}22a, ZZW^{+}21, ZYTT21b, AAW^{+}23, CACJR^{+}21, CDG^{+}21, dHJTGO21, JL21b, LCC20, NAN^{+}24, THT^{+}23, UOO20]. evidence-based [QLU^{+}23]. evidence-informed [GAD^{+}23]. evidencing [CMS^{+}21]. Evolution [DSJ^{+}23, LXH^{+}22, LHPJ23, QYD20, SS23, YYL20, AWX21, CVA22b, LXSX21, LJ22, PN20, QZZX20, RS24, Sah22, SCY24, YYD23, ZWG^{+}21, ZWC^{+}23, HLM^{+}20, YH20, ZSZ20]. Evolutionary [WXLD21, Alf20, Bir20, DLX21, JWX^{+}20, SVHV20, WLQD21, WL22b, XHZ^{+}21]. Evolutions [DPP^{+}20, FGC21]. evolving [KY20, NEF23, VYB20]. ex [AMIM21]. ex-vessel [AMIM21]. examination [TAK^{+}22, ZD20a]. examine [CLB^{+}20, Ren21a]. Examining [AL20, CCCC21, NHD23, SCSR^{+}23, SPJB22, SRH^{+}22, AELR^{+}20, SCSR^{+}22]. example [AB20, BGMCG24, COJE24, Gir22, KMS20, LC20b, MMNC22, OSV21, Sam22, Wan23a]. Examples [BI21, GA20]. Exclusive [SOLN20, CLP21, OKFN24]. exempted [BBJ^{+}22]. Exemptions [Dom23]. exhaustion [TNY23]. exist [Gar20]. existence [WJB^{+}20]. Existing [COJS^{+}22, CSHO^{+}20, FGK^{+}20, FMS^{+}22, GRPRVG23, RB20]. expanding [BvBN^{+}20, CVA22b]. expansion [AHAB^{+}22, CFM^{+}21, Cha20a, Dun21a, LHTR23, POM24, Tra23b]. expectation [RHC20]. expedited [FXC^{+}22]. experience [ACG21, HWY22, KKO^{+}22, LCMBGP22, MKS21, SBT21, ZLZ21, DFTK20, OGGC20, Tse21, YYL20]. experienced [AO20]. Experiences [YS23, INW20, LXH^{+}22, OBNK24, PGSFRMHH24, RB20, ZTW23, LEVET21]. experiment [CDD^{+}23, CLL20, CKY20, GdACT20, KNY22, MZT^{+}23, RXX24, SNP21, SL20, TOAPJ21, WLX21, WYD^{+}22]. Experimental [NRYG20]. Expert [CPV^{+}21, KHLS23, BVT^{+}21, OC20]. expertise [ZD20b]. Explaining [IAB^{+}20, VM21]. explains [SOS20]. explanatory [BSM23]. Exploitation [ML22, HSM^{+}23, LWX^{+}20, NOL^{+}23, PGSS22, War21, YS21]. exploited [MN23]. exploiting [BDDE20]. Exploration [BDC23, LNRM23, BL22, NEF23, VJL^{+}21]. Exploratory [SPT^{+}20, BvdBvdFK20, FS21, NGP^{+}21, TMT20]. Exploring [CGP23, DGMF^{+}20, GFD21, HWT23, HMG^{+}21b, Kas21, LTSL22, LJ23, ML21, PSD23, PL24b, Pra24, QRK^{+}23, RRR^{+}23, SBP24, VWJMAJ20, WSD21, XZ23, YW20, dSdLSGB22, QCM^{+}20, GSA23, KK22, KL21b]. export [FZYF23, VKN20]. expose [SFN^{+}21]. exposure [CBR21, KDH^{+}21]. extended [DJC^{+}23, DLYM22, HLCM21, MLM23]. extension [MM23, ZJ22]. extent [CMOJ^{+}24, LW22]. external [KA21, PGK22, SCLM^{+}21]. extinct [JG20, VCC^{+}20]. extractive [VVR^{+}22]. extractivism [Gal21]. extreme [SPA^{+}21]. eye [BHEE22, NBB^{+}23].

Fabinyi [AFHM23]. face [AH23, LL24]. faced [DSL^{+}21, DSC^{+}23]. facilitate [AANM^{+}22, Han21, STIS20, SBJF^{+}21]. facilitated [LC22]. facilitating [KMT^{+}21]. facilities [Osm23, PB20]. Facing [AHB^{+}23, DC20, DP23, MKM^{+}20]. Fact [HGSD23]. factor [DL22, RZ21b]. Factors [HMFH20, KWMM20b, LSY20, MLMD21, PPK23, AAKI21, BC20b, Dam24, DLX21, GLADGGVL23, HP23, IGO21, JZW21, KKDK20, NEF23, PGSFRMHH24, Pra24, RE21, SGL^{+}21, TMS^{+}22, WFG22a, XHY22, XXLW24, ZWXL21, ZSLY20]. Factory [RBNE20]. Factory-made [RBNE20]. FAD [MSZ^{+}23, MZT^{+}23]. FADs [MZT^{+}23]. fail [CBR21, HM22c, MS21]. failed [Åm21, OSV21]. failure [GVL22, MF23, Pol23, SOS20]. failures [SSP^{+}21a]. fair [Dau23]. faire [KK22, YL21]. fairways [PYVF24]. Falklands [AM22]. Falklands/Malvinas [AM22]. fall [GABSD21, SCF^{+}22]. fallout [HS22]. Family [MLVT21, PNS^{+}23, TNY23]. Family-run [MLVT21]. fan [SPCCvT21]. FAO [SFGEPC22]. farm [BST23, BBO21, CSLGFV20, FAC^{+}24, HRR^{+}20, RBNE20, SBT21, TJZ^{+}21]. farm-induced [FAC^{+}24]. farmed [CAHR^{+}24, EHLT21, SLMG^{+}21, STG^{+}20]. farmer [AAW^{+}23]. farmers [RKR20]. farming [BE20a, FGPPPV21, Gir22, GVR20, HSN^{+}21, KHdSM^{+}22, KWY^{+}22, LJ22, Mrd23, SCLM^{+}21, SLS^{+}20, YK20]. farms [BvdBvdFK20, BLY^{+}24, CCW21, DWC^{+}20, PKMT23, SVKP21, TYS24]. Faroe [Chr20, DAA21, EB24, Mrd23]. fatalities [NSPRSS20]. Fatigue [RE23, ZTW23, ZTM^{+}23, PGK22, ZWJvL20]. fauna [SS23]. favourable [Che21b]. fear [Pol23]. fears [WM21]. Feasibility [WDG^{+}20, ZFGL20]. feature [AH23]. features [PSTW21, YWY20]. February [Ano20n, Ano21q, Ano22p, Ano23p, Ano24f]. federal [ALML23, NH20, WEF^{+}21]. feedback [FTW^{+}23]. Feeding [AWX21]. ferruginea [OVSVPP^{+}22]. fiction [HGSD23]. field [Wil22]. fifth [TMDN23]. fifty [Smi23a]. fight [Gar22, GY23, Oan22]. Fiji [RGK^{+}20, TPG^{+}23]. Fijian [YS23]. fill [SSOP^{+}21]. Filtering [Wil22]. fin [TOAPJ21]. final [HHD20, MTN23]. finance [SLY21, WSD21]. finances [KJS^{+}22]. Financial [Ero22, PALB20, WLY21, AAW^{+}23, BDHI21, EP20, KB20, KWMM20b, KTI^{+}24, PPK23, Tan22, ZZW^{+}21, ZLZ21]. financialization [SS22]. financing [SWHT22, TT22]. Finding [Gug24, hC24]. Findings [KNY22, PBD^{+}23, TMT20]. finfish [CGGA20, HHB^{+}21, KHdSM^{+}22, LHTR23]. Finland [PKH22]. Finnish [HvT22]. Fins [BMF20]. Firm [SSK23, BXV20, ZYTT21a]. firms [KLSG21, QZZ20]. First [DPG^{+}22, APF22, BBK21a, BGK^{+}22, BMR23, GPH^{+}21, GBC21a, LKE^{+}22, MZT^{+}23, PBD^{+}23, SBT21, ZLD^{+}21]. fiscal [Chr20, JL21b]. Fish [MAK22, MMW^{+}23, SBD^{+}23, SGZH^{+}21, SOS^{+}21b, AED20, ABRA20, AP23, AFMdSdTdCC22, AAW^{+}23, BEM^{+}20, Bea22, BHdMCM^{+}21, BLM24, BUM^{+}24, BTE21, COJE24, CMNH21, CDG^{+}21, DPB^{+}23, hBDDF22, EMH^{+}23, Fox22, GSG^{+}24, GCV^{+}21, GVR20, GAA^{+}20, Han20, HMG^{+}21a, HdM23, HGH^{+}20, IO23, JY20, cL22, MFG^{+}23, MSO23, Mar21b, MN23, MDG^{+}23, MMK^{+}22, Nal24, NTH^{+}21, NAS^{+}22, OOSG21, OVA22, PT23, PDJ22, RAFLRP21, SCBB^{+}22, SKL^{+}20, SOS20, SHFP21, TCC^{+}22, TJZ^{+}21, UISM^{+}23, VM21, WLRH^{+}20, YFD^{+}21, YHU22, ZTZ^{+}21, ZMM^{+}23, CMT^{+}23, MSZ^{+}23, PMT^{+}20]. Fisher [AGK22, AFMdSdTdCC22, BUM^{+}24, HRMPCRRA22, MNGM24, PBS^{+}23, SMLRB^{+}21, AACdlT23, AS23a, BEM^{+}20, VBd^{+}21, GSAF21, GRPRVG23, GMRSP23, KTI^{+}24, Li23b, OCZ^{+}22, PGSS22, RPK23, RCM^{+}22, SPG20, SBFHD^{+}23, Sol23b, SCBdA20, TSH22b, VDP^{+}22, YCL22a]. fisherfolk [RBA23]. Fisheries [ASGS23, ALLS23, AGA^{+}21, Bra22, CTH^{+}23, DL21, DMG^{+}23, EB24, EHM^{+}20, EHM^{+}22, FSKJ23, FLB^{+}21, FMCT^{+}21, Gal21, Gar24, GML^{+}24, INW20, JC23, Kim21, KMJ24, Li23b, NSZ^{+}21, NCE21, PV23, Run24, SOT^{+}22, SKW^{+}22, VDP^{+}22, VCC^{+}21, VYB20, WKH22, ZL20, APV^{+}20, AMIM21, Abo23, AANM^{+}22, ASV^{+}21, AOO^{+}23b, Alm21a, Álv23, ACV^{+}21, AEL^{+}23, ABRA20, AFMdSdTdCC22, AVPAFM^{+}22, Aqu23, AMR21b, ALF^{+}22, AR23a, AR23b, AEJ24, Art20, ASGZ22, AH21, AL20, BSBD21, BPA^{+}22, BKW^{+}24, Bal22, BMB^{+}24, BDM20, BBC24, BEM^{+}20, BB21, BJGG^{+}24, Bea22, BSSL^{+}20, BHdMCM^{+}21, BHKK23, BBS^{+}21, BDD^{+}21, BBJ^{+}22, BLY^{+}24, BASM22, BGMCG24, BK23, BSK^{+}20, BDC^{+}21, BAD21, CPR21, CdLN^{+}21, CPBO20, CCGM21, CTH^{+}21, CAVJ^{+}20, CAM^{+}23, CHB^{+}20, CDG^{+}21, CDHY22, CMOB^{+}20, CMSN^{+}22, CES20, CHC^{+}21, CRCP^{+}23, CDB20]. fisheries [CSL21, DLJ^{+}20, DKKS21, DQJ23, DH22, DDF^{+}20, DAN^{+}23, DSL20, DGS^{+}24, DCC^{+}20, DDGdS^{+}21, DFF23, DHR23, EAAG21, EAGV20, EB23, EMS^{+}23, FARBV^{+}24, FS21, For22, FS22, FWL^{+}22, Fuj21, FHBS23, GSA23, GLAVL21, GLADGGVL23, Gar22, GSAF21, GLS^{+}21, GRA^{+}22, GDLC20, GCB^{+}23, GM21, GCL21, GBSC22, GMRSP23, GZW^{+}20, GMCMJA^{+}21, Gug22, GJS22, GRPE^{+}21, GVH^{+}23, GV22, Guo20, GFL^{+}21, HMFH20, HDH21, HDCH21, HGSD23, HWU^{+}23, HLW^{+}20, HYN^{+}23, Han21, HWT23, Har22, HS22, HvdBH23, HF20, HW22, HZ23, HBR21, HBR^{+}23, HGLR22, HCSBM^{+}23, HTP^{+}23, HSdT21, HHH^{+}23, HT23, HvPPV20b, How20, Hsi20, HCS21, HBHJ20, HBL^{+}21, IKB21, IRK^{+}21, IAB^{+}20, JY20, JTR^{+}20, JSW^{+}22a, Jen23, JAdS^{+}20, JKKT20, KWS20, KWF^{+}21, KA21, KYMW21, KMS20, Kao23, KFB^{+}20, KB20, KFP20, KSB^{+}23a, KSB^{+}23b, Ken20, KPN20]. fisheries [Kim23, KHH^{+}23, KWMM20a, KWMM20b, KLSG21, Kob22, Koe22, KAN^{+}23, KSV^{+}23, KJS^{+}22, LBL21, LRN^{+}22, LBWS23, LLLF22, LCAX21, Liu23, LA23, LSGG^{+}20, LGW^{+}21, Lov23, LLS^{+}20, LSD^{+}22, MMM^{+}22, MMK^{+}23, MKM23, ML21, MLVT21, MF22, MMB^{+}23, MVR21a, MZ22, MCPS20, MWM^{+}23a, McC20, McC21, MAK22, MM22, MH20a, MKS21, MHN^{+}20, MMSB^{+}20, MAG^{+}22, MS21, MPSF20, Mrd23, MIH^{+}24, MSGAR20, NH20, NNM^{+}21, NHA23, NSK20, NAS^{+}22, OCZ^{+}22, OR21, OYKBA22, OOGS20a, OOGS20b, OOGK^{+}24, OI23, OCD^{+}23, OCCSC23, OAM21, OMC^{+}22, OA20, PCA21, PRM^{+}20, PFG^{+}21, PSPR20, PL20a, PVH^{+}21, PGSFRMHH24, Pen22, PWL^{+}23a, Pet22, PCK22, PJN24, PGGA22, PALB20, PCCOS24, PLOG24, PKS20, PFWM23, PKMT23, QTPL21, RLS20, RYS^{+}21, RMU^{+}22, RROP23, RTMCVMDE22, RKS^{+}24, RHO24, RDLA^{+}20, SAO^{+}21]. fisheries [SPFA^{+}20, SOO20, SK20a, SK20b, SIW^{+}21, SWI^{+}22, SLB21, SCJ22, SGR21, SBD^{+}23, SVV24, SSZ^{+}24, SMYH22, Seu22, SWF^{+}24, SSH^{+}23, Shi23b, SKE22, SS22, SS21a, SSV^{+}23, SHP^{+}21, Smi23a, SS21b, SPMN23, SGL^{+}21, SG21, SH22, SH23, SRH23, STKC22, SRH^{+}22, SHFP21, SSP^{+}21b, SG20c, SKL23, STF21, Szy20, TTA^{+}20, Tel23b, THP22a, THGR^{+}20, TSH22b, TLT^{+}22, TOL^{+}22, TKP23, TEA^{+}22, Tra23a, TMTC20, TK22, TDP20, UOO20, Urb22, VMD^{+}24, VSPMS^{+}22, VGC^{+}22, VSDR^{+}22, VGAH^{+}21, VLS^{+}23, WS21, WHPF20, WLRH^{+}20, WH20, WWC23, WM22b, ZV20, Zha21, ZJ20, ZCCQ22b, vHSSK20, vdGD21, AMR21a, DDGdS^{+}21, DFTK20, EAAG21, FS22, GWAPT21, LJL22, MPSF20, RM21b, WBPG20, Zho23, Fla21]. fisheries-management [BEM^{+}20]. fisherman [LSY20]. Fishermen [PLOG24, BC20b, BMK^{+}22, GWCD22, dSJNS20, LSY20, Mat20, SKW^{+}22, TYS24, ZWXL21, dOPM^{+}20]. Fishers [CAVJ^{+}20, CWMM^{+}21, Gue23, ONN^{+}20, BSI^{+}21, CPR21, CJ23, CRJ22, DMd^{+}20, DSB21, DME23, Har21, IOHC22, JS23, KPN20, LBWS23, LEERRM^{+}21, MRKM23, Mar21b, MM23, MS21, Muk21, Nal24, OCD^{+}23, PRM^{+}20, Pra20, PDJ22, RTMCVMDE22, SSAP24, VPB^{+}24, YS23]. fisherwomen [SSB^{+}22]. Fishery [CMC^{+}23, JAH^{+}24, JK22, Mot20, WS23, AWX21, ABDRV^{+}23, AE21, ASM21a, AP23, AMR21a, AONF21, BCBS22, BLH^{+}20, BDHI21, Ber22, Ber23, BCJ21, BFMS21, BM21, Bjø23, Bot20, BRR^{+}22, CC21, COJE24, CMR21, CMC22, CdLN^{+}21, CCB^{+}21, Cha20a, CCKK22, CSHO^{+}20, CNB^{+}23, CHC^{+}21, CSCGC20, DGG22, DSL^{+}21, DRCdMCJ21, Doe21, DFF23, DAAP21, DD22, EVPMKP22, EMH^{+}23, FGPPGG22, FGPPGG23, FLB^{+}21, FOH^{+}20, GMPT^{+}20, GCC^{+}21a, GCS21, GHW^{+}22, GSG20, GABSD21, GWAPRPRU22, GCM^{+}23, GNMGWM21, HGG^{+}22a, HYN^{+}23, HZ23, HvPPV20a, HC21, JGWM^{+}23, JL21a, JF20, KMJ24, KN21, LCC20, LLM^{+}23, LMC^{+}23a, LBWS23, LLL^{+}24, LGW^{+}21, LJRHH21, LJ21, LFLF22, LCMBGP22, MRKM23, MVR21a, MVR21b, MJ20, MAG^{+}22, MH20b, MEG^{+}21, MM20, NAN^{+}24, OKFN24, OAM20, PL24a, PHF20, PL20b, PZZ^{+}21, Pra20, PWL^{+}23b]. fishery [QWBY21, RAB22, RQB^{+}20, RMBR22, Ros22, RHO24, RHST21, SII22, Sak21, SFGEPC22, SC21, SCBB^{+}22, SDL^{+}22, SC22, SCGC20, SMKS22, SCCW21, SCF^{+}22, SdCMN^{+}22, SGZH^{+}21, SGH22, SBW^{+}21, SG20a, SWXM21, SLK^{+}23, TM23a, TSS^{+}24, TMT20, Tra23b, TYS22, TPG^{+}23, UISM^{+}23, VJO23, WJB^{+}20, WW21b, WW22, WWG^{+}23, Wen20, WNMS^{+}24, WNBRSG22, WLW^{+}22, YTBM^{+}22, YML^{+}21, YYD23, ZXGL21, ZCCQ22a, ZSZ20, ZZW^{+}21, ZLZ21, ZLD^{+}21, Bra22, CDG^{+}21, DAA21, Mot22]. fishery-based [HC21]. fishery-independent [CdLN^{+}21]. fishery-specific [GHW^{+}22]. fishes [AGN^{+}20, LLLF22, SO20]. Fishing [AM22, AFHM23, BST23, Lin23, MB22, NGRM^{+}22, OAAMS24, SFGEPC22, AGN^{+}20, AGS20, ÁCT20, BV21, BJH21, BSI^{+}21, BCJ21, BBJ^{+}22, CRP^{+}20, CBR21, CDNS23, CLS^{+}23, CSGC22, DGG22, DT23, hBDF22, DXS^{+}21, DGIS^{+}22, DLX21, DA23, DP23, DAHS^{+}23, DS22b, Faj22, Fen23, FARBV^{+}24, FRM22, Fla21, Fro21, GRS^{+}23, GCF^{+}23, GÁPSAP^{+}23, dlFGFFR^{+}20, Gar24, GSG^{+}24, GHW^{+}22, GAD^{+}23, GAD^{+}20, GE23, GASLCP20, GPTASM23, GY23, GR20b, GABA^{+}22, GSB20, GAA^{+}20, HCWC21, HDAPG^{+}24, HBP^{+}21, HM22c, JS23, KL21a, Kar22c, Kas21, KgJqY^{+}24, Kim21, KLSG21, KDVR20, KAN^{+}23, LMC^{+}23a, Lov23, Ma20, MLMD21, MMM^{+}22, MMC^{+}22, MCBN22, MKM23, MJMAA21, MM23, MGF20, Mil22, MLAS23, MdCLL^{+}23, MMGAA^{+}21, MHEW20, NBB^{+}23, NSPRSS20, Oan22, OGS24, OVA22, OA20, Pat22, Pen22, PMC24, PPK23, PMT^{+}20, QCM^{+}20]. fishing [Ree20, RWVH21, RC22, RM21b, Ros22, RDD^{+}21, SPFA^{+}20, SL23, SIW^{+}21, SDI^{+}20, SCD21, SLL^{+}24, SCSR^{+}22, SCSR^{+}23, SSZ^{+}24, Sha22, SNS23, SCGC20, SSB^{+}22, SKL^{+}20, SKS21, SCH^{+}24, SBT21, SGH22, SSP^{+}21a, SH23, Sta23b, SRY24, SCBdA20, SSF23, TOAPJ21, Tei23, TSH^{+}22a, TSLK23, Tse21, TK22, TGM22, VMM21, VSD^{+}23, VAMR^{+}20, VPRR^{+}21, VVR^{+}22, WKO^{+}22, WNMS^{+}24, WBH^{+}23, WGP^{+}21, WLW^{+}22, YS23, YL21, dMLB^{+}20, dlPdlLLC^{+}22, Smi23a]. Fishpath [CDB20]. fit [GNJR24, MBL24, WS23]. fit-for-purpose [GNJR24]. Fits [Mag23]. Fittest [OYKBA22]. five [CMGF20, cL22, CMGF20]. fix [Abo23, FDF^{+}21]. Fixed [FDF^{+}21, OAM21]. flag [BISSBdlMM22, DBZ24, Fel22, Kar22a, PSMR20, dlÁPAFGO20]. flagged [BLH^{+}20, HVN21]. flagging [ZD20a]. flagging-out [ZD20a]. Flags [PSMR20, Fel22]. flapper [GPH^{+}21]. flavescens [PBH^{+}23]. fleet [ATF^{+}21, ASGZ22, BDCJ22, CWMM^{+}21, Fel22, dHJTGO21, HCWC21, SPG20, SLL^{+}24, VSD^{+}23, ZD20a]. fleets [HBJ21, MPL^{+}23, MLAS23, NSPRSS20]. flexibilities [May22]. flexible [DBS22a, PVH^{+}21]. floating [ADR^{+}21, CSLGFV20, OR21]. flood [CBAPM20]. Florianópolis [FSF^{+}21]. Florida [AP23, BDMW22, FMWW21]. Flows [GMG^{+}23, JKKT20]. flying [GWAPRPRU22]. FMEA [CAC23]. focus [BSSL^{+}20, LBWS23, BC20b]. Focusing [YK20]. Follow [Hor23, BGMCG24]. following [FAGS21, LCL20]. follows [Tan24]. FONDEPES [AACdlT23]. FONOP [Tod22a]. Food [BvdBvdFK20, DMG^{+}23, YH20, BMF20, BRM^{+}21a, CTIH21, CGHV21, FTM^{+}22, JG23, LLS^{+}20, MF22, NRYG20, OOGK^{+}24, PBT^{+}20, RBNE20, SNK^{+}20, SN23, SSP^{+}21b, TFH^{+}23, TGHF^{+}22, TCC^{+}22, TMTC20, WMS21, ZWC^{+}23, Gug22]. footprint [DGG22, SLL^{+}24]. footprints [Kas21]. forage [THGR^{+}20]. force [Cha21, Ngu23]. forced [SRY24, TOL^{+}22, YS21]. forces [TJ21a]. foreign [ABI24, AED20, Kao23, Mar20a, OOGS20a, PSMR20]. foresight [PMT^{+}20]. Foresighting [KFS^{+}22]. forest [HGG^{+}22b, QTBL23]. forests [CKY20]. Form [Tan24]. formal [GBSC22, PL20b, SPHO23]. forming [Zou23]. Forsskål [SGPF21]. fortunes [BMF20]. forum [NT20]. forward [CHB^{+}20, Chr20, DSJ^{+}23, JPM^{+}23, VBH^{+}23, VGC^{+}22, Wil21b]. fossil [HSM^{+}23, vLM22]. foster [GM21, GMG^{+}23]. fostering [GABA^{+}23]. fosters [HSM^{+}23]. found [Lev20]. Foundation [ZD20b]. foundations [GEB^{+}21, Sta23a]. four [Ken20, SDL^{+}22]. fourth [MDJ^{+}22, War23]. FPIs [AE21]. FQA [FDF^{+}21]. Fragmented [Jen21]. frames [TDC21]. Framework [ANS^{+}20, CJL^{+}23, GKP^{+}22, GABA^{+}23, MPD^{+}23, MFRQ23, PKB^{+}21, PDB^{+}23, SLP^{+}22, VRBC^{+}21, VBH^{+}23, ASV^{+}21, BN20, BGLTC21, CMT^{+}23, CDRS22, CSVC23, COJZ^{+}21, DQJ23, DBS22a, Dbo22, DJC^{+}23, FFGJ22, GE23, HJ21, HGL^{+}20, KF23, LZL23, LSGG^{+}20, LJRHH21, LLS^{+}20, LLL23, MAR^{+}20b, Mor20, O’D21, OGGC20, PL23, Puh23, Qi22, QUK21, QLU^{+}23, SWHT22, SS22, SG20b, SLH22, SRSGMGG21a, STF21, TJ21b, VÖB20a, VÖB20b, WPWM^{+}22, ZCCQ22b, ZLY23, ZCA^{+}23, vTHS21]. frameworks [BUR^{+}22, GLS^{+}21, NCE21, OC20, OT23, PY20, SRvNN21]. Framing [TDC21, BVvKL21]. France [CBAPM20, DPP^{+}20, LCL20, TLYC20]. Francis [Ngu23]. Francisco [THGR^{+}20]. fraud [JG23, LFVCF24]. fraudulent [SFN^{+}21]. Free [FXC^{+}22]. Freedom [Ngu23]. French [SO20, BG20a, DMG^{+}23, FLB^{+}21, KLSG21, TMT20]. frequently [CMI^{+}22]. fresh [RBNE20]. Fried [Lv20]. frigates [Dun24]. frontier [KGW20]. frustration [Pol23]. FTCs [VSD^{+}23]. Fuel [BAD21, DGG22, GZW^{+}20, HKMS21, HSM^{+}23, MF23, RAB22, SC22, VSD^{+}23, ZY21, ZGY21]. fuels [HRS^{+}24, vLM22]. Fukushima [LWZ23, MXZ23, PL24a, WL22a, XTY22]. Fulfiling [LWZ23]. fulfilment [KL22]. Full [PFG^{+}21]. function [OAM20, Tan24]. functional [Lam21, TZZ^{+}21]. functioning [Kiz24]. Fund [VDP^{+}22]. funded [MPSF20]. funding [CMC22, JV20]. further [KOBS21]. Future [KSB^{+}23a, MZ24, AKFY22, AS23b, CDC21b, DC20, FGL^{+}21, Gun21, HRR^{+}20, JTR^{+}20, Kas21, Kaw22, KFS^{+}22, KSB^{+}23b, KgJqY^{+}24, Liu21, LSD^{+}22, MMB^{+}23, MCRR23, MG22, Nal24, NJBM^{+}23, NT20, PSD23, RB20, SC21, SGL^{+}21, TSH^{+}22a, Ton22a, Tra23a, WNDZ23, Wan23a, WM22b, XT23, YH20, ZCZ20, ZLY23, Zhu22, ZMM^{+}23, dKMK21, vdBSvdV^{+}23, Fuj21]. futures [PCA^{+}23, RHST21, SDL^{+}22]. fuzzy [CMR21].

G [Smi20]. Gakkel [KY20]. Galapagos [CD23, TOAPJ21, ZMR20, CAVJ^{+}20]. Galicia [HPC^{+}23, dlÁPAFGO20]. gallina [BSBD21]. Gambia [NRF21, SN23, AANM^{+}22]. game [BM21, DLX21, GSB20, HZ23, JWX^{+}20, JCMD20, XHZ^{+}21, ZLGX20, ZHL^{+}23, ZY21, ZGY21]. gamma [Kaw22]. gap [AMR21a, HSSM24, JV20, LTSL22, Li23b]. gaps [AGM^{+}22, CLP21, DA23, FAC^{+}24, FBC^{+}23, JAH^{+}24, Lov23, MAG^{+}22, NAP^{+}21, PMM^{+}21, QK24, SSOP^{+}21, SRH23, XZ23, YU20, ZSWJ23]. Gargoor [HGG^{+}22a]. Gas [Cha20b, Che21b, HKMS21, HSSM24, Zhu22, Zhu23]. Gastropoda [OVSVPP^{+}22]. Gata [HSGNMGC21]. Gathering [SCCW21]. GATT [WL20]. gayi [RQB^{+}20]. Gaza [Qaf24]. GCC [Als20, AED20]. Gdańsk [COJZ^{+}21]. gear [ABDRV^{+}23, DT23, DSB21, DA23, DFF23, Fen23, GRS^{+}23, GHW^{+}22, GAD^{+}23, HGG^{+}22a, LTJ^{+}23, Lov23, McC21, MPH^{+}23, RWVH21, SCSR^{+}22, SCSR^{+}23, TSLK23]. gears [SDG^{+}23, SKL^{+}20]. Gender [AOO^{+}23b, BGLTC21, GFL^{+}21, LRN^{+}22, CN23, LTSL22, MLE^{+}24, ML21, MM22, MMK^{+}22, OKNB22, OBNK24, PGSFRMHH24, PCSS^{+}21]. gender-blind [CN23]. Genderizing [Szy20]. general [DC23, Koj23, Shi23a, WHPF20, WWD23, PCSS^{+}21]. generalised [vTHS21]. genetic [HGL^{+}20, LR20]. genus [SFGEPC22]. geo [PMT^{+}20]. geo-foresight [PMT^{+}20]. geoduck [LCMBGP22]. Geographic [dSOW^{+}21, FLC^{+}23, FAC^{+}23, JMB^{+}20, KBD^{+}20, VKS23]. Geographical [HKKY22, Gir22]. geometric [CMR21]. European [LBWS23]. geopolitical [BSY^{+}20]. Geopolitics [Smi23b, Mar20a, MLAS23, WRL23]. Georges [WCF^{+}21]. Georgia [AM22, CHC^{+}21, Tra23a]. Geospatial [FDGM24, MDLA21, SBP24]. German [GAD^{+}24, GABSD21, ZD20a, ZD20b]. Germany [ZD20b]. GES [HvT22]. Getting [Obu20, SSDQ20]. Ghana [AE21, AVF^{+}23a, ABRA20, AAW^{+}23, AONF21, GE23, GAA^{+}20, KDVR20, NSZ^{+}21, OR21, OCD^{+}23, OA20, SAO^{+}21, SCBB^{+}22]. Ghanaian [TJZ^{+}21]. GHG [CMC^{+}23, HRS^{+}24]. Ghost [DA23, LTJ^{+}23, MPH^{+}23, TSLK23, YSEK22]. Giang [Han21]. giant [CJCS^{+}22, LZZ22]. Gibraltar [STS^{+}23]. Gili [PN20]. gillnet [Ber23, TK22]. gillnets [SGL20]. Gilt [LHW21]. GIS [CA21, CSLGFV20, JKV^{+}22]. Giving [Har21]. glacialis [MM20]. glass [KYMW21]. glauca [DSL20, SSA^{+}23]. gleaning [PJN24]. Global [FWGM24, HBR^{+}23, HGLR22, KKA^{+}21, Pyć23, SKDK23, Ada22, BGK^{+}22, BABC20, BvBN^{+}20, Cla20, CNG^{+}23, Dun21b, EAAG21, FLC^{+}23, FFGJ22, FGL^{+}21, GPA^{+}22, Gar24, HCW^{+}23, Han21, Has20, HHBC20, HBR21, HMY^{+}23, KKDK20, Kaw22, LDAC20, LMC^{+}23b, MSP23, PBS^{+}23, Pet22, PCK22, QF20, QUK21, RU20, RP23a, RP23b, RWVH21, Seu22, SGL^{+}23, SSH^{+}23, SS21a, SSV^{+}23, VGS^{+}20, VHS^{+}20, WvdMG^{+}21, WW21a, WBH^{+}23, CLZ23, HDHM^{+}20, SBDC24, TCZ21, GZW^{+}20]. globally [Gal22, QLU^{+}23]. glocal [SVF20]. goal [JV20, PFWM23, STS^{+}23]. Goals [BCN23, HSN^{+}21, NSN22, WCM^{+}24, YW20, TML^{+}21, vSHHT20]. goers [Lev20]. going [SCF^{+}22]. gold [FTG^{+}23]. golden [Dbo22]. Goliath [CNB^{+}23]. Gomez [CMWN^{+}23, DBH^{+}22, WvdMG^{+}21]. Gone [FAC^{+}24]. Good [PK20b, CKY20, DPB^{+}23, DUYEA21, GG21, KMT^{+}21, MH24, NNRS24, Puh23, GKP^{+}22, MMGAA^{+}21]. goodbye [BBBS22]. Gorringe [CLS^{+}23]. got [ALML23]. govern [Smi22]. Governance [CJ21, GCC^{+}21b, HWU^{+}23, KJE21b, LJRHH21, Rus20, Smi20, SWGZ22, Sta23a, SGL20, SG20b, VHSV20, YK21, YC21b, dPJ21, ASV^{+}21, AVF^{+}23a, Alf20, ANR^{+}22, BL23a, BPA^{+}22, BJGG^{+}24, BCB20, Bir20, BOW^{+}22, BHKJT23, Boy21, CC23, CVH^{+}20, CGHV21, CJ22, COSS21, CGP23, CMOJ^{+}24, DPP^{+}20, De 22, DGS^{+}24, DB20, DCC^{+}20, DHR23, EdLP^{+}21, EFFP23, FRM22, GGP^{+}22, Gal21, GCL21, GBC21a, GBC21b, Gro20, GJS22, GFL^{+}21, HIP^{+}21, HJ21, HM22a, HvdBH23, HSdT21, HSGNMGC21, HNMSG21, HGL^{+}20, JCMD20, JL21c, Jon21, dOJCSLdSB21, KA21, Kat20, KY20, LLT^{+}20, LJ23, LSH21, LTJ^{+}21, LJ21, LSY20, MM21a, MAR^{+}20b, Mba22, MMSB^{+}20, Mur21b, OPL20, OYKBA22, OI23, OCD^{+}23, OMC^{+}22, PFG^{+}21, PN20, PKH22, PMN^{+}22, Pet22, dlÁPAFGO20, Pri22, PPK23, PMT^{+}20, Pyć23, QLZ^{+}24, RU20, RKS^{+}24, SMDM20]. governance [SA21, SVHV20, SLB21, SPHO23, SMYH22, SNS23, SJ21, SKW^{+}22, SOLN20, SSP^{+}21a, SMCW22, SCA^{+}22, SHZC21, TUBA22, Tea20, TSA^{+}22, TFH^{+}23, TJ21a, TMT20, Tod22b, VTRS20, VHD20, VHS^{+}20, VYB20, Wan23c, WA24, WKBS20, Wu22, XLLS21, XXMZ23, YCL22b, YC21a, ZL20, ZLPT23, dZ20, Dun21a, KJE21a]. Governing [STG^{+}20, SRvNN21, STF21, Sos21a]. Government [YZG23, CdLN^{+}21, CS21, CVA22a, Fla21, HZ23, HWY22, KWMM20a, LLCH21, THP22a, WYH21, WXLD21, WLQD21]. governmental [CMS20, HWY22, KL22, Pet22, dOSTB^{+}20, WMRB22]. governments [JL21b]. grabbing [Mal21]. Grand [Ken20, LA23]. Grande [Ken20, dS21]. Granger [dHJTGO21]. grata [RROP23]. graying [HCWC21]. great [MCL22, DR22, Wat22]. Greater [WKH22, SAO^{+}21]. Greece [GPM20, Mar21a, Sak21]. Greek [GAD^{+}20, TDP20]. green [LWG23, LC22, RZ21b, Ren21b, RXX24, RKS^{+}23, Tho22, TC23]. greenhouse [Che21b, HKMS21, HSSM24, Zhu22, Zhu23]. Greenland [LJ21, SOS20, SSAP24]. grey [SLK^{+}23]. grid [RQB^{+}20]. grief [Pol23]. Grieg [MF20]. Griffiths [HH21]. griseus [NGT^{+}23]. groundfish [ASGZ22, LFLF22, MSGAR20, SCCW21]. Grounds [NSPRSS20]. group [GPH^{+}21, HW22, LSY20]. Grouper [CNB^{+}23, ASM21a, HDAPG^{+}24, RYS^{+}21, Squ20]. groupers [dMLB^{+}20]. groups [BC20b, DSL^{+}21, FS22, MPSF20]. Growing [AML^{+}21, HWMO20, SLK^{+}23]. growth [AB20, BBWC21, BHB^{+}24, EAA^{+}21, FZYF23, IAB^{+}20, JK22, LXSX21, LXW^{+}21, MCKZ21, SDI^{+}20, Sha20, SGK21, SKE22, WW21b, ZT21, ZZD^{+}23, HvT22]. Guaceto [Rus20]. Guadalupe [MAMCHP^{+}20, SMHPMO^{+}21]. Guadeloupean [MLJ^{+}22]. Guamuchal [GBC21b]. guaranteeing [WS23]. Guard [Lin24]. guards [GC22]. Guatemala [GBC21a, GBC21b, OWCFDC21]. Guérinière [CBAPM20]. Guianas [LFLF22]. Guidance [DSC^{+}23, vdBSvdV^{+}23]. guide [BBBS22, GCM^{+}23, Wan23c, IMA22]. guidelines [AdSMS21, NH20, NCE21]. guides [SKS21]. Guilds [HRMPCRRA22]. Guinea [HFD^{+}20, Kas21, OYECdOP24, SMKS22]. guitarfish [SCBB^{+}22]. Guitarfishes [SSH^{+}23, CJCS^{+}22, SCBB^{+}22]. Gujarat [KSV^{+}23]. Gulf [ABDRV^{+}23, HGG^{+}22a, MLMD21, MMC^{+}22, QTBL23, WLW^{+}22, ALLR^{+}24, SONB^{+}22, BCMCM21, CLL20, CAGBM22, CCKK22, GMPT^{+}20, GTA^{+}20, GBSC22, GWCD22, GHK20, Hon20, Kas21, OYECdOP24, PKH22, SLP^{+}22, SSF23, TLT^{+}22, VCC^{+}21]. gull [TDC21]. Gunther [Cha20b]. GVC [Seu22].

Habitat [MFRQ23, PBH^{+}23, CLD^{+}24, GA20, HLW^{+}23, Lot22, MKAF21, WD20]. habitats [DGIS^{+}22, FARBV^{+}24, MPD^{+}23, NJBM^{+}23, SSMG21, TBPC20]. Hadjimichael [Sta23a]. Haeundae [CLC21]. Hainan [ZJ20]. hake [BDCJ22, GNMGWM21, May22, MKM^{+}20, RQB^{+}20]. Hakodate [TM23a]. halibut [LJ21, SSAP24]. Halophytes [CLCV21]. hammerhead [ALLR^{+}24, Mag23]. hamper [OCD^{+}23]. Handbook [Tan23b]. Handl [Cha20b]. handling [Okş23]. happen [NGRM^{+}22]. Harassment [MLE^{+}24, ÖB22, PK20a]. Harbour [JAH^{+}24, OGGC20, GAA^{+}20]. harbours [VGS^{+}20]. Hardback [Rub24]. Hardcover [Sta23a]. harmful [CMSN^{+}22, EHLT21, FR23, SS21a, SSV^{+}23, SLMG^{+}21, VSDR^{+}22].

Harnessing [PRM^{+}20]. Harris [Smi20]. Harvest [HM22b, CC21, CCK^{+}21, CLL20, GML^{+}24, Mot20, Mot22, Muk21, SBD^{+}23]. harvester [MFG^{+}23, SVP^{+}23]. harvesting [BADO20, EHLT21, SVP^{+}23]. hashtag [WFR23]. hatches [SMF^{+}23]. Hauraki [QTBL23]. havens [HMY^{+}23]. Hawai’i [Nal24, SC21, ALML23, Cha20a, Lev20]. Hawaiian [BMR20]. Hazard [TSLK23, KWG23, MDLA21]. hazards [IRK^{+}21, TFMT^{+}23]. Head [Cat20, SEG^{+}22]. Health [MK24, Sha22, BDM21, Fen23, GWCD22, LCL20, PGK22, PK20a, Pyć23, SNS23, SHZ^{+}21, SKW^{+}22, TAK^{+}22, TSH22b, VDP^{+}22, WW21a, dO23, YCBW21]. healthier [dO23]. heart [HJ21]. heavily [GCV^{+}20]. heavily-used [GCV^{+}20]. hedging [HJPL23]. Hedonic [NTH^{+}21, Cat20]. heel [BD23]. hegemony [SHPW21]. help [CBR21, CTC^{+}23, Fen23, MZ24, SLMG^{+}21]. helped [SSB^{+}22]. Helping [Bea22]. herald [GN22]. Heritage [LC23, BEH^{+}21, CSH23, DBH^{+}22, GCL21, GW21, KOCL23, LHW21, LKB^{+}23, Lin24, ML22, PT23, SRS23, dS21, WM22a, War23]. hermeneutical [SCJ22]. herring [Hor23, THGR^{+}20]. heterogeneity [IGO21, JK22, SGK21, SSK23, WNM22]. heterogeneous [NNRS24]. Hexanchus [NGT^{+}23]. hidden [CvHF^{+}24, DBH^{+}22]. hierarchical [DL22, MF23, PBT^{+}20, YYD23]. hierarchy [GCB^{+}23, HF20, JL21c]. High [TCZ21, Alf20, AFHHD22, BMR20, CMD^{+}20, DBH^{+}22, GR20b, Gul23, Lam21, MCPS20, MLAS23, MSGAR20, PCA^{+}23, Ros22, SSH^{+}23, TSWC23, WvdMG^{+}21, WLY21, vdGD21]. high-modernist [Alf20]. high-quality [WLY21]. high-seas [AFHHD22, vdGD21]. higher [MM20, WSD21]. Highlighting [Nal24, VBC^{+}20]. highlights [PSTW21]. highly [BBK^{+}21b, VKN20]. hilsa [INW20, DAAP21]. Hilton [Cat20]. hinder [PGSFRMHH24]. hindering [Bea22]. hindrance [CTC^{+}23]. hippurus [MAG^{+}22]. hired [CSGC22]. historic [BMK^{+}22, SMKS22]. historical [BMMF22, GABA^{+}22, GMVM23, SPHO23, WA24, dPJ21]. historically [MVR21b]. history [ADB^{+}22, BLH^{+}20, Laj21, LLLF22, MN23, WWC23, vHSSK20]. holdings [BK23, SS22]. holistic [FOH^{+}20, FMCT^{+}21, SEG^{+}22]. Homarus [MM20]. Homer [Bir20]. Hondo [MTC^{+}23]. Honduras [CAGBM22]. Hong [Zhu22]. Hooks [MUR^{+}21a]. hope [Pol23, TMDN23]. hopes [Gul23]. horizons [HRS^{+}24]. Hormigas [HSGNMGC21]. Hornidge [Sta23a]. horse [MKM^{+}20, SGR21]. Horticulture [CTIH21]. host [YNZ22]. hostility [PGP22]. Hot [SHFP21]. Hotspot [CRP^{+}20, FSF^{+}21, MN23, WvdMG^{+}21]. hotspots [RXN23]. hours [BBK21a]. Household [KJS^{+}22, KPN20, NRYG20, TVN23]. households [KDVR20, SPA^{+}21, TSS^{+}24]. housing [CBAPM20]. Huaura [JTR^{+}20]. hub [HLLK22]. Human [BDM20, KHMM20, MGF20, SRY24, AVF^{+}23a, BDM21, CK23, CLB^{+}20, CWR20, EMN^{+}24, GCF^{+}23, GE23, HJPL23, HACCM22, MIMS^{+}20, NOB22, OAAMS24, PCA^{+}23, RSAP22, SMC^{+}21, SSMG21, SPJB22, TLT^{+}22, TDC21, WPWM^{+}22, Who23, WKBS20, WBH^{+}23, BAP21, PDJ22]. human-animal [HACCM22]. Human-crocodile [KHMM20]. human-nature [PCA^{+}23]. human-ocean [HJPL23, WPWM^{+}22]. human-rights [OAAMS24]. humanitarian [De 21]. humanity [TSA^{+}22]. Humanizing [CHB23]. Humpback [GCV^{+}20, FMCD20, GHK20, MCW^{+}22b]. humphead [HdM23]. Humphreys [De 20]. hundred [Bau20, Liu21]. hundred-year [Bau20]. hunting [EVB21]. hybrid [Cla20]. Hydrate [PY20]. hydrothermal [GCG^{+}21]. hydrothermally [VCC^{+}20]. hypoxia [Muk21].

ice [TBV^{+}23, WCT^{+}20]. ice-breeding [WCT^{+}20]. Iceberg [WD22]. icebreaking [BOW^{+}22]. Iceland [GV22, SDI^{+}20, VJO23]. Icelandic [BDDE20, BADO20, BAD21, EAGV20, Mat20, Run24]. ICES [LNRM23]. ichthyofauna [AVF23b]. ICJ [Shi23a]. ICZM [BMM^{+}24, OGGC20]. Identification [DA23, DL22, FDGM24, Kuł22]. identify [ES22, HLW^{+}23, MFG^{+}23, dSOW^{+}21]. Identifying [FHBS23, GWAPT21, KVL^{+}23, OCCSC23, PWL^{+}23a, QCM^{+}20, WFG22a, GCG^{+}21, JHHG23]. ignore [JGWM^{+}23]. II [HdM23, Sel20]. II-listed [HdM23]. II-Listing [Sel20]. III [Fra23]. ilisha [DAAP21]. Illegal [GASLCP20, KgJqY^{+}24, SH23, TSH^{+}22a, ABDRV^{+}23, AGS20, BJH21, DGG22, GAD^{+}20, JGWM^{+}23, KL21a, KAN^{+}23, Ma20, MJMAA21, NBB^{+}23, Pat22, SL23, SCD21, SBFHD^{+}23, TGM22, VVR^{+}22]. Illuminating [TOL^{+}22]. illusions [RCH22]. illustrate [KBD^{+}20]. Imanga [KDH^{+}21]. imbalance [SL23]. IMO [BH23a, BAP21, BÖ24, CR23, DG24b, FFE^{+}24, Kar22b]. Impact [IKB21, MMM^{+}22, MCBN22, PBD^{+}23, RAFLRP21, SC22, SLY21, TSS^{+}24, WHL22, AED20, Álv23, AS23a, AH21, BG20a, CHEC21, CAVJ^{+}20, CWKW20, CHC^{+}21, CM23, DGG22, DBS22a, FR23, FS22, GMGLV22, GMCMJA^{+}21, GWCD22, Guo20, HJLN24, JL21a, JACD20, Kaw22, LDL21, LJL22, MLVT21, MSSS22, MPL^{+}23, MX21, Muk21, OYKBA22, PGK22, RXX24, RDD^{+}21, SDG^{+}23, SC21, SSP^{+}21a, SRSGMGG21a, Tan22, TSH22b, UOO20, WCC^{+}20, WRL23, Wan23c, YFD^{+}21, Zha22, ZT23a, dOPM^{+}20, dlPdlLLC^{+}22, CDC21b]. impacting [NEF23, SMKS22]. Impacts [AEL^{+}23, CWKE20, DAHS^{+}23, GNMGWM21, HHB^{+}21, LKE^{+}22, MLE^{+}24, NAS^{+}22, RKR20, SdCMN^{+}22, ZTZ^{+}21, AELR^{+}20, AR23a, AR23b, BMR20, BRM^{+}21a, BSY^{+}20, BRP^{+}23, BASM22, CAM^{+}23, Cha20a, CLD^{+}24, DDGdS^{+}21, Dun21a, Ero22, EP20, EAGV20, GAD^{+}23, HCWC21, HPC^{+}23, HBP^{+}21, Kob22, KHLS23, LSB^{+}21, MMB^{+}23, MCPS20, MN23, Mil22, MMK^{+}22, MZT^{+}23, Nal24, NVP^{+}23, OR21, PCG^{+}23, PV23, RGK^{+}20, RLRS21, SCSR^{+}22, SCSR^{+}23, SOP^{+}22, SBFHD^{+}23, SBT21, SGFM^{+}21, SNK^{+}20, SRH^{+}22, SSP^{+}21b, SRSGMGG21b, TSH^{+}22a, TLT^{+}22, TKP23, TGHF^{+}22, WABR^{+}22, WNMS^{+}24, WLXY22, ZD20a, ZCS^{+}23, ZNA^{+}23, dO23, LEERRM^{+}21]. impairs [RPY^{+}24]. impede [KS22]. impedes [TUBA22]. Impediments [WH20, QLZ^{+}24]. impetus [KK22]. implement [GTA^{+}20, HC22, MFRQ23, PŚ22, Tra23b]. Implementation [ANS^{+}20, COJS^{+}22, KWY^{+}22, AMR21a, ALF^{+}22, BKW^{+}24, DG23, DLX21, EdLP^{+}21, GBCNCL^{+}21, HWU^{+}23, HZ23, INW20, KCCC^{+}21, LZRC23, MLMD21, MMC^{+}22, MB22, NAP^{+}21, OPL20, OWS20, OSV21, PS21, QTD21, RR20, SBP24, SS21b, SGL20, VBH^{+}23, VJO23, VvSvH^{+}22, YHU22, ZSLY20]. Implementing [BJH21, Cra21, DFTK20, LXH^{+}22, Moh23, TZZ^{+}21, THGR^{+}20, DLJ^{+}20, NCE21, RBB^{+}24]. Implication [NK21, IMA22]. Implications [CNB^{+}23, DER^{+}21, FMS^{+}22, Kra22, LBWS23, Puh23, TKP23, WL20, Wat22, AKL^{+}22, BHdMCM^{+}21, BTV^{+}22, CDP22, CNG^{+}23, CLCV21, DBH^{+}22, DPG^{+}22, DMS^{+}24, Dua22, EB23, FRVB22, GJY23, Gun21, GHK20, HPS23, Has20, HLLK22, JGWM^{+}23, KL21a, KSS23, Liu23, MMW^{+}23, NUD^{+}23, PL24a, PKC23, Pra24, PMFR^{+}22, RLRS21, RB22, SDI^{+}20, STS^{+}23, SKW^{+}22, SGL20, TKC23, TCC^{+}22, TBV^{+}23, THS^{+}24, VSD^{+}23, YBW20, YCBW21, ZPX^{+}22, ZSLY20, ZY21, AGN^{+}20, Boy21, GMPT^{+}20, HWMO20, OWCFDC21, RGK^{+}20, TMTC20]. imply [SFN^{+}21]. Import [BGK^{+}22]. importance [AAFK^{+}21, CLB^{+}23, CBD^{+}23, DGS^{+}24, MWM^{+}23b, NEH^{+}23, TN21, WvdMG^{+}21]. important [COJS^{+}22, QLU^{+}23]. imports [AWX21, WDG^{+}20]. imposition [Ros22]. improve [AACdlT23, AAW^{+}23, BGK^{+}22, CCS23, CDC21b, DRCdMCJ21, FLB^{+}21, MPH^{+}23, MT20, OYKBA22, WKH22]. Improved [HRC^{+}24, AGK22, BGMCG24, BDC^{+}21, DJC^{+}23, HBP^{+}21, YHU22]. improvement [AHAB^{+}22, GWAPT21, KOBS21, WS23, ZGY21]. Improvements [AO20]. improves [SDG^{+}21, SDG^{+}22]. Improving [BHCL23, CJ22, GPM23, HCT^{+}21, SCBdA20, AFHHD22, BDD^{+}21, GBB^{+}21, MPD^{+}23, VHD20]. in-stream [CA20]. inactive [JG20, VCC^{+}20]. inadvertently [SLL^{+}24]. incentive [JL21b, JZW21]. incentive-compatible [JZW21]. incentives [BC21, CGGA20, HWT23, OSV21, PGGA22]. incident [SCGC20]. incidental [BUM^{+}24]. incidents [JG23, PKC23, Shi22]. Including [DSQ^{+}22, FTW^{+}23, TCZ21]. inclusion [AFHHD22, ML21, PALB20, RBA23, RMF^{+}21, TEA^{+}22, dS21]. inclusive [AC23, LHTR23]. inclusivity [JAK^{+}23]. incoherence [EB24]. income [AS23a, IOHC22, KTI^{+}24, MDG^{+}23, Seu22, TVN23]. incomplete [GOH^{+}20]. inconsistency [hBDF22]. Incorporating [Kar22c]. increase [CSH23, MM22, MWB^{+}21, SKE22]. Increased [Boy21, LJL22, SHP^{+}21]. increases [MJ20, MHN^{+}20]. Increasing [CVA22a, TJZ^{+}21, AH23, Sar23]. increasingly [TJ21a]. independent [CdLN^{+}21]. Index [EV22, RP23a, Smi22, Smi23b, FMGV20, Sta23b, WD23, COJS^{+}22, LMSLMQ21, PDJ22, RP23b, SGWL21]. India [BPA^{+}22, CDP22, DT23, DMS^{+}24, DDD22, DAAP21, GSA23, KHMM20, KSV^{+}23, NRF21, PRM^{+}20, RBA23, Sol23b, SPA^{+}21, TJS22]. Indian [KWF^{+}21, Alf20, BLH^{+}20, DXS^{+}21, GSA23, Kar23, MDG^{+}23, McC24, RPK23, SBS21]. Indicating [BE20a]. Indications [Gir22]. Indicator [Liu23, COJE24, CD24, TBV^{+}23]. Indicators [BRR^{+}22, MVR21a, MKB21, WNMS^{+}24, DAA21, DDGdS^{+}21, EAAG21].

Indigenising [RRM^{+}24]. Indigenous [BMM^{+}20, LMMT23, MWM^{+}23b, AHB^{+}23, SPN^{+}21, WCM^{+}24, BMB^{+}21, DBH^{+}21, MRD^{+}20, VHS^{+}20, WJLSH21]. Individual [DP23, HvPPV20b, DBB^{+}22, HWT23, SSAP24, Zha22]. individuals [SOP^{+}22]. Indonesia [ABI24, AAKI21, AHAB^{+}22, APF22, AGSI22, ANR^{+}22, CLRW22, DDAB^{+}23, DHR23, EAA^{+}22, EABA^{+}22, HLW^{+}20, HYN^{+}23, HF20, JSW^{+}22a, JPM^{+}23, KMG^{+}20, KgJqY^{+}24, LLLF22, MKAF21, NRYG20, PN20, PMM^{+}21, QLU^{+}23, RBEA^{+}22, RYS^{+}21, SMDM20, SIW^{+}21, SWI^{+}22, SLW^{+}23, SWH^{+}23, SGL^{+}21, TUBA22, TSS^{+}24, THT^{+}23, TEA^{+}22, VKS23, WMRB22, YC21b, Zou23]. Indonesian [AFS^{+}23, DSWB20, Dun21b, LWSS21, LTW23, Lin20, PFWM23, RMU^{+}22, VKS23]. induced [FAC^{+}24]. industrial [AE21, CQ20, CCKK22, KLSG21, LJL22, NSPRSS20, RHO24, Tra23a, ZW21]. industrialization [AML^{+}21]. industries [LXSX21, TM23a, TFF22, YZH22]. Industry [TYS22, AOO^{+}23a, BDR20, BMH^{+}22, Cha22, CQ20, CGGA20, Fla21, FR23, Fro21, FBC^{+}23, Gar24, GG21, HPS23, HPG20, HJLN24, HHB^{+}21, HPH22, HKKY22, JFB^{+}24, JLY23, JD20, KCCC^{+}21, KSS23, Kiz24, KF23, Koe22, LWSS21, LTW23, LLHG23, LJ22, LJL22, LC22, LFPBD^{+}20, LKE^{+}22, MGF20, MMO^{+}24, NSN22, PAK22, PPK23, RU20, Ren21a, Ren21b, RDM^{+}20, RAB22, SDI^{+}20, SLY21, SZ20, SLS^{+}20, SLS^{+}21, Ton22a, WL22b, XHY22, ZCCQ22b, ZZW^{+}21, ZWC^{+}23, ZSLY20, ZCS^{+}23, vLPBK21, SDI^{+}20, LJL22]. Industry-initiated [TYS22]. inequalities [IOHC22]. inequality [Mag23, MLE^{+}24, TVN23]. inequities [MTC^{+}23]. inertia [SGR21]. infectious [Bov22]. infer [CAVJ^{+}20]. infinite [PCA^{+}23]. Influence [KTI^{+}24, WZZC21, ZW21, AB20, BAP21, BMR23, HT23, QF20, WDG^{+}20, WWD23]. influenced [YNZ22]. influences [YSEK22]. influencing [AAKI21, Dam24, HMFH20, JZW21, KWMM20b, MLMD21, Pra24, PPK23, XHY22, XXLW24]. inform [BBB23, BEM^{+}20, MMNC22, Nal24, PBS^{+}23, SMLRB^{+}21, SSOP^{+}21, SMW^{+}21, VCC^{+}20, WD20, WNBRSG22, ZCG^{+}23]. Informal [MJ23, Pau21, GBSC22, IO23, PL20b]. informalisation [YS23]. informality [GWAPRPRU22]. Information [dSOW^{+}21, ATF^{+}21, CMR21, DQJ23, DSC^{+}23, FLC^{+}23, GSB20, LR20, NM23, RPY^{+}24, SDG^{+}21, SDG^{+}22, WM21]. informatization [JL21a]. Informed [AAA^{+}21, GAD^{+}23]. informing [WtDD^{+}21]. informs [BL23a, MDG^{+}23]. infrastructure [AH23, BL23b, BS22, CDRS22, HLLK22, RKS^{+}23, ZCG^{+}23]. infrastructures [Alf20]. infringement [Gar22]. ingredients [Lev20]. Ingvild [Niu22]. Initial [RDD^{+}21, MMNC22, PWL^{+}23b, Tan24, ZTZ^{+}21]. initiated [TYS22]. initiative [LMC^{+}23b, TSLK23, Wan23c, Ber23, DFTK20, TBPC20]. initiatives [MGDJ22, VVR^{+}22]. injuries [Dev22]. injustice [WWCS21]. inland [OOSG21, PPK23, Zha22]. Innovation [AOO^{+}23a, PVH^{+}21, BBGHR23, FWL^{+}22, GLS^{+}21, GVR20, NWZ^{+}21, PSD23, Pen22, RJ21, Sha20, SLY21, SKDK23, VMD^{+}24, WSD21, ZW21]. innovations [HCW^{+}23, LMMT23, SDG^{+}23, SR22]. Innovative [ABVPM20, TT22]. inorganic [TAP^{+}23]. Input [BSSL^{+}20, dlFGFFR^{+}20, BG20a, SG21, SRSGMGG21b, TM23a, ZCCQ22a, ZT21, ZZD^{+}23]. Input-Output [dlFGFFR^{+}20, BG20a, SRSGMGG21b, ZT21, ZZD^{+}23]. inquiries [CVA22a]. insecurity [NRYG20]. inshore [BCJ21, CWMM^{+}21, SKS21, SSAP24]. insight [OC20]. Insights [AAA^{+}24, BOOO23, Bir20, DT23, DBBL21, MKM23, MZ22, MPH^{+}23, OCD^{+}23, PMC24, CGHV21, GPG^{+}21, GCM^{+}23, OGS24, RDD^{+}21, SKDK23, Tho22, VWJMAJ20, BCBS22, BBS^{+}21, CBO^{+}21, GRA^{+}22, HMG^{+}21b, IRK^{+}21, JZW21, KK22, KDVR20, SGR21, SL20, SSB^{+}22, SSMG21]. Inspection [XWL^{+}20, Öİ24, Şan20]. inspections [Kar22a]. inspiration [Ada22, LMT21]. instability [BLM24]. instigating [FGK^{+}20]. Institute [HH21]. Institution [BV21]. Institution-based [BV21]. Institutional [AHM22, BDHI21, Cla20, Kra22, SGR21, SG20a, YZH22, AGSI22, BPA^{+}22, BSM23, BUR^{+}22, CDG^{+}21, Cra21, DP23, GLADGGVL23, RRR^{+}23, SL23, SVF20, SOS20, WH20]. Institutionalising [GNJR24]. Institutionalization [ONN^{+}20]. institutions [Ber22, GBSC22, HZ23, Kat20, PL20b]. instrument [CWR20, Gar20, LR20, Lou22, MFH^{+}20, Pri22]. Instruments [PŚ22, Aqu23, BSSL^{+}20, Kar22b, PMFR^{+}22, RLS20]. insurance [JF20, Set23, ZSZ20]. integral [BRR^{+}22]. Integrated [HIP^{+}21, KSB^{+}23a, AHM22, BBGG22, EdLP^{+}21, KABT21, LLCH21, LZL23, PMR23, SGFM^{+}21, TCS^{+}20, TFMT^{+}23, ZCA^{+}23, dSdLSGB22, CAGBM22]. Integrating [APV^{+}20, BEH^{+}21, LKB^{+}23, MM23, PHF20, PKS20, SSF23, HvT22, SRR^{+}21, MLVT21, PFG^{+}21]. Integration [AS23b, JKV^{+}22, Koj23, Kon20, BC20a, GGP^{+}22, IOHC22, QYX21, RLS20, VHSV20, XXLW24]. integrity [FTW^{+}23, FOKM23, PSTW21]. Intelligent [ZLWY22, THP22a, TC23]. intensity [SS21a, TBV^{+}23, BAD21]. intensive [KMM21]. intention [DLYM22]. intentions [GPM20, MLM23]. Inter [Har22, HSdT21, RKS^{+}24, SRY24]. inter-organizational [HSdT21, RKS^{+}24]. Inter-state [Har22, SRY24]. interacting [POM24]. interaction [Kar23, TDC21, YWYL22]. Interactions [RTMCVMDE22, TH22, BPA^{+}22, BL22, HTP^{+}23, HACCM22, OPL20, RHST21, SVHV20, SMC^{+}21, WPWM^{+}22, BMM^{+}24]. interactive [CRP^{+}20, LJ23]. interacts [BOW^{+}22]. Interconnection [AB20]. interdisciplinary [SRH^{+}22]. interest [BAP21, Mar21a]. interests [Has23]. interface [CMS^{+}21, GSA23, PKH22]. interfere [CMOJ^{+}24]. intergovernmental [DMNT20, MDJ^{+}22, TMDN23]. intermedius [GPH^{+}21]. International [CCX23, Cha20b, HCW^{+}23, Har22, HH21, KMT^{+}21, LNRM23, QLZ^{+}24, Shi23b, Tan23b, TCZ21, WL20, Wan23b, War21, ZY23, BBK^{+}21b, CZW20, Cha22, CMWN^{+}23, Che21b, CWR20, Dom23, DC23, Gar22, GY23, Gra21, Gue23, HCSBM^{+}23, Kle21, KL22, Koj23, MAG^{+}22, Pri22, Sam22, SSH^{+}23, Shi23a, SSK23, SK24, SLH22, TKC23, Ton22a, TC23, Urb22, VMV^{+}24, WRL23, Wan23a, WdSRY23, ZT23b, vdGD21, AR22, BDC23, BHKJT23, BRM21b, GSM21, Kuł22, Mar21a, MK24, PZPB21, RJAT23, Sin21]. internationally [BLM24]. Internet [AAW^{+}23, PGK22, TJZ^{+}21]. interoperable [LHV^{+}20]. interplay [Hor23, PL20b, SLB21]. Interpretation [PL23, WL20]. interpretations [BDM20]. intersection [GPA^{+}22, vdGD21]. Interstate [PGP22]. Intertidal [PJN24]. intertwined [Tru23]. intervention [Fen23]. Interventionist [LTW23]. interventions [ALML23, BOOO23, FMWW21, GHW^{+}22, MS21, PBG23, SLB21, VWJMAJ20]. intra [Har22]. intra-state [Har22]. introduced [SPN^{+}21]. Introducing [Liu22, TCT21, WCC^{+}20, Åm21, Jen23, NHA23, RAB22]. Introduction [Ehl21a, GAD^{+}23, PZPB21, SNS23, EAAG21]. intrusion [RKR20]. Inuit [BOW^{+}22, KDH^{+}21, SCF^{+}22]. invasions [SMF^{+}23]. Invasive [MMB^{+}23, AC21, CAHR^{+}24, GTTM21, SSDQ20, SMC^{+}21]. invertebrate [CCK^{+}21]. invertebrates [OVSVPP^{+}22]. Investigating [GTTM21, TBV^{+}23, SS22]. investigation [FFE^{+}24, Han20]. investment [AED20, FGPPPV21]. investments [BXV20, SDI^{+}20]. invisible [IO23]. involved [MMK^{+}22]. involvement [LEVET21, STIS20, vRB24]. Involving [DSL20, HS20, DBBL21]. iObserver [VAMR^{+}20]. IOCD [Bov22]. IPCC [CTH^{+}21]. Iran [AB20, PAK22]. Iranian [HGG^{+}22a]. Ireland [HFD^{+}20, AH23, AHM22, CCW21, CM23, MMNC22, OCD^{+}20, OGGC20, dlÁPAFGO20, vRB24]. Irish [CPR21, OC20]. Irregular [MSSS22]. ISA [MB21]. ISAV [BVvKL21]. ISBN [Cha20b, HH21, Lv20, Ngu23, Niu22, Rub24, Sta23a, Tan23a, Tan23b]. ISIS [PMT^{+}20]. ISIS-Fish [PMT^{+}20]. Isla [MTC^{+}23, GASLCP20, dPJ21]. Island [BMH^{+}22, Cat20, CTIH21, DRY22, KWG23, MAMCHP^{+}20, Moh23, SMHPMO^{+}21, SOP^{+}22, AHM22, AL20, COSS21, KMS20, NRYG20, NSN22, NNR^{+}21, PMR23, RMU^{+}22, SMKS22, TFH^{+}23, ASM^{+}21b, IGS^{+}23, Kak20, KK22, Los23, Los24, MZ24, OAAMS24]. island-based [RMU^{+}22]. Islands [LTJ^{+}23, PN20, SWW^{+}21, TJS22, APF22, DČFW23, THP22a, THT^{+}23, AGSI22, AM22, BDMW22, CTIH21, Chr20, DAA21, Dun21a, EB24, GBFMT23, GMGLV22, GMRSP23, GMCMJA^{+}21, MBORGN24, Mrd23, RGH^{+}22, SW22, Tod22b, Tra23a, YTBM^{+}22, ZMR20]. Islas [HSGNMGC21]. Israel [PC20, YNZ22]. Issue [EB23, BR22, Ehl21a, GAD^{+}23, PZPB21, SNS23]. issued [BRM21b]. Issues [BKW^{+}24, AR22, CHB^{+}20, CLL22a, GTMP20, HCW^{+}23, MMNC22, PGA^{+}20, Lv20]. Isurus [DSL20, Kai21]. Italian [VMD^{+}24]. Italy [Mar21a, Mor20, RBB^{+}24, VMC^{+}21]. ITQ [GV22]. ITQ-managed [GV22]. IUCN [ABM21, Kai21, PNS^{+}23]. IUU [Faj22, GÁPSAP^{+}23, GY23, HM22c, HMY^{+}23, JS23, KL21a, KA21, Oan22, Tei23]. IWC [Gal22]. ix [Smi23b].

Jaeckel [Tan23b]. Jakobsen [Niu22]. Jamaica [LZRC23]. James [Tan23a]. Janeiro [SdACdSLdAS24]. January [Ano20o, Ano21r, Ano22q, Ano23q, Ano24g]. Japan [Kim21, AKFY22, BI21, FMS^{+}22, HWU^{+}23, HT23, HWMO20, IMA22, JY20, KS23, Kim20, Kob22, KWY^{+}22, MM21b, OKH^{+}24, OI23, PL24a, SII22, TYS24, TM23a, TTA^{+}20, Tru23, TYS22, WM21, WNM22, XTY22, ZTA23]. Japanese [HWT23, KYMW21, KB20, KN21, ONA21, SRH^{+}22, TZM^{+}22, TSWC23, WS21]. Jariego [TGM22, SCD21]. Jelly [MSZ^{+}23]. Jelly-FAD [MSZ^{+}23]. Jellyfish [KSV^{+}23, BCMCM21]. Jiangsu [XWH^{+}22]. Jiaozhou [SL20, WLX21]. Job [PSPR20, ZWXL21]. Johansen [Niu22]. John [De 20]. Joint [Lin24, CHB^{+}20, Kim20, CHB^{+}20]. Jolla [KL21b]. Jor [AAKI21]. Joseph [Lv20]. judgement [Pen22]. Judicial [CCX23, Li23b]. July [Ano20p, Ano21s, Ano22r, Ano23r]. jumbo [GWAPRPRU22]. June [Mar21a, Ano20q, Ano21t, Ano22s, Ano23s]. juridical [Gul23]. jurisdiction [BLL22, CWR20, DMNT20, DBZ24, FMS^{+}22, GNJR24, Gar20, HJ21, HM22a, HGL^{+}20, MDJ^{+}22, MH24, MCR22, OKR^{+}20, Pat22, Pri22, TD21, TMDN23, VHD20, VBC^{+}20, WGJ^{+}21, YG20, ZJ22, dO23, Has23, MTN23, Teo23, TM23b, TCD^{+}20]. jurisdictional [dO23]. jurisdictions [SHP^{+}21, YT20]. jus [Kuł22]. just [AC23, GABA^{+}23, KL21a, Lot23, SS21b, WLW^{+}22, SCF^{+}22]. Justice [Pen22, BC20a, BCN23, BBWC21, BAF^{+}23, Dau23, DUYEA21, GFL^{+}21, SCJ22].

Ka’an [MMSB^{+}20]. Kadavu [RGK^{+}20]. Kaohsiung [TCT21]. KAP [KCCC^{+}21]. Kate [AFHM23]. kayaking [AH23]. Kazakhstan [AB20]. keeping [vLPBK21]. Kei [THT^{+}23]. kelp [CTC^{+}23, GRPRVG23, GTTM21, HGG^{+}22b, HKKY22, QTBL23, TDC21]. Kenya [AOO^{+}23b, KWS20, KWMM20a, KWMM20b, LSB^{+}21, OWS20, OKNB22, OBNK24, OKFN24]. Kepulauan [HF20]. Kerala [DT23, NRF21, RBA23]. Kerguelen [TH22]. Key [HCW^{+}23, KKDK20, KSS23, MF20, NEF23, PGA^{+}20, AFHHD22, BN20, CCCC21, GCP^{+}20, KJS^{+}22, LSD^{+}22, MN23, SdACdSLdAS24, WFG22a, ZSLY20, Sin21]. Kil [Tan23a]. killer [Dam24, FMCD20, Shi22]. Killing [HM22c]. Kin [DCC^{+}20]. king [BRR^{+}22, DD22, VKN20, BRR^{+}22]. Kingdom [Cro21, FS21, KSB^{+}23b, SSA^{+}23, SPH^{+}20, VSD^{+}23]. Kiribati [Mal21]. Klang [RGH^{+}22]. Klein [Tan23b]. Knowledge [HBP^{+}21, MN23, MWM^{+}23b, AAKI21, Alm21a, AELR^{+}20, AHN^{+}23, ALLR^{+}24, BC20a, BB21, BMK^{+}22, BVT^{+}21, CPV^{+}21, CAVJ^{+}20, CSVC23, DMd^{+}20, DSQ^{+}22, DG24b, DA23, FBC^{+}23, GNJR24, GSAF21, GCP^{+}20, GSM21, HTP^{+}23, JK22, dSJNS20, KCCC^{+}21, Kar22c, KDH^{+}21, MMB^{+}23, MFG^{+}23, MWM^{+}23a, MMS24, MFH^{+}20, NGT^{+}23, OCZ^{+}22, PMM^{+}21, SBFHD^{+}23, SCCW21, SVP^{+}23, Sta23a, ZYTT21a, ZYTT21b]. knowledge-base [GSM21]. knowledge-based [CPV^{+}21]. Known [HSGS^{+}23]. Kong [Zhu22]. Korea [Kim21, CKY20, CLC21, CKY22, HS22, JD20, KPN20, KNY20, Kim20, KNY22, Kim23, KSS23, KLK23, KGP^{+}22, LrS22, PL24a, PKR^{+}22, PKC23, YK20]. Korea-Radionuclides [PL24a]. Korean [hC24, KLS22, KHH^{+}23]. Kraska [Tan23a]. krill [CMC22, Tra23b]. Kristoffer [Cha20b]. Kuwait [AGS20]. label [AH21, BdLAdC^{+}20, CMR21]. labeling [GVL22, WEF^{+}21]. labelling [CMNH21, JAK^{+}23, SFN^{+}21, WWC23]. Labels [PMK^{+}21]. labor [LXSX21, WKBS20]. Labour [CK23, PGSS22, SMCW22, SRY24, TOL^{+}22, WS23, YS21, YL21, KL21a]. Labrador [NGRM^{+}22]. Lacepède [dSJNS20]. lack [TMDN23, VKN20]. lagoon [AC21, BASM22, Han21, RB22]. Lagos [BRR^{+}22]. laissez [KK22, YL21]. laissez-faire [KK22, YL21]. Lancaster [KDH^{+}21]. Land [SVHV20, CACJR^{+}21, Gug24, JKV^{+}22, LSH21, MX21, OPL20, PMAVM20, SVF20, SL20, SCF^{+}22, BMM^{+}24]. land- [CACJR^{+}21, SVF20]. land-based [Gug24]. Land-sea [SVHV20, OPL20]. Land-Sea-Interactions [BMM^{+}24]. landing [CLB^{+}20, FGPPGG23, GMPT^{+}20, OSV21, PMAVM20, PV23, SSZ^{+}24, SOLN21, YU20]. landing-ban [CLB^{+}20]. landings [AFMdSdTdCC22, IGO21, SOLN20, TDP20]. Landsat [XWH^{+}22]. landscape [DBH^{+}22]. landscapes [HM21, PFM20]. lanes [VKS23]. language [ASKI20]. Lanka [AEL^{+}23, CLB^{+}20, CLB^{+}23, GRS^{+}23, JAH^{+}24, Kul23, NEH^{+}23, PDJ22]. Lankan [Pra20]. Lao [KJE21b]. Laoang [PCA21]. Large [ASGZ22, SLP^{+}22, WWG^{+}23, ABM21, DLJ^{+}20, HYN^{+}23, HLW^{+}23, HSH^{+}21, HLO^{+}21, KNY20, KLK23, MZT^{+}23, PBG23, Sol23a, WD23, dO23]. Large-scale [ASGZ22, DLJ^{+}20, HYN^{+}23, HLW^{+}23, KNY20, MZT^{+}23, PBG23]. largest [CLP21, EMH^{+}23, NBB^{+}23]. last [KGW20, LCMBGP22, VMC^{+}21, WEF^{+}21]. lasting [NSPRSS20]. late [GMVM23]. later [Ege23]. latest [LC23]. Latin [AE23, MBC^{+}24]. latissima [HKKY22]. launches [LP21]. Law [Kuł22, Ngu23, Niu22, Set23, Smi23b, Tan23a, TKC23, Álv23, BMR23, CLS20, CS20, Cha21, DC23, Ege23, Gue23, Hon20, KS22, Koj23, LC20a, Li23a, Lin24, NH20, SK20b, Shi23a, TZC23, Tse21, Wan23b, Wan23a, WM22a, War21, WSDB20, ZT23b, Gul23, Koj23, Mar21a, TH22, ZY23, Tan23b]. law-making [Wan23b, Wan23a]. Laws [RS24, CMNH21, CSHO^{+}20, Kle21, KGP^{+}22, LHW21, PK20a, Wil20, WEF^{+}21]. Lawscape [O’D21]. Lawships [GC22]. laying [KS22]. layout [LLHG23]. lead [MMK^{+}23]. leader [HWT23]. leaders [SSB^{+}22]. leadership [GFD21, HGG22c]. leading [DHR23, RE21]. learn [RMBR22]. learned [ADR^{+}21, BDMW22, BBJ^{+}22, CHC^{+}21, CNG^{+}23, DSL^{+}21, GLAVL21, GABA^{+}23, MF22, MCW^{+}22a, MCR22, NJBM^{+}23, SOP^{+}22, Squ21, VWFW22, VVR^{+}22]. Learning [BMH^{+}22, VBH^{+}23, CRJ22, OVA22, RB20, TCT21, YHU22, ZYTT21a, JF20]. learning-experiences [RB20]. learnt [ALF^{+}22, KSB^{+}23b, MZT^{+}23, PKB^{+}21, UISM^{+}23]. leasing [EP20]. led [DBS^{+}22b, Lin24, MGDJ22, RBA23, WdSRY23]. left [DMNT20]. Legal [Bov22, Cha20b, CWKW20, CLL22a, FSF^{+}21, MSSS22, Pie21, TN21, dS22a, BN20, BUR^{+}22, CC23, CZW20, CDRS22, CMD^{+}20, Gal21, Gar22, GMVM23, JMB^{+}20, Kar22b, Kar23, KL22, Kuł22, LC20a, Ma20, NCE21, OT23, PY20, PL23, RFPS21, SRvNN21, TK23, Urb22, XZ23, Cra21]. legalism [NW20]. legally [CWR20, LR20, MB22, Pri22]. Legislation [Lin24, TCZ21, GBC^{+}20, Kak20, KAN^{+}23, PMK^{+}21, Qaf24, SS23, Wil20, YT20, YD22]. Legislative [Li23b, DPB^{+}23, FFD^{+}23, KF23, Lov23, WCM^{+}24]. legitimate [Dau23]. length [ASM21a]. length-based [ASM21a]. lens [DDGdS^{+}21, FFE20, SK20a]. lenses [PDJ22]. Leone [OOGS20a, OOGS20b, OOSG21]. Lepidorhombus [APV^{+}20]. less [SMC^{+}21, SMCW22]. Lessonia [CBO^{+}21]. Lessons [ABDRV^{+}23, ALF^{+}22, BDMW22, CNG^{+}23, DSL^{+}21, FOKM23, GABA^{+}23, HNMSG21, KSB^{+}23b, MF22, MCW^{+}22a, MCR22, MZT^{+}23, NJBM^{+}23, QTBL23, RCM^{+}22, SOP^{+}22, UISM^{+}23, VVR^{+}22, BBJ^{+}22, CMGF20, DPP^{+}20, FSKJ23, KHdSM^{+}22, VWFW22, VBH^{+}23, ZT23a, ADR^{+}21, BC20a, CHC^{+}21, GLAVL21, KGW20, Kle21, OA20, PKB^{+}21, PGSFRMHH24, Squ21, SVKP21, SB20]. Lesvos [Sak21]. lethal [AMLC^{+}23, HSH^{+}21, SON21]. level [BADO20, CHEC21, DSC^{+}23, HJLN24, IAB^{+}20, JY20, JLY23, Kim23, Mal21, Mba22, MCKZ21, NSPRSS20, Ora23, QF20, RXX24, SPG20, SPMN23]. levelled [HvdBH23]. levels [AFMdSdTdCC22, AdSMS21, HLCM21, MMW^{+}23, ZWJvL20]. Leveraging [DKW20]. lewini [ALLR^{+}24]. liability [FY24, Fei21, MH23, TN21]. Liberia [AAA^{+}24, JKKT20, LA23, NAS^{+}22, WM22b]. lice [GVR20]. licence [KFM^{+}20, LMMT23, DN20]. license [Ber23, SS22]. licenses [OOGA22]. licensing [Her22]. Life [GCC^{+}21a, ADB^{+}22, BBBS22, EAA^{+}21, Fro21, LLLF22, MN23, MS21, WWC23, AP23, Gul20]. light [Mar21a, NH20, Pri22, Sar23, TLT^{+}22]. Like [Ree20]. likely [SPN^{+}21, VPRR^{+}21]. limit [GHK20, TYS22]. limited [CNB^{+}23, CDB20, Fuj21, GCS21, LLLF22, PWL^{+}23b]. limits [Aqu23, AMR21b, AS23a, Bau20, MBDG23, Sam22, SRY24]. Lindi [YMG23]. line [DAA21, LCAX21, MVR21b]. linear [MEG^{+}21]. liner [BHEE22]. lines [KKDK20]. link [CCK^{+}21, HHH^{+}23, RMBR22]. linkages [Chr20, CSL21, EB23]. linked [SRSGMGG21b]. Linking [GABA^{+}22, dOJCSLdSB21, MWM^{+}23a, MFM21, OMC^{+}22, SCA^{+}22, WGR^{+}22]. Lion [VCC^{+}21]. lionfish [YTBM^{+}22]. lions [PBH^{+}23, dOPM^{+}20]. Lisbon [SCA^{+}22]. listed [HdM23]. listing [Kai21, Sel20]. lists [CMNH21]. literacy [CSH23, LZL23, TSWC23]. literature [ABM21, DGMF^{+}20, HLLK22, Jen23, LMT21, LRC24, ÖB22, SM23]. Litter [MB22, AAA^{+}24, ADR^{+}21, Che21a, DPB^{+}23, HPC^{+}23, KLK23, MTC^{+}23, Mor20, ONN^{+}20, PKB^{+}23, PBG^{+}22]. littleneck [BRR^{+}22]. Liu [Sos21a]. live [CRJ22]. lived [UISM^{+}23]. livelihood [HP23, LRN^{+}22, Pha20, WLW^{+}22]. livelihoods [BOW^{+}22, CTH^{+}23, Han21, HBP^{+}21, LFLF22, MF22, Mba22, NAS^{+}22, SF22, SGL^{+}21, SNK^{+}20, YMG23, dlPdlLLC^{+}22, AFHM23]. livestock [TOHK20]. Living [GB21, Sos21a, Liu20, PCA^{+}23, SCCW21]. LMMA [LTJ^{+}21]. LNG [HPS23, MF23]. Lobos [dPJ21]. Lobster [KKO^{+}22, Bot20, DMS^{+}24, Doe21, Gor21, HS20, MM20, OAM20, OAM21, PLOG24, SOP^{+}22, TPG^{+}23]. Local [Alm21a, ALLR^{+}24, FTM^{+}22, FS22, MMB^{+}23, MPSF20, OCZ^{+}22, WJB^{+}20, YML^{+}21, YD22, ALML23, AELR^{+}20, BC20a, BMK^{+}22, CAVJ^{+}20, CS21, CSVC23, DB20, Dun21b, EVB21, HRR^{+}20, HTP^{+}23, JL21b, dSJNS20, KYMW21, Kar22c, Kat20, KHMM20, Kim23, LTW23, LRN^{+}22, LLCH21, Los23, NHD23, NAN^{+}24, NGT^{+}23, PCA21, Pha20, PJN24, SCCW21, SVP^{+}23, TCT21, VRBC^{+}21, VHS^{+}20, QUK21]. local-level [Kim23]. local-scale [PJN24]. Locally [LTJ^{+}21, MGDJ22, RGK^{+}20, MB22]. locals [ZNA^{+}23]. located [Lin20]. location [BLY^{+}24, COJZ^{+}21]. locations [RNGRC22]. lockdown [GS21, MCBN22]. logbook [SIW^{+}21]. logbooks [CLS^{+}23]. logistic [GSB20]. Logistical [BSY^{+}20]. logo [GVL22]. Lombok [AAKI21]. London [Ngu23]. Long [BSBD21, AFMdSdTdCC22, BCBS22, BMH^{+}22, CVA22b, Jon21, McC21, MM20, NSPRSS20, SOP^{+}22, SGFM^{+}21, Wat22, SOP^{+}22, Fla21]. long-lasting [NSPRSS20]. Long-term [BSBD21, AFMdSdTdCC22, CVA22b, Jon21, McC21, MM20, SOP^{+}22, SGFM^{+}21, Wat22, Fla21]. longer [Bra22]. longirostris [BDCJ22]. longitudinal [WFG22c]. Longline [OKFN24, BDC^{+}21, CD23, Cha20a, CHC^{+}21, DSL20, HCS21, KB20, LLL^{+}24, SCSR^{+}22, SCSR^{+}23, SBW^{+}21]. Looking [TTA^{+}20]. losers [BCMCM21, PKMT23]. loss [RWVH21, Tel23b]. losses [GAA^{+}20, YSWW21]. Lost [SG21, DAHS^{+}23, GRS^{+}23, GHW^{+}22, GAD^{+}23, HGG^{+}22a, Lov23, MDG^{+}23]. Low [ÁFFSC20, Wan23c, ZY21, AH21, DG24b, dlPdlLLC^{+}22]. low-carbon [DG24b]. low-environmental-impact [AH21]. Low-impact [Wan23c, dlPdlLLC^{+}22]. Low-quality [ÁFFSC20]. Low-sulfur [ZY21]. Luce [LTJ^{+}21]. lucens [LSY20]. Lucia [TJ21a]. Lyme [SJ21].

M [Sos21a, Sta23a]. Māori [RR20]. Macaronesia [GSGOP^{+}21, ANS^{+}20]. Macaronesian [GMGLV21]. mackerel [BDDE20, Gra21, MKM^{+}20, SLL^{+}24, SGR21, RHST21]. Macmillan [AFHM23]. Macro [BMC20]. Macro-plastic [BMC20]. macroalgae [KWY^{+}22]. macroalgal [KWY^{+}22]. Madagascar [LTJ^{+}21, LJRHH21, RRR^{+}23]. made [RBNE20, ZCA^{+}23]. Madeira [DDGdS^{+}21]. magnitude [GAD^{+}23]. main [BN20, CM23, GSR21]. Maine [BLS21, MM22]. mainstream [HM22c]. maintaining [AH23, McC21, ZD20b]. maintenance [KS22]. Major [SAA^{+}21, SFGEPC22, HCSBM^{+}23, LHPJ23, PL20a]. make [PK20b, RM21b]. makers [RB20]. Making [Men23, NGRM^{+}22, AMR21b, AH21, BL23a, BBBS22, DJC^{+}23, DBBL21, FRM22, GCP^{+}20, GSM21, GSG20, HDH21, HDCH21, KL22, KVL^{+}23, MWM^{+}23b, MF23, MKB21, MMGAA^{+}21, SWW^{+}21, TDC21, Wan23b, Wan23a, YG20]. mako [DSL20, Kai21, Sel20]. Malacca [POO23]. Malawi [MMK^{+}22]. Malaysia [KCCC^{+}21, MSBA22, Pra24]. Malaysian [ALM^{+}21b, Lin20]. Maldives [YFD^{+}21]. Mali [AANM^{+}22, NSZ^{+}21]. Maluku [LLLF22]. Malvinas [AM22]. Mammal [BGK^{+}22, BC20b, BC21, SMDM20]. mammals [BEM^{+}20, BGK^{+}22, KDH^{+}21, MGBS24, RTMCVMDE22]. man [Wil21a, ZCA^{+}23]. man-made [ZCA^{+}23]. Manabique [GBC21a]. manage [BMGDS^{+}20, KLS22]. Managed [LTJ^{+}21, OKS^{+}22, CdLN^{+}21, DTBP21, Dun21b, GV22, LJRHH21, MN23, RGK^{+}20, RDLA^{+}20, AGK22]. Management [AS23a, BMR23, COJE24, DB20, EHM^{+}20, EHM^{+}22, FTW^{+}23, GML^{+}24, KYMW21, MCW^{+}22a, NNM^{+}21, PCA21, Rub24, TJ21a, TK22, APV^{+}20, AKFY22, Abo23, ÁFFSC20, AHAB^{+}22, AGM^{+}22, ABVPM20, AML^{+}21, ALF^{+}22, AGSI22, Art20, BE20a, BD20, BMB^{+}24, BDM20, BEM^{+}20, BISSBdlMM22, Bea22, BSSL^{+}20, BMM^{+}20, BM21, BLM24, BDD^{+}21, BBJ^{+}22, BGMCG24, BBGG22, CPR21, CBO^{+}21, CMT^{+}23, CCK^{+}21, CCB^{+}21, CCGM21, CTH^{+}21, CLRW22, CAM^{+}23, CHB^{+}20, CDP22, hC24, CAHR^{+}24, CWKE20, CES20, Cra21, CMD^{+}20, CDB20, CSL21, DL21, DT23, DBS22a, DH22, Dbo22, DSL^{+}21, DRY22, DSL20, DAAP21, EdLP^{+}21, EAGV20, EB23, EV22, EAA^{+}21, FDGM24, FTGD^{+}20, FLB^{+}21, FS21, FMLMM23, FSF^{+}21, Fuj21, GSAF21, GLS^{+}21, GN22, GCP^{+}20, GCC^{+}21a, GAD^{+}23, GRPRVG23, GWAPRPRU22, GABA^{+}23, GTTM21, GW21, HMFH20, HDH21, HDCH21, HWU^{+}23, HLW^{+}20]. management [HYN^{+}23, HGG^{+}22b, HWT23, Har22, HdM23, HW22, HHBC20, HMG^{+}21b, HHH^{+}23, Hor23, HvPPV20b, How20, HCS21, INW20, JS23, JSW^{+}22a, JAH^{+}24, JPM^{+}23, JL21d, KWF^{+}21, KS23, Kai21, KFB^{+}20, Kar22c, Kaw22, KSB^{+}23a, KSB^{+}23b, Kim23, KKO^{+}22, KGL22, KN21, KTC^{+}22, Kul23, KJS^{+}22, LDAC20, LLT^{+}20, LMC^{+}23a, LBWS23, LDC^{+}20, LW22, LYHL23, LLCH21, LDL^{+}23, Liu23, LA23, LSGG^{+}20, LGW^{+}21, Los24, LSY20, LLL23, LCMBGP22, MUR^{+}21a, MCRPC20, ML21, MDLA21, MMB^{+}23, MVR21a, MN21, MFG^{+}23, MSBA22, MRD^{+}20, MAK22, MWM^{+}23b, MPH^{+}23, MMW^{+}23, MSGAR20, MFF^{+}23, Nal24, NAP^{+}21, NUD^{+}23, OEV23, ORSB22, OT23, OWCFDC21, OVSVPP^{+}22, Öİ24, PCD^{+}23, PKB^{+}23, PVH^{+}21, PL20b, PK23, PMAVM20, PJN24, PMR23, PLOG24, PMFR^{+}22, PFWM23, PKMT23, RMF^{+}21, RLS20, RR20, RYS^{+}21, RAFLRP21, RM21b, Ros22, RB22, RKS^{+}23]. management [SRvNN21, SK20a, SWI^{+}22, STIS20, STS^{+}23, SSZ^{+}24, SPHO23, SSOP^{+}21, SON21, SLW^{+}23, SMC^{+}21, SCCW21, SSMG21, SH23, Stø21, SLH22, SCBdA20, STKC22, SLK^{+}23, SLS^{+}21, TSH22b, TK23, TYS22, UISM^{+}23, VWFW22, VCC^{+}20, VDP^{+}22, VMC^{+}21, WJB^{+}20, WHPF20, Wat22, Wen20, WMRB22, WLCG23, WH20, WNBRSG22, WWC23, XHZ^{+}21, YWY20, YHU22, YW20, YCBW21, ZJ22, ZWJvL20, ZPX^{+}22, ZHL^{+}23, dSdLSGB22, dMLB^{+}20, CAGBM22, CTH^{+}23, De 20]. managers [Dun21b, GABA^{+}23, OCT^{+}22]. Managing [AVPAFM^{+}22, BDC23, CSLGFV20, FRM22, Lov23, MT20, RKS^{+}24, VTRS20, HIP^{+}21, HPPH22, MKAF21, SG20c, YBW20, BMF20, Cha20b]. Manchón [GBC21b]. mandatory [CMR21, HBHJ20, MPL^{+}23]. maneuver [Lou22]. manganese [HKMS21]. Mangrove [RGH^{+}22, BNF21, dFVCGdS21, MKAF21, MTC^{+}23, NHD23, QLU^{+}23, SLW^{+}23, TMS^{+}22]. mangroves [AS23b, IRK^{+}21, NEH^{+}23, RRR^{+}23]. mangroves-based [IRK^{+}21]. Manila [dZ20]. mankind [dS21]. manning [Ste23]. manta [GRPE^{+}21, SEG^{+}22, DRCdMCJ21]. mantis [BMGDS^{+}20]. manufacturing [CQ20, FZYF23, HPH22]. Map [APF22, WP23]. MaPP [SSB^{+}23, DBS^{+}22b]. Mapping [AANM^{+}22, BE20b, BS22, FFE20, GBFMT23, HP20, OTK^{+}21, Qi24, SSV^{+}23, AC21, BCJ21, CRP^{+}20, Kar22c, MFRQ23, QRK^{+}23, VGS^{+}20, WPWM^{+}22, WD20, CSCGC20]. maps [TZM^{+}22]. Mar [EHM^{+}22, Mal21, Mot22, RP23a, SCSR^{+}23]. March [Ano20r, Ano21u, Ano22t, Ano23t, Ano24h]. marginalisation [OCD^{+}23]. marginalise [SMCW22]. marginalization [Mba22]. mariculture [BHB^{+}24, DMS^{+}24, FAC^{+}23, HFD^{+}20, LHTR23, LDL^{+}23, MMW^{+}23, OWS20, PBD^{+}23, RZ21b, Ren21a, Tel23a, XHY22, YYL20, YH20, YZG23]. marina [VWJMAJ20]. Marine [SONB^{+}22, CD23, Cha20a, COSS21, CJ21, DBS^{+}22b, EAA^{+}22, EABA^{+}22, VBd^{+}21, FRM22, FMLMM23, GGP^{+}22, GSGOP^{+}21, GSR21, Gar20, GB21, GMVM23, HSGNMGC21, Isr22, JMB^{+}20, Jon21, KJE21b, KAC22, LTJ^{+}21, Mah22, MFRQ23, McC21, MBB^{+}22, MTC^{+}23, NJBM^{+}23, ONN^{+}20, PCG^{+}23, PKB^{+}21, PDB^{+}23, PKB^{+}23, QTBL23, RSFvH^{+}22, RNGRC22, SMW^{+}21, SSB^{+}23, SPCCvT21, Smi23b, SV22, Sos21a, SLP^{+}22, Tan23b, TJ21a, TGM22, VMC^{+}21, WGJ^{+}21, ZGY21, AHBAKH24, AAA^{+}24, AWBC20, ACA^{+}21, AED20, AHAB^{+}22, AELR^{+}20, AE23, AP23, ADR^{+}21, AVPAFM^{+}22, ACG21, AHM22, ABVPM20, ASM^{+}21b, AFS^{+}23, AR23a, ADB^{+}22, BMR20, BPA^{+}22, BC20a, BSM23, BDM20, BEM^{+}20, BB21, BUR^{+}22, BGK^{+}22, Ber23, BSI^{+}21, BCJ21, BC20b, BC21, BBGG22, BMB^{+}21, CMS20, CMGF20, CCS23, CMI^{+}22, CLRW22, CTCD21, CS20, CVH^{+}20, CQ20, CDHY22]. marine [CJL^{+}23, CCKK22, CJ22, DDAB^{+}23, CES20, Col20, CWR20, CNG^{+}23, CHB23, CLCV21, CMS^{+}21, DLJ^{+}20, DKW20, DFC^{+}20, DBS22a, DPB^{+}23, DHC22, DBBL21, DBH^{+}21, DLX21, DMS^{+}24, DSC^{+}23, DMG^{+}23, DG20, DL22, Dua22, DGMF^{+}20, EdLP^{+}21, Ehl21a, EV22, FTGD^{+}20, FFE^{+}24, FZYF23, FWGM24, FLB^{+}21, FR23, FWL^{+}22, FMS^{+}22, GFD21, GPA^{+}22, GNJR24, GPG^{+}21, dlFGFFR^{+}20, GSG^{+}24, GCC^{+}21a, GCMT20, GPM20, GCL21, GBC21a, GBC21b, GZW^{+}20, GABA^{+}23, Gro20, GEB^{+}21, Gug24, GW21, GTMP20, Gul23, Guo20, HWM22, HS20, HJ21, HLCM21, HM22a, HvdBH23, HPH22, HZ23, HBR21, HPPH22, HMG^{+}21b, HPC^{+}23, Hor23, How20, HLM^{+}20, HGL^{+}20, HCT^{+}21, HWY22, HBHJ20, IOHC22, JDF^{+}23, JHHG23, JLY23, JCMD20, JL21b, JF22, JPM^{+}23, JL21c, JSW22b, JL21d, KWF^{+}21, Kar22c, KFM^{+}20, KSB^{+}23a, KSB^{+}23b, KLK23, KWMM20a, KWMM20b, KvTNL20]. marine [KDM^{+}20, KGP^{+}22, KDH^{+}21, Kra22, KTC^{+}22, KHdSM^{+}22, KVL^{+}23, Kuł22, Kul23, KHLS23, LR20, LHV^{+}20, LTSL22, LGF24, LJ20, LXSX21, LW22, LJ23, LWG23, LMC^{+}23b, LYL^{+}22, LWX^{+}20, Liu20, LXW^{+}21, LC22, LHPJ23, LP21, LZRC23, LRC24, Lot22, MM21a, MUR^{+}21a, MFG^{+}22, MCRPC20, MFT^{+}22, MBDG23, MSSS22, MN21, MSO23, MSBA22, MRD^{+}20, MFM21, MWM^{+}23b, Men23, MGBS24, MMO^{+}24, Mor20, MBL24, NNM^{+}21, NW20, OYKBA22, OOSG21, OKFN24, OCT^{+}22, OVSVPP^{+}22, PKC23, PKH22, PBG23, PT23, PK23, Pha20, PS21, PCCOS24, PC20, Pri22, PDJ22, PBH^{+}23, PMFR^{+}22, Puh23, PKMT23, QZZ20, QYD20, QS21, QTPL21, QYX21, RBA23, RSAP22, RLS20, RSS^{+}20, RDT^{+}23, RP23a, RJ21, Ren21b, RXN23, RXX24, RHC20, RGK^{+}20, RTMCVMDE22, RPY^{+}24, RKS^{+}23, Rus20, SAO^{+}21, SMDM20, SGB21, SAA^{+}21, SC21, SBP24, SPHO23, SGLF^{+}23]. marine [SL20, SSDQ20, Sha20, SGK21, SLY21, SON21, SBJF^{+}21, SBDC24, SS23, SJ21, SDG^{+}21, SDG^{+}22, Sol23a, SZBA20, SPH^{+}20, SG20b, SRH23, SO20, SRS23, SWXM21, SB20, TUBA22, TLYC20, TZZ^{+}21, THP22a, THT^{+}23, TBH^{+}21, TD21, Tho22, TV21, Ton22b, TEA^{+}22, Tra23a, TJ21b, Tru23, TDP20, TBPC20, DR22, VWFW22, VHD20, VBC^{+}20, VWJMAJ20, WD23, WYH21, WXLD21, WLQD21, WSD21, WLY21, WCZS21, WW22, WHL22, WRL23, WNDZ23, WA24, War23, WD20, WMRB22, WFG^{+}22b, WBPG20, WSDB20, WCM^{+}24, WL22b, Wu22, XHZ^{+}21, XXMZ23, XXLW24, YT20, YZH22, YW20, YML^{+}21, YCL22b, YD22, Yu22, YC21b, ZK21, ZJ22, ZW21, ZLPT23, ZPX^{+}22, ZHL^{+}23, ZWXL21, ZWC^{+}23, ZLH21, Zhu22, Zhu23, ZCA^{+}23, ZBF^{+}22, dSOW^{+}21, dS22a, dO23, dV21, vTHS21, ANS^{+}20, AIG^{+}20, ÁFFSC20, AP23, AVF23b, BI21, BvBN^{+}20, BEH^{+}21, CR23, CS21]. Marine [CLW^{+}21, DBH^{+}21, Ehl21b, FGK^{+}20, GKP^{+}22, GABA^{+}22, JAK^{+}23, KMT^{+}21, LJ21, MCRPC20, MPD^{+}23, MDLA21, MPH^{+}23, Mor20, NRF21, ORSB22, QTD21, QTBL23, RA20, RP23b, RCM^{+}22, Rub24, SK20a, SSB^{+}23, SRSGMGG21a, SRSGMGG21b, TOL^{+}22, VRBC^{+}21, VBH^{+}23, WWD23, WJLSH21, ZY21, De 20, GAD^{+}23]. marine-related [QZZ20]. marine/maritime [ACG21, Ehl21a, ZK21]. Maritime [AB20, BG20a, BMRM23, CLZ23, DBH^{+}22, GJY23, HH21, HRMPCRRA22, Kar22b, Kle21, LP20, Lv20, MK24, NM23, Øst20, PDB^{+}23, RJAT23, Smi23b, SW22, TKC23, Tse21, VDP^{+}22, YC21a, ZYTT21b, Ada22, ASKI20, ACG21, BN20, BRP^{+}23, Boy21, BL23b, CPV^{+}21, CVCV^{+}22, CSQ^{+}24, CG22, CZW20, Cha21, CN23, CGP23, Day24, Dom23, Dun23, DGMF^{+}20, Ehl21a, EVPMKP22, FFE20, Fra23, FGC21, GSR21, GAD^{+}24, GCL21, GCC^{+}21b, GW21, HPS23, HS22, IGS^{+}23, JF24, JD20, Kar22a, KSS23, KF23, LKB^{+}23, LC20a, LLW22, LEVET21, MLJ^{+}22, MCL22, Mar21a, MPL^{+}23, MBORGN24, OYECdOP24, POO23, PB20, PGP22, PGA^{+}20, PK20b, PKS20, Qaf24, QF20, RLRS21, SF22, SSZ^{+}22, SPT^{+}20, SHZC21, Tod20, Wan23a, WOG20, WFG22c, XT23, ZK21, ZD20b, ZMZ23, ZT23a, ZSLY20, ZYTT21a, Zou23, vLPBK21, vdBSvdV^{+}23, ANS^{+}20, AMK^{+}21]. Maritime [BAP21, BRP^{+}23, CK23, COJZ^{+}21, DBZ24, FTW^{+}23, GBCNCL^{+}21, RBB^{+}24, SWGZ22, SGRW22, ZWG^{+}21]. markers [Los24]. Market [PGGA22, YWY20, AMIM21, CMI^{+}22, Cat20, CMNH21, Cro21, FGPPPV21, HCSBM^{+}23, MMM^{+}22, NTH^{+}21, PFM20, PLOG24, SGZH^{+}21, VKN20, YZH22, YLSC23]. Market-based [YWY20]. marketing [GM21, GPM23]. markets [GAD^{+}20, Hor23, KKA^{+}21, SGA^{+}21, SOS20, SFN^{+}21]. Markov [AS23b]. MARPOL [VSV^{+}23]. Marshall [CTIH21]. MARTEC18 [ABVPM20]. MASS [JD20]. massive [JG20, Kuł22]. master [CTC^{+}23]. match [RHC20]. matched [ZTM^{+}23]. Matching [GHW^{+}22]. materialities [Teo23]. Mathematical [vLPBK21]. matter [HdM23, JL21b, Øst20, TJZ^{+}21, WZZC21]. matters [BAP21, HRR^{+}20, LP20, OTK^{+}21, Sin21]. maturity [TPG^{+}23]. Mau [NHD23]. Mauritania [BGMCG24, TGHF^{+}22]. Mauritius [KMS20]. maw [BHdMCM^{+}21]. Maxent [AS23b]. maximisation [ASGS23]. Maximizing [PT23]. May [Mot22, MM20, Ano20s, Ano21v, Ano22u, Ano23u]. Maya [TGM22, SCD21]. Maya-Jariego [TGM22, SCD21]. Mbour [CDNS23]. MCE [Day24]. MCS [FOKM23]. meadows [KWY^{+}22]. mean [Bal22, CMR21, Kim20, QZZ20, SCF^{+}22]. means [MZ22]. measurable [KTC^{+}22]. measure [KYMW21, SPMN23]. measured [GS21]. Measurement [CDHY22]. measurements [VSV^{+}23]. Measures [BJH21, APV^{+}20, Ada22, CPR21, CHB^{+}20, CMWN^{+}23, DH22, DG23, DER^{+}21, EAA^{+}22, FTGD^{+}20, HRS^{+}24, Ma20, MBDG23, NNR^{+}21, PL24a, Ros22, SMLRB^{+}21, TAK^{+}22, WJB^{+}20, GKP^{+}22]. Measuring [ADR^{+}21, Che21a, CLW^{+}21, Gul20, POM24, Run24, SBW^{+}21, TMM^{+}23, WW21b, PWL^{+}23a, RBNE20]. meat [BBHN20, SCBB^{+}22]. mechanism [BDD^{+}21, JCMD20, Lin24, LC20b, SWHT22, TT22, XHZ^{+}21]. mechanisms [GBCNCL^{+}21, OPL20, TOL^{+}22]. Media [HPG20, PBG23, CVA22b, ES22, LDL21, Los23, RM21a, SCD21, SMW^{+}21, TMM^{+}23, WYH21, WXLD21, ZCG^{+}23, ZLWY22]. mediating [LM24, TNY23]. medicine [GLT23]. Medinipur [CDP22]. Mediterranean [CMC^{+}23, HRMPCRRA22, VCC^{+}21, ATF^{+}21, BSBD21, BSSL^{+}20, Bjø23, BMM^{+}24, CCS23, CCB^{+}21, CDD^{+}23, DL21, DFC^{+}20, DDF^{+}20, DGIS^{+}22, ES22, GM21, GCL21, GKP^{+}22, HNMSG21, Kle21, KGL22, LSGG^{+}20, MMB^{+}23, May22, NGT^{+}23, OCCSC23, OVSVPP^{+}22, SPCCvT21, Tes20, VPB^{+}24, VGK23]. medium [HYN^{+}23]. medium-scale [HYN^{+}23]. meet [GDLC20, RSFvH^{+}22, VDP^{+}22]. meeting [GPH^{+}21, YK21]. megafauna [BUR^{+}22]. Megalops [GCM^{+}23]. Megaptera [MCW^{+}22b]. Mekong [NHD23]. Melanesia [ML21]. member [Puh23]. membership [TSS^{+}24]. Memorializing [TCD^{+}20]. men [MMK^{+}22]. MENA [BEH^{+}21]. Mental [FAGS21, MEG^{+}21, PGK22, TAK^{+}22]. Merchant [Fel22, HVN21, PZPB21, XZ23]. Merluccius [BDCJ22, RQB^{+}20]. mesh [GMM^{+}20]. mesopelagic [SG20a, SG21]. Mesopelagics [FTG^{+}23]. meta [BCB20]. meta-organization [BCB20]. metabarcoding [CCS23]. meteorology [FRM22]. method [BMGDS^{+}20, BTV^{+}22, CA21, DXS^{+}21, LrS22, Okş23, PMC24, WLQD21, OTK^{+}21]. Methodological [IO23]. methodology [CGGA20, GKP^{+}22, POM24]. methods [AMLC^{+}23, GAD^{+}23]. metrics [KTC^{+}22, SS21a]. metropolitan [SdCMN^{+}22]. Mexican [CSCGC20, LEERRM^{+}21, LCMBGP22, MVSP^{+}21, PGSFRMHH24, PL20b, RTMCVMDE22, SMLRB^{+}21, VSPMS^{+}22]. Mexico [SLP^{+}22, AAFK^{+}21, BCMCM21, CLL20, GÁPSAP^{+}23, GWCD22, HDAPG^{+}24, MMSB^{+}20, MAMCHP^{+}20, SMHPMO^{+}21]. MFAD [WLRH^{+}20]. Michael [AFHM23]. Michel [Gir22]. micro [OVSVPP^{+}22]. micro-reserves [OVSVPP^{+}22]. microbeads [MLM23]. microbiological [PL23]. Micronesian [WGP^{+}21]. microplastic [BR22, DDD22, Pra24]. Mid [CSGC22]. Mid-Atlantic [CSGC22]. middle [DMNT20, TCD^{+}20]. migrant [CJ23, YL21]. migrants [ALM^{+}21b]. migrating [MCW^{+}22b]. migration [DPG^{+}22, Kle21, MSSS22]. migratory [ABM21, MWM^{+}23b]. mile [WEF^{+}21]. militarized [HGLR22]. Military [CLD^{+}24, ABI24, Dun24]. milkfish [MGS21, SGPF21]. Mind [HSSM24]. mine [ATF^{+}21]. mineral [KGW20, WRV21]. minimising [HRC^{+}24]. minimize [Dam24, SCSR^{+}22, SCSR^{+}23]. minimum [FMCD20, YU20]. Mining [Sin21, WRV21, AGM^{+}22, AR22, BHKJT23, BRM21b, CGHV21, CDC21a, CDC21b, CL21, Cra21, DMd^{+}20, Fei21, GG21, HKMS21, HSSM24, HSK^{+}23, Jae21, JACD20, Kak20, KGW20, MCR22, MM21b, PCG^{+}23, PL24b, RFPS21, Squ21, TML^{+}21, WABR^{+}22, WLCC23, Wil20, Wil21a, Wil22, ZTA23, vdGD21]. Ministerial [ALLS23]. Miramar [HM21]. Mis [BMF20]. Misalignment [MS21]. miscellaneous [BBB23]. misinform [SMW^{+}21]. mislabelling [GAD^{+}20]. mismatch [BXV20]. misperceptions [RWVH21]. misuse [HdM23]. mitigate [BKW^{+}24, BRP^{+}23, JGWM^{+}23, MZT^{+}23, TOL^{+}22, VÖB20a, VÖB20b]. Mitigating [CHC^{+}21, WS23]. mitigation [ABVPM20, BEM^{+}20, CLD^{+}24, DSB21, DBB^{+}22, GCB^{+}23, GAD^{+}23, HSGS^{+}23, MDLA21, MSGAR20, NHD23, SNP21, SAB^{+}20, WLW^{+}22, ZWJvL20, ZCSK23, Rub24]. mixed [CPR21, COJE24, GSG20, MH20a, NGP^{+}21, NSK20]. Miyakojima [Tru23]. MLS [YU20]. mm [GMM^{+}20]. mobile [FARBV^{+}24]. Mobilising [HPG20]. mobility [ATF^{+}21]. Mobilizing [WGP^{+}21]. model [BCB20, DLYM22, HLW^{+}23, HRR^{+}20, JWX^{+}20, Kaw22, KDH^{+}21, MDG^{+}23, MLM23, Obu20, PHF20, PGSS22, PMT^{+}20, SGK21, VPRR^{+}21, WHPF20, WCZS21, Zha21, ZHL^{+}23, ZYTT21b, GMG^{+}23]. Modeling [OR21, SA21, Cat20, SGFM^{+}21, WGR^{+}22]. Modelling [ZCG^{+}23, BVT^{+}21, CAM^{+}23, MDLA21, PB20, THT^{+}23, vLPBK21]. models [AS23b, BDR20, DKW20, FAGS21, Gal21, GCL21, KPN20, MEG^{+}21, PKH22, PBT^{+}20, Pra20, WL22b]. moderating [NRYG20]. modern [Ege23, WKBS20]. modernist [Alf20]. modes [VHSV20, KA21]. molidae [PNS^{+}23]. Molidea [PNS^{+}23]. mollusc [AOO^{+}23b, PL23]. momentum [LXSX21]. Mondego [SCBdA20]. money [MJMAA21, SCD21, TGM22]. monitor [GMG^{+}23, TLT^{+}22]. Monitoring [BLH^{+}20, LWSS21, NJBM^{+}23, PKB^{+}21, Rub24, SIW^{+}21, SG20c, WCF^{+}21, WCM^{+}24, XWH^{+}22, vdBSvdV^{+}23, AIG^{+}20, BCJ21, BDC^{+}21, CCS23, CWR20, EAA^{+}21, EHM^{+}20, EHM^{+}22, GCC^{+}21a, GDLC20, GAD^{+}23, KMS20, KDM^{+}20, MGBS24, OWS20, PSTW21, SCA^{+}20, WPWM^{+}22, WD20, ZLD^{+}21]. Mont [Gir22]. Mont-Saint-Michel [Gir22]. Monterrico [GBC21b]. Monument [Cha20a]. Moore [Ngu23]. Moorea [SO20]. moored [WLRH^{+}20]. moratorium [DLX21]. morio [HDAPG^{+}24]. Mortality [HdM23, OKS^{+}22, SLMG^{+}21]. mosaic [THP^{+}22b]. most [Men23, OVSVPP^{+}22, VPRR^{+}21]. mother [SMKS22]. mother-of-pearl [SMKS22]. motivate [WLQD21]. motivation [EVPMKP22]. Motivations [GRPE^{+}21, MFM21]. MoU [XWL^{+}20]. Mount [LA23]. move [JPM^{+}23]. movement [Lot22]. movements [GCV^{+}20]. Moving [RDM^{+}20, DUYEA21, VBH^{+}23, Cla20]. Mozambique [RDM^{+}20]. MPA [VMC^{+}21, ABM21, BBK^{+}21b, CLS^{+}23, dOJCSLdSB21, NBB^{+}23, PSTW21, SRR^{+}21, WD20]. MPAG [JL21c]. MPAs [CWB^{+}20, EAA^{+}22, HNMSG21, Kat20, Mur21b, RCH22, TCZ21, VvSvH^{+}22, dPJ21]. MSC [BI21, JAK^{+}23, SK20a, DQJ23, GML^{+}24, JC23, KMJ24, LLM^{+}23, PCK22, SPG20, WS21]. MSC-certified [SPG20]. MSE [COJE24]. MSP [AMK^{+}21, BMM^{+}24, CMOJ^{+}24, HvT22, QRK^{+}23]. MSY [TM23a]. MSY-based [TM23a]. Much [OSV21, DMNT20, SMHPMO^{+}21]. mud [DMS^{+}24, RKR20, HCT^{+}21]. mullet [MKM^{+}20]. Multi [LRD^{+}22, NUD^{+}23, ZLPT23, ASV^{+}21, BCB20, COJE24, CPBO20, CvHF^{+}24, CDB20, hBDDF22, EAGV20, HvdBH23, HW22, HBR^{+}23, KLSG21, Mba22, RBB^{+}24, SSZ^{+}22, TMS^{+}22, VTRS20]. multi-criteria [ASV^{+}21, CPBO20, EAGV20, HW22, TMS^{+}22]. Multi-dimensional [LRD^{+}22]. multi-level [Mba22]. multi-levelled [HvdBH23]. multi-risk [HBR^{+}23]. multi-scalar [RBB^{+}24]. Multi-scale [ZLPT23, VTRS20]. multi-specific [COJE24]. multi-stakeholder [BCB20, CDB20]. multi-usage [hBDDF22]. multi-use [CvHF^{+}24, SSZ^{+}22]. multi-vessel [KLSG21]. Multi-year [NUD^{+}23]. multiagent [Qi22]. multidimensional [HIP^{+}21, RMBR22]. multifaceted [ACG21]. Multilateral [POO23]. multilayered [ORSB22]. multilevel [SOLN21]. multilingual [Pau21]. multinational [CRP^{+}20]. Multiple [DČFW23, DB20, PFWM23, BDM21, GAD^{+}24, GAD^{+}20, HvPPV20b, MM21a, MGF20, OCT^{+}22, PCD^{+}23, SHP^{+}21, WGR^{+}22, GBC21b, Has20]. Multiple-goal [PFWM23]. multiple-use [MM21a]. multispecies [CTH^{+}23, GMPT^{+}20, GVH^{+}23]. Multivariate [MDG^{+}23]. municipality [IAB^{+}20]. Murcia [ASV^{+}21]. murky [Wil22]. mussel [SPCCvT21]. mussels [Gir22]. mutual [JF20]. mutualism [CMI^{+}22]. Myanmar [HBJ21]. Myeik [HBJ21].

Nam [Han21]. Namibia [FGK^{+}20]. naming [CMNH21]. Napoleon [AGSI22]. Nariño [PCCOS24]. narrative [CMOB^{+}20, MCL22, MSO23, Obu20]. narratives [BDMW22, LLT^{+}20, TTA^{+}20]. narrow [GCV^{+}20]. Nash [Han20]. Natalie [Tan23b]. Nation [BMB^{+}21, HHH^{+}23]. National [Cha20a, COSS21, CJ21, GASLCP20, Has23, HNMSG21, KJE21a, MTN23, Teo23, TM23b, TCD^{+}20, YCL22b, BLL22, CWR20, DMNT20, Fox22, FMS^{+}22, GNJR24, Gar20, HJ21, HM22a, Hon20, HGL^{+}20, KAN^{+}23, LrS22, Mar21a, MDJ^{+}22, MH24, MCR22, ONA21, OKR^{+}20, PBD^{+}23, Pri22, QZZX20, Qi22, Qi24, SL23, TD21, TMDN23, VHD20, VBC^{+}20, Wil20, WGJ^{+}21, YG20, YML^{+}21, ZJ22, ZD20a, ZT21, dKMK21, ASGS23, QTD21, SLP^{+}22]. nationality [HVN21]. nationally [Moh23]. nations [BDDE20, TFH^{+}23, Gar20, KS22, MH24]. native [GTTM21]. natural [AML^{+}21, BDDE20, BEH^{+}21, EMN^{+}24, OGS24, RXX24, RMBR22, SMLRB^{+}21, SOP^{+}22, JL21d]. Nature [Sta23a, DRY22, PCA^{+}23, RSFvH^{+}22, SVKP21, SRH^{+}22, HDHM^{+}20]. nature-based [RSFvH^{+}22]. nautical [GBFMT23]. naval [Dun23, Dun24]. navigate [WCC^{+}20]. Navigating [DBZ24, GAD^{+}24, MRD^{+}20, SLS^{+}21, SS21b]. Navigation [Ngu23]. navigator [KP24]. Nazca [CMWN^{+}23, DBH^{+}22, WvdMG^{+}21]. near [WPWM^{+}22]. nearly [LJ20]. nearshore [DLJ^{+}20, SCCW21]. necessary [BTE21, GR20a, VvSvH^{+}22]. necessity [Dau23]. need [AGK22, BBB23, BB20, Cho22, FDF^{+}21, GCG^{+}21, MH24, Öİ24, Sin21]. needed [MJMAA22, McC20, SOO20, SCD21, Tra23b, VCC^{+}20]. needs [ASKI20, GPG^{+}21, HGG22c, MAMCHP^{+}20, MKSG20, UISM^{+}23, VDP^{+}22, WM22b]. negative [HACCM22, SCSR^{+}22, SCSR^{+}23]. negligible [QF20]. Negotiating [Alf20]. Negotiations [TMDN23, BLL22, De 22, Dua22, HW22, HHD20, MTN23, SCGC20]. neighbours [PK20b]. Nemo [HM22c]. Nengye [Sos21a]. Neoliberal [Ree20]. Neostructural [BBGHR23]. Nephrops [BMGDS^{+}20]. Nerlovian [OAM20]. nesting [LTRA20]. net [SKL^{+}20, WLW^{+}22]. Netherlands [dV21]. Netnography [SCD21, TGM22, MJMAA21]. nets [DAHS^{+}23]. Network [ANR^{+}22, MFG^{+}22, SWH^{+}23, WJLSH21, CCS23, CCCC21, FFD^{+}23, GPM23, GMRSP23, GGL22, KABT21, LZRC23, MF20, ORSB22, SLS^{+}20, VWFW22, VMD^{+}24, WW21a, WGR^{+}22, WMRB22, XHZ^{+}21, YWYL22, ZLGX20, ZWG^{+}21, ZHL^{+}23]. networks [AOO^{+}23a, ABM21, ACV^{+}21, AHB^{+}23, CS20, DTBP21, HRC^{+}24, MIMS^{+}20, PN20, RKS^{+}24, WMS21, KA21]. neutral [GCB^{+}23]. neutrality [HZ23]. never [Gal22]. new-type [Zha22]. Newfoundland [DC20, EVB21, Fro21, NGRM^{+}22]. news [CVA22b, LDL21, RM21a, TMM^{+}23]. Next [YK21, BÖ24, Set23]. nexus [KSB^{+}23a]. NFDS [ASGS23]. Nha [KJE21b]. Nicaragua [LZRC23]. Nicole [Lv20]. Niger [NSZ^{+}21, UOO20]. Nigeria [BOOO23, KTI^{+}24, MAA^{+}24, NW20, UOO20]. nighttime [TLT^{+}22]. Níjar [HSGNMGC21]. Ningaloo [Jon21]. Niño [HGLR22]. NIS [Kul23]. NL [MF20]. No [GVL22, BBB^{+}20, Bra22, Che21b, GV22, HLCM21, LC20a, QTPL21, QTBL23, WBPG20]. no-take [BBB^{+}20, QTPL21, QTBL23, WBPG20]. noa [MAR^{+}20b]. noa/we [MAR^{+}20b]. nodes [HRC^{+}24]. nodule [HKMS21, JACD20, WABR^{+}22]. noise [AdSMS21, BRP^{+}23, CLD^{+}24, KDH^{+}21, MG22, VÖB20a, VÖB20b, VRBC^{+}21]. Non [Cat20, SPN^{+}21, BPA^{+}22, CMS20, CC23, EMH^{+}23, HvT22, HWY22, KL22, Lin23, MJ23, MEG^{+}21, OAM21, Pet22, PCCOS24, QF20, RROP23, SCSR^{+}22, SCSR^{+}23, SON21, WMRB22, WCM^{+}24, ZT21]. non-binding [HvT22]. non-climate [PCCOS24]. non-competitive [ZT21]. non-compliance [Lin23, MJ23]. non-discretionary [OAM21]. non-entangling [EMH^{+}23]. non-governmental [CMS20, HWY22, KL22, Pet22, WMRB22]. Non-indigenous [SPN^{+}21, WCM^{+}24]. non-lethal [SON21]. non-linear [MEG^{+}21]. Non-market [Cat20]. non-negligible [QF20]. non-state [BPA^{+}22, CC23]. non-target [SCSR^{+}22, SCSR^{+}23]. nondegradable [VJL^{+}21]. nonindigenous [Kul23]. Nonlinear [Sha20]. norm [AH21]. Normative [Fro21, Gra21]. norms [MJ23]. North [KABT21, NEH^{+}23, SZBA20, vHSSK20, DXS^{+}21, HNMSG21, Mag23, MFRQ23, ALML23, ÁFFSC20, BG20b, BvdBvdFK20, BDCJ22, BM21, BB20, BLY^{+}24, BGMCG24, DBS^{+}22b, DMS^{+}24, FSKJ23, FFE^{+}24, GLADGGVL23, Gra21, HS22, HvdBH23, Kai21, MCRR23, MBB^{+}22, MM20, PS21, PKMT23, RHL20, SOT^{+}22, SNS23, SV22, SR22, TCS^{+}20, TYS22, VBH^{+}23, VGK23, YK20, dV21]. North-East [ÁFFSC20, FFE^{+}24, VBH^{+}23]. North-Eastern [DMS^{+}24]. north-west [DXS^{+}21]. North-western [NEH^{+}23, HNMSG21, MFRQ23, GLADGGVL23]. Northeast [CCGM21, CDRS22, dFVCGdS21, JHL21, MKM23, MDLA21, WD23, Bjø23, BUM^{+}24, CPV^{+}21, EB24, FTSG21, HC21, MKSG20, RHST21]. Northeastern [WW21b]. Northern [PCA21, CBO^{+}21, CC21, CDHY22, EVPMKP22, GPTASM23, GRPRVG23, GNMGWM21, HDAPG^{+}24, JGWM^{+}23, dSJNS20, OKH^{+}24, WLW^{+}22, dPJ21, BISSBdlMM22, EMN^{+}24, EMS^{+}23, Gun21, HLLK22, KvTNL20, ONN^{+}20, RZ21a, Tod22a, WDG^{+}20, Wan23c]. Northwest [dlFGFFR^{+}20, DC20, MSGAR20, HLW^{+}23, SBW^{+}21, ZNA^{+}23]. Northwestern [WCF^{+}21, BMR20]. norvegicus [BMGDS^{+}20]. Norway [ASGZ22, DNR21, EMS^{+}23, GBLBV20, GVL22, HJLN24, KKO^{+}22, ONN^{+}20, SDI^{+}20, SLL^{+}24, SGZH^{+}21, SGH22, SGL20, SG20a, SH23, TSH22b]. Norwegian [Åm21, AOO^{+}23a, BE20a, BE20b, BDR20, BDDE20, Fla21, GVR20, Her22, RAB22, SZ20, STG^{+}20, SAO^{+}24, TFF22, YHU22]. note [AEJ24, Tes20, ZT23b]. Notes [Smi22]. nothing [OSV21]. notification [For22]. notification-driven [For22]. notion [Urb22]. Nova [CA20, TFF22]. novaeangliae [MCW^{+}22b]. novel [CAC23, CLCV21, DQJ23, IKB21, LJRHH21, PVH^{+}21]. November [Ano20t, Ano21w, Ano22v, Ano23v]. nuclear [LWZ23, WL22a, XTY22]. number [KPN20]. numbers [LLM^{+}23, WD23]. Nunavut [KDH^{+}21, BOW^{+}22]. nursery [MCW^{+}22b]. Nusa [RYS^{+}21, YC21b]. nutrient [RMF^{+}21]. nutrition [TCC^{+}22]. nuts [SG20a]. Nuxalk [BMM^{+}20]. NW [Sol23a, VCC^{+}21].

object [NM23, SHPW21, dZ20]. objectives [GDLC20, HGH^{+}20, HvPPV20b, LBL21, Mur21b, OGS24, PCD^{+}23, PFWM23, YW20, TML^{+}21]. objects [PKH22]. obligation [FGPPGG23, GMPT^{+}20, Guo20, KS22, OSV21, PMAVM20, PV23, SOLN21]. obligations [Har22, KL22]. observation [MAMCHP^{+}20, TZM^{+}22, VHSV20]. Observations [BMR20]. observer [EHM^{+}20, EHM^{+}22]. observers [FSKJ23, Gar24]. obstacles [AMLC^{+}23, VJO23]. obtained [KMG^{+}20]. obvious [NBB^{+}23]. occupation [HVN21]. occupational [PK20a, SHZ^{+}21, Sha22]. occur [CMS^{+}21]. Occurrence [Kul23, PY20]. Ocean [GMG^{+}23, HDHM^{+}20, HGG22c, Kar23, MZ24, MLAS23, Pol23, SBS21, Smi20, SWGZ22, Teo23, TSWC23, VYB20, YK21, vSHHT20, AAA^{+}24, AAA^{+}21, AFHHD22, AHN^{+}23, BLH^{+}20, BH23b, BBWC21, BAF^{+}23, BCB20, BDM21, CC23, Cha24, CMWZ24, CSH23, CNG^{+}23, COJS^{+}22, EFFP23, FDGM24, GPA^{+}22, GGP^{+}22, GBC^{+}20, HJPL23, HLO^{+}21, JV20, KvTNL20, KN21, LLT^{+}20, LZL23, MLE^{+}24, MCL22, MSP23, MFG^{+}23, MGDJ22, MMNC22, Moh23, MM21b, NMCL24, OKNB22, OBNK24, PMN^{+}22, PNS^{+}23, PMC24, Pyć23, SdACdSLdAS24, Shi20b, SWHT22, SMC^{+}21, SDG^{+}21, SDG^{+}22, SCA^{+}22, TZM^{+}22, TSA^{+}22, Tho22, VHS^{+}20, WPWM^{+}22, XT23, YG20, YWYL22, YC21a, ZJ22, ZCZ20, ZT21, ZZD^{+}23, ZTA23, Alf20, BUM^{+}24, CTH^{+}21, DSL20, GSGOP^{+}21, Gar20, GB21, HLW^{+}23, KWF^{+}21, Kai21, KY20, LCMBGP22, MFT^{+}22, MDG^{+}23, McC24, MAG^{+}22, MSGAR20, NOL^{+}23]. Ocean [NT20, Pri22, SGWL21, Sta23a, Tra23b, TMTC20, TYS22, VYB20, WCF^{+}21, Wen20, Smi23b]. ocean-based [HLO^{+}21, Moh23]. oceangoing [ASGZ22]. oceanic [LP20]. Oceanographic [TK23]. oceanography [Yu22]. Oceans [BCN23, BR22, BDM21, CSLGFV20, CWKW20, DKKS21, KFS^{+}22, RSFvH^{+}22, STF21, Smi22]. October [Ano20u, Ano21x, Ano22w, Ano23w]. octopus [CSCGC20, DHR23, LMC^{+}23a, SPG20]. octopuses [BBB23]. Ocypode [YSEK22]. odd [Wil21a]. Odisha [KHMM20]. OECD [LCC20]. OECMs [EAA^{+}22, MBDG23]. Off [TMDN23, AML^{+}21, BH23a, CHEC21, Dun24, HDAPG^{+}24, SOP^{+}22, Tod20, WWG^{+}23]. off-the-shelf [Dun24]. offer [HSH^{+}21]. officers [KZR22, Tan22]. official [PKB^{+}21]. offs [BCN23, GASLCP20, MB22, PKMT23, ZGY21]. offset [KWY^{+}22]. offsetting [MRKM23]. Offshore [Cha20b, Laj21, MCRR23, BvdBvdFK20, BST23, BLY^{+}24, CSLGFV20, CLL22a, CCW21, DLJ^{+}20, DWC^{+}20, DBHT22, FBC^{+}23, HRR^{+}20, HBJ21, JFB^{+}24, JF22, KNY20, KABT21, LJ21, MBORGN24, OC20, ONA21, OKH^{+}24, PRM^{+}20, PKMT23, SBT21, SZBA20, SVKP21, TYS24, WNMS^{+}24, YK20, vRB24]. often [MS21]. OHS [Sha22]. Oil [Cha20b, JHHG23, MFT^{+}22, dOSTB^{+}20, SRvNN21, SBT^{+}23, SGFM^{+}21, WDG^{+}20, PKH22]. Okinawa [Tru23]. old [VBd^{+}21, Gal22, ASM^{+}21b]. Olympics [HWMO20]. Oman [ASGS23, MDLA21]. on-board [ASKI20, MRD^{+}20]. on-land [PMAVM20]. onboard [DG23, PGK22, VAMR^{+}20, dOPM^{+}20]. One [Mag23, PBG^{+}22, SCA^{+}22, dO23, Liu21, OVSVPP^{+}22, PT23, APF22, Her22]. One-Size-Fits-All [Mag23]. ongoing [SEG^{+}22]. online [MGS21, RPY^{+}24, STIS20]. only [WFR23]. onshore [KSS23]. ontologies [Teo23]. Open [WOG20, AFHHD22, AR22, Art20, OA20, SOS20]. open-air [SOS20]. open-ocean [AFHHD22]. Opening [GEB^{+}21, BSY^{+}20, CC22, CNB^{+}23, GGL22, Tod20]. operate [LMMT23]. Operating [BBBS22, BCJ21, DSL20, HBJ21, KZR22]. operation [DG24a, HKMS21, PWL^{+}23b, SC22]. Operational [KABT21, Yu22, BTV^{+}22, DG23, Guo20, TBV^{+}23, WD20]. Operationalisation [GPG^{+}21]. operationalising [GG21]. operationalization [MZ22]. operations [MBL24, Stø21]. operators [BHEE22, EP20, WW21a]. opinion [GLT23, KSB^{+}23a, WFG22a]. opinions [CPR21, DDD22, KHLS23]. Opportunities [Hsi20, LGW^{+}21, OYECdOP24, SMF^{+}23, SSZ^{+}24, AMLC^{+}23, BI21, SONB^{+}22, BAD21, CWB^{+}20, CWR20, DS22b, JF24, KWF^{+}21, KFS^{+}22, OGS24, QWBY21, SPFA^{+}20, SLB21, SCGC20, SF22, SGL^{+}21, TBPC20, WvdMG^{+}21, WGR^{+}22, WRV21]. opportunity [Dbo22, KFP20, KMM21, MZ24, SSZ^{+}22, Sta23b]. Optimal [EHLT21, XTY22, PHF20, SCSR^{+}22, SCSR^{+}23, SKE22]. optimise [CRP^{+}20]. Optimistic [SBT21]. Optimization [GSB20, BMC20, Mot20, Mot22, TBV^{+}23]. Options [RJAT23, Tod22a, BKW^{+}24, DBHT22, FGPPPV21, IBC^{+}20, KGW20, Shi20a]. orange [Squ20]. orange-spotted [Squ20]. orca [Dam24]. Orchid [KWG23]. Orcinus [Dam24]. order [BN20, SHZC21]. ordering [TK23]. Oregon [HCWC21, SCCW21]. Organic [SOK^{+}22]. organisation [Ros22]. organisational [SKL23]. Organisations [GML^{+}24, GLAVL21, HWY22, KL22]. organised [BE20c]. Organization [EHM^{+}22, BCB20, GMRSP23, dOJCSLdSB21, SLB21, SCBdA20, CMSN^{+}22, EHM^{+}20, Gug22, RJAT23]. Organizational [BCB20, HSdT21, KLSG21, RKS^{+}24, ZYTT21a, SOS^{+}21b]. organizations [CMS20, HMFH20, LSY20, Pet22, WMRB22]. orientation [LDAC20, VKN20]. oriented [MLVT21, PCCOS24]. ornamental [CMI^{+}22, PDJ22]. Oslo [Qaf24]. OSPAR [TCZ21, RCH22]. Osthagen [Smi23b]. Otaria [PBH^{+}23]. other [EAA^{+}22, GCL21, MBDG23, SGA^{+}21, SZBA20, YG20, dS21]. otherwise [GRS^{+}23, Lov23]. otter [NSK20]. our [JV20, LMSLMQ21, RSFvH^{+}22, Cha24]. outbreak [NK21]. outbreaks [HCW^{+}23]. outcome [JDF^{+}23]. outcomes [BDD^{+}21, GABA^{+}22, GABA^{+}23, Has23, MFM21, VPRR^{+}21]. outdated [CMNH21]. outer [dS21, dS22a]. outermost [CVCV^{+}22]. Outlook [SK20b, FGC21]. Output [dlFGFFR^{+}20, BG20a, BSSL^{+}20, SRSGMGG21b, TM23a, WW22, ZCCQ22a, ZT21, ZZD^{+}23]. outstanding [Sin21]. Ouvea [LDAC20]. Over-fishing [dMLB^{+}20]. overall [AGK22]. Overcoming [JAH^{+}24]. overfishing [FTSG21]. overlap [SSMG21]. overlapping [SWH^{+}23]. Overlaps [DS22b]. overlooked [Gar24]. Overseas [DHC22]. Overview [LZRC23, CDRS22, CL21, DG24a, WA24, NSZ^{+}21]. own [FGK^{+}20, Ste23]. owner [EP20, PYVF24]. owner-operators [EP20]. ownership [EP20, HvPPV20a]. Oxford [Smi22]. oxyrinchus [DSL20, Kai21]. oyster [BABC20, LCL20, PZZ^{+}21, SN23, WM21].

P [BRR^{+}22]. P. [BRR^{+}22]. Pêche [Ken20]. Pacific [Abo23, AFHM23, CDG^{+}21, DBS^{+}22b, HLW^{+}23, Kai21, LCMBGP22, MBB^{+}22, SOT^{+}22, TYS22, War23, Abo23, AL20, BBC24, BBS^{+}21, Ber23, BDC^{+}21, CMWN^{+}23, CLD^{+}24, DPG^{+}22, FSKJ23, FTM^{+}22, FOKM23, GBC21b, GC22, HWT23, JHHG23, JF24, Kak20, KFB^{+}20, Lou22, MZ24, NNR^{+}21, OWCFDC21, OAAMS24, RROP23, RTMCVMDE22, SMYH22, SMKS22, SGRW22, SW22, STF21, WKH22, YS21]. PADDD [ACA^{+}21]. Pagellus [SFGEPC22]. pair [LXW^{+}21]. Pairing [MLJ^{+}22]. Palau [DLJ^{+}20]. Palawan [MFG^{+}22]. Palestine [Qaf24]. Palgrave [AFHM23]. Palos [HSGNMGC21]. Palos-Islas [HSGNMGC21]. Pan [DB20]. panacea [OOSG21]. Panama [GHK20]. pandemic [AEL^{+}23, ASGZ22, BSI^{+}21, CDNS23, CRCP^{+}23, DME23, Ero22, FTM^{+}22, GPTASM23, HDH21, HCW^{+}23, IKB21, Kiz24, LKE^{+}22, MÇ23, MK24, PGK22, PAK22, SVV24, Set23, SSB^{+}22, SRH^{+}22, Tan22, TAK^{+}22, VMM21, VDP^{+}22, WHL22, ZTZ^{+}21, ZTW23, ZTM^{+}23, GMCMJA^{+}21]. Pandora [Tod20]. panel [PNS^{+}23, SGK21, XSCL21, Sha20]. panels [SKL^{+}20]. Pangkajene [HF20]. Papahānaumokuākea [Cha20a]. Paper [RP23a, RP23b]. Papua [HFD^{+}20, SWI^{+}22, SMKS22]. Papuan [SEG^{+}22]. paradigm [CCB^{+}21, MSZ^{+}23, SLW^{+}23]. paradise [Tea20]. paradox [DPB^{+}23]. Paralympics [HWMO20]. Parapenaeus [BDCJ22]. Paris [Moh23]. Park [COSS21, CJ21, KJE21a, QTBL23, RP23a, RP23b, Tan23a, CLW^{+}21, GASLCP20, HNMSG21, MCRPC20, YCL22b]. Parks [Jon21, ALML23, HWM22]. Parnaíba [dFVCGdS21]. part [RB22]. Partelow [Sta23a]. participants [HACCM22, WNMS^{+}24]. participated [SCBdA20]. participating [Cha22]. participation [BDC23, CR23, Dua22, GTMP20, HvdBH23, Kak20, MMC^{+}22, MFM21, MCR22, MMGAA^{+}21, NEF23, OBNK24, ORSB22, RYS^{+}21, RPY^{+}24, SAO^{+}21, SBD^{+}23, SCA^{+}20, SN23, Szy20, WXLD21, WLQD21]. Participatory [AC21, DHR23, PMC24, ASV^{+}21, BJGG^{+}24, GSGOP^{+}21, Kar22c, MGS21, MM23, PMT^{+}20, RA20, dSOW^{+}21]. Particularly [Cho22, RJAT23]. particulate [WZZC21]. parties [Bov22, FY24]. Partnership [DBS^{+}22b, MBB^{+}22, SSB^{+}23, HC22, Liu22, DBH^{+}21]. partnerships [LTJ^{+}23, LGW^{+}21, WJLSH21]. party [Kim21]. passage [GCV^{+}20, CDRS22, FFE^{+}24, JHL21, TCD^{+}20, ZNA^{+}23]. Passive [MGBS24, KDM^{+}20]. Past [AS23b, MCRR23, DL21, Liu21]. Patagonia [AHB^{+}23]. Patagonian [CHC^{+}21]. patchwork [Wil20]. Patella [OVSVPP^{+}22]. Patellidae [OVSVPP^{+}22]. Path [TUBA22, Ber22, PMM^{+}21]. paths [DČFW23, GY23]. pathway [ZWXL21]. Pathways [DTBP21, ADB^{+}22, CMWN^{+}23, DP23, HSGS^{+}23, LWZ23, NNR^{+}21]. Patrol [Laj21]. Patron [RMU^{+}22]. Patron-client [RMU^{+}22]. pattern [TC23]. Patterns [BSK^{+}20, BMK^{+}22, CHEC21, CLS^{+}23, FLC^{+}23, GS21, LCC20, MKM^{+}20]. Paul [Smi20]. Paulo [PMC24]. pay [BR22, CLL20, CLC21, CKY22, LC20a, XLLS21, XXLW24]. Payment [RRR^{+}23]. payments [JZW21]. pays [Zhu23]. peace [hC24, CSHO^{+}20]. pearl [SMKS22, WHPF20]. peer [BSK^{+}20, GR20a, PGK22]. peer-review [GR20a]. pelagic [AGN^{+}20, AONF21, BCBS22, CC21, COJE24, CDC21a, hBDDF22, MPD^{+}23, OKFN24, PFWM23, SBW^{+}21, TSA^{+}22]. pelagics [ABM21, BGMCG24, MGS21, TGHF^{+}22]. Penghu [LTJ^{+}23]. penguin [TD21]. Penida [YC21b]. Peninsula [HDAPG^{+}24, hC24, TDC21]. people [AHB^{+}23, BEM^{+}20, BS22, DBBL21, EVB21, TKC23]. Peoples [VHS^{+}20]. Perceived [LHTR23, CD24, NGP^{+}21, SBDC24, SKS21, TAK^{+}22, TCS^{+}20, ZWJvL20]. Perception [CDP22, FTGD^{+}20, NGA23, OCT^{+}22, BMMF22, CCW21, HDAPG^{+}24, Hon21, LDL21, RPK23, SVP^{+}23, WFR23, dOPM^{+}20]. Perceptions [GLADGGVL23, McC24, MTC^{+}23, Tru23, WNBRSG22, AGK22, AMLC^{+}23, CGGA20, CLB^{+}20, EVPMKP22, VBd^{+}21, GTA^{+}20, GPM20, GRPRVG23, GCM^{+}23, HPC^{+}23, HACCM22, HHEJB20, HvPPV20a, KWG23, KK22, KMG^{+}20, KNY22, KWMM20a, OMC^{+}22, PVH^{+}21, PBS^{+}23, PLOG24, RP23a, RP23b, RHC20, RKS^{+}24, SMLRB^{+}21, SDI^{+}20, SON21, SCH^{+}24, SPJB22, SAB^{+}20, SLS^{+}21, TDC21, VWJMAJ20, WLCG23]. perceptual [QRK^{+}23]. perfect [Wil20]. Performance [BDDE20, DAA21, DDGdS^{+}21, EAAG21, AE21, AACdlT23, AAW^{+}23, AGA^{+}21, BJGG^{+}24, CDG^{+}21, DMS^{+}24, DMG^{+}23, EP20, GLS^{+}21, GABA^{+}22, HMFH20, HMY^{+}23, HvPPV20a, Kar22a, KB20, KWMM20a, KWMM20b, LFPBD^{+}20, MVR21a, NSZ^{+}21, OMC^{+}22, PKR^{+}22, QTD21, SGZH^{+}21, SZBA20, WD23, ZYTT21a, YW20]. performances [ZZW^{+}21]. period [HSHS24]. peripheral [HM21]. permanent [vHSSK20]. permits [BBJ^{+}22]. perpetuated [Mba22]. Persian [HGG^{+}22a, SONB^{+}22]. persist [NAP^{+}21]. perspective [ADR^{+}21, BM21, BMM^{+}24, SONB^{+}22, CWKW20, CKY20, CLCV21, Doe21, EVB21, FRM22, FWL^{+}22, Guo20, HWT23, Hon20, Kak20, KKDK20, KLS22, LLW22, LSH21, MSO23, MB21, OEV23, OC20, OAAMS24, PYVF24, Qi22, QYD20, QUK21, RR20, Ren21a, RTMCVMDE22, SVHV20, Set23, Shi23a, SCGC20, TTA^{+}20, TD21, TK23, WDG^{+}20, War23, WD20, ZTZ^{+}21, ZW21, ZWG^{+}21, ZYTT21a]. Perspectives [BL23a, DAAP21, Pra24, BCB20, BLS21, CWMM^{+}21, FMLMM23, GNJR24, Has23, HC21, Kas21, MGBS24, MHEW20, PZPB21, PMN^{+}22, QWBY21, STS^{+}23, Sol23b, SKW^{+}22, VJO23, WFG22c, HDHM^{+}20]. pertaining [DDD22]. Peru [GWAPRPRU22, GNMGWM21, JGWM^{+}23, JTR^{+}20, dPJ21]. Peruvian [CC21, GPTASM23, GWAPT21, GRPE^{+}21]. perverse [hBDF22]. PES [RRR^{+}23, SGA^{+}21]. PESTLE [THS^{+}24]. PGIS [dSOW^{+}21]. phase [HCS21]. philanthropy [GEB^{+}21]. Philippine [SLS^{+}20, SLS^{+}21]. Philippines [AR23a, BMB^{+}24, HJPL23, MLMD21, MMC^{+}22, MMK^{+}23, MNGM24, MH20b, MIH^{+}24, OT23, PCA21, PCSS^{+}21, WKO^{+}22, dZ20]. Phu [KK22]. Physiological [BMGDS^{+}20]. phytoplankton [SCA^{+}20]. Pier [QCM^{+}20]. Pierucci [LLM^{+}23]. pilot [GRS^{+}23, GDLC20, ZLD^{+}21]. pilots [ZT23a]. Pingtan [SWW^{+}21]. pinnipeds [WCT^{+}20]. pioneering [Gar22]. pipelines [KS22]. piracy [FLC^{+}23, GJY23, POO23]. place [CD24, SCF^{+}22, SWW^{+}21, VJL^{+}21]. Placentia [MF20]. plaice [NSK20]. Plan [DBH^{+}21, DBS^{+}22b, MBB^{+}22, SSB^{+}23, GAD^{+}24, PKB^{+}23, DL21]. planet [HJ21]. Planned [DLYM22, LYQM21, MLM23, XXLW24]. Planning [DKW20, LXH^{+}22, PDB^{+}23, SV22, TBH^{+}21, WJLSH21, AHBAKH24, ACG21, AHM22, AML^{+}21, BN20, BMMF22, Bir20, BMB^{+}21, CPV^{+}21, CVCV^{+}22, CSQ^{+}24, CTCD21, CHB23, DBH^{+}21, DBS^{+}22b, DGMF^{+}20, Ehl21a, EVPMKP22, FGC21, GPA^{+}22, GGP^{+}22, GPG^{+}21, GSR21, GOH^{+}20, GCC^{+}21b, HLW^{+}23, HLLK22, JDF^{+}23, Jae21, JHHG23, JMB^{+}20, KvTNL20, KVL^{+}23, LKB^{+}23, LJ20, LJ23, LEVET21, Mah22, MBORGN24, OEV23, OWS20, Öİ24, PBS^{+}23, PGA^{+}20, PS21, PMC24, PKS20, RSAP22, RLS20, RDT^{+}23, RYS^{+}21, RHO24, SGB21, SAA^{+}21, SBP24, SSOP^{+}21, SZBA20, SSZ^{+}22, SG20b, SSF23, TJ21b, WMRB22, WGJ^{+}21, ZK21, ZJ22, ZBF^{+}22, dKMK21, dV21, vdBSvdV^{+}23, vTHS21, ANS^{+}20, AMK^{+}21, BRP^{+}23, COJZ^{+}21, Ehl21b, FGK^{+}20, FTW^{+}23, GSGOP^{+}21, GBCNCL^{+}21, KMT^{+}21, RBB^{+}24, SRSGMGG21a, SRSGMGG21b, WJLSH21]. Plans [SGLF^{+}23, BD20, BRP^{+}23, CBO^{+}21, IKB21, NM23]. plant [FGPPPV21]. Plastic [WWD23, AAA^{+}24, AWBC20, AVF^{+}23a, AFS^{+}23, BMC20, SONB^{+}22, Dau23, DBS22a, DDD22, Gug24, HLCM21, JAH^{+}24, KK22, KLK23, MFT^{+}22, MGDJ22, MTC^{+}23, Men23, MLM23, NW20, Osm23, WYH21, WHL22, Wan23b, WFG^{+}22b, WNBRSG22, Wu22, XHZ^{+}21, ZMR20, ZHL^{+}23, NW20]. plastics [AVF^{+}23a, RU20]. Plateau [TH22]. platessa [NSK20]. platform [GS21, GVH^{+}23, HCT^{+}21, WLQD21, WCC^{+}20]. platforms [Cha24, TZM^{+}22]. players [MF20]. Plenty [TGM22, MJMAA21, SCD21]. Pleuronectes [NSK20]. plight [JS23]. plucked [SSH^{+}23]. plume [WABR^{+}22]. plundered [LHW21]. plural [Mba22]. PMS [VCC^{+}20]. poachers [GRA^{+}22]. poaching [MJ23]. points [BHB^{+}24, SAO^{+}21]. poisoning [Nal24]. Polar [Sos21a, Geh23, KOBS21, Liu20, FFE20, KOBS21]. POLARIS [FFE20]. polarisation [CVA22a]. Policies [AWBC20, AS23a, DRCdMCJ21, hBDF22, GBC^{+}20, GSB20, HT23, HSHS24, KGP^{+}22, Kon20, LBWS23, MDLA21, MVSP^{+}21, McC20, OKNB22, PK23, PK20a, QS21, RBEA^{+}22, RXX24, Shi23b, TDC21, WYH21, WYD^{+}22, XXMZ23, YW20, YYL20, YYD23, ZV20, dSdLSGB22]. Policy [DBHT22, EHM^{+}22, FZYF23, FGC21, GZW^{+}20, HPS23, KSS23, KL22, Mal21, Mot22, PCCOS24, PMFR^{+}22, QTD21, RP23a, RJAT23, SCSR^{+}23, SBJF^{+}21, TGM22, VMV^{+}24, WM22a, YWYL22, ZWC^{+}23, Åm21, AOO^{+}23a, ALML23, AMR21a, ALF^{+}22, AFS^{+}23, ADB^{+}22, AC23, BH23a, BDMW22, BB21, BISSBdlMM22, BUR^{+}22, BOW^{+}22, BDM21, CA20, CMWN^{+}23, CNG^{+}23, CLCV21, CMS^{+}21, DL21, DC20, DKKS21, Dbo22, De 20, DMS^{+}24, Ero22, FXC^{+}22, FAC^{+}23, FFD^{+}23, GSA23, GR20a, GJY23, Gor21, GVR20, GABA^{+}23, Guo20, HLCM21, HS22, HLLK22, Hon21, HSHS24, HHD20, HWY22, Isr22, JC23, JKV^{+}22, JAH^{+}24, JSW22b, KL21a, KA21, KCCC^{+}21, Kao23, KFP20, KSB^{+}23a, KSB^{+}23b, KvTNL20, LTW23, LDAC20, LDC^{+}20, LXSX21, Lov23, LSD^{+}22, LFVCF24, Mar20a, MPH^{+}23, MF23, Mor20, NSZ^{+}21, NH20, NCE21, NAP^{+}21, OBNK24, OOGA22, PKR^{+}22, PKC23, PKH22, Pra24]. policy [QYD20, RRR^{+}23, RA20, RHC20, RS24, SK20b, SDI^{+}20, SSDQ20, Smi20, SS21b, Sta23a, SRH23, SWXM21, SCY24, TYS24, TM23a, TKP23, TMM^{+}23, TMTC20, THS^{+}24, TAP^{+}23, VÖB20a, VÖB20b, VMD^{+}24, VSPMS^{+}22, WXLD21, War23, WGR^{+}22, WtDD^{+}21, WCM^{+}24, WFG22c, XXMZ23, YU20, YH20, YML^{+}21, YCBW21, ZLPT23, ZSZ20, ZLZ21, ZY21, ZLH21, ZTA23, APF22, BTE21, DMG^{+}23, Gun21, HBP^{+}21, KGW20, PV23, SPFA^{+}20, WH20, YBW20, ZSLY20, Smi23a, GAD^{+}23]. Policy-oriented [PCCOS24]. policymaking [DSQ^{+}22, QLU^{+}23]. polish [BN20, CMOJ^{+}24, PKS20, RB22]. political [JL21b, Lou22, Mal21, SDL^{+}22, Tse21]. politics [Abo23, KGP^{+}20, Mar21a, SBS21, Smi20]. pollutant [VGK23]. pollutants [YZG23]. polluter [Zhu23]. Pollution [Rub24, AWBC20, ALML23, AFS^{+}23, BG20a, BMC20, CAC23, CLD^{+}24, Dau23, DBS22a, DPB^{+}23, DDD22, Gug24, HLCM21, HKMS21, JL21b, KK22, KSS23, Kuł22, LTJ^{+}23, MGDJ22, MTC^{+}23, Men23, NW20, Osm23, Pra24, RMF^{+}21, SGB21, Sha20, Tru23, VÖB20a, VÖB20b, VMV^{+}24, WYH21, WHL22, WRL23, Wan23b, WFG^{+}22b, WNBRSG22]. polymetallic [JACD20, VCC^{+}20, WABR^{+}22]. Polynesia [FLB^{+}21, SO20]. Pool [KL21b]. pooling [SII22]. Poor [RPY^{+}24, Alm21a, AML^{+}21, PBS^{+}23, SCBdA20, YTG^{+}23]. population [CNB^{+}23, IAB^{+}20, KYMW21, KTC^{+}22, MGBS24]. populations [BMF20, PMFR^{+}22, SSH^{+}23, SO20, SLK^{+}23, YNZ22]. porcelain [LHW21]. Port [AE23, BHEE22, COJZ^{+}21, Hon20, OEV23, Osm23, ZZH21, ZSWJ23, BDMW22, CMOJ^{+}24, Dun21a, FXC^{+}22, GAMR20, HPS23, HLLK22, HMY^{+}23, KKDK20, KSS23, KABT21, Okş23, PGSS22, RC22, Şan20, SPMN23, WP23, WdSRY23, ZWG^{+}21, BJH21, BHCL23, HNMSG21, Kar22a, LMSLMQ21, XWL^{+}20]. Port-based [ZZH21, ZSWJ23, PGSS22]. Port-Cros [HNMSG21]. port-level [SPMN23]. portion [DAHS^{+}23]. Portofino [VMC^{+}21]. portrayals [LDL21]. Ports [ABME22, BGLTC21, CDNS23, CM23, GGL22, HMY^{+}23, LMSLMQ21, Öİ24, PYVF24, SOK^{+}22, VÖB20b, VGK23, ZZH21]. Portugal [MCRPC20, ACV^{+}21, BASM22, CSQ^{+}24, CACJR^{+}21, CLCV21, MFRQ23, RFPS21, SVP^{+}23, SCBdA20]. Portuguese [CHEC21]. posed [Wen20]. Poseidon [Smi22]. poses [SBT^{+}23]. positional [dSOW^{+}21]. Positive [BHB^{+}24, YBW20]. positively [YNZ22]. possibilities [ASM^{+}21b, SOK^{+}22]. Possibility [KLK23, IGS^{+}23]. Possible [Kao23, CJ22, ONA21, MCPS20]. post [BM21, DGS^{+}24, FFGJ22, KFP20, LLHG23, YCBW21]. post-2020 [FFGJ22]. post-Brexit [DGS^{+}24, KFP20]. post-epidemic [LLHG23]. postharvest [GAA^{+}20]. posthuman [CTH^{+}23]. pot [HGG^{+}22a]. Potential [CDC21a, DD22, FOKM23, GGL22, KWS20, LWZ23, WKO^{+}22, WWG^{+}23, YK20, ZZD^{+}23, vdGD21, AS23b, BSM23, CVH^{+}20, Chr20, GMGLV21, GMGLV22, HSH^{+}21, HKMS21, LZRC23, MdCLL^{+}23, ONA21, OKH^{+}24, PCG^{+}23, PL20b, Pra20, RSFvH^{+}22, SGB21, SHP^{+}21, Squ20, Tes20, Tod22a, VCC^{+}21, YH20, dSOW^{+}21, EV22]. potentially [CAHR^{+}24]. Potentials [HIP^{+}21]. poverty [OWS20, RMBR22, Sol23b, ZZW^{+}21]. Power [AH21, BHCL23, DBHT22, Har21, KNY20, KSS23, MCL22, MMK^{+}22, Sta23b]. powerhouse [Wil21a]. Pp [Smi23b, Niu22, Rub24, Smi22, Sta23a]. pp232 [HH21]. Practical [HHD20, CSL21, MB22, NOB22, AAA^{+}21, SG21]. Practical- [SG21]. Practice [BKW^{+}24, ALF^{+}22, CLL22a, CJ22, GG21, HBP^{+}21, JSW22b, Lin24, MJMAA22, Oan22, SWI^{+}22, SVKP21, TCZ21, TAP^{+}23, WM22a, YD22]. practices [AAKI21, ABM21, BLY^{+}24, BSK^{+}20, CRP^{+}20, CWKW20, DT23, DPB^{+}23, DRY22, FTM^{+}22, FGK^{+}20, GMPT^{+}20, HT23, KCCC^{+}21, KMT^{+}21, MF22, MWM^{+}23a, MMGAA^{+}21, ONN^{+}20, Pra20, SBFHD^{+}23, SPHO23, Squ21, TOAPJ21, Tse21, WH20, WSDB20, dlPdlLLC^{+}22, vdBSvdV^{+}23]. practitioners [HZ23, SBP24]. pre [MÇ23, BM21]. pre-pandemic [MÇ23]. preamble [Lot23]. precarious [YS23]. Precarity [CJ23]. precaution [KGL22, Pen22]. precautionary [MXZ23, PL24b, WH20]. precedents [hC24]. Preconditions [COJZ^{+}21]. precursor [GR20a]. predators [MLAS23]. predict [GABA^{+}22, HHEJB20, VPRR^{+}21]. predictions [SR22]. Prefecture [Tru23]. preference [CMR21, MMW^{+}23, WLX21, WYD^{+}22]. Preferences [DSB21, CDD^{+}23, DWC^{+}20, GTTM21, MM21b, POM24, Pra20, Sol23a, WMS21, XXMZ23, YFD^{+}21, vRB24, BR22]. Preferential [MMO^{+}24]. Preliminary [GMCMJA^{+}21, MH20b, DDD22]. premium [SPG20]. preparation [Set23]. prepare [IGS^{+}23]. preparedness [McC24, dKMK21]. Preparing [MN23]. presence [PCD^{+}23, Hon21]. present [BMMF22, DL21, Fra23, Liu21, MCRR23]. present-day [Fra23]. presentist [TDP20]. presents [OGS24]. preservation [JACD20]. preserve [GPM20, KGW20]. preserving [FWGM24]. presidential [TLYC20]. Press [Niu22, Smi20, Smi22, Tan23a]. pressure [DXS^{+}21, GABA^{+}22, Mil22, SH22]. pressures [AH23, MLJ^{+}22]. Prevalence [HHH^{+}23]. prevent [MM20]. prevention [CLC21, LDAC20, LC20b, LFVCF24]. previous [hC24]. Previously [TJ21a]. Price [Gor21, JG23, Cat20, FR23, MMM^{+}22, May22, SPG20, SVV24, SKE22]. Priced [EP20]. prices [NTH^{+}21]. pricing [RLRS21]. primer [HSK^{+}23]. principal [MF23]. principal-agent [MF23]. principle [Che21b, LC20a, MXZ23, Zhu23]. principled [Dbo22]. principles [Dbo22, DC23, JSW^{+}22a, Koj23, Shi23a, War21]. Prionace [DSL20, SSA^{+}23]. prior [DLJ^{+}20, Lev20]. priorities [BDM21, CLRW22, FWGM24, GAD^{+}23, HPPH22, LJ20, Mur21b]. prioritization [MCTN23]. prioritizing [OCCSC23]. priority [GWAPT21, VBC^{+}20]. prisoner [Zha21]. Private [WKBS20, CRJ22, DSWB20, GEB^{+}21, GBC21b]. private-boat [CRJ22]. privatization [CAHR^{+}24]. Privy [RROP23]. pro [CCK^{+}21, WFG^{+}22b]. pro-active [CCK^{+}21]. pro-environmental [WFG^{+}22b]. proactive [HM22a]. problem [Art20, Fen23, GVR20, MF23, OOGS20b, Pie21, TCT21]. problem/project [TCT21]. Problematizing [Lot22]. Problems [JPM^{+}23, Gro20, HIP^{+}21, Pau21, SOLN21, AGS20, ZL20]. procedures [Kak20, BBBS22]. process [ANS^{+}20, ANR^{+}22, BD20, hC24, DBS^{+}22b, HF20, JDF^{+}23, JC23, ML22, MPH^{+}23, MKB21, SS21b, SB20, Wan23b, Wan23a, ZK21, WJLSH21]. processed [GSG^{+}24]. processes [GM21, RA20, vHSSK20]. processing [CBD^{+}23, DSWB20, HRC^{+}24, PCSS^{+}21]. processors [IOHC22]. produced [GEB^{+}21, WXLD21]. Producer [CGGA20, HLCM21]. product [QZZX20, QZZ20, TOHK20]. production [AOO^{+}23a, AED20, BABC20, CSCGC20, EMS^{+}23, FXC^{+}22, FTM^{+}22, HCSBM^{+}23, IGO21, JY20, LGF24, LLW22, LJ22, LSY20, PL23, SCJ22, SHP^{+}21, SOS^{+}21b, SB20, YCL22a]. production-based [LLW22]. productivity [FWGM24, LXSX21, NHA23, TJZ^{+}21, WLY21, WW21b]. products [CMR21, CLCV21, LHV^{+}20, MLM23, PMK^{+}21, QZZ20, SPG20, SFN^{+}21, WM21]. professionals [MLE^{+}24]. Profit [OAM20]. Profitability [CCGM21, JHL21, TCS^{+}20, BV21, RAFLRP21]. profits [MM20]. program [AACdlT23, AGK22, IMA22, TCT21]. Programme [FFGJ22]. Programmes [GKP^{+}22, OKNB22]. programming [PFWM23]. programs [AACdlT23, CTC^{+}23, EHM^{+}20, EHM^{+}22, STIS20, ZLD^{+}21]. Progress [FTSG21, MG22, MDJ^{+}22, RHO24, CTCD21, EdLP^{+}21, Ehl21b, GMG^{+}23, GKP^{+}22, Mah22, QLZ^{+}24, RZ21b, Ren21a, Shi20b, VSV^{+}23, ZL20]. prohibited [CMI^{+}22]. prohibiting [VJL^{+}21]. Prohibition [Kuł22, NW20]. Project [Smi22, BHCL23, CTIH21, GBCNCL^{+}21, KNY20, LJ20, QTD21, TCT21, BMM^{+}20, MF20]. projections [Gor21]. projects [BDMW22, LLSY22, MPSF20, OWS20, RBEA^{+}22, TT22, WS23]. prolifera [XLLS21]. proliferation [SMCW22]. prominent [WNBRSG22]. promote [CSHO^{+}20, GCV^{+}21, Gir22, GMRSP23, HPH22, Liu22, NWZ^{+}21, TLYC20]. Promoting [HC22, cL22, WFG^{+}22b, HWMO20, OCD^{+}23, WKO^{+}22, WP23]. promotion [JL21b]. propagation [ZXGL21]. Property [Run24, AELR^{+}20, RDLA^{+}20]. Proposal [SLP^{+}22]. Proposals [MPD^{+}23]. proposed [Dun21a, HGL^{+}20, LR20, MK24, SWHT22, TSH^{+}22a]. propositions [OCD^{+}23]. propulsion [KDS22, THS^{+}24]. prosecution [Álv23, Pat22]. prospect [Wu22]. prospective [CHEC21, DC20]. Prospects [TCC^{+}22, CCX23, GCC^{+}21b, HLM^{+}20, Wan23b, YW20, ZL20, ZLZ21, ZMM^{+}23]. prosperity [LP20]. Protect [RJAT23, ALML23, ADB^{+}22, DMG^{+}23, Mag23, MH24, SMF^{+}23, SMCW22]. Protected [AIG^{+}20, ACA^{+}21, ÁFFSC20, AVF23b, BvBN^{+}20, BEH^{+}21, CCS23, VBd^{+}21, FMLMM23, GABA^{+}22, HSGNMGC21, KJE21b, NJBM^{+}23, ORSB22, QTBL23, RA20, RP23b, RCM^{+}22, VMC^{+}21, WJLSH21, AHAB^{+}22, ASM^{+}21b, BDM20, BD23, BCJ21, BUM^{+}24, BBK^{+}21b, CMGF20, CLRW22, CJ22, DDAB^{+}23, Col20, DLJ^{+}20, DKW20, De 20, DB20, DSC^{+}23, Dua22, EV22, FTGD^{+}20, GFD21, Gar20, GCMT20, GCL21, GBC21a, GBC21b, GMVM23, HJ21, HPPH22, HLM^{+}20, HCT^{+}21, JPM^{+}23, JL21c, KFM^{+}20, KDM^{+}20, KOCL23, LZRC23, MM21a, MCRPC20, MSBA22, MMO^{+}24, NNM^{+}21, PBG23, Pha20, PBH^{+}23, PKMT23, RP23a, RPY^{+}24, Rus20, SAO^{+}21, SMLRB^{+}21, SPHO23, SJ21, SDG^{+}21, SDG^{+}22, SBT^{+}23, Sol23a, SO20, SB20, TLYC20, THT^{+}23, TBH^{+}21, TEA^{+}22, Tra23a, TJ21b, VWFW22, WD20, WMRB22, YT20, YCL22b, YD22, YC21b, ZPX^{+}22, dS22a]. Protecting [AH23, CMWN^{+}23, HJ21, SB20, TD21, Tod22b, dS22a, AMLC^{+}23, Li23b, Pyć23, WvdMG^{+}21]. Protection [BGK^{+}22, CR23, MSP23, RCM^{+}22, WM22a, BNF21, BC20b, BC21, BL23b, BS22, CJ23, CLP21, COJS^{+}22, DSC^{+}23, FFYE21, GCG^{+}21, GW21, Gul23, Guo20, GMVM23, HVS22, KWG23, KN21, LHW21, LWG23, Lin20, PT23, PC20, RHC20, RCH22, SMDM20, SDG^{+}21, SDG^{+}22, SOS^{+}21b, SRLLG23, Tra23a, VTRS20, DR22, VBC^{+}20, VvSvH^{+}22, WYD^{+}22, WBPG20, YL21, LC23]. protein [Tel23a]. provide [GCM^{+}23]. provides [KFP20]. Province [HFD^{+}20, LYL^{+}22, RYS^{+}21, JTR^{+}20, NHD23, QS21, XWH^{+}22]. provinces [CQ20, RZ21b, RJ21]. Provincial [Zho23]. provision [CWKE20]. provisioning [SWH^{+}23]. Provisions [BGK^{+}22, ML22]. proviso [Tan24]. prowess [PRM^{+}20]. proxies [DR22]. PSMA [HMY^{+}23]. PSSA [Cho22]. psychology [YBW20, YCBW21]. Public [CKY20, CCW21, hBDF22, GPM20, HPC^{+}23, HHEJB20, KNY20, KNY22, MCR22, VDP^{+}22, WLX21, AACdlT23, Bea22, CKY22, DSWB20, EV22, Fen23, GBB^{+}21, GTTM21, Hon21, Kak20, KSB^{+}23a, LDL21, MVSP^{+}21, MBB^{+}22, OKNB22, Pra24, RHC20, RS24, RPY^{+}24, STIS20, VSPMS^{+}22, WWD23, WFG22a, vRB24]. publications [AMR21b]. Publishing [Sos21a, Sta23a, Tan23b]. Puerto [MTC^{+}23, PSPR20, VGAH^{+}21]. Puget [CTC^{+}23]. pulse [KGP^{+}20, Pen22]. Punta [GBC21a, HNMSG21]. Punta-Coda [HNMSG21]. Purba [CDP22]. purchase [DLYM22, MLM23]. purpose [GNJR24, Laj21, WS23]. purposes [SSMG21]. Purse [RROP23, BLH^{+}20, EMH^{+}23, TYS22]. Pursuing [OI23]. pursuit [BLL22]. Putting [Jon21]. puzzle [RE23].

Q [CGGA20]. Q-methodology [CGGA20]. Qatar [Hon20]. Qatar-Gulf [Hon20]. Qin [Sos21a]. Qingdao [XLLS21]. QRE [ZLGX20]. qualify [WS21]. qualitative [KDS22, MKB21, NEF23, QTD21, QUK21]. qualitatively [QLU^{+}23]. Quality [Run24, SGH22, ASV^{+}21, ÁFFSC20, BDR20, Cat20, Kon20, LXW^{+}21, MKAF21, RB22, SGZH^{+}21, WLY21]. Quality-enhancing [SGH22, BDR20]. Quantification [BAP21, OVA22]. quantify [MEG^{+}21, VAMR^{+}20]. Quantifying [HKMS21, SRS23]. Quantitative [HSHS24, KCCC^{+}21, KDS22, MKB21, OMC^{+}22, PMT^{+}20]. quasi [RXX24]. queen [MVR21b]. question [SGA^{+}21, AM22]. questionnaire [Zhu22]. questionnaires [Öİ24]. questions [HMG^{+}21a, WLCC23]. quiet [Kar23]. quo [Zhu22, Wan23b]. Quoc [KK22]. Quota [FDF^{+}21, AGN^{+}20, BFMS21, BADO20, Cro21, HWT23, HvPPV20a, KV20, SS22, EP20]. quotas [BXV20, HvPPV20b, SSAP24].

race [BCMCM21]. Racing [AR22]. radio [BMRM23]. radioactive [MXZ23, WM21]. Radionuclides [PL24a]. rahui [FLB^{+}21]. Raja [SWI^{+}22, WMRB22]. Rallying [Fel22]. ranching [DG20, DL22, HdM23, QYD20, QS21, WXLD21, WLQD21, YW20, YWYL22]. ranchings [QTD21]. range [CMS^{+}21]. rankings [FMGV20]. Rapa [DUYEA21]. rapid [AHAB^{+}22, GTMP20, HF20, MIH^{+}24, SNK^{+}20]. rapidly [MCW^{+}22b]. raster [CA21]. Rate [HGG^{+}22a]. rating [IMA22]. ray [BdLAdC^{+}20, GRPE^{+}21, PMM^{+}21, SEG^{+}22]. Rayfuse [Tan23b]. rays [HMG^{+}21a, RPK23, TJS22]. RBM [ALF^{+}22]. RCEP [Cha22]. Re [BVvKL21, OGS24, Ren21a, LMMT23, Öİ24]. re-centre [LMMT23]. Re-examine [Ren21a]. Re-framing [BVvKL21]. re-planning [Öİ24]. Re-thinking [OGS24]. reached [MTN23]. reaching [SPCCvT21]. reacted [SRH^{+}22]. Ready [FAB^{+}23, GTA^{+}20, Gar20, LGW^{+}21, MWM^{+}23a, RDT^{+}23, SSAP24]. Real [FGPPPV21, TZM^{+}22, WPWM^{+}22]. real-time [TZM^{+}22, WPWM^{+}22]. Realising [SDI^{+}20, NSN22]. realities [AVF^{+}23a, SPHO23]. reality [HBL^{+}21]. realize [Ren21a]. Really [Gue23]. realm [BSM23]. reasons [GWAPRPRU22]. Reassessing [SLK^{+}23]. rebuilding [AMR21a, Ber22, DSL^{+}21, DGS^{+}24]. Rebuttal [TGM22]. ReCAAP [POO23]. recall [GVH^{+}23]. reception [Osm23]. reciprocal [KTI^{+}24]. Reclamation [Squ21, LDC^{+}20, LYHL23, MX21, SL20, XWH^{+}22]. Recognising [PJN24]. recognition [BMMF22, EAA^{+}22, OVA22]. recognized [GBC21a, VvSvH^{+}22]. Recommendations [GPH^{+}21, PNS^{+}23, ALF^{+}22, SONB^{+}22, DBBL21, Ero22, FMLMM23, GPG^{+}21, GSR21, HBP^{+}21, Isr22, NCD23, OBNK24, PKB^{+}23, TYS24, VWFW22, VMV^{+}24, LHV^{+}20, MT20]. Reconciling [Che21b, LMC^{+}23b]. Reconfiguration [VKS23]. Reconsidering [CD23]. reconstruction [GÁPSAP^{+}23, VKS23]. recordings [BBB23]. records [BBK21a, OWCFDC21, SSZ^{+}24, WWCS21]. recover [KYMW21]. recovery [BMGDS^{+}20, CNB^{+}23, DS22b, IKB21, Isr22, WH20]. recreation [WFR23]. Recreational [DWC^{+}20, SVP^{+}23, VMC^{+}21, AVPAFM^{+}22, CACJR^{+}21, CAM^{+}23, CRJ22, ES22, dlFGFFR^{+}20, GSG20, GCL21, GMCMJA^{+}21, GVH^{+}23, KWS20, KWF^{+}21, MRKM23, MJMAA21, MJMAA22, MMO^{+}24, OGS24, PHF20, QWBY21, RDD^{+}21, SCD21, SSA^{+}23, SBT21, TGM22, ZCCQ22a, ZCCQ22b]. Recreationist [SON21]. recreationists [PC20]. recruitment [CC21, CCK^{+}21, QTPL21, UISM^{+}23]. Recycling [DN20, HP20]. Red [EMN^{+}24, HDAPG^{+}24, BMK^{+}22, DD22, May22, MKM^{+}20, SPCCvT21, VKN20, YSEK22, YNZ22]. Redistribution [MRKM23]. reduce [AWBC20, BAD21, CBR21, CPR21, FMWW21, HLCM21, HSH^{+}21, KSS23, NSPRSS20, NW20, PBG^{+}22, SDG^{+}23, SMLRB^{+}21, SNP21, SSB^{+}22, SLMG^{+}21, WYH21]. reduced [SWF^{+}24]. reduces [Shi22]. Reducing [DPB^{+}23, GMM^{+}20, MGDJ22, MM20]. Reduction [SGLF^{+}23, AAA^{+}24, AFS^{+}23, HLCM21, HRS^{+}24, MPL^{+}23, MM23, OOGK^{+}24, RQB^{+}20, Sol23b, ZZW^{+}21]. Reef [VBd^{+}21, AKFY22, BBC24, CWB^{+}20, CLRW22, CWKE20, HVS22, HF20, LRD^{+}22, Lev20, McC20, McC21, NTH^{+}21, PDJ22, RBEA^{+}22, SKE22, SO20, YFD^{+}21, DR22, Wat22]. reef-building [AKFY22]. reefs [AHN^{+}23, BDMW22, BS22, Dun21a, Dun21b, GA20, GN22, MDG^{+}23, SVF20]. reference [JACD20, TCZ21]. refining [ZGY21]. Reflections [EFFP23, AAA^{+}21, XT23, ZY23, BG20b, DPP^{+}20]. reform [CMOB^{+}20, SK20b, ZPX^{+}22]. reforms [MVR21a, Sak21]. Refuge [GBC21a]. refugia [CWB^{+}20]. regard [Lin24, MS20]. regarding [CLL22a, EV22, EVPMKP22, GLADGGVL23, Pen22, SCGC20, SS23, Tan24]. Regency [HF20]. regime [CC21, Gar22, KY20, LV23, OA20, Tel23b, TBH^{+}21, TN21, YK21, XWL^{+}20]. regimes [GSAF21, GRPRVG23, Kar23]. Region [BRR^{+}22, GBC21b, OA20, YMG23, Als20, ABRA20, ADB^{+}22, BEH^{+}21, CVCV^{+}22, CBO^{+}21, CRCP^{+}23, Kas21, LEVET21, PKB^{+}23, TVN23, ZFGL20, EMN^{+}24, GAA^{+}20, Kiz24, MÇ23, Şan20, SR22, Tod22b]. Regional [EHM^{+}22, FFGJ22, HDH21, HDCH21, IOHC22, LXH^{+}22, LSH21, LFLF22, RCH22, Ros22, War23, BLL22, BDD^{+}21, BMM^{+}24, CS21, CCCC21, Cla20, COM20, EVPMKP22, dlFGFFR^{+}20, HvT22, HMFH20, HJPL23, JLY23, JK22, LJ22, Mar21a, Osm23, PWL^{+}23a, PWL^{+}23b, QYX21, SOLN20, SLH22, SHFP21, SHZC21, EHM^{+}20, GML^{+}24]. regions [Alm21a, FMGV20, HM22c, KOBS21, Liu20, NWZ^{+}21, Sha20, SGK21, Sos21a]. register [HVN21]. register-based [HVN21]. registration [PKR^{+}22]. registries [WOG20]. regression [JHL21]. regulate [ABI24]. regulated [SKL23]. Regulating [Doe21, HSSM24, WMKG21, Che21b]. Regulation [AKL^{+}22, WL20, AKFY22, CQ20, DG24b, DER^{+}21, GG21, KMM21, Kon20, LC23, LC22, NWZ^{+}21, OOGA22, RJ21, Sam22, SKE22, SDG^{+}21, SDG^{+}22, TBV^{+}23, TML^{+}21, YL21, YZG23, ZW21, Zhu22, KA21]. regulation-centered [SDG^{+}21, SDG^{+}22]. Regulations [MK24, Álv23, AP23, CLL20, Dam24, FMCD20, JKV^{+}22, NSPRSS20, SFN^{+}21, VMV^{+}24, YU20, ML22]. regulators [MCR22]. Regulatory [STS^{+}23, AR22, BTV^{+}22, DBZ24, DSC^{+}23, FAC^{+}24, Lov23, OC20, Squ21, Wu22]. Reichelt [Rub24]. Reichelt-Brushett [Rub24]. related [AH23, AGM^{+}22, CDRS22, FTGD^{+}20, MN23, Pau21, QZZ20, SAO^{+}21, SWH^{+}23, TM23a, WMS21]. relation [CMR21, Dev22, IOHC22]. relational [HC22, MUR^{+}21a, PPK23]. relations [Gra21, KA21, LJ23, MMK^{+}22, Pau21]. relationship [CLW^{+}21, DC23, SSA^{+}23, Shi23a, Tse21, WSD21, ZY23]. relationships [CCK^{+}21, MEG^{+}21, PCA^{+}23, RMU^{+}22, TJ21b, WP23]. relative [HGH^{+}20, IAB^{+}20]. release [WL22a]. releasing [MXZ23]. relevance [TMTC20]. relevant [BCBS22, Gal22, GHW^{+}22, KVL^{+}23, War21]. Reliability [KMG^{+}20]. Remaking [LLT^{+}20]. remedies [FBC^{+}23]. remote [ASM^{+}21b, HMG^{+}21a, KZR22, LWSS21, VSV^{+}23, YCL22b]. remote-controlled [KZR22]. remotely [Stø21]. removal [ADR^{+}21]. removing [VSDR^{+}22]. remuneration [Mat20]. renaissance [FLB^{+}21]. renewable [GSR21, WNMS^{+}24, vRB24]. Renewal [DD22]. Rent [AMIM21, Åm21]. Reorganising [Bot20]. Reorganization [LCL20]. reorienting [CFM^{+}21]. reply [SCD21]. report [BUM^{+}24, CLS20, CS20, DRY22, DČFW23, DC23, GMG^{+}23, LLHG23, LSH21, MdCLL^{+}23, MH20b, QYX21, RFPS21, SWGZ22, SCA^{+}22, SHZC21, TZC23, Ton22a, Ton22b, TC23, YC21a, ZCZ20]. reporting [HBHJ20, ZLD^{+}21]. Reports [SBP24, PKB^{+}21]. representations [HPG20, NGP^{+}21]. representative [CES20]. representativeness [MPD^{+}23]. reproductive [LLLF22, MWB^{+}21]. Republic [TKC23, KPN20, Kim20, Kim21, KLK23, PSPR20]. requirements [BRM21b, KMJ24, MPH^{+}23, RSFvH^{+}22]. requires [AHAB^{+}22, WLW^{+}22]. rescue [BMRM23, CDRS22, Sar23, CGP23]. Research [CMC22, HSHS24, Ngu23, Ren21b, RXX24, VCC^{+}20, XHZ^{+}21, XHY22, ZHL^{+}23, ZWXL21, BMM^{+}20, CKY22, GEB^{+}21, HZ23, LMT21, LSD^{+}22, MIMS^{+}20, MH24, MM23, MZT^{+}23, OWCFDC21, PSD23, SCA^{+}22, STF21, TV21, TMM^{+}23, WHL22, WRL23, WSDB20, Yu22, ZLY23, Tan23b]. researches [AFS^{+}23]. reservations [SBT21]. Reserve [GBC21b, ADB^{+}22, McC21, QTPL21, CD23, MMSB^{+}20]. Reserves [RNGRC22, DTBP21, KAC22, KKO^{+}22, OVSVPP^{+}22, WBPG20]. Reshaping [AOO^{+}23a]. residency [NUD^{+}23]. resident [XLLS21, XXMZ23, XXLW24]. residents [HP23, KWG23, NNRS24, LYQM21]. residual [TN21]. Resilience [ABME22, BSI^{+}21, HCWC21, LMSLMQ21, CSVC23, CBD^{+}23, FTM^{+}22, GABSD21, Jon21, KAC22, LHPJ23, MKSG20, PGSS22, PALB20, Pyć23, SCLM^{+}21, SPMN23, TCS^{+}20, VSPMS^{+}22, VLS^{+}23, YG20, YFD^{+}21, ZLH21]. resiliency [CSGC22]. resilient [TC23, YG20]. Resolution [Yu22, Kar23]. resolved [Sin21]. Resolving [COSS21, GVR20]. Resort [HM21]. Resounding [SEG^{+}22]. Resource [MM21b, Åm21, AML^{+}21, BDDE20, hBDDF22, DCC^{+}20, GSAF21, GRPRVG23, GABA^{+}23, Hor23, HGL^{+}20, Isr22, Kaw22, KGW20, LR20, MJ20, MWM^{+}23b, PK23, SOP^{+}22, SCCW21, VKN20, WWG^{+}23, YMG23]. Resources [Smi23b, Sos21a, BH23b, BM21, CCK^{+}21, CHB^{+}20, FR23, JKV^{+}22, KN21, KTI^{+}24, LDL^{+}23, Liu20, MN21, Ren21b, SCJ22, THT^{+}23, WRV21, ZSWJ23, GB21]. respect [Ros22]. responding [Dun21b, LC20b]. Response [MJMAA22, MJ23, JHHG23, Kar22b, Muk21, PL24a, PKH22, SGK21, SPJB22, SHFP21]. Responses [DSWB20, BC21, BvdH23, CLB^{+}20, LEERRM^{+}21, MAK22, Zou23]. responsibilities [Che21b]. Responsibility [TCT21, HLCM21, LZZ22, Liu22, RFG24, SMCW22, UOO20]. Responsible [MJMAA22, SCD21, FBC^{+}23, LLJZ22, TK22]. responsiveness [DKKS21]. rest [BBK21a]. Restoration [GN22, DRY22, GTTM21, LRD^{+}22, QTBL23, RBEA^{+}22, SBJF^{+}21, SPCCvT21, WLX21, XXLW24]. restoring [Wat22]. restraint [KMM21]. Restricted [VCC^{+}21, OCCSC23]. restriction [McC21, WYD^{+}22]. restrictions [KL21b, SLL^{+}24, SNK^{+}20]. restructuring [Ton22a]. resulting [CWKE20, Fei21, JF22]. Results [ALF^{+}22, DAA21, RZ21b, CLL20]. Results-based [ALF^{+}22]. retention [JAK^{+}23, WFG22c]. Rethinking [AVF^{+}23a, CTH^{+}23, MM21a, RU20, DCC^{+}20]. retirements [Ber23]. retrieval [ABDRV^{+}23]. Reunion [Los23, Los24]. reveal [AGK22, AMLC^{+}23, BC20b, PVH^{+}21, WWCS21]. revealed [GPTASM23, TSLK23]. Revealing [LMC^{+}23a]. reveals [ALLR^{+}24, EdLP^{+}21, OCZ^{+}22, SWH^{+}23]. revenue [Gor21]. revenues [Dom23]. Reverse [BB21]. Review [AFHM23, Cha20b, De 20, DSJ^{+}23, HH21, Lv20, Ngu23, Niu22, Rub24, Smi20, Smi22, Smi23a, Smi23b, Sos21a, Sta23a, Tan23a, Tan23b, BDM20, BOW^{+}22, BRP^{+}23, BABC20, BSK^{+}20, CMWN^{+}23, CJL^{+}23, CNG^{+}23, DGMF^{+}20, EdLP^{+}21, GPA^{+}22, GR20a, GMVM23, HHBC20, HLLK22, Jen23, LRC24, OT23, ÖB22, PNS^{+}23, PCK22, PV23, QLU^{+}23, QK24, RBEA^{+}22, SMDM20, SM23, SRH23, TSH^{+}22a, WvdMG^{+}21, WNDZ23, YL21, ZLY23, vdBSvdV^{+}23, JC23]. Reviewing [DR22]. Revillagigedo [CWB^{+}20]. revised [DMG^{+}23]. revisited [Zha21]. Revisiting [VSD^{+}23, Sel20]. revitalisation [CS21]. revitalization [CCCC21, HC22, LLSY22, TCT21]. rewards [PY20]. RFMOs [GML^{+}24, Fis22, FOKM23, Kao23]. Ria [BASM22]. Rica [RJAT23, Teo23, PSPR20]. Richard [HH21]. Richest [OYKBA22]. Rico [PSPR20, VGAH^{+}21]. Ridge [BMR20, KY20]. ridges [DBH^{+}22, WvdMG^{+}21, CMWN^{+}23]. right [Bea22, Gal21, MM20, NM23, SPFA^{+}20, SHAC23, TDC21, RCM^{+}22, RM21b]. Rights [FMCT^{+}21, HDHM^{+}20, Run24, AR23a, AR23b, AEJ24, BBS^{+}21, BOW^{+}22, CdLN^{+}21, CK23, FFYE21, GE23, HLW^{+}20, Jen23, Li23b, MSP23, Mrd23, NOB22, NHA23, OAAMS24, Pyć23, RDLA^{+}20, SCGC20, SRY24, SRLLG23, WYD^{+}22, Wan23c, Who23, WKBS20, WS23, YL21, YLSC23]. rights-based [CdLN^{+}21]. rigid [RQB^{+}20]. Rio [SdACdSLdAS24, dS21]. rise [Fla21, GABSD21, Liu20, Mal21, Ora23, dS21]. Risk [Cha20b, CLL^{+}22b, CL21, JF20, AMR21a, BB21, CAC23, DDAB^{+}23, CGP23, DBS22a, GJY23, GWAPRPRU22, GHK20, HHBC20, HBR21, HBR^{+}23, HVN21, HSdT21, HMY^{+}23, JSW22b, LDL21, LMT21, LTRA20, Los23, Los24, LLL23, MDLA21, MCW^{+}22b, MHN^{+}20, OKH^{+}24, PB20, RKS^{+}24, SWF^{+}24, SLS^{+}21, VGS^{+}20, VDP^{+}22, VRBC^{+}21, VBH^{+}23, WCT^{+}20, ZCSK23, SLS^{+}21]. risk-based [VRBC^{+}21, VBH^{+}23]. Risk-sharing [JF20]. risks [BBWC21, CBAPM20, DS22b, FBC^{+}23, GCS21, LDAC20, PL24b, SCH^{+}24, SBT^{+}23, SLS^{+}21, TOL^{+}22, WWG^{+}23, PY20]. River [JZW21, NHD23, QYX21, SCBdA20, WHPF20, SCA^{+}22, DDGdS^{+}21]. RMSI [POO23]. Road [HH21, BMH^{+}22, SF22, TBPC20, ZWG^{+}21]. Roadmap [DN20]. Robert [De 20]. Roberts [Lv20]. rock [OAM20, OAM21]. rockfish [GBSC22]. rocks [Gul23]. Role [BBJ^{+}22, GA20, ABRA20, ALM^{+}21b, Bea22, Ber22, BGMCG24, CMI^{+}22, FFGJ22, HRMPCRRA22, HACCM22, IGO21, Kat20, KL22, LMC^{+}23a, LGF24, LJ21, MS20, MBL24, NSN22, OOGS20a, OAM21, Pet22, PK20a, RU20, RFG24, RC22, SGL^{+}23, SFB24, SOLN20, SPJB22, TN21, TNY23, THS^{+}24, TAP^{+}23, VÖB20b, WMRB22, ZLH21]. Roles [DG23, LNRM23, LLSY22, OBNK24, SWH^{+}23, ZZH21]. Romania [VTRS20]. Romanian [Tea20]. rookery [KL21b]. roots [BV21]. rose [BDCJ22]. Rosemary [Tan23b]. Ross [ADB^{+}22, BBK^{+}21b]. rough [LL24]. Route [TBV^{+}23, RZ21a, ZLGX20, ZXGL21, BSY^{+}20, Gun21, HLLK22, RZ21a, Tod22a, WDG^{+}20]. routes [Boy21, GSB20, GGL22]. Routledge [Ngu23, Smi23a]. Rule [CS20, KP24, BGK^{+}22]. rules [CC21, CLL^{+}22b, GML^{+}24, Mot20, Mot22, NT20, Pra20]. run [BCBS22, MLVT21]. rural [DCC^{+}20, FTM^{+}22, Fro21, RC22]. rush [FTG^{+}23]. Russia [AB20, EB24, Tod20]. Russian [DD22, FFE20, KV20, Tod22a]. RVSPM [GBC21a].

S [Fla21, Lv20]. S. [Sta23a]. Saccharina [HKKY22]. Safe [AFHHD22, HMY^{+}23, Ste23]. safely [NGRM^{+}22]. Safety [AKL^{+}22, Sha22, Stø21, VJL^{+}21, YBW20, BvdBvdFK20, BISSBdlMM22, FRM22, JG23, Kar22a, NSPRSS20, PK20a, SNS23, TSH22b, WM21, WW21a, BAP21]. Saharan [UOO20]. sailed [KOBS21]. Sailing [LL24, KP24]. Saint [Gir22, TJ21a]. Sainte [LTJ^{+}21]. Sakuraebi [LSY20]. Salas [CMWN^{+}23, DBH^{+}22, WvdMG^{+}21]. Sale [May22, Lev20, NTH^{+}21, PSMR20]. Sales [CCX23]. salinity [RKR20]. Salish [DAHS^{+}23, FMCD20]. Salmon [Ber23, MF20, Åm21, BVvKL21, BE20a, BBGHR23, BMH^{+}22, CAHR^{+}24, EHLT21, GVL22, KS23, LJ22, LKE^{+}22, Mrd23, POM24, RFG24, SZ20, SCLM^{+}21, SLMG^{+}21, STG^{+}20, TFF22, WFG22a, YK20]. salmon-farming [Mrd23]. saltfish [Ken20]. salutogenic [CHB23]. salvage [FY24, LC20a]. salvors [FY24]. Sam [GAA^{+}20]. Samar [PCA21]. same [MAR^{+}20b, dOPM^{+}20]. sample [HACCM22]. samples [ZTM^{+}23]. Sampling [GCC^{+}21a]. San [PB20, PCCOS24, THGR^{+}20]. sanctions [Fel22, Oan22, Ros22, Tei23, WOG20]. sanctuaries [INW20]. Sanctuary [KHMM20, MLJ^{+}22, Mar20a]. Sandwich [Tra23a]. Santa [QCM^{+}20]. Santay [MTC^{+}23]. Santos [PCSS^{+}21]. São [PMC24]. saratan [YSEK22]. sardine [MIH^{+}24, WWG^{+}23]. satisfaction [CLW^{+}21, LSY20, PSPR20, QWBY21]. satisfactory [HWT23]. Satisfying [UISM^{+}23]. Saving [PT23, JV20]. say [BBBS22]. Sbragaglia [MJMAA22]. Scalar [Abo23, RBB^{+}24]. Scale [CMC^{+}23, ÁCT20, ACV^{+}21, AFMdSdTdCC22, Art20, ASGZ22, AH21, BPA^{+}22, BMB^{+}24, BDM20, BJGG^{+}24, BBS^{+}21, BSI^{+}21, BASM22, CPBO20, CAVJ^{+}20, CCKK22, DLJ^{+}20, DAN^{+}23, DDGdS^{+}21, FS22, GLADGGVL23, GSG^{+}24, GPTASM23, GMRSP23, GNMGWM21, GMCMJA^{+}21, GFL^{+}21, HLW^{+}20, HYN^{+}23, Han21, HLW^{+}23, HMG^{+}21a, HTP^{+}23, HBP^{+}21, JAdS^{+}20, JKKT20, KMS20, KB20, KNY20, KWMM20a, KWMM20b, KJS^{+}22, LRN^{+}22, LBWS23, LJRHH21, LEERRM^{+}21, Lov23, LLS^{+}20, MMK^{+}23, ML21, MB22, MF22, MMB^{+}23, MM23, Mil22, MIH^{+}24, MMGAA^{+}21, MZT^{+}23, NH20, NAS^{+}22, OCZ^{+}22, OYKBA22, OOGS20b, OCD^{+}23, OMC^{+}22, OA20, PSPR20, PBG23, PGSFRMHH24, PL20b, PJN24, PALB20, PMC24, PKS20, RTMCVMDE22, RDLA^{+}20, RKS^{+}23, SPFA^{+}20, SIW^{+}21, SLB21, SCJ22, SSZ^{+}24, SdCMN^{+}22, SS21b, SSP^{+}21a, SPA^{+}21, SGL^{+}21, SRH23, SSP^{+}21b, TTA^{+}20, TPG^{+}23, UOO20, VTRS20]. scale [VGC^{+}22, VGAH^{+}21, WKO^{+}22, YCL22a, ZLPT23, ZJ20, dlPdlLLC^{+}22, NCE21]. scales [KBD^{+}20, KVL^{+}23, OPL20, SHPW21, VRBC^{+}21]. Scaling [Obu20, LRD^{+}22]. scallop [BDD^{+}21, YML^{+}21, YTG^{+}23]. scalloped [ALLR^{+}24, Mag23]. scandals [VM21]. SCAR [BL23a]. scarse [CPBO20]. Scenario [IGS^{+}23, YHU22, dKMK21, GOH^{+}20, HM22b, PMC24, SMC^{+}21]. scenarios [BCBS22, CPV^{+}21, CVCV^{+}22, KDS22, PJN24, VPRR^{+}21]. scenery [CDP22]. schedules [vLPBK21]. scheme [BE20a, GHK20, HCS21, dlÁPAFGO20, RRR^{+}23, WLXY22]. schemes [AO20, BMB^{+}24, BI21, Sah22, WWC23]. Scholarly [XT23, ABM21]. school [TSWC23]. schoolchildren [WFG^{+}22b]. schools [cL22]. sci [SLY21]. sci-tech [SLY21]. Science [De 20, KGP^{+}20, MZ24, Pen22, AFHHD22, AMR21b, CMT^{+}23, CMS^{+}21, FLB^{+}21, GNJR24, GSA23, Geh23, GS21, GVH^{+}23, HWU^{+}23, HLCM21, HMG^{+}21b, HGG22c, KDH^{+}21, Kra22, LNRM23, LGW^{+}21, MFM21, MH24, OKNB22, OBNK24, PKH22, PZPB21, Ren21b, STKC22, SB20, TSLK23, TMM^{+}23, WSD21, WP23, WGR^{+}22, WFG^{+}22b, vSHHT20, Pol23]. science-based [CMT^{+}23, HGG22c, STKC22]. science-policy [GSA23, PKH22]. Scientific [MIMS^{+}20, RFPS21, SLMG^{+}21, AGM^{+}22, BD20, CBR21, DGMF^{+}20, GDLC20, GCG^{+}21, HTP^{+}23, Los24, MH24, MKSG20, SCCW21, SG21, WRL23, WSDB20, Yu22, CGHV21, CMD^{+}20, GNJR24]. scientists [DMG^{+}23]. scores [KMJ24]. Scoring [Fox22]. Scotia [CA20]. Scotian [TFF22]. Scotland [Cro21]. Scottish [OC20]. scrappage [WWCS21]. Scuba [AHN^{+}23, LM24, MMO^{+}24, NSN22]. SDG [CMOB^{+}20, Gul20, Liu23]. SDG14b [SPFA^{+}20]. SDGs [Obu20]. SEA [HM22a, ASM^{+}21b, BSY^{+}20, BMM^{+}24, BK23, BBK^{+}21b, Cho22, DMS^{+}24, EMN^{+}24, GÁPSAP^{+}23, Gul23, Gun21, HvdBH23, HLLK22, Kiz24, LYHL23, LXH^{+}22, LNRM23, MÇ23, Ngu23, Niu22, RZ21a, RJAT23, RCH22, Şan20, SR22, TBH^{+}21, Tod22a, WDG^{+}20, AB20, AR22, BEM^{+}20, Bau20, BHKJT23, Boy21, BE20c, CHEC21, CDD^{+}23, CAM^{+}23, CLS20, CJ23, CKY20, CvHF^{+}24, CSHO^{+}20, CDC21b, CL21, De 21, Ege23, EAA^{+}21, EHM^{+}20, EHM^{+}22, FMWW21, GVR20, HFD^{+}20, Har21, HSSM24, HVN21, Hon20, HSHS24, Jen21, JACD20, Kak20, KS22, LMSLMQ21, LDC^{+}20, Li23a, Lin24, LSH21, MXZ23, Mal21, MMB^{+}23, Mat20, MK24, MSGAR20, OPL20, Ora23, OKR^{+}20, ÖB22, PL24b, PMR23, Pra20, PBH^{+}23, QF20, RE21, RE23, RFPS21, Sar23, SVF20, SVHV20, SSA^{+}23, SCBB^{+}22, Set23, Shi23a, SBW^{+}21, SZBA20, SSZ^{+}22, SRY24, SRLLG23]. sea [TZC23, VJL^{+}21, VPB^{+}24, WLCC23, Wil20, Wil21a, Wil22, WDC^{+}20, WSDB20, YWY20, YNZ22, YLSC23, YBW20, ZL20, ZTM^{+}23, dOPM^{+}20, vdGD21, KGW20, ATF^{+}21, ADB^{+}22, BSBD21, BDC23, BvdBvdFK20, BFMS21, BM21, BMM^{+}24, BLY^{+}24, CRP^{+}20, CMC^{+}23, CHB^{+}20, CWKE20, DGG22, DFC^{+}20, DDF^{+}20, DUYEA21, DAHS^{+}23, DD22, EVPMKP22, ES22, FGPPGG22, FGPPGG23, FMCD20, GC22, Guo20, HJPL23, Hsi20, KGL22, Koj23, LSH21, LEVET21, MCRR23, NGT^{+}23, PKH22, PL20a, Pie21, PS21, PWL^{+}23b, PKMT23, RKS^{+}23, SDL^{+}22, SPN^{+}21, SV22, SZBA20, SSZ^{+}22, SW22, SHZC21, Tes20, TH22, VCC^{+}21, VKS23, VPB^{+}24, VGK23, YSEK22, YNZ22, Zha21, ZFGL20, ZLGX20, Zou23, dV21, vHSSK20, HH21, Smi23b, Tan23a]. sea-based [SVF20]. sea-level [Mal21]. sea-use [LDC^{+}20]. sea-uses [SZBA20]. seabass [LFPBD^{+}20]. Seabed [AR22, BHKJT23, GSM21, Sin21, AGM^{+}22, BRM21b, CGHV21, CDC21a, Cra21, Fei21, GG21, HSK^{+}23, Jae21, Mal21, MCR22, PSTW21, SDG^{+}23, SSMG21, Squ21, TML^{+}21, TCD^{+}20, VvSvH^{+}22, WRV21]. seabird [BKW^{+}24]. seabirds [CHC^{+}21, LTRA20, OKH^{+}24]. seabream [LFPBD^{+}20]. Seafarer [BBK21a, Dev22, TAK^{+}22, DG24a, DG24b, NEF23, PGK22, Tan22, ZSWJ23]. Seafarers [VMM21, ASKI20, DG23, Gue23, HVN21, LL24, PK20a, Sam22, SHZ^{+}21, TNY23, YTÇ22, YBW20, ZWJvL20, ZZH21, ZTW23, ZTM^{+}23]. seafloor [JG20]. Seafood [BI21, EMS^{+}23, IMA22, MFF^{+}23, CMNH21, CBD^{+}23, CSH23, CM23, DLYM22, EB23, FFYE21, FGL^{+}21, GM21, GPM23, HPG20, HT23, HRC^{+}24, HWMO20, LGF24, LJ22, MAA^{+}24, NOB22, NGA23, PMK^{+}21, SFB24, SMCW22, SVKP21, SAO^{+}24, TOHK20, VVR^{+}22, WKBS20, WEF^{+}21, WMS21]. seagrass [BMMF22, GABA^{+}23, KWY^{+}22, SGA^{+}21, SWH^{+}23, SLH22]. seagrass-data [SWH^{+}23]. seagrasses [HTP^{+}23]. Seal [EVB21, KL21b, SLK^{+}23]. seals [PBH^{+}23, WJB^{+}20]. Seamount [BMR20, DGIS^{+}22]. seamounts [ALLR^{+}24, BMR20]. Seaport [SKDK23, OCD^{+}20]. seaports [CMOJ^{+}24]. search [BMRM23, CDRS22, CGP23]. Searching [FS21]. Seas [FFGJ22, TCZ21, AFHHD22, BMR20, CMD^{+}20, DBH^{+}22, GR20b, Gul23, Lam21, LL24, MCPS20, MLAS23, MSGAR20, PCA^{+}23, Ros22, WvdMG^{+}21, WCT^{+}20, vdGD21, HGLR22, LTRA20, PKB^{+}23, SSMG21, Smi20]. Seascape [EABA^{+}22, AL20, CACJR^{+}21, LFLF22, SEG^{+}22]. season [MLMD21, MMM^{+}22, MMC^{+}22, OA20, SGPF21]. Seasonal [JKKT20, DLX21]. Seasonality [SGPF21]. seatrout [SKS21]. seawater [SCA^{+}20]. Seaweed [Chr20, FAB^{+}23, ALM^{+}21b, BvdBvdFK20, KCCC^{+}21, LWSS21, LTW23, MM22, MBC^{+}24, RMF^{+}21, SLS^{+}20, SLS^{+}21, TCS^{+}20]. Seaweeds [HSN^{+}21, OR21]. seaworthiness [SM23]. second [CLP21]. section [EAAG21]. sector [AEL^{+}23, AVPAFM^{+}22, BADO20, CBD^{+}23, CM23, DT23, DRCdMCJ21, Fen23, dHJTGO21, HS20, HvdBH23, KMM21, LEVET21, MCR22, MMGAA^{+}21, NVP^{+}23, OCD^{+}20, PCSS^{+}21, SN23, VWJMAJ20, WLXY22, dV21, vLM22, vRB24]. sectors [WM22b]. secure [GWAPRPRU22, dlPdlLLC^{+}22]. Securing [CCX23, FGL^{+}21, NH20]. security [Als20, CN23, CTIH21, DG20, DL22, HS22, Hon20, Jen21, JF24, MF22, MM21b, OYECdOP24, OOGK^{+}24, RBA23, SNK^{+}20, SHZC21, Tel23a, TFH^{+}23, TGHF^{+}22, TCC^{+}22, WCZS21, YH20, ZWC^{+}23, Zou23, Lv20]. Sediment [RB22]. see [BHEE22]. seed [ZWC^{+}23]. Seeing [Ree20]. seeking [SRLLG23]. segmentation [PFM20]. segments [HRS^{+}24]. seine [BLH^{+}20, EMH^{+}23, NSK20, SII22, TYS22]. Seiner [RROP23]. seizures [GÁPSAP^{+}23]. selectable [SKE22]. Selected [Smi22, NNR^{+}21, dKMK21]. selecting [MKB21]. selection [CA21, CSLGFV20, CMWZ24, HPS23, JKV^{+}22, MMO^{+}24, vTHS21]. selective [MKM^{+}20]. selectivity [GMM^{+}20]. self [LSY20, SLB21]. self-governance [LSY20, SLB21]. self-organization [SLB21]. selfishness [AAA^{+}21]. Senegal [CDNS23, hBDF22, hBDDF22, KA21, NSZ^{+}21, SPHO23]. sensation [SRLLG23]. sense [CD24]. sensing [ASM^{+}21b, LWSS21]. Sensitive [Cho22, RJAT23]. sensitivity [KMJ24]. sensors [TZM^{+}22]. sentiment [ZLWY22]. sentinels [WtDD^{+}21]. separation [GHK20, RQB^{+}20]. September [Ano20v, Ano21y, Ano22x, Ano23x]. sequence [LR20]. sequential [CVA22a, GCB^{+}23, NGP^{+}21]. sequestration [TMTC20]. Sergia [LSY20]. series [SFGEPC22]. serious [Kuł22]. service [HPPH22, SL20, YCBW21]. Services [RRR^{+}23, CDD^{+}23, CWKE20, DJC^{+}23, DA23, FRVB22, GPG^{+}21, HPC^{+}23, JZW21, LW22, LRC24, MKAF21, MFC^{+}24, MWWS21, POM24, QK24, RFG24, RGH^{+}22, SAO^{+}21, ZZH21, ZCG^{+}23, dSOW^{+}21, vSHHT20, vTHS21]. serving [IGS^{+}23]. session [MTN23]. set [LXW^{+}21]. setback [vRB24]. Setting [TML^{+}21]. settings [DJC^{+}23]. Settlement [Shi20a, HC22, Li23b]. Settlements [PK20b]. seven [ZV20]. severe [DDAB^{+}23]. severity [BMK^{+}22]. Sexual [PK20a]. Seychelles [BC20a, BCN23, COSS21, CJ21, DFTK20]. SFPA [KA21]. SFPAs [THP^{+}22b]. shad [INW20]. shadow [JL21c]. Shake [TMDN23]. Shandong [QS21]. Shanghai [ZCSK23]. Shanyuan [HM21]. shape [BDMW22, RMU^{+}22]. shaped [SWF^{+}24]. shapes [SLB21]. share [MCKZ21, STIS20]. Shared [MZ24, BM21, BLM24, CDG^{+}21, Fox22, Han20, Wan23a, XT23, ZCZ20]. Sharing [BL22, CMS20, SII22, BFMS21, BTE21, GSB20, JF20, WLCC23]. Shark [Jon21, Los23, MCW^{+}22a, PMM^{+}21, BMF20, BdLAdC^{+}20, BL22, BBHN20, CLB^{+}20, CRJ22, DSL20, GMGLV21, GMGLV22, LDL21, Los24, MAMCHP^{+}20, MT20, NUD^{+}23, OT23, PGGA22, SMHPMO^{+}21, SBDC24, SMC^{+}21, SAB^{+}20, TMM^{+}23, HACCM22]. shark-bite [SAB^{+}20]. shark-diving [GMGLV21, GMGLV22]. shark-human [SMC^{+}21]. shark-marine [SBDC24]. Sharks [LM24, PMFR^{+}22, RM21a, Alm21a, AMLC^{+}23, ALLR^{+}24, BBHN20, CLB^{+}23, GBB^{+}21, HMG^{+}21a, HACCM22, KL22, LDL21, MH20b, NGP^{+}21, SSA^{+}23, Sel20, SON21, TJS22]. shelf [Dun24, KS22, MFRQ23, WCF^{+}21, dS21, dS22a, AML^{+}21, LFLF22]. shelled [AOO^{+}23b]. shellfish [BHB^{+}24, CWKE20, Gir22, LDL^{+}23, MM22, MWWS21, MFF^{+}23]. shift [MCKZ21, MSZ^{+}23]. shift-share [MCKZ21]. shifted [Tel23b]. Shifting [TOAPJ21, YFD^{+}21, CCB^{+}21, CD24, GTMP20, KSB^{+}23a, VLS^{+}23]. shifts [CC21, CSGC22, NGA23, VGC^{+}22]. shimmering [Gul23]. Ship [BBK21a, COJS^{+}22, DN20, MCW^{+}22b, WWCS21, BVT^{+}21, CLL^{+}22b, Dun24, Gun21, HSH^{+}21, HP20, KDH^{+}21, KDS22, KMM21, LMT21, MÇ23, MTN23, NK21, NNR^{+}21, Öİ24, PKR^{+}22, PYVF24, RZ21a, SNP21, Ste23, GHK20]. ship-derived [Öİ24]. ship-source [KDH^{+}21]. Shipping [Tod22b, Wan23c, ZD20b, BH23a, BHEE22, BTV^{+}22, Cha22, Che21b, DG24b, Gun21, GGL22, GCV^{+}20, HRS^{+}24, JHL21, JD20, KKDK20, KLS22, Kiz24, Kon20, KABT21, LrS22, MF23, MCTN23, NCD23, NNR^{+}21, PZPB21, QLZ^{+}24, RZ21a, SGB21, Sar23, SSK23, SPN^{+}21, Ton22a, TC23, THS^{+}24, VGK23, WDG^{+}20, WA24, WCT^{+}20, WLXY22, XSCL21, ZWJvL20, ZXGL21, ZLWY22, ZLY23, ZNA^{+}23, vLM22]. Ships [Laj21, ASKI20, AKL^{+}22, CCX23, CAC23, DG23, DG24a, HVN21, VMV^{+}24, XZ23, CCX23, DBZ24]. Shipyard [CA21]. Shoal [FMLMM23]. Shocks [SMKS22, Gor21, SPMN23]. shopping [NT20]. shore [BHCL23, FHBS23, MTN23]. shore-based [FHBS23]. Shoreline [BD20, BS22, CTC^{+}23]. shorescapes [BvdH23]. short [UISM^{+}23]. short-lived [UISM^{+}23]. shortcomings [SS23]. shortfin [DSL20, Kai21]. Should [ABI24, CMOJ^{+}24]. shows [HLCM21]. Shrimp [AFMdSdTdCC22, BDCJ22, CPBO20, DME23, GABSD21, HCT^{+}21, KTI^{+}24, LFLF22, MVR21a, May22, PAK22, TCS^{+}20]. shrimp-tilapia-seaweed [TCS^{+}20]. Sian [MMSB^{+}20]. Sicilian [DPB^{+}23]. side [GLT23]. side-effects [GLT23]. sides [dOPM^{+}20]. SIDS [CTIH21]. Sierra [OOGS20a, OOGS20b, OOSG21]. signals [WKO^{+}22]. Signe [Niu22]. significance [Wan23a]. significant [BMR20]. Silk [HH21, ZWG^{+}21, SF22]. similarities [BMB^{+}21]. simmering [TMDN23]. SIMNORAT [GBCNCL^{+}21]. Simulating [RHST21]. simulation [CDHY22, JHL21, PBT^{+}20, ZLGX20]. simulation-based [PBT^{+}20]. since [GMVM23, PL20a, SWXM21]. Singapore [CJCS^{+}22, ZSLY20]. single [AAA^{+}24, AWBC20, Fra23]. single-use [AAA^{+}24, AWBC20]. sink [HZ23]. sinks [WL22b]. site [CA21, JKV^{+}22, vTHS21, CSLGFV20]. situation [QLZ^{+}24, CLS20]. situational [DAN^{+}23]. situations [CPBO20, RE21]. Size [Mag23, OCZ^{+}22, SKE22, TPG^{+}23, YU20, GMM^{+}20]. Skagerrak [NSK20]. skate [GPH^{+}21]. Skepticism [Pol23]. skill [GCP^{+}20]. skills [ASKI20]. skipjack [PWL^{+}23b]. sky [FTG^{+}23]. slavery [VM21, WKBS20]. sleep [vLPBK21]. sleepiness [vLPBK21]. Sleuthing [KDM^{+}20]. slow [cL22, SVF20]. Sluggish [Kar22b]. Small [ACV^{+}21, APF22, Art20, BMB^{+}24, CPBO20, CMC^{+}23, hBDDF22, GSG^{+}24, MF22, Moh23, SPFA^{+}20, TTA^{+}20, TFH^{+}23, YCL22a, AGN^{+}20, ÁCT20, AFMdSdTdCC22, AONF21, AH21, BCBS22, BPA^{+}22, BDM20, BJGG^{+}24, BBS^{+}21, BSI^{+}21, BCJ21, BASM22, BGMCG24, CC21, COJE24, CAVJ^{+}20, CCKK22, COSS21, DAN^{+}23, DBB^{+}22, DDGdS^{+}21, FGPPPV21, FMCD20, FS22, GLADGGVL23, GPTASM23, GMRSP23, GNMGWM21, GMCMJA^{+}21, GFL^{+}21, HLW^{+}20, Han21, HTP^{+}23, HBP^{+}21, JAdS^{+}20, JKKT20, KMS20, KB20, KWMM20a, KWMM20b, KJS^{+}22, LRN^{+}22, LBWS23, LJRHH21, LEERRM^{+}21, Lov23, LLS^{+}20, MGS21, MMK^{+}23, ML21, MMB^{+}23, MM23, Mil22, MIH^{+}24, MMGAA^{+}21, NRYG20, NSN22, NAS^{+}22, OCZ^{+}22, OYKBA22, OOGS20b, OCD^{+}23, OMC^{+}22, OA20, PSPR20, PGSFRMHH24, PL20b, PALB20, PMC24, PKS20, PFWM23, RTMCVMDE22, RDLA^{+}20, SIW^{+}21, SLB21, SCJ22, SSZ^{+}24, SdCMN^{+}22]. small [SS21b, SSP^{+}21a, SPA^{+}21, SGL^{+}21, SRH23, SSP^{+}21b, THP22a, THT^{+}23, TGHF^{+}22, TPG^{+}23, UOO20, VGC^{+}22, VGAH^{+}21, WKO^{+}22, ZJ20, dlPdlLLC^{+}22, CTIH21, NH20, NCE21]. small-island [NRYG20]. Small-Scale [CMC^{+}23, ACV^{+}21, Art20, BMB^{+}24, CPBO20, GSG^{+}24, MF22, SPFA^{+}20, TTA^{+}20, YCL22a, ÁCT20, AFMdSdTdCC22, AH21, BPA^{+}22, BDM20, BJGG^{+}24, BBS^{+}21, BSI^{+}21, BASM22, CAVJ^{+}20, CCKK22, DAN^{+}23, DDGdS^{+}21, FS22, GLADGGVL23, GPTASM23, GMRSP23, GMCMJA^{+}21, GFL^{+}21, HLW^{+}20, Han21, HTP^{+}23, HBP^{+}21, JAdS^{+}20, JKKT20, KMS20, KWMM20a, KWMM20b, KJS^{+}22, LRN^{+}22, LBWS23, LJRHH21, LEERRM^{+}21, Lov23, LLS^{+}20, MMK^{+}23, ML21, MMB^{+}23, MM23, Mil22, MIH^{+}24, MMGAA^{+}21, NAS^{+}22, OCZ^{+}22, OYKBA22, OOGS20b, OCD^{+}23, OMC^{+}22, OA20, PSPR20, PGSFRMHH24, PL20b, PALB20, PMC24, PKS20, RTMCVMDE22, RDLA^{+}20, SIW^{+}21, SLB21, SCJ22, SSZ^{+}24, SdCMN^{+}22, SS21b, SSP^{+}21a, SPA^{+}21, SGL^{+}21, SRH23, SSP^{+}21b, TPG^{+}23, UOO20, VGC^{+}22, VGAH^{+}21, WKO^{+}22, ZJ20, dlPdlLLC^{+}22, NCE21, NH20]. smart [OOGK^{+}24, THP22a]. snapper [QTPL21, RYS^{+}21]. snow [Gor21]. Social [FHBS23, HFD^{+}20, KFM^{+}20, MVSP^{+}21, NGP^{+}21, PN20, RDLA^{+}20, Sol23a, TCT21, VGC^{+}22, AE21, BMMF22, CJ23, CCCC21, CLB^{+}23, DUYEA21, Doe21, ES22, FFYE21, FAGS21, GPM23, GMRSP23, GABA^{+}23, GTMP20, HRR^{+}20, HHEJB20, JDF^{+}23, JMB^{+}20, dOJCSLdSB21, KBD^{+}20, LZZ22, Liu22, LMMT23, MCW^{+}22a, MF20, MTC^{+}23, ONA21, ORSB22, PFG^{+}21, PWL^{+}23a, PV23, RBA23, RFG24, RC22, SOO20, SCD21, SCF^{+}22, SOLN20, SMCW22, Squ20, SLS^{+}20, STF21, TMM^{+}23, UOO20, XXLW24, ZCG^{+}23, ZLWY22, ZCA^{+}23, dlPdlLLC^{+}22, PZPB21, Sol23b, WGP^{+}21]. Social-ecological [RDLA^{+}20, VGC^{+}22, FAGS21, GABA^{+}23, HRR^{+}20, PWL^{+}23a, SOO20]. societal [MLVT21, RSFvH^{+}22]. society [HRR^{+}20]. Socio [BDCJ22, PDJ22, SSA^{+}23, SCBB^{+}22, ZNA^{+}23, CVA22b, VBd^{+}21, FMLMM23, JTR^{+}20, PBS^{+}23, RGK^{+}20, SRR^{+}21, SDL^{+}22, TMS^{+}22]. Socio-cultural [SSA^{+}23]. socio-ecological [CVA22b, JTR^{+}20, PBS^{+}23, SRR^{+}21]. Socio-economic [BDCJ22, PDJ22, SCBB^{+}22, ZNA^{+}23, FMLMM23, RGK^{+}20, TMS^{+}22]. socio-environmental [VBd^{+}21]. socio-political [SDL^{+}22]. sociocultural [CG22, WNMS^{+}24]. Socioecology [CAHR^{+}24]. Socioeconomic [CDNS23, BSY^{+}20, CDG^{+}21, CM23, HWT23, JAdS^{+}20, MLVT21, MFC^{+}24, WNMS^{+}24, ZCCQ22a]. Softcover [Sta23a]. soil [SOK^{+}22]. sole [MKM^{+}20]. Solution [hC24, GOH^{+}20]. Solutions [Niu22, Ada22, AGS20, ADR^{+}21, BN20, BBWC21, CJ22, DQJ23, DBZ24, Gug24, HHD20, PZZ^{+}21, RSFvH^{+}22, SNP21, SBJF^{+}21, WLW^{+}22]. solve [Fen23]. solved [Pie21]. solving [PVH^{+}21]. Some [Zhu23, ZY23, VMC^{+}21]. Songhees [BMB^{+}21]. Soufriere [TJ21a]. Sound [KDH^{+}21, KDM^{+}20, CTC^{+}23, SOP^{+}22]. source [AAA^{+}24, BHdMCM^{+}21, KDH^{+}21, SGLF^{+}23]. sources [GAD^{+}20, Gug24, KTI^{+}24, MCTN23, SSK23]. South [BBGG22, FMLMM23, HGLR22, MCW^{+}22a, SHAC23, YMG23, ADR^{+}21, CACJR^{+}21, CSHO^{+}20, Mag23, WWG^{+}23, WDC^{+}20, ZL20, Abo23, AM22, BBHN20, CHB^{+}20, CDG^{+}21, CKY20, CKY22, CES20, CHC^{+}21, DPG^{+}22, GC22, Guo20, HWM22, HJPL23, Han21, Hsi20, KFB^{+}20, KNY20, KNY22, Kim23, KSS23, MFT^{+}22, Mar20a, MGDJ22, OAM20, OAM21, PL24a, PKC23, PKB^{+}23, PBG23, PL20a, PGP22, PK20b, PWL^{+}23b, PBH^{+}23, SRvNN21, SRR^{+}21, SON21, SBDC24, dOSTB^{+}20, SBT^{+}23, SS21b, SSP^{+}21a, SW22, SHZC21, Tra23a, WLCG23, Zha21, ZFGL20, ZLGX20, Zou23, dOPM^{+}20]. south-east [ADR^{+}21]. South-Eastern [YMG23, WWG^{+}23]. Southeast [MGDJ22, LTJ^{+}21, CMWN^{+}23]. Southern [Isr22, MH20b, RCM^{+}22, CC21, DSQ^{+}22, FMCT^{+}21, cL22, Mar21b, MJMAA21, MIH^{+}24, PBH^{+}23, PKMT23, SCD21, SVP^{+}23, SHAC23, TDC21, TGM22, AFMdSdTdCC22, BdLAdC^{+}20, CTH^{+}21, Gar20, GB21, KTI^{+}24, MMS24, NEH^{+}23, SBFHD^{+}23, SPA^{+}21, Tra23b, TMTC20, Wen20]. Southwest [BOOO23]. Southwestern [CMI^{+}22, ALLR^{+}24, RRR^{+}23, PCG^{+}23]. Sovereignty [Gug24, BMF20, Mar21b, SW22, WD22]. Sowing [MdCLL^{+}23]. Space [LP21, APF22, QZZX20, QZZ20, Qi24]. Spain [BSBD21, CRCP^{+}23, dlFGFFR^{+}20, HPC^{+}23, dlÁPAFGO20, SPG20, TGM22, ASV^{+}21, BDLO23, BISSBdlMM22, FGPPPV21, GLADGGVL23, GSR21, Gar22, HSGNMGC21, MJMAA21, RS24, Sol23a, Tei23, SCD21]. span [Dun23]. Spanish [Álv23, BLH^{+}20, DL21, DSL20, FGPPGG22, FGPPGG23, FR23, dHJTGO21, HRMPCRRA22, JG23, Oan22, Pat22, RAFLRP21]. Spatial [ANS^{+}20, AMK^{+}21, AGN^{+}20, BRP^{+}23, CLS^{+}23, COJZ^{+}21, DGIS^{+}22, Ehl21b, FLC^{+}23, FGK^{+}20, GBCNCL^{+}21, GASLCP20, KMT^{+}21, LYL^{+}22, PDB^{+}23, RBB^{+}24, SLL^{+}24, SV22, SSMG21, SRSGMGG21a, SRSGMGG21b, TZM^{+}22, TMS^{+}22, APV^{+}20, AFMdSdTdCC22, ACG21, AHM22, BN20, CPV^{+}21, CVCV^{+}22, CSQ^{+}24, CSLGFV20, Cat20, CTCD21, CMD^{+}20, CHB23, DBH^{+}21, DGMF^{+}20, Ehl21a, EVPMKP22, FTW^{+}23, FGC21, GPA^{+}22, GGP^{+}22, GPG^{+}21, GSR21, GAD^{+}24, GABA^{+}22, GCC^{+}21b, GGL22, HLW^{+}23, JDF^{+}23, JKV^{+}22, JMB^{+}20, KvTNL20, KVL^{+}23, LKB^{+}23, LJ20, LJ23, LEVET21, MFG^{+}22, Mah22, MBORGN24, NM23, PHF20, PBT^{+}20, PGA^{+}20, PS21, PKS20, PKMT23, RSAP22, RLS20, RDT^{+}23, RDLA^{+}20, SGB21, SAA^{+}21, SBP24, SZBA20, SSZ^{+}22, SPT^{+}20, TJ21b, VPRR^{+}21, WNM22, WGJ^{+}21, ZK21, ZJ22, ZBF^{+}22, dV21, vdBSvdV^{+}23, vdGD21, CHEC21, GSGOP^{+}21]. Spatial-temporal [LYL^{+}22].

Spatio [FARBV^{+}24, RHST21, WL22b]. Spatio-temporal [FARBV^{+}24, RHST21, WL22b]. Spatiotemporal [GVH^{+}23, JK22]. spawning [SGPF21]. spearfishermen [MJ23]. spearfishing [MJMAA22]. special [EAAG21, Ehl21a, GAD^{+}23, LBWS23, PZPB21, SNS23]. specialization [ÁCT20]. Species [BADO20, AC21, BBB23, BdLAdC^{+}20, BUM^{+}24, CHEC21, COJE24, CMI^{+}22, CAHR^{+}24, DUYEA21, GÁPSAP^{+}23, GMM^{+}20, GTTM21, JSW22b, Kul23, MN23, MWM^{+}23b, OTK^{+}21, OVA22, SCSR^{+}22, SCSR^{+}23, SSDQ20, SMW^{+}21, SPN^{+}21, TBPC20, WCM^{+}24]. species-at-risk [JSW22b]. Species-level [BADO20]. species-specific [BBB23]. specific [BBB23, COJE24, GHW^{+}22]. spectrum [PFG^{+}21]. speed [GHK20, MPL^{+}23]. Spermonde [HF20]. Sphyrna [ALLR^{+}24]. spill [JHHG23, MFT^{+}22, PKH22, SRvNN21, dOSTB^{+}20, SBT^{+}23]. spillover [JK22]. spills [SGFM^{+}21]. spin [MB22]. spin-offs [MB22]. spiny [DMS^{+}24, Doe21, PLOG24]. Spiraling [RC22]. Spitsbergen [Tod20]. Splitting [Mat20]. sponges [MSGAR20]. sport [KWF^{+}21]. spot [CBR21]. spotted [SKS21, Squ20]. spp [APV^{+}20, MH20b, YNZ22]. Spratlys [Mar21b]. spring [GS21]. Springer [Rub24, Sta23a]. Sputc [BMM^{+}20]. square [GMM^{+}20]. squid [GWAPRPRU22, PLOG24, TM23a]. squid-related [TM23a]. Squids [MLAS23]. Squilla [BMGDS^{+}20]. Sri [JAH^{+}24, NEH^{+}23, AEL^{+}23, CLB^{+}20, CLB^{+}23, GRS^{+}23, Kul23, Pra20, PDJ22]. SSF [BMB^{+}24]. St [CJ21]. St. [GMRSP23]. Stability [JZW21, GC22, LGF24]. stage [HHD20]. stages [Pol23, YC21b]. Stakeholder [AMLC^{+}23, BOOO23, CdLN^{+}21, CR23, FOH^{+}20, FMLMM23, GNJR24, HDAPG^{+}24, MKSG20, OCD^{+}20, RYS^{+}21, WLCG23, BDC23, BCB20, BDM21, CDB20, DSL^{+}21, Doe21, EVPMKP22, FTW^{+}23, GRA^{+}22, HvdBH23, HPPH22, dOJCSLdSB21, KMG^{+}20, MHEW20, OMC^{+}22, Öİ24, POM24, RA20, RP23a, RP23b, STS^{+}23, ZLWY22, PVH^{+}21]. Stakeholder-defined [MKSG20]. Stakeholder-government [CdLN^{+}21]. stakeholders [CLW^{+}21, DSL20, FDGM24, HS20, HC21, JCMD20, MBB^{+}22, PMFR^{+}22, SWH^{+}23, XTY22, ZK21, ZXGL21, GTA^{+}20]. stalked [GRA^{+}22]. Standard [JC23, SK20a, BBBS22, KMJ24]. standards [BGMCG24, GLS^{+}21, HPG20, Sha22, SGA^{+}21]. State [AP23, BJH21, HMY^{+}23, Kar22a, RBA23, Sam22, WL22a, XWL^{+}20, BPA^{+}22, CC23, CLL20, COSS21, DBZ24, GBC^{+}20, Har22, JAdS^{+}20, KS22, MS21, RCT20, Şan20, SFB24, SRY24, WdSRY23, YT20, LWZ23, Shi22, Wil21b]. state-led [WdSRY23]. statement [Dbo22]. States [CTIH21, Ngu23, Hon20, HLO^{+}21, KMS20, Kar22a, NSN22, Puh23, TN21, WdSRY23, dS21, CNB^{+}23, Col20, FBC^{+}23, HWU^{+}23, JFB^{+}24, Moh23, MS20, NGA23, PWL^{+}23a, SBT21, SRH23, WW21b, Wil21a]. static [CMD^{+}20, DFF23, FARBV^{+}24, FWL^{+}22]. statistical [RZ21a]. Status [KSV^{+}23, MPD^{+}23, Tho22, WLRH^{+}20, ASM21a, BGK^{+}22, BHKJT23, BRM21b, BRR^{+}22, CZW20, CES20, DFC^{+}20, DUYEA21, FZYF23, Fox22, FMCT^{+}21, GCM^{+}23, HF20, HGH^{+}20, Kas21, Kul23, LC20a, LYHL23, MIH^{+}24, NGT^{+}23, OKR^{+}20, PCG^{+}23, Pie21, Puh23, TYS24, WNDZ23, WM22b, Zhu22, ZMM^{+}23, GKP^{+}22, Wan23b]. statute [Shi23a]. Step [GCMT20, APF22, CHB^{+}20, COJS^{+}22, GJY23, LC23, NCE21, SPFA^{+}20, SCA^{+}22]. steps [YK21]. Stewards [Har21]. Stewardship [BI21, JAK^{+}23, LJ21, MPH^{+}23, NRF21, SK20a, TOL^{+}22, GBSC22, HBCA^{+}24, HDHM^{+}20]. still [BCBS22, HWT23, Sin21]. Stock [CCK^{+}21, BBB23, CDG^{+}21, CD24, HHH^{+}23]. Stock-recruitment-environment [CCK^{+}21]. stocks [BLM24, Fox22, Han20, HGH^{+}20, OOSG21]. Stopping [SGLF^{+}23]. storage [QTBL23]. Storm [HBR21, YSWW21]. story [MB22, NW20]. Strait [Tod22b, STS^{+}23]. straits [Lot22, Tod22a, POO23]. stranded [BBBS22]. stranding [OWCFDC21, SPJB22]. Strategic [BXV20, CVCV^{+}22, Jae21, KV20, PS21, TML^{+}21, AHBAKH24, OEV23, YZH22, HM22a]. Strategies [ABME22, COJE24, DDF^{+}20, WYH21, AFMdSdTdCC22, BB21, CMS20, CDNS23, CCK^{+}21, CBAPM20, DSB21, DSL20, FAC^{+}23, GML^{+}24, HSH^{+}21, HM22b, MM21b, Nal24, Okş23, PCCOS24, Rus20, RHL20, SIW^{+}21, SLS^{+}21, WNBRSG22, XTY22, ZTZ^{+}21, ZCSK23]. Strategy [ANS^{+}20, ASGS23, Ber23, GKP^{+}22, MPD^{+}23, MFRQ23, Mor20, PKB^{+}21, PDB^{+}23, VRBC^{+}21, VBH^{+}23, VJO23, BÖ24, DFC^{+}20, GW21, HJPL23, Pre21, Puh23, RKR20, TCT21, VSV^{+}23, VPRR^{+}21, WLQD21, YZG23, ZD20a, ZMZ23, ZWC^{+}23, MCW^{+}22a]. stream [CA20, CKY22]. Strengthening [BMM^{+}20, CWR20, GSM21, GOH^{+}20, VSDR^{+}22, HJ21]. stressor [WGR^{+}22]. stressors [Dun21b, MNGM24, OCT^{+}22, PCCOS24, SCLM^{+}21]. stricter [BH23a]. strike [MCW^{+}22b]. strikes [FMWW21]. striped [BSBD21]. Strong [AEJ24, AR23a, AR23b, Jen23, NHA23]. Structural [NHA23, ZZD^{+}23]. structure [CQ20, SZ20, SAO^{+}24, SKL23, TMTC20, YWYL22]. structured [ÖB22]. structures [Gal21, MCRR23, SVF20, VTRS20, ZCA^{+}23]. struggle [Smi22, SCF^{+}22]. Stuck [DMNT20]. students [DDD22, TSWC23]. Studies [HH21, AANM^{+}22, BMMF22, CJ23, CES20, GPA^{+}22, GCC^{+}21b, JL21c, LLLF22, SRSGMGG21b, ZWJvL20]. Study [ZLGX20, Åm21, APV^{+}20, ASV^{+}21, AED20, BvdBvdFK20, BMGDS^{+}20, BJGG^{+}24, CWB^{+}20, CMC^{+}23, CCCC21, CSHO^{+}20, CVA22b, dFVCGdS21, Cro21, COJZ^{+}21, CMOJ^{+}24, DRCdMCJ21, DSQ^{+}22, DXS^{+}21, DMS^{+}24, DG20, FGPPPV21, Fro21, FMWW21, FMS^{+}22, GRS^{+}23, GLADGGVL23, GAA^{+}20, Han21, HLW^{+}23, HRR^{+}20, HHB^{+}21, HdM23, HZ23, HRS^{+}24, HVN21, Hor23, HCT^{+}21, IBC^{+}20, JAH^{+}24, JAdS^{+}20, JMB^{+}20, KMG^{+}20, KGP^{+}22, KABT21, LDAC20, cL22, LWG23, LTJ^{+}23, LLSY22, LJ21, LSY20, MLJ^{+}22, MBDG23, MMSB^{+}20, MGF20, MBL24, NSZ^{+}21, NCD23, NHD23, NNRS24, NEH^{+}23, OCD^{+}20, Okş23, PCA21, PKC23, PC20, PPK23, Qi24, RBB^{+}24, RP23a, RP23b, RDM^{+}20, RNGRC22, SMHPMO^{+}21, SVV24, SSZ^{+}24, SMKS22, SMC^{+}21, SKL^{+}20, SRH^{+}22, TM23a, THGR^{+}20, TSH22b, TLT^{+}22, TCT21, TMT20, TPG^{+}23, UISM^{+}23, WD23, WJB^{+}20, WLQD21, WSD21, WFG^{+}22b, WDC^{+}20, WFG22c, WLW^{+}22, Wu22]. study [WWC23, YFD^{+}21, YS21, YTÇ22, YC21b, ZJ20, ZCCQ22b, ZTM^{+}23, ZCSK23, ZCS^{+}23, Zhu22, vRB24]. sturgeon [BUM^{+}24]. sub [KKA^{+}21, UOO20, WCT^{+}20]. sub-Arctic [KKA^{+}21, WCT^{+}20]. sub-Saharan [UOO20]. Subarea [CHC^{+}21]. subjected [CBAPM20]. submarine [KS22]. submissions [BL23a]. subregional [VRBC^{+}21]. subsidies [BCBS22, BHKK23, CMOB^{+}20, CMSN^{+}22, FR23, For22, KGP^{+}20, LBL21, MVSP^{+}21, SC22, SS21a, SSV^{+}23, VSD^{+}23, VSDR^{+}22, ALLS23]. subsidy [DGG22, SS21a, WLQD21]. subsistence [Gal21, QCM^{+}20]. substitution [SKS21]. subsystems [LJL22]. success [DKW20, MB22, SEG^{+}22, ZSLY20]. successes} [CMS^{+}21]. successful [EdLP^{+}21, Geh23, Rus20, VvSvH^{+}22]. sufficient [CLD^{+}24, KYMW21]. sugar [HKKY22]. suggestions [KOBS21]. suitable [OKH^{+}24, RZ21b]. sulfide [JG20, VCC^{+}20]. sulfur [WZZC21, ZY21]. sulphur [Kon20, Tes20]. Sulu [MIH^{+}24]. Sunda [EABA^{+}22]. sunfishes [PNS^{+}23]. sunken [Lin20]. Sunscreen [Lev20, ALML23]. sunscreens [Lev20]. super [MT20]. super-shark [MT20]. supervision [WXLD21]. supply [HCSBM^{+}23, HRC^{+}24, How20, HP20, KSS23, MMM^{+}22, Tan22, TCC^{+}22, VVR^{+}22, WKBS20, WEF^{+}21, YZH22, ZCSK23]. support [AGK22, BST23, CWB^{+}20, CTH^{+}21, CVH^{+}20, CDB20, DKW20, hBDF22, DB20, FDGM24, Fla21, GPG^{+}21, GLS^{+}21, HS20, JKV^{+}22, KWMM20a, KVL^{+}23, LTW23, LDL21, LMC^{+}23a, MLMD21, MMC^{+}22, MF23, MM20, PGK22, RLS20, RSS^{+}20, RBNE20, RKS^{+}23, THP22a, TEA^{+}22, TFMT^{+}23, ZLZ21]. Supporting [LHV^{+}20, PWL^{+}23b, SSZ^{+}22, BGMCG24, KGP^{+}22, Koe22, MH24]. supportive [TAK^{+}22]. surf [MMS24]. Surface [DBZ24, Squ21, BK23]. surfer [BL22]. surge [YSWW21]. Surigao [MMK^{+}23, MNGM24]. Suriname [DB20]. surrounding [MSO23]. surveillance [AIG^{+}20, CWR20, DGMF^{+}20, SAB^{+}20]. survey [CC21, DDD22, KCCC^{+}21, KOBS21, KP24, MM21b, MSGAR20, NK21, DNR21]. survey-based [CC21]. surveys [CdLN^{+}21, GVH^{+}23, PBD^{+}23, SSOP^{+}21]. Survival [NSK20, OYKBA22]. suspects [TV21]. suspended [TMDN23]. Sustain [HCS21, MJ23, SWHT22]. Sustainability [ASM21a, GV22, HRMPCRRA22, LCAX21, NOL^{+}23, STS^{+}23, WKH22, AANM^{+}22, AVPAFM^{+}22, AMR21a, AH21, BMF20, BGLTC21, BDR20, BBO21, Cha24, Doe21, FOH^{+}20, FMCT^{+}21, GdACT20, GCV^{+}21, HPG20, HHH^{+}23, KgJqY^{+}24, Kob22, MMK^{+}23, MDG^{+}23, MAMCHP^{+}20, MM20, NRF21, OR21, OI23, OGS24, PFG^{+}21, PMM^{+}21, RC22, Sah22, TUBA22, TOHK20, UISM^{+}23, VCC^{+}21, ZLWY22, Als20]. Sustainable [BH23b, BCN23, CTIH21, KA21, LHTR23, LBL21, LJ22, MZ24, NH20, Ton22b, ZLY23, vSHHT20, AAA^{+}24, AAA^{+}21, AO20, Aqu23, BE20a, BCB20, BB20, BDM21, BHB^{+}24, BRR^{+}22, CBO^{+}21, CMC^{+}23, CJL^{+}23, Dbo22, DFF23, EMN^{+}24, EAA^{+}21, FGL^{+}21, FMS^{+}22, GNJR24, GLAVL21, Gir22, GMG^{+}23, GMRSP23, Gug22, HS20, HPH22, HT23, HWMO20, HSN^{+}21, HC21, HP23, HM21, HBL^{+}21, IKB21, IMA22, JV20, KWF^{+}21, KS23, KFP20, Koe22, LZZ22, LLL^{+}24, MZ22, MSBA22, McC21, MMNC22, MBC^{+}24, MCTN23, NSN22, OCD^{+}23, PSD23, PN20, PKB^{+}23, Pen22, Pri22, PŚ22, RJ21, Ren21a, RDM^{+}20, SdACdSLdAS24, SHP^{+}21, SHZ^{+}21, SGL^{+}21, TOAPJ21, Tel23a, THP22a, Tho22, Tra23b, WP23, ZJ20, ZWG^{+}21, ZXGL21, Zha22, dlPdlLLC^{+}22]. sustaining [FWGM24, NSPRSS20]. Svalbard [Tod20, Liu21, NT20]. Svendsen [Cha20b]. Swedish [DME23, PYVF24, WJB^{+}20]. swimmer [OCZ^{+}22]. Switzerland [Sta23a]. sword [Gug24]. swordfish [NOL^{+}23]. SWOT [AHBAKH24]. SWOT-AHP [AHBAKH24]. symbiosis [WCZS21]. symbolism [BST23]. Symes [Smi23a]. Symposium [SWGZ22]. sync [WEF^{+}21]. Synergetic [GW21]. synergies [ANS^{+}20, BPA^{+}22, PT23]. Synergistic [GGP^{+}22, ZLH21]. synergy [CCB^{+}21, LJL22]. synthesis [BB21, HMG^{+}21a, TFMT^{+}23, WNMS^{+}24]. synthesizes [DBS22a]. system [ADB^{+}22, CLL22a, CDHY22, CJL^{+}23, DLX21, DSJ^{+}23, EMN^{+}24, FLC^{+}23, FFE20, FDF^{+}21, GDLC20, HYN^{+}23, HCW^{+}23, HJPL23, HLM^{+}20, JKV^{+}22, JF22, KL21a, KABT21, LDC^{+}20, LXH^{+}22, PVH^{+}21, PBS^{+}23, Qi22, Qi24, RMU^{+}22, SII22, SK20a, SIW^{+}21, SSZ^{+}24, SdCMN^{+}22, SCA^{+}22, SSP^{+}21b, THP22a, YLSC23, AM22, Gar20]. systematic [CJL^{+}23, EdLP^{+}21, HLLK22, LRC24, QK24, TSH^{+}22a]. Systems [dSOW^{+}21, BISSBdlMM22, BCJ21, BSK^{+}20, FTM^{+}22, FAGS21, GM21, GPM23, GTMP20, HWT23, HRR^{+}20, Her22, JL21d, LLS^{+}20, Mba22, MdCLL^{+}23, MEG^{+}21, PWL^{+}23a, RDLA^{+}20, SOO20, Sta23a, STKC22, SN23, STF21, VGS^{+}20, GMG^{+}23].

T [HH21]. table [TM23a]. TAC [DSJ^{+}23, ZLD^{+}21, MH20a]. tacit [Zou23]. tackle [BR22]. Tackling [GE23, MG22]. tactical [CPR21, CBR21]. tag [MFG^{+}23]. Tagus [SCA^{+}22]. Taiwan [CS21, CCCC21, CJ22, HC21, HP23, HM21, HCS21, HCT^{+}21, HWY22, Kao23, KWG23, KOCL23, LHW21, LTSL22, cL22, LTJ^{+}23, Liu22, LLL^{+}24, LSY20, Shi20b, Tse21, TK22, TK23, VM21, WYH21, WWC23, YL21]. Taiwanese [LLCH21, LLSY22, SCGC20, TSWC23, YCL22a]. take [BBB^{+}20, NOB22, QTPL21, QTBL23, WBPG20]. Taking [CLD^{+}24, MZ24, TCZ21, VKN20, Wan23a, LC20b]. tale [PL24b]. Tallurutiup [KDH^{+}21]. Tam [Han21]. Tangerang [MKAF21]. tangled [OPL20]. Tanzania [MFC^{+}24, YMG23]. Tanzanian [IOHC22]. Target [GA20, CHEC21, DUYEA21, HGH^{+}20, SCSR^{+}22, SCSR^{+}23]. targeting [BLH^{+}20, BDCJ22, GBC^{+}20]. targets [SPCCvT21, VvSvH^{+}22, Mor20]. tarpon [GCM^{+}23]. task [MBL24, RE21]. tasks [Dun23]. Tasmania [POM24]. Tasting [CSH23]. Taurua [MRD^{+}20]. Tavolara [HNMSG21]. Tawi [MH20b]. Tawi-Tawi [MH20b]. tax [HYN^{+}23, RAB22, VSD^{+}23]. taxation [Åm21]. taxes [MVSP^{+}21]. taxonomy [FFE^{+}24]. Taylor [Ngu23]. Tears [SWW^{+}21]. tech [SLY21]. Technical [CPR21, GNJR24, DER^{+}21, GAMR20, RFPS21]. technique [KCCC^{+}21]. techniques [DBB^{+}22, SRSGMGG21b]. Technological [SDG^{+}23, VMD^{+}24, GLADGGVL23, KSB^{+}23a, RZ21b, RJ21, Ren21a, Sha20]. technologies [ABVPM20, JD20, SBP24, SG20c]. Technology [KZR22, OYECdOP24, Tan23a, GLS^{+}21, HPH22, MBL24, NNR^{+}21, Ren21b, RBNE20, VHD20, WSD21, ZGY21]. technology-compatible [GLS^{+}21]. Tel [PC20]. Telecouplings [SGL^{+}23]. temperate [QTPL21]. temperature [BK23]. temporal [FARBV^{+}24, GASLCP20, LYL^{+}22, RHST21, WL22b]. temporary [BBB^{+}20, LBL21]. ten [Her22, BBWC21]. tendencies [NGP^{+}21]. Tenggara [RYS^{+}21]. tensions [COSS21]. Tenualosa [DAAP21]. tenure [GWAPRPRU22, Li23b]. Teouma [SGPF21]. term [AFMdSdTdCC22, BSBD21, CVA22b, Jon21, Kim20, McC21, MM20, SOP^{+}22, SGFM^{+}21, Wat22, Fla21]. terminal [BHEE22, COJZ^{+}21, WW21a]. terminating [BNF21]. terminologies [HGSD23]. Terms [ABME22, CMC^{+}23]. terrestrial [MCR22]. territorial [HLW^{+}20, Lin20, MPSF20, ONA21, FMCT^{+}21, RM21b]. territorial-based [HLW^{+}20]. territorialisation [SW22]. territorialization [Lam21]. territory [YT20]. test [LM24, SCLM^{+}21]. tested [HVS22]. Testing [KMJ24, RKS^{+}23, ALML23]. Texas [SKS21]. Text [BDLO23, ZTA23]. texts [HSHS24]. textual [YYD23]. Thai [CCKK22]. Thailand [CJ23, CCKK22, Isr22, KL21a, KA21, VM21, WKBS20]. Thames [BMC20]. Their [Laj21, AANM^{+}22, BMB^{+}24, CCK^{+}21, CDC21b, CNG^{+}23, DA23, GBB^{+}21, Has20, HSK^{+}23, Kim20, NGP^{+}21, PMK^{+}21, PBG^{+}22, PMFR^{+}22, RLRS21, TAK^{+}22, WMS21, WL22b, FMCT^{+}21, ZZH21]. them [Gal21, NT20]. thematic [Sta23a]. themes [RSS^{+}20]. theoretic [BM21, ZGY21]. Theoretical [YLSC23, CSL21, MWB^{+}21]. theories [BSM23, EFFP23]. theorise [SG20b]. Theory [ABME22, Alf20, Gra21, GSB20, LYQM21, MLM23, PHF20, PFG^{+}21, XXLW24, ZY21, DLYM22, HVS22]. theory-based [GSB20]. there [NT20, OOGS20b]. Thermal [RJAT23, Teo23]. thinking [JL21d, OGS24, STKC22]. Third [Kim21, DMNT20, FY24]. thirty [Szy20]. Thoothoor [PRM^{+}20]. those [AO20, KOBS21]. thought [CGHV21, DMG^{+}23]. thoughts [Zhu23]. threat [Kar22b, SSOP^{+}21]. threaten [dMLB^{+}20]. threatened [BdLAdC^{+}20, BUM^{+}24, Jon21, Smi20, TBPC20]. threatens [CMI^{+}22, KgJqY^{+}24]. threats [FTGD^{+}20, GCM^{+}23, KHLS23, LTRA20, MIMS^{+}20, McC24, OTK^{+}21]. Three [HM22c, SAO^{+}21, TUBA22, BI21, FOKM23, GPM20, HNMSG21, LXSX21, Mur21b, SRSGMGG21b, BH23a, STG^{+}20]. thresher [CLB^{+}20, MH20b]. threshold [Sha20]. Thresholds [HSK^{+}23, PWL^{+}23a, TPG^{+}23]. throughout [VKN20]. thwarted [OSV21]. Tianbao [Sos21a]. tidal [AH23, BMC20, CA20, CKY22, HHEJB20, ZCG^{+}23]. tide [BMK^{+}22, RCH22]. tiered [HGL^{+}20, YLSC23]. Tierra [dPJ21]. tilapia [TCS^{+}20]. timber [VVR^{+}22]. Time [vLM22, Bau20, Bov22, BASM22, DMNT20, Dev22, Fox22, HRS^{+}24, SFGEPC22, TZM^{+}22, WPWM^{+}22, WL22b]. time-dependent [WL22b].

Timeline [BLH^{+}20]. timely [ZLD^{+}21]. times [GABSD21, LBL21, LEERRM^{+}21]. tipping [BHB^{+}24]. title [CCX23]. Toda [dHJTGO21]. Togo [NSZ^{+}21]. Tokyo [HWMO20, XWL^{+}20]. tolerable [AS23a]. too [SMHPMO^{+}21]. tool [BCJ21, BRP^{+}23, CRP^{+}20, CDB20, JHHG23, KKO^{+}22, OVSVPP^{+}22, PS21, Sol23b, SAB^{+}20, WS23]. Toolkit [MFC^{+}24]. tools [GGP^{+}22, QYD20, SMCW22]. toothfish [ADB^{+}22, CHC^{+}21]. top [SLB21]. top-down [SLB21]. topography [MLJ^{+}22]. TOPSIS [DG20, Kar22a]. Torre [Rus20]. Total [Sta23b, DSJ^{+}23, ZLD^{+}21]. tour [Dev22]. Tourism [Pha20, SWW^{+}21, AVPAFM^{+}22, AAFK^{+}21, BBC24, DHC22, GMGLV21, GMGLV22, HHBC20, Jon21, JK22, LLJZ22, LM24, MAMCHP^{+}20, Mil22, Mur21b, NUD^{+}23, NAN^{+}24, PFM20, PN20, RDM^{+}20, SBDC24, SCY24, Ton22b, TEA^{+}22]. tourist [LLJZ22, NNRS24, QWBY21]. tours [HACCM22]. toxic [Lev20]. toxicity [Lev20]. TPB [LYQM21]. trabeculata [CBO^{+}21]. traceability [DSWB20, HRC^{+}24, LFVCF24, THP^{+}22b]. traceable [DLYM22]. track [MKB21]. Tracking [CSGC22]. Trade [BCN23, CMSN^{+}22, EB23, FXC^{+}22, PKMT23, AED20, BdLAdC^{+}20, CJCS^{+}22, DC20, GAD^{+}20, GASLCP20, IOHC22, IO23, LP20, LCC20, LWG23, Ma20, MAA^{+}24, MMK^{+}22, PMM^{+}21, SMF^{+}23, SGL^{+}23, SSH^{+}23, SK24, SFB24, SFN^{+}21, SAO^{+}24, TJS22, XSCL21, ZTZ^{+}21, ZLGX20, ZT21, ZGY21, vRB24]. Trade-offs [BCN23, PKMT23, GASLCP20, ZGY21]. traders [IOHC22]. trades [CMI^{+}22]. Trading [WLXY22, WXLD21, WLQD21]. Traditional [DMd^{+}20, MFH^{+}20, DSQ^{+}22, FLB^{+}21, HLW^{+}20, HTP^{+}23, Li23b]. Traffic [MCW^{+}22b, TKC23, Boy21, Gun21, GHK20, MLJ^{+}22, SMF^{+}23, SSOP^{+}21]. trafficking [MGF20]. tragedy [Zha21]. training [CLD^{+}24, DG24a, KZR22, RKR20]. traits [WWC23]. trammel [WLW^{+}22]. trammel-net [WLW^{+}22]. Trang [KJE21b]. Trans [hC24, RB22]. Trans-boundary [hC24, RB22]. transaction [YLSC23]. Transboundary [GBCNCL^{+}21, LJ20, Abo23, CC21, DAAP21, FDGM24, HSdT21, KMT^{+}21, LJ23, PGA^{+}20, RKS^{+}24, SMYH22, Tel23b, VLS^{+}23]. Transdisciplinary [GCP^{+}20, STF21, SONB^{+}22, FMCT^{+}21, GTMP20]. transfer [VHD20, ZWXL21]. transferable [HvPPV20b, SSAP24]. transference [JPM^{+}23]. transfers [NRYG20]. transformation [BJGG^{+}24, EFFP23, LC22, HCS21]. Transformational [IRK^{+}21]. transformations [BSM23]. transformative [GM21, MdCLL^{+}23, VGC^{+}22]. transformed [PCA^{+}23]. Transforming [BBGG22, CCS23, HLW^{+}20, PMC24, SLB21]. Transition [ZLD^{+}21, AC23, Dom23, EMH^{+}23, GdACT20, GWAPRPRU22, KZR22, MF23, NNR^{+}21, SS21b, THS^{+}24, WP23, ZTA23, vHSSK20]. Transitions [KDS22, MM22, CMOB^{+}20]. transits [RZ21a]. Translating [SWI^{+}22]. translation [Fla21, SG21]. transmission [JG23]. transnational [BE20c, CAGBM22, TK23, Wu22]. Transparency [DH22, DFTK20, Gug22, GJS22, Pet22, TFF22, WKH22, Wil22, AMR21b, CMNH21, For22, Har22, Koe22, VVR^{+}22]. transparent [BD20, Fis22, HGG22c, SMYH22]. Transpolar [BSY^{+}20]. transport [Ada22, AB20, ADB^{+}22, BG20a, Dom23, LLW22, RLRS21, VGK23]. transportation [IGS^{+}23]. Transposition [BN20]. Transshipment [SOT^{+}22, HLLK22, SMYH22]. traps [ADR^{+}23, VGC^{+}22]. trawl [AE21, BDCJ22, DDF^{+}20, GMPT^{+}20, JGWM^{+}23, KGP^{+}20, MKM^{+}20, NSK20, RQB^{+}20, SC22, SCCW21, ZV20]. trawlers [DXS^{+}21]. Trawling [DAA21, ATF^{+}21, BMGDS^{+}20, CHEC21, DL21, FARBV^{+}24]. treatment [Che21b]. Treaty [AM22, Gar20, Dau23, HHD20, Kra22, MH24, MS20, MK24, TM23b, Liu21, TH22]. trend [RXN23, SWXM21, PZZ^{+}21]. Trends [KFB^{+}20, LBWS23, Mah22, QK24, Wu22, BSBD21, BMC20, BRR^{+}22, CLRW22, DDF^{+}20, DAAP21, FARBV^{+}24, GZW^{+}20, Gun21, HBJ21, LSD^{+}22, MAG^{+}22, MIH^{+}24, PL20a, SKDK23, Tho22, TDP20, WLRH^{+}20, LJ20, SGL^{+}21]. Tri [SLP^{+}22]. Tri-National [SLP^{+}22]. Trial [CBR21]. trials [CBR21, CC22]. Trindade [dS22a]. tripartite [HZ23]. Triple [MVR21b, DAA21, LCAX21]. trips [DWC^{+}20]. tropical [BLH^{+}20, GCC^{+}21b, LLLF22, MDLA21, NOL^{+}23, RTMCVMDE22, RHO24, SdCMN^{+}22, DMS^{+}24]. tropics [MFG^{+}22]. Trouble [Tea20, BL23b]. Troubled [Smi23a, FS21]. trunculus [SVP^{+}23]. trust [DGS^{+}24, EVPMKP22, FOH^{+}20, FS21, HSdT21, How20, KJS^{+}22, RKS^{+}24]. Trusts [HBCA^{+}24]. Tudela [MJ23]. Tumaco [PCCOS24]. Tuna [Kar23, SBS21, Abo23, BLH^{+}20, Bjø23, BDC^{+}21, DSWB20, EMH^{+}23, dHJTGO21, GSG20, GSB20, HWT23, HCS21, JG23, Kao23, KB20, Kaw22, KMG^{+}20, KgJqY^{+}24, Koe22, LLL^{+}24, PCSS^{+}21, PWL^{+}23b, RHO24, SMYH22, STF21, TOAPJ21, THP^{+}22b]. tunas [NOL^{+}23]. Tunisia [BDCJ22]. turbot [FGPPPV21]. TURF [FMCT^{+}21]. TURFs [AS23a]. Turkey [CA21, YU20]. Turkiye [Okş23, Ero22]. Turkmenistan [AB20]. Turning [NBB^{+}23, RCH22]. turtle [HBHJ20, SBW^{+}21, VPB^{+}24]. turtles [FMWW21]. Twitter [BDLO23]. Two [Ehl21b, KKO^{+}22, AFMdSdTdCC22, BBB^{+}20, CES20, DKW20, DBB^{+}22, Fro21, GJY23, GBC21b, GRPRVG23, JDF^{+}23, LJ20, LLSY22, MKM^{+}20, Pau21, PT23, RNGRC22, SON21, Sin21, ZWJvL20, ZTM^{+}23, dOPM^{+}20, dPJ21]. two-step [GJY23]. two-year [Sin21]. Tying [ADB^{+}22]. type [MB22, Zha22]. types [MÇ23, NNR^{+}21]. typology [BDC23, KLSG21, MPSF20]. Tyrrhenian [PBG^{+}22].

U.S. [AP23, AL20, BGK^{+}22, BBJ^{+}22, DAHS^{+}23, GLS^{+}21, HKKY22, LGF24, MKSG20, MM20, RMF^{+}21, RBNE20, Squ21]. Ucides [dFVCGdS21]. UK [EB23, Che21a, Dbo22, DFF23, FDF^{+}21, Fox22, GVL22, KFP20, KSB^{+}23a, LP21, RSS^{+}20, WtDD^{+}21, WCM^{+}24]. Ulrikke [Niu22]. Ulva [XLLS21]. umbilicals [VJL^{+}21]. unable [Han21]. Unambiguously [VvSvH^{+}22]. unavoidable [IBC^{+}20]. Unbound [Wil21a]. uncertain [BvdH23, DC20]. uncertainties [GSM21, SLS^{+}21]. uncertainty [NMCL24, SPCCvT21, UISM^{+}23]. uncharted [DBZ24]. UNCLOS [DBZ24, DC23, Ege23, Fra23, Guo20, HGL^{+}20, LR20, Men23, Ora23, Tel23b, Who23, Yu22, ZY23, ZT23b]. Uncovering [WFR23]. Uncrewed [Dun23, Ste23]. undelimited [Kim21]. under-management [dMLB^{+}20]. undercut [LWG23]. underdeveloped [GCP^{+}20, NWZ^{+}21]. Underlying [RE21]. Undermanaged [SSH^{+}23]. undermine [CMNH21]. undermining [BLL22, LV23, Tan24]. underperforming [MVR21b]. underprotected [RCH22]. undersized [NSK20]. understand [BMK^{+}22, GMRSP23, MFM21]. Understanding [BSM23, CMS^{+}21, DLYM22, DDGdS^{+}21, GBB^{+}21, HPS23, HP23, KP24, LLJZ22, NEH^{+}23, OBNK24, SBDC24, SKL23, XXMZ23, Zha21, ZCCQ22b, vRB24, AVF^{+}23a, FHBS23, IO23, LLM^{+}23, NHD23, O’D21, SOO20, SCCW21, STF21, Zou23, KDM^{+}20]. understandings [Has20]. Underwater [LC23, War23, AIG^{+}20, BRP^{+}23, HMG^{+}21a, KOCL23, LHW21, Lin24, ML22, MLJ^{+}22, MG22, MJMAA21, PT23, SCD21, TGM22, VÖB20a, VÖB20b, VRBC^{+}21, WM22a]. UNDRIP [Wan23c]. undulatus [HdM23]. UNECA [MFC^{+}24]. UNEP [FFGJ22]. UNFCCC [Tel23b]. unified [ZLY23]. Unilateral [Li23a, BGK^{+}22]. Union [Ada22, Álv23, Aqu23, FGC21, KGP^{+}20, LFPBD^{+}20, NVP^{+}23, PŚ22, SPFA^{+}20, SOLN21, VSDR^{+}22]. unique [Tse21]. United [Cro21, WW21b, CNB^{+}23, Col20, FS21, FBC^{+}23, Gar20, Gul20, HWU^{+}23, JFB^{+}24, KS22, KSB^{+}23b, MH24, NGA23, PWL^{+}23a, SSA^{+}23, SBT21, Squ20, SPH^{+}20, SRH23, VSD^{+}23, Wil21a]. universities [LLSY22, OKNB22]. University [LTJ^{+}23, Liu22, Niu22, Smi20, Smi22, Tan23a, TCT21]. University-driven [LTJ^{+}23]. unjust [SSV^{+}23]. unknown [DSC^{+}23]. unlikely [HSGS^{+}23]. unmanaged [OKS^{+}22]. unmanned [CZW20, XZ23]. Unmasking [SSP^{+}21a]. unofficial [SSZ^{+}24]. Unpacking [HvdBH23, NCE21, THP^{+}22b, HGSD23, LV23]. Unprecedented [AM22]. Unraveling [CJCS^{+}22, MAMCHP^{+}20]. Unravelling [CBO^{+}21, Mba22, SLS^{+}20]. unregulated [AGS20, BJH21, DGG22, KL21a, Ma20, TSH^{+}22a]. Unreported [SOLN20, AGS20, BJH21, DGG22, KL21a, Ma20, SSH^{+}23, TSH^{+}22a]. unsolved [RE23]. Untangling [CMGF20, EFFP23]. Untimely [AMR21b]. unveiled [BdLAdC^{+}20]. Unveiling [DG24b]. unwanted [CPR21, IBC^{+}20]. upcoming [BRP^{+}23]. uphold [NOB22]. upon [CCCC21]. Upper [ABDRV^{+}23]. ups [CSQ^{+}24]. upsurge [SK24]. upward [TDP20]. urban [KL21b, LHPJ23, MCW^{+}22b, NWZ^{+}21, QCM^{+}20]. urbanization [SGK21, Zha22]. urgency [SHZ^{+}21]. Uruguay [BJGG^{+}24]. usage [hBDDF22, RZ21a, YCBW21]. Use [BMK^{+}22, DB20, FMCT^{+}21, GBC21b, LXH^{+}22, Liu23, Ngu23, PCD^{+}23, RM21b, VAMR^{+}20, AKFY22, AAA^{+}24, AWBC20, AAW^{+}23, BH23b, BMRM23, Cha21, CvHF^{+}24, FMS^{+}22, GNJR24, GSAF21, HdM23, Her22, Kle21, Lev20, LDC^{+}20, LJRHH21, MM21a, MM23, OVA22, Pri22, PBH^{+}23, SCD21, SS21a, SSZ^{+}22, TLT^{+}22, TJZ^{+}21, DR22, VPRR^{+}21, WYH21, XZ23, YMG23, YWY20, YLSC23]. used [DQJ23, GCV^{+}20, HGSD23, LW22]. useful [PMC24]. User [Mrd23, Run24, AR23a, AR23b, AEJ24, Jen23, NHA23, SAB^{+}20]. users [SMC^{+}21]. Uses [Los24, ASM^{+}21b, DGMF^{+}20, SV22, SZBA20]. Using [CCS23, CAVJ^{+}20, CLB^{+}20, EAGV20, ES22, HLW^{+}23, HLCM21, KDH^{+}21, MEG^{+}21, PMT^{+}20, STIS20, SNP21, SSOP^{+}21, SKL^{+}20, STKC22, ASV^{+}21, BVT^{+}21, CRP^{+}20, CSH23, DLX21, DLYM22, EAAG21, EV22, FLC^{+}23, GMM^{+}20, GCS21, HF20, HTP^{+}23, JKV^{+}22, KCCC^{+}21, Kar22a, Kaw22, KPN20, KDM^{+}20, LWSS21, LrS22, LYQM21, MVR21a, MFC^{+}24, MLM23, NGT^{+}23, Okş23, RQB^{+}20, SGK21, SGWL21, TM23a, TPG^{+}23, WD23, WSD21, XSCL21, XWH^{+}22, ZTA23, Cat20, FGK^{+}20]. Ustica [HNMSG21]. usual [TV21]. Utilising [CTH^{+}21]. Utilization [TJS22, Hor23, SOK^{+}22]. Utilizing [CSHO^{+}20].

Vague [CMNH21]. Valdés [TDC21]. validation [LMSLMQ21]. valorisation [IBC^{+}20, PMAVM20]. Valuable [CLB^{+}23, dMLB^{+}20, KVL^{+}23, VKN20]. Valuation [MFC^{+}24, Cat20, CKY20, LrS22, LW22, NNRS24, QTPL21, ZMR20]. Value [dFVCGdS21, JAdS^{+}20, LLL^{+}24, YTG^{+}23, AANM^{+}22, ALM^{+}21b, BRM^{+}21a, BBJ^{+}22, Bot20, CLW^{+}21, CLB^{+}23, CSCGC20, DLJ^{+}20, FFYE21, GSG^{+}24, GNMGWM21, HPG20, HKKY22, JKKT20, KWMM20a, KWMM20b, LTW23, LCL20, LDL^{+}23, LRC24, MGS21, MBDG23, MWB^{+}21, MMW^{+}23, NAS^{+}22, PSTW21, PCSS^{+}21, QTBL23, RMU^{+}22, Seu22, SdCMN^{+}22, THT^{+}23, THP^{+}22b, TGHF^{+}22, TOHK20, TPG^{+}23, DR22, VKN20, WNM22, WCC^{+}20, WMS21, ZW21]. value-added [WCC^{+}20]. value-adding [LTW23]. value-chain [ALM^{+}21b]. value-satisfaction-demand [CLW^{+}21]. values [FRVB22, GCL21, TW22, ZWC^{+}23, ZCA^{+}23]. Valuing [DJC^{+}23, MKAF21, QWBY21, RNGRC22, SL20, WFR23]. Vancouver [IGS^{+}23]. Vanuatu [SGPF21, SNK^{+}20]. variability [CAVJ^{+}20]. variables [Kao23, OAM21]. variation [EMS^{+}23, GABA^{+}22]. Variations [HPPH22, Zho23]. various [LWX^{+}20, vLPBK21]. vary [CLL20]. varying [CDG^{+}21]. vector [SdACdSLdAS24, SGK21]. vehicles [CZW20, Kle21, VJL^{+}21, WABR^{+}22]. Veierud [Niu22]. venus [BSBD21]. Verde [Aqu23]. versa [AFMdSdTdCC22]. versus [BSSL^{+}20, CCKK22, COM20, DMd^{+}20, dOPM^{+}20]. Vessel [BCJ21, AMIM21, AdSMS21, CSGC22, Dam24, FMWW21, KLS22, KLSG21, KDM^{+}20, NSPRSS20, SSOP^{+}21, Shi22, Stø21]. vessels [BV21, Dun23, FMCD20, KZR22, MCW^{+}22b, OVA22, PSMR20, Set23, Ste23, SRY24, TBV^{+}23, Tse21, VÖB20a, VÖB20b, VAMR^{+}20, YS23]. VI [VSV^{+}23]. via [BDLO23, JHL21, MWB^{+}21, VSPMS^{+}22, ZLWY22]. viability [DAN^{+}23, YTBM^{+}22]. viable [RZ21a]. vice [AFMdSdTdCC22]. vice-versa [AFMdSdTdCC22]. vicinity [Shi22]. vicious [YS21]. victories [CMS20]. video [AIG^{+}20, HMG^{+}21a, SKL^{+}20]. Viet [Han21]. Vietnam [DBHT22, HS20, KJE21a, KJE21b, NHD23, TCS^{+}20, TVN23]. view [Kuł22]. views [ZLWY22]. VIII [WL20]. violence [Bra22]. Virgin [GMRSP23, YTBM^{+}22]. Virtual [HDH21]. visceral [CSH23]. vision [RFPS21, VAMR^{+}20, XT23, ZMZ23]. visitor [DHC22]. Vistula [RB22]. visualization [STIS20]. Visualizing [KBD^{+}20, ZTA23, WCF^{+}21]. Vitoria [dS22a]. VMS [CLS^{+}23]. voice [LEERRM^{+}21]. Vol [Mot22]. Volatility [VSPMS^{+}22]. volcanoes [SNK^{+}20]. volume [SVV24, SSH^{+}23]. Voluntary [Cha24, MM23, SMCW22, WS23, NH20]. volunteers [MFM21]. voting [Sta23b]. voyage [Bov22]. vs [GABSD21, KLSG21]. vu [SBT^{+}23]. Vulnerabilities [GPTASM23, HSM^{+}23]. Vulnerability [CBAPM20, KHH^{+}23, MGS21, CD24, CSGC22, DAN^{+}23, DGIS^{+}22, FMGV20, Gar24, HP23, JTR^{+}20, KDVR20, RDLA^{+}20, SMLRB^{+}21, SWF^{+}24, SCLM^{+}21, SPA^{+}21, VSPMS^{+}22]. vulnerable [BMR20, HRC^{+}24, JHHG23, dMLB^{+}20].

W [De 20]. Wadden [CWKE20]. Wades [RROP23]. wages [YTÇ22]. waka [MAR^{+}20b, MRD^{+}20]. Waka-Taurua [MRD^{+}20]. wake [FGL^{+}21]. wakes [FAC^{+}24]. Wales [BBGG22, MCW^{+}22a, BD20, BHB^{+}24, SSDQ20]. wall [dZ20]. war [Gra21]. warming [CTH^{+}21, KS23, LDAC20]. warning [SK24, WCZS21]. warnings [SLMG^{+}21]. Warships [Ngu23, GC22, Lin20]. Washington [Shi22]. waste [DMd^{+}20, DBS22a, JAH^{+}24, MB22, Öİ24, XTY22, XHZ^{+}21, ZMR20, ZHL^{+}23]. wastewater [PL24a]. watch [AdSMS21, vLPBK21]. watching [AdSMS21, STS^{+}23, Shi22, SHAC23]. Water [Zho23, BDCJ22, hC24, HCS21, JKV^{+}22, KLS22, LWZ23, LYQM21, MXZ23, MSGAR20, Shi23b, WL22a, XTY22, YS23, YL21, NSN22, Gul20]. watercraft [SMF^{+}23]. Waters [Smi23a, BBK^{+}21b, CMWN^{+}23, CMOJ^{+}24, DWC^{+}20, Dbo22, DER^{+}21, DBZ24, DD22, FS21, GBC^{+}20, KKO^{+}22, Lin20, MLJ^{+}22, MGBS24, ONA21, OOGS20a, OKFN24, SCGC20, TZM^{+}22, Tod22b, TYS22, Wil22, WtDD^{+}21, vdGD21]. watershed [JZW21]. wave [MMS24]. waves [BL22]. way [DSJ^{+}23, DHR23, FGK^{+}20, JDF^{+}23, MS21, VGC^{+}22, Wil21b]. ways [JPM^{+}23]. weakfish [dSJNS20]. Wealth [ASGS23, GV22]. weather [FRM22, GAMR20, SPA^{+}21]. web [OPL20, PBT^{+}20, SLS^{+}20, THP22a, TMTC20]. web-based [THP22a]. Weddell [TBH^{+}21]. wedge [SVP^{+}23]. wedgefishes [CJCS^{+}22]. Welfare [Gro20, ASGS23, CAM^{+}23, DBB^{+}22, STG^{+}20, SOS^{+}21b, TSS^{+}24, ZZH21, ZSWJ23]. well [AVF^{+}23a, GCF^{+}23, GSA23, MNGM24, MTC^{+}23, MH20a, QTPL21, RGH^{+}22, SCF^{+}22, SKL23]. well-being [AVF^{+}23a, GCF^{+}23, MNGM24, MTC^{+}23, RGH^{+}22, SCF^{+}22]. well-established [QTPL21]. well-regulated [SKL23]. wellbeing [CFM^{+}21, dlPdlLLC^{+}22]. West [BG20b, BGMCG24, GSG^{+}24, NRF21, RHL20, SWI^{+}22, SN23, DXS^{+}21, GSA23, AAA^{+}24, AWBC20, BCBS22, DP23, FTGD^{+}20, KABT21, RYS^{+}21]. Western [Jon21, ABRA20, BSBD21, BDC^{+}21, HNMSG21, KDH^{+}21, LSGG^{+}20, MFRQ23, McC24, NWZ^{+}21, NEH^{+}23, OCCSC23, Abo23, CCK^{+}21, GLADGGVL23, GAA^{+}20, KWF^{+}21, MAG^{+}22, OA20, RDD^{+}21]. wetlands [WLX21, ZCG^{+}23]. whacking [YML^{+}21]. Whale [AdSMS21, Mar20a, RCM^{+}22, GCV^{+}20, MM20, OKS^{+}22, STS^{+}23, SNP21, Shi22, SHAC23, TDC21]. whale-ship [SNP21]. Whale-watch [AdSMS21]. whale-watching [AdSMS21, Shi22, SHAC23]. whales [Dam24, FMCD20, GHK20, HSH^{+}21, MCW^{+}22b, Shi22, SHAC23]. Whaling [WL20, WNM22]. whelk [DUYEA21]. Where [BÖ24, RSFvH^{+}22]. whether [HRR^{+}20]. which [RE21]. white [AS23b, BBHN20, HACCM22, MAMCHP^{+}20, SMHPMO^{+}21, HACCM22]. Who [CSCGC20, KOBS21, LV23, Pre21, MK24]. Whose [MB21]. wicked [Gro20, OOGS20b]. wide [HMG^{+}21a]. widespread [RZ21a]. Wild [Geh23, MM22]. wild-caught [MM22]. Wilderness [McC20]. Wildlife [KHMM20, Lot22, NUD^{+}23, GBC21a]. Will [BGK^{+}22, PK23, Bal22, Ber23, JD20, KZR22]. Willingness [CLC21, BR22, BUM^{+}24, CLL20, CKY22, XLLS21, XXLW24]. willingness-to-pay [CLL20]. win [Ton22b]. win-win [Ton22b]. wind [BvdBvdFK20, BST23, BLY^{+}24, CSLGFV20, CCW21, DWC^{+}20, DBHT22, FAC^{+}24, HRR^{+}20, KNY20, MBORGN24, OC20, ONA21, OKH^{+}24, PKMT23, SBT21, SZBA20, SVKP21, TYS24, vRB24, FAC^{+}24]. Windfarms [DS22b]. winding [BMH^{+}22]. window [Lot23]. Winners [PKMT23, BCMCM21]. withdraw [PCK22]. withdrawals [LLM^{+}23]. within [BCJ21, CTH^{+}21, CG22, Gar24, MPD^{+}23, NBB^{+}23, SMKS22, TOL^{+}22, VHD20, dS21]. without [MLM23]. WMO [Yu22]. Women [FS22, MMC^{+}22, MMK^{+}23, ABRA20, FFYE21, GE23, LL24, MMK^{+}22, NEF23, OBNK24, PK20a, Szy20, UOO20]. words [HBL^{+}21]. Work [TNY23, ASKI20, BBK21a, JSW^{+}22a, Kuł22, LJ20, LFLF22, LSD^{+}22, MWWS21, Pau21, SGA^{+}21, YS23, YTÇ22, NSN22]. Work-family [TNY23]. work-related [Pau21]. Worker [SMCW22]. Worker-less [SMCW22]. workers [SMCW22, VM21, WS23]. workforce [JD20]. Working [Bau20, AFHHD22, GPH^{+}21, JSW^{+}22a, RAFLRP21, VM21]. Workplace [ÖB22, NEF23, WW21a]. workplaces [NOB22]. works [CHB^{+}20]. Workshop [MGBS24, GRA^{+}22, HMG^{+}21a]. world [BLM24, BMR23, BS22, CLP21, EMH^{+}23, HM22c, KFP20, Seu22, Smi22, ZJ22, CMSN^{+}22]. worldwide [CMGF20]. worth [Tod20]. Would [BBHN20, ALLS23]. wrasse [HdM23, AGSI22]. wreck [PT23]. WTO [ALLS23, PL24a]. WWII [Lin20].

x [Smi22]. Xávega [ACV^{+}21]. xiii [Smi20]. Xin’an [JZW21]. xv [Rub24]. XX [WL20, Sta23a].

yacht [VGK23]. Yaldad [DSQ^{+}22]. Yamamoto [dHJTGO21]. Yangtze [QYX21]. Yantai [DG20]. year [Bau20, VBd^{+}21, MFT^{+}22, NUD^{+}23, PBG^{+}22, Sin21, VMM21]. Years [FSKJ23, BMR23, Ege23, FFE^{+}24, Gal22, Gul23, KSB^{+}23b, Liu21, NAP^{+}21, Smi23a, Szy20]. yellow [TOAPJ21, Cho22, LSH21]. yellow-fin [TOAPJ21]. yields [McC21]. Yong [TCT21]. Yong-An [TCT21]. York [Ngu23, Smi23a]. Young [Tan23a, DBBL21]. youth [SN23]. Yucatan [HDAPG^{+}24].

Zambia [MMK^{+}22]. Zealand [LLT^{+}20, MAR^{+}20b, QTBL23, RR20, RM21a, RNGRC22, RRM^{+}24, SPJB22, WJLSH21]. zero [GCMT20]. Zhejiang [LYL^{+}22]. Zhikong [YML^{+}21]. Zhoushan [YD22]. Zone [BMR23, CAGBM22, BGMCG24, CDHY22, CLP21, FRVB22, OKFN24, RXX24, dSdLSGB22, FXC^{+}22]. zones [BBB^{+}20, JACD20, PMR23, QF20, dO23, SOLN20]. zoning [MFG^{+}22, ONA21, SRR^{+}21, SSB^{+}23, TZZ^{+}21, NAP^{+}21].

References

Addo:2021:ISP

[AAA^{+}21] Kwasi Appeaning Addo, Olanike Adeyemo, Ibukun Jacob Adewumi, Robert Blasiak, S. Kim Juniper, Dwight Owens, Tessa Owens, and U. Rashid Sumaila. Informed selfishness — Practical reflections on building a sustainable ocean economy. Marine Policy, 133(??):Article 104735, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003468.

Abrokwah:2024:ASS

[AAA^{+}24] Sika Abrokwah, Ivy Serwaa Gyimah Akuoko, Margaret Fafa Awushie Akwetey, Mike Izava Olendo, Peter Kershaw, and Denis Worlanyo Aheto. Achieving sustainable source reduction of marine litter for ocean conservation in West Africa: Insights from single-use plastic consumers in Liberia. Marine Policy, 159(??):Article 105937, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004700.

Alvarez:2023:CPC

[AACdlT23] Lourdes Alvarez, Gregory Arcaya, Francisco Cueto, and Renato de la Torre. Can public credit programs improve artisanal fisher performance? The case of FONDEPES a credit program. Marine Policy, 156(??):Article 105792, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003251.

Arcos-Aguilar:2021:DTM

[AAFK^{+}21] Ramiro Arcos-Aguilar, Fabio Favoretto, Joy A. Kumagai, Victoria Jiménez-Esquivel, Adán L. Martínez-Cruz, and Octavio Aburto-Oropeza. Diving tourism in Mexico — Economic and conservation importance. Marine Policy, 126(??):Article 104410, April 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000208.

AlAmin:2021:CKA

[AAKI21] M. A. Al Amin, L. Adrianto, T. Kusumastanto, and Z. Imran. Community knowledge, attitudes and practices towards environmental conservation: Assessing influencing factors in Jor Bay Lombok Indonesia. Marine Policy, 129(??):Article 104521, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001329.

Acosta-Alba:2022:MFV

[AANM^{+}22] Ivonne Acosta-Alba, Gian Nicolay, Adama Mbaye, Moustapha Dème, Ludovic Andres, Marc Oswald, Henri Zerbo, Joseph Ndenn, and Angel Avadí. Mapping fisheries value chains to facilitate their sustainability assessment: Case studies in The Gambia and Mali. Marine Policy, 135(??):Article 104854, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004656.

Ankrah:2023:CUI

[AAW^{+}23] Martinson Twumasi Ankrah, Dennis Asante, Pengcheng Wang, Evans Brako Ntiamoah, and Yuansheng Jiang. Can the use of the Internet improve fish farmers’ financial performance? Evidence from Ghana. Marine Policy, 149(??):Article 105494, March 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000210.

Akbulaev:2020:MTE

[AB20] Nurkhodzha Akbulaev and Gadir Bayramli. Maritime transport and economic growth: Interconnection and influence (an example of the countries in the Caspian sea coast; Russia, Azerbaijan, Turkmenistan, Kazakhstan and Iran). Marine Policy, 118(??):Article 104005, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302726.

Aceves-Bueno:2023:DGI

[ABDRV^{+}23] E. Aceves-Bueno, L. Davids, J. A. Rodriguez-Valencia, A. M. Jaramillo-Legorreta, E. Nieto-Garcia, G. Cárdenas-Hinojosa, E. Hidalgo-Pla, A. Bonilla-Garzón, A. J. Diaz de Leon, L. Rojas-Bracho, and M. A. Cisneros-Mata. Derelict gear from an illegal fishery: Lessons from gear retrieval efforts in the Upper Gulf of California. Marine Policy, 147(??):Article 105387, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004341.

Afriansyah:2024:SIR

[ABI24] Arie Afriansyah, Leonardo Bernard, and Christou Imanuel. Should Indonesia regulate foreign military activities in its EEZ? Marine Policy, 159(??):Article 105931, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004645.

Allan:2021:ECD

[ABM21] Jamie C. Allan, Karen F. Beazley, and Anna Metaxas. Ecological criteria for designing effective MPA networks for large migratory pelagics: Assessing the consistency between IUCN best practices and scholarly literature. Marine Policy, 127(??):Article 104219, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308654.

Ayaz:2022:RSP

[ABME22] İlke Sezin Ayaz, Umur Bucak, Mahmut Mollaoğlu, and Soner Esmer. Resilience strategies of ports against Covid-19 in terms of chaos theory. Marine Policy, 146(??):Article 105323, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003700.

Abolhassani:2023:SPT

[Abo23] Angela Abolhassani. Scalar politics in transboundary fisheries management: the Western and Central Pacific Fisheries Commission as an eco-scalar fix for South Pacific albacore tuna management. Marine Policy, 152(??):Article 105583, June 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001100.

Ameyaw:2020:FCA

[ABRA20] Anita B. Ameyaw, Annette Breckwoldt, Hauke Reuter, and Denis W. Aheto. From fish to cash: Analyzing the role of women in fisheries in the western region of Ghana. Marine Policy, 113(??):Article 103790, March 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306541.

Antelo:2020:IMT

[ABVPM20] Luis T. Antelo, Jose M. Bellido, Julio Valeiras, and Ricardo I. Pérez-Martín. Innovative marine technologies applied to discard mitigation and management: the MARTEC18 conference. Marine Policy, 116(??):Article 103911, June 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20302013.

Azzurro:2021:PMI

[AC21] Ernesto Azzurro and Jacopo Cerri. Participatory mapping of invasive species: a demonstration in a coastal lagoon. Marine Policy, 126(??):Article 104412, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000221.

Axon:2023:BBE

[AC23] Stephen Axon and Sammy Collier. Breaking blue: Establishing comprehensive policy for a just and inclusive transition for the Blue Economy. Marine Policy, 147(??):Article 105343, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003906.

Albrecht:2021:PAD

[ACA^{+}21] Renee Albrecht, Carly N. Cook, Olive Andrews, Kelsey E. Roberts, Martin F. J. Taylor, Michael B. Mascia, and Rachel E. Golden Kroner. Protected area downgrading, downsizing, and degazettement (PADDD) in marine protected areas. Marine Policy, 129(??):Article 104437, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21000476.

Ansong:2021:MAB

[ACG21] J. Ansong, H. Calado, and P. M. Gilliland. A multifaceted approach to building capacity for marine/maritime spatial planning based on European experience. Marine Policy, 132(??):Article 103422, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304056.

Alvarez:2020:DSA

[ÁCT20] Antonio Álvarez, Lorena Couce, and Lourdes Trujillo. Does specialization affect the efficiency of small-scale fishing boats? Marine Policy, 113(??):Article 103796, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301721.

Alves:2021:SSA

[ACV^{+}21] Fátima Alves, Paula Casaleiro, Cristina Sá Valentim, Paula Castro, and Miguel Pardal. Small-scale artisanal fisheries and its networks: the case of Arte Xávega in Portugal. Marine Policy, 128(??):Article 104484, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000956.

Adamowicz:2022:DMT

[Ada22] Magdalena Adamowicz. Decarbonisation of maritime transport — European Union measures as an inspiration for global solutions? Marine Policy, 145(??):Article 105085, November 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001324.

Ashford:2022:TPS

[ADB^{+}22] Julian Ashford, Michael Dinniman, Cassandra Brooks, Lian Wei, and Guoping Zhu. Tying policy to system: Does the Ross Sea region marine reserve protect transport pathways connecting the life history of Antarctic toothfish? Marine Policy, 136(??):Article 104903, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005145.

Andres:2021:MCS

[ADR^{+}21] Marga Andrés, Matthias Delpey, Irene Ruiz, Amandine Declerck, Caroline Sarrade, Peggy Bergeron, and Oihane C. Basurko. Measuring and comparing solutions for floating marine litter removal: Lessons learned in the southeast coast of the Bay of Biscay from an economic perspective. Marine Policy, 127(??):Article 104450, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000609.

Antonelis:2023:DED

[ADR^{+}23] Kyle Antonelis, Joan Drinkwin, Paul Rudell, Jason Morgan, James Selleck, Donald Velasquez, and Don P. Rothaus. Determining effectiveness of Dungeness crab escapement in derelict traps. Marine Policy, 149(??):Article 105499, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2300026X.

Arranz:2021:WWV

[AdSMS21] Patricia Arranz, Natacha Aguilar de Soto, Peter T. Madsen, and Kate R. Sprogis. Whale-watch vessel noise levels with applications to whale-watching guidelines and conservation. Marine Policy, 134(??):Article 104776, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003870.

Akpalu:2021:ESE

[AE21] Wisdom Akpalu and Håkan Eggert. The economic, social and ecological performance of the industrial trawl fishery in Ghana: Application of the FPIs. Marine Policy, 125(??):Article 104241, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20308873.

Anbleyth-Evans:2023:PDM

[AE23] Jeremy Anbleyth-Evans. Port developments and marine democracy in Latin America. Marine Policy, 157 (??):Article 105767, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003007.

Alshubiri:2020:IFP

[AED20] Faris Alshubiri, Mohamed Elheddad, and Nadia Doytch. The impact of fish production on marine trade balance and foreign direct investment: an empirical study of the GCC economies. Marine Policy, 116(??):Article 103660, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19301320.

Arnason:2024:SUR

[AEJ24] Ragnar Arnason, Håkan Eggert, and Frank Jensen. Strong user rights in fisheries: an editorial note. Marine Policy, 161(??):Article 106009, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000071.

Amaralal:2023:ICP

[AEL^{+}23] K. H. M. L. Amaralal, E. M. A. P. Edirimanna, W. A. S. W. Lakmini, K. K. D. Chamodi, A. U. Kuragodage, R. G. Sanuja, E. G. K. Y. C. Bandara, M. M. A. S. Maheepala, M. N. D. F. Abeykoon, K. P. G. L. Sadaruwan, S. Kuganathan, and K. H. M. Ashoka Deepananda. Impacts of COVID-19 pandemic on the fisheries sector of Sri Lanka. Marine Policy, 147(??):Article 105339, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003864.

Anbleyth-Evans:2020:TMD

[AELR^{+}20] Jeremy Anbleyth-Evans, Francisco Araos Leiva, Francisco Ther Rios, Ricardo Segovia Cortés, Vreni Häussermann, and Carolina Aguirre-Munoz. Toward marine democracy in Chile: Examining aquaculture ecological impacts through common property local ecological knowledge. Marine Policy, 113(??):Article 103690, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19303513.

Alvarez-Fernandez:2020:LQM

[ÁFFSC20] Inmaculada Álvarez-Fernández, Juan Freire, and Noela Sánchez-Carnero. Low-quality management of Marine Protected Areas in the North-East Atlantic. Marine Policy, 117(??):Article 103922, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19303847.

Amon:2022:SWE

[AFHHD22] Diva J. Amon, Zoleka Filander, Lindsey Harris, and Harriet Harden-Davies. Safe working environments are key to improving inclusion in open-ocean, deep-ocean, and high-seas science. Marine Policy, 137(??):Article 104947, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005583.

Azzahra:2023:BRA

[AFHM23] Tami Azzahra, Muh. Farhan, Muh Aldhy Hatmar, and Mamnunah. Book review: Asia–Pacific Fishing Livelihoods, Michael Fabinyi, Kate Barclay. Palgrave Macmillan, Cham (2021). Marine Policy, 155(??):Article 105704, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002373.

Andriguetto-Filho:2022:SFV

[AFMdSdTdCC22] José Milton Andriguetto-Filho, Rodrigo Pereira Medeiros, André Martins Vaz dos Santos, and Paulo de Tarso da Cunha Chaves. Shrimp up, fish down, and vice-versa: Fishers’ strategies and long-term changes in small-scale fisheries landings at two spatial levels in Southern Brazil. Marine Policy, 143(??):Article 105184, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002317.

Arifin:2023:IPR

[AFS^{+}23] Zainal Arifin, Dede Falahudin, Hiroaki Saito, Tuti Hendrawati Mintarsih, Muhammad Hafizt, and Yulianto Suteja. Indonesian policy and researches toward 70% reduction of marine plastic pollution by 2025. Marine Policy, 155(??):Article 105692, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002191.

Asche:2021:FPA

[AGA^{+}21] Frank Asche, Taryn M. Garlock, Wisdom Akpalu, Ebele Chinelo Amaechina, Robert Botta, Nnaemeka Andegbe Chukwuone, Håkan Eggert, Ken Hutchings, Razack Lokina, Byela Tibesigwa, and Jane K. Turpie. Fisheries performance in Africa: an analysis based on data from 14 countries. Marine Policy, 125(??):Article 104263, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030909X.

Alves:2022:FPB

[AGK22] Catherine L. Alves, Olga Denise Garcia, and Randall A. Kramer. Fisher perceptions of Belize’s Managed Access program reveal overall support but need for improved enforcement. Marine Policy, 143(??):Article 105192, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002391.

Amon:2022:ASG

[AGM^{+}22] Diva J. Amon, Sabine Gollner, Telmo Morato, Craig R. Smith, Chong Chen, Sabine Christiansen, Bronwen Currie, Jeffrey C. Drazen, Tomohiko Fukushima, Matthew Gianni, Kristina M. Gjerde, Andrew J. Gooday, Georgina Guillen Grillo, Matthias Haeckel, Thembile Joyini, Se-Jong Ju, Lisa A. Levin, Anna Metaxas, Kamila Mianowicz, Tina N. Molodtsova, Ingo Narberhaus, Beth N. Orcutt, Alison Swaddling, Joshua Tuhumwire, Patricio Urueña Palacio, Michelle Walker, Phil Weaver, Xue-Wei Xu, Clement Yow Mulalap, Peter E. T. Edwards, and Chris Pickens. Assessment of scientific gaps related to the effective environmental management of deep-seabed mining. Marine Policy, 138(??):Article 105006, April 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000537.

Aedo:2020:SDS

[AGN^{+}20] Gustavo Aedo, Cristóbal Garcés, Edwin Niklitschek, Selim Musleh, Luis A. Cubillos, and Renato A. Quiñones. Spatial distribution of small pelagic fishes: Implications for fishing quota allocations. Marine Policy, 120(??):Article 104147, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19304956.

Alqattan:2020:IUU

[AGS20] Mohammad E. A. Alqattan, Tim S. Gray, and Selina M. Stead. The illegal, unreported and unregulated fishing in Kuwait: Problems and solutions. Marine Policy, 116(??):Article 103775, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19302672.

Arieta:2022:CIA

[AGSI22] Siti Arieta, Budhi Gunawan, R. Widya Setiabudi Sumadinata, and Yudi Nurul Ihsan. Calling for institutional arrangements in Napoleon Wrasse conservation and management in Indonesia: the Anambas Islands case. Marine Policy, 145(??):Article 105258, November 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003050.

Autzen:2021:WSB

[AH21] Mathilde Højrup Autzen and Troels Jacob Hegland. When ‘sustainability’ becomes the norm: Power dynamics in the making of a new eco-label for low-environmental-impact, small-scale fisheries. Marine Policy, 133(??):Article 104742, November 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003535.

Alducci:2023:PCI

[AH23] Beatrice Alducci and Stephen Hynes. Protecting coastal infrastructure in the face of increasing climate related pressures: a cost-benefit analysis of maintaining a tidal kayaking feature in Ireland. Marine Policy, 150(??):Article 105550, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000775.

Amkieltiela:2022:REI

[AHAB^{+}22] Amkieltiela, Christian N. Handayani, Dominic A. Andradi-Brown, Estradivari, Amanda K. Ford, Maria Beger, Amehr Hakim, Dominic K. Muenzel, Eleanor Carter, Firdaus Agung, Laura Veverka, Mohamad Iqbal, Muhammad E. Lazuardi, Muhammad N. Fauzi, Sylvie N. Tranter, and Gabby N. Ahmadia. The rapid expansion of Indonesia’s marine protected area requires improvement in management effectiveness. Marine Policy, 146(??):Article 105257, December 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003049.

Araos:2023:FBA

[AHB^{+}23] Francisco Araos, Carlos Hidalgo, Francisco Brañas, Jeremy Anbleyth-Evans, Florencia Diestre, and Allan Yu Iwama. Facing the blue Anthropocene in Patagonia by empowering indigenous peoples’ action networks. Marine Policy, 147(??):Article 105397, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004444.

Abdel-Hady:2024:CAS

[AHBAKH24] Mahmoud M. Abdel-Hady, Ahmed A. Barrania, Zeinab M. Abdel-Khalek, and Shaimaa M. Haggag. A comprehensive approach to strategic planning for marine aquaculture in Egypt: SWOT-AHP analysis. Marine Policy, 162(??):Article 106057, April 2024. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000551.

Ansong:2022:IBI

[AHM22] Joseph Onwona Ansong, Ritchie Heather, and Linda McElduff. Institutional barriers to integrated marine spatial planning on the island of Ireland. Marine Policy, 141(??):Article 105082, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001294.

Apps:2023:SDC

[AHN^{+}23] Kirin Apps, Elizabeth Heagney, Quach Thi Khanh Ngoc, Kay Dimmock, and Kirsten Benkendorff. Scuba divers, coral reefs, and knowledge of ocean acidification. Marine Policy, 155(??):Article 105779, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003123.

Aguzzi:2020:EVM

[AIG^{+}20] J. Aguzzi, N. Iveša, M. Gelli, C. Costa, A. Gavrilovic, N. Cukrov, M. Cukrov, Nu Cukrov, D. Omanovic, M. Štifanić, S. Marini, M. Piria, E. Azzurro, E. Fanelli, and R. Danovaro. Ecological video monitoring of Marine Protected Areas by underwater cabled surveillance cameras. Marine Policy, 119(??):Article 104052, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302669.

Abe:2022:DUM

[AKFY22] Hiroya Abe, Yuko F. Kitano, Tomohiro Fujita, and Hiroya Yamano. Distribution, use, management, regulation, and future concerns of reef-building corals based on administrative documents in Japan. Marine Policy, 141(??):Article 105090, July 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001373.

Arguello:2022:RSA

[AKL^{+}22] Gabriela Argüello, Niels Krabbe, David Langlet, Ida-Maja Hassellöv, Claes Martinson, and Astrid Helmstad. Regulation of ships at anchor: Safety and environmental implications. Marine Policy, 140(??):Article 105052, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000999.

Ayers:2020:ESC

[AL20] Adam L. Ayers and Kirsten Leong. Examining the seascape of compliance in U.S. Pacific island fisheries. Marine Policy, 115(??):Article 103820, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305573.

Alff:2020:NCI

[Alf20] Henryk Alff. Negotiating coastal infrastructures: an evolutionary governance theory (EGT) approach to Chinese high-modernist development along the Indian Ocean. Marine Policy, 112(??):Article 103545, February 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303348.

Arias:2022:RBM

[ALF^{+}22] Juliana Arias, Unn Laksá, Benvido D’Oliveira Fonseca, Mamadou Diallo, Khallahi Brahim, Alexandre Rodríguez, Sonia Martin Doblado, Juliana Antunes Galvão, Jónas R. Vidarsson, and Ragnhildur Fridriksdóttir. Results-based management in practice: Lessons learnt and policy recommendations from the implementation of RBM in European fisheries outside Europe. Marine Policy, 139(??):Article 105038, May 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000859.

Ayres:2024:LDK

[ALLR^{+}24] Kathryn A. Ayres, Frida Lara-Lizardi, Callum M. Roberts, Walter Pisco-Limones, Peter Klimley, Salvador J. Jorgensen, Felipe Galván-Magaña, Mauricio Hoyos-Padilla, and James T. Ketchum. Local diver knowledge reveals decline in scalloped hammerhead sharks (Sphyrna lewini) at seamounts in the southwestern Gulf of California. Marine Policy, 159(??):Article 105915, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004487.

Alger:2023:WWA

[ALLS23] Justin Alger, Philippe Le Billon, Eric Leinberger, and U. Rashid Sumaila. What would Article 5.1 of the 2022 WTO Ministerial Agreement on Fisheries Subsidies accomplish? Marine Policy, 153(??):Article 105641, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001689.

Almojil:2021:LEK

[Alm21a] Dareen Almojil. Local ecological knowledge of fisheries charts decline of sharks in data-poor regions. Marine Policy, 132(??):Article 104638, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002499.

Asri:2021:RMM

[ALM^{+}21b] Azam Asri, Virginie Le Masson, Valéria Montalescot, Phaik Eem Lim, Adibi M. Nor, Hanafi Hussin, and Louise Shaxson. The role of migrants in the Malaysian seaweed value-chain. Marine Policy, 134(??):Article 104812, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004231.

Akerlof:2023:WLP

[ALML23] K. L. Akerlof, Jacqueline Loevenich, Sara Melena, and Christine A. Lipsky. What’s local policy got to do with it? Testing the efficacy of federal interventions to protect coastal parks from sunscreen chemical pollution in Hawaii and North Carolina. Marine Policy, 157(??):Article 105849, November 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003822.

Alshubiri:2020:EBE

[Als20] Faris Alshubiri. Editorial: the blue economy of the GCC region: Sustainability and security. Marine Policy, 116(??):Article 103843, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20300816.

Alvarez:2023:PFC

[Álv23] Inma Valeije Álvarez. The prosecution of fisheries crime in Spanish criminal law: the impact of European Union regulations. Marine Policy, 147(??):Article 105327, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003748.

Aam:2021:CPS

[Åm21] Heidrun Åm. A critical policy study on why introducing resource rent taxation in Norwegian salmon aquaculture failed. Marine Policy, 131(??):Article 104692, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003031.

Arpi:2022:FAS

[AM22] Bruno Arpi and Jeffrey McGee. Fishing around the South Georgia Islands and the “Question of the Falklands/Malvinas”: Unprecedented challenges for the Antarctic Treaty System. Marine Policy, 143(??):Article 105201, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002482.

Abe:2021:RDE

[AMIM21] Keita Abe, Yusuke Matsuki, Gakushi Ishimura, and Shunsuke Managi. Rent distribution in an ex-vessel auction market of fisheries. Marine Policy, 133(??):Article 104771, November 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003821.

Abspoel:2021:CMS

[AMK^{+}21] Lodewijk Abspoel, Igor Mayer, Xander Keijser, Harald Warmelink, Rhona Fairgrieve, Malena Ripken, Andrej Abramic, Andreas Kannen, Roland Cormier, and Sue Kidd. Communicating Maritime Spatial Planning: the MSP Challenge approach. Marine Policy, 132(??):Article 103486, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18304044.

Araujo:2021:GIP

[AML^{+}21] Laís S. Araujo, Ulises Rodrigo Magdalena, Tayana S. Louzada, Paulo S. Salomon, Fernando C. Moraes, Beatrice P. Ferreira, Eduardo T. C. Paes, Alex C. Bastos, Renato C. Pereira, Leonardo T. Salgado, Maria Lucia Lorini, Patricia Yager, and Rodrigo L. Moura. Growing industrialization and poor conservation planning challenge natural resources’ management in the Amazon Shelf off Brazil. Marine Policy, 128(??):Article 104465, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000762.

Atkins:2023:SPR

[AMLC^{+}23] Shanan Atkins, Judy Mann-Lang, Geremy Cliff, Neville Pillay, and Mauricio Cantor. Stakeholder perceptions reveal obstacles and opportunities to change lethal methods of protecting bathers from sharks. Marine Policy, 155(??):Article 105762, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002956.

Archibald:2021:IGC

[AMR21a] D. W. Archibald, R. McIver, and R. Rangeley. The implementation gap in Canadian fishery policy: Fisheries rebuilding and sustainability at risk. Marine Policy, 129(??):Article 104490, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001019.

Archibald:2021:UPD

[AMR21b] D. W. Archibald, R. McIver, and R. Rangeley. Untimely publications: Delayed Canadian fisheries science advice limits transparency of decision-making. Marine Policy, 132(??):Article 104690, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003018.

Anonymous:2020:A

[Ano20a] Anonymous. August 2020. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:D

[Ano20b] Anonymous. December 2020. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:EBa

[Ano20c] Anonymous. Editorial Board. Marine Policy, 111(??): Article 103762, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19308681.

Anonymous:2020:EBb

[Ano20d] Anonymous. Editorial Board. Marine Policy, 112 (??):Article 103811, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20300063.

Anonymous:2020:EBc

[Ano20e] Anonymous. Editorial Board. Marine Policy, 113(??): Article 103838, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X20300865.

Anonymous:2020:EBd

[Ano20f] Anonymous. Editorial Board. Marine Policy, 115(??): Article 103916, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X20302384.

Anonymous:2020:EBe

[Ano20g] Anonymous. Editorial Board. Marine Policy, 116(??): Article 103995, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X20303687.

Anonymous:2020:EBf

[Ano20h] Anonymous. Editorial Board. Marine Policy, 117(??): Article 104035, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X2030470X.

Anonymous:2020:EBg

[Ano20i] Anonymous. Editorial Board. Marine Policy, 118(??): Article 104078, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X20305388.

Anonymous:2020:EBh

[Ano20j] Anonymous. Editorial Board. Marine Policy, 119 (??):Article 104167, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20307831.

Anonymous:2020:EBi

[Ano20k] Anonymous. Editorial Board. Marine Policy, 120(??): Article 104204, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X20308496.

Anonymous:2020:EBj

[Ano20l] Anonymous. Editorial Board. Marine Policy, 121 (??):Article 104234, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20308800.

Anonymous:2020:EBk

[Ano20m] Anonymous. Editorial Board. Marine Policy, 122 (??):Article 104309, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309544.

Anonymous:2020:F

[Ano20n] Anonymous. February 2020. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:Ja

[Ano20o] Anonymous. January 2020. Marine Policy, 111(??):??, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:Jc

[Ano20p] Anonymous. July 2020. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:Jb

[Ano20q] Anonymous. June 2020. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:Ma

[Ano20r] Anonymous. March 2020. Marine Policy, 113(??):??, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:Mb

[Ano20s] Anonymous. May 2020. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:N

[Ano20t] Anonymous. November 2020. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:O

[Ano20u] Anonymous. October 2020. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:S

[Ano20v] Anonymous. September 2020. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:Aa

[Ano21a] Anonymous. April 2021. Marine Policy, 126(??):??, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:Ab

[Ano21b] Anonymous. August 2021. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:D

[Ano21c] Anonymous. December 2021. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:EBa

[Ano21d] Anonymous. Editorial Board. Marine Policy, ??(??):Article 103902, ???? 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20301937.

Anonymous:2021:EBb

[Ano21e] Anonymous. Editorial Board. Marine Policy, 123(??): Article 104345, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X20309957.

Anonymous:2021:EBc

[Ano21f] Anonymous. Editorial Board. Marine Policy, 124 (??):Article 104395, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21000051.

Anonymous:2021:EBd

[Ano21g] Anonymous. Editorial Board. Marine Policy, 125(??): Article 104417, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X21000270.

Anonymous:2021:EBe

[Ano21h] Anonymous. Editorial Board. Marine Policy, 126(??): Article 104457, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X21000683.

Anonymous:2021:EBf

[Ano21i] Anonymous. Editorial Board. Marine Policy, 127(??): Article 104495, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X21001068.

Anonymous:2021:EBg

[Ano21j] Anonymous. Editorial Board. Marine Policy, 128(??): Article 104538, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X21001494.

Anonymous:2021:EBh

[Ano21k] Anonymous. Editorial Board. Marine Policy, 129(??): Article 104597, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X21002086.

Anonymous:2021:EBi

[Ano21l] Anonymous. Editorial Board. Marine Policy, 130(??): Article 104628, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X21002396.

Anonymous:2021:EBj

[Ano21m] Anonymous. Editorial Board. Marine Policy, 131 (??):Article 104718, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003298.

Anonymous:2021:EBk

[Ano21n] Anonymous. Editorial Board. Marine Policy, 132(??): Article 104758, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X21003699.

Anonymous:2021:EBl

[Ano21o] Anonymous. Editorial Board. Marine Policy, 133 (??):Article 104818, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004292.

Anonymous:2021:EB

[Ano21p] Anonymous. Editorial Board. Marine Policy, 134 (??):Article 104859, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2100470X.

Anonymous:2021:F

[Ano21q] Anonymous. February 2021. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:Ja

[Ano21r] Anonymous. January 2021. Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:Jc

[Ano21s] Anonymous. July 2021. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:Jb

[Ano21t] Anonymous. June 2021. Marine Policy, 128(??):??, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:Ma

[Ano21u] Anonymous. March 2021. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:Mb

[Ano21v] Anonymous. May 2021. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:N

[Ano21w] Anonymous. November 2021. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:O

[Ano21x] Anonymous. October 2021. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:S

[Ano21y] Anonymous. September 2021. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:Aa

[Ano22a] Anonymous. April 2022. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:Ab

[Ano22b] Anonymous. August 2022. Marine Policy, 142(??):??, August 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:D

[Ano22c] Anonymous. December 2022. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:EBa

[Ano22d] Anonymous. Editorial Board. Marine Policy, 135(??): Article 104895, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X21005066.

Anonymous:2022:EBb

[Ano22e] Anonymous. Editorial Board. Marine Policy, 136 (??):Article 104965, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000124.

Anonymous:2022:EBc

[Ano22f] Anonymous. Editorial Board. Marine Policy, 137(??): Article 105001, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X22000483.

Anonymous:2022:EBd

[Ano22g] Anonymous. Editorial Board. Marine Policy, 138(??): Article 105028, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X22000756.

Anonymous:2022:EBe

[Ano22h] Anonymous. Editorial Board. Marine Policy, 139(??): Article 105063, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X22001105.

Anonymous:2022:EBf

[Ano22i] Anonymous. Editorial Board. Marine Policy, 140(??): Article 105114, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X22001610.

Anonymous:2022:EBg

[Ano22j] Anonymous. Editorial Board. Marine Policy, 141(??): Article 105150, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X2200197X.

Anonymous:2022:EBh

[Ano22k] Anonymous. Editorial Board. Marine Policy, 142(??): Article 105166, August 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X22002135.

Anonymous:2022:EBi

[Ano22l] Anonymous. Editorial Board. Marine Policy, 143 (??):Article 105246, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002937.

Anonymous:2022:EBj

[Ano22m] Anonymous. Editorial Board. Marine Policy, 144(??): Article 105270, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X22003177.

Anonymous:2022:EBk

[Ano22n] Anonymous. Editorial Board. Marine Policy, 145 (??):Article 105317, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003645.

Anonymous:2022:EB

[Ano22o] Anonymous. Editorial Board. Marine Policy, 146 (??):Article 105369, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200416X.

Anonymous:2022:F

[Ano22p] Anonymous. February 2022. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:Ja

[Ano22q] Anonymous. January 2022. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:Jc

[Ano22r] Anonymous. July 2022. Marine Policy, 141(??):??, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:Jb

[Ano22s] Anonymous. June 2022. Marine Policy, 140(??):??, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:Ma

[Ano22t] Anonymous. March 2022. Marine Policy, 137(??):??, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:Mb

[Ano22u] Anonymous. May 2022. Marine Policy, 139(??):??, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:N

[Ano22v] Anonymous. November 2022. Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:O

[Ano22w] Anonymous. October 2022. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:S

[Ano22x] Anonymous. September 2022. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:Aa

[Ano23a] Anonymous. April 2023. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:Ab

[Ano23b] Anonymous. August 2023. Marine Policy, 154(??):??, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:D

[Ano23c] Anonymous. December 2023. Marine Policy, 158(??):??, December 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:EBa

[Ano23d] Anonymous. Editorial Board. Marine Policy, 147(??): Article 105416, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X22004638.

Anonymous:2023:EBb

[Ano23e] Anonymous. Editorial Board. Marine Policy, 148 (??):Article 105487, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000143.

Anonymous:2023:EBc

[Ano23f] Anonymous. Editorial Board. Marine Policy, 149(??): Article 105541, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X23000684.

Anonymous:2023:EBd

[Ano23g] Anonymous. Editorial Board. Marine Policy, 150(??): Article 105565, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X23000921.

Anonymous:2023:EBe

[Ano23h] Anonymous. Editorial Board. Marine Policy, 151(??): Article 105617, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X23001446.

Anonymous:2023:EBf

[Ano23i] Anonymous. Editorial Board. Marine Policy, 152(??): Article 105651, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X23001781.

Anonymous:2023:EBg

[Ano23j] Anonymous. Editorial Board. Marine Policy, 153(??): Article 105677, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X2300204X.

Anonymous:2023:EBih

[Ano23k] Anonymous. Editorial Board. Marine Policy, 154(??): Article 105717, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X23002506.

Anonymous:2023:EBj

[Ano23l] Anonymous. Editorial Board. Marine Policy, 155 (??):Article 105811, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003445.

Anonymous:2023:EBk

[Ano23m] Anonymous. Editorial Board. Marine Policy, 156(??): Article 105823, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X23003561.

Anonymous:2023:EBl

[Ano23n] Anonymous. Editorial Board. Marine Policy, 157 (??):Article 105886, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004190.

Anonymous:2023:EBm

[Ano23o] Anonymous. Editorial Board. Marine Policy, 158 (??):Article 105903, December 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004360.

Anonymous:2023:F

[Ano23p] Anonymous. February 2023. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:Ja

[Ano23q] Anonymous. January 2023. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:Jc

[Ano23r] Anonymous. July 2023. Marine Policy, 153(??):??, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:Jb

[Ano23s] Anonymous. June 2023. Marine Policy, 152(??):??, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:Ma

[Ano23t] Anonymous. March 2023. Marine Policy, 149(??):??, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:Mb

[Ano23u] Anonymous. May 2023. Marine Policy, 151(??):??, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:N

[Ano23v] Anonymous. November 2023. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:O

[Ano23w] Anonymous. October 2023. Marine Policy, 156(??):??, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:S

[Ano23x] Anonymous. September 2023. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2024:Aa

[Ano24a] Anonymous. April 2024. Marine Policy, 162(??):??, April 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2024:EBa

[Ano24b] Anonymous. Editorial Board. Marine Policy, 159(??): Article 105971, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X23005043.

Anonymous:2024:EBb

[Ano24c] Anonymous. Editorial Board. Marine Policy, 160 (??):Article 106022, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000204.

Anonymous:2024:EBc

[Ano24d] Anonymous. Editorial Board. Marine Policy, 161(??): Article 106067, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X24000654.

Anonymous:2024:EBd

[Ano24e] Anonymous. Editorial Board. Marine Policy, 162(??): Article 106091, April 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X24000897.

Anonymous:2024:F

[Ano24f] Anonymous. February 2024. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2024:Ja

[Ano24g] Anonymous. January 2024. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2024:Ma

[Ano24h] Anonymous. March 2024. Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Ayostina:2022:NAB

[ANR^{+}22] Ines Ayostina, Lucentezza Napitupulu, Barakalla Robyn, Cynthia Maharani, and Daniel Murdiyarso. Network analysis of blue carbon governance process in Indonesia. Marine Policy, 137(??):Article 104955, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000021.

Abramic:2020:IMS

[ANS^{+}20] A. Abramic, N. Nogueira, P. Sepulveda, M. Cavallo, Y. Fernández-Palacios, C. Andrade, S. Kaushik, and R. Haroun. Implementation of the Marine Strategy Framework Directive in Macaronesia and synergies with the Maritime Spatial Planning process. Marine Policy, 122(??):Article 104273, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309192.

Amundsen:2020:BCB

[AO20] Vilde Steiro Amundsen and Tonje Cecilie Osmundsen. Becoming certified, becoming sustainable? Improvements from aquaculture certification schemes as experienced by those certified. Marine Policy, 119(??):Article 104097, September 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2030169X.

Asiedu:2021:FDA

[AONF21] Berchie Asiedu, Paulina Okpei, Francis Kofi Ewusie Nunoo, and Pierre Failler. A fishery in distress: an analysis of the small pelagic fishery of Ghana. Marine Policy, 129(??):Article 104500, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001111.

Afewerki:2023:IPN

[AOO^{+}23a] Samson Afewerki, Tonje Osmundsen, Marit Schei Olsen, Kristine Vedal Størkersen, Andreas Misund, and Trine Thorvaldsen. Innovation policy in the Norwegian aquaculture industry: Reshaping aquaculture production innovation networks. Marine Policy, 152(??):Article 105624, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001513.

Alati:2023:GAF

[AOO^{+}23b] Victor Mwakha Alati, Kennedy Osuka, Levy Michael Otwoma, Paul Tuda, and Lina Mtwana Nordlund. Gender analysis in fisheries: the case of the shelled mollusc fisheries in Kenya. Marine Policy, 158(??):Article 105863, December 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003962.

Anderson:2023:CAU

[AP23] Paul A. Anderson and Robert S. Pomeroy. Comparative assessment of the U.S. State of Florida’s Marine Life regulations for its commercial marine aquarium fish fishery. Marine Policy, 149(??):Article 105523, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000507.

Andrefouet:2022:IIN

[APF22] Serge Andréfouët, Mégane Paul, and A. Riza Farhan. Indonesia’s 13558 islands: a new census from space and a first step towards a One Map for Small Islands Policy. Marine Policy, 135(??):Article 104848, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004590.

Abad:2020:ISM

[APV^{+}20] Esther Abad, Maria Grazia Pennino, Julio Valeiras, Raul Vilela, José María Bellido, Antonio Punzón, and Francisco Velasco. Integrating spatial management measures into fisheries: the Lepidorhombus spp. case study. Marine Policy, 116(??):Article 103739, June 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308509.

Aquino:2023:LEU

[Aqu23] Mateus Lopes Aquino. The limits of the European Union’s fisheries agreements as sustainable development instruments: the case of Cape Verde. Marine Policy, 148(??):Article 105455, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22005024.

Ardito:2022:RCR

[AR22] Giovanni Ardito and Marzia Rovere. Racing the clock: Recent developments and open environmental regulatory issues at the International Seabed Authority on the eve of deep-sea mining. Marine Policy, 140(??):Article 105074, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2200121X.

Arnason:2023:AIS

[AR23a] Ragnar Arnason and Birgir Th. Runolfsson. Assessing the impacts of strong user rights in fisheries: the case of the Philippines marine fisheries. Marine Policy, 152(??):Article 105602, June 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2300129X.

Arnason:2023:DIS

[AR23b] Ragnar Arnason and Birgir Th. Runolfsson. Delineating the impacts of strong user rights in fisheries. Marine Policy, 155(??):Article 105666, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001938.

Arthur:2020:SSF

[Art20] Robert Ian Arthur. Small-scale fisheries management and the problem of open access. Marine Policy, 115(??):Article 103867, May 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19307043.

Arias:2023:MPI

[AS23a] Natalia M. Arias and Wolfgang B. Stotz. Management policies impact fishers’ income beyond tolerable limits: the case of the Chilean TURFs in Central Chile. Marine Policy, 157(??):Article 105834, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003676.

Asgarian:2023:PPF

[AS23b] Ali Asgarian and Alireza Soffianian. Past and potential future distribution of white mangroves in an arid estuarine environment: Integration of Maxent and CA-Markov models. Marine Policy, 147(??):Article 105345, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200392X.

Al-Subhi:2023:EON

[ASGS23] Khalid Al-Subhi, Tim Gray, and Selina M. Stead. An evaluation of Oman’s National Fisheries Development Strategy (NFDS) 2013–2020: Wealth or welfare maximisation? Marine Policy, 155(??):Article 105687, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002142.

Asche:2022:LSF

[ASGZ22] Frank Asche, Geir Sogn-Grundvåg, and Dengjun Zhang. Large-scale fisheries during the COVID-19 pandemic: the case of the oceangoing groundfish fleet in Norway. Marine Policy, 144(??):Article 105223, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002706.

Ahmmed:2020:NAM

[ASKI20] Raju Ahmmed, Begum Shahnaz Sinha, Rubina Khan, and Dewan Mazharul Islam. A needs analysis of maritime English language skills for Bangladeshi seafarers to work on-board ships. Marine Policy, 119(??):Article 104041, September 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20301044.

Amorim:2021:SSG

[ASM21a] Patrícia Amorim, Pedro Sousa, and Gui M. Menezes. Sustainability status of the grouper fishery in the Azores archipelago: a length-based approach. Marine Policy, 130(??):Article 104562, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001731.

Appleby:2021:SPO

[ASM^{+}21b] Thomas Appleby, Matthew Studley, Brendon Moorhouse, Judith Brown, Chad Staddon, and Emma Bean. Sea of possibilities: Old and new uses of remote sensing data for the enforcement of the Ascension Island marine protected area. Marine Policy, 127(??):Article 103184, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300289.

Aguado:2021:QFG

[ASV^{+}21] S. H. Aguado, I. S. Segado, M. E. S. Vidal, T. J. Pitcher, and M. E. Lam. The quality of fisheries governance assessed using a participatory, multi-criteria framework: a case study from Murcia, Spain. Marine Policy, 124(??):Article 104280, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030926X.

Armelloni:2021:ADM

[ATF^{+}21] Enrico Nicola Armelloni, Anna Nora Tassetti, Carmen Ferrà, Alessandro Galdelli, Martina Scanu, Adriano Mancini, Gianna Fabi, and Giuseppe Scarcella. AIS data, a mine of information on trawling fleet mobility in the Mediterranean Sea. Marine Policy, 129(??):Article 104571, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001822.

Akuoko:2023:RPR

[AVF^{+}23a] Ivy Serwaa Gyimah Akuoko, Jessica Vandenberg, Jill C. Falman, Karin Otsuka, Gabriel K. Lau, Marlena Skrobe, Suzy An, Elaine M. Faustman, and Yoshitaka Ota. Rethinking plastic realities in Ghana: a call for a wellbeing approach to understanding human–plastics entanglements for more equitable plastics governance. Marine Policy, 158(??):Article 105856, December 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003895.

Arcas:2023:EEC

[AVF23b] Esther Arcas, Carlos Valle, and Aitor Forcada. Evaluation of the economic crisis on the conservation of the ichthyofauna in Marine Protected Areas. Marine Policy, 147(??):Article 105347, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003943.

Angulo-Valdes:2022:MMR

[AVPAFM^{+}22] Jorge Angulo-Valdes, Fabián Pina-Amargos, Tamara Figueredo-Martin, Rod Fujita, Sepp Haukebo, Valerie Miller, Eduardo Boné-Morón, and Daniel Whittle. Managing marine recreational fisheries in Cuba for sustainability and economic development with emphasis on the tourism sector. Marine Policy, 145(??):Article 105254, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003013.

Adam:2020:PRS

[AWBC20] Issahaku Adam, Tony R. Walker, Joana Carlos Bezerra, and Andrea Clayton. Policies to reduce single-use plastic marine pollution in West Africa. Marine Policy, 116(??):Article 103928, June 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19304865.

Abbott:2021:FDE

[AWX21] Joshua K. Abbott, Daniel Willard, and Jintao Xu. Feeding the dragon: the evolution of China’s fishery imports. Marine Policy, 133(??):Article 104733, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003444.

Botta:2020:RGO

[BABC20] Robert Botta, Frank Asche, J. Scott Borsum, and Edward V. Camp. A review of global oyster aquaculture production and consumption. Marine Policy, 117(??):Article 103952, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19305433.

Byrne:2021:FII

[BAD21] Conor Byrne, Sveinn Agnarsson, and Brynhildur Davidsdottir. Fuel Intensity in Icelandic fisheries and opportunities to reduce emissions. Marine Policy, 127(??):Article 104448, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000580.

Byrne:2020:SLQ

[BADO20] Conor Byrne, Sveinn Agnarsson, Brynhildur Davidsdottir, and Maartje Oostdijk. Species-level quota concentration in the Icelandic harvesting sector. Marine Policy, 121(??):Article 104108, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20303110.

Bennett:2023:EJA

[BAF^{+}23] Nathan J. Bennett, Juan José Alava, Caroline E. Ferguson, Jessica Blythe, Elisa Morgera, David Boyd, and Isabelle M. Côté. Environmental (in)justice in the Anthropocene ocean. Marine Policy, 147(??):Article 105383, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004304.

Balton:2022:WWB

[Bal22] David Balton. What will the BBNJ agreement mean for the Arctic fisheries agreement? Marine Policy, 142(??):Article 103745, August 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X1930449X.

Baumler:2021:QII

[BAP21] Raphael Baumler, Maria Carrera Arce, and Anne Pazaver. Quantification of influence and interest at IMO in Maritime Safety and Human Element matters. Marine Policy, 133(??):Article 104746, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003572.

Braga:2022:CCT

[BASM22] Heitor O. Braga, Ulisses M. Azeiteiro, Alexandre Schiavetti, and Luísa Magalhães. Checking the changes over time and the impacts of COVID-19 on cockle (Cerastoderma edule) small-scale fisheries in Ria de Aveiro coastal lagoon, Portugal. Marine Policy, 135(??):Article 104843, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004541.

Baumler:2020:WTL

[Bau20] Raphael Baumler. Working time limits at sea, a hundredyear construction. Marine Policy, 121(??):Article 104101, November 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20301998.

Bonnett:2020:CCC

[BB20] Nicole Bonnett and S. Jeff Birchall. Coastal communities in the Circumpolar North and the need for sustainable climate adaptation approaches. Marine Policy, 121(??):Article 104175, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2030302X.

Bastardie:2021:RDC

[BB21] Francois Bastardie and Elliot John Brown. Reverse the declining course: a risk assessment for marine and fisheries policy strategies in Europe from current knowledge synthesis. Marine Policy, 126(??):Article 104409, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000191.

Bostedt:2020:BCT

[BBB^{+}20] Göran Bostedt, Charlotte Berkström, Runar Brännlund, Ola Carlén, Ann-Britt Florin, Lars Persson, and Ulf Bergström. Benefits and costs of two temporary no-take zones. Marine Policy, 117(??):Article 103883, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308108.

Barrett:2023:CCM

[BBB23] C. J. Barrett, K. Bradley, and A. Brazier. Common, curled or miscellaneous: the need for species-specific recordings of octopuses to inform stock assessments. Marine Policy, 153(??):Article 105632, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001598.

Boys:2022:WHS

[BBBS22] Rebecca M. Boys, Ngaio J. Beausoleil, Emma L. Betty, and Karen A. Stockin. When and how to say goodbye: an analysis of Standard Operating Procedures that guide end-of-life decision-making for stranded cetaceans in Australasia. Marine Policy, 138(??):Article 104949, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005601.

Bartelet:2024:ECD

[BBC24] Henry A. Bartelet, Michele L. Barnes, and Graeme S. Cumming. Estimating and comparing the direct economic contributions of reef fisheries and tourism in the Asia–Pacific.

Marine Policy, 159(??):Article 105939, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004724.

Brooks:2022:TCM

[BBGG22] Kate Brooks, Kate Barclay, R. Quentin Grafton, and Natalie Gollan. Transforming coastal and marine management: Deliberative democracy and integrated management in New South Wales, Australia. Marine Policy, 139(??):Article 104053, May 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20301470.

Barton:2023:NID

[BBGHR23] Jonathan R. Barton, Sebastián Baeza-González, Joaquín Zenteno Hopp, and Álvaro Román. Neostructural innovation and directionality in Chilean salmon aquaculture. Marine Policy, 150(??):Article 105518, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000453.

Braccini:2020:WES

[BBHN20] Matias Braccini, Nick Blay, Alastair Harry, and Stephen J. Newman. Would ending shark meat consumption in Australia contribute to the conservation of white sharks in South Africa? Marine Policy, 120(??):Article 104144, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20306370.

Bonito:2022:UEF

[BBJ^{+}22] Lindsay Bonito, Lyall Bellquist, Alexis M. Jackson, Kate Kauer, Mary G. Gleason, Jono Wilson, and Stuart Sandin. U.S. exempted fishing permits: Role, value, and lessons learned for adaptive fisheries management. Marine Policy, 138(??):Article 104992, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000392.

Baumler:2021:SFS

[BBK21a] Raphael Baumler, Bikram Singh Bhatia, and Momoko Kitada. Ship first: Seafarers’ adjustment of records on work and rest hours. Marine Policy, 130(??):Article 104186, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19309388.

Brooks:2021:RSA

[BBK^{+}21b] Cassandra M. Brooks, Evan Bloom, Andrea Kavanagh, Emily S. Nocito, George M. Watters, and John Weller. The Ross Sea, Antarctica: a highly protected MPA in international waters. Marine Policy, 134(??):Article 104795, December 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004061.

Bottema:2021:AAS

[BBO21] Mariska J. M. Bottema, Simon R. Bush, and Peter Oosterveer. Assuring aquaculture sustainability beyond the farm. Marine Policy, 132(??):Article 104658, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002694.

Bennett:2021:ARC

[BBS^{+}21] Nathan J. Bennett, Natalie C. Ban, Anna Schuhbauer, Dacotah-Victoria Splichalova, Megan Eadie, Kiera Vandeborne, Jim McIsaac, Eric Angel, Joshua Charleson, Erika R. Gavenus, Sarah Harper, Terre Satterfield, Tasha Sutcliffe, and Rashid Sumaila. Access rights, capacities and benefits in small-scale fisheries: Insights from the Pacific Coast of Canada. Marine Policy, 130(??):Article 104581, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001925.

Bennett:2021:BGB

[BBWC21] Nathan James Bennett, Jessica Blythe, Carole Sandrine White, and Cecilia Campero. Blue growth and blue justice: Ten risks and solutions for the ocean economy. Marine Policy, 125(??):Article 104387, March 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310381.

Baker:2020:EJI

[BC20a] Susan Baker and Natasha Louise Constant. Epistemic justice and the integration of local ecological knowledge for marine conservation: Lessons from the Seychelles. Marine Policy, 117(??):Article 103921, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304920.

Bisack:2020:CMM

[BC20b] Kathryn Bisack and Patricia M. Clay. Compliance with marine mammal protection: Focus groups reveal factors in commercial fishermen’s decisions. Marine Policy, 115(??):Article 103789, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19301794.

Bisack:2021:BRC

[BC21] Kathryn D. Bisack and Patricia M. Clay. Behavioral responses to competing incentives and disincentives: Compliance with marine mammal protection. Marine Policy, 132(??):Article 104674, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002852.

Berkowitz:2020:OPS

[BCB20] Heloise Berkowitz, Larry B. Crowder, and Cassandra M. Brooks. Organizational perspectives on sustainable ocean governance: a multi-stakeholder, meta-organization model of collective action. Marine Policy, 118(??):Article 104026, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19305159.

Ba:2022:SSR

[BCBS22] Aliou Ba, Christian Chaboud, Patrice Brehmer, and Jörn O. Schmidt. Are subsidies still relevant in West African artisanal small pelagic fishery? Insights from long run bioeconomic scenarios. Marine Policy, 146(??):Article 105294, December 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003414.

Birchenough:2021:VMS

[BCJ21] Sarah Elizabeth Birchenough, Patrick Arthur Cooper, and Antony Charles Jensen. Vessel monitoring systems as a tool for mapping fishing effort for a small inshore fishery operating within a marine protected area. Marine Policy, 124(??):Article 104325, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309726.

Brotz:2021:RJW

[BCMCM21] Lucas Brotz, Andrés M. Cisneros-Montemayor, and Miguel Ángel Cisneros-Mata. The race for jellyfish: Winners and losers in Mexico’s Gulf of California. Marine Policy, 134(??):Article 104775, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003869.

Baker:2023:OJT

[BCN23] Susan Baker, Natasha Constant, and Poppy Nicol. Oceans justice: Trade-offs between sustainable development goals in the Seychelles. Marine Policy, 147(??):Article 105357, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004043.

Ballinger:2020:SMP

[BD20] Rhoda C. Ballinger and Wendy Dodds. Shoreline management plans in England and Wales: a scientific and transparent process? Marine Policy, 111(??):??, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301422.

Bergseth:2023:CLA

[BD23] Brock Bergseth and Jon C. Day. Compliance — the ‘Achilles heel’ of protected areas. Marine Policy, 155 (??):Article 105728, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002610.

Brown:2021:EMI

[BDC^{+}21] Christopher J. Brown, Amelia Desbiens, Max D. Campbell, Edward T. Game, Eric Gilman, Richard J. Hamilton, Craig Heberer, David Itano, and Kydd Pollock. Electronic monitoring for improved accountability in western Pacific tuna longline fisheries. Marine Policy, 132(??):Article 104664, October 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2100275X.

Ballesteros:2023:MPA

[BDC23] Marta Ballesteros and Mark Dickey-Collas. Managing participation across boundaries: a typology for stakeholder engagement in the International Council for the Exploration of the Sea. Marine Policy, 147(??):Article 105389, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004365.

BenArfa:2022:SEA

[BDCJ22] Yessine Ben Arfa, Antonio Di Cintio, Luca Ceriola, and Othmen Jarboui. Socio-economic analysis of the trawl fleet targeting deep-water rose shrimp (Parapenaeus longirostris) and European hake (Merluccius merluccius) in North Tunisia (2015–2017). Marine Policy, 137(??):Article 104952, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21005637.

Bloor:2021:BBC

[BDD^{+}21] Isobel S. M. Bloor, Peter F. Duncan, Samuel P. Dignan, Jack Emmerson, David Beard, Fiona R. Gell, Karen McHarg, and Michel J. Kaiser. Boom not bust: Cooperative management as a mechanism for improving the commercial efficiency and environmental outcomes of regional scallop fisheries. Marine Policy, 132(??):Article 104649, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002608.

Bertheussen:2020:PDB

[BDDE20] Bernt Arne Bertheussen, Bent Magne Dreyer, Sofie Dreyer, and Susann Evenseth. Performance differences between nations exploiting a common natural resource: the Icelandic–Norwegian mackerel case. Marine Policy, 122(??):Article 104269, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309155.

Bertheussen:2021:IFE

[BDHI21] Bernt Arne Bertheussen, Bent Magne Dreyer, Øystein Hermansen, and John Roald Isaksen. Institutional and financial entry barriers in a fishery. Marine Policy, 123(??):Article 104303, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309489.

Bernardo:2020:LCA

[BdLAdC^{+}20] Cristina Bernardo, Aisni Mayumi Corrêa de Lima Adachi, Vanessa Paes da Cruz, Fausto Foresti, Robin H. Loose, and Hugo Bornatowski. The label “Cação” is a shark or a ray and can be a threatened species! Elasmobranch trade in Southern Brazil unveiled by DNA barcoding. Marine Policy, 116(??):Article 103920, June 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309261.

Baraibar-Diez:2023:TEA

[BDLO23] Elisa Baraibar-Diez, Ignacio Llorente, and María D. Odriozola. Text emotion analysis in aquaculture communication via Twitter: the case of Spain. Marine Policy, 152(??):Article 105605, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2300132X.

Barreto:2020:HDM

[BDM20] Giovanna C. Barreto, Maikon Di Domenico, and Rodrigo Pereira Medeiros. Human dimensions of marine protected areas and small-scale fisheries management: a review of the interpretations. Marine Policy, 119(??):Article 104040, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19301381.

Britton:2021:ASO

[BDM21] Easkey Britton, Christine Domegan, and Patricia McHugh. Accelerating sustainable ocean policy: the dynamics of multiple stakeholder priorities and actions for oceans and human health. Marine Policy, 124(??):Article 104333, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309829.

Bailey:2022:HNS

[BDMW22] Sabine Bailey, Kelly Heber Dunning, Daniel Morris, and Ryan D. Williamson. How narratives shape policy: Lessons learned from port projects adjacent to coral reefs in Florida and the Cayman Islands. Marine Policy, 144(??):Article 105233, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002809.

Bertheussen:2020:ESQ

[BDR20] Bernt Arne Bertheussen, Bent Dreyer, and Audun Reiertsen. Economic sustainability of quality-enhancing business models in the Norwegian cod industry. Marine Policy, 113(??):Article 103821, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19307778.

Bailey:2020:ISS

[BE20a] Jennifer L. Bailey and Sigrid Sandve Eggereide. Indicating sustainable salmon farming: the case of the new Norwegian aquaculture management scheme. Marine Policy, 117(??):Article 103925, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930452X.

Bailey:2020:MAA

[BE20b] Jennifer L. Bailey and Sigrid Sandve Eggereide. Mapping actors and arguments in the Norwegian aquaculture debate. Marine Policy, 115(??):Article 103898, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302702.

Bueger:2020:BCC

[BE20c] Christian Bueger and Timothy Edmunds. Blue crime: Conceptualising transnational organised crime at sea. Marine Policy, 119(??):Article 104067, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20300270.

Bean:2022:HHR

[Bea22] Emma Bean. Helping or hindering? The role of the public right to fish in the management of coastal fisheries. Marine Policy, 138(??):Article 104991, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000380.

Breen:2021:ICN

[BEH^{+}21] Colin Breen, Crystal El Safadi, Harmen Huigens, Sophie Tews, Kieran Westley, Georgia Andreou, Rodrigo Ortiz Vazquez, Julia Nikolaus, and Lucy Blue. Integrating cultural and natural heritage approaches to Marine Protected Areas in the MENA region. Marine Policy, 132(??):Article 104676, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002876.

Barz:2020:BDF

[BEM^{+}20] Fanny Barz, Josefa Eckardt, Steffi Meyer, Sarah B. M. Kraak, and Harry V. Strehlow. ‘Boats don’t fish, people do’ — how fishers’ agency can inform fisheries-management on bycatch mitigation of marine mammals and sea birds. Marine Policy, 122(??):Article 104268, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309143.

Bertheussen:2022:RPD

[Ber22] Bernt Arne Bertheussen. The role of path–dependent institutions during the collapse and rebuilding of a fishery. Marine Policy, 136(??):Article 104944, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005558.

Bertram:2023:PSS

[Ber23] Douglas F. Bertram. Pacific Salmon Strategy Initiative will conserve marine birds with fishery closures and gillnet license retirements. Marine Policy, 150(??):Article 105551, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000787.

Bjorndal:2021:BCQ

[BFMS21] Trond Bjørndal, Torben Foss, Gordon R. Munro, and Mogens Schou. Brexit and consequences for quota sharing in the Barents Sea cod fishery. Marine Policy, 131(??):Article 104622, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002335.

Bagoulla:2020:MTF

[BG20a] Corinne Bagoulla and Patrice Guillotreau. Maritime transport in the French economy and its impact on air pollution: an input-output analysis. Marine Policy, 116(??):Article 103818, June 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19305408.

Ballinger:2020:ERC

[BG20b] Rhoda C. Ballinger and Jeremy Gault. Editorial: Reflections on coastal climate change adaptation in North West Europe. Marine Policy, 111(??):??, January 2020.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309679.

Bering:2022:WUA

[BGK^{+}22] Janet Bering, Henry Gargan, Jess Kuesel, Maria Morrison, Claire Mullaney, Andrew J. Read, Stephen E. Roady, and Addyson Rowe. Will unilateral action improve the global conservation status of marine mammals? A first analysis of the U.S. Marine Mammal Protection Act’s Import Provisions Rule. Marine Policy, 135(??):Article 104832, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004437.

Barreiro-Gen:2021:GES

[BGLTC21] María Barreiro-Gen, Rodrigo Lozano, Melis Temel, and Angela Carpenter. Gender equality for sustainability in ports: Developing a framework. Marine Policy, 131 (??):Article 104593, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002049.

Braham:2024:RCS

[BGMCG24] Cheikh-Baye Braham, Jo Gascoigne, Carlos Montero-Castaño, and Christophe Godard. The role of certification standards in supporting improved fisheries management for small pelagics in Mauritania: an example to follow in the North West African zone? Marine Policy, 161(??):Article 106054, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000526.

Bach:2023:ICD

[BH23a] Hanna Bach and Teis Hansen. IMO off course for decarbonisation of shipping? Three challenges for stricter policy. Marine Policy, 147(??):Article 105379, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004262.

Bailey:2023:SUO

[BH23b] David M. Bailey and Charlotte R. Hopkins. Sustainable use of ocean resources. Marine Policy, 154(??):Article 105672, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001999.

Brown:2024:PTP

[BHB^{+}24] A. Ross Brown, Tim Huntington, Lisa K. Bickley, Tom Powell, and Charles R. Tyler. Positive tipping points for the sustainable growth of bivalve shellfish mariculture in England and Wales. Marine Policy, 162(??):Article 106058, April 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24000563.

Bullock:2023:ISP

[BHCL23] Simon Bullock, Eliott Higgins, James Crossan, and Alice Larkin. Improving shore power project economics at the Port of Aberdeen. Marine Policy, 152(??):Article 105625, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001525.

Ben-Hasan:2021:CFM

[BHdMCM^{+}21] Abdulrahman Ben-Hasan, Yvonne Sadovy de Mitcheson, Miguel A. Cisneros-Mata, Érica Antunes Jimenez, Moslem Daliri, Andrés M. Cisneros-Montemayor, Rekha J. Nair, Sangeetha A. Thankappan, Carl J. Walters, and Villy Christensen. China’s fish maw demand and its implications for fisheries in source countries. Marine Policy, 132(??):Article 104696, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003079.

Bastug:2022:PCD

[BHEE22] Sedat Baştuğ, Hercules Haralambides, Soner Esmer, and Enes Eminoğlu. Port competitiveness: Do container terminal operators and liner shipping companies see eye to eye? Marine Policy, 135(??):Article 104866, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004772.

Blanchard:2023:CSD

[BHKJT23] Catherine Blanchard, Ellycia Harrould-Kolieb, Emily Jones, and Michelle L. Taylor. The current status of deepsea mining governance at the International Seabed Authority. Marine Policy, 147(??):Article 105396, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004432.

Benediktsdottir:2023:AFS

[BHKK23] Brynhildur Benediktsdóttir, Gunnar Haraldsson, Helga Kristjánsdóttir, and Ögmundur Knútsson. Assessing fisheries subsidies. Marine Policy, 157(??):Article 105835, November 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003688.

Blandon:2021:SCS

[BI21] Abigayil Blandon and Hiroe Ishihara. Seafood certification schemes in Japan: Examples of challenges and opportunities from three Marine Stewardship Council (MSC) applicants. Marine Policy, 123(??):Article 104279, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309258.

Birchall:2020:CCA

[Bir20] S. Jeff Birchall. Coastal climate adaptation planning and evolutionary governance: Insights from Homer, Alaska. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18303919.

Basterretxea-Iribar:2022:BMP

[BISSBdlMM22] Imanol Basterretxea-Iribar, Iranzu Sotés, Javier Sanchez-Beaskoetxea, and María de las Mercedes Maruri. Beach management policy analysis concerning safety flag systems in Northern Spain. Marine Policy, 144(??):Article 105226, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002731.

Bausero-Jorcin:2024:APP

[BJGG^{+}24] Sofía Bausero-Jorcin, Stefan Gelcich, Ignacio Gianelli, Gabriela Jorge-Romero, Cecilia Lezama, and Omar Defeo. Assessing the performance of a participatory governance transformation in small-scale fisheries: a case study from Uruguay. Marine Policy, 160(??):Article 105964, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004979.

Bethel:2021:IPS

[BJH21] Lonna Bethel, Henning Jessen, and Johan Hollander. Implementing the Port State Measures Agreement to combat illegal, unreported and unregulated fishing in the Caribbean. Marine Policy, 132(??):Article 104643, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002542.

Bjorndal:2023:NAM

[Bjø23] Trond Bjørndal. The Northeast Atlantic and Mediterranean bluefin tuna fishery: Back from the brink. Marine Policy, 157(??):Article 105848, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003810.

Brahmana:2023:SST

[BK23] Rayenda Khresna Brahmana and Maria Kontesa. Sea surface temperature anomalies and cash holdings: Evidence from fisheries companies. Marine Policy, 148 (??):Article 105452, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004997.

Baker:2024:IAB

[BKW^{+}24] G. Barry Baker, Valeriya Komyakova, Alexia Wellbelove, Nicola Beynon, and Marcus Haward. The implementation of ACAP Best Practice Advice to mitigate seabird bycatch in fisheries: Issues and options. Marine Policy, 160(??):Article 105879, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004128.

Boyle:2022:SWE

[BL22] Apryl Boyle and Brianna Le Busque. Sharing the waves: an exploration of surfer and shark interactions. Marine Policy, 145(??):Article 105260, November 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003074.

Bai:2023:HSI

[BL23a] Jiayu Bai and Xiaoyu Li. How SCAR informs decisionmaking for Antarctic governance: Perspectives from SCAR submissions. Marine Policy, 155(??):Article 105757, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002907.

Bueger:2023:CMI

[BL23b] Christian Bueger and Tobias Liebetrau. Critical maritime infrastructure protection: What’s the trouble? Marine Policy, 155(??):Article 105772, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003056.

Baez:2020:MSF

[BLH^{+}20] José Carlos Báez, María Lourdes Ramos, Miguel Herrera, Hilario Murua, José Luis Cort, Santiago Déniz, Vanessa Rojo, Jon Ruiz, Pedro J. Pascual-Alayón, Anertz Muniategi, Alejandro Pérez San Juan, Javier Ariz, Felipe Fernández, and Francisco Abascal. Monitoring of Spanish flagged purse seine fishery targeting tropical tuna in the Indian ocean: Timeline and history. Marine Policy, 119(??):Article 104094, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307559.

Beringen:2022:APU

[BLL22] Ethan Beringen, Nengye Liu, and Michelle Lim. Australia and the pursuit of “not undermining” regional bodies at the biodiversity beyond national jurisdiction negotiations. Marine Policy, 136(??):Article 104929, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21005406.

Bjorndal:2024:CMI

[BLM24] Trond Bjørndal, Marko Lindroos, and Gordon Munro. The cooperative management of internationally shared fish stocks in a world of instability. Marine Policy, 160(??):Article 105961, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004943.

Britsch:2021:DPA

[BLS21] Melissa L. Britsch, Heather M. Leslie, and Joshua S. Stoll. Diverse perspectives on aquaculture development in Maine. Marine Policy, 131(??):Article 104697, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003080.

Bonsu:2024:CLF

[BLY^{+}24] Prince Owusu Bonsu, Jonas Letschert, Katherine L. Yates, Jon C. Svendsen, Jörg Berkenhagen, Marcel J. C. Rozemeijer, Thomas R. H. Kerkhove, Jennifer Rehren, and Vanessa Stelzenmüller. Co-location of fisheries and offshore wind farms: Current practices and enabling conditions in the North Sea. Marine Policy, 159(??):Article 105941, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004748.

Bjorndal:2021:GTP

[BM21] Trond Bjørndal and Gordon R. Munro. A game theoretic perspective on the management of shared North Sea fishery resources: Pre and post Brexit. Marine Policy, 132(??):Article 104669, October 2021. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002803.

Buscher:2021:DSB

[BMB^{+}21] Elena Buscher, Darcy L. Mathews, Cheryl Bryce, Kathleen Bryce, Darlene Joseph, and Natalie C. Ban. Differences and similarities between Indigenous and conventional marine conservation planning: the case of the Songhees Nation, Canada. Marine Policy, 129(??):Article 104520, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001317.

Barboza:2024:SSF

[BMB^{+}24] Anna Barboza, Edison D. Macusi, Elaine Q. Borazon, Mudjekeewis D. Santos, Richard N. Muallil, and Ivy M. Nallos. Small-scale fisheries (SSF) management and conservation schemes and their application in the Philippines. Marine Policy, 161(??):Article 106018, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000162.

Bernardini:2020:MPP

[BMC20] Giulia Bernardini, A. J. McConville, and Arturo Castillo Castillo. Macro-plastic pollution in the tidal Thames: an analysis of composition and trends for the optimization of data collection. Marine Policy, 119(??):Article 104064, September 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20300397.

Baker-Medard:2020:FMF

[BMF20] Merrill Baker-Médard and Jake Faber. Fins and (Mis)fortunes: Managing shark populations for sustainability and food sovereignty. Marine Policy, 113(??):Article 103805, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19301617.

Barragan-Mendez:2020:PRA

[BMGDS^{+}20] C. Barragán-Méndez, M. M. González-Duarte, I. Sobrino, Y. Vila, J. M. Mancera, and I. Ruiz-Jarabo. Physiological recovery after bottom trawling as a method to manage discards: the case study of Nephrops norvegicus and Squilla mantis. Marine Policy, 116(??):Article 103895, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308157.

Billi:2022:LCL

[BMH^{+}22] Marco Billi, Aldo Mascareño, Pablo A. Henríquez, Ignacia Rodríguez, Felipe Padilla, and Gonzalo A. Ruz. Learning from crises? The long and winding road of the salmon industry in Chiloé Island, Chile. Marine Policy, 140(??):Article 105069, June 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001166.

Blake:2022:UFL

[BMK^{+}22] Suzana Dumitrita Blake, Matthew McPherson, Mandy Karnauskas, Skyler Rose Sagarese, Adyan Rios, Amanda Daria Stoltz, Anthony Mastitski, and Michael Jepson. Use of fishermen’s local ecological knowledge to understand historic red tide severity patterns. Marine Policy, 145 (??):Article 105253, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003001.

Beveridge:2020:NSE

[BMM^{+}20] Rachelle Beveridge, Megan Moody, Grant Murray, Chris Darimont, and Bernie Pauly. The Nuxalk Sputc (Eulachon) Project: Strengthening indigenous management authority through community-driven research. Marine Policy, 119(??):Article 103971, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304506.

Bocci:2024:LSI

[BMM^{+}24] Martina Bocci, Marina Markovic, Aleš Mlakar, Margarita Stancheva, Michelle Borg, Fabio Carella, Andrea Barbanti, and Emiliano Ramieri. Land-Sea-Interactions in MSP and ICZM: a regional perspective from the Mediterranean and the Black Sea. Marine Policy, 159(??):Article 105924, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004578.

Baranano:2022:ERS

[BMMF22] Carlota Barañano, Gonzalo Méndez, María Isabel Míguez, and Emilio Fernández. Enhancing recognition of seagrass ecosystems through communication planning based on historical and present social perception studies. Marine Policy, 146(??):Article 105296, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003438.

Baco:2020:OVM

[BMR20] Amy R. Baco, Nicole B. Morgan, and E. Brendan Roark. Observations of vulnerable marine ecosystems and significant adverse impacts on high seas seamounts of the northwestern Hawaiian Ridge and Emperor Seamount Chain. Marine Policy, 115(??):Article 103834, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302611.

Botero:2023:YCZ

[BMR23] Camilo M. Botero, Celene B. Milanes, and Sebastian Robledo. 50 years of the Coastal Zone Management Act: the bibliometric influence of the first coastal management law on the world. Marine Policy, 150(??):Article 105548, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000751.

Brown:2023:MSR

[BMRM23] Robert Brown, David Molyneux, Joshua Ryan, and Alexandria Major. Maritime search and rescue in Canada and the use of emergency radio beacons. Marine Policy, 147(??):Article 105261, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003086.

Bakowski:2020:TDE

[BN20] Tomasz Bakowski and Justyna Nawrot. Transposition of the directive establishing a framework for maritime spatial planning into the polish legal order (main aspects and key solutions). Marine Policy, 117(??):Article 103946, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19303276.

Bernardino:2021:CTM

[BNF21] Angelo F. Bernardino, Gabriel N. Nóbrega, and Tiago O. Ferreira. Consequences of terminating mangrove’s protection in Brazil. Marine Policy, 125(??):Article 104389, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2031040X.

Bilgili:2024:ISW

[BÖ24] Levent Bilgili and Aykut I. Ölçer. IMO 2023 strategy — where are we and what’s next? Marine Policy, 160(??):Article 105953, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004864.

Badru:2023:SEC

[BOOO23] Gbolahan Badru, Shakirudeen Odunuga, Ademola Omojola, and Emmanuel Oladipo. Stakeholder engagement in coastal adaptation interventions: Insights from the Southwest coast of Nigeria. Marine Policy, 155 (??):Article 105690, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002178.

Botha:2020:RLF

[Bot20] Mark Botha. Reorganising the lobster fishery value chain. Marine Policy, 120(??):Article 104149, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303781.

Boviatsis:2022:LAB

[Bov22] M. Boviatsis. Legal assessment of BIMCO’s infectious or contagious diseases (IOCD) clauses for voyage and time charter parties. Marine Policy, 144(??):Article 105206, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002536.

Bishop:2022:HIG

[BOW^{+}22] Breanna Bishop, Jade Owen, Lisette Wilson, Tagalik Eccles, Aldo Chircop, and Lucia Fanning. How icebreaking governance interacts with Inuit rights and livelihoods in Nunavut: a policy review. Marine Policy, 137(??):Article 104957, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000045.

Boylan:2021:IMT

[Boy21] Brandon M. Boylan. Increased maritime traffic in the Arctic: Implications for governance of Arctic sea routes. Marine Policy, 131(??):Article 104566, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001779.

Baiju:2022:AGS

[BPA^{+}22] K. K. Baiju, Shinoj Parappurathu, S. Abhilash, C. Ramachandran, P. S. Swathi Lekshmi, Praveen Padmajan, Shelton Padua, and Thomson Kaleekal. Achieving governance synergies through institutional interactions among non-state and state actors in small-scale marine fisheries in India. Marine Policy, 138(??):Article 104990, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000379.

Borriello:2022:IMO

[BR22] Antonio Borriello and John M. Rose. The issue of microplastic in the oceans: Preferences and willingness to pay to tackle the issue in Australia. Marine Policy, 135(??):Article 104875, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004863.

Braithwaite:2022:NLB

[Bra22] Jeremy Braithwaite. This is no longer a Bristol Bay Fishery”: Fisheries dispossession and colonial violence in Bristol Bay, Alaska. Marine Policy, 143(??):Article 105172, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002196.

Belton:2021:CIA

[BRM^{+}21a] Ben Belton, Leah Rosen, Lucinda Middleton, Saadiah Ghazali, Abdullah-Al Mamun, Jacqueline Shieh, Hamia S. Noronha, Goutam Dhar, Mohammod Ilyas, Christopher Price, Ahmed Nasr-Allah, Ibrahim Elsira, Bikram K. Baliarsingh, Arun Padiyar, Suresh Rajendran, A. B. C. Mohan, Ravi Babu, Michael Joseph Akester, Ei Ei Phyo, Khin Maung Soe, Ajibola Olaniyi, Sunil N. Siriwardena, John Bostock, David C. Little, Michael Phillips, and Shakuntala H. Thilsted. COVID-19 impacts and adaptations in Asia and Africa’s aquatic food value chains. Marine Policy, 129(??):Article 104523, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001342.

Brager:2021:CSE

[BRM21b] Stefan Bräger, Gabriela Q. Romero Rodriguez, and Sandor Mulsow. The current status of environmental requirements for deep seabed mining issued by the International Seabed Authority. Marine Policy, ??(??):Article 103258, ???? 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304329.

Bosi:2023:MSP

[BRP^{+}23] Sofia Bosi, Emiliano Ramieri, Marta Picciulin, Stefano Menegon, Michol Ghezzo, Antonio Petrizzo, Thomas Folegot, Fantina Madricardo, and Andrea Barbanti. Is Maritime Spatial Planning a tool to mitigate the impacts of underwater noise? A review of adopted and upcoming maritime spatial plans in Europe. Marine Policy, 155(??):Article 105725, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002580.

Bustos:2022:IDS

[BRR^{+}22] Eduardo Bustos, Hugo Robotham, Guillermo Rodríguez, Francisco Ther-Ríos, Jorge Muñoz, Lorenzo Andrade, Nancy. Barahona, and Alejandro Clement. Indicators for the determination of the status and trends of king’s littleneck clam fishery (Ameghinomya antiqua, P. P. King, 1832) in Los Lagos Region, Chile: an integral sustainable approach. Marine Policy, 143(??):Article 105128, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001750.

Burke:2022:SPW

[BS22] Lauretta Burke and Mark Spalding. Shoreline protection by the world’s coral reefs: Mapping the benefits to people, assets, and infrastructure. Marine Policy, 146(??):Article 105311, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200358X.

Baeta:2021:LTT

[BSBD21] Marc Baeta, Marco Antonio Solís, Manuel Ballesteros, and Omar Defeo. Long-term trends in striped venus clam (Chamelea gallina) fisheries in the western Mediterranean Sea: the case of Ebro Delta (NE Spain). Marine Policy, 134(??):Article 104798, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004097.

Bhowmik:2021:RSS

[BSI^{+}21] Joy Bhowmik, Samiya Ahmed Selim, Haseeb Md. Irfanullah, Jannat Shancharika Shuchi, Rumana Sultana, and Shaikh Giasuddin Ahmed. Resilience of smallscale marine fishers of Bangladesh against the COVID-19 pandemic and the 65-day fishing ban. Marine Policy, 134(??):Article 104794, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2100405X.

Brown:2020:PPF

[BSK^{+}20] Stephen K. Brown, Manoj Shivlani, Roberto F. Koeneke, David Agnew, Julia Byrd, Martin Cryer, Cathy Dichmont, David Die, William Michaels, Keith Reid, Jake Rice, Henrik Sparholt, and James Weinberg. Patterns and practices in fisheries assessment peer review systems. Marine Policy, 117(??):Article 103880, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301241.

Banikoi:2023:UTM

[BSM23] Hudu Banikoi, Achim Schlüter, and Aisa O. Manlosa. Understanding transformations in the marine coastal realm: the explanatory potential of theories of institutional change. Marine Policy, 155(??):Article 105791, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300324X.

Bellido:2020:IVO

[BSSL^{+}20] Jose M. Bellido, U. Rashid Sumaila, Jose Luis Sánchez-Lizaso, Maria Lourdes Palomares, and Daniel Pauly. Input versus output controls as instruments for fisheries management with a focus on Mediterranean fisheries. Marine Policy, 118(??):Article 103786, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303720.

Bidwell:2023:ASO

[BST23] David Bidwell, Tiffany Smythe, and Grant Tyler. Anglers’ support for an offshore wind farm: Fishing effects or clean energy symbolism. Marine Policy, 151(??):Article 105568, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000957.

Bennett:2020:OTS

[BSY^{+}20] Mia M. Bennett, Scott R. Stephenson, Kang Yang, Michael T. Bravo, and Bert De Jonghe. The opening of the Transpolar Sea Route: Logistical, geopolitical, environmental, and socioeconomic impacts. Marine Policy, 121(??):Article 104178, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2030453X.

Brugere:2021:MTF

[BTE21] Cecile Brugere, Max Troell, and Hampus Eriksson. More than fish: Policy coherence and benefit sharing as necessary conditions for equitable aquaculture development. Marine Policy, 123(??):Article 104271, January 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309179.

Browne:2022:MEO

[BTV^{+}22] Thomas Browne, Trung Tien Tran, Brian Veitch, Doug Smith, Faisal Khan, and Rocky Taylor. A method for evaluating operational implications of regulatory constraints on Arctic shipping. Marine Policy, 135(??):Article 104839, January 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004504.

Breve:2024:FWR

[BUM^{+}24] Niels W. P. Brevé, Kateryna Urbanovych, AlberTinka J. Murk, Paul A. M. van Zwieten, Leopold A. J. Nagelkerke, and Marloes Kraan. Fishers’ willingness to report incidental bycatches of endangered, threatened and protected fish species: the case of European sturgeon in the Northeast Atlantic Ocean. Marine Policy, 162(??):Article 106056, April 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2400054X.

Begum:2022:APL

[BUR^{+}22] Amany Begum, Md Kutub Uddin, Md Mizanur Rahman, Md Mostafa Shamsuzzaman, and Mohammad Mahmudul Islam. Assessing policy, legal and institutional frameworks of marine megafauna conservation in Bangladesh. Marine Policy, 143(??):Article 105187, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002342.

Bertheussen:2021:IBR

[BV21] Bernt Arne Bertheussen and Terje Vassdal. Institutionbased roots to fishing vessels profitability. Marine Policy, 123(??):Article 104286, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309325.

Brander:2020:GCB

[BvBN^{+}20] Luke M. Brander, Pieter van Beukering, Lynn Nijsten, Alistair McVittie, Corinne Baulcomb, Florian V. Eppink, and Jorge Amrit Cado van der Lelij. The global costs and benefits of expanding Marine Protected Areas. Marine Policy, 116(??):Article 103953, June 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302386.

Banach:2020:FSD

[BvdBvdFK20] J. L. Banach, S. W. K. van den Burg, and H. J. van der Fels-Klerx. Food safety during seaweed cultivation at offshore wind farms: an exploratory study in the North Sea. Marine Policy, 120(??):Article 104082, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301500.

Brugnach:2023:EAC

[BvdH23] Marcela Brugnach and Ronald van den Hoek. Embracing ambiguity in climate change adaptation for more effective responses to new uncertain shorescapes conditions. Marine Policy, 152(??):Article 105626, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001537.

Browne:2021:CMA

[BVT^{+}21] Thomas Browne, Brian Veitch, Rocky Taylor, Jennifer Smith, Doug Smith, and Faisal Khan. Consequence modelling for Arctic ship evacuations using expert knowledge. Marine Policy, 130(??):Article 104582, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001937.

Bachmann-Vargas:2021:RFS

[BVvKL21] Pamela Bachmann-Vargas, C. S. A. (Kris) van Koppen, and Machiel Lamers. Re-framing salmon aquaculture in the aftermath of the ISAV crisis in Chile. Marine Policy, 124(??):Article 104358, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310095.

Bertheussen:2020:SIC

[BXV20] Bernt Arne Bertheussen, Jinghua Xie, and Terje Vassdal. Strategic investments in catch capacity and quotas: How costly is a mismatch for a firm? Marine Policy, 117(??):Article 103874, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307092.

Carlson:2020:ACS

[CA20] Jordan Thomas Carlson and Michelle Adams. Assessing the consistency of in-stream tidal energy development policy in Nova Scotia, Canada. Marine Policy, 113(??):Article 103743, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X1930048X.

Caner:2021:SSS

[CA21] Halil Ibrahim Caner and Cevdet Coşkun Aydin. Shipyard site selection by raster calculation method and AHP in GIS environment, İskenderun, Turkey. Marine Policy, 127(??):Article 104439, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100049X.

Ceylan:2023:NFA

[CAC23] Bulut Ozan Ceylan, Demir Ali Akyar, and Mehmet Serdar Celik. A novel FMEA approach for risk assessment of air pollution from ships. Marine Policy, 150(??):Article 105536, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000635.

Cardoso-Andrade:2021:ALS

[CACJR^{+}21] Mariana Cardoso-Andrade, Frederico Cruz-Jesus, Francisco Castro Rego, Mafalda Rangel, and Henrique Queiroga. Assessing the land- and seascape determinants of recreational diving: Evidence for Portugal’s south coast. Marine Policy, 123(??):Article 104285, January 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309313.

Caviedes:2022:ICZ

[CAGBM22] Verónica Caviedes, Pedro Arenas-Granados, and Juan Manuel Barragán-Muñoz. Integrated Coastal Zone Management on a transnational area: the Gulf of Honduras. Marine Policy, 136(??):Article 104931, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2100542X.

Cid-Aguayo:2024:SFS

[CAHR^{+}24] Beatriz Eugenia Cid-Aguayo, Alfonso Henríquez, Loreto Arias Ramírez, Marcelo Durán Pérez, Chris Harrod, and Daniel Gomez-Uchida. Socioecology of farmed salmon escapes:

’commons’ and de-privatization of escaped salmon for better management of a potentially invasive species. Marine Policy, 160(??):Article 105935, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004682.

Cevenini:2023:AWI

[CAM^{+}23] Fabio Cevenini, Barnaby Andrews, Angela Muench, Philip Lamb, Silvia Ferrini, and Kieran Hyder. Assessing the welfare impacts of changes in recreational fisheries management: a modelling approach for European sea bass. Marine Policy, 148(??):Article 105408, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004559.

Catma:2020:NMV

[Cat20] Serkan Catma. Non-market valuation of beach quality: Using spatial hedonic price modeling in Hilton Head Island, SC. Marine Policy, 115(??):Article 103866, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303100.

Cavole:2020:ULE

[CAVJ^{+}20] Leticia Maria Cavole, Solange Andrade-Vera, José R. Marin Jarrin, Daniela Faggiani Dias, Octavio Aburto-Oropeza, and María José Barrágan-Paladines. Using local ecological knowledge of fishers to infer the impact of climate variability in Galápagos’ small-scale fisheries. Marine Policy, 121(??):Article 104195, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20300439.

Creach:2020:VCA

[CBAPM20] Axel Creach, Emilio Bastidas-Arteaga, Sophie Pardo, and Denis Mercier. Vulnerability and costs of adaptation strategies for housing subjected to flood risks: Application to La Guérinière France. Marine Policy, 117(??):Article 103438, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16306303.

Cramer:2023:ISP

[CBD^{+}23] Lori A. Cramer, Jennifer Beaullieu, Jamie Doyle, Marta Maldonado, Hillary Egna, Maria Johnson, and Flaxen D. L. Conway. The importance of the seafood processing sector to coastal community resilience. Marine Policy, 156(??):Article 105797, October 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003305.

Campos:2021:ULT

[CBO^{+}21] Leonardo Campos, Fernando Berrios, Rómulo Oses, Jorge E. González, and Estefanía Bonnail. Unravelling Lessonia trabeculata management in coastal areas of the Atacama region of northern Chile through a DPSIR approach: Insights for sustainable plans. Marine Policy, 133(??):Article 104737, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003481.

Calderwood:2021:TET

[CBR21] Julia Calderwood, Guillaume Bal, and David G. Reid. Trial and error: Tactical changes in fishing behaviour can help reduce discards and exposure to chokes, but scientific trials can fail to spot this. Marine Policy, 124(??):Article 104365, February 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20310162.

Canales:2021:ESB

[CC21] Cristian M. Canales and Luis A. Cubillos. Empirical survey-based harvest control rules in a transboundary small pelagic fishery under recruitment regime shifts: the case of the northern Chilean–southern Peruvian anchovy. Marine Policy, 134(??):Article 104784, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100395X.

Cao:2022:OCC

[CC22] Xingguo Cao and Yen-Chiang Chang. The opening of cabotage: China’s trials and challenges. Marine Policy, 143(??):Article 105174, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002214.

Cao:2023:DLB

[CC23] Xingguo Cao and Yen-Chiang Chang. Developing the legal basis for non-state actors in China’s ocean governance. Marine Policy, 155(??):Article 105727, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002609.

Cardinale:2021:MFM

[CCB^{+}21] Massimiliano Cardinale, Francesco Colloca, Angelo Bonanno, Giuseppe Scarcella, Enrico Arneri, Angélique Jadaud, Claire Saraux, Salvatore Aronica, Simona Genovese, Marco Barra, Gualtiero Basilone, Silvia Angelini, Fabio Falsone, Vita Gancitano, Alberto Santojanni, Fabio Fiorentino, Giacomo Milisenda, Matteo Murenu, Tommaso Russo, Piera Carpi, Beatriz Guijarro, José Luis Pérez Gil, María González, Pedro Torres, Ana Giráldez, Cristina García, Antonio Esteban, Encarnación García, Miguel Vivas, Enric Massutí, Francesc Ordines, Antoni Quetglas, and Juan Gil Herrera. The Mediterranean fishery management: a call for shifting the current paradigm from duplication to synergy. Marine Policy, 131(??):Article 104612, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002232.

Chuang:2021:EKD

[CCCC21] Min-Ta Chuang, Wen-Hsuan Chou, Chia-Hsuan Chang, and Wen-Lun Chou. Examining the key drivers for regional revitalization based upon social network analysis: a case study of Badouzi in Taiwan. Marine Policy, 133(??):Article 104754, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003651.

Carvalho:2021:PMC

[CCGM21] Natacha Carvalho, John Casey, Jordi Guillen, and Jann Th. Martinsohn. Profitability and management costs in the EU Northeast Atlantic fisheries. Marine Policy, 123(??):Article 104281, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309271.

Caputi:2021:SRE

[CCK^{+}21] Nick Caputi, Arani Chandrapavan, Mervi Kangas, Simon de Lestang, Anthony Hart, Danielle Johnston, and James Penn. Stock-recruitment-environment relationships of invertebrate resources in Western Australia and their link to pro-active management harvest strategies. Marine Policy, 133(??):Article 104728, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003390.

Chumchuen:2022:CTM

[CCKK22] Watcharapong Chumchuen, Shiela Villamor Chumchuen, Krisada Kajonrit, and Kraison Krueajun. COVID-19 and Thai marine capture fishery in the Gulf of Thailand: a case of small-scale fishery versus industrial fishery. Marine Policy, 144(??):Article 105240, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002871.

Capurso:2023:TMM

[CCS23] Ginevra Capurso, Brendan Carroll, and Kathryn A. Stewart. Transforming marine monitoring: Using eDNA metabarcoding to improve the monitoring of the Mediterranean Marine Protected Areas network. Marine Policy, 156(??):Article 105807, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003408.

Cronin:2021:PPO

[CCW21] Yvonne Cronin, Valerie Cummins, and Eric Wolsztynski. Public perception of offshore wind farms in Ireland. Marine Policy, 134(??):Article 104814, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004255.

Cao:2023:SCT

[CCX23] Xingguo Cao, Leyi Chang, and Houqun Xing. Securing the clean title of ships: the prospects and achievements of the Convention on the International Effects of Judicial Sales of Ships. Marine Policy, 154(??):Article 105685, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002129.

Castrejon:2023:RLB

[CD23] Mauricio Castrejón and Omar Defeo. Reconsidering the longline ban in the Galapagos Marine Reserve. Marine Policy, 151(??):Article 105589, May 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001161.

Conley:2024:ASP

[CD24] Elizabeth D. Conley and Emily P. Diamond. Angler’s sense of place as an indicator for perceived vulnerability to shifting stock distributions. Marine Policy, 161(??):Article 106007, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000058.

Crosman:2020:EFM

[CDB20] Katherine M. Crosman, Natalie A. Dowling, and Ann Bostrom. The effects of Fishpath, a multi-stakeholder decision-support tool, on stakeholder buy-in to management in data-limited fisheries. Marine Policy, 122 (??):Article 104215, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20308617.

Christiansen:2021:PED

[CDC21a] Bernd Christiansen, Anneke Denda, and Sabine Christiansen. Potential effects of deep seabed mining on pelagic and benthopelagic biota. Marine Policy, ??(??):Article 103442, ???? 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X18306407.

Clark:2021:EIA

[CDC21b] Malcolm R. Clark, Jennifer M. Durden, and Sabine Christiansen. Environmental Impact Assessments for deep-sea mining: Can we improve their future effectiveness? Marine Policy, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307309.

Carlesi:2023:EPM

[CDD^{+}23] Lorenzo Carlesi, Emilia Cubero Dudinskaya, Roberto Danovaro, Gianfranco D'Onghia, Serena Mandolesi, Simona Naspetti, and Raffaele Zanoli. Estimating preferences for Mediterranean deep-sea ecosystem services: a discrete choice experiment. Marine Policy, 151(??):Article 105593, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001203.

Chavez:2021:PSF

[CDG^{+}21] Carlos Chávez, Jorge Dresdner, Nuria González, Mauricio Leiva, and Miguel Quiroga. The performance of shared fish stock fisheries under varying institutional and socioeconomic conditions: Evidence from the South Eastern Pacific Anchoveta Fishery. Marine Policy, 125(??):Article 104262, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309088.

Chen:2022:MCE

[CDHY22] Xiaolong Chen, Qianbin Di, Zhiwen Hou, and Zhe Yu. Measurement of carbon emissions from marine fisheries and system dynamics simulation analysis: China’s northern marine economic zone case. Marine Policy, 145(??):Article 105279, November 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003268.

Caputi:2021:SGC

[CdLN^{+}21] Nick Caputi, Simon de Lestang, Stephen J. Newman, Gary Jackson, and Kim Smith. Stakeholder-government collaboration in developing cost-effective fishery-independent surveys in rights-based and co-managed fisheries. Marine Policy, 128(??):Article 104510, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001214.

Camara:2023:SEC

[CDNS23] Fatoumata Camara, Idrissa Diedhiou, Waly Ndiaye, and Alassane Sarr. Socioeconomic effects of the COVID-19 pandemic on artisanal fishing in Senegal: Adaptation strategies of actors in the fishing ports of Dakar and Mbour. Marine Policy, 153(??):Article 105633, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001604.

Chatterjee:2022:BBB

[CDP22] Poushali Chatterjee, Rajarshi Dasgupta, and Ashis Kumar Paul. Beach beauty in Bengal: Perception of scenery and its implications for coastal management in Purba Medinipur district, eastern India. Marine Policy, 139(??):Article 105034, May 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000811.

Christodoulou:2022:OLS

[CDRS22] Anastasia Christodoulou, Dimitrios Dalaklis, Peter Raneri, and Rebecca Sheehan. An overview of the legal search and rescue framework and related infrastructure along the Arctic Northeast Passage. Marine Policy, 138(??):Article 104985, April 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2200032X.

Cochrane:2020:DSE

[CES20] Kevern L. Cochrane, Jessica Eggers, and Warwick H. H. Sauer. A diagnosis of the status and effectiveness of marine fisheries management in South Africa based on two representative case studies. Marine Policy, 112(??):Article 103774, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19307006.

Campbell:2021:BEB

[CFM^{+}21] Lisa M. Campbell, Luke Fairbanks, Grant Murray, Joshua S. Stoll, Linda D’Anna, and Julia Bingham. From Blue Economy to Blue Communities: reorienting aquaculture expansion for community wellbeing. Marine Policy, 124(??):Article 104361, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20310125.

Ch:2022:ESC

[CG22] María Catalina García Ch. and Joyeeta Gupta. Environmental and sociocultural claims within maritime boundary disputes. Marine Policy, 139(??):Article 105043, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000902.

Chikudza:2020:PPI

[CGGA20] Lynnlee Chikudza, Claire Gauzente, Patrice Guillotreau, and Karen A. Alexander. Producer perceptions of the incentives and challenges of adopting ecolabels in the European finfish aquaculture industry: a Q-methodology approach. Marine Policy, 121(??):Article 104176, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20303146.

Christiansen:2021:EGD

[CGHV21] S. Christiansen, H. Ginzky, K. Houghton, and A. Vink. Environmental governance of deep seabed mining — Scientific insights and food for thought. Marine Policy, ??(??):Article 103827, ???? 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20300646.

Cucinelli:2023:ERG

[CGP23] Jessica Cucinelli, Floris Goerlandt, and Ronald Pelot. Exploring risk governance deficits of maritime Search and Rescue in Canada. Marine Policy, 149(??):Article 105511, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000386.

Chan:2020:EIP

[Cha20a] Hing Ling Chan. Economic impacts of Papahānaumokuākea Marine National Monument expansion on the Hawaii longline fishery. Marine Policy, 115(??):Article 103869, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19300090.

Chang:2020:BRM

[Cha20b] Yen-Chiang Chang. Book review: Managing the Risk of Offshore Oil and Gas Accidents — The International Legal Dimension, Gunther Handl, Kristoffer Svendsen (Eds.). Edward Elgar (2019), 482, ISBN 978-1-78643-673-3. Marine Policy, 111(??):Article 103698, January 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307122.

Chang:2021:UFD

[Cha21] Yen-Chiang Chang. The use of force during law enforcement in disputed maritime areas. Marine Policy, 124(??):Article 104382, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20310332.

Chang:2022:ADI

[Cha22] Yen-Chiang Chang. Accelerating domestic and international dual circulation in the Chinese shipping industry — an assessment from the aspect of participating RCEP. Marine Policy, 135(??):Article 104862, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004735.

Chan:2024:VCO

[Cha24] Nicholas Chan. Voluntary commitments for ocean sustainability: Comparing the UN Ocean Conference and Our Ocean Conference platforms. Marine Policy, 160 (??):Article 105956, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2300489X.

Chang:2020:SFJ

[CHB^{+}20] Shui-Kai Chang, Nien-Tsu Alfred Hu, Samsudin Basir, Hai Vu Duyen, Praulai Nootmorn, Mudjekeewis D. Santos, and Fayakun Satria. A step forward to the joint management of the South China Sea fisheries resources: Joint works on catches, management measures and conservation issues. Marine Policy, 116(??):Article 103716, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305500.

Curran:2023:HMS

[CHB23] Kaitlyn Curran, Kayla M. Hamelin, and Megan Bailey. Humanizing marine spatial planning: a salutogenic approach. Marine Policy, 154(??):Article 105660, August 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001872.

Collins:2021:MIL

[CHC^{+}21] M. A. Collins, P. R. Hollyman, J. Clark, M. Soeffker, O. Yates, and R. A. Phillips. Mitigating the impact of longline fisheries on seabirds: Lessons learned from the South Georgia Patagonian toothfish fishery (CCAMLR Subarea 48.3). Marine Policy, 131(??):Article 104618, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002293.

Chen:2021:MLD

[Che21a] Yong Chen. Measuring litter distribution on UK beaches. Marine Policy, 130(??):Article 104592, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002037.

Chen:2021:RCD

[Che21b] Yuli Chen. Reconciling common but differentiated responsibilities principle and no more favourable treatment principle in regulating greenhouse gas emissions from international shipping. Marine Policy, 123(??):Article 104317, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309647.

Campos:2021:DST

[CHEC21] Aida Campos, Victor Henriques, Karim Erzini, and Margarida Castro. Deep-sea trawling off the Portuguese continental coast — Spatial patterns, target species and impact of a prospective EU-level ban. Marine Policy, 128(??):Article 104466, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21000774.

Choi:2022:AND

[Cho22] Junghwan Choi. Assessing the need for the designation of the Yellow Sea Particularly Sensitive Sea Area (PSSA). Marine Policy, 137(??):Article 104971, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000185.

Christensen:2020:SCF

[Chr20] Lotte Dalgaard Christensen. Seaweed cultivation in the Faroe Islands: Analyzing the potential for forward and fiscal linkages. Marine Policy, 119(??):Article 104015, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20301342.

Cockerell:2021:GAS

[CJ21] Lucy M. Cockerell and Peter J. S. Jones. Governance analysis of St Anne Marine National Park, Seychelles. Marine Policy, 127(??):Article 103912, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306426.

Chung:2022:IMP

[CJ22] Huey-Shian Elly Chung and Juei-Cheng Jao. Improving marine protected area governance: Concerns and possible solutions from Taiwan’s practice. Marine Policy, 140(??):Article 105078, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001257.

Chantavanich:2023:PSP

[CJ23] Supang Chantavanich and Waranya Jitpong. Precarity and social protection of migrant fishers in Thailand: Case studies of death and disappearance at sea. Marine Policy, 155(??):Article 105743, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002762.

Choy:2022:UTW

[CJCS^{+}22] Christina Pei Pei Choy, Rima W. Jabado, Naomi Clark-Shen, Danwei Huang, Min Yi Choo, and Madhu Rao. Unraveling the trade in wedgefishes and giant guitarfishes in Singapore. Marine Policy, 136(??):Article 104914, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2100525X.

Chen:2023:FSM

[CJL^{+}23] Fenggui Chen, Yuhuan Jiang, Zhenghua Liu, Ruijuan Lin, and Wei Yang. Framework system of marine sustainable development assessment based on systematic review. Marine Policy, 154(??):Article 105689, August 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002166.

Chang:2023:MLC

[CK23] Yen-Chiang Chang and Muneeb Khan. The Maritime Labour Convention 2006 in human rights context: an appraisal. Marine Policy, 154(??):Article 105688, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002154.

Choi:2020:PPC

[CKY20] Kyung-Ran Choi, Ju-Hee Kim, and Seung-Hoon Yoo. Public perspective on constructing sea forests as a public good: a contingent valuation experiment in South Korea. Marine Policy, 120(??):Article 104146, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302979.

Choi:2022:PWP

[CKY22] Kyung-Ran Choi, Ju-Hee Kim, and Seung-Hoon Yoo. The public willingness to pay for the research and demonstration of tidal stream energy in South Korea. Marine Policy, 138(??):Article 104981, April 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000288.

Cormier:2021:RAD

[CL21] Roland Cormier and Jemma Londsdale. Risk assessment for deep sea mining: an overview of risk. Marine Policy, ??(??):Article 103485, ???? 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306109.

Lee:2022:PSF

[cL22] Hsiao chien Lee. Promoting slow fish education in southern Taiwan coastal areas: an empirical case study of five elementary schools. Marine Policy, 138(??):Article 104995, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000422.

Clark:2020:IAN

[Cla20] Nichola A. Clark. Institutional arrangements for the new BBNJ agreement: Moving beyond global, regional, and hybrid. Marine Policy, 122(??):Article 104143, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19308796.

Collins:2020:UPE

[CLB^{+}20] Claire Collins, Tom Bech Letessier, Annette Broderick, Isuru Wijesundara, and Ana Nuno. Using perceptions to examine human responses to blanket bans: the case of the thresher shark landing-ban in Sri Lanka. Marine Policy, 121(??):Article 104198, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20303973.

Collins:2023:VBE

[CLB^{+}23] Claire Collins, Tom B. Letessier, Aloka Benaragama, Annette Broderick, Isuru Wijesundara, Dilhara Wijetunge, and Ana Nuno. Valuable bycatch: Eliciting social importance of sharks in Sri Lanka through value chain analysis. Marine Policy, 157(??):Article 105832, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003652.

Choi:2021:WPP

[CLC21] Eun Chul Choi, Joo Suk Lee, and Jeong-In Chang. Willingness to pay for the prevention of beach erosion in Korea: the case of Haeundae beach. Marine Policy, 132(??):Article 104667, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002785.

Custodio:2021:HNM

[CLCV21] Marco Custódio, Ana I. Lillebø, Ricardo Calado, and Sebastián Villasante. Halophytes as novel marine products — a consumers’ perspective in Portugal and policy implications. Marine Policy, 133(??):Article 104731, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003420.

Cox:2024:MTG

[CLD^{+}24] Kieran D. Cox, Audrey Looby, Hailey L. Davies, Kelsie A. Murchy, Brittnie Spriel, Aaron N. Rice, Francis Juanes, and Isabelle M. Côté. Military training in the Canadian Pacific: Taking aim at critical habitat or sufficient mitigation of noise pollution impacts? Marine Policy, 160(??):Article 105945, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004785.

Carter:2020:DAW

[CLL20] David W. Carter, Sabrina J. Lovell, and Christopher Liese. Does angler willingness-to-pay for changes in harvest regulations vary by state? Results from a choice experiment in the Gulf of Mexico. Marine Policy, 121(??):Article 104196, November 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20302487.

Chang:2022:LIR

[CLL22a] Yen-Chiang Chang, Xinyi Liu, and Shuang Liu. Legal issues regarding the establishment of an offshore data collection system — a practice from China. Marine Policy, 140(??):Article 105077, June 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001245.

Chen:2022:RAS

[CLL^{+}22b] Yuli Chen, Naiyuan Lou, Guangli Liu, Yang Luan, and Han Jiang. Risk analysis of ship detention defects based on association rules. Marine Policy, 142(??):Article 105123, August 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001701.

Claudet:2021:CGP

[CLP21] Joachim Claudet, Charles Loiseau, and Antoine Pebayle. Critical gaps in the protection of the second largest exclusive economic zone in the world. Marine Policy, 124(??):Article 104379, February 2021. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20310307.

Ceccarelli:2022:EMP

[CLRW22] Daniela M. Ceccarelli, Asri P. Lestari, Rudyanto, and Alan T. White. Emerging marine protected areas of eastern Indonesia: Coral reef trends and priorities for management. Marine Policy, 141(??):Article 105091, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001385.

Chang:2020:CSC

[CLS20] Yen-Chiang Chang, Hefei Liu, and Yue Sun. Current Situation and Challenges in the law of armed conflict at sea: Conference report. Marine Policy, 113(??):Article 103779, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19308814.

Campos:2023:SPF

[CLS^{+}23] Aida Campos, Pedro Leitão, Luís Sousa, Patrícia Gaspar, and Victor Henriques. Spatial patterns of fishing activity inside the Gorringe bank MPA based on VMS, AIS and elogbooks data. Marine Policy, 147(??):Article 105356, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004031.

Chen:2021:MRA

[CLW^{+}21] Fenggui Chen, Jinwen Liu, Jianwei Wu, Nurul Dhewani Mirah Sjafrie, Puji Rahmadi, and Respaty Yudha Putranto. Measuring the relationship among stakeholders from valuesatisfaction-demand in the development of ecotourism of Marine Park. Marine Policy, 129(??):Article 104519, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001305.

Chang:2023:CGM

[CLZ23] Yen-Chiang Chang, Shuang Liu, and Xu Zhang. The construction of Global Maritime Capital — current development in China. Marine Policy, 151(??):Article 105576, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001033.

Curtin:2023:SIS

[CM23] Richard Curtin and Neil McCullough. The socioeconomic impact of the seafood sector at Ireland’s main ports. Marine Policy, 152(??):Article 105627, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001549.

Cappell:2022:RFE

[CMC22] Rod Cappell, Graeme MacFadyen, and Andrew Constable. Research funding and economic aspects of the Antarctic krill fishery. Marine Policy, 143(??):Article 105200, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002470.

Cavraro:2023:SSF

[CMC^{+}23] Francesco Cavraro, Marco Anelli Monti, Alberto Caccin, Fabio Fiori, Fabio Grati, Elisabetta Russo, Giuseppe Scarcella, Dario Vrdoljak, Sanja Matić-Skoko, and Fabio Pranovi. Is the small-scale fishery more sustainable in terms of GHG emissions? A case study analysis from the Central Mediterranean Sea. Marine Policy, 148(??):Article 105474, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000015.

Crespo:2020:BSS

[CMD^{+}20] Guillermo Ortuño Crespo, Joanna Mossop, Daniel Dunn, Kristina Gjerde, Elliott Hazen, Gabriel Reygondeau, Robin Warner, Derek Tittensor, and Patrick Halpin. Beyond static spatial management: Scientific and legal considerations for dynamic management in the high seas. Marine Policy, 122(??):Article 104102, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309248.

Canovas-Molina:2020:UWC

[CMGF20] Almudena Cánovas-Molina and Eduardo García-Frapolli. Untangling worldwide conflicts in marine protected areas: Five lessons from the five continents. Marine Policy, 121(??):Article 104185, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20304097.

Carvalho:2022:MOM

[CMI^{+}22] Laila M. Carvalho, Miguel Mies, Kelly Y. Inagaki, Eduardo G. Sanches, Marcelo R. Souza, Guilherme O. Longo, and Acácio R. G. Tomás. The marine ornamental market in Brazil (Southwestern Atlantic) frequently trades prohibited and endangered species, and threatens the ecosystem role of cleaning mutualism. Marine Policy, 146(??):Article 105305, December 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003529.

Cawthorn:2021:VLL

[CMNH21] Donna-Mareè Cawthorn, Torie E. Murphy, Amanda M. Naaum, and Robert H. Hanner. Vague labelling laws and outdated fish naming lists undermine seafood market transparency in Canada. Marine Policy, 125(??):Article 104335, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309842.

Cisneros-Montemayor:2020:CNF

[CMOB^{+}20] Andrés M. Cisneros-Montemayor, Yoshitaka Ota, Megan Bailey, Christina C. Hicks, Ahmed S. Khan, Anthony Rogers, U. Rashid Sumaila, John Virdin, and Kevin K. He. Changing the narrative on fisheries subsidies reform: Enabling transitions to achieve SDG 14.6 and beyond. Marine Policy, 117(??):Article 103970, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302529.

Czermanski:2024:MPA

[CMOJ^{+}24] Ernest Czermański, Magdalena Matczak, Aneta Oniszczuk-Jastrzabek, Karolina Krośnicka, and Joanna Witkowska.

MSP for port areas — to what extent should we interfere with governance of ports’ waters? Case study of Polish seaports. Marine Policy, 160(??):Article 105985, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23005183.

Cantillo:2021:ALM

[CMR21] Javier Cantillo, Juan Carlos Martín, and Concepción Román. Assessing the label’s mandatory information for fishery and aquaculture products in the EU28. A consumer approach based on a consistent fuzzy preference relation with geometric Bonferroni mean. Marine Policy, 128(??):Article 104515, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001263.

Cadman:2020:SVC

[CMS20] Rachael Cadman, Bertrum H. MacDonald, and Suzuette S. Soomai. Sharing victories: Characteristics of collaborative strategies of environmental non-governmental organizations in Canadian marine conservation. Marine Policy, 115(??):Article 103862, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19304038.

Cvitanovic:2021:UEB

[CMS^{+}21] C. Cvitanovic, M. Mackay, R. J. Shellock, E. I. van Putten, D. B. Karcher, and M. Dickey-Collas. Understanding and evidencing a broader range of ‘successes’ that can occur at the interface of marine science and policy. Marine Policy, 134(??):Article 104802, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004139.

Cisneros-Montemayor:2022:CCW

[CMSN^{+}22] Andrés M. Cisneros-Montemayor, Hussain Sinan, Tu Nguyen, José María Da Rocha, U. Rashid Sumaila, Daniel J. Skerritt, Anna Schuhbauer, Enrique Sanjurjo, and Megan Bailey. A constructive critique of the World Trade Organization draft agreement on harmful fisheries subsidies. Marine Policy, 135(??):Article 104872, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004838.

Capello:2023:DSB

[CMT^{+}23] Manuela Capello, Gorka Merino, Mariana Tolotti, Hilario Murua, and Laurent Dagorn. Developing a science-based framework for the management of drifting Fish Aggregating Devices. Marine Policy, 153(??):Article 105657, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001847.

Chavez-Molina:2023:PSG

[CMWN^{+}23] Vasco Chavez-Molina, Daniel Wagner, Emily S. Nocito, Michelle Benedum, Carlos F. Gaymer, Duncan Currie, Emily Golden Beam, and Cassandra M. Brooks. Protecting the Salas y Gomez and Nazca Ridges: a review of policy pathways for creating conservation measures in the international waters of the Southeast Pacific. Marine Policy, 152(??):Article 105594, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001215.

Chen:2024:BDN

[CMWZ24] Dongxing Chen, Guizhen Mu, Xutao Wang, and Yafeng Zhang. Brief discussion on the new approach of selection and designation of ocean dumping areas in China. Marine Policy, 160(??):Article 105949, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004827.

Chapsos:2023:MSG

[CN23] Ioannis Chapsos and Elizabeth Ann Norman. Is maritime security gender-blind? Marine Policy, 147(??): Article 105399, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X22004468.

Coleman:2023:COL

[CNB^{+}23] Felicia C. Coleman, José A. C. C. Nunes, Áthila A. Bertoncini, Leonardo S. Bueno, Matheus O. Freitas, Maíra Borgonha, Jonas R. Leite, Márcio J. C. A. Lima-Júnior, Beatrice Ferreira, Bianca Bentes, Christopher C. Koenig, Christopher R. Malinowski, Eduardo G. Sanches, Maurício Hostim-Silva, and Cláudio L. S. Sampaio. Controversial opening of a limited fishery for Atlantic goliath grouper in the United States: Implications for population recovery. Marine Policy, 155(??):Article 105752, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002853.

Cummins:2023:GRO

[CNG^{+}23] Gabrielle H. Cummins, Matthew L. Navarro, Kingsley Griffin, Julian Partridge, and Tim J. Langlois. A global review of ocean ecosystem accounts and their data: Lessons learned and implications for marine policy. Marine Policy, 153(??):Article 105636, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300163X.

Canales:2024:MSE

[COJE24] Cristian M. Canales, Gonzalo Olea, Viviana Jurado, and Miguel Espíondola. Management strategies evaluation (MSE) in a mixed and multi-specific fishery based on indicator species: an example of small pelagic fish in Ecuador. Marine Policy, 162(??):Article 106044, April 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000423.

Czermanski:2022:IEE

[COJS^{+}22] Ernest Czermański, Aneta Oniszczuk-Jastrzabek, Eugen F. Spangenberg, Łukasz Kozłowski, Magdalena Adamowicz, Jakub Jankiewicz, and Giuseppe T. Cirella. Implementation of the Energy Efficiency Existing Ship Index: an important but costly step towards ocean protection. Marine Policy, 145(??):Article 105259, November 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003062.

Czermanski:2021:PNC

[COJZ^{+}21] Ernest Czermański, Aneta Oniszczuk-Jastrzabek, Jacek Zaucha, Barbara Pawłowska, Magdalena Matczak, and Łukasz Szydłowski. Preconditions of new container terminal location in the Maritime Spatial Planning framework. A case study for the Central Port Concept in Gdańsk. Marine Policy, 130(??):Article 104585, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001962.

Collier:2020:ECC

[Col20] Chantal E. Collier. Enabling conditions for communitybased comanagement of marine protected areas in the United States. Marine Policy, 122(??):Article 104244, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20308903.

Cooper:2020:CDV

[COM20] J. A. G. Cooper, M. C. O’Connor, and S. McIvor. Coastal defences versus coastal ecosystems: a regional appraisal. Marine Policy, 111(??):Article 102332, January 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300331.

Clifton:2021:RCD

[COSS21] Julian Clifton, Eslam O. Osman, David J. Suggett, and David J. Smith. Resolving conservation and development tensions in a small island state: a governance analysis of Curieuse Marine National Park, Seychelles. Marine Policy, 127(??):Article 103617, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304737.

Carvalho:2020:SSS

[CPBO20] Adriana Rosa Carvalho, Maria Grazia Pennino, Jose Maria Bellido, and George Olavo. Small-scale shrimp fisheries bycatch: a multi-criteria approach for data-scarse situations. Marine Policy, 116(??):Article 103613, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307164.

Calderwood:2021:TTM

[CPR21] Julia Calderwood, Debbi Pedreschi, and David G. Reid. Technical and tactical measures to reduce unwanted catches in mixed fisheries: Do the opinions of Irish fishers align with management advice? Marine Policy, 123(??):Article 104290, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309362.

Calado:2021:EKB

[CPV^{+}21] H. Calado, C. Pegorelli, M. Vergílio, C. Hipólito, A. Campos, F. Moniz, A. C. Costa, C. Pereira da Silva, C. Fonseca, C. Frazão Santos, D. Gabriel, J. Guerreiro, A. J. F. Gil, D. Johnson, K. Ng, M. M. Monwar, M. A. Ventura, J. L. Suárez de Vivero, M. Pinho, P. Borges, M. Caña-Varona, and E. A. Papaioannou. Expert knowledgebased co-development of scenarios for maritime spatial planning in the Northeast Atlantic. Marine Policy, 133(??):Article 104741, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003523.

Chen:2020:EME

[CQ20] Xuan Chen and Weiwen Qian. Effect of marine environmental regulation on the industrial structure adjustment of manufacturing industry: an empirical analysis of China’s eleven coastal provinces. Marine Policy, 113(??):Article 103797, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19303744.

Cariou:2023:SPI

[CR23] Pierre Cariou and Laingo M. Randrianarisoa. Stakeholders’ participation at the IMO Marine Environmental Protection Committee. Marine Policy, 149(??):Article 105506, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000337.

Craik:2021:IAM

[Cra21] Neil Craik. Implementing adaptive management in deep seabed mining: Legal and institutional challenges. Marine Policy, ??(??):??, ???? 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18303178.

Cousido-Rocha:2023:DEF

[CRCP^{+}23] Marta Cousido-Rocha, Marta González Carballo, Maria Grazia Pennino, Marta Coll, and José C. Báez. Differential effect of fisheries to the COVID-19 pandemic in the region of Andalusia (Spain). Marine Policy, 148(??):Article 105442, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22004894.

Coulson:2022:CPB

[CRJ22] Peter G. Coulson, Karina L. Ryan, and Gary Jackson. Are charter and private-boat recreational fishers learning to live with shark depredation? Marine Policy, 141(??):Article 105096, July 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001439.

Cross:2021:CQM

[Cro21] Helen Cross. Challenging quota market efficiencies: a case-study of Scotland (United Kingdom). Marine Policy, 126(??):Article 104384, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20310356.

Calderwood:2020:HMC

[CRP^{+}20] Julia Calderwood, Marianne Robert, Lionel Pawlowski, Youen Vermard, Zachary Radford, Thomas L. Catchpole, and David G. Reid. Hotspot mapping in the Celtic Sea: an interactive tool using multinational data to optimise fishing practices. Marine Policy, 116(??):Article 103511, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18307693.

Chang:2020:RLM

[CS20] Yen-Chiang Chang and Xinxiang Shi. Rule of law and the marine environmental networks: Conference report. Marine Policy, 113(??):Article 103778, March 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309315.

Chen:2021:BEB

[CS21] Tai-An Peter Chen and Yi-Che Shih. Blue economy based on local DNA in Taiwan: Marine regional revitalisation under the collaboration of the local and central government. Marine Policy, 132(??):Article 104668, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002797.

Coronado:2020:WWV

[CSCGC20] Eva Coronado, Silvia Salas, María Fernanda Cepeda-González, and Ratana Chuenpagdee. Who’s who in the value chain for the Mexican octopus fishery: Mapping the production chain. Marine Policy, 118(??):Article 104013, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19308784.

Cutler:2022:TSV

[CSGC22] Matthew Cutler, Angela Silva, Lauren Gentile, and Lisa Colburn. Tracking shifts in the vulnerability and resiliency of commercial fishing vessel crew and hired captains in New England and the Mid-Atlantic. Marine Policy, 138(??):Article 104980, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000276.

Cretella:2023:TOH

[CSH23] Agnese Cretella, Cordula Scherer, and Poul Holm. Tasting the ocean: How to increase ocean literacy using seafood heritage with a visceral approach. Marine Policy, 149(??):Article 105476, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000039.

Clark-Shen:2020:CSF

[CSHO^{+}20] Naomi Clark-Shen, Amanda Hsiao, Leon Oei, Madhu Rao, Ning-Yi Sun, and Alexander Douglas. A comparative study of fishery and environment laws in the south China sea: Utilizing existing laws to promote peace. Marine Policy, 121(??):Article 103568, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303756.

Cucuzza:2021:ETP

[CSL21] Marina Cucuzza, Joshua S. Stoll, and Heather M. Leslie. Evaluating the theoretical and practical linkages between ecosystem-based fisheries management and fisheries comanagement. Marine Policy, 126(??):Article 104390, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20310411.

Castro-Santos:2020:MOS

[CSLGFV20] Laura Castro-Santos, María Isabel Lamas-Galdo, and Almudena Filgueira-Vizoso. Managing the oceans: Site selection of a floating offshore wind farm based on GIS spatial analysis. Marine Policy, 113(??):Article 103803, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19303707.

Calado:2024:UDM

[CSQ^{+}24] H. Calado, C. Frazão Santos, A. Quintela, C. Fonseca, and D. Gutierrez. The ups and downs of maritime spatial planning in Portugal. Marine Policy, 160(??):Article 105984, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23005171.

Costa:2023:FAR

[CSVC23] H. S. Carvalho Costa, R. Scachetti, M. Vivacqua, and R. Choueri. A framework to assess resilience attributes based on local knowledge. Marine Policy, 155 (??):Article 105766, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002993.

Conroy:2023:SMP

[CTC^{+}23] Katherine Conroy, Daniel Tonnes, Max Calloway, Nicole Naar, Kalloway Page, and Betsy Peabody. Shoreline master programs: a hindrance or a help to kelp conservation in Puget Sound? Marine Policy, 155(??):Article 105771, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003044.

Chalastani:2021:BAP

[CTCD21] Vasiliki I. Chalastani, Vasiliki K. Tsoukala, Harry Coccossis, and Carlos M. Duarte. A bibliometric assessment of progress in marine spatial planning. Marine Policy, 127(??):Article 104329, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309763.

Cavanagh:2021:UIA

[CTH^{+}21] Rachel D. Cavanagh, Philip N. Trathan, Simeon L. Hill, Jess Melbourne-Thomas, Michael P. Meredith, Philip Hollyman, Björn A. Krafft, Monica M. C. Muelbert, Eugene J. Murphy, Martin Sommerkorn, John Turner, and Susie M. Grant. Utilising IPCC assessments to support the ecosystem approach to fisheries management within a warming Southern Ocean. Marine Policy, 131(??):Article 104589, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002001.

Copeland:2023:RCF

[CTH^{+}23] Kellen Copeland, Bastian Thomsen, Michael Harte, Shaozeng Zhang, David Fennell, and David Lewis. Rethinking 21st Century Fisheries Management: posthuman multispecies livelihoods approach. Marine Policy, 152(??):Article 105623, June 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001501.

Cheng:2021:SFS

[CTIH21] Yan-Tzong Cheng, Yun-Ching Tseng, Yoko Iwaki, and Michael C. Huang. Sustainable food security in Small Island Developing States (SIDS): a case of Horticulture project in Marshall Islands. Marine Policy, 128(??):Article 104378, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20310290.

Condie:2022:IPA

[CVA22a] Corrine M. Condie, Joanna Vince, and Karen A. Alexander. Increasing polarisation in attitudes to aquaculture: Evidence from sequential government inquiries. Marine Policy, 136(??):Article 104867, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004784.

Condie:2022:LTE

[CVA22b] Corrine M. Condie, Joanna Vince, and Karen A. Alexander. The long-term evolution of news media in defining socio-ecological conflict: a case study of expanding aquaculture. Marine Policy, 138(??):Article 104988, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000355.

Calado:2022:SSM

[CVCV^{+}22] Helena Calado, Marta Vergílio, Mario Caña-Varona, Camila Pegorelli, Cláudia Hipólito, Aida Silva, Gilberto Carreira, Maria Paramio, and Eva Papaioannou. Strategic scenarios for maritime spatial planning in an European outermost region — the case of the Azores. Marine Policy, 145(??):Article 105255, November 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003025.

Chen:2020:EAP

[CVH^{+}20] Wenting Chen, Kristof Albert M. Van Assche, Stephen Hynes, Trine Bekkby, Hartvig C. Christie, and Hege Gundersen. Ecosystem accounting’s potential to support coastal and marine governance. Marine Policy, 112(??):Article 103758, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19302933.

Ciravegna:2024:HCM

[CvHF^{+}24] E. Ciravegna, L. van Hoof, C. Frier, F. Maes, H. B. Rasmussen, A. Soete, and S. W. K. van den Burg. The hidden costs of multi-use at sea. Marine Policy, 161(??):Article 106017, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000150.

Carter:2020:AOS

[CWB^{+}20] Amber L. Carter, A. Meriwether W. Wilson, Maximiliano Bello, E. Mauricio Hoyos-Padilla, Mark E. Inall, James T. Ketchum, Andrew Schurer, and Alexander W. Tudhope. Assessing opportunities to support coral reef climate change refugia in MPAs: a case study at the Revillagigedo Archipelago. Marine Policy, 112(??):Article 103769, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19301654.

Cobacho:2020:ISR

[CWKE20] Sara P. Cobacho, Sonja Wanke, Zoi Konstantinou, and Ghada El Serafy. Impacts of shellfish reef management on the provision of ecosystem services resulting from climate change in the Dutch Wadden Sea. Marine Policy, 119(??):Article 104058, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306876.

Chang:2020:LPC

[CWKW20] Yen-Chiang Chang, Chuanliang Wang, Mehran Idris Khan, and Nannan Wang. Legal practices and challenges in addressing climate change and its impact on the oceans — a Chinese perspective. Marine Policy, 111(??):Article 103355, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18304925.

Conejo-Watt:2021:FPB

[CWMM^{+}21] Heather Conejo-Watt, Angela Muench, Stephen C. Mangi, Keith Jeffery, and Kieran Hyder. Fishers perspectives on the barriers for the English inshore fleet to diversify into aquaculture. Marine Policy, 131(??):Article 104610, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002219.

Cremers:2020:SMC

[CWR20] Klaudija Cremers, Glen Wright, and Julien Rochette. Strengthening monitoring, control and surveillance of human activities in marine areas beyond national jurisdiction: Challenges and opportunities for an international legally binding instrument. Marine Policy, 122(??):Article 103976, December 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19309534.

Chang:2020:ILS

[CZW20] Yen-Chiang Chang, Chao Zhang, and Nannan Wang. The international legal status of the unmanned maritime vehicles. Marine Policy, 113(??):Article 103830, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304774.

Do:2023:GFG

[DA23] Huu-Luat Do and Claire W. Armstrong. Ghost fishing gear and their effect on ecosystem services — identification and knowledge gaps. Marine Policy, 150(??):Article 105528, April 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000556.

Danielsen:2021:TTB

[DAA21] Rannvá Danielsen, Christopher M. Anderson, and Sveinn Agnarsson. Trawling for triple bottom line results: Applying the Fishery Performance Indicators in the Faroe Islands. Marine Policy, 125(??):Article 104250, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308964.

Dutta:2021:BET

[DAAP21] Sachinandan Dutta, Ibtisam Al-Abri, and Sourav Paul. Bioeconomic trends of Hilsa (Tenualosa ilisha) fishery: Perspectives of transboundary management between India and Bangladesh. Marine Policy, 128(??):Article 104483, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21000944.

Drinkwin:2023:ILF

[DAHS^{+}23] Joan Drinkwin, Kyle Antonelis, Madison Heller-Shipley, Paul Rudell, Michael Etnier, Tom Good, Anna Elz, and Jason Morgan. Impacts of lost fishing nets in the U.S. portion of the Salish Sea. Marine Policy, 148(??):Article 105430, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22004778.

Damato:2024:EFI

[Dam24] Nikol Damato. Evaluating factors influencing compliance with vessel regulations to minimize disturbances to endangered killer whales (Orcinus orca). Marine Policy, 161(??):Article 106036, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000344.

Dias:2023:VVS

[DAN^{+}23] Ana Carolina Esteves Dias, Derek Armitage, Prateep Kumar Nayak, Shehu Latunji Akintola, Evans Kwasi Arizi, Ratana Chuenpagdee, Basanta Kumar Das, Sabiha Ahmed Diba, Raktima Ghosh, Moenieba Isaacs, Gazi Md Nurul Islam, Alioune Kane, Yinji Li, Moffat Mzama Manase, Ahmadou Aly Mbaye, Paul Onyango, Sarmistha Pattanaik, Aliou Sall, Indah Susilowati, Sika Abrokwah, Denis Worlanyo Aheto, Olajumoke Akiode, Afful Chineboaba Araba, Foluke Omotayo Areola, Frederick Ato Armah, Esinam Attipoe, Aliyu Hamza Danagalan, Piyashi DebRoy, Alyne Delaney, Mariama Dieng, Kafayat A. Fakoya, Aini Nur Furoida, Cheikh Amadou Bamba Gaye, Adama Gueye, Fama Gueye, Mafaniso Hara, Haseeb Md. Irfanullah, Taofeek Iyanda, Kungwan Juntarashote, Ayorinde Kappo, Ishmael B. M. Kosamu, Hapsari Ayu Kusumawardhani, Tracey Lee, Fatou Gueye, Jenia Mukherjee, Tamano Namikawa, Sachiko Harada, Friday J. Njaya, Pranaya Kumar Parida, Betty Ngui Chiew Pieng, Made Ika Prasetyadewi, Erwin Prayogi, Daniel Quilter, Syed Tauheed Raihan, Aparna Roy, Alberta Ama Sagoe, Madalo Esther Samati, Clement Sambou, Revarunan Sammogam, Alassane Sarr, Khady Yama Sarr, Suvaluck Satumanatpan, Izumi Seki, Samiya A. Selim, Amrita Sen, Oumar Sow, I. Gusti Lanang Tantra, Suphakarn Traesupap, Vanessa Warren, Batuli Mohammed Yahya, Thamasak Yeemin, Sevil Berenji, Jessica Blythe, Alida Bundy, and Simron Singh. From vulnerability to viability: a situational analysis of small-scale fisheries in Asia and Africa. Marine Policy, 155(??):Article 105731, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002646.

Dauvergne:2023:NJF

[Dau23] Peter Dauvergne. The necessity of justice for a fair, legitimate, and effective treaty on plastic pollution. Marine Policy, 155(??):Article 105785, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003184.

Dayan:2024:CMC

[Day24] Eleanor Dayan. Conceptualizing maritime constabulary effectiveness (MCE). Marine Policy, 161(??):Article 106047, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24000459.

Djosetro:2020:BLS

[DB20] Marijem Djosetro and Jelle Hendrik Behagel. Building local support for a coastal protected area: Collaborative governance in the Bigi Pan Multiple Use Management Area of Suriname. Marine Policy, 112(??):Article 103746, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18306286.

Dolman:2022:IWC

[DBB^{+}22] Sarah J. Dolman, Charlotte N. Breen, Philippa Brakes, Andy Butterworth, and Simon J. Allen. The individual welfare concerns for small cetaceans from two bycatch mitigation techniques. Marine Policy, 143(??):Article 105126, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001737.

Devenport:2021:IRI

[DBBL21] Emilie Devenport, Esther Brooker, Adam Brooker, and Chris Leakey. Insights and recommendations for involving young people in decision making for the marine environment. Marine Policy, 124(??):Article 104312, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309593.

Diggon:2021:MPP

[DBH^{+}21] Steve Diggon, Caroline Butler, Aaron Heidt, John Bones, Russ Jones, and Craig Outhet. The Marine Plan Partnership: Indigenous community-based marine spatial planning. Marine Policy, 132(??):Article 103510, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305268.

Delgado:2022:HLM

[DBH^{+}22] James P. Delgado, Michael L. Brennan, Sergio A. Rapu Haoa, Julianna H. Rapu Leong, Carlos F. Gaymer, Diego Carabias, Emily Stokes, and Daniel Wagner. The hidden landscape: Maritime cultural heritage of the Salas y Gómez and Nazca ridges with implications for conservation on the high seas. Marine Policy, 136(??):Article 104877, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004887.

Do:2022:POO

[DBHT22] Thang Nam Do, Paul J. Burke, Llewelyn Hughes, and Ta Dinh Thi. Policy options for offshore wind power in Vietnam. Marine Policy, 141(??):Article 105080, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001270.

Dbouk:2022:EBP

[Dbo22] Wassim Dbouk. The Environment Bill’s policy statement on environmental principles: a golden opportunity for a principled framework for the sustainable management of UK waters. Marine Policy, 136(??):Article 104920, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005315.

Davis:2022:FRA

[DBS22a] Kaleigh E. Davis, Fiona Beaty, and Carolina Sánchez. A flexible risk assessment framework for marine plastic pollution that synthesizes waste management and ecological impact data. Marine Policy, 137(??):Article 104833, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004449.

Diggon:2022:MPP

[DBS^{+}22b] Steve Diggon, John Bones, Charles J. Short, Joanna L. Smith, Megan Dickinson, Kelly Wozniak, Karen Topelko, and Kylee A. Pawluk. The marine plan partnership for the North Pacific Coast — MaPP: a collaborative and coled marine planning process in British Columbia. Marine Policy, 142(??):Article 104065, August 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306657.

Dong:2024:NUW

[DBZ24] Bingying Dong, Lowell Bautista, and Ling Zhu. Navigating uncharted waters: Challenges and regulatory solutions for flag state jurisdiction of Maritime Autonomous Surface Ships under UNCLOS. Marine Policy, 161(??):Article 106039, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2400037X.

Daly:2020:FUF

[DC20] Jack Daly and Ratana Chuenpagdee. Facing an uncertain future: a prospective analysis of trade policy on coastal communities in northwest Newfoundland. Marine Policy, 117(??):Article 103890, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306475.

Duan:2023:RBG

[DC23] Xingyi Duan and Yen-Chiang Chang. The relationship between the general principles of international law and UN-CLOS: Conference report. Marine Policy, 150(??):Article 105552, April 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000799.

Donkersloot:2020:KCD

[DCC^{+}20] Rachel Donkersloot, Jesse Coleman, Courtney Carothers, Danielle Ringer, and Paula Cullenberg. Kin, community, and diverse rural economies: Rethinking resource governance for Alaska rural fisheries. Marine Policy, 117(??):Article 103966, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19300910.

Deng:2023:MPT

[DČFW23] Yuncheng Deng, Anica Čuka, Yanyan Fu, and Jiangyue Wu. Multiple paths towards Eco islands and blue development: Conference report. Marine Policy, 149(??):Article 105526, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000532.

Dvoretsky:2022:RAR

[DD22] Alexander G. Dvoretsky and Vladimir G. Dvoretsky. Renewal of the amateur red king crab fishery in Russian waters of the Barents Sea: Potential benefits and costs. Marine Policy, 136(??):Article 104916, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005273.

Clippele:2023:EAS

[DDAB^{+}23] Laurence H. De Clippele, Laura Alonso Díaz, Dominic A. Andradi-Brown, Muhammad Erdi Lazuardi, Mohamad Iqbal, Imam Musthofa Zainudin, Derta Prabuning, Ruben van Hooidonk, Amehr Hakim, Firdaus Agung, Agus Dermawan, and Sebastian J. Hennige. Evaluating annual severe coral bleaching risk for marine protected areas across Indonesia. Marine Policy, 148(??):Article 105428, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004754.

Dowarah:2022:PSA

[DDD22] Kaushik Dowarah, Hemashree Duarah, and Suja P. Devipriya. A preliminary survey to assess the awareness, attitudes/behaviours, and opinions pertaining to plastic and microplastic pollution among students in India. Marine Policy, 144(??):Article 105220, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002676.

DeAngelis:2020:STB

[DDF^{+}20] Paola De Angelis, Lorenzo D’Andrea, Simone Franceschini, Stefano Cataudella, and Tommaso Russo. Strategies and trends of bottom trawl fisheries in the Mediterranean Sea. Marine Policy, 118(??):Article 104016, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301263.

Doria:2021:UID

[DDGdS^{+}21] Carolina Rodriguesda Costa Doria, Jynessa Dutka-Gianelli, Suelen Taciane Brasil de Sousa, Jingjie Chu, and Taryn Marie Garlock. Understanding impacts of dams on the smallscale fisheries of the Madeira River through the lens of the Fisheries Performance Indicators. Marine Policy, 125(??):Article 104261, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309076.

DeSanto:2020:BRM

[De 20] Elizabeth De Santo. Book review: Marine protected areas: Science, policy and management, John Humphreys, Robert W. E. Clark (Eds.). Elsevier, Amsterdam (2020). Marine Policy, 121(??):Article 104151, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20307181.

DeBeukelaer:2021:CBC

[De 21] Christiaan De Beukelaer. COVID-19 border closures cause humanitarian crew change crisis at sea. Marine Policy, 132(??):Article 104661, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002724.

DeLucia:2022:BNE

[De 22] Vito De Lucia. The BBNJ negotiations and ecosystem governance in the Arctic. Marine Policy, 142(??): Article 103756, August 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19306025.

Dolman:2021:INT

[DER^{+}21] S. J. Dolman, P. G. H. Evans, F. Ritter, M. P. Simmonds, and J. Swabe. Implications of new technical measures regulation for cetacean bycatch in European waters. Marine Policy, 124(??):Article 104320, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309672.

Devereux:2022:SIR

[Dev22] Helen Devereux. Seafarer injuries in relation to time into tour of duty. Marine Policy, 135(??):Article 104865, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004760.

Danovaro:2020:TMS

[DFC^{+}20] R. Danovaro, E. Fanelli, M. Canals, T. Ciuffardi, M.-C. Fabri, M. Taviani, M. Argyrou, E. Azzurro, S. Bianchelli, A. Cantafaro, L. Carugati, C. Corinaldesi, W. P. de Haan, A. Dell’Anno, J. Evans, F. Foglini, B. Galil, M. Gianni, M. Goren, S. Greco, J. Grimalt, Q. Güell-Bujons, A. Jadaud, L. Knittweis, J. L. Lopez, A. Sanchez-Vidal, P. J. Schembri, P. Snelgrove, S. Vaz, L. Angeletti, M. Barsanti, J. A. Borg, M. Bosso, A. Brind’Amour, G. Castellan, F. Conte, I. Delbono, F. Galgani, G. Morgana, S. Prato, A. Schirone, and E. Soldevila. Towards a marine strategy for the deep Mediterranean Sea: Analysis of current ecological status. Marine Policy, 112(??):Article 103781, February 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305524.

Drakeford:2023:BSF

[DFF23] Benjamin M. Drakeford, Andy Forse, and Pierre Failler. Biodegradability and sustainable fisheries: the case for static gear in the UK Channel fishery. Marine Policy, 155(??):Article 105774, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2300307X.

Drakeford:2020:IFT

[DFTK20] Benjamin M. Drakeford, Pierre Failler, Ben Toorabally, and Elyes Kooli. Implementing the Fisheries Transparency Initiative: Experience from the Seychelles. Marine Policy, 119(??):Article 104060, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309285.

Crespo:2021:VCM

[dFVCGdS21] Maria de Fátima Vieira Crespo, Jaíra Maria Alcobaça Gomes, and Regina Oliveira da Silva. Value chain of the mangrove crab (Ucides cordatus): a case study of the Parnaíba Delta, Northeast Brazil. Marine Policy, 131(??):Article 104642, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002530.

Du:2020:ESE

[DG20] Yuan-Wei Du and Kun Gao. Ecological security evaluation of marine ranching with AHP-entropy-based TOP-SIS: a case study of Yantai, China. Marine Policy, 122(??):Article 104223, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20308691.

Dewan:2023:RCS

[DG23] Mohammud Hanif Dewan and Radu Godina. Roles and challenges of seafarers for implementation of energy efficiency operational measures onboard ships. Marine Policy, 155(??):Article 105746, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002798.

Dewan:2024:OSE

[DG24a] Mohammud Hanif Dewan and Radu Godina. An overview of seafarers’ engagement and training on energy efficient operation of ships. Marine Policy, 160(??):Article 105980, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23005134.

Dewan:2024:USA

[DG24b] Mohammud Hanif Dewan and Radu Godina. Unveiling seafarers’ awareness and knowledge on energy-efficient and low-carbon shipping: a decade of IMO regulation enforcement. Marine Policy, 161(??):Article 106037, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000356.

Dagtekin:2022:CAI

[DGG22] Murat Dağtekin, Ali Cemal Gücü, and Yaşar Genç. Concerns about illegal, unreported and unregulated fishing, carbon footprint, and the impact of fuel subsidy — an economic analysis of the Black Sea anchovy fishery. Marine Policy, 140(??):Article 105067, June 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001142.

Diez:2022:SAB

[DGIS^{+}22] A. De la Torriente Diez, J. M. González-Irusta, A. Serrano, R. Aguilar, F. Sánchez, M. Blanco, and A. Punzón. Spatial assessment of benthic habitats vulnerability to bottom fishing in a Mediterranean seamount. Marine Policy, 135(??):Article 104850, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004619.

Dupont:2020:EUM

[DGMF^{+}20] Clément Dupont, Françoise Gourmelon, Catherine Meur-Ferec, Frédérick Herpers, and Christophe Le Visage. Exploring uses of maritime surveillance data for marine spatial planning: a review of scientific literature. Marine Policy, 117(??):Article 103930, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308930.

Dixon:2024:IRT

[DGS^{+}24] Maximilian Dixon, Gaetano Grilli, Bryce D. Stewart, Rosalind H. Bark, and Silvia Ferrini. The importance of rebuilding trust in fisheries governance in post-Brexit England. Marine Policy, 161(??):Article 106034, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000320.

Davis:2022:TFC

[DH22] Ruth A. Davis and Quentin Hanich. Transparency in fisheries conservation and management measures. Marine Policy, 136(??):Article 104088, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301172.

Deely:2022:OVD

[DHC22] John Deely, Stephen Hynes, and Mary Cawley. Overseas visitor demand for marine and coastal tourism. Marine Policy, 143(??):Article 105176, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002238.

Garcia-del-Hoyo:2021:GCB

[dHJTGO21] Juan José García del Hoyo, Ramón Jiménez-Toribio, and Félix García-Ordaz. Granger causality between the canning sector and the Spanish tuna fleet: Evidence from the Toda–Yamamoto approach. Marine Policy, 132(??):Article 104701, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003122.

Dudayev:2023:PFG

[DHR23] Rayhan Dudayev, Lugas Lukmanul Hakim, and Indah Rufiati. Participatory fisheries governance in Indonesia: Are octopus fisheries leading the way? Marine Policy, 147(??):Article 105338, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003852.

DeValck:2023:VES

[DJC^{+}23] Jeremy De Valck, Diane Jarvis, Anthea Coggan, Ella Schirru, Petina Pert, Victoria Graham, and Maxine Newlands. Valuing ecosystem services in complex coastal settings: an extended ecosystem accounting framework for improved decision-making. Marine Policy, 155 (??):Article 105761, September 2023. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002944.

Daly:2021:CCB

[DKKS21] Jack Daly, Christine Knott, Poppy Keogh, and Gerald G. Singh. Changing climates in a blue economy: Assessing the climate-responsiveness of Canadian fisheries and oceans policy. Marine Policy, 131(??):Article 104623, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002347.

deKlerk:2021:SPA

[dKMK21] Yvette de Klerk, Michael Ekow Manuel, and Momoko Kitada. Scenario planning for an autonomous future: a comparative analysis of national preparedness of selected countries. Marine Policy, 127(??):Article 104428, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000385.

Dahood:2020:PSL

[DKW20] Adrian Dahood, Emily S. Klein, and George M. Watters. Planning for success: Leveraging two ecosystem models to support development of an Antarctic marine protected area. Marine Policy, 121(??):Article 104109, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309145.

Dahlet:2021:FCM

[DL21] Lol Iana Dahlet and José Luis Sánchez Lizaso. Fisheries co-management, past and present: From the Plan Castellón (1961–1966) for Spanish Mediterranean trawling fisheries to the current EU fisheries policy. Marine Policy, 128(??):Article 104480, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000919.

Du:2022:CFI

[DL22] Yuan-Wei Du and Xiao-Xue Li. Critical factor identification of marine ranching ecological security with hierarchical DEMATEL. Marine Policy, 138(??):Article 104982, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2200029X.

Pineiro-Antelo:2020:FSG

[dlÁPAFGO20] María de los Ángeles Piñeiro-Antelo, Jesús Felicidades-García, and Brendan O’Keeffe. The FLAG scheme in the governance of EU coastal areas. The cases of Ireland and Galicia (Spain). Marine Policy, 112(??):Article 103424, February 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X18302045.

Garcia-de-la-Fuente:2020:CCC

[dlFGFFR^{+}20] Laura García de-la Fuente, Lucía García-Flórez, Ma. Pino Fernández-Rueda, Jorge Alcázar-Álvarez, Arturo Colina-Vuelta, Esteban Fernández-Vázquez, and Carmen Ramos-Carvajal. Comparing the contribution of commercial and recreational marine fishing to regional economies in Europe. An input-output approach applied to Asturias (Northwest Spain). Marine Policy, 118(??):Article 104024, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19306980.

Dacks:2020:DBV

[DLJ^{+}20] Rachel Dacks, Staci A. Lewis, Philip A. S. James, Lincy L. Marino, and Kirsten L. L. Oleson. Documenting baseline value chains of Palau’s nearshore and offshore fisheries prior to implementing a large-scale marine protected area. Marine Policy, 117(??):Article 103754, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301782.

delaPuente:2022:ASL

[dlPdlLLC^{+}22] Santiago de la Puente, Rocío López de la Lama, Camila Llerena-Cayo, Benny R. Martínez, Gonzalo Rey-Cama, Villy Christensen, María Rivera-Ch, and Armando Valdés-Velasquez. Adoption of sustainable low-impact fishing practices is not enough to secure sustainable livelihoods and social wellbeing in small-scale fishing communities. Marine Policy, 146(??):Article 105321, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003682.

Ding:2021:DFI

[DLX21] Lili Ding, Mengtong Lu, and Yuemei Xue. Driving factors on implementation of seasonal marine fishing moratorium system in China using evolutionary game. Marine Policy, 133(??):Article 104707, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003183.

Ding:2022:UCC

[DLYM22] Lili Ding, Mengxi Liu, Ying Yang, and Wen Ma. Understanding Chinese consumers’ purchase intention towards traceable seafood using an extended Theory of Planned Behavior model. Marine Policy, 137(??):Article 104973, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000203.

DaCostaOliveira:2020:TKF

[DMd^{+}20] Pablo Da Costa Oliveira, Ana Paula Madeira Di Beneditto, Valéria da Silva Quaresma, Alex Cardoso Bastos, and Camilah Antunes Zappes. Traditional knowledge of fishers versus an environmental disaster from mining waste in Central Brazil. Marine Policy, 120(??):Article 104129, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20301585.

Dutto:2023:AEE

[DME23] Davide Dutto, Krister Mars, and Håkan Eggert. Assessing the economic effects of the COVID-19 pandemic on Swedish shrimp fishers. Marine Policy, 150(??):Article 105496, April 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000234.

Drouineau:2023:FTF

[DMG^{+}23] Hilaire Drouineau, Fabien Moullec, Didier Gascuel, Francis Laloë, Sterenn Lucas, Nicolas Bez, Patrice Guillotreau, Jérôme Guitton, Pierre-Yves Hernvann, Martin Huret, Sigrid Lehuta, Marc Léopold, Stéphanie Mahévas, Marianne Robert, Mathieu Woillez, and Youen Vermard. Food for thought from French scientists for a revised EU Common Fisheries Policy to protect marine ecosystems and enhance fisheries performance. Marine Policy, 148(??):Article 105460, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22005073.

deMitcheson:2020:VVF

[dMLB^{+}20] Yvonne J. Sadovy de Mitcheson, Christi Linardich, João Pedro Barreiros, Gina M. Ralph, Alfonso Aguilar-Perera, Pedro Afonso, Brad E. Erisman, David A. Pollard, Sean T. Fennessy, Athila A. Bertoncini, Rekha J. Nair, Kevin L. Rhodes, Patrice Francour, Thierry Brulé, Melita A. Samoilys, Beatrice P. Ferreira, and Matthew T. Craig. Valuable but vulnerable: Over-fishing and under-management continue to threaten groupers so what now? Marine Policy, 116(??):Article 103909, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19306116.

DeSanto:2020:SMY

[DMNT20] Elizabeth M. De Santo, Elizabeth Mendenhall, Elizabeth Nyman, and Rachel Tiller. Stuck in the middle with you (and not much time left): the third intergovernmental conference on biodiversity beyond national jurisdiction. Marine Policy, 117(??):Article 103957, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308334.

Divu:2024:EPM

[DMS^{+}24] Damodaran Nair Divu, Suresh Kumar Mojjada, Pratheesh Omana Sudhakaran, Swathi Lekshmi Perumal Sundaram, Muktha Menon, Ramesh Kumar Mojjada, Mayur Shivdas Tade, Vinuja Syamala Vishwambharan, Jai Shree, Aarsha Subramanian, Boby Ignatius, and Achamveetil Gopalakrishnan. Economic performance and marine policy implications of mud spiny lobster mariculture in Tropical Sea Cages, North-Eastern Arabian Sea, India: an empirical study in marine economics. Marine Policy, 161(??):Article 106041, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24000393.

Devaux:2020:DES

[DN20] Caroline Devaux and Jean-Philippe Nicolaï. Designing an EU Ship Recycling Licence: a Roadmap. Marine Policy, 117(??):Article 103826, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301757.

Diekert:2021:CAC

[DNR21] Florian Diekert, Linda Nøstbakken, and Andries Richter. Control activities and compliance behavior — Survey evidence from Norway. Marine Policy, 125(??):Article 104381, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20310320.

deOliveira:2023:OJH

[dO23] César Soares de Oliveira. One jurisdiction away from a healthier ecosystem? The impacts of jurisdictional zones on the health of large marine ecosystems. Marine Policy, 155(??):Article 105698, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002257.

Doerr:2021:RSE

[Doe21] Angee N. Doerr. Regulating for social and environmental sustainability: a stakeholder perspective from the Bahamian spiny lobster fishery. Marine Policy, 124(??):Article 104366, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20310174.

Junior:2021:LSO

[dOJCSLdSB21] José Gilmar Cavalcante de Oliveira Júnior, João V. Campos-Silva, Richard J. Ladle, and Vandick da Silva Batista. Linking social organization, attitudes, and stakeholder empowerment in MPA governance. Marine Policy, 130(??):Article 104543, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001548.

Dominioni:2023:TET

[Dom23] Goran Dominioni. Towards an equitable transition in the decarbonization of international maritime transport: Exemptions or carbon revenues? Marine Policy, 154(??):Article 105669, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23001963.

deOliveira:2020:AEI

[dOPM^{+}20] Larissa Rosa de Oliveira, Ana Carolina Pont, Rodrigo Machado, Mônica Tais Engel, Paulo Henrique Ott, Enrique A. Crespo, and Silvio Marchini. Assessing the economic impact caused by South American sea lions based on onboard check versus fishermen’s perception: the two sides of the same coin. Marine Policy, 121(??):Article 104193, November 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19306517.

Soares:2020:OSS

[dOSTB^{+}20] Marcelo de Oliveira Soares, Carlos Eduardo Peres Teixeira, Luís Ernesto Arruda Bezerra, Sandra Vieira Paiva, Tallita Cruz Lopes Tavares, Tatiane Martins Garcia, Jorge Thé de Araújo, Carolina Coelho Campos, Sarah Maria Cavalcante Ferreira, Helena Matthews-Cascon, Alice Frota, Tarin Cristino Frota Mont’Alverne, Solange Teles Silva, Emanuelle Fontenele Rabelo, Cristiane Xerez Barroso, João Eduardo Pereira de Freitas, Mauro de Melo Júnior, Renata Polyana de Santana Campelo, Claudeilton Severino de Santana, Pedro Bastos de Macedo Carneiro, Antônio Jeovah Meirelles, Bráulio Almeida Santos, André Henrique Barbosa de Oliveira, Paulo Horta, and Rivelino Martins Cavalcante. Oil spill in South Atlantic (Brazil): Environmental and governmental disaster. Marine Policy, 115(??):Article 103879, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308346.

Drakopulos:2023:FCI

[DP23] Lauren Drakopulos and Melissa Poe. Facing change: Individual and institutional adaptation pathways in West Coast fishing communities. Marine Policy, 147(??):Article 105363, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004109.

DeDomenico:2023:RML

[DPB^{+}23] Francesca De Domenico, Cristina Pedà, Claudio Berti, Pierpaolo Consoli, Francesco Longo, Silvestro Greco, and Teresa Romeo. Reducing marine litter pollution from anchored fish aggregating devices: Sicilian good practices and the legislative paradox. Marine Policy, 154(??):Article 105695, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002221.

Delord:2022:FEM

[DPG^{+}22] Karine Delord, Timothée Poupart, Nicolas Gasco, Henri Weimerskirch, and Christophe Barbraud. First evidence of migration across the South Pacific in endangered Amsterdam albatross and conservation implications. Marine Policy, 136(??):Article 104921, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005327.

delPozo:2021:GAT

[dPJ21] Daniela Laínez del Pozo and Peter J. S. Jones. Governance analysis of two historical MPAs in northern Peru: Isla Lobos de Tierra and Isla Lobos de Afuera. Marine Policy, 127(??):Article 104096, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309492.

Daniell:2020:EEG

[DPP^{+}20] Katherine A. Daniell, Roel Plant, Victoria Pilbeam, Catherine Sabinot, Nicolas Paget, Karen Astles, Ruben Steffens, Olivier Barreteau, Severine Bouard, Peter Coad, Anna Gordon, Nils Ferrand, Pierre-Yves Le Meur, Caroline Lejars, Pierre Maurel, Ana Rubio, Jean-Emmanuel Rougier, and Ian White. Evolutions in estuary governance? Reflections and lessons from Australia, France and New Caledonia. Marine Policy, 112(??):Article 103704, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303336.

Davies:2023:NFE

[DQJ23] Tim K. Davies, Elise Quinn, and Ernesto Jardim. A novel framework to evaluate the accuracy of information used in MSC fisheries assessments: Development challenges and solutions. Marine Policy, 158(??):Article 105869, December 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004025.

Valck:2022:RUP

[DR22] Jeremy De Valck and John Rolfe. Reviewing the use of proxies to value coastal and marine biodiversity protection: the Great Barrier Reef in Australia. Marine Policy, 136(??):Article 104890, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005017.

Delgado:2021:AFE

[DRCdMCJ21] Erick Rodolfo Menéndez Delgado, Elvira Bernardita Rodríguez Ríos, Darío Bernal Casasola, and María del Mar Cerbán Jiménez. Artisanal fishery in Ecuador. A case study of Manta city and its economic policies to improve competitiveness of the sector. Marine Policy, 124(??):Article 104313, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030960X.

Deng:2022:IER

[DRY22] Yuncheng Deng, James Randall, and Fang Ye. Island ecological restoration and management practices based on nature: Conference report. Marine Policy, 143(??):Article 105188, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002354.

daSilva:2021:BAA

[dS21] Alexandre Pereira da Silva. Brazil advances over the Area: the inclusion of the Rio Grande Rise within the Brazilian outer continental shelf and its consequences for other states and for the common heritage of mankind. Marine Policy, 125(??):Article 104399, March 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000099.

daSilva:2022:PVT

[dS22a] Alexandre Pereira da Silva. Protecting the Vitoria– Trindade Chain: Legal clarifications and alternatives toward a marine protected area in the Brazilian outer continental shelf. Marine Policy, 136(??):Article 104926, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21005376.

Dunkley:2022:WFB

[DS22b] Frith Dunkley and Jean-Luc Solandt. Windfarms, fishing and benthic recovery: Overlaps, risks and opportunities. Marine Policy, 145(??):Article 105262, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003098.

DelBene:2021:PDG

[DSB21] James A. DelBene, Andrew M. Scheld, and Donna M. Bilkovic. Preferences for derelict gear mitigation strategies by commercial fishers. Marine Policy, 132(??):Article 104662, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X21002736.

Driedger:2023:GMP

[DSC^{+}23] Alex Driedger, Jennifer Sletten, Claire Colegrove, Timothé Vincent, Virgil Zetterlind, Joachim Claudet, and Barbara Horta e Costa. Guidance on marine protected area protection level assignments when faced with unknown regulatory information. Marine Policy, 148(??):Article 105441, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22004882.

deLima:2022:ECC

[dSdLSGB22] Andre de Souza de Lima, Marinez Eymael Garcia Scherer, Tiago Borges Ribeiro Gandra, and Jarbas Bonetti. Exploring the contribution of climate change policies to integrated coastal zone management in Brazil. Marine Policy, 143(??):Article 105180, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002275.

Ding:2023:ECT

[DSJ^{+}23] Qi Ding, Xiujuan Shan, Xianshi Jin, Harry Gorfine, Lisha Guan, and Tao Yang. Evolution of China’s Total Allowable Catch (TAC) system: Review and way forward. Marine Policy, 147(??):Article 105390, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004377.

Junior:2020:BEB

[dSJNS20] Oswaldo Gomes de Souza Junior, José Leocyvan Gomes Nunes, and Renato Azevedo Matias Silvano. Biology, ecology and behavior of the acoupa weakfish Cynoscion acoupa (Lacepède, 1801) according to the local knowledge of fishermen in the northern coast of Brazil. Marine Policy, 115(??):Article 103870, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302593.

Dinkel:2020:ISE

[DSL20] T. M. Dinkel and J. L. Sánchez-Lizaso. Involving stakeholders in the evaluation of management strategies for shortfin mako (Isurus oxyrinchus) and blue shark (Prionace glauca) in the Spanish longline fisheries operating in the Atlantic Ocean. Marine Policy, 120(??):Article 104124, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20301883.

Deith:2021:LLC

[DSL^{+}21] Mairin C. M. Deith, Daniel J. Skerritt, Roberto Licandeo, Daniel E. Duplisea, Caroline Senay, Divya A. Varkey, and Murdoch K. McAllister. Lessons learned for collaborative approaches to management when faced with diverse stakeholder groups in a rebuilding fishery. Marine Policy, 130(??):Article 104555, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001664.

daSilva:2021:PGI

[dSOW^{+}21] Carina Vieira da Silva, Margarida Ortigão, Tom Willaert, Renato Rosa, Luís Catela Nunes, and Maria A. Cunha e Sá. Participatory Geographic Information Systems (PGIS): Alternative approaches to identify potential conflicts and positional accuracy in marine and coastal ecosystem services. Marine Policy, 131(??):Article 104650, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100261X.

Delgado:2022:ITK

[DSQ^{+}22] Luisa E. Delgado, Camila Sandoval, Paola Quintanilla, David Quiñones-Guerrero, Ignacio A. Marín, and Víctor H. Marín. Including traditional knowledge in coastal policymaking: Yaldad Bay (Chiloé, southern Chile) as a case study. Marine Policy, 143(??):Article 105181, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002287.

Doddema:2020:RIT

[DSWB20] Mandy Doddema, Gert Spaargaren, Budy Wiryawan, and Simon R. Bush. Responses of Indonesian tuna processing companies to enhanced public and private traceability. Marine Policy, 119(??):Article 104100, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302207.

Daniel:2023:DFG

[DT23] Damaris Benny Daniel and Saly N. Thomas. Derelict fishing gear abundance, its causes and debris management practices — insights from the fishing sector of Kerala, India. Marine Policy, 148(??):Article 105429, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004766.

Domondon:2021:PEM

[DTBP21] Paolo Roberto Domondon, Raquel Sanchez Tirona, Steve Box, and Robert Pomeroy. Pathways to establishing managed access and networks of reserves. Marine Policy, 130(??):Article 104580, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001913.

Duan:2022:CPD

[Dua22] Wen Duan. China’s participation in the discussion on marine protected areas in the BBNJ negotiations and its implications. Marine Policy, 145(??):Article 105266, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2200313X.

Dunning:2021:AGP

[Dun21a] Kelly Heber Dunning. Adaptive Governance: the proposed port expansion in the Cayman Islands and its impacts to coral reefs. Marine Policy, 124(??):Article 104248, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20308940.

Dunning:2021:HMR

[Dun21b] Kelly Heber Dunning. How are managers responding to local and global ecological stressors? The case of Indonesian co-managed coral reefs in the Anthropocene. Marine Policy, 131(??):Article 104560, September 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001718.

Dunley:2023:UNV

[Dun23] Richard Dunley. Uncrewed naval vessels and the span of maritime tasks. Marine Policy, 149(??):Article 105481, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000088.

Dunley:2024:AMS

[Dun24] Richard Dunley. Australianised military off-the-shelf: Australia’s naval ship design ethos and the Anzac Class frigates. Marine Policy, 161(??):Article 106003, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000010.

Demirel:2021:MTA

[DUYEA21] Nazli Demirel, Aylin Ulman, Taner Yıldız, and Pinar Ertör-Akyazi. A moving target: Achieving good environmental status and social justice in the case of an alien species, Rapa whelk in the Black Sea. Marine Policy, 132(??):Article 104687, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002980.

deVrees:2021:AMS

[dV21] Leo de Vrees. Adaptive marine spatial planning in the Netherlands sector of the North Sea. Marine Policy, 132(??):Article 103418, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18304597.

Dalton:2020:RBP

[DWC^{+}20] Tracey Dalton, Michael Weir, Aislyne Calianos, Nelle D’Aversa, and Julia Livermore. Recreational boaters’ preferences for boating trips associated with offshore wind farms in US waters. Marine Policy, 122(??):Article 104216, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20308629.

Devi:2021:EPA

[DXS^{+}21] Manoharmayum Shaya Devi, K. A. Martin Xavier, Asem Sanjit Singh, Leela Edwin, Veerendra Veer Singh, and Latha Shenoy. Environmental pressure of active fishing method: a study on carbon emission by trawlers from north-west Indian coast. Marine Policy, 127(??):Article 104453, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000646.

Zoysa:2020:BWD

[dZ20] Rapti Siriwardane de Zoysa. Beyond the wall: Dyking as an object of everyday governance in the Bay of Manila, Philippines. Marine Policy, 112(??):Article 103661, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18303567.

Estes:2021:EML

[EAA^{+}21] Maurice Estes, Clarissa Anderson, Ward Appeltans, Nic Bax, Nina Bednaršek, Gabrielle Canonico, Samy Djavidnia, Elva Escobar, Peer Fietzek, Marilaure Gregoire, Elliott Hazen, Maria Kavanaugh, Franck Lejzerowicz, Fabien Lombard, Patricia Miloslavich, Klas O. Möller, Jacquomo Monk, Enrique Montes, Hassan Moustahfid, Monica M. C. Muelbert, Frank Muller-Karger, Lindsey E. Peavey Reeves, Erin V. Satterthwaite, Jörn O. Schmidt, Ana M. M. Sequeira, Woody Turner, and Lauren V. Weatherdon. Enhanced monitoring of life in the sea is a critical component of conservation management and sustainable economic growth. Marine Policy, 132(??):Article 104699, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003109.

Estradivari:2022:MCB

[EAA^{+}22] Estradivari, Muh. Firdaus Agung, Dedi Supriadi Adhuri, Sebastian C. A. Ferse, Ita Sualia, Dominic A. Andradi-Brown, Stuart J. Campbell, Mohamad Iqbal, Harry D. Jonas, Muhammad Erdi Lazuardi, Hellen Nanlohy, Fitryanti Pakiding, Ni Kadek Sri Pusparini, Hikmah C. Ramadhana, Toni Ruchimat, I Wayan Veda Santiadji, Natelda R. Timisela, Laura Veverka, and Gabby N. Ahmadia. Marine conservation beyond MPAs: Towards the recognition of other effective area-based conservation measures (OECMs) in Indonesia. Marine Policy, 137(??):Article 104939, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21005509.

Eggert:2021:AGF

[EAAG21] Håkan Eggert, Christopher M. Anderson, James L. Anderson, and Taryn M. Garlock. Assessing global fisheries using Fisheries Performance Indicators: Introduction to special section. Marine Policy, 125(??):Article 104253, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030899X.

Estradivari:2022:MCS

[EABA^{+}22] Estradivari, Dominic A. Andradi-Brown, Amkieltiela, Christian N. Handayani, Fikri F. Sjahruddin, Muh. Firdaus Agung, Stuart J. Campbell, Kelly Claborn, Matheus De Nardo, Helen E. Fox, Louise Glew, Amehr Hakim, Muhammad Erdi Lazuardi, Hellen Nanlohy, Wira Sanjaya, Edy Setyawan, Natelda R. Timisela, Laura Veverka, Nara Wisesa Wiwardhana, Marthen Welly, Imam Musthofa Zainudin, and Gabby N. Ahmadia. Marine conservation in the Sunda Banda Seascape, Indonesia. Marine Policy, 138(??):Article 104994, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000410.

Elvarsson:2020:UMC

[EAGV20] Bjarki Tór Elvarsson, Sveinn Agnarsson, Saga Gudmundsdóttir, and Jónas Vidarsson. Using multi-criteria analysis the assess impacts of change in ecosystem-based fisheries management: the case of the Icelandic cod. Marine Policy, 116(??):Article 103825, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18308303.

Elvestad:2023:EUT

[EB23] Christel Elvestad and Trond Bjørndal. The EU–UK Trade and Cooperation Agreement: Issue linkages and its implications for fisheries management and trade in seafood. Marine Policy, 148(??):Article 105380, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004274.

Ellefsen:2024:FIF

[EB24] Hans Ellefsen and Daniel W. Bromley. Faroe Islands fisheries in the Northeast Atlantic: Brexit, Russia, and economic incoherence. Marine Policy, 159(??):Article 105895, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004281.

Eger:2021:SRI

[EdLP^{+}21] S. L. Eger, R. C. de Loë, J. Pittman, G. Epstein, and S. C. Courtenay. A systematic review of integrated coastal and marine management progress reveals core governance characteristics for successful implementation. Marine Policy, 132(??):Article 104688, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002992.

Evans:2023:UTT

[EFFP23] Tegan Evans, Stephen Fletcher, Pierre Failler, and Jonathan Potts. Untangling theories of transformation: Reflections for ocean governance. Marine Policy, 155(??):Article 105710, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002439.

Egede:2023:UAC

[Ege23] Edwin Egede. UNCLOS 82: Africa’s contributions to the development of modern law of the sea 40 years later. Marine Policy, 148(??):Article 105463, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22005103.

Ehler:2021:ISI

[Ehl21a] Charles Ehler. Introduction to a special issue on marine/maritime spatial planning. Marine Policy, 132(??): Article 104117, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X20306424.

Ehler:2021:TDP

[Ehl21b] Charles N. Ehler. Two decades of progress in Marine Spatial Planning. Marine Policy, 132(??):Article 104134, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18303816.

Engehagen:2021:OHF

[EHLT21] Sondre Engehagen, Marius Hornslien, Maria Lavrutich, and Stian Tønnessen. Optimal harvesting of farmed salmon during harmful algal blooms. Marine Policy, 129(??):Article 104528, July 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001391.

Ewell:2020:ERF

[EHM^{+}20] Christopher Ewell, John Hocevar, Elizabeth Mitchell, Samantha Snowden, and Jennifer Jacquet. An evaluation of Regional Fisheries Management Organization atsea compliance monitoring and observer programs. Marine Policy, 115(??):Article 103842, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S0308597X19305913. See corrigendum [EHM^{+}22].

Ewell:2022:CER

[EHM^{+}22] Christopher Ewell, John Hocevar, Elizabeth Mitchell, Samantha Snowden, and Jennifer Jacquet. Corrigendum to “An evaluation of Regional Fisheries Management Organization at-sea compliance monitoring and observer programs” [Mar. Policy 115 (2020) 103842]. Marine Policy, 139(??):Article 105023, May 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000707. See [EHM^{+}20].

Escalle:2023:TNE

[EMH^{+}23] Lauriane Escalle, Jennyfer Mourot, Paul Hamer, Steven R. Hare, Naiten Bradley Phillip, and Graham M. Pilling. Towards non-entangling and biodegradable drifting fish aggregating devices — baselines and transition in the world’s largest tuna purse seine fishery. Marine Policy, 149(??):Article 105500, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000271.

Eladawy:2024:TCH

[EMN^{+}24] Ahmed Eladawy, Neil C. Mitchell, Takashi Nakamura, Momen El-Husseiny, Yuta A. Takagi, Nabil Elhady, Brook Muller, Sara Abdel-Hamid, Asmaa Mohammed, Kazuo Nadaoka, and J. P. Walsh. Towards a combined human– natural system approach in the Northern Red Sea Region: Ecological challenges, sustainable development, and community engagement. Marine Policy, 159(??):Article 105917, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004505.

Espinasse:2023:SPN

[EMS^{+}23] Marina Espinasse, Eirik Mikkelsen, Sigrunn Holbek Sørbye, Mette Skern-Mauritzen, Jannike Falk-Andersson, and Per Fauchald. Seafood production in Northern Norway: Analyzing variation and co-development in aquaculture and coastal fisheries. Marine Policy, 155(??):Article 105777, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2300310X.

Edwards:2020:POQ

[EP20] D. N. Edwards and E. Pinkerton. Priced out of ownership: Quota leasing impacts on the financial performance of owner-operators. Marine Policy, 111(??):Article 103718, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19304105.

EROL:2022:FEI

[Ero22] Sercan Erol. Financial and economic impacts of the COVID-19 pandemic on aquaculture in Türkiye and financial policy recommendations. Marine Policy, 146 (??):Article 105313, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003608.

Eryasar:2022:USM

[ES22] Ahmet Raif Eryaşar and İsmet Saygu. Using social media to identify recreational bluefish angling in the Mediterranean and Black Sea. Marine Policy, 135(??):Article 104834, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004450.

Engel:2022:BPA

[EV22] Monica T. Engel and Jerry J. Vaske. Balancing public acceptability and consensus regarding marine protected areas management using the Potential for Conflict Index_{2}. Marine Policy, 139(??):Article 105042, May 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000896.

Engel:2021:SHN

[EVB21] Monica T. Engel, Jerry J. Vaske, and Alistair J. Bath. Seal hunting in Newfoundland from the perspective of local people. Marine Policy, 128(??):Article 104491, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001020.

Erkkila-Valimaki:2022:CFS

[EVPMKP22] Anne Erkkilä-Välimäki, Mari Pohja-Mykrä, Jenny Katila, and Riitta Pöntynen. Coastal fishery stakeholders’ perceptions, motivation, and trust regarding maritime spatial planning and regional development: the case in the Bothnian Sea of the northern Baltic Sea. Marine Policy, 144(??):Article 105205, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002524.

Fujita:2023:SBC

[FAB^{+}23] Rod Fujita, Simona Augyte, Jennifer Bender, Poppy Brittingham, Alejandro H. Buschmann, Max Chalfin, Jamie Collins, Kristen A. Davis, John Barry Gallagher, Rebecca Gentry, Rebecca L. Gruby, Kristin Kleisner, Monica Moritsch, Nichole Price, Loretta Roberson, John Taylor, and Charles Yarish. Seaweed blue carbon: Ready? Or not? Marine Policy, 155(??):Article 105747, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002804.

Fisher:2023:ECC

[FAC^{+}23] Jeffrey Fisher, Dror Angel, Myriam Callier, Daniel Cheney, Ramon Filgueira, Bobbi Hudson, Christopher W. McKindsey, Lisa Milke, Heather Moore, Francis O’Beirn, Jack O’Carroll, Berit Rabe, Trevor Telfer, and Carrie J. Byron. Ecological carrying capacity in mariculture: Consideration and application in geographic strategies and policy. Marine Policy, 150(??):Article 105516, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300043X.

Finseraas:2024:GWW

[FAC^{+}24] Eirik Finserås, Ignacio Herrera Anchustegui, Etienne Cheynet, Cristian Guillermo Gebhardt, and Joachim Reuder. Gone with the wind? Wind farm-induced wakes and regulatory gaps. Marine Policy, 159(??):Article 105897, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2300430X.

Furman:2021:MMA

[FAGS21] Kelsi L. Furman, Payam Aminpour, Steven A. Gray, and Steven B. Scyphers. Mental models for assessing coastal social-ecological systems following disasters. Marine Policy, 125(??):Article 104334, March 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309830.

Fajardo:2022:CCI

[Faj22] Teresa Fajardo. To criminalise or not to criminalise IUU fishing: the EU’s choice. Marine Policy, 144(??): Article 105212, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X22002597.

Fernandez-Arcaya:2024:BFB

[FARBV^{+}24] Ulla Fernandez-Arcaya, Augusto Rodríguez-Basalo, Patricia Verísimo, José Rodriguez, Elvira Ceballos, Jose Manuel Gonzalez-Irusta, Ana García-Alegre, Maider Plaza-Morlote, Alberto Serrano, and Antonio Punzón. Bottom fishing beyond trawling. spatio-temporal trends of mobile and static bottom fisheries on benthic habitats. Marine Policy, 159(??):Article 105805, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300338X.

Fujita:2023:TER

[FBC^{+}23] Rod Fujita, Poppy Brittingham, Ling Cao, Halley Froehlich, Matt Thompson, and Taylor Voorhees. Toward an environmentally responsible offshore aquaculture industry in the United States: Ecological risks, remedies, and knowledge gaps. Marine Policy, 147(??):Article 105351, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003980.

Forse:2021:BBU

[FDF^{+}21] Andy Forse, Benjamin Drakeford, Pierre Failler, Jonathan Potts, and Negar Akbari. Beyond Brexit — Is the UK’s Fixed Quota Allocation (FQA) system in need of a fix? Marine Policy, 129(??):Article 104563, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001743.

Fabbri:2024:GIS

[FDGM24] Federico Fabbri, Jean-Noël Druon, Brendan J. Godley, and David March. Geospatial identification of stakeholders to support dynamic ocean management in transboundary areas. Marine Policy, 161(??):Article 106035, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000332.

Feichtner:2021:CLE

[Fei21] Isabel Feichtner. Contractor liability for environmental damage resulting from deep seabed mining activities in the area. Marine Policy, ??(??):Article 103502, ???? 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307115.

Feldman:2022:RAF

[Fel22] Nizan Feldman. Rallying around the “flag of convenience”: Merchant fleet flags and sanctions effectiveness. Marine Policy, 143(??):Article 105129, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001762.

Fennell:2023:CBC

[Fen23] Hannah L. Fennell. Changing behavior: Can intervention design from the public health sector help solve the problem of fishing gear conflict? Marine Policy, 151(??):Article 105527, May 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000544.

Frohlich:2023:NAA

[FFD^{+}23] Miguel Frohlich, Pedro Fidelman, Ian Dutton, Marcus Haward, Brian W. Head, Dianne Maynard, David Rissik, and Joanna Vince. A network approach to analyse Australia’s blue economy policy and legislative arrangements. Marine Policy, 151(??):Article 105588, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300115X.

Fedi:2020:MAM

[FFE20] Laurent Fedi, Olivier Faury, and Laurent Etienne. Mapping and analysis of maritime accidents in the Russian Arctic through the lens of the Polar Code and POLARIS system. Marine Policy, 118(??):Article 103984, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030230X.

Fedi:2024:AIT

[FFE^{+}24] Laurent Fedi, Olivier Faury, Laurent Etienne, Ali Cheaitou, and Patrick Rigot-Muller. Application of the IMO taxonomy on casualty investigation: Analysis of 20 years of marine accidents along the North-East Passage. Marine Policy, 162(??):Article 106061, April 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000599.

Ferreira:2022:RUR

[FFGJ22] Maria Adelaide Ferreira, Christopher Barrio Froján, Vikki Gunn, and David E. Johnson. A role for UNEP’s Regional Seas Programme under the post-2020 global biodiversity framework. Marine Policy, 136(??):Article 104930, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21005418.

Finkbeiner:2021:CPW

[FFYE21] Elena M. Finkbeiner, Juno Fitzpatrick, and Whitney Yadao-Evans. A call for protection of women’s rights and economic, social, cultural (ESC) rights in seafood value chains. Marine Policy, 128(??):Article 104482, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000932.

Friess:2021:POR

[FGC21] Bernhard Friess and Marie Grémaud-Colombier. Policy outlook: Recent evolutions of maritime spatial planning in the European Union. Marine Policy, 132(??):Article 103428, October 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X18304652.

Finke:2020:NWM

[FGK^{+}20] Gunnar Finke, Kira Gee, Anja Kreiner, Maria Amunyela, and Rodney Braby. Namibia’s way to Marine Spatial Planning — Using existing practices or instigating its own approach? Marine Policy, 121(??):Article 104107, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2030333X.

Froehlich:2021:SSF

[FGL^{+}21] Halley E. Froehlich, Rebecca R. Gentry, Sarah E. Lester, Richard S. Cottrell, Gavin Fay, Trevor A. Branch, Jessica A. Gephart, Easton R. White, and Julia K. Baum. Securing a sustainable future for US seafood in the wake of a global crisis. Marine Policy, 124(??):Article 104328, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309751.

Fernandez-Gonzalez:2022:CSC

[FGPPGG22] Raquel Fernández-González, Marcos I. Pérez-Pérez, and María Dolores Garza-Gil. COVID-19 and the Spanish Celtic Sea fishery: an economic analysis. Marine Policy, 143(??):Article 105204, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002512.

Fernandez-Gonzalez:2023:AEE

[FGPPGG23] Raquel Fernández-González, Marcos I. Pérez-Pérez, and María Dolores Garza-Gil. Assessing the economic effects of the European landing obligation on the Spanish fishery in the Celtic Sea. Marine Policy, 148(??):Article 105453, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22005000.

Fernandez-Gonzalez:2021:ROS

[FGPPPV21] Raquel Fernández-González, Marcos I. Pérez-Pérez, and Raisa Pérez-Vas. Real options for a small company in a context of market concentration: a case study of investment in a turbot farming plant in Spain. Marine Policy, 134(??):Article 104828, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004395.

Furman:2023:SES

[FHBS23] Kelsi L. Furman, Sharon L. Harlan, Luiz Barbieri, and Steven B. Scyphers. Social equity in shore-based fisheries: Identifying and understanding barriers to access. Marine Policy, 148(??):Article 105355, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200402X.

Fischer:2022:HTR

[Fis22] Johanne Fischer. How transparent are RFMOs? Achievements and challenges. Marine Policy, 136(??):Article 104106, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19307158.

Flaaten:2021:RDF

[Fla21] Ola Flaaten. The rise and decline of fishing industry support — with a translation from Norwegian of Bjørn S. Brochmann’s 1981 article ‘Long-term effects of government support to the fisheries’. Marine Policy, 126(??):Article 104112, April 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20304541.

Filous:2021:FSM

[FLB^{+}21] Alexander Filous, Robert J. Lennox, Jean Pierre Beaury, Hinano Bagnis, Mathew Mchugh, Alan M. Friedlander, Eric E. G. Clua, Steven J. Cooke, Todd K. Fuller, and Andy J. Danylchuk. Fisheries science and marine education catalyze the renaissance of traditional management (rahui) to improve an artisanal fishery in French Polynesia. Marine Policy, 123(??):Article 104291, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309374.

Fan:2023:SPC

[FLC^{+}23] Hanwen Fan, Jing Lyu, Zheng Chang, Xuzhuo He, and Shu Guo. Spatial patterns and characteristics of global piracy analyzed using a geographic information system. Marine Policy, 157(??):Article 105816, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003494.

Fraser:2020:CSV

[FMCD20] Molly D. Fraser, Lauren H. McWhinnie, Rosaline R. Canessa, and Chris T. Darimont. Compliance of small vessels to minimum distance regulations for humpback and killer whales in the Salish Sea. Marine Policy, 121(??):Article 104171, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20306709.

Franco-Melendez:2021:TUR

[FMCT^{+}21] Milagros Franco-Meléndez, Luis A. Cubillos, Jorge Tam, Simón Hernández Aguado, Renato A. Quiñones, and Aldo Hernández. Territorial Use Rights for Fisheries (TURF) in central-southern Chile: Their sustainability status from a transdisciplinary holistic approach. Marine Policy, 132(??):Article 104644, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002554.

Fernandez-Macho:2020:AAV

[FMGV20] Javier Fernández-Macho, Pilar González, and Jorge Virto. Assessing anthropogenic vulnerability of coastal regions: DEA-based index and rankings for the European Atlantic Area. Marine Policy, 119(??):Article 104030, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20302694.

Francolini:2023:SPS

[FMLMM23] E. M. Francolini, J. B. Mann-Lang, E. McKinley, and B. Q. Mann. Stakeholder perspectives on socio-economic challenges and recommendations for better management of the Aliwal Shoal Marine Protected Area in South Africa. Marine Policy, 148(??):Article 105470, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22005218.

Fujii:2022:IEC

[FMS^{+}22] Iwao Fujii, Miko Maekawa, Nozomi Shimizu, Naohisa Kanda, Nariaki Mikuni, Kazunobu Suzuki, Izumi Tsurita, Miriam C. Balgos, and Marjo K. Vierros. Implications of existing capacity building efforts for the conservation and sustainable use of marine biological diversity of areas beyond national jurisdiction: a case study of Japan. Marine Policy, 138(??):Article 105004, April 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000513.

Fuentes:2021:CIR

[FMWW21] Mariana M. P. B. Fuentes, Zoë A. Meletis, Natalie E. Wildermann, and Matthew Ware. Conservation interventions to reduce vessel strikes on sea turtles: a case study in Florida. Marine Policy, 128(??):Article 104471, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000828.

Fleming:2020:STH

[FOH^{+}20] Aysha Fleming, Emily Ogier, Alistair J. Hobday, Linda Thomas, Jason R. Hartog, and Bianca Haas. Stakeholder trust and holistic fishery sustainability assessments. Marine Policy, 111(??):Article 103719, January 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306335.

Fujii:2023:PCR

[FOKM23] Iwao Fujii, Yumi Okochi, Hajime Kawamura, and Mitsutaku Makino. Potential cooperation of RFMOs for the integrity of MCS: Lessons from the three RFMOs in the Asia– Pacific. Marine Policy, 155(??):Article 105748, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002816.

Formenti:2022:ATF

[For22] Lorenzo Formenti. Assessing transparency in fisheries subsidies: a notification-driven analysis. Marine Policy, 136(??):Article 104152, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19307766.

Fox:2022:SSU

[Fox22] Clive J. Fox. Scoring the status of UK shared and national fish stocks around the time of Brexit. Marine Policy, 135(??):Article 104851, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004620.

Florido:2023:ECP

[FR23] Alberto Roca Florido and Emilio Padilla Rosa. Everything comes at a price: the impact of eliminating harmful subsidies in the Spanish marine resources industry. Marine Policy, 153(??):Article 105661, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001884.

Franckx:2023:SMB

[Fra23] Erik Franckx. A single maritime boundary: From UN-CLOS III to present-day developments. Marine Policy, 148(??):Article 105425, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004729.

Finnis:2022:MWF

[FRM22] Joel Finnis and Emily Reid-Musson. Managing weather & fishing safety: Marine meteorology and fishing decisionmaking from a governance and safety perspective. Marine Policy, 142(??):Article 105120, August 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001671.

Frost:2021:NBE

[Fro21] Rebecca Frost. Normative beliefs and economic life: a case study of the fishing industry in two communities in rural Newfoundland. Marine Policy, 127(??):Article 104423, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21000336.

Filho:2022:ESV

[FRVB22] Luiz Magalhães Filho, Peter Roebeling, Sebastián Villasante, and Maria Isabel Bastos. Ecosystem services values and changes across the Atlantic coastal zone: Considerations and implications. Marine Policy, 145 (??):Article 105265, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003128.

Ford:2021:SBT

[FS21] Eleanor Ford and Bryce D. Stewart. Searching for a bridge over troubled waters: an exploratory analysis of trust in United Kingdom fisheries management. Marine Policy, 132(??):Article 104686, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002979.

Freeman:2022:WES

[FS22] Richard Freeman and Kristina Svels. Women’s empowerment in small-scale fisheries: the impact of Fisheries Local Action Groups. Marine Policy, 136(??):Article 104907, February 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21005182.

Frohlich:2021:LBA

[FSF^{+}21] Miguel F. Frohlich, Timothy F. Smith, Pedro Fidelman, Claudia Baldwin, Chris Jacobson, and R. W. (Bill) Carter. Legal barriers to adaptive coastal management at a coastal erosion hotspot in Florianópolis, Brazil. Marine Policy, 127(??):Article 104436, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000464.

Faunce:2023:FOE

[FSKJ23] Craig H. Faunce, Jaclyn Smith, Andy Kingham, and Dennis Jaszka. Fisheries observers as enforcement assets: 21 years of lessons from the North Pacific. Marine Policy, 158(??):Article 105868, December 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004013.

Fjeld:2023:MNG

[FTG^{+}23] Kristian Fjeld, Rachel Tiller, Eduardo Grimaldo, Leif Grimsmo, and Inger-Beate Standal. Mesopelagics — new gold rush or castle in the sky? Marine Policy, 147(??):Article 105359, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004067.

Failler:2020:PTR

[FTGD^{+}20] Pierre Failler, Grégoire Touron-Gardic, Benjamin Drakeford, Oumar Sadio, and Marie-Suzanne Traoré. Perception of threats and related management measures: the case of 32 marine protected areas in West Africa. Marine Policy, 117(??):Article 103936, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304452.

Ferguson:2022:LPP

[FTM^{+}22] Caroline E. Ferguson, Teri Tuxson, Sangeeta Mangubhai, Stacy Jupiter, Hugh Govan, Victor Bonito, Semese Alefaio, Maxine Anjiga, Jonathan Booth, Tracey Boslogo, Delvene Boso, Ambroise Brenier, Akanisi Caginitoba, Ana Ciriyawa, Joeli Bili Fahai’ono, Margaret Fox, Andy George, Hampus Eriksson, Alec Hughes, Eugene Joseph, Sean Kadannged, Eferemo Kubunavanua, Sesimani Loni, Semisi Meo, Fiorenza Micheli, Elizah Nagombi, Rebecca Omaro, Anouk Ride, Annisah Sapul, Ann Singeo, Karen Stone, Margaret Tabunakawai-Vakalalabure, Marama Tuivuna, Caroline Vieux, Vutaieli B. Vitukawalu, and McKenzie Waide. Local practices and production confer resilience to rural Pacific food systems during the COVID-19 pandemic. Marine Policy, 137(??):Article 104954, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200001X.

Froese:2021:PTE

[FTSG21] Rainer Froese, Athanassios C. Tsikliras, Giuseppe Scarcella, and Didier Gascuel. Progress towards ending overfishing in the Northeast Atlantic. Marine Policy, 125(??):Article 104282, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309283.

Flynn:2023:MSD

[FTW^{+}23] Sarah Flynn, Elizabeth Tray, Tom Woolley, Adam Leadbetter, Kellie Heney, Deirdre O’Driscoll, Caitriona Nic Aonghusa, and Andrew Conway. Management of spatial data integrity including stakeholder feedback in Maritime Spatial Planning. Marine Policy, 156(??):Article 105799, October 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003329.

Fujita:2021:AMD

[Fuj21] Rod Fujita. The assessment and management of data limited fisheries: Future directions. Marine Policy, 133(??):Article 104730, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003419.

Fermepin:2024:GMC

[FWGM24] Solange Fermepin, James E. M. Watson, Hedley S. Grantham, and Martin Mendez. Global marine conservation priorities for sustaining marine productivity, preserving biodiversity and addressing climate change. Marine Policy, 161(??):Article 106016, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000149.

Fu:2022:EIA

[FWL^{+}22] Xiu-Mei Fu, Li-Xia Wang, Chun-Yu Lin, Wan-Yu Wu, Hong-Li Ku, Shan-Shan Jiang, and Ying Liu. Evaluation of the innovation ability of China’s marine fisheries from the perspective of static and dynamic. Marine Policy, 139(??):Article 105032, May 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000793.

Fan:2022:CFT

[FXC^{+}22] Guohao Fan, Xiaoke Xie, Jihong Chen, Zheng Wan, Mingzhu Yu, and Jia Shi. Has China’s Free Trade Zone policy expedited port production and development? Marine Policy, 137(??):Article 104951, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005625.

Faure:2024:LST

[FY24] Michael G. Faure and Haiyang Yu. The liability of salvors towards third parties in case of environmental salvage: Economic analysis. Marine Policy, 160(??):Article 105914, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004475.

Feng:2023:CMM

[FZYF23] Fan Feng, Chenqi Zhai, Zhenguo Yang, and Kuo Feng. China’s marine manufacturing export growth: Policy, status, and driving components. Marine Policy, 149(??):Article 105495, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000222.

Gairin:2020:RHD

[GA20] Emma Gairin and Serge Andréfouët. Role of habitat definition on Aichi Target 11: Examples from New Caledonian coral reefs. Marine Policy, 116(??):Article 103951, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19306505.

Gyan:2020:APF

[GAA^{+}20] Watson Ray Gyan, Elliot Haruna Alhassan, Amos Asase, Daniel Nsoh Akongyuure, and Yang Qi-Hui. Assessment of postharvest fish losses: the case study of Albert Bosomtwi-Sam fishing harbour, Western Region, Ghana. Marine Policy, 120(??):Article 104120, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301536.

Griffiths:2022:LHF

[GABA^{+}22] Laura L. Griffiths, Dominic A. Andradi-Brown, Gabby N. Ahmadia, Purwanto, Awaludinnoer, Dale Bryan-Brown, and Christopher J. Brown. Linking historical fishing pressure to biodiversity outcomes to predict spatial variation in Marine Protected Area performance. Marine Policy, 139(??):Article 105024, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000719.

Grimm:2023:FFJ

[GABA^{+}23] Kerry E. Grimm, Jessica L. Archibald, Sara E. Bonilla-Anariba, Nadia Bood, and Steven W. J. Canty. Framework for fostering just and equitable seagrass policy, management, and social-ecological outcomes: Lessons learned from Belizean marine resource managers. Marine Policy, 152(??):Article 105606, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001331.

Goti-Aralucea:2021:EVR

[GABSD21] Leyre Goti-Aralucea, Jörg Berkenhagen, Erik Sulanke, and Ralf Döring. Efficiency vs resilience: the rise and fall of the German brown shrimp fishery in times of COVID 19. Marine Policy, 133(??):Article 104675, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002864.

Giovos:2020:AMS

[GAD^{+}20] Ioannis Giovos, Marco Arculeo, Nikolaos Doumpas, Dimitra Katsada, Mary Maximiadi, Eleni Mitsou, Vangelis Paravas, Roxani Naasan Aga-Spyridopoulou, Vasilis-Orestis Stoilas, Francesco Tiralongo, Ioannis E. Tsamadias, Luca Vecchioni, and Dimitrios K. Moutopoulos. Assessing multiple sources of data to detect illegal fishing, trade and mislabelling of elasmobranchs in Greek markets. Marine Policy, 112(??):Article 103730, February 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302027.

Gilman:2023:IMP

[GAD^{+}23] Eric Gilman, Kyle Antonelis, Joan Drinkwin, Saeid Gorgin, Petri Suuronen, Saly N. Thomas, and Jono Wilson. Introduction to the Marine Policy special issue on abandoned, lost and discarded fishing gear: Causes, magnitude, impacts, mitigation methods and priorities for monitoring and evidence-informed management. Marine Policy, 155(??):Article 105738, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002713.

Gee:2024:NMS

[GAD^{+}24] Kira Gee, Philipp Arndt, Marie Dahmen, Nele-Kristin Geitel, Anna Hunke, Bettina Käppeler, Dominic Plug, Ulrich Scheffler, and Kai Trümpler. A new maritime spatial plan for the German EEZ: Navigating multiple aspirations and spatial developments. Marine Policy, 161(??):Article 106006, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000046.

Galligan:2021:FER

[Gal21] Bryan P. Galligan. Fisheries extractivism and the right to subsistence: Conflicting governance models and the legal structures that enact them. Marine Policy, 133 (??):Article 104729, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003407.

Gales:2022:YOI

[Gal22] Nick Gales. At 75 years old, the IWC has never been more globally relevant. Marine Policy, 141(??):Article 105089, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001361.

Garcia-Alonso:2020:EWC

[GAMR20] Lorena Garcia-Alonso, Ticiana Grecco Zanon Moura, and David Roibas. The effect of weather conditions on port technical efficiency. Marine Policy, 113(??):Article 103816, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19305895.

Gamboa-Alvarez:2023:SCC

[GÁPSAP^{+}23] Miguel Á. Gamboa-Álvarez, Alicia V. Poot-Salazar, Alfonso Aguilar-Perera, R. Isaac Rojas-González, and Cuauhtémoc Ruiz-Pineda. Sea cucumber catch reconstruction and species composition as determined by seizures analyses from the IUU fishing in the Campeche Bank, Mexico. Marine Policy, 147(??):Article 105291, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003384.

Gardiner:2020:MPA

[Gar20] Natasha B. Gardiner. Marine protected areas in the Southern Ocean: Is the Antarctic Treaty System ready to co-exist with a new United Nations instrument for areas beyond national jurisdiction? Marine Policy, 122(??):Article 104212, December 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20308587.

Garcia:2022:SPC

[Gar22] Isabel Artime García. Spain: a pioneering country in the fight against the infringement of the international legal regime for fisheries. Marine Policy, 144(??):Article 105230, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002779.

Garcia:2024:FOO

[Gar24] Eleanor L. Garcia. Fisheries observers: an overlooked vulnerability for crime and corruption within the global fishing industry. Marine Policy, 161(??):Article 106029, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24000277.

Gonzalez-Andres:2020:IFI

[GASLCP20] Cristina González-Andrés, José Luis Sánchez-Lizaso, Jorge Cortés, and Maria Grazia Pennino. Illegal fishing in Isla del Coco National Park: Spatial–temporal distribution and the economic trade-offs. Marine Policy, 119(??):Article 104023, September 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19302192.

Goldsworthy:2021:CCS

[GB21] Lynda Goldsworthy and Eaven Brennan. Climate change in the Southern Ocean: Is the Commission for the Convention for the Conservation of Antarctic Marine Living Resources doing enough? Marine Policy, 130(??):Article 104549, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001603.

Giovos:2021:UPA

[GBB^{+}21] Ioannis Giovos, Adi Barash, Monica Barone, Claudio Barría, Diego Borme, Cecile Brigaudeau, Anastasia Charitou, Cristina Brito, Jock Currie, Mareike Dornhege, Lara Endrizzi, Kerstin Forsberg, Armelle Jung, Periklis Kleitou, Alison MacDiarmid, Dimitrios K. Moutopoulos, Shotaro Nakagun, João Neves, Flávia L. D. Nunes, Dorien Schröder, Ruth H. Thurstan, Malcolm Tull, Sezginer Tuncer, and Carlotta Mazzoldi. Understanding the public attitude towards sharks for improving their conservation. Marine Policy, 134(??):Article 104811, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100422X.

Galdies:2020:EPL

[GBC^{+}20] Charles Galdies, Richard Bellerby, Donata Canu, Wenting Chen, Enrique Garcia-Luque, Blaženka Gašparović, Jelena Godrijan, Paul J. Lawlor, Frank Maes, Alenka Malej, Dionisios Panagiotaras, Beatriz Martinez Romera, Claire E. Reymond, Julien Rochette, Cosimo Solidoro, Robert Stojanov, Rachel Tiller, Isabel Torres de Noronha, Grzegorz Uścinowicz, Nataşa Vaidianu, Cormac Walsh, and Roberta Guerra. European policies and legislation targeting ocean acidification in European waters — Current state. Marine Policy, 118(??):Article 103947, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309054.

Gonzalez-Bernat:2021:GAG

[GBC21a] Mariá José González-Bernat and Julian Clifton. A governance analysis of Guatemala’s first recognized marine protected area: the Wildlife Refuge of Punta de Manabique (RVSPM). Marine Policy, 127(??):Article 103626, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19304646.

Gonzalez-Bernat:2021:GAT

[GBC21b] María José González-Bernat and Julian Clifton. A governance analysis of two marine protected areas in the Pacific Region of Guatemala: the Multiple Use Area of Monterrico and the Private Reserve La Chorrera–Manchón Guamuchal. Marine Policy, 127(??):Article 103625, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304610.

Gomez-Ballesteros:2021:TCM

[GBCNCL^{+}21] M. Gómez-Ballesteros, C. Cervera-Núñez, M. Campillos-Llanos, A. Quintela, L. Sousa, M. Marques, F. L. Alves, C. Murciano, N. Alloncle, P. Sala, A. Lloret, A. P. Simão, A. C. Costa, D. Carval, D. Bailly, C. Nys, H. Sybill, and J. Dilasser. Transboundary cooperation and mechanisms for Maritime Spatial Planning implementation. SIMNORAT project. Marine Policy, 127(??):Article 104434, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000440.

Gomez:2023:MNC

[GBFMT23] Aina G. Gómez, Pau Balaguer, Àngels Fernández-Mora, and Joaquín Tintoré. Mapping the nautical carrying capacity of anchoring areas of the Balearic Islands’ coast. Marine Policy, 155(??):Article 105775, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003081.

Gjertsen:2020:AAA

[GBLBV20] Arild Gjertsen, Ingrid Bay-Larsen, Hilde Bjørkhaug, and Bjørn Vidar Vangelsten. Access to areas for algae cultivation in Norway. Marine Policy, 115(??):Article 103853, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19306050.

Gordon:2022:EFI

[GBSC22] Jesse Y. Gordon, Anne H. Beaudreau, Emma M. Saas, and Courtney Carothers. Engaging formal and informal institutions for stewardship of rockfish fisheries in the Gulf of Alaska. Marine Policy, 143(??):Article 105170, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002172.

Guilfoyle:2022:LWC

[GC22] Douglas Guilfoyle and Edward Sing Yue Chan. Lawships or warships? Coast guards as agents of (in)stability in the Pacific and South and East China Sea. Marine Policy, 140(??):Article 105048, June 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000951.

Gilman:2023:BNF

[GCB^{+}23] Eric Gilman, Milani Chaloupka, Hollie Booth, Martin Hall, Hilario Murua, and Jono Wilson. Bycatch-neutral fisheries through a sequential mitigation hierarchy. Marine Policy, 150(??):Article 105522, April 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000490.

Gilbey:2021:LDS

[GCC^{+}21a] John Gilbey, Gary Carvalho, Rita Castilho, Ilaria Coscia, Mark W. Coulson, Geir Dahle, Sofie Derycke, Sara M. Francisco, Sarah J. Helyar, Torild Johansen, Claudia Junge, Kara K. S. Layton, Jann Martinsohn, Iveta Matejusova, Joana I. Robalo, Naiara Rodríguez-Ezpeleta, Gonçalo Silva, Ilona Strammer, Anti Vasemägi, and Filip A. M. Volckaert. Life in a drop: Sampling environmental DNA for marine fishery management and ecosystem monitoring. Marine Policy, 124(??):Article 104331, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309787.

Guerreiro:2021:GPM

[GCC^{+}21b] José Guerreiro, Ana Carvalho, Daniela Casimiro, Marie Bonnin, Helena Calado, Hilde Toonen, Philippe Fotso, Ibrahima Ly, Osvaldina Silva, and Solange Teles da Silva. Governance prospects for maritime spatial planning in the tropical Atlantic compared to EU case studies. Marine Policy, 123(??):Article 104294, January 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309404.

Gamarra:2023:CFH

[GCF^{+}23] N. C. Gamarra, A. C. L. Costa, M. A. C. Ferreira, L. M. Diele-Viegas, A. P. O. Santos, R. J. Ladle, A. C. Malhado, and J. V. Campos-Silva. The contribution of fishing to human well-being in Brazilian coastal communities. Marine Policy, 150(??):Article 105521, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000489.

Gollner:2021:ASC

[GCG^{+}21] S. Gollner, A. Colaço, A. Gebruk, P. N. Halpin, N. Higgs, E. Menini, N. C. Mestre, P.-Y. Qian, J. Sarrazin, K. Szafranski, and C. L. Van Dover. Application of scientific criteria for identifying hydrothermal ecosystems in need of protection. Marine Policy, 132(??):Article 104641, October 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002529.

Gomez:2021:CHE

[GCL21] Sílvia Gómez, Arnau Carreño, and Josep Lloret. Cultural heritage and environmental ethical values in governance models: Conflicts between recreational fisheries and other maritime activities in Mediterranean marine protected areas. Marine Policy, 129(??):Article 104529, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001408.

Griffin:2023:AGP

[GCM^{+}23] Lucas P. Griffin, Grace A. Casselberry, Ezra M. Markowitz, Jacob W. Brownscombe, Aaron J. Adams, Bill Horn, Steven J. Cooke, and Andy J. Danylchuk. Angler and guide perceptions provide insights into the status and threats of the Atlantic tarpon (Megalops atlanticus) fishery. Marine Policy, 151(??):Article 105569, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000969.

Giraldi-Costa:2020:SZM

[GCMT20] Ana Clara Giraldi-Costa, Rodrigo Pereira Medeiros, and Liliani Marília Tiepolo. Step zero of marine protected areas of Brazil. Marine Policy, 120(??):Article 104119, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19306529.

Giebels:2020:TKM

[GCP^{+}20] Diana Giebels, Jana Carus, Maike Paul, Michael Kleyer, Bernd Siebenhüner, Arne Arns, Alexander Bartholomä, Vanessa Carlow, Jürgen Jensen, Britta Tietjen, Achim Wehrmann, and Boris Schröder. Transdisciplinary knowledge management: a key but underdeveloped skill in EBM decision-making. Marine Policy, 119(??):Article 104020, September 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X1930644X.

Gilman:2021:ERD

[GCS21] Eric Gilman, Milani Chaloupka, and Chrissie Sieben. Ecological risks of a data-limited fishery using an ensemble of approaches. Marine Policy, 133(??):Article 104752, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003638.

Guzman:2020:HWM

[GCV^{+}20] Héctor M. Guzman, Juan J. Capella, Carlos Valladares, Jorge Gibbons, and Richard Condit. Humpback whale movements in a narrow and heavily-used shipping passage, Chile. Marine Policy, 118(??):Article 103990, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302049.

Giacomarra:2021:AFE

[GCV^{+}21] Marcella Giacomarra, Maria Crescimanno, Demetris Vrontis, Lluís Miret Pastor, and Antonino Galati. The ability of fish ecolabels to promote a change in the sustainability awareness. Marine Policy, 123(??):Article 104292, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309386.

Gerhardinger:2020:CST

[GdACT20] Leopoldo Cavaleri Gerhardinger, Mariana Martins de Andrade, Marina Ribeiro Corrêa, and Alexander Turra. Crafting a sustainability transition experiment for the Brazilian blue economy. Marine Policy, 120(??):Article 104157, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20305509.

Gilman:2020:CPF

[GDLC20] Eric Gilman, Vicente De Ramón Castejón, Eparama Loganimoce, and Milani Chaloupka. Capability of a pilot fisheries electronic monitoring system to meet scientific and compliance monitoring objectives. Marine Policy, 113(??):Article 103792, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19306608.

Golo:2023:TCC

[GE23] Harrison K. Golo and Bolanle Erinosho. Tackling the challenges confronting women in the Elmina fishing community of Ghana: a human rights framework. Marine Policy, 147(??):Article 105349, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003967.

Gruby:2021:OBB

[GEB^{+}21] Rebecca L. Gruby, Ashley Enrici, Michele Betsill, Elodie Le Cornu, and Xavier Basurto. Opening the black box of conservation philanthropy: a co-produced research agenda on private foundations in marine conservation. Marine Policy, 132(??):Article 104645, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002566.

Gehrke:2023:WPB

[Geh23] Charlotte Gehrke. Wild polar bear conservation: a case of successful Arctic science diplomacy? Marine Policy, 155(??):Article 105783, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003160.

Gabela-Flores:2021:ELD

[GFD21] María Virginia Gabela-Flores and Amy Diedrich. Exploring leadership dynamics in community-based marine protected areas. Marine Policy, 129(??):Article 104569, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001809.

Gustavsson:2021:GBJ

[GFL^{+}21] Madeleine Gustavsson, Katia Frangoudes, Lars Lindström, María Catalina Ávarez, and Maricela de la Torre Castro. Gender and Blue Justice in small-scale fisheries governance. Marine Policy, 133(??):Article 104743, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003547.

Gerber:2021:COG

[GG21] Leonardus J. Gerber and Renée L. Grogan. Challenges of operationalising good industry practice and best environmental practice in deep seabed mining regulation. Marine Policy, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304639.

Guo:2022:PSE

[GGL22] Jianke Guo, Shu Guo, and Jing Lv. Potential spatial effects of opening Arctic shipping routes on the shipping network of ports between China and Europe. Marine Policy, 136(??):Article 104885, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004966.

Gacutan:2022:MSP

[GGP^{+}22] Jordan Gacutan, Ibon Galparsoro, Kemal Pınarbaşı, Arantza Murillas, Ibukun J. Adewumi, Teerapong Praphotjanaporn, Emma L. Johnston, Ken P. Findlay, and Ben M. Milligan. Marine spatial planning and ocean accounting: Synergistic tools enhancing integration in ocean governance. Marine Policy, 136(??):Article 104936, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005479.

Guzman:2020:SCT

[GHK20] Hector M. Guzman, Natasha Hinojosa, and Stefanie Kaiser. Ship’s compliance with a traffic separation scheme and speed limit in the Gulf of Panama and implications for the risk to humpback whales. Marine Policy, 120(??):Article 104113, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030021X.

Gilman:2022:MFS

[GHW^{+}22] Eric Gilman, Jenn Humberstone, Jono R. Wilson, Emmanuel Chassot, Alexis Jackson, and Petri Suuronen. Matching fishery-specific drivers of abandoned, lost and discarded fishing gear to relevant interventions. Marine Policy, 141(??):Article 105097, July 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001440.

Girard:2022:CGI

[Gir22]  Sophie Girard.  Can Geographical Indications pro-

mote sustainable shellfish farming? The example of Bay of Mont-Saint-Michel mussels. Marine Policy, 135(??):Article 104845, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004565.

Guggisberg:2022:TFG

[GJS22] Solène Guggisberg, Aline Jaeckel, and Tim Stephens. Transparency in fisheries governance: Achievements to date and challenges ahead. Marine Policy, 136 (??):Article 104639, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002505.

Gong:2023:MPR

[GJY23] Xiaoxing Gong, Huilin Jiang, and Dong Yang. Maritime piracy risk assessment and policy implications: a two-step approach. Marine Policy, 150(??):Article 105547, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2300074X.

Gorjanc:2022:EPA

[GKP^{+}22] Sašo Gorjanc, Katja Klančnik, Nadia K. Papadopoulou, Arantza Murillas-Maza, Klara Jarni, Theodora Paramana, Mišo Pavičić, Francesca Ronchi, Maria C. Uyarra, Špela Koren, Manos Dassenakis, Olja Vidjak, Christopher J. Smith, and Sanda Skejić. Evaluating the progress in achieving Good Environmental Status in the Mediterranean: a methodology to assess the effectiveness of Marine Strategy Framework Directive’s Programmes of Measures. Marine Policy, 136(??):Article 104889, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005005.

Garcia-Lorenzo:2023:PSS

[GLADGGVL23] Iria García-Lorenzo, Lucy Amigo-Dobaño, María Dolores Garza-Gil, and Manuel Varela-Lafuente. Perceptions in small-scale fisheries regarding institutional, economic, technological and environmental factors. Case study in North-Western Spain. Marine Policy, 150(??):Article 105530, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2300057X.

Garcia-Lorenzo:2021:CBF

[GLAVL21] Iria García-Lorenzo, Dewan Ahsan, and Manuel Varela-Lafuente. Community-based fisheries organisations and sustainable development: Lessons learned from a comparison between European and Asian countries. Marine Policy, 132(??):Article 104672, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002839.

Garren:2021:HPS

[GLS^{+}21] Melissa Garren, Forrest Lewis, Laura Sanchez, Daniella Spina, and Annie Brett. How performance standards could support innovation and technology-compatible fisheries management frameworks in the U.S. Marine Policy, 131(??):Article 104631, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002426.

Guo:2023:CAO

[GLT23] Jianping Guo, Wanqiang Li, and Haoyu Tian. The climate advisory opinion: a medicine with side-effects? Marine Policy, 156(??):Article 105817, October 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003500.

Gomez:2021:ASM

[GM21] Sílvia Gómez and Francesc Maynou. Alternative seafood marketing systems foster transformative processes in Mediterranean fisheries. Marine Policy, 127(??):Article 104432, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000427.

Guerra-Marrero:2021:PAI

[GMCMJA^{+}21] Airam Guerra-Marrero, Lorena Couce-Montero, David Jiménez-Alvarado, Ana Espino-Ruano, Raibel Núñez-González, Airam Sarmiento-Lezcano, Ángelo Santana del Pino, and José J. Castro. Preliminary assessment of the impact of Covid-19 Pandemic in the small-scale and recreational fisheries of the Canary Islands. Marine Policy, 133(??):Article 104712, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003237.

GoncalvesLoureiro:2023:OAA

[GMG^{+}23] Tainã Gonçalves Loureiro, Ben Milligan, Jordan Gacutan, Ibukun J. Adewumi, and Ken Findlay. Ocean accounts as an approach to foster, monitor, and report progress towards sustainable development in a changing ocean — the Systems and Flows Model. Marine Policy, 154(??):Article 105668, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001951.

Gonzales-Mantilla:2021:CCP

[GMGLV21] Pedro G. Gonzáles-Mantilla, Austin J. Gallagher, Carmelo J. León, and Gabriel M. S. Vianna. Challenges and conservation potential of shark-diving tourism in the Macaronesian archipelagos. Marine Policy, 131(??):Article 104632, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002438.

Gonzales-Mantilla:2022:EIC

[GMGLV22] Pedro G. Gonzáles-Mantilla, Austin J. Gallagher, Carmelo J. León, and Gabriel M. S. Vianna. Economic impact and conservation potential of shark-diving tourism in the Azores Islands. Marine Policy, 135(??):Article 104869, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004802.

Gutteridge:2024:MAD

[GML^{+}24] A. N. Gutteridge, M. C. Melnychuk, C. Longo, R. J. C. Currey, and E. Jardim. The MSC approach for developing harvest strategies and harvest control rules for Regional Fisheries Management Organisations (RFMOs). Marine Policy, 161(??):Article 106004, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000022.

Gatti:2020:RDD

[GMM^{+}20] Paul Gatti, Sonia Méhault, Fabien Morandeau, Marie Morfin, and Marianne Robert. Reducing discards of demersal species using a 100 mm square mesh cylinder: Size selectivity and catch comparison analysis. Marine Policy, 116(??):Article 103777, June 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307358.

Gamaza-Marquez:2020:DPG

[GMPT^{+}20] Maria Angeles Gamaza-Márquez, Maria Grazia Pennino, Maria Angeles Torres, Juan Jose Acosta, Karim Erzini, and Ignacio Sobrino. Discard practices in the Gulf of Cadiz multispecies trawl fishery. Implications for the EU ‘landing obligation’. Marine Policy, 118(??):Article 104008, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19302349.

Grace-McCaskey:2023:CSN

[GMRSP23] Cynthia A. Grace-McCaskey, Maria C. Ramos, Anja Sjostrom, and Sarah E. Page. Combining social network analysis and ethnography to better understand fishers’ organization and promote sustainable small-scale fisheries in St. Croix, US Virgin Islands. Marine Policy, 152(??):Article 105573, June 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001008.

Guzman:2023:MPA

[GMVM23] Diego Hernández Guzmán, Roberto Lastra Mier, Alba Vergara, and Celene B. Milanes. Marine protected areas in Colombia: a historical review of legal marine protection since the late 1960s to 2023. Marine Policy, 155(??):Article 105726, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002592.

Gibbs:2022:RHN

[GN22] Mark T. Gibbs and Maxine Newlands. Restoration heralds’ new management challenges for coral reefs. Marine Policy, 136(??):Article 104911, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005224.

Gaebel:2024:ISK

[GNJR24] Christine Gaebel, Paula Novo, David E. Johnson, and J. Murray Roberts. Institutionalising science and knowledge under the agreement for the conservation and sustainable use of marine biodiversity of areas beyond national jurisdiction (BBNJ): Stakeholder perspectives on a fit-for-purpose Scientific and Technical Body. Marine Policy, 161(??):Article 105998, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23005316.

Grillo-Nunez:2021:ICV

[GNMGWM21] J. Grillo-Núñez, T. Mendo, R. Gozzer-Wuest, and J. Mendo. Impacts of COVID-19 on the value chain of the hake small scale fishery in northern Peru. Marine Policy, 134(??):Article 104808, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2100419X.

Gray:2020:SCA

[GOH^{+}20] S. Gray, C. O’Mahony, J. Hills, B. O’Dwyer, R. Devoy, and J. Gault. Strengthening coastal adaptation planning through scenario analysis: a beneficial but incomplete solution. Marine Policy, 111(??):Article 102391, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302275.

Gordon:2021:PRP

[Gor21] Daniel V. Gordon. Price and revenue projections under alternative policy shocks due to the coronavirus: Canadian lobster and snow crab. Marine Policy, 130(??): Article 104556, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X21001676.

Gacutan:2022:EIB

[GPA^{+}22] Jordan Gacutan, Kemal Pınarbaşı, Messan Agbaglah, Crystal Bradley, Ibon Galparsoro, Arantza Murillas, Ibukun Adewumi, Teerapong Praphotjanaporn, Michael Bordt, Kenneth Findlay, Coulson Lantz, and Ben M. Milligan. The emerging intersection between marine spatial planning and ocean accounting: a global review and case studies. Marine Policy, 140(??):Article 105055, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001026.

Galparsoro:2021:OES

[GPG^{+}21] Ibon Galparsoro, Kemal Pınarbaşı, Elena Gissi, Fiona Culhane, Jordan Gacutan, Jonne Kotta, David Cabana, Sonja Wanke, Robert Aps, Daniel Bazzucchi, Giacomo Cozzolino, Marco Custodio, Mihhail Fetissov, Miguel Inácio, Susanna Jernberg, Alessandro Piazzi, Keshav Prasad Paudel, Alex Ziemba, and Daniel Depellegrin. Operationalisation of ecosystem services in support of ecosystem-based marine spatial planning: insights into needs and recommendations. Marine Policy, 131(??):Article 104609, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002207.

Garbett:2021:CEF

[GPH^{+}21] Amy Garbett, Natasha D. Phillips, Jonathan D. R. Houghton, Paulo Prodöhl, James Thorburn, Sophie L. Loca, Lawrence E. Eagling, Gary Hannon, Daniel Wise, Liz Pothanikat, Cat Gordon, Maurice Clarke, Peter Williams, Rebecca Hunter, Ronan McShane, Aafke Brader, Jane Dodd, Chris McGonigle, Heidi McIlvenny, Olivia Daly, Ronald Surgenor, Sarah Varian, Peter Verhoog, Gijs Van Zonneveld, Lylian R. Burke, Ian Davies, Terri A. Souster, Paul A. Mayo, Tanja N. Schwanck, Catherine S. Jones, and Patrick C. Collins. The critically endangered flapper skate (Dipturus intermedius): Recommendations from the first flapper skate working group meeting. Marine Policy, 124(??):Article 104367, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310186.

Gkargkavouzi:2020:PPM

[GPM20] Anastasia Gkargkavouzi, Stefanos Paraskevopoulos, and Steriani Matsiori. Public perceptions of the marine environment and behavioral intentions to preserve it: the case of three coastal cities in Greece. Marine Policy, 111(??):Article 103727, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19301290.

Gomez:2023:ISS

[GPM23] Sílvia Gómez, Beatriz Patraca, and José Luis Molina. Improving seafood systems with social network analysis: the case of cuttlefish marketing in Catalonia. Marine Policy, 150(??):Article 105517, April 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000441.

Gonzalez-Pestana:2023:VNP

[GPTASM23] Adriana Gonzalez-Pestana, Daniela C. S. Thorne, Joanna Alfaro-Shigueto, and Jeffrey C. Mangel. Vulnerabilities of northern Peruvian small-scale fishing communities revealed by the COVID-19 pandemic. Marine Policy, 149(??):Article 105503, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000301.

Gerwing:2020:CCP

[GR20a] Travis G. Gerwing and Joshua A. Rash. Constructive and collegial peer-review as a necessary precursor to data-driven environmental policy. Marine Policy, 111(??):Article 103721, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19303367.

Green:2020:CHS

[GR20b] Jessica F. Green and Bryce Rudyk. Closing the high seas to fishing: a club approach. Marine Policy, 115(??):Article 103855, May 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19303161.

Gray:2021:NTI

[Gra21] Tim Gray. Normative theory of international relations and the ‘mackerel war’ in the North East Atlantic. Marine Policy, 131(??):Article 104620, September 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002311.

Geiger:2022:CPE

[GRA^{+}22] Katja J. Geiger, Antonella Rivera, Alba Aguión, Jorge Álvarez, Julio Arrontes, Yaisel Juan Borrell, Teresa Cruz, Dominique Davoult, Jesús Dubert, Marieke E. Feis, Joana N. Fernandes, Consolación Fernández, Lucía García-Flórez, David Jacinto, Didier Jollivet, Gonzalo Macho, Elena Mateo, David Mateus, Paloma Morán, Carlota Muñiz, Amandine Nicolle, Rita Nolasco, Marina Parrondo, Henrique Queiroga, José Rico, Alina Sousa, Salvador Román, Teresa Silva, Eric Thiébaut, Elsa Vázquez, and José Luis Acuña. Coping with poachers in European stalked barnacle fisheries: Insights from a stakeholder workshop. Marine Policy, 135(??):Article 104826, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004371.

Groeneveld:2020:WEW

[Gro20] Rolf A. Groeneveld. Welfare economics and wicked problems in coastal and marine governance. Marine Policy, 117(??):Article 103945, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19303951.

Guirkinger:2021:MCP

[GRPE^{+}21] Lucie Guirkinger, Stefany Rojas-Perea, Isabel Ender, Mark Ramsden, Charley Lenton-Lyons, and Jonas Geldmann. Motivations for compliance in Peruvian manta ray fisheries. Marine Policy, 124(??):Article 104315, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309623.

Gonzalez-Roca:2023:AFP

[GRPRVG23] Fernando González-Roca, Ángel Pérez-Ruzafa, Julio A. Vásquez, and Stefan Gelcich. Artisanal fishers’ perceptions of two co-existing kelp resource management regimes in northern Chile. Marine Policy, 153(??):Article 105610, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001379.

Gallagher:2023:ALO

[GRS^{+}23] Anthony Gallagher, Peter Randall, David Sivyer, Umberto Binetti, Gayathri Lokuge, and Mohamed Munas. Abandoned, lost or otherwise discarded fishing gear (ALDFG) in Sri Lanka — a pilot study collecting baseline data. Marine Policy, 148(??):Article 105386, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200433X.

Gundelund:2021:CAD

[GS21] Casper Gundelund and Christian Skov. Changes in angler demography and angling patterns during the Covid-19 lockdown in spring 2020 measured through a citizen science platform. Marine Policy, 131(??):Article 104602, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2100213X.

Gangal:2023:HWD

[GSA23] Mayuresh Gangal, Venkatachalam Suri, and Rohan Arthur. How well does Indian fisheries policy engage with fisheries biology? Exploring the science-policy interface of coastal capture fisheries along the west coast of India. Marine Policy, 156(??):Article 105796, October 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003299.

Garmendia:2021:UFK

[GSAF21] Vladimir Garmendia, Maria Dulce Subida, Ainara Aguilar, and Miriam Fernández. The use of fishers’ knowledge to assess benthic resource abundance across management regimes in Chilean artisanal fisheries. Marine Policy, 127(??):Article 104425, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100035X.

Groba:2020:OTF

[GSB20] Carlos Groba, Antonio Sartal, and Gustavo Bergantiño. Optimization of tuna fishing logistic routes through information sharing policies: a game theory-based approach.

Marine Policy, 113(??):Article 103795, March 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306699.

Goldsmith:2020:DMM

[GSG20] William M. Goldsmith, Andrew M. Scheld, and John E. Graves. Decision-making in a mixed commercial-recreational fishery for Atlantic bluefin tuna. Marine Policy, 115(??): Article 103848, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19303495.

Gbedomon:2024:SSM

[GSG^{+}24] Rodrigue Castro Gbedomon, Kolawolé Valère Salako, Setondé Constant Gnansounou, Kisito Gandji, Pierre Failler, Achille Ephrem Assogbadjo, and Romain Glèlè Kakaï. Small-scale marine fishing in Benin, West Africa: a comprehensive assessment of the processed fish value chain. Marine Policy, 161(??):Article 106032, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000307.

Garcia-Sanabria:2021:MSP

[GSGOP^{+}21] Javier García-Sanabria, Javier García-Onetti, Víctor Cordero Penín, María de Andrés, Claudia Millán Caravaca, Eleonora Verón, and Cristina Pallero-Flores. Marine Spatial Planning cross-border cooperation in the ‘European Macaronesia Ocean’: a participatory approach. Marine Policy, 132(??):Article 104671, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002827.

Ginzky:2021:SIS

[GSM21] Harald Ginzky, Pradeep A. Singh, and Till Markus. Strengthening the International Seabed Authority’s knowledgebase: Addressing uncertainties to enhance decision-making. Marine Policy, ??(??):Article 103823, ???? 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19300661.

Garcia:2021:MRE

[GSR21] Pablo Quero García, Javier García Sanabria, and Juan Adolfo Chica Ruiz. Marine renewable energy and maritime spatial planning in Spain: Main challenges and recommendations. Marine Policy, 127(??):Article 104444, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000543.

Gamaza:2020:WRI

[GTA^{+}20] MariAngeles Gamaza, Maria Angeles Torres, Juan Jose Acosta, Karim Erzini, and Ignacio Sobrino. Are we ready to implement a discard ban in the Gulf of Cádiz? Stakeholders’ perceptions. Marine Policy, 116(??):Article 103711, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18307528.

Guillotreau:2020:ATP

[GTMP20] Patrice Guillotreau, Brice Trouillet, Stéphanie Mahévas, and Sophie Pardo. Addressing transdisciplinary and participation issues to cope with rapid changes shifting marine social ecological systems. Marine Policy, 117(??):Article 103929, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20302463.

Grover:2021:IPP

[GTTM21] Isobella M. Grover, Mark S. Tocock, Dugald R. Tinch, and Darla Hatton MacDonald. Investigating public preferences for the management of native and invasive species in the context of kelp restoration. Marine Policy, 132(??):Article 104680, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002918.

Guelker:2023:FSI

[Gue23] Daphne Guelker. Fishers and seafarers in international law — really so different? Marine Policy, 148(??):Article 105473, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22005243.

Guggisberg:2022:TAF

[Gug22] Solène Guggisberg. Transparency in the activities of the Food and Agriculture Organization for sustainable fisheries. Marine Policy, 136(??):Article 104498, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001093.

Guggisberg:2024:FES

[Gug24] Solène Guggisberg. Finding equitable solutions to the landbased sources of marine plastic pollution: Sovereignty as a double-edged sword. Marine Policy, 159(??):Article 105960, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004931.

Gulseven:2020:MAT

[Gul20] Osman Gulseven. Measuring achievements towards SDG 14, life below water, in the United Arab Emirates. Marine Policy, 117(??):Article 103972, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20300488.

Gullett:2023:CLS

[Gul23] Warwick Gullett. The contribution of the Law of the Sea Convention to marine environmental protection: 40 years of shimmering seas or high hopes dashed against juridical rocks? Marine Policy, 149(??):Article 105498, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000258.

Gunnarsson:2021:RST

[Gun21] Björn Gunnarsson. Recent ship traffic and developing shipping trends on the Northern Sea Route — Policy implications for future Arctic shipping. Marine Policy, 124(??):Article 104369, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310204.

Guo:2020:DME

[Guo20] Jianping Guo. The developments of marine environmental protection obligation in article 192 of UNCLOS and the operational impact on China’s marine policy — a South China Sea fisheries perspective. Marine Policy, 120(??):Article 104140, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19308917.

Gunnlaugsson:2022:SWC

[GV22] Stefan B. Gunnlaugsson and Hreidar Valtysson. Sustainability and wealth creation, but no consensus: Recent decades in Iceland’s ITQ-managed fisheries. Marine Policy, 135(??):Article 104836, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004474.

Gundelund:2023:SMC

[GVH^{+}23] Casper Gundelund, Paul Venturelli, Bruce W. Hartill, Kieran Hyder, Hans Jakob Olesen, and Christian Skov. Spatiotemporal and multispecies comparisons between a citizen science platform and recall surveys in recreational fisheries. Marine Policy, 155(??):Article 105780, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003135.

Gulbrandsen:2022:NLF

[GVL22] Lars H. Gulbrandsen, Irja Vormedal, and Mari Lie Larsen. No logo? The failure of ASC salmon labeling in Norway and the UK. Marine Policy, 138(??):Article 104987, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000343.

Greaker:2020:EPI

[GVR20] Mads Greaker, Irja Vormedal, and Kristin Rosendal. Environmental policy and innovation in Norwegian fish farming: Resolving the sea lice problem? Marine Policy, 117(??):Article 103942, July 2020. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19309686.

Guilan:2021:SMS

[GW21] Lin Guilan and Dong Weiwei. Synergetic management strategy for maritime cultural heritage protection and marine development in China. Marine Policy, 125(??):Article 104383, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20310344.

Gozzer-Wuest:2022:WRD

[GWAPRPRU22] Renato Gozzer-Wuest, Enrique Alonso-Población, Stefany Rojas-Perea, and Rubén H. Roa-Ureta. What is at risk due to informality? Economic reasons to transition to secure tenure and active co-management of the jumbo flying squid artisanal fishery in Peru. Marine Policy, 136(??):Article 104886, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004978.

Gozzer-Wuest:2021:IPA

[GWAPT21] Renato Gozzer-Wuest, Enrique Alonso-Población, and Geoffrey A. Tingley. Identifying priority areas for improvement in Peruvian Fisheries. Marine Policy, 129(??):Article 104545, July 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001561.

Guillot-Wright:2022:IED

[GWCD22] Shannon Guillot-Wright, Ellie Cherryhomes, and Lacy Davis. The impact of economic deregulation for health disparities among Gulf of Mexico commercial fishermen. Marine Policy, 141(??):Article 105099, July 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001464.

Gou:2023:DPI

[GY23] Yiping Gou and Cuibai Yang. Dilemmas and paths of international cooperation in China’s fight against IUU fishing analysis. Marine Policy, 155(??):Article 105789, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003226.

Greer:2020:CGT

[GZW^{+}20] Krista Greer, Dirk Zeller, Jessika Woroniak, Angie Coulter, Maeve Winchester, M. L. Deng Palomares, and Daniel Pauly. Corrigendum to ‘Global trends in carbon dioxide (CO_{2}) emissions from fuel combustion in marine fisheries from 1950 to 2016’ [Marine Policy, 107 (2019), 103382]. Marine Policy, 121(??):Article 104188, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308149.

Hoenicka:2022:RCH

[HACCM22] Ma. Alejandra Koeneke Hoenicka, Sara Andreotti, Humberto Carvajal-Chitty, and Conrad A. Matthee. The role of controlled human-animal interactions in changing the negative perceptions towards white sharks, in a sample of White Shark cage diving tours participants. Marine Policy, 143(??):Article 105130, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001774.

Hannesson:2020:NCA

[Han20] Rögnvaldur Hannesson. The Nash–Cournot approach to shared fish stocks: an empirical investigation. Marine Policy, 118(??):Article 103978, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309431.

Hanh:2021:WFA

[Han21] Tong Thi Hai Hanh. Why are fisheries agencies unable to facilitate the development of alternative livelihoods in small-scale fisheries and aquaculture in the global South? A case study of the Tam Giang lagoon, Viet Nam. Marine Policy, 133(??):Article 104778, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003894.

Hart:2021:SSG

[Har21] Paul J. B. Hart. Stewards of the sea. Giving power to fishers. Marine Policy, 126(??):Article 104421, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000312.

Harrison:2022:ITO

[Har22]  James Harrison.  International transparency obliga-

tions in fisheries conservation and management: Interstate and intra-state dimensions. Marine Policy, 136 (??):Article 104105, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19306153.

Hassanali:2020:CBE

[Has20] Kahlil Hassanali. CARICOM and the blue economy — Multiple understandings and their implications for global engagement. Marine Policy, 120(??):Article 104137, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20301123.

Hassanali:2023:ABA

[Has23] Kahlil Hassanali. The Agreement on Biodiversity in Areas Beyond National Jurisdiction — Caribbean Community perspectives on interests, asks and outcomes. Marine Policy, 156(??):Article 105800, October 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003330.

Hernandez-Blanco:2024:CAT

[HBCA^{+}24] Marcello Hernández-Blanco, Robert Costanza, Tundi Agardy, Joshua Farley, Elizabeth Fulton, Ida Kubiszewski, and Rifaee Rasheed. Common Asset Trusts for blue commons stewardship. Marine Policy, 159(??):Article 105957, January 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004906.

Deme:2022:SPF

[hBDDF22] El hadj Bara Deme, Moustapha Deme, and Pierre Failler. Small pelagic fish in Senegal: a multi-usage resource. Marine Policy, 141(??):Article 105083, July 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001300.

Deme:2022:PPS

[hBDF22] El hadj Bara Dème and Pierre Failler. Public policies to support artisanal fishing in Senegal: Between inconsistency and perverse effects. Marine Policy, 138(??):Article 105012, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000598.

Hurtubise:2020:EMR

[HBHJ20] Jessica A. Hurtubise, Emily P. Bond, Kelly E. Hall, and Michael C. James. Evaluating mandatory reporting of marine turtle bycatch in Atlantic Canadian fisheries. Marine Policy, 121(??):Article 104084, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930942X.

Hosch:2021:CET

[HBJ21] Gilles Hosch, Ben Belton, and Gareth Johnstone. Catch and effort trends in Myanmar’s offshore fleets operating out of Myeik — 2009–2018. Marine Policy, 123(??):Article 104298, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2030943X.

Huse:2021:CSF

[HBL^{+}21] Geir Huse, Åsmund Bjordal, Harald Loeng, Kari Østervold Toft, and Reidar Toresen. Commitments to sustainable fisheries: Empty words or reality? Marine Policy, 124(??):Article 104360, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310113.

Hoque:2021:KPE

[HBP^{+}21] Md. Sazedul Hoque, Despena Andrioti Bygvraa, Kate Pike, Md. Mahmudul Hasan, Muhammad Arifur Rahman, Shaida Akter, Darren Mitchell, and Eric Holliday. Knowledge, practice, and economic impacts of COVID-19 on small-scale coastal fishing communities in Bangladesh: Policy recommendations for improved livelihoods. Marine Policy, 131(??):Article 104647, September 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100258X.

Heck:2021:SRM

[HBR21] Nadine Heck, Michael W. Beck, and Borja Reguero. Storm risk and marine fisheries: a global assessment. Marine Policy, 132(??):Article 104698, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003092.

Heck:2023:GCC

[HBR^{+}23] Nadine Heck, Michael W. Beck, Borja Reguero, Kerstin Pfliegner, Max Ricker, and Ruben Prütz. Global climate change risk to fisheries — a multi-risk assessment. Marine Policy, 148(??):Article 105404, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004511.

Hsiao:2021:DPS

[HC21] Yao-Jen Hsiao and Jyun-Long Chen. Different perspectives of stakeholders on the sustainable development of fishery-based communities in Northeast Taiwan. Marine Policy, 130(??):Article 104576, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001871.

Ho:2022:PRC

[HC22] Li-Te Ho and Kuei-Chao Chang. Promoting relational collaborative partnership to implement coastal settlement revitalization. Marine Policy, 146(??):Article 105242, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002895.

Choi:2024:TBW

[hC24] Jee hyun Choi. Trans-boundary water management during the peace process in Korean peninsula: Solution finding from previous precedents. Marine Policy, 160(??):Article 105966, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004992.

Huang:2021:SPT

[HCS21] Chao-Chin Huang, Shui-Kai Chang, and Shiahn-Wern Shyue. Sustain or phase out: Transformation of Taiwan’s management scheme on distant water tuna longline fisheries. Marine Policy, 123(??):Article 104297, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309428.

Hernandez-Casas:2023:ASD

[HCSBM^{+}23] Sergio Hernández-Casas, Juan Carlos Seijo, Luis Felipe Beltrán-Morales, Álvaro Hernández-Flores, Francisco Arreguín-Sánchez, and Germán Ponce-Díaz. Analysis of supply and demand in the international market of major abalone fisheries and aquaculture production. Marine Policy, 148(??):Article 105405, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004523.

Hung:2021:IMP

[HCT^{+}21] Yi-Ping Hung, Yi Chang, Minh-Phuong N. Truong, Jen-Han Yang, Tung-Yao Hsu, and Shih-Chun Hsiao. Improving marine protected area with coordination platform: Mud shrimp conservation in Taiwan case study. Marine Policy, 131(??):Article 104607, September 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002189.

Han:2023:ICC

[HCW^{+}23] Wenjie Han, Jihong Chen, Kui Wei, Jia Shi, and Guoling Jia. International crew changes amid global pandemic outbreaks: Key issues and system innovations. Marine Policy, 147(??):Article 105342, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200389X.

Haugen:2021:RAC

[HCWC21] Brianna I. Haugen, Lori A. Cramer, George G. Waldbusser, and Flaxen D. L. Conway. Resilience and adaptive capacity of Oregon’s fishing community: Cumulative impacts of climate change and the graying of the fleet. Marine Policy, 126(??):Article 104424, April 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000348.

Hernandez-Delgado:2024:SPC

[HDAPG^{+}24] Fredy Hernández-Delgado, Alfonso Aguilar-Perera, Vinicius J. Giglio, Virginia Nóh-Quiñones, Jorge I. Euán-Ávila, Wilian de Jesús Aguilar-Cordero, and Celia I. Sélem-Salas. Stakeholders’ perception on consumption, fishing, and conservation of red grouper, Epinephelus morio, off the northern coast of the Yucatan Peninsula, Mexico. Marine Policy, 161(??):Article 105999, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23005328.

Haas:2021:RFMb

[HDCH21] Bianca Haas, Ruth Davis, Brooke Campbell, and Quentin Hanich. Regional fisheries management: COVID-19 calendars and decision making. Marine Policy, 128(??):Article 104474, June 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000853.

Haas:2021:RFMa

[HDH21] Bianca Haas, Ruth Davis, and Quentin Hanich. Regional fisheries management: Virtual decision making in a pandemic. Marine Policy, 125(??):Article 104288, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309349.

Harden-Davies:2020:RNP

[HDHM^{+}20] Harriet Harden-Davies, Fran Humphries, Michelle Maloney, Glen Wright, Kristina Gjerde, and Marjo Vierros. Rights of Nature: Perspectives for Global Ocean Stewardship. Marine Policy, 122(??):Article 104059, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20303365.

Hau:2023:MMM

[HdM23] C. Y. Hau. and Y. J. Sadovy de Mitcheson. Mortality and management matter: Case study on use and misuse of ‘ranching’ for a CITES Appendix II-listed fish, humphead wrasse (Cheilinus undulatus). Marine Policy, 149(??):Article 105515, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000428.

Hersoug:2022:OCT

[Her22] Bjørn Hersoug. “One country, ten systems” — the use of different licensing systems in Norwegian aquaculture. Marine Policy, 137(??):Article 104902, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005133.

Haya:2020:ACR

[HF20] La Ode Muhammad Yasir Haya and Masahiko Fujii. Assessment of coral reef ecosystem status in the Pangkajene and Kepulauan Regency, Spermonde Archipelago, Indonesia, using the rapid appraisal for fisheries and the analytic hierarchy process. Marine Policy, 118(??):Article 104028, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1930524X.

Hair:2020:SEC

[HFD^{+}20] Cathy Hair, Simon Foale, Nicholas Daniels, Peter Minimulu, John Aini, and Paul C. Southgate. Social and economic challenges to community-based sea cucumber mariculture development in New Ireland Province, Papua New Guinea. Marine Policy, 117(??):Article 103940, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307316.

Haghighatjou:2022:RAA

[HGG^{+}22a] N. Haghighatjou, S. Gorgin, R. Ghorbani, E. Gilman, R. A. Naderi, H. Raeisi, and Sh Farrukhbin. Rate and amount of abandoned, lost and discarded gear from the Iranian Persian Gulf Gargoor pot fishery. Marine Policy, 141(??):Article 105100, July 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001476.

Hamilton:2022:EBM

[HGG^{+}22b] Sara L. Hamilton, Mary G. Gleason, Natalio Godoy, Norah Eddy, and Kirsten Grorud-Colvert. Ecosystembased management for kelp forest ecosystems. Marine Policy, 136(??):Article 104919, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005303.

HortaeCosta:2022:OCN

[HGG22c] Barbara Horta e Costa, Jorge M. S. Gonçalves, and Emanuel J. Gonçalves. UN Ocean Conference needs transparent and science-based leadership on ocean conservation. Marine Policy, 143(??):Article 105197, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002445.

Hill:2020:ASA

[HGH^{+}20] Nicholas J. Hill, Caleb Gardner, Malcolm Haddon, Klaas Hartmann, L. Richard Little, Jeremy M. Lyle, and Bradley R. Moore. Assessing the status of Australia’s fish stocks relative to target objectives. Marine Policy, 112(??):Article 103741, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19301952.

Humphries:2020:TAM

[HGL^{+}20] Fran Humphries, Hiroko Muraki Gottlieb, Sarah Laird, Rachel Wynberg, Charles Lawson, Michelle Rourke, Morten Walløe Tvedt, Maria Julia Oliva, and Marcel Jaspars. A tiered approach to the marine genetic resource governance framework under the proposed UNCLOS agreement for biodiversity beyond national jurisdiction (BBNJ). Marine Policy, 122(??):Article 103910, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308218.

Hendrix:2022:GCN

[HGLR22] Cullen S. Hendrix, Sarah M. Glaser, Joshua E. Lambert, and Paige M. Roberts. Global climate, El Niño, and militarized fisheries disputes in the East and South China Seas. Marine Policy, 143(??):Article 105137, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001841.

Haas:2023:FFU

[HGSD23] Bianca Haas, Camille Goodman, Hussain Sinan, and Ruth Davis. Fact or fiction? Unpacking the terminologies used in fisheries allocation discussions. Marine Policy, 152(??):Article 105630, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001574.

Hebbar:2021:BRM

[HH21] Anish Arvind Hebbar and Shantanu Anish Hebbar. Book review: The Maritime Silk Road: China’s Belt and Road at Sea, Richard T. Griffiths. International Institute for Asian Studies (2020), pp232, ISBN: 978-90-823810-4-7(e). Marine Policy, 129(??):Article 104504, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001159.

Hasan:2021:ICF

[HHB^{+}21] Neaz A. Hasan, Richard D. Heal, Abul Bashar, Alif Layla Bablee, and Mohammad Mahfujul Haque. Impacts of COVID-19 on the finfish aquaculture industry of Bangladesh: a case study. Marine Policy, 130(??):Article 104577, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X21001883.

Healy:2020:GRE

[HHBC20] Teleah Joy Healy, Nicholas James Hill, Adam Barnett, and Andrew Chin. A global review of elasmobranch tourism activities, management and risk. Marine Policy, 118(??):Article 103964, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19302143.

Humphries:2020:PPS

[HHD20] Fran Humphries and Harriet Harden-Davies. Practical policy solutions for the final stage of BBNJ treaty negotiations. Marine Policy, 122(??):Article 104214, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20308605.

Hooper:2020:PPT

[HHEJB20] Tara Hooper, Caroline Hattam, Andrew Edwards-Jones, and Nicola Beaumont. Public perceptions of tidal energy: Can you predict social acceptability across coastal communities in England? Marine Policy, 119(??):Article 104057, September 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20303791.

Hill:2023:PEF

[HHH^{+}23] Nicholas J. Hill, Malcolm Haddon, Klaas Hartmann, L. Richard Little, Jeremy M. Lyle, Bradley R. Moore, and Simon Nicol. Prevalence of effective fisheries management in a developed nation and its link to stock sustainability. Marine Policy, 151(??):Article 105596, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001239.

Haapasaari:2021:IGM

[HIP^{+}21] Päivi Haapasaari, Suvi Ignatius, Mia Pihlajamäki, Andreas Bryhn, Simo Sarkki, Jouni Tuomisto, Lauri Nevalainen, Annukka Lehikoinen, Timo Assmuth, Atso Romakkaniemi, Heikki Peltonen, and Sakari Kuikka. Integrated governance for managing multidimensional problems: Potentials, challenges, and arrangements. Marine Policy, 123(??):Article 104276, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309222.

Hammond:2021:PBH

[HJ21] Amy Hammond and Peter J. S. Jones. Protecting the ‘blue heart of the planet’: Strengthening the governance framework for marine protected areas beyond national jurisdiction. Marine Policy, 127(??):Article 104260, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309064.

Hagspiel:2024:CLB

[HJLN24] Verena Hagspiel, Markus Bjørkli Jansen, Maria Lavrutich, and Gaute Nepstad. Company level biodiversity impact assessment: an application to the aquaculture industry in Norway. Marine Policy, 159(??):Article 105899, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004323.

Han:2023:PHS

[HJPL23] Zenglin Han, Kaiping Jiang, Fei Peng, and Shuqin Li. The Philippines’ hedging strategy against China in the South China Sea dispute: Based on the human-ocean regional system. Marine Policy, 151(??):Article 105578, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001057.

Heidkamp:2022:GCC

[HKKY22] C. Patrick Heidkamp, Louie Vincent Krak, Miriah M. Russo Kelly, and Charles Yarish. Geographical considerations for capturing value in the U.S. sugar kelp (Saccharina latissima) industry. Marine Policy, 144(??):Article 105221, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002688.

Heinrich:2021:QFC

[HKMS21] Luise Heinrich, Andrea Koschinsky, Till Markus, and Pradeep Singh. Quantifying the fuel consumption, greenhouse gas emissions and air pollution of a potential commercial manganese nodule mining operation. Marine Policy, ??(??):Article 103678, ???? 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19300818.

Harris:2021:UCS

[HLCM21] Lucas Harris, Max Liboiron, Louis Charron, and Charles Mather. Using citizen science to evaluate extended producer responsibility policy to reduce marine plastic debris shows no reduction in pollution levels. Marine Policy, 123(??):Article 104319, January 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309660.

Hermann:2022:ATH

[HLLK22] Roberto Rivas Hermann, Ning Lin, Julien Lebel, and Alina Kovalenko. Arctic transshipment hub planning along the Northern Sea Route: a systematic literature review and policy implications of Arctic port infrastructure. Marine Policy, 145(??):Article 105275, November 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003220.

Hu:2020:CMP

[HLM^{+}20] Wenjia Hu, Jie Liu, Zhiyuan Ma, Yuyu Wang, Dian Zhang, Weiwei Yu, and Bin Chen. China’s marine protected area system: Evolution, challenges, and new prospects. Marine Policy, 115(??):Article 103780, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303987.

Hume:2021:TOB

[HLO^{+}21] Andrew Hume, Jim Leape, Kirsten L. L. Oleson, Emily Polk, Kevin Chand, and Robert Dunbar. Towards an ocean-based large ocean states country classification. Marine Policy, 134(??):Article 104766, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003778.

Halim:2020:TTM

[HLW^{+}20] Abdul Halim, Neil R. Loneragan, Budy Wiryawan, Rod Fujita, Dedi S. Adhuri, Adrian R. Hordyk, and M. Fedi A. Sondita. Transforming traditional management into contemporary territorial-based fisheries management rights for smallscale fisheries in Indonesia. Marine Policy, 116(??):Article 103923, June 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19309297.

Hao:2023:UHC

[HLW^{+}23] Huang Hao, Cai Lei, Lei Wang, Weiwen Li, Ou Danyun, Wu Lijun, Ou Rimei, and Lin Lianghua. Using habitat classification and large-scale spatial planning model to identify environmental areas: a case study in the area of the Northwest Pacific Ocean. Marine Policy, 147(??):Article 105360, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004079.

Huang:2021:CLS

[HM21] Yi-Chen Huang and Leslie Mabon. Coastal landscapes, sustainable consumption and peripheral communities: Evaluating the Miramar Resort controversy in Shanyuan Bay, Taiwan. Marine Policy, 123(??):Article 104283, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309295.

Hassanali:2022:EPG

[HM22a] Kahlil Hassanali and Robin Mahon. Encouraging proactive governance of marine biological diversity of areas beyond national jurisdiction through Strategic Environmental Assessment (SEA). Marine Policy, 136(??):Article 104932, February 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21005431.

Holmes:2022:HSA

[HM22b] Glen Holmes and Shana Miller. Harvest strategies and allocation — a chicken and egg scenario? Marine Policy, 135(??):Article 104871, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004826.

Hosch:2022:KNT

[HM22c] G. Hosch and G. Macfadyen. Killing Nemo: Three world regions fail to mainstream combatting of IUU fishing. Marine Policy, 140(??):Article 105073, June 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001208.

Haas:2020:FIP

[HMFH20] Bianca Haas, Jeffrey McGee, Aysha Fleming, and Marcus Haward. Factors influencing the performance of regional fisheries management organizations. Marine Policy, 113(??):Article 103787, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19303653.

Harvey:2021:BWA

[HMG^{+}21a] Euan S. Harvey, Dianne L. McLean, Jordan S. Goetze, Benjamin J. Saunders, Tim J. Langlois, Jacquomo Monk, Neville Barrett, Shaun K. Wilson, Thomas H. Holmes, Daniel Ierodiaconou, Alan R. Jordan, Mark G. Meekan, Hamish A. Malcolm, Michelle R. Heupel, David Harasti, Charlie Huveneers, Nathan A. Knott, David V. Fairclough, Leanne M. Currey-Randall, Michael J. Travers, Ben T. Radford, Matthew J. Rees, Conrad W. Speed, Corey B. Wakefield, Mike Cappo, and Stephen J. Newman. The BRUVs workshop — an Australia-wide synthesis of baited remote underwater video data to answer broad-scale ecological questions about fish, sharks and rays. Marine Policy, 127(??):Article 104430, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000403.

Hermoso:2021:EDE

[HMG^{+}21b] María Isabel Hermoso, Victoria Y. Martin, Stefan Gelcich, Wolfgang Stotz, and Martin Thiel. Exploring diversity and engagement of divers in citizen science: Insights for marine management and conservation. Marine Policy, 124(??):Article 104316, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309635.

Hosch:2023:ISH

[HMY^{+}23] Gilles Hosch, Nathan A. Miller, Yann Yvergniaux, Elaine Young, and Tim Huntington. IUU safe havens or PSMA ports: a global assessment of port state performance and risk. Marine Policy, 155(??):Article 105751, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002841.

Hogg:2021:LTN

[HNMSG21] Katie Hogg, Pedro Noguera-Méndez, and María Semitiel-García. Lessons from three north-western Mediterranean MPAs: a governance analysis of Port-Cros National Park, Tavolara Punta-Coda Cavallo and Ustica. Marine Policy, 127(??):Article 102943, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307388.

Honniball:2020:PSC

[Hon20] Arron N. Honniball. Port states, coastal states and national security: a law of the sea perspective on the 2017 Qatar-Gulf crisis. Marine Policy, 116(??):Article 103817, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19304889.

Hong:2021:CAP

[Hon21] Nong Hong. China and the Antarctic: Presence, policy, perception, and public diplomacy. Marine Policy, 134(??):Article 104779, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003900.

Hornborg:2023:FHC

[Hor23] Sara Hornborg. Follow the herring — a case study on the interplay between management and markets for marine resource utilization. Marine Policy, 158(??):Article 105874, December 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004074.

Howson:2020:BTE

[How20] Peter Howson. Building trust and equity in marine conservation and fisheries supply chain management with blockchain. Marine Policy, 115(??):Article 103873, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19307067.

Hsuan:2020:MSC

[HP20] Juliana Hsuan and Cristiana Parisi. Mapping the supply chain of ship recycling. Marine Policy, 118(??): Article 103979, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19305858.

Hsu:2023:UVS

[HP23] Kang Hsu and Li-Pei Peng. Understanding vulnerability and sustainable livelihood factors from coastal residents in Taiwan. Marine Policy, 155(??):Article 105793, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003263.

Herrera:2023:PPM

[HPC^{+}23] Mariana Herrera, Pablo Pita, Daniela Castelo, C. Marisa R. Almeida, Sandra Ramos, and Sebastián Villasante. Public perceptions of marine litter and impacts on coastal ecosystem services in Galicia (Spain). Marine Policy, 155(??):Article 105742, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002750.

Haas:2020:MRS

[HPG20] Bianca Haas, Michelle Phillipov, and Fred Gale. Media representations of seafood certification in Australia: Mobilising sustainability standards to attack or defend the value of an industry. Marine Policy, 120(??):Article 104126, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20303341.

He:2022:DDT

[HPH22] Xinhua He, Qiuyun Ping, and Wenfa Hu. Does digital technology promote the sustainable development of the marine equipment manufacturing industry in China? Marine Policy, 136(??):Article 104868, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004796.

Heck:2022:VSE

[HPPH22] Nadine Heck, Adina Paytan, Donald Potts, and Brent Haddad. Variations in stakeholders’ ecosystem service priorities for managing a marine protected area. Marine Policy, 146(??):Article 105330, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003773.

Ha:2023:UCD

[HPS23] Min-Ho Ha, Ho Park, and Young-Joon Seo. Understanding core determinants in LNG bunkering port selection: Policy implications for the maritime industry. Marine Policy, 152(??):Article 105608, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001355.

Hopkins:2024:ITS

[HRC^{+}24] C. R. Hopkins, S. I. Roberts, A. J. Caveen, C. Graham, and N. M. Burns. Improved traceability in seafood supply chains is achievable by minimising vulnerable nodes in processing and distribution networks. Marine Policy, 159(??):Article 105910, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004438.

Herrera-Racionero:2022:RSM

[HRMPCRRA22] Paloma Herrera-Racionero, Luís Miret-Pastor, Roberto Cervelló-Royo, and Miguel Rodilla-Alama. The role of the Spanish Mediterranean Fisher’s Guilds in maritime sustainability. Marine Policy, 140(??):Article 105058, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001051.

Haraldsson:2020:HMS

[HRR^{+}20] Matilda Haraldsson, Aurore Raoux, Fabien Riera, Julien Hay, Jeffrey M. Dambacher, and Nathalie Niquil. How to model social-ecological systems? — A case study on the effects of a future offshore wind farm on the local society and ecosystem, and whether social compensation matters. Marine Policy, 119(??):Article 104031, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306530.

Hellstrom:2024:GER

[HRS^{+}24] Magnus Hellström, Rodrigo Rabetino, Henry Schwartz, Anastasia Tsvetkova, and Syed Hammad Ul Haq. GHG emission reduction measures and alternative fuels in different shipping segments and time horizons — a Delphi study. Marine Policy, 160(??):Article 105997, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23005304.

Hai:2020:ISS

[HS20] Au Ton Nu Hai and Stijn Speelman. Involving stakeholders to support sustainable development of the marine lobster aquaculture sector in Vietnam. Marine Policy, 113(??):Article 103799, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19305883.

Hastings:2022:MSF

[HS22] Justin V. Hastings and Annie Young Song. The maritime security fallout from North Korea’s fisheries policy. Marine Policy, 136(??):Article 104915, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005261.

Hickey:2021:IOT

[HSdT21] Gordon M. Hickey, Hunter T. Snyder, Jasper R. deVries, and Owen Temby. On inter-organizational trust, control and risk in transboundary fisheries governance. Marine Policy, 134(??):Article 104772, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003833.

Hogg:2021:GAC

[HSGNMGC21] Katie Hogg, María Semitiel-García, Pedro Noguera-Méndez, and Jose Antonio García-Charton. A governance analysis of Cabo de Palos-Islas Hormigas and Cabo de Gata–Níjar Marine Protected Areas, Spain. Marine Policy, 127(??):Article 102944, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X17307327.

Howard:2023:BCP

[HSGS^{+}23] Jennifer Howard, Ariana E. Sutton-Grier, Lindsey S. Smart, Christian C. Lopes, Jill Hamilton, Joan Kleypas, Stefanie Simpson, Jennifer McGowan, Albert Pessarrodona, Heidi K. Alleway, and Emily Landis. Blue carbon pathways for climate mitigation: Known, emerging and unlikely. Marine Policy, 156(??):Article 105788, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003214.

Hausner:2021:DSO

[HSH^{+}21] Arjun Hausner, Jameal F. Samhouri, Elliott L. Hazen, Delgerzaya Delgerjargal, and Briana Abrahms. Dynamic strategies offer potential to reduce lethal ship collisions with large whales under changing climate conditions. Marine Policy, 130(??):Article 104565, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001767.

Huan:2024:REE

[HSHS24] Xiaomeng Huan, Jingzhu Shan, Limin Han, and Haiming Song. Research on the efficacy and effect assessment of deep-sea aquaculture policies in China: Quantitative analysis of policy texts based on the period 2004–2022. Marine Policy, 160(??):Article 105963, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004967.

Hsiao:2020:OFE

[Hsi20] Amanda Hsiao. Opportunities for fisheries enforcement cooperation in the South China Sea. Marine Policy, 121(??):Article 103569, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19303689.

Hitchin:2023:TDS

[HSK^{+}23] B. Hitchin, S. Smith, K. Kröger, DOB Jones, A. Jaeckel, NC Mestre, J. Ardron, E. Escobar, J. van der Grient, and T. Amaro. Thresholds in deep-seabed mining: a primer for their development. Marine Policy, 149(??):Article 105505, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000325.

Horta:2023:BFF

[HSM^{+}23] Paulo Horta, Marina Sissini, Carolina Melissa Mueller, Fernanda M. M. Soares, Paulo Pagliosa, Leonardo Rörig, José Bonomi-Barufi, Flavio Berchez, Leticia Cotrim da Cunha, Rodrigo Kerr, Sergio Rossi, Marcelo O. Soares, Jorge L. Rodrigues-Filho, Hudson T. Pinheiro, Luciano Henning, Marcos Aurélio Espindola, Nicole Figueiredo de Oliveira, and Alessandra Larissa Fonseca. Brazil fosters fossil fuel exploitation despite climate crises and the environmental vulnerabilities. Marine Policy, 148(??):Article 105423, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004705.

Hossain:2021:SFS

[HSN^{+}21] M. Shahadat Hossain, S. M. Sharifuzzaman, M. Nur Nobi, M. Shah Nawaz Chowdhury, Subrata Sarker, M. Alamgir, Sheikh Aftab Uddin, Sayedur Rahman Chowdhury, M. Mizanur Rahman, M. Shajjadur Rahman, Faisal Sobhan, and Showmitra Chowdhury. Seaweeds farming for sustainable development goals and blue economy in Bangladesh. Marine Policy, 128(??):Article 104469, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21000804.

Heinrich:2024:MGC

[HSSM24] Luise Heinrich, Pradeep Singh, Karen Smith Stegen, and Till Markus. Mind the gap and close it: Regulating greenhouse gas emissions from deep-sea mining in the area. Marine Policy, 160(??):Article 105929, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004621.

Hirokawa:2023:INS

[HT23] Toya Hirokawa and Benjamin S. Thompson. The influence of new sustainable fisheries policies on seafood company practices and consumer awareness in Japan. Marine Policy, 157(??):Article 105819, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003524.

Herrera:2023:DIB

[HTP^{+}23] Mariana Herrera, Ana Tubío, Pablo Pita, Elsa Vázquez, Celia Olabarria, Andrés Simón, José Carlos Mariño Balsa, Liliana Solís, Ignacio Gianelli, and Sebastian Villasante. Disentangling interactions between seagrasses and smallscale fisheries using scientific and local traditional knowledge. Marine Policy, 155(??):Article 105741, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002749.

Hatenboer:2023:DFS

[HvdBH23] Constant Hatenboer, Caspar van den Berg, and Ronald Holzhacker. The Dutch fisheries sector and the North Sea Accord: Unpacking stakeholder participation in multilevelled marine governance. Marine Policy, 147(??):Article 105364, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X22004110.

Herttua:2021:AAO

[HVN21] Kimmo Herttua, Jan Vork, and Jesper Bo Nielsen. Associations of age, occupation and nationality of seafarers with risk of evacuations at sea: a register-based study on Danish-flagged merchant ships. Marine Policy, 130(??):Article 103989, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X1930908X.

Hoshino:2020:DQO

[HvPPV20a] Eriko Hoshino, Ingrid van Putten, Sean Pascoe, and Simon Vieira. Does quota ownership affect perceptions of fishery performance? Marine Policy, 120(??):Article 104155, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20304589.

Hoshino:2020:ITQ

[HvPPV20b] Eriko Hoshino, Ingrid van Putten, Sean Pascoe, and Simon Vieira. Individual transferable quotas in achieving multiple objectives of fisheries management. Marine Policy, 113(??):Article 103744, March 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930346X.

Halik:2022:DCR

[HVS22] Abdul Halik, Marco Verweij, and Achim Schlüter. Deliberating coral reef protection — Cultural Theory tested. Marine Policy, 139(??):Article 105036, May 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000835.

Haapasaari:2022:FRN

[HvT22] Päivi Haapasaari and Jan P. M. van Tatenhove. A Finnish regional non-binding MSP approach: What are the consequences for integrating Blue Growth and GES? Marine Policy, 141(??):Article 105101, July 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001488.

Hayes:2022:CMC

[HW22] Christopher Hayes and Lisa Wainger. Comparing multicriteria decision analysis to group negotiations in fisheries co-management. Marine Policy, 138(??):Article 104997, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000446.

Haensch:2022:CAT

[HWM22] Juliane Haensch, Sarah Ann Wheeler, and Stephanie McWhinnie. Community attitudes towards marine parks in South Australia. Marine Policy, 139(??):Article 104121, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20301457.

Hori:2020:CAS

[HWMO20] Juri Hori, Hiroki Wakamatsu, Tsutom Miyata, and Yoshioki Oozeki. Has the consumers awareness of sustainable seafood been growing in Japan? Implications for promoting sustainable consumerism at the Tokyo 2020 Olympics and Paralympics. Marine Policy, 115(??):Article 103851, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19304816.

Hanzawa:2023:SSA

[HWT23] Yudai Hanzawa, Shingo Watari, and Shion Takemura. Not satisfactory, but still acceptable: Exploring socioeconomic incentives of individual quota systems in Japanese Pacific bluefin tuna fisheries from the management leaders’ perspective. Marine Policy, 150(??):Article 105501, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000283.

Hakala:2023:GSI

[HWU^{+}23] Siri Hakala, Shingo Watari, Shinji Uehara, Yujiro Akatsuka, Richard Methot, and Yoshi Oozeki. Governance and science implementation in fisheries management in Japan as it compares to the United States. Marine Policy, 155(??):Article 105670, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001975.

Hung:2022:CBG

[HWY22] Ling-Ya Hung, Shun-Mei Wang, and Ting-Kuang Yeh. Collaboration between the government and environmental non-governmental organisations for marine debris policy development: the Taiwan experience. Marine Policy, 135(??):Article 104849, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004607.

Halimatussadiah:2023:CML

[HYN^{+}23] Alin Halimatussadiah, Satoshi Yamazaki, Rachel Nichols, Umi Muawanah, Fachry Abdul Razak Afifi, and Muhammad Adriansyah. Co-management of large-scale and medium-scale fisheries: an assessment of the fishery tax system in Indonesia. Marine Policy, 148(??):Article 105458, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2200505X.

He:2023:GSI

[HZ23] Yixiong He and Fengxuan Zhang. A game study on the implementation of marine carbon sink fisheries in the context of carbon neutrality — analysis of the tripartite behavior of fishery practitioners, research institutions, and the government. Marine Policy, 147(??):Article 105365, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004122.

Iversen:2020:GDF

[IAB^{+}20] Audun Iversen, Frank Asche, Marcus Buck, Edgar Henriksen, Jonas Stein, and Sigbjørn Svalestuen. The growth and decline of fisheries communities: Explaining relative population growth at municipality level. Marine Policy, 112(??):Article 103776, February 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304555.

Inarra:2020:EUU

[IBC^{+}20] Bruno Iñarra, Carlos Bald, Marta Cebrián, Irene Peral, Raquel Llorente, and Jaime Zufía. Evaluation of unavoidable unwanted catches valorisation options: the Bay of Biscay case study. Marine Policy, 116(??):Article 103680, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18308455.

Inguanzo:2021:RPF

[IGO21] Belén Inguanzo, María-José Gutiérrez, and Susan Orbe. The role of production factors on landings heterogeneity between EU countries. Marine Policy, 132(??):Article 104679, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002906.

Islam:2023:DDS

[IGS^{+}23] Samsul Islam, Floris Goerlandt, Quazi Mohammed Habibus Sakalayen, Yangyan Shi, and V. G. Venkatesh. Developing a “Disaster Scenario” to prepare for the possibility of disruptions to maritime transportation serving coastal communities of Vancouver Island. Marine Policy, 150(??):Article 105531, April 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000581.

Islam:2021:INC

[IKB21] Md Monirul Islam, Makidul Islam Khan, and Aparna Barman. Impact of novel coronavirus pandemic on aquaculture and fisheries in developing countries and sustainable recovery plans: Case of Bangladesh. Marine Policy, 131(??):Article 104611, September 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002220.

Iue:2022:DBS

[IMA22] Minako Iue, Mitsutaku Makino, and Misuzu Asari. The development of “Blue Seafood Guide,” a sustainable seafood rating program, and its implication in Japan. Marine Policy, 137(??):Article 104945, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100556X.

Islam:2020:FCM

[INW20] Mohammad Mahmudul Islam, Md. Nahiduzzaman, and Md. Abdul Wahab. Fisheries co-management in hilsa shad sanctuaries of Bangladesh: Early experiences and implementation challenges. Marine Policy, 117(??):Article 103955, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19303380.

Ibengwe:2023:CIM

[IO23] Lilian J. Ibengwe and Paul O. Onyango. Capturing the invisible: Methodological approach in understanding informal cross border fish trade. Marine Policy, 149(??):Article 105482, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2300009X.

Ibengwe:2022:RTI

[IOHC22] Lilian J. Ibengwe, Paul O. Onyango, Aloyce S. Hepelwa, and Martin J. Chegere. Regional trade integration and its relation to income and inequalities among Tanzanian marine dagaa fishers, processors and traders. Marine Policy, 137(??):Article 104975, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000227.

Islam:2021:TAC

[IRK^{+}21] Md. Monirul Islam, Md. Asadur Rahman, Mohammad Shahneawz Khan, Gouri Mondal, and Makidul Islam Khan. Transformational adaptations to climatic hazards: Insights from mangroves-based coastal fisheries dependent communities of Bangladesh. Marine Policy, 128(??):Article 104475, June 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000865.

Israngkura:2022:MRR

[Isr22] Adis Israngkura. Marine resource recovery in southern Thailand during COVID-19 and policy recommendations. Marine Policy, 137(??):Article 104972, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000197.

Jones:2020:ECI

[JACD20] Daniel O. B. Jones, Jeff A. Ardron, Ana Colaço, and Jennifer M. Durden. Environmental considerations for impact and preservation reference zones for deep-sea polymetallic nodule mining. Marine Policy, 118(??):??, August 2020.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301374.

Jimenez:2020:VCD

[JAdS^{+}20] Érica Antunes Jimenez, Marilu Teixeira Amaral, Pauliana Leão de Souza, Maria de Nazaré Ferreira Costa, Alex Souza Lira, and Flávia Lucena Frédou. Value chain dynamics and the socioeconomic drivers of small-scale fisheries on the Amazon coast: a case study in the state of Amapá, Brazil. Marine Policy, 115(??):Article 103856, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301428.

Jaeckel:2021:SEP

[Jae21] Aline Jaeckel. Strategic environmental planning for deep seabed mining in the area. Marine Policy, ??(??):Article 103423, ???? 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X1830722X.

Jayasinghe:2024:OPG

[JAH^{+}24] Rohantha Rukshan Jayasinghe, Wasudha Prabodhani Abeyrathna, Manuja Promodya Hendawitharana, Champika Liyanage, Karl Williams, and Rangika Umesh Halwatura. Overcoming policy gaps for effective plastic waste management: a case study of Dikkowita Fishery Harbour, Sri Lanka. Marine Policy, 161(??):Article 106028, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000265.

Jones:2023:ERI

[JAK^{+}23] Stuart T. Jones, Edward H. Allison, Kailin Kroetz, Yoshitaka Ota, and Sunny L. Jardine. Enrollment, retention, and inclusivity of Marine Stewardship Council (MSC) eco-labelling certifications. Marine Policy, 155(??):Article 105734, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002671.

Jardim:2023:MFS

[JC23] Ernesto Jardim and Rohan J. C. Currey. The MSC Fisheries Standard Review policy development process. Marine Policy, 157(??):Article 105855, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003883.

Jiang:2020:GMS

[JCMD20] Daokui Jiang, Zhuo Chen, Levi McNeil, and Guilin Dai. The game mechanism of stakeholders in comprehensive marine environmental governance. Marine Policy, 112(??):Article 103728, February 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309485.

Jo:2020:DTS

[JD20] Sohyun Jo and Enrico D’agostini. Disrupting technologies in the shipping industry: How will MASS development affect the maritime workforce in Korea. Marine Policy, 120(??):Article 104139, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20304668.

Jacob:2023:TWP

[JDF^{+}23] Céline Jacob, Sereno DuPrey Diederichsen, Liam Fullbrook, Amanda T. Lombard, Siân E. Rees, Nina Rivers, Bernadette Snow, Mia Strand, Rachel Zuercher, and Holly J. Niner. A two way process — social capacity as a driver and outcome of equitable marine spatial planning. Marine Policy, 149(??):Article 105507, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000349.

Jensen:2021:FDC

[Jen21] Rikke Bjerg Jensen. Fragmented digital connectivity and security at sea. Marine Policy, 130(??):Article 104289, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309350.

Jensen:2023:ADI

[Jen23] Frank Jensen. Advantages and disadvantages of introducing strong user rights in fisheries. A literature review. Marine Policy, 157(??):Article 105839, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300372X.

Jiang:2020:RSC

[JF20] Minzhen Jiang and Michael Faure. Risk-sharing in the context of fishery mutual insurance: Learning from China. Marine Policy, 121(??):Article 104191, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030405X.

Jiang:2022:CSM

[JF22] Minzhen Jiang and Michael Faure. The compensation system for marine ecological damage resulting from offshore drilling in China. Marine Policy, 143(??):Article 105132, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001798.

Jin:2024:GAP

[JF24] Yongming Jin and Yuan Feng. Asia–Pacific maritime security cooperation: Challenges and opportunities. Marine Policy, 160(??):Article 105942, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300475X.

Jacquet:2024:OAI

[JFB^{+}24] Jennifer Jacquet, Becca Franks, Heather Browning, Alfonso Donoso, Matthew Hayek, Kathy Hessler, and Laurie Sellars. Away from an offshore aquaculture industry in the United States. Marine Policy, 160(??):Article 105986, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23005195.

Jamieson:2020:DAI

[JG20] J. W. Jamieson and A. Gartman. Defining active, inactive, and extinct seafloor massive sulfide deposits. Marine Policy, 117(??):Article 103926, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307845.

Jin:2023:PTD

[JG23] Yan Jin and José María Gil. Price transmission during food safety incidents: the case of Spanish tuna fraud. Marine Policy, 156(??):Article 105794, October 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003275.

James:2023:IME

[JGWM^{+}23] M. A. James, R. Gozzer-Wuest, T. Mendo, I. Gomez, J. Grillo-Núñez, and J. Mendo. To ignore or mitigate — economic implications of an illegal artisanal trawl fishery in northern Peru. Marine Policy, 158(??):Article 105865, December 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003986.

Jeffery:2023:OSR

[JHHG23] Sharon Jeffery, Lucie C. Hannah, Leif-Matthias Herborg, and Candice St Germain. Oil spill response planning in Pacific Canada: a tool for identifying vulnerable marine biota. Marine Policy, 148(??):Article 105466, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22005176.

Jiang:2021:PCS

[JHL21] Miaomiao Jiang, Maixiu Hu, and Markus Leibrecht. Profitability of container shipping via the Arctic Northeast Passage: a simulation and regression analysis. Marine Policy, 133(??):Article 104738, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003493.

Juan:2022:FKS

[JK22] Wang Juan and Wang Kexin. Fishery knowledge spillover effects on tourism economic growth in China — spatiotemporal effects and regional heterogeneity. Marine Policy, 139(??):Article 105019, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000665.

Jueseah:2020:SFE

[JKKT20] Alvin Slewion Jueseah, Ogmundur Knutsson, Dadi Mar Kristofersson, and Tumi Tómasson. Seasonal flows of economic benefits in small-scale fisheries in Liberia: a value chain analysis. Marine Policy, 119(??):Article 104042, September 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19305330.

Jayanthi:2022:ILW

[JKV^{+}22] M. Jayanthi, M. Kumaran, S. Vijayakumar, M. Duraisamy, P. R. Anand, M. Samynathan, S. Thirumurthy, S. Kabiraj, K. P. K. Vasagam, A. Panigrahi, and M. Muralidhar. Integration of land and water resources, environmental characteristics, and aquaculture policy regulations into site selection using GIS based spatial decision support system. Marine Policy, 136(??):Article 104946, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005571.

Ji:2021:DCF

[JL21a] Jianyue Ji and Yanming Li. The development of China’s fishery informatization and its impact on fishery economic efficiency. Marine Policy, 133(??):Article 104711, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003225.

Jiang:2021:DPP

[JL21b] Shan-Shan Jiang and Jing-Mei Li. Do political promotion incentive and fiscal incentive of local governments matter for the marine environmental pollution? Evidence from China’s coastal areas. Marine Policy, 128(??):Article 104505, June 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001160.

Jones:2021:ADR

[JL21c] Peter J. S. Jones and Stephen D. Long. Analysis and discussion of 28 recent marine protected area governance (MPAG) case studies: Challenges of decentralisation in the shadow of hierarchy. Marine Policy, 127(??):Article 104362, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20310137.

Judd:2021:AST

[JL21d] Adrian Judd and Jemma-Anne Lonsdale. Applying systems thinking: the Ecosystem Approach and Natural Capital Approach — Convergent or divergent concepts in marine management? Marine Policy, 129(??):Article 104517, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001287.

Ji:2023:ERD

[JLY23] Jianyue Ji, Huimin Liu, and Xingmin Yin. Evaluation and regional differences analysis of the marine industry development level: the case of China. Marine Policy, 148(??):Article 105445, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004924.

Johnson:2020:MSP

[JMB^{+}20] Ayana Elizabeth Johnson, William James McClintock, Ogden Burton, Wayde Burton, Andrew Estep, Kathryn Mengerink, Read Porter, and Stephanie Tate. Marine spatial planning in Barbuda: a social, ecological, geographic, and legal case study. Marine Policy, 113(??):Article 103793, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18309928.

Jones:2021:GAN

[Jon21] Peter J. S. Jones. A governance analysis of Ningaloo and Shark Bay Marine Parks, Western Australia: Putting the ‘eco’ in tourism to build resilience but threatened in longterm by climate change? Marine Policy, 127(??):Article 103636, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19303458.

Jompa:2023:TMP

[JPM^{+}23] Jamaluddin Jompa, Amanda Pricella Putri, Abigail Mary Moore, Hartati Tamti, and Siti Haerani. The transference of marine protected area management authority in Indonesia: Problems encountered, consequences and ways to move forward. Marine Policy, 155(??):Article 105756, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002890.

Jaleel:2023:PFM

[JS23] Azmath Jaleel and Hance D. Smith. The plight of the fishers and the management of IUU fishing. Marine Policy, 150(??):Article 105557, April 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000842.

Jaya:2022:WPF

[JSW^{+}22a] Indra Jaya, Fayakun Satria, Wudianto, Duto Nugroho, Lilis Sadiyah, Eny A. Buchary, Alan T. White, Erik C. Franklin, Catherine A. Courtney, Gina Green, and Stuart J. Green. “Are the working principles of fisheries management at work in Indonesia?”. Marine Policy, 140(??):Article 105047, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2200094X.

Jubinville:2022:PPA

[JSW22b] Isabelle Jubinville, Nancy L. Shackell, and Boris Worm. From policy to practice: Addressing bycatch for marine species-at-risk in Canada. Marine Policy, 146 (??):Article 105300, December 2022. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003475.

Jara:2020:CFS

[JTR^{+}20] Hans J. Jara, Jorge Tam, Borja G. Reguero, Francisco Ganoza, Gladis Castillo, Carlos Y. Romero, Manon Gévaudan, and Américo A. Sánchez. Current and future socio-ecological vulnerability and adaptation of artisanal fisheries communities in Peru, the case of the Huaura province. Marine Policy, 119(??):Article 104003, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19305597.

Johansen:2020:CSO

[JV20] Despina F. Johansen and Rolf A. Vestvik. The cost of saving our ocean — estimating the funding gap of sustainable development goal 14. Marine Policy, 112(??):Article 103783, February 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19305111.

Jiang:2020:EGM

[JWX^{+}20] Bao Jiang, Xiaoqiong Wang, Hailiang Xue, Jian Li, and Yu Gong. An evolutionary game model analysis on emission control areas in China. Marine Policy, 118(??):Article 104010, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20301408.

Jang:2020:CLA

[JY20] Ho Geun Jang and Satoshi Yamazaki. Communitylevel analysis of correlated fish production in fisheries and aquaculture: the case of Japan. Marine Policy, 122(??):Article 104240, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20308861.

Jiang:2021:SIF

[JZW21] Ke Jiang, Xinxin Zhang, and Yusheng Wang. Stability and influencing factors when designing incentive-compatible payments for watershed services: Insights from the Xin’an River Basin, China. Marine Policy, 134(??):Article 104824, December 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004358.

Kadfak:2021:SNM

[KA21] Alin Kadfak and Anna Antonova. Sustainable Networks: Modes of governance in the EU’s external fisheries policy relations under the IUU Regulation in Thailand and the SFPA with Senegal. Marine Policy, 132(??):Article 104656, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002670.

Kurt:2021:OCA

[KABT21] Ismail Kurt, Murat Aymelek, Evangelos Boulougouris, and Osman Turan. Operational cost analysis for a container shipping network integrated with offshore container port system: a case study on the West Coast of North America. Marine Policy, 126(??):Article 104400, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000105.

King:2022:MRR

[KAC22] Chloe King, Dedi S. Adhuri, and Julian Clifton. Marine reserves and resilience in the era of COVID-19. Marine Policy, 141(??):Article 105093, July 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001403.

Kai:2021:CIC

[Kai21] Mikihiko Kai. Are the current IUCN category and CITES listing appropriate for the conservation and management of shortfin mako, Isurus oxyrinchus, in the North Pacific Ocean? Marine Policy, 134(??):Article 104790, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004012.

Kakee:2020:DSM

[Kak20] Tomoko Kakee. Deep-sea mining legislation in Pacific Island countries: From the perspective of public participation in approval procedures. Marine Policy, 117(??):Article 103881, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19305718.

Kuemlangan:2023:EAA

[KAN^{+}23] Blaise Kuemlangan, Elizabeth-Rose Amidjogbe, Julia Nakamura, Alessandra Tomassi, Rudolph Hupperts, Buba Bojang, and Teresa Amador. Enforcement approaches against illegal fishing in national fisheries legislation. Marine Policy, 149(??):Article 105514, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000416.

Kao:2023:ATF

[Kao23] Shih-Ming Kao. Analysis of Taiwan’s foreign fisheries policy in tuna RFMOs: Possible variables and determinants. Marine Policy, 151(??):Article 105598, May 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001252.

Kara:2022:DMS

[Kar22a] Esma Gül Emecen Kara. Determination of maritime safety performance of flag states based on the Port State Control inspections using TOPSIS. Marine Policy, 143(??):Article 105156, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002032.

Karim:2022:MCI

[Kar22b] Md Saiful Karim. Maritime cybersecurity and the IMO legal instruments: Sluggish response to an escalating threat? Marine Policy, 143(??):Article 105138, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001853.

Karnad:2022:ILE

[Kar22c] Divya Karnad. Incorporating local ecological knowledge aids participatory mapping for marine conservation and customary fishing management. Marine Policy, 135(??):Article 104841, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004528.

Karim:2023:IOT

[Kar23] Md Saiful Karim. Indian Ocean Tuna Commission Climate Change Resolution: a quiet interaction of ocean and climate change legal regimes. Marine Policy, 148(??):Article 105434, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200481X.

Kassouri:2021:EDF

[Kas21] Yacouba Kassouri. Exploring the dynamics of fishing footprints in the Gulf of Guinea and Congo Basin region: Current status and future perspectives. Marine Policy, 133(??):Article 104739, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100350X.

Katikiro:2020:RLI

[Kat20] Robert Eliakim Katikiro. The role of local institutions in governance of MPAs. Marine Policy, 117(??):Article 103906, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19303823.

Kawamoto:2022:CEG

[Kaw22] Taro Kawamoto. A challenge to estimate global canned tuna demand and its impact on future tuna resource management using the gamma model. Marine Policy, 139(??):Article 105016, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200063X.

Kawamoto:2020:CFP

[KB20] Taro Kawamoto and Osamu Baba. Comparison of financial performance of Japanese and Australian small scale tuna longline fisheries. Marine Policy, 115(??):Article 103824, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1930733X.

Krause:2020:VSA

[KBD^{+}20] Gesche Krause, Suzannah-Lynn Billing, John Dennis, Jon Grant, Lucia Fanning, Ramón Filgueira, Molly Miller, José Antonio Pérez Agúndez, Nardine Stybel, Selina M. Stead, and Wojciech Wawrzynski. Visualizing the social in aquaculture: How social dimension components illustrate the effects of aquaculture across geographic scales. Marine Policy, 118(??):Article 103985, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307481.

Kambey:2021:EBP

[KCCC^{+}21] Cicilia S. B. Kambey, Iona Campbell, Elizabeth J. Cottier-Cook, Adibi R. M. Nor, Azhar Kassim, Ahemad Sade, and Phaik-Eem Lim. Evaluating biosecurity policy implementation in the seaweed aquaculture industry of Malaysia, using the quantitative knowledge, attitude, and practices (KAP) survey technique. Marine Policy, 134(??):Article 104800, December 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004115.

Kochanowicz:2021:UWS

[KDH^{+}21] Zuzanna Kochanowicz, Jackie Dawson, William D. Halliday, Michael Sawada, Luke Copland, Natalie Ann Carter, Adrian Nicoll, Steven H. Ferguson, Mads Peter Heide-Jørgensen, Marianne Marcoux, Cortney Watt, and David J. Yurkowski. Using western science and Inuit knowledge to model ship-source noise exposure for cetaceans (marine mammals) in Tallurutiup Imanga (Lancaster Sound), Nunavut, Canada. Marine Policy, 130(??):Article 104557, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001688.

Kline:2020:SSU

[KDM^{+}20] Logan R. Kline, Annamaria I. DeAngelis, Candace McBride, Giverny G. Rodgers, Timothy J. Rowell, Jeremy Smith, Jenni A. Stanley, Andrew D. Read, and Sofie M. Van Parijs. Sleuthing with sound: Understanding vessel activity in marine protected areas using passive acoustic monitoring. Marine Policy, 120(??):Article 104138, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309005.

Kohler:2022:TSP

[KDS22] Jonathan Köhler, Ewa Dönitz, and Frank Schätter. Transitions for ship propulsion to 2050: the AHOY combined qualitative and quantitative scenarios. Marine Policy, 140(??):Article 105049, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000963.

Koomson:2020:CVF

[KDVR20] Daniel Koomson, Katherine Sian Davies-Vollum, and Debadayita Raha. Characterising the vulnerability of fishing households to climate and environmental change: Insights from Ghana. Marine Policy, 120(??):Article 104142, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19308590.

Kenchington:2020:GPD

[Ken20] Trevor J. Kenchington. La Grande Pêche: Drivers of change in the saltfish fisheries on Grand Bank through four centuries. Marine Policy, 115(??):Article 103868, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19300132.

Knapp:2023:EAE

[KF23] Sabine Knapp and Philip Hans Franses. Empirical analysis of the effectiveness of the legislative framework in the maritime industry. Marine Policy, 147(??):Article 105401, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004481.

Karcher:2020:TSP

[KFB^{+}20] Denis B. Karcher, Elodie Fache, Annette Breckwoldt, Hugh Govan, Xochitl E. Elías Ilosvay, Juliette Kon Kam King, Léa Riera, and Catherine Sabinot. Trends in South Pacific fisheries management. Marine Policy, 118(??):Article 104021, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19306499.

Kelly:2020:SLM

[KFM^{+}20] Rachel Kelly, Aysha Fleming, Mary Mackay, Carolina García, and Gretta T. Pecl. Social licence for marine protected areas. Marine Policy, 115(??):Article 103782, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19305196.

Kemp:2020:CPO

[KFP20] Paul S. Kemp, Rainer Froese, and Daniel Pauly. COVID-19 provides an opportunity to advance a sustainable UK fisheries policy in a post-Brexit brave new world. Marine Policy, 120(??):Article 104114, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301032.

Kelly:2022:FFO

[KFS^{+}22] Rachel Kelly, Paul Foley, Robert L. Stephenson, Alistair J. Hobday, Gretta T. Pecl, Fabio Boschetti, Christopher Cvitanovic, Aysha Fleming, E. A. Fulton, Kirsty L. Nash, Barbara Neis, Gerald G. Singh, and E. Ingrid van Putten. Foresighting future oceans: Considerations and opportunities. Marine Policy, 140(??):Article 105021, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000689.

Khan:2024:IFT

[KgJqY^{+}24] Alexander M. A. Khan, Ming guo Jiang, Xiao qiang Yang, Izza Mahdiana Apriliani, Noir Primadona Purba, Budy Wiryawan, Am Azbas Taurusman, and Buntora Pasaribu. Illegal fishing threatens the sustainability of future tuna commodities in Indonesia. Marine Policy, 159(??):Article 105936, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004694.

Koehler:2022:APE

[KGL22] Lydia Koehler, Ioannis Giovos, and Jason Lowther. The application of precaution in elasmobranch conservation and management in the Mediterranean Sea. Marine Policy, 135(??):Article 104830, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004413.

Kraan:2020:SSP

[KGP^{+}20] Marloes Kraan, Rolf Groeneveld, Annet Pauwelussen, Tim Haasnoot, and Simon R. Bush. Science, subsidies and the politics of the pulse trawl ban in the European Union. Marine Policy, 118(??):Article 103975, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309881.

Ko:2022:CSL

[KGP^{+}22] Dong-Hui Ko, Yunzheng Ge, Jin-Soon Park, Lei Liu, Changlei Ma, Fengyun Chen, Jingping Peng, Seung-Koo Kang, Jin-Hak Yi, and Weimin Liu. A comparative study of laws and policies on supporting marine energy development in China and Korea. Marine Policy, 141(??):Article 105057, July 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2200104X.

Kirkham:2020:DSM

[KGW20] Nicholas R. Kirkham, Kristina M. Gjerde, and A. Meriwether W. Wilson. DEEP-SEA mining: Policy options to preserve the last frontier — Lessons from Antarctica’s mineral resource convention. Marine Policy, 115(??):Article 103859, May 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19305615.

Kuhnen:2022:CLM

[KHdSM^{+}22] Vanessa V. Kuhnen, Kevin Hopkins, Lucas dos Santos Mota, Otavio M. Sousa, and Eduardo G. Sanches. Challenges and lessons from marine finfish farming in Brazil. Marine Policy, 138(??):Article 104979, April 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000264.

Kim:2023:VAK

[KHH^{+}23] Moo-Jin Kim, Jae-Beum Hong, In-Seong Han, Joon-Soo Lee, and Do-Hoon Kim. Vulnerability assessment of Korean fisheries to climate change. Marine Policy, 155(??):Article 105735, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002683.

Kvamsdal:2023:EOT

[KHLS23] Sturla Kvamsdal, Arnt O. Hopland, Yuanhao Li, and Simon Selle. Expert opinions on threats and impacts in the marine environment. Marine Policy, 147(??):Article 105382, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004298.

Khan:2020:HCC

[KHMM20] Wajiha Khan, Upamanyu Hore, Soham Mukherjee, and Gowri Mallapur. Human-crocodile conflict and attitude of local communities toward crocodile conservation in Bhitarkanika Wildlife Sanctuary, Odisha, India. Marine Policy, 121(??):Article 104135, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19300739.

Kim:2020:WDR

[Kim20] Hyun Jung Kim. What did the Republic of Korea and Japan mean by the term “joint development” in their 1974 agreement? Marine Policy, 117(??):Article 103932, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20300154.

Kim:2021:TPF

[Kim21] Hyun Jung Kim. Third party fishing in an undelimited area under the Republic of Korea–Japan Fisheries Agreement. Marine Policy, 125(??):Article 104402, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000129.

Kim:2023:ACL

[Kim23] Sojeong Kim. Adaptive capacity and local-level fisheries co-management activities: a case of South Korea. Marine Policy, 154(??):Article 105665, August 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001926.

Kizielewicz:2024:ECC

[Kiz24] Joanna Kizielewicz. Economic consequences of the COVID’19 pandemic for functioning of the cruise shipping industry in the Baltic Sea Region. Marine Policy, 159(??):Article 105925, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2300458X.

Khuu:2021:GAD

[KJE21a] Duong T. Khuu, Peter J. S. Jones, and Paul Ekins. A Governance analysis of Con Dao National Park, Vietnam. Marine Policy, 127(??):Article 103986, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305421.

Khuu:2021:GAN

[KJE21b] Duong T. Khuu, Peter J. S. Jones, and Paul Ekins. Governance analysis of Nha Trang Bay and Cu Lao Cham Marine Protected Areas, Vietnam. Marine Policy, 127(??):Article 104330, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20309775.

Kushardanto:2022:HFT

[KJS^{+}22] Hari Kushardanto, Raymond Jakub, Wahid Suherfian, Tarlan Subarno, Ahmad Isa Ansyori, La Sara, Naslina Alimina, Fajriah, La Ode Kardini, Emilio de la Rosa, Ade Yuliani, Eva Medianti, Imanda Pradana, Haris Setiawan, Yoni Muhammad, Lely Fajriah Djafar, Stephen Box, Courtney Cox, and Stuart J. Campbell. Household finances and trust are key determinants of benefits from small-scale fisheries co-management. Marine Policy, 145(??):Article 105284, November 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003311.

Kerber:2022:LFA

[KK22] Heide Kerber and Johanna Kramm. From laissez-faire to action? Exploring perceptions of plastic pollution and impetus for action. Insights from Phu Quoc Island. Marine Policy, 137(??):Article 104924, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005352.

Kaiser:2021:GEE

[KKA^{+}21] Brooks A. Kaiser, Melina Kourantidou, Dewan Ahsan, Sergey Bakanev, AnnDorte Burmeister, Ginny Eckert, Linda M. Fernandez, Hyun Pyo Hong, Albert Alexandre Monsalve, Darrell Mullowney, Bjørg Helen Nøstvold, Hojeong Park, Edward Poulsen, Lars Ravn-Jonsen, Chris Siddon, Jan H. Sundet, Kanae Tokunaga, and Masashi Yamamoto. Global ecological and economic connections in Arctic and sub-Arctic crab markets. Marine Policy, 127(??):Article 104442, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2100052X.

Kaliszewski:2020:KFC

[KKDK20] Adam Kaliszewski, Arkadiusz Kozłowski, Janusz Dabrowski, and Hanna Klimek. Key factors of container port competitiveness: a global shipping lines perspective. Marine Policy, 117(??):Article 103896, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19307146.

Knutsen:2022:LRM

[KKO^{+}22] Jan Atle Knutsen, Alf Ring Kleiven, Esben Moland Olsen, Halvor Knutsen, Sigurd Heiberg Espeland, Tonje Knutsen Sørdalen, Susanna Huneide Thorbjørnsen, Jeffrey A. Hutchings, Albert Fernández-Chacón, Mats Huserbråten, David Villegas-Ríos, Kim Tallaksen Halvorsen, Portia Joy Nillos Kleiven, Thomas Kiland Langeland, and Even Moland. Lobster reserves as a management tool in coastal waters: Two decades of experience in Norway. Marine Policy, 136(??):Article 104908, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005194.

Kadfak:2021:MTJ

[KL21a] Alin Kadfak and Sebastian Linke. More than just a carding system: Labour implications of the EU’s illegal, unreported and unregulated (IUU) fishing policy in Thailand. Marine Policy, 127(??):Article 104445, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000555.

Konrad:2021:CBA

[KL21b] Leilani Konrad and Arielle Levine. Controversy over beach access restrictions at an urban coastal seal rookery: Exploring the drivers of conflict escalation and endurance at Children’s Pool Beach in La Jolla, CA. Marine Policy, 132(??):Article 104659, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002700.

Koehler:2022:PMS

[KL22] Lydia Koehler and Jason Lowther. Policy making for sharks and the role and contribution of non-governmental organisations in the fulfilment of international legal obligations. Marine Policy, 144(??):Article 105228, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002755.

Klein:2021:MAV

[Kle21] Natalie Klein. Maritime autonomous vehicles and international laws on boat migration: Lessons from the use of drones in the Mediterranean. Marine Policy, 127(??):Article 104447, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21000579.

Kim:2023:PAC

[KLK23] Boram Kim, Hyejin Lee, and Taeil Kim. Possibility analysis of complex marine accidents caused by large marine plastic litter in the Republic of Korea. Marine Policy, 152(??):Article 105591, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001185.

Kim:2022:HCM

[KLS22] A-Rom Kim, Sung-Woo Lee, and Young-Joon Seo. How to control and manage vessels’ ballast water: the perspective of Korean shipping companies. Marine Policy, 138(??):Article 105007, April 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000549.

Kinds:2021:CLA

[KLSG21] Arne Kinds, Pascal Le Floc’h, Stijn Speelman, and Olivier Guyader. Challenging the ‘artisanal vs. industrial’ dichotomy in French Atlantic fisheries: an organizational typology of multi-vessel fishing firms. Marine Policy, 134(??):Article 104753, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100364X.

Khan:2020:RDT

[KMG^{+}20] Alexander M. A. Khan, Aileen C. Mill, Tim S. Gray, Mingguo Jiang, Hazmi Arief, Arthur Brown, Amirul Karman, and Nicholas V. C. Polunin. Reliability of the data on tuna catches obtained from the dockside in Indonesia: a study of stakeholders’ perceptions. Marine Policy, 122(??):Article 104242, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308885.

Koerner:2024:TSF

[KMJ24] Lauren M. Koerner, Michael C. Melnychuk, and Ernesto Jardim. Testing the sensitivity of fishery assessment scores to changes in the MSC Fisheries Standard requirements. Marine Policy, 161(??):Article 106005, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000034.

Kokosalakis:2021:ERE

[KMM21] George Kokosalakis, Anna Merika, and Xakousti-Afroditi Merika. Environmental regulation on the energy-intensive container ship sector: a restraint or opportunity? Marine Policy, 125(??):Article 104278, March 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309246.

Kalidin:2020:BEM

[KMS20] Bhavnah Komul Kalidin, Carlo Mattone, and Marcus Sheaves. Barriers to effective monitoring and evaluation of small-scale fisheries in small island developing states: an example from Mauritius. Marine Policy, 118(??):Article 103845, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X1930541X.

Kull:2021:IGP

[KMT^{+}21] M. Kull, J. R. Moodie, H. L. Thomas, S. Mendez-Roldan, A. Giacometti, A. Morf, and I. Isaksson. International good practices for facilitating transboundary collaboration in Marine Spatial Planning. Marine Policy, 132(??):Article 103492, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304330.

Komatsu:2021:EMP

[KN21] Masayuki Komatsu and Tomoko Nakamura. Ecosystem management for protection of Japanese ocean and fishery resources. Marine Policy, 132(??):Article 104682, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002931.

Kim:2020:PAL

[KNY20] Ju-Hee Kim, Jungho Nam, and Seung-Hoon Yoo. Public acceptance of a large-scale offshore wind power project in South Korea. Marine Policy, 120(??):Article 104141, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20302499.

Kim:2022:PPB

[KNY22] Ju-Hee Kim, Jungho Nam, and Seung-Hoon Yoo. Public perceptions of blue carbon in South Korea: Findings from a choice experiment. Marine Policy, 144(??):Article 105236, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002834.

Kobayashi:2022:CIC

[Kob22] Masanori Kobayashi. The COVID-19 impacts and challenges to achieving sustainability in Japan’s fisheries and aquaculture. Marine Policy, 143(??):Article 105161, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002081.

Karahalil:2021:EPC

[KOBS21] Meric Karahalil, Burcu Ozsoy, Ersan Basar, and Tanzer Satır. The evaluation of the Polar Code by the survey conducted with those who have sailed in polar regions, and suggestions for further improvement. Marine Policy, 128(??):Article 104502, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001135.

Kuo:2023:CCD

[KOCL23] Ching-Ching Kuo, Ching-Hsiewn Ou, Wen-Yan Chiau, and Chao-Shing Lee. Challenges and conflicts in the designation of Taiwan’s underwater cultural heritage protected areas. Marine Policy, 153(??):Article 105656, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001835.

Koehler:2022:CIT

[Koe22] Holly Koehler. The case for industry transparency in supporting sustainable tuna fisheries. Marine Policy, 136(??):Article 104087, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19306384.

Kojima:2023:IGP

[Koj23] Chie Kojima. Integration of general principles of international environmental law into the Law of the Sea: Assessment and challenges. Marine Policy, 149(??):Article 105497, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X23000246.

Kontovas:2020:IAQ

[Kon20] Christos A. Kontovas. Integration of air quality and climate change policies in shipping: the case of sulphur emissions regulation. Marine Policy, 113(??):Article 103815, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19304683.

Kim:2024:USR

[KP24] Jong-Kwan Kim and Deuk-Jin Park. Understanding of sailing rule based on COLREGs: Comparison of navigator survey and automated collision–avoidance algorithm. Marine Policy, 159(??):Article 105894, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300427X.

Kim:2020:DCN

[KPN20] Tae-Hyun Kim, Cheol-Hyung Park, and Jongoh Nam. The determinants of changes in the number of fishers employed by fisheries household in the Republic of Korea using count data models. Marine Policy, 117(??):Article 103974, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1930569X.

Kraabel:2022:IAB

[Kra22] Kristine Dalaker Kraabel. Institutional arrangements in a BBNJ treaty: Implications for Arctic marine science. Marine Policy, 142(??):Article 103807, August 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305081.

Kaminski:2022:CSO

[KS22] Tomasz Kamiński and Rafał Szewczyk. The coastal state obligation not to impede the laying or maintenance of submarine pipelines on the continental shelf according to United Nations convention on the law of the sea. Marine Policy, 143(??):Article 105086, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001336.

Kaeriyama:2023:EBS

[KS23] Masahide Kaeriyama and Isao Sakaguchi. Ecosystem-based sustainable management of chum salmon in Japan’s warming climate. Marine Policy, 157(??):Article 105842, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003755.

Kemp:2023:FAU

[KSB^{+}23a] Paul S. Kemp, Gowshika Subbiah, Richard Barnes, Kristina Boerder, Bethan C. O’Leary, Bryce D. Stewart, and Chris Williams. Future advances in UK marine fisheries policy: Integrated nexus management, technological advance, and shifting public opinion. Marine Policy, 147(??):Article 105335, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003827.

Kemp:2023:FMF

[KSB^{+}23b] Paul S. Kemp, Gowshika Subbiah, Richard Barnes, Kristina Boerder, Bethan C. O’Leary, Bryce D. Stewart, and Chris Williams. The future of marine fisheries management and conservation in the United Kingdom: Lessons learnt from over 100 years of biased policy. Marine Policy, 147(??):Article 105075, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001221.

Kim:2023:KBA

[KSS23] A-Rom Kim, Jungyong Seo, and Young-Joon Seo. Key barriers to adopting onshore power supply to reduce port air pollution: Policy implications for the maritime industry in South Korea. Marine Policy, 157(??):Article 105866, November 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003998.

Kumawat:2023:JFA

[KSV^{+}23] Tarachand Kumawat, Raju Saravanan, Kavungal Vinod, Ashok Kumar Jaiswar, Geetanjali Deshmukhe, Latha Shenoy, Damodaran Divu, Kunnummel Krishnan Joshi, and Achamveetil Gopalakrishnan. Jellyfish fisheries along the Gujarat coast, India: Status and challenges. Marine Policy, 150(??):Article 105554, April 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000817.

Krueck:2022:BMP

[KTC^{+}22] Nils C. Krueck, Colm Tong, Courtney Cox, Eric A. Treml, Kay Critchell, Iliana Chollett, Dedi S. Adhuri, Maria Beger, Dominic Muenzel, Daniel Holstein, Stuart J. Campbell, and Raymond Jakub. Benefits of measurable population connectivity metrics for area-based marine management. Marine Policy, 144(??):Article 105210, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002573.

Komi:2024:IRF

[KTI^{+}24] Gentle Wilson Komi, Nelson Turyahabwe, Prossy Isubikalu, Olufemi Martins Adesope, Stephen Okodudu, and Nkopuyo Udoekpo Sunday. Influence of reciprocal financial resources on shrimp fishers’ income sources in Southern Nigeria. Marine Policy, 162(??):Article 106050, April 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000484.

Kulaga:2022:PMS

[Kuł22] Łukasz Kułaga. Prohibition of massive and serious pollution of marine environment as a jus cogens. Identification, legal bases and consequences in view of the recent work of the International Law Commission. Marine Policy, 144(??):Article 105217, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002640.

Kularatne:2023:OMN

[Kul23] Ranil Kavindra Asela Kularatne. Occurrence of marine nonindigenous (NIS) species: Current status, management approaches and challenges in Sri Lanka. Marine Policy, 149(??):Article 105477, March 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000040.

Kalinin:2020:SAR

[KV20] Nikita Kalinin and Mark Vershinin. Strategic analysis of the Russian crab quota auction in 2019. Marine Policy, 122(??):Article 104266, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030912X.

Kuismanen:2023:IEV

[KVL^{+}23] Lauri M. J. Kuismanen, Elina A. Virtanen, Juho Lappalainen, Lasse Kurvinen, Penina Blankett, and Markku Viitasalo. Identifying ecologically valuable marine areas to support conservation and spatial planning at scales relevant for decision making. Marine Policy, 158(??):Article 105890, December 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004232.

Kirkfeldt:2020:OAN

[KvTNL20] Trine Skovgaard Kirkfeldt, Jan P. M. van Tatenhove, Helle Nedergaard Nielsen, and Sanne Vammen Larsen. An ocean of ambiguity in Northern European marine spatial planning policy designs. Marine Policy, 119 (??):Article 104063, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20303213.

Kadagi:2021:COS

[KWF^{+}21] Nelly Isigi Kadagi, Nina Wambiji, Sean Thomas Fennessy, Micheal S. Allen, and Robert N. M. Ahrens. Challenges and opportunities for sustainable development and management of marine recreational and sport fisheries in the Western Indian Ocean. Marine Policy, 124(??):Article 104351, February 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20310010.

Kao:2023:CDP

[KWG23] Shih-Ming Kao, Wei-Ning Wu, and Meng-En Gu. Coastal disaster protection and hazard perceptions of the residents of Orchid Island, Taiwan. Marine Policy, 148 (??):Article 105459, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22005061.

Kimani:2020:APG

[KWMM20a] Patrick Kimani, Andrew Wamukota, Julius O. Manyala, and Chrisestom Mwatete Mlewa. Actors’ perceptions of government performance in support of value chain development in marine small-scale fisheries in Kenya. Marine Policy, 122(??):Article 104221, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308678.

Kimani:2020:FIF

[KWMM20b] Patrick Kimani, Andrew Wamukota, Julius O. Manyala, and Chrisestom Mwatete Mlewa. Factors influencing financial performance in marine small-scale fisheries value chain in Kenya. Marine Policy, 122(??):Article 104218, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20308642.

Kadagi:2020:PCR

[KWS20] N. I. Kadagi, N. Wambiji, and M. E. Swisher. Potential for conflicts in recreational and artisanal billfish fisheries on the coast of Kenya. Marine Policy, 117(??):Article 103960, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19307341.

Kuwae:2022:IBC

[KWY^{+}22] Tomohiro Kuwae, Atsushi Watanabe, Satoru Yoshihara, Fujiyo Suehiro, and Yoshihisa Sugimura. Implementation of blue carbon offset crediting for seagrass meadows, macroalgal beds, and macroalgae farming in Japan. Marine Policy, 138(??):Article 104996, April 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000434.

Kullerud:2020:AGR

[KY20] Lars Kullerud and Oran R. Young. Adding a Gakkel Ridge regime to the evolving Arctic Ocean governance complex. Marine Policy, 122(??):Article 104270, December 2020.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309167.

Kaifu:2021:MGE

[KYMW21] Kenzo Kaifu, Kazuki Yokouchi, Michael J. Miller, and Izumi Washitani. Management of glass eel fisheries is not a sufficient measure to recover a local Japanese eel population. Marine Policy, 134(??):Article 104806, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004176.

Kennard:2022:TTH

[KZR22] Alice Kennard, Pengfei Zhang, and Sriram Rajagopal. Technology and training: How will deck officers transition to operating autonomous and remote-controlled vessels? Marine Policy, 146(??):Article 105326, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003736.

Livingstone:2023:DFC

[LA23] Olivia Livingstone and Brandon P. Anthony. Documenting fisheries co-management in action: Case of the Grand Cape Mount County Co-Management Association, Liberia. Marine Policy, 157(??):Article 105844, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003779.

Lajeunesse:2021:CAO

[Laj21] Adam Lajeunesse. Canada’s Arctic Offshore and Patrol Ships (AOPS): Their history and purpose. Marine Policy, 124(??):Article 104323, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309702.

Lambach:2021:FTH

[Lam21] Daniel Lambach. The functional territorialization of the high seas. Marine Policy, 130(??):Article 104579, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001901.

LeBrenne:2021:SOC

[LBL21] Valérie Le Brenne, Laetitia Bisiaux, and Frédéric Le Manach. Sustainable objectives and commitments deceived by fisheries subsidies for ‘temporary cessations’ in times of COVID. Marine Policy, 132(??):Article 104670, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002815.

Lewin:2023:TEC

[LBWS23] Wolf-Christian Lewin, Fanny Barz, Marc Simon Weltersbach, and Harry V. Strehlow. Trends in a

European coastal fishery with a special focus on smallscale fishers — implications for fisheries policies and management. Marine Policy, 155(??):Article 105680, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002075.

Li:2020:LSE

[LC20a] Xintong Li and Yen-Chiang Chang. The legal status of the “employment salvage contract” in the Chinese maritime salvage law: a challenge to the principle of ‘no cure, no pay’. Marine Policy, 118(??):Article 104011, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20300014.

Liu:2020:ERM

[LC20b] Xiaofei Liu and Yen-Chiang Chang. An emergency responding mechanism for cruise epidemic prevention — taking COVID-19 as an example. Marine Policy, 119 (??):Article 104093, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20302451.

Liu:2022:ERF

[LC22] Xintian Liu and Suisui Chen. Has environmental regulation facilitated the green transformation of the marine industry? Marine Policy, 144(??):Article 105238, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002858.

Li:2023:SCC

[LC23] Xintong Li and Yen-Chiang Chang. A step closer to the convention on the Protection of Underwater Cultural Heritage 2001: China’s latest efforts in regulation. Marine Policy, 147(??):Article 105346, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003931.

Liu:2021:SCT

[LCAX21] Lijun Liu, Jingjie Chu, James L. Anderson, and Jintao Xu. Sustainability comparisons in the triple bottom line for Chinese fisheries. Marine Policy, 125(??):Article 104259, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309052.

Lee:2020:DDF

[LCC20] Tzu-Ming Lee, Pei-Yu Chi, and Kuo-I. Chang. Duration and determinants of fishery trade patterns: Evidence from OECD countries. Marine Policy, 118(??):Article 103977, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19309583.

LeBihan:2020:RVC

[LCL20] Véronique Le Bihan, Marie Catalo, and Jeanine Le Bihan. Reorganization of the value chain activities of oyster companies on the Atlantic coast following health crises in France (2006–2013). Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306868.

Luquin-Covarrubias:2022:LGE

[LCMBGP22] Marlene Anaid Luquin-Covarrubias, Enrique Morales-Bojórquez, and Sergio Scarry González-Peláez. The last geoduck: the experience of geoduck clam fishery management in the Mexican Pacific Ocean. Marine Policy, 143(??):Article 105145, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001920.

LeDuff:2020:OCD

[LDAC20] Matthieu Le Duff, Pascal Dumas, Michel Allenbach, and Olivier Cohen. An orientation for coastal disaster risks management and prevention policy in a global warming context: Case study in Ouvea (New Caledonia). Marine Policy, 117 (??):??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X16307084.

Li:2020:CSR

[LDC^{+}20] Feixue Li, Dongdong Ding, Zhenjie Chen, Honghua Chen, Ting Shen, Qianlian Wu, and Chenye Zhang. Change of sea reclamation and the sea-use management policy system in China. Marine Policy, 115(??):Article 103861, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302738.

LeBusque:2021:INM

[LDL21] Brianna Le Busque, Jillian Dorrian, and Carla Litchfield. The impact of news media portrayals of sharks on public perception of risk and support for shark conservation. Marine Policy, 124(??):Article 104341, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030991X.

Lin:2023:EVM

[LDL^{+}23] Chun-Yu Lin, Gui-Lin Dai, Yang Liu, Meng-Qi Zhang, Ying Liu, Wen Jiang, Xiu-Mei Fu, and Han-Xue Chen. Ecological value of mariculture shellfish resources in China: Assessment and management. Marine Policy, 148(??):Article 105406, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004535.

Lopez-Ercilla:2021:VMS

[LEERRM^{+}21] Ines Lopez-Ercilla, Maria Jose Espinosa-Romero, Francisco J. Fernandez Rivera-Melo, Stuart Fulton, Rebeca Fernández, Jorge Torre, Araceli Acevedo-Rosas, Arturo J. Hernández-Velasco, and Imelda Amador. The voice of Mexican small-scale fishers in times of COVID-19: Impacts, responses, and digital divide. Marine Policy, 131(??):Article 104606, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002177.

Levine:2020:SUA

[Lev20] Arielle Levine. Sunscreen use and awareness of chemical toxicity among beach goers in Hawaii prior to a ban on the sale of sunscreens containing ingredients found to be toxic to coral reef ecosystems. Marine Policy, 117(??):Article 103875, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19309029.

Luhtala:2021:BSI

[LEVET21] Hanna Luhtala, Anne Erkkilä-Välimäki, Søren Qvist Eliasen, and Harri Tolvanen. Business sector involvement in maritime spatial planning — Experiences from the Baltic Sea region. Marine Policy, 123(??):Article 104301, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309465.

Lout:2022:DWS

[LFLF22] Gabrielle Lout, Juno Fitzpatrick, Alejandro J. Garcia Lozano, and Elena Finkbeiner. Decent work in a seascape of livelihoods: Regional evaluation of the shrimp and groundfish fishery of the Guianas–Brazil Shelf. Marine Policy, 144(??):Article 105231, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002780.

Llorente:2020:AEP

[LFPBD^{+}20] Ignacio Llorente, José Fernández-Polanco, Elisa Baraibar-Diez, María D. Odriozola, Trond Bjørndal, Frank Asche, Jordi Guillen, Lamprakis Avdelas, Rasmus Nielsen, Maria Cozzolino, Manuel Luna, José L. Fernández-Sánchez, Ladislao Luna, Cristóbal Aguilera, and Bernardo Basurco. Assessment of the economic performance of the seabream and seabass aquaculture industry in the European Union. Marine Policy, 117(??):Article 103876, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307304.

Luna:2024:BBA

[LFVCF24] Manuel Luna, Simon Fernandez-Vazquez, Emilio Tereñes Castelao, and Álvaro Arias Fernández. A blockchain-based approach to the challenges of EU’s environmental policy compliance in aquaculture: From traceability to fraud prevention. Marine Policy, 159(??):Article 105892, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004256.

Lester:2024:RMA

[LGF24] Sarah E. Lester, Rebecca R. Gentry, and Halley E. Froehlich. The role of marine aquaculture in contributing to the diversity and stability of U.S. seafood production. Marine Policy, 160(??):Article 105994, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23005274.

Lomonico:2021:OFP

[LGW^{+}21] Serena Lomonico, Mary G. Gleason, Jono R. Wilson, Darcy Bradley, Kate Kauer, Richard J. Bell, and Thomas Dempsey. Opportunities for fishery partnerships to advance climate-ready fisheries science and management. Marine Policy, 123(??):Article 104252, January 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308988.

Liu:2023:ECC

[LHPJ23] Yang Liu, Limin Han, Zhaobin Pei, and Yiying Jiang. Evolution of the coupling coordination between the marine economy and urban resilience of major coastal cities in China. Marine Policy, 148(??):Article 105456, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22005036.

Larson:2023:SID

[LHTR23] Silva Larson, Sreynov Hoy, Somony Thay, and Michael A. Rimmer. Sustainable and inclusive development of finfish mariculture in Cambodia: Perceived barriers to engagement and expansion. Marine Policy, 148(??):Article 105439, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22004869.

Lear:2020:SER

[LHV^{+}20] D. Lear, P. Herman, G. Van Hoey, L. Schepers, N. Tonné, M. Lipizer, F. E. Muller-Karger, W. Appeltans, W. D. Kissling, N. Holdsworth, M. Edwards, E. Pecceu, H. Nygård, G. Canonico, S. Birchenough, G. Graham, K. Deneudt, S. Claus, and P. Oset. Supporting the essential — Recommendations for the development of accessible and interoperable marine biological data products. Marine Policy, 117(??):Article 103958, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19301915.

Leary:2021:GDP

[LHW21] David Leary, Nien-Tsu Alfred Hu, and Pei-Fu Wu. Gilt dragons, porcelain and plundered doubloons: a comparative analysis of Taiwan and Australian underwater cultural heritage protection laws. Marine Policy, 134(??):Article 104787, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003985.

Li:2023:UAD

[Li23a] Shuo Li. Unilateral actions in the development of the law of the sea. Marine Policy, 153(??):Article 105658, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001859.

Li:2023:FTA

[Li23b] Tao Li. Fisheries tenure arrangement in China: Legislative ambiguity, judicial settlement conflicts, and the gap in protecting traditional fishers’ rights. Marine Policy, 152(??):Article 105607, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23001343.

Lin:2020:PSW

[Lin20] Zhen Lin. The protection of sunken WWII warships located in Indonesian or Malaysian territorial waters. Marine Policy, 113(??):Article 103804, March 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303240.

Lindley:2023:FNC

[Lin23] Jade Lindley. Fishing non-compliance and culture. Marine Policy, 152(??):Article 105581, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001082.

Lin:2024:JLE

[Lin24] Zhen Lin. Joint law enforcement mechanism with regard to underwater cultural heritage at sea led by China Coast Guard: Legislation and practice. Marine Policy, 160(??):Article 105979, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23005122.

Liu:2020:RCC

[Liu20] Nengye Liu. The rise of China and conservation of marine living resources in the polar regions. Marine Policy, 121(??):Article 104181, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20303158.

Liu:2021:COH

[Liu21] Nengye Liu. China and one hundred years of the Svalbard Treaty: Past, present and future. Marine Policy, 124(??):Article 104354, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310046.

Liu:2022:USR

[Liu22] Wen-Hong Liu. University social responsibility to promote coastal partnership: Introducing Taiwan coastal communities. Marine Policy, 146(??):Article 105303, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003505.

Liu:2023:USI

[Liu23] Suyu Liu. Use of SDG Indicator 14.7.1 for ecosystembased fisheries management: Challenges and implications. Marine Policy, 151(??):Article 105560, May 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000878.

Li:2020:TMS

[LJ20] Shenghui Li and Stephen Jay. Transboundary marine spatial planning across Europe: Trends and priorities in nearly two decades of project work. Marine Policy, 118(??):Article 104012, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20301676.

Long:2021:GOG

[LJ21] Stephen Long and Peter J. S. Jones. Greenland’s offshore Greenland halibut fishery and role of the Marine Stewardship Council certification: a governance case study. Marine Policy, 127(??):Article 104095, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302736.

Lindfors:2022:SRI

[LJ22] Emil Tomson Lindfors and Stig-Erik Jakobsen. Sustainable regional industry development through co-evolution — the case of salmon farming and cell-based seafood production. Marine Policy, 135(??):Article 104855, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004668.

Li:2023:ATC

[LJ23] Shenghui Li and Stephen Jay. Addressing transboundary challenges: Exploring the interactive relations between collaborative governance and transboundary marine spatial planning in Europe. Marine Policy, 158(??):Article 105880, December 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2300413X.

Liu:2022:DIS

[LJL22] Chenxuan Liu, Qijun Jiang, and Markus Leibrecht. Does an increased synergy in an industry’s subsystems impact on industrial development? The case of the Fisheries industry in China. Marine Policy, 135(??):Article 104837, January 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004486.

Long:2021:GAC

[LJRHH21] Stephen Long, Peter J. S. Jones, Zoavina Randriana, and Jeneen Hadj-Hammou. Governance analysis of a community managed small-scale crab fishery in Madagascar: novel use of an empirical framework. Marine Policy, 127(??):Article 102974, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X17307522.

Lees:2023:IMC

[LKB^{+}23] Liisi Lees, Krista Karro, Francisco R. Barboza, Ann Ideon, Jonne Kotta, Triin Lepland, Maili Roio, and Robert Aps. Integrating maritime cultural heritage into maritime spatial planning in Estonia. Marine Policy, 147(??):Article 105337, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003840.

Lorena:2022:ICC

[LKE^{+}22] Romero Lorena, Yévenes Karina, Pokrant Ekaterina, Baumberger Cecilia, Zavala Sebastián, Burgos José Miguel, Wacyk Jurij, and Cornejo Javiera. Impacts of COVID-19 on the Chilean salmon: a first approach to the effects of the pandemic in the industry. Marine Policy, 144(??):Article 105185, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002329.

Lyu:2024:SRS

[LL24] Wenyu Lyu and Lianbo Li. Sailing into rough seas: the difficulties and challenges Chinese women seafarers face. Marine Policy, 160(??):Article 105955, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004888.

Lin:2021:CAI

[LLCH21] Ta-Li Lin, Wen-Hong Liu, Yi Chang, and Shih-Chun Hsiao. Capacity assessment of integrated coastal management for Taiwanese local government. Marine Policy, 134(??):Article 104769, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003808.

Li:2023:DLC

[LLHG23] Hua Li, Jing Lu, Shanqi Hu, and Shengjun Gan. The development and layout of China’s cruise industry in the post-epidemic era: Conference report. Marine Policy, 149(??):Article 105510, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000374.

Liu:2022:UTE

[LLJZ22] Jia Liu, Jing Li, SooCheong (Shawn) Jang, and Yichen Zhao. Understanding tourists’ environmentally responsible behavior at coastal tourism destinations. Marine Policy, 143(??):Article 105178, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002251.

Luna:2023:CFR

[LLL23] Manuel Luna, Ignacio Llorente, and Ladislao Luna. A conceptual framework for risk management in aquaculture. Marine Policy, 147(??):Article 105377, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004249.

Liu:2024:TSD

[LLL^{+}24] Wen-Hong Liu, Chih-Cheng Lin, Cheng-Feng Lee, Chien-Shu Tsai, and Tsung-Yu Yang. Toward sustainable development of tuna longline fishery in Taiwan: Value chain analysis. Marine Policy, 161(??):Article 106010, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000083.

Limmon:2022:CDR

[LLLF22] Gino V. Limmon, Ken Longenecker, Ross Langston, and Erik C. Franklin. Capacity development in reproductive life history studies of tropical fishes in Ambon, Maluku, Indonesia for data limited fisheries. Marine Policy, 136(??):Article 104948, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005595.

Lees:2023:CNA

[LLM^{+}23] Samantha Lees, Catherine Longo, Michael C. Melnychuk, Polly Burns, and Beth Polidoro. Corrected numbers and assumptions for understanding MSC certified fishery withdrawals: Commentary on Pierucci et al. Marine Policy, 155(??):Article 105683, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002105.

Lowitt:2020:ESS

[LLS^{+}20] Kristen Lowitt, Charles Z. Levkoe, Andrew Spring, Colleen Turlo, Patricia L. Williams, Sheila Bird, Chief Dean Sayers, and Melaine Simba. Empowering small-scale, communitybased fisheries through a food systems framework. Marine Policy, 120(??):Article 104150, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20300373.

Liu:2022:RTU

[LLSY22] Wen-Hong Liu, Hsiao-Chien Lee, Wei-Ying Sung, and Tsung-Yu Yang. The roles of Taiwanese universities in coastal revitalization: a study of two case projects. Marine Policy, 139(??):Article 105050, May 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000975.

LeHeron:2020:ROG

[LLT^{+}20] Erena Le Heron, Richard Le Heron, Lara Taylor, Carolyn J. Lundquist, and Alison Greenaway. Remaking ocean governance in Aotearoa, New Zealand through boundarycrossing narratives about ecosystem-based management. Marine Policy, 122(??):Article 104222, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030868X.

Li:2022:EMT

[LLW22] Rongrong Li, Yi Liu, and Qiang Wang. Emissions in maritime transport: a decomposition analysis from the perspective of production-based and consumption-based emissions. Marine Policy, 143(??):Article 105125, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001725.

Lucrezi:2024:STC

[LM24] Serena Lucrezi and Tafadzwa Matiza. Sharks, tourism and conservation: a test of causative and mediating effects on scuba divers’ attitude. Marine Policy, 160(??):Article 105996, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23005298.

Leitao:2023:RRC

[LMC^{+}23a] Francisco Leitão, João N. Monteiro, Pedro Cabral, Maria A. Teodósio, and Ruben H. Roa-Ureta. Revealing the role of crab as bait in octopus fishery: an ecological and fishing approach to support management decisions. Marine Policy, 158(??):Article 105878, December 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004116.

Li:2023:RCD

[LMC^{+}23b] Yunzhou Li, Jian Ma, Abigail Costigan, Xiangyan Yang, Ellen Pikitch, and Yong Chen. Reconciling China’s domestic marine conservation agenda with the global 30×30 initiative. Marine Policy, 156(??):Article 105790, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003238.

Lyons:2023:EII

[LMMT23] Peci Lyons, Sara Mynott, and Jess Melbourne-Thomas. Enabling indigenous innovations to re-centre social licence to operate in the Blue Economy. Marine Policy, 147(??):Article 105384, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004316.

Leon-Mateos:2021:AOS

[LMSLMQ21] Fernando León-Mateos, Antonio Sartal, Lucas López-Manuel, and María A. Quintás. Adapting our sea ports to the challenges of climate change: Development and validation of a Port Resilience Index. Marine Policy, 130(??):Article 104573, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001846.

Li:2021:HCC

[LMT21] Hua Li, Shuhan Meng, and Helong Tong. How to control cruise ship disease risk? Inspiration from the research literature. Marine Policy, 132(??):Article 104652, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002633.

Linke:2023:RAS

[LNRM23] Sebastian Linke, Kåre Nolde Nielsen, and Paulina Ramírez-Monsalve. Roles for advisory science in the international council for the exploration of the sea (ICES). Marine Policy, 148(??):Article 105469, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22005206.

Losen:2023:SAR

[Los23] Barbara Losen. Shark attack risk on Reunion Island: Emphasis on local media construction. Marine Policy, 157(??):Article 105851, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003846.

Losen:2024:USD

[Los24] Barbara Losen. Uses of scientific discourses and markers in shark attack risk management on Reunion Island. Marine Policy, 161(??):Article 106011, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000095.

Lott:2022:BWM

[Lot22] Alexander Lott. Barriers to wildlife movement in straits: Problematizing habitat connectivity across marine ecosystems. Marine Policy, 141(??):Article 105107, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001543.

Lothian:2023:BPM

[Lot23] Sarah Lothian. The BBNJ preamble: More than just window dressing. Marine Policy, 153(??):Article 105642, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001690.

Louey:2022:PBE

[Lou22] Philippa Louey. The Pacific blue economy: an instrument of political maneuver. Marine Policy, 135(??):Article 104880, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004917.

Lovell:2023:MAL

[Lov23] Tricia A. Lovell. Managing abandoned, lost and otherwise discarded fishing gear (derelict gear) in Eastern Caribbean small-scale fisheries: an assessment of legislative, regulatory and policy gaps. Marine Policy, 148(??):Article 105432, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22004791.

Lane:2020:MDE

[LP20] Jesse M. Lane and Michael Pretes. Maritime dependency and economic prosperity: Why access to oceanic trade matters. Marine Policy, 121(??):Article 104180, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20301743.

Lonsdale:2021:SLU

[LP21] Jemma-Anne Lonsdale and Claire Phillips. Space launches and the UK marine environment. Marine Policy, 129(??): Article 104479, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X21000907.

Lawson:2020:DSI

[LR20] Charles Lawson and Michelle Rourke. Digital sequence information as a marine genetic resource under the proposed UNCLOS legally binding instrument. Marine Policy, 122(??):Article 103878, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308061.

Lopez-Rivas:2024:WEV

[LRC24] Jose D. Lopez-Rivas and Juan-Camilo Cardenas. What is the economic value of coastal and marine ecosystem services? A systematic literature review. Marine Policy, 161(??):Article 106033, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000319.

Lamont:2022:MDA

[LRD^{+}22] Timothy A. C. Lamont, Tries B. Razak, Rili Djohani, Noel Janetski, Saipul Rapi, Frank Mars, and David J. Smith. Multi-dimensional approaches to scaling up coral reef restoration. Marine Policy, 143(??):Article 105199, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002469.

Lekshmi:2022:GSS

[LRN^{+}22] P. S. Swathi Lekshmi, Kalidoss Radhakrishnan, R. Narayanakumar, V. P. Vipinkumar, Shinoj Parappurathu, Shyam S. Salim, B. Johnson, and Phalguni Pattnaik. Gender and small-scale fisheries: Contribution to livelihood and local economies. Marine Policy, 136(??):Article 104913, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21005248.

Lee:2022:EVN

[LrS22] Taylor Tae-Hwee Lee and Ae rim So. Economic valuation of a national shipping company using a contingent valuation method (CVM): the case of Korea. Marine Policy, 135(??):Article 104870, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004814.

Lau:2021:CIC

[LSB^{+}21] Jacqueline Lau, Sarah Sutcliffe, Michele Barnes, Emmanuel Mbaru, Innocent Muly, Nyawira Muthiga, Stephen Wanyonyi, and Joshua E. Cinner. COVID-19 impacts on coastal communities in Kenya. Marine Policy, 134 (??):Article 104803, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004140.

Lozano:2022:DWF

[LSD^{+}22] Alejandro J. Garcia Lozano, Jessica L. Decker Sparks, Davina P. Durgana, Courtney M. Farthing, Juno Fitzpatrick, Birgitte Krough-Poulsen, Gavin McDonald, Sara McDonald, Yoshitaka Ota, Nicole Sarto, Andrés M. Cisneros-Montemayor, Gabrielle Lout, Elena Finkbeiner, and John N. Kittinger. Decent work in fisheries: Current trends and key considerations for future research and policy. Marine Policy, 136(??):Article 104922, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005339.

Lizaso:2020:NMF

[LSGG^{+}20] José Luis Sánchez Lizaso, Ivan Sola, Elena Guijarro-García, José Maria Bellido, and Ramón Franquesa. A new management framework for western Mediterranean demersal fisheries. Marine Policy, 112(??):Article 103772, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19306104.

Liu:2021:REG

[LSH21] Chen-Hong Liu, Yue Sun, and Yang Han. Regional environmental governance of the Yellow Sea and Bohai Sea from the perspective of land and sea coordination: Conference report. Marine Policy, 127(??):Article 104446, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000567.

Lu:2020:FAF

[LSY20] Yu-Heng Lu, Takahiro Sajiki, and Nobuyuki Yagi. Factors affecting fisherman satisfaction with fishermen’s selfgovernance organizations: a case study of the Taiwan Donggang Sakuraebi (Sergia lucens) production and management group. Marine Policy, 115(??):Article 103819, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19304129.

Long:2021:CAG

[LTJ^{+}21] Stephen Long, Grace Thurlow, Peter J. S. Jones, Andrew Turner, Sylvestre M. Randrianantenaina, Tom Gammage, Jessica Savage, and Jeremie R. Ndriamanja. Critical analysis of the governance of the Sainte Luce Locally Managed Marine Area (LMMA), southeast Madagascar. Marine Policy, 127(??):Article 103691, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302635.

Lin:2023:UDC

[LTJ^{+}23] Chih-Cheng Lin, Kuo-Huan Ting, Hao-Tang Jhan, Chung-Lun Liu, Li-Shu Chen, and Wen-Hong Liu. Universitydriven coastal partnerships to address ghost gear pollution: a case study in Penghu Islands of Taiwan. Marine Policy, 155(??):Article 105732, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002658.

Lieske:2020:SRA

[LTRA20] David J. Lieske, Laura McFarlane Tranquilla, Robert A. Ronconi, and Sue Abbott. “Seas of risk”: Assessing the threats to colonial-nesting seabirds in Eastern Canada. Marine Policy, 115(??):Article 103863, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305263.

Lee:2022:EGG

[LTSL22] Hsiao-Chien Lee, Ying-Lee Tseng, Wei-Ying Sung, and Wen-Hong Liu. Exploring the gender gap in marine affairs in Taiwan. Marine Policy, 137(??):Article 104937, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21005480.

Langford:2023:IIP

[LTW23] Alexandra Langford, Welem Turupadang, and Scott Waldron. Interventionist industry policy to support local valueadding: Evidence from the Eastern Indonesian seaweed industry. Marine Policy, 151(??):Article 105561, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300088X.

Lv:2020:BRI

[Lv20] Fang-Yuan Lv. Book review: Issues in Maritime Cyber Security, Joseph Direnzo, Nicole K. Drumhiller, Fried S. Roberts (Eds.) (2017), 568, ISBN: 978-1-78643-673-3. Marine Policy, 117(??):Article 103963, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301639.

Langlet:2023:LUW

[LV23] Arne Langlet and Alice B. M. Vadrot. Not ‘undermining’ who? Unpacking the emerging BBNJ regime complex. Marine Policy, 147(??):Article 105372, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004195.

Li:2022:HWE

[LW22] Jing-Mei Li and Na Wang. How and to what extent is ecosystem services economic valuation used in coastal and marine management in China? Marine Policy, 138(??):Article 104976, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000239.

Li:2023:DTP

[LWG23] Rongrong Li, Qiang Wang, and Yunfei Ge. Does trade protection undercut the green efficiency of the marine economy? A case study. Marine Policy, 157(??):Article 105864, November 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003974.

Langford:2021:MCA

[LWSS21] Alexandra Langford, Scott Waldron, Sulfahri, and Hasnawati Saleh. Monitoring the COVID-19-affected Indonesian seaweed industry using remote sensing data. Marine Policy, 127(??):Article 104431, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000415.

Liu:2020:ECB

[LWX^{+}20] Baiqiong Liu, Jing Wang, Min Xu, Zaifeng Wang, and Lin Zhao. Evaluation of the comprehensive benefit of various marine exploitation activities in China. Marine Policy, 116(??):Article 103924, June 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930315X.

Li:2023:FSD

[LWZ23] Fenghua Li, Leifan Wang, and Xiaoyu Zhou. Fulfiling the State duty to cooperate on the discharge of Fukushima nuclear contaminated water: Potential pathways. Marine Policy, 150(??):Article 105546, April 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000738.

Lin:2022:IRS

[LXH^{+}22] Yanhong Lin, Xiongzhi Xue, Wenjia Hu, Qiuming Chen, Yang Zhang, Faming Huang, Jinliang Huang, and Chao Wang. Implementing regional sea use planning system in China: Evolution, drivers and experiences. Marine Policy, 146(??):Article 105298, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003451.

Li:2021:ECM

[LXSX21] Fuzhu Li, Wenxiu Xing, Meng Su, and Juan Xu. The evolution of China’s marine economic policy and the labor productivity growth momentum of marine economy and its three economic industries. Marine Policy, 134(??):Article 104777, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003882.

Liu:2021:ECM

[LXW^{+}21] Baiqiong Liu, Min Xu, Jing Wang, Zaifeng Wang, and Lin Zhao. Evaluation of China’s marine economic growth quality based on set pair analysis. Marine Policy, 126(??):Article 104405, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21000154.

Li:2023:CSE

[LYHL23] Wei Li, Qi Yue, Heng Hu, and Liang Liu. Current status of ecological compensation for bohai sea reclamation and management countermeasures. Marine Policy, 155(??):Article 105778, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003111.

Lin:2022:STE

[LYL^{+}22] Ying Lin, Yiqun Yang, Peiliang Li, Cuicui Feng, Jieqiong Ding, Junyu Zhou, Qutu Jiang, and Guanqiong Ye. Spatial-temporal evaluation of marine ecological civilization of Zhejiang Province, China. Marine Policy, 135(??):Article 104835, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004462.

Lili:2021:RAU

[LYQM21] Ding Lili, Yang Ying, Hu Qiuhui, and Liu Mengxi. Residents’ acceptance of using desalinated water in China based on the theory of planned behaviour (TPB). Marine Policy, 123(??):Article 104293, January 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309398.

Liu:2023:TIF

[LZL23] Shiyu Liu, Wei Zeng, and Xiyao Li. Toward an integrated framework of ocean literacy: a Delphi-AHP approach. Marine Policy, 157(??):Article 105830, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003639.

Lopera:2023:OPI

[LZRC23] Luisa Lopera, Paula A. Zapata-Ramírez, and Yuley Cardona. Overview and potential implementation of a marine protected area network between Colombia, Nicaragua, and Jamaica. Marine Policy, 150(??):Article 105411, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004584.

Li:2022:CEG

[LZZ22] Hua Li, Li Zhu, and Mengyuan Zheng. Is the cruise enterprise giant also a giant of social responsibility for sustainable development? Marine Policy, 138(??):Article 105009, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000562.

Ma:2020:ELA

[Ma20] Xuechan Ma. An economic and legal analysis of trade measures against illegal, unreported and unregulated fishing. Marine Policy, 117(??):Article 103980, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309455.

Marin:2024:DST

[MAA^{+}24] Charlotte Marin, Olugbenga M. Adewumi, Frank Asche, Taryn M. Garlock, Dadi M. Kristofersson, Kai Lorenzen, and Bixuan Yang. Does seafood trade enhance seafood availability in developing countries? The case of Nigeria. Marine Policy, 161(??):Article 106030, March 2024.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000289.

Merten:2022:CIF

[MAG^{+}22] Wessley Merten, Richard Appeldoorn, Abby Grove, Alfonso Aguilar-Perera, Freddy Arocha, and Roberto Rivera. Condition of the international fisheries, catch and effort trends, and fishery data gaps for dolphinfish (Coryphaena hippurus) from 1950 to 2018 in the Western Central Atlantic Ocean. Marine Policy, 143(??):Article 105189, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002366.

Magnuson:2023:OSF

[Mag23] Briana Magnuson. One-size-fits-all? How north–south inequality challenges CITES’ ability to protect the scalloped hammerhead. Marine Policy, 158(??):Article 105867, December 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004001.

Mahadeo:2022:MSP

[Mah22] Sarah Mahadeo. Marine spatial planning in the Eastern Caribbean: Trends and progress. Marine Policy, 145(??):Article 105277, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003244.

McClanahan:2022:FCR

[MAK22] T. R. McClanahan, M. K. Azali, and J. K. Kosgei. Fish catch responses to COVID-19 disease curfews dependent on compliance, fisheries management, and environmental contexts. Marine Policy, 144(??):Article 105239, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200286X.

Mallin:2021:CSL

[Mal21] Marc-Andrej Felix Mallin. Corrigendum to: From sealevel rise to seabed grabbing: the political economy of climate change in Kiribati [Mar. Policy 97 (2018) 244– 252]. Marine Policy, 129(??):Article 104509, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001202.

Meza-Arce:2020:UWS

[MAMCHP^{+}20] Mercedes I. Meza-Arce, Luis Malpica-Cruz, Mauricio E. Hoyos-Padilla, Francisco J. Mojica, María Concepción Arredondo-García, Claudia Leyva, Rebeca Zertuche-Chanes, and Omar Santana-Morales. Unraveling the white shark observation tourism at Guadalupe Island, Mexico: Actors, needs and sustainability. Marine Policy, 119(??):Article 104056, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19307353.

Marcondes:2020:CGB

[Mar20a] Danilo Marcondes. Conservationist geopolitics: Brazilian foreign policy and the South Atlantic Whale Sanctuary. Marine Policy, 120(??):Article 104054, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301666.

Maxwell:2020:WEN

[MAR^{+}20b] Kimberley Maxwell, Shaun Awatere, Kelly Ratana, Kate Davies, and Caine Taiapa. He waka eke noa/we are all in the same boat: a framework for co-governance from Aotearoa, New Zealand. Marine Policy, 121(??):Article 104213, November 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20308599.

Marghelis:2021:MDA

[Mar21a] Aris Marghelis. The maritime delimitation agreement between Greece and Italy of 9 June 2020: an analysis in the light of International Law, national interest and regional politics. Marine Policy, 126(??):Article 104403, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000130.

Martinson:2021:CSF

[Mar21b]  Ryan Martinson.  Catching sovereignty fish: Chi-

nese fishers in the southern Spratlys. Marine Policy, 125(??):Article 104372, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2031023X.

Matthiasson:2020:SCR

[Mat20] Thorolfur Matthiasson. Splitting the catch, remuneration of Icelandic fishermen in a sea of change. Marine Policy, 117(??):Article 103959, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307729.

Maynou:2022:SPF

[May22] Francesc Maynou. Sale price flexibilities of Mediterranean hake and red shrimp. Marine Policy, 136 (??):Article 104904, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21005157.

Moses:2021:WBC

[MB21] Jonathon W. Moses and Anne Margrethe Brigham. Whose benefit? A comparative perspective for the ISA. Marine Policy, 131(??):Article 104550, September 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001615.

Mannaart:2022:FLI

[MB22] Michael Mannaart and Arabelle Bentley. Fishing for litter: From the implementation of practical actions locally, to its spin-offs and the adoption of a new legally adopted waste type at continental scale, a success story. Marine Policy, 145(??):Article 105256, November 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003037.

Mbatha:2022:UPM

[Mba22] Philile Mbatha. Unravelling the perpetuated marginalization of customary livelihoods on the coast by plural and multi-level conservation governance systems. Marine Policy, 143(??):Article 105143, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001907.

McGee:2022:MPP

[MBB^{+}22] Gord McGee, Josie Byington, John Bones, Sally Cargill, Megan Dickinson, Kelly Wozniak, and Kylee A. Pawluk. Marine plan partnership for the North Pacific Coast: Engagement and communication with stakeholders and the public. Marine Policy, 142(??):Article 104613, August 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002244.

Mendez:2024:BBS

[MBC^{+}24] Cynthia Mendez, Danilo E. Bustamante, Martha S. Calderon, Cecilia Gauna, Leila Hayashi, Daniel Robledo, Claudia Tapia-Larios, Iona Campbell, Renato Westermeier, and Pedro Murúa. Biosecurity baseline for a sustainable development of seaweed aquaculture in Latin America. Marine Policy, 159(??):Article 105933, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004669.

Maini:2023:CVL

[MBDG23] Bani Maini, Jessica L. Blythe, Emily S. Darling, and Georgina G. Gurney. Charting the value and limits of other effective conservation measures (OECMs) for marine conservation: a Delphi study. Marine Policy, 147(??):Article 105350, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003979.

Mumtaz:2024:ARB

[MBL24] Ubaid Ullah Mumtaz, Paul Bergey, and Nicholas Letch. Assessing the role of blockchain technology for marine bunkering operations — a case study of task technology fit. Marine Policy, 159(??):Article 105909, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004426.

Martin-Betancor:2024:EMS

[MBORGN24] M. Martín-Betancor, J. Osorio, A. Ruiz-García, and I. Nuez. Evaluation of maritime spatial planning for offshore wind energy in the Canary Islands: a comparative analysis. Marine Policy, 161(??):Article 106046, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000447.

Masalaci:2023:DDT

[MÇ23] Burcu Çelik Maşalacı and Erkan Çakır. Differences in deficiency types causing ship detentions at the Black Sea Region during the Covid-19 pandemic and pre-pandemic. Marine Policy, 151(??):Article 105553, May 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000805.

Mallik:2022:ICL

[MCBN22] Abhijit Mallik, Puja Chakraborty, Shashi Bhushan, and Binaya Bhusan Nayak. Impact of COVID-19 lockdown on aquatic environment and fishing community: Boon or bane? Marine Policy, 141(??):Article 105088, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200135X.

McClanahan:2020:WCP

[McC20] Timothy Rice McClanahan. Wilderness and conservation policies needed to avoid a coral reef fisheries crisis. Marine Policy, 119(??):Article 104022, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307225.

McClanahan:2021:MRM

[McC21] Tim R. McClanahan. Marine reserve more sustainable than gear restriction in maintaining long-term coral reef fisheries yields. Marine Policy, 128(??):Article 104478, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000890.

McClanahan:2024:PPA

[McC24] Tim R. McClanahan. Perceptions of preparedness to address climate change threats in the western Indian Ocean. Marine Policy, 162(??):Article 106055, April 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000538.

Mogila:2021:BBG

[MCKZ21] Zbigniew Mogila, Dorota Ciolek, Jakub M. Kwiatkowski, and Jacek Zaucha. The Baltic blue growth — a country-level shift-share analysis. Marine Policy, 134 (??):Article 104799, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004103.

Mallory:2022:COC

[MCL22] Tabitha Grace Mallory, Andrew Chubb, and Sallie Lau. China’s ocean culture and consciousness: Constructing a maritime great power narrative. Marine Policy, 144(??):Article 105229, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002767.

Martinell:2020:CHS

[MCPS20] Daniel Peñalosa Martinell, Tim Cashion, Robert Parker, and U. Rashid Sumaila. Closing the high seas to fisheries: Possible impacts on aquaculture. Marine Policy, 115(??):Article 103854, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19306931.

Menini:2022:PPS

[MCR22] Elisabetta Menini, Anindita Chakraborty, and Stephen E. Roady. Public participation in seabed mining in areas beyond national jurisdiction: Lessons learned from national regulators in the terrestrial mining sector. Marine Policy, 146(??):Article 105308, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003554.

Maestro:2020:AMP

[MCRPC20] María Maestro, Juan Adolfo Chica-Ruiz, and María Luisa Pérez-Cayeiro. Analysis of marine protected area management: the Marine Park of the Azores (Portugal). Marine Policy, 119(??):Article 104104, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930819X.

Martins:2023:OES

[MCRR23] Maria Clara Iruzun Martins, Matt ID Carter, Sally Rouse, and Debbie JF Russell. Offshore energy structures in the North Sea: Past, present and future. Marine Policy, 152(??):Article 105629, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001562.

Munim:2023:TPA

[MCTN23] Ziaul Haque Munim, Mohammed Mojahid Hossain Chowdhury, Hasan Mahbub Tusher, and Theo Notteboom. Towards a prioritization of alternative energy sources for sustainable shipping. Marine Policy, 152(??):Article 105579, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001069.

Martin:2022:SDN

[MCW^{+}22a] Carol L. Martin, Belinda Curley, Kim Wolfenden, Marcel Green, and Natalie A. Moltschaniwskyj. The social dimension to the New South Wales Shark Management Strategy, 2015–2020, Australia: Lessons learned.

Marine Policy, 141(??):Article 105079, July 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001269.

Mayaud:2022:TNS

[MCW^{+}22b] Raphael Mayaud, Juliana Castrillon, Craig Wilson, David Peel, Joshua N. Smith, Greta Dalle Luche, Jenny Allen, and Susan Bengtson Nash. Traffic in a nursery: Ship strike risk from commercial vessels to migrating humpback whales (Megaptera novaeangliae) in a rapidly developing Australian urban embayment. Marine Policy, 146(??):Article 105332, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003797.

Moranta:2023:SPT

[MdCLL^{+}23] Joan Moranta, David Florido del Corral, Lucía López-López, Sílvia Gómez, Julio Agujetas, Marta Albo-Puigserver, Zoe W. Brent, Lamine Bathily, Rita Calvário, Marta Cavallé, Marta Coll, Annya Crane, Ericka C. L. de Oliveira, Irmak Ertör, Paloma Herrera-Racionero, Francisco Leitão, Lluís Miret-Pastor, Beatriz Patraca, Miquel Ortega, José J. Pascual-Fernández, Jaime Ramón-Bruquetas, Jorge Sáez-Jiménez, Abdoulaye Seck-Papalaye, and Sebastián Villasante. Sowing potential transformative changes in the fishing and agrifood systems — conference report. Marine Policy, 154(??):Article 105673, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002002.

McClanahan:2023:MEF

[MDG^{+}23] Timothy R. McClanahan, Stephanie D’Agata, N. A. J. Graham, Maxwell Azali Kodia, and Joseph M. Maina. Multivariate environment–fish biomass model informs sustainability and lost income in Indian Ocean coral reefs. Marine Policy, 152(??):Article 105590, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001173.

Mendenhall:2022:DDP

[MDJ^{+}22] Elizabeth Mendenhall, Elizabeth De Santo, Mathias Jankila, Elizabeth Nyman, and Rachel Tiller. Direction, not detail: Progress towards consensus at the fourth intergovernmental conference on biodiversity beyond national jurisdiction. Marine Policy, 146(??):Article 105309, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003566.

Mansour:2021:GMT

[MDLA21] Shawky Mansour, Stephen Darby, Julian Leyland, and Peter M. Atkinson. Geospatial modelling of tropical cyclone risk along the northeast coast of Oman: Marine hazard mitigation and management policies. Marine Policy, 129(??):Article 104544, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100155X.

Murphy:2021:UMM

[MEG^{+}21] Robert Murphy, Austin Estabrooks, John Gauvin, Steven Gray, Anita C. Kroska, Nathan Wolf, and Bradley P. Harris. Using mental models to quantify linear and non-linear relationships in complex fishery systems. Marine Policy, 132(??):Article 104695, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003067.

Mendenhall:2023:MMW

[Men23] Elizabeth Mendenhall. Making the most of what we already have: Activating UNCLOS to combat marine plastic pollution. Marine Policy, 155(??):Article 105786, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003196.

Maxwell:2020:KPG

[MF20] R. J. Maxwell and R. Filgueira. Key players in the Grieg NL Placentia Bay Atlantic Salmon Aquaculture Project: a social network analysis. Marine Policy, 113(??):Article 103800, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19305184.

March:2022:SSF

[MF22] Antaya March and Pierre Failler. Small-scale fisheries development in Africa: Lessons learned and best practices for enhancing food security and livelihoods. Marine Policy, 136(??):Article 104925, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005364.

Monios:2023:PAP

[MF23] Jason Monios and Laurent Fedi. The principal-agent problem in hierarchical policy making: a failure of policy to support the transition to LNG as an alternative shipping fuel. Marine Policy, 157(??):Article 105846, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003792.

Maskaeva:2024:ASE

[MFC^{+}24] Asiya Maskaeva, Pierre Failler, Honita Cowaloosur, Philippe Lallemand, and Jerry Mang’ena. Assessment of socioeconomic and ecosystem services of the blue economy in Tanzania using the UNECA’s Blue Economy Valuation Toolkit. Marine Policy, 159(??):Article 105920, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004530.

Murray:2023:SCM

[MFF^{+}23] Grant D. Murray, Robin Fail, Luke Fairbanks, Lisa M. Campbell, Linda D’Anna, and Joshua Stoll. Seafood consumption and the management of shellfish aquaculture. Marine Policy, 150(??):Article 105534, April 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000611.

Madarcos:2022:DMS

[MFG^{+}22] K. Madarcos, M. Fortnam, L. Gajardo, T. Chaigneau, RJ. Manucan, G. Cadigal, J. Matulac, L. Creencia, B. Gonzales, and L. Evans. Doing marine spatial zoning in coastal marine tropics: Palawan’s Environmental Critical Areas Network (ECAN). Marine Policy, 145(??):Article 105207, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002548.

Marshall:2023:CET

[MFG^{+}23] Rachel C. Marshall, Jonathan A. D. Fisher, Paul Gatti, Erin H. Carruthers, Dominique Robert, and Arnault Le Bris. Combining electronic tag data and fish harvesters’ knowledge to identify ocean management concerns. Marine Policy, 147(??):Article 105393, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004407.

Mulalap:2020:TKB

[MFH^{+}20] Clement Yow Mulalap, Tekau Frere, Elise Huffer, Edvard Hviding, Kenneth Paul, Anita Smith, and Marjo K. Vierros. Traditional knowledge and the BBNJ instrument. Marine Policy, 122(??):Article 104103, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308620.

McAteer:2021:LMO

[MFM21] Benedict McAteer, Wesley Flannery, and Brendan Murtagh. Linking the motivations and outcomes of volunteers to understand participation in marine community science. Marine Policy, 124(??):Article 104375, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310265.

Mamede:2023:HMP

[MFRQ23] Renato Mamede, Rosa Freitas, Ana Maria Rodrigues, and Victor Quintino. Habitat mapping of Portugal’s northwestern continental shelf and its application to implement the European Marine Strategy Framework Directive. Marine Policy, 155(??):Article 105765, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002981.

Magalhaes:2022:OPS

[MFT^{+}22] Karine Matos Magalhães, José Souto Rosa Filho, Carlos Eduardo Peres Teixeira, Clemente Coelho-Jr, Maria Cecilia Santana Lima, Ana Maria Costa Souza, and Marcelo O. Soares. Oil and plastic spill: 2021 as another challenging year for marine conservation in the South Atlantic Ocean. Marine Policy, 140(??):Article 105076, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001233.

Maruf:2022:TAU

[MG22] Maruf and Warwick Gullett. Tackling anthropogenic underwater noise through the Convention on Biological Diversity: Progress and future development. Marine Policy, 146(??):Article 105293, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003402.

Michel:2024:PAM

[MGBS24] Mathilde Michel, Benjamin Guichard, Julie Béesau, and Flore Samaran. Passive acoustic monitoring for assessing marine mammals population in European waters: Workshop conclusions and perspectives. Marine Policy, 160(??):Article 105983, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300516X.

Mathis:2022:ROP

[MGDJ22] Jennifer Estelle Mathis, Marianne Carliez Gillet, Henri Disselkoen, and Jenna R. Jambeck. Reducing ocean plastic pollution: Locally led initiatives catalyzing change in South and Southeast Asia. Marine Policy, 143(??):Article 105127, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001749.

Mileski:2020:HTC

[MGF20] Joan P. Mileski, Cassia Bomer Galvao, and Zaida Denise Forester. Human trafficking in the commercial fishing industry: a multiple case study analysis. Marine Policy, 116(??):Article 103616, June 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19300533.

Macusi:2021:VDS

[MGS21] Edison D. Macusi, Rollan C. Geronimo, and Mudjekeewis D. Santos. Vulnerability drivers for small pelagics and milkfish aquaculture value chain determined through online participatory approach. Marine Policy, 133(??):Article 104710, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003213.

McQuaw:2020:WCM

[MH20a] Kristin McQuaw and Ray Hilborn. Why are catches in mixed fisheries well below TAC? Marine Policy, 117(??):Article 103931, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19300909.

Muallil:2020:PRA

[MH20b] Richard N. Muallil and Mohammad Gafor N. Hapid. Preliminary report on an artisanal fishery for thresher sharks (Alopias spp) in Tawi-Tawi, Southern Philippines. Marine Policy, 117(??):Article 103894, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304713.

Mendenhall:2023:BAL

[MH23] Elizabeth Mendenhall and Kahlil Hassanali. The BBNJ agreement and liability. Marine Policy, 150(??):Article 105549, April 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000763.

Mendenhall:2024:WBA

[MH24] Elizabeth Mendenhall and Rebecca R. Helm. When the “best available science” is not good enough: the need for supporting scientific research in the United Nations treaty to protect biodiversity beyond national jurisdiction. Marine Policy, 161(??):Article 105940, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004736.

Murphy:2020:CSP

[MHEW20] Robert Murphy, Bradley Harris, Austin Estabrooks, and Nathan Wolf. Capturing stakeholder perspectives through a collaboration with a commercial fishing cooperative. Marine Policy, 117(??):Article 103948, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308425.

Mendenhall:2020:CCI

[MHN^{+}20] Elizabeth Mendenhall, Cullen Hendrix, Elizabeth Nyman, Paige M. Roberts, John Robison Hoopes, James R. Watson, Vicky W. Y. Lam, and U. Rashid Sumaila. Climate change increases the risk of fisheries conflict. Marine Policy, 117(??):Article 103954, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304622.

Muallil:2024:RAS

[MIH^{+}24] Richard N. Muallil, Roger A. Irilis, Argamar A. Habibuddin, Nur aisa S. Abdulmajid, Jaro O. Ajik, and Ronaldo A. Ancheta. A rapid assessment of the status, trends and challenges in small-scale commercial sardine fisheries in the Sulu Archipelago, southern Philippines. Marine Policy, 160(??):Article 105965, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004980.

Miller:2022:DTI

[Mil22] Karly Marie Miller. Disentangling tourism impacts on small-scale fishing pressure. Marine Policy, 137(??): Article 104960, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X22000070.

Marega-Imamura:2020:SCN

[MIMS^{+}20] Marcela Marega-Imamura, Fernanda Michalski, Kátia Silva, Alexandre Schiavetti, Yvonnick Le Pendu, and Leonardo de Carvalho Oliveira. Scientific collaboration networks in research on human threats to cetaceans in Brazil. Marine Policy, 112(??):Article 103738, February 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305876.

Mazur:2020:EIF

[MJ20] Mackenzie Dale Mazur and Teresa R. Johnson. Effects of increases in fishery resource abundance on conservation compliance. Marine Policy, 122(??):Article 104217, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20308630.

Maya-Jariego:2023:INS

[MJ23]  Isidro Maya-Jariego.  Informal norms sustain non-

compliance and poaching among spearfishermen. A response to Tudela et al. (2022). Marine Policy, 148 (??):Article 105472, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22005231. See [TGM22].

Maya-Jariego:2021:PBM

[MJMAA21] Isidro Maya-Jariego, Inmaculada Martínez-Alba, and Deniza Alieva. “Plenty of black money”: Netnography of illegal recreational underwater fishing in southern Spain. Marine Policy, 126(??):Article 104411, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2100021X. See reply [SCD21].

Maya-Jariego:2022:RPR

[MJMAA22] Maya-Jariego, Martínez-Alba, and D. Alieva. Responsible practice of recreational spearfishing is needed. A response to sbragaglia et al. (2021). Marine Policy, 139(??):Article 105022, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000690. See [SCD21].

Mulder:2024:WSM

[MK24] Thomas Mulder and Natalie Klein. WHO at sea: Maritime dimensions of the proposed pandemic treaty and changes to the International Health Regulations 2005. Marine Policy, 159(??):Article 105918, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004517.

Marlianingrum:2021:VHQ

[MKAF21] Peggy Ratna Marlianingrum, Tridoyo Kusumastanto, Luky Adrianto, and Achmad Fahrudin. Valuing habitat quality for managing mangrove ecosystem services in coastal Tangerang District, Indonesia. Marine Policy, 133(??):Article 104747, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003584.

Montenero:2021:QQD

[MKB21] Kelly Montenero, Chris Kelble, and Kathy Broughton. A quantitative and qualitative decision-making process for selecting indicators to track ecosystem condition. Marine Policy, 129(??):Article 104489, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001007.

Mehault:2020:EPR

[MKM^{+}20] Sonia Méhault, Dorothée Kopp, Marie Morfin, Jean Pierre Cuillandre, Pascal Larnaud, Fabien Morandeau, Jean Philippe Vacherot, and Thomas Rimaud. Escapement patterns of red mullet, sole, horse mackerel and hake facing two trawl selective devices. Marine Policy, 116(??):Article 103844, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18307930.

Maltby:2023:BEC

[MKM23] Katherine M. Maltby, Sabrina Kerin, and Katherine E. Mills. Barriers and enablers of climate adaptation in fisheries: Insights from northeast US fishing communities. Marine Policy, 147(??):Article 105331, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003785.

McShane:2021:AAF

[MKS21] Paul McShane, Ian Knuckey, and Sevaly Sen. Access and allocation in fisheries: the Australian experience. Marine Policy, 132(??):Article 104702, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003134.

Molino:2020:SDS

[MKSG20] Grace D. Molino, Melissa A. Kenney, and Ariana E. Sutton-Grier. Stakeholder-defined scientific needs for coastal resilience decisions in the Northeast U.S. Marine Policy, 118(??):Article 103987, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306955.

Mangubhai:2021:EGI

[ML21] Sangeeta Mangubhai and Sarah Lawless. Exploring gender inclusion in small-scale fisheries management and development in Melanesia. Marine Policy, 123(??):Article 104287, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309337.

Ma:2022:APC

[ML22] Mingfei Ma and Zhe Li. Amendment process and comments of the provisions of underwater cultural heritage in the Draft Exploitation Regulations in the area. Marine Policy, 140(??):Article 105081, June 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001282.

Montecalvo:2023:OPS

[MLAS23] Isabella Montecalvo, Philippe Le Billon, Chris Arsenault, and Milko Schvartzman. Ocean predators: Squids, Chinese fleets and the geopolitics of high seas fishing. Marine Policy, 152(??):Article 105584, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001112.

Maia:2024:HBA

[MLE^{+}24] Michele Cristina Maia, Gabriela Lamego, Carla I. Elliff, Jana M. Del Favero, Juliana Leonel, and Catarina R. Marcolin. Harassment and bullying aboard: Impacts of gender inequality on ocean professionals. Marine Policy, 160(??):Article 105946, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004797.

Madon:2022:PAD

[MLJ^{+}22] Bénédicte Madon, Damien Le Guyader, Jean-Luc Jung, Benjamin De Montgolfier, Pascal Jean Lopez, Eric Foulquier, Laurent Bouveret, and Iwan Le Berre. Pairing AIS data and underwater topography to assess maritime traffic pressures on cetaceans: Case study in the Guadeloupean waters of the Agoa sanctuary. Marine Policy, 143(??):Article 105160, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200207X.

Mu:2023:ECC

[MLM23] Manning Mu, Mengxi Liu, and Wen Ma. Evaluating Chinese consumers’ purchase intentions for daily chemical products without plastic microbeads using an extended theory of planned behavior model. Marine Policy, 154(??):Article 105671, August 2023. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23001987.

Macusi:2021:FIC

[MLMD21] Edison D. Macusi, Arianne Kay O. Liguez, Erna S. Macusi, and Larry N. Digal. Factors influencing catch and support for the implementation of the closed fishing season in Davao Gulf, Philippines. Marine Policy, 130(??):Article 104578, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001895.

Mantziaris:2021:FRB

[MLVT21] Stamatis Mantziaris, Angelos Liontakis, George Valakas, and Irene Tzouramani. Family-run or business-oriented fisheries? Integrating socioeconomic and environmental aspects to assess the societal impact. Marine Policy, 131(??):Article 104591, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002025.

Myers:2020:REU

[MM20] Hannah J. Myers and Michael J. Moore. Reducing effort in the U.S. American lobster (Homarus americanus) fishery to prevent North Atlantic right whale (Eubalaena glacialis) entanglements may support higher profits and long-term sustainability. Marine Policy, 118(??):Article 104017, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19309157.

Macedo:2021:RGB

[MM21a] Heitor Schulz Macedo and Rodrigo Pereira Medeiros. Rethinking governance in a Brazilian multiple-use marine protected area. Marine Policy, 127(??):Article 103235, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18305529.

Motoori:2021:RSS

[MM21b] Ran Motoori and Benjamin C. McLellan. Resource security strategies and preferences for deep ocean mining from a community survey in Japan. Marine Policy, 128(??):Article 104511, June 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001226.

McClenachan:2022:TWC

[MM22] Loren McClenachan and Allie Moulton. Transitions from wild-caught fisheries to shellfish and seaweed aquaculture increase gender equity in Maine. Marine Policy, 146(??):Article 105312, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003591.

Mesquita:2023:IRF

[MM23] Isabeli Cristina Gomes Mesquita and Rodrigo Pereira Medeiros. Integrating research and fishing extension approaches to engage small-scale fishers in the participatory evaluation and voluntary use of bycatch reduction devices. Marine Policy, 152(??):Article 105599, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001264.

Marchessaux:2023:IBC

[MMB^{+}23] Guillaume Marchessaux, Maria Cristina Mangano, Sergio Bizzarri, Charaf M’Rabet, Elena Principato, Nicola Lago, Dimitri Veyssiere, Marie Garrido, Steven B. Scyphers, and Gianluca Sarà. Invasive blue crabs and smallscale fisheries in the Mediterranean sea: Local ecological knowledge, impacts and future management. Marine Policy, 148(??):Article 105461, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22005085.

Macusi:2022:WPS

[MMC^{+}22] Edison D. Macusi, Erna S. Macusi, Charlen Mae G. Canales, Anna Barboza, and Larry N. Digal. Women’s participation and support for the implementation of the closed fishing season in Davao Gulf, Philippines. Marine Policy, 143(??):Article 105133, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001804.

Murillas-Maza:2021:EPS

[MMGAA^{+}21] Arantza Murillas-Maza, Antonio García-Allut, Martin Aranda, César Pazos, Ane Iriondo, and Miriam Montero. Enhancing the participation of small-scale fishing sector in decision-making: Good practices from Atlantic in EU. Marine Policy, 129(??):Article 104533, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001445.

Mudege:2022:ICG

[MMK^{+}22] Netsayi Noris Mudege, Catherine Mawia Mwema, Keagan Kakwasha, Andrew Chisopo, Chikondi Manyungwa-Pasani, Lisungu Banda, Emmanuel Kaunda, and Pamela Marinda. The impacts of Covid-19 on gender dynamics and power relations among men and women involved in cross border fish trade in Zambia and Malawi. Marine Policy, 146(??):Article 105322, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003694.

Macusi:2023:WLS

[MMK^{+}23] Edison D. Macusi, Imee S. Maynawang, Robert E. Katikiro, Jee Grace Suyo, and Erna S. Macusi. Women lead smallscale fisheries to sustainability in Surigao del Sur, Philippines. Marine Policy, 155(??):Article 105768, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003019.

Macusi:2022:ICF

[MMM^{+}22] Edison D. Macusi, Ilah Dianne G. Morales, Erna S. Macusi, Aijelyn Pancho, and Larry N. Digal. Impact of closed fishing season on supply, catch, price and the fisheries market chain. Marine Policy, 138(??):Article 105008, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000550.

McManus:2022:ICD

[MMNC22] Catherine McManus, Alistair McIlgorm, Rachel Nichols, and Andrew Cooper. An initial consideration of data availability issues in downscaling ocean accounting to inform sustainable aquaculture development: the example of Clew Bay, Ireland. Marine Policy, 145(??):Article 105286, November 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003335.

Morse:2024:PSM

[MMO^{+}24] Molly Morse, Douglas McCauley, Sara Orofino, Keenan Stears, Samantha Mladjov, Jenn Caselle, Tyler Clavelle, and Ryan Freedman. Preferential selection of marine protected areas by the recreational scuba diving industry. Marine Policy, 159(??):Article 105908, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004414.

Matoso:2024:WKD

[MMS24] Thomaz Bosquetto Matoso, Gustavo Goulart Moreira Moura, and Thiago Zagonel Serafini. A wave knowledge description of a surf break in Southern Brazil. Marine Policy, 159(??):Article 105928, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300461X.

Mendez-Medina:2020:ACD

[MMSB^{+}20] Crisol Méndez-Medina, Birgit Schmook, Xavier Basurto, Stuart Fulton, and Alejandro Espinoza-Tenorio. Achieving coordination of decentralized fisheries governance through collaborative arrangements: a case study of the Sian Ka’an Biosphere Reserve in Mexico. Marine Policy, 117(??):Article 103939, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18306377.

Mirera:2023:FPD

[MMW^{+}23] David Oersted Mirera, Esther Wairimu Magondu, Miriam Wambui Wainaina, Brendan Muli, Douglas Okemwa, Rose Angulu, Irene Heba, and Hellen Moyoni. Fish preference at different value chain levels and implications for management of mariculture. Marine Policy, 157(??):Article 105845, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003780.

Marques:2021:ECM

[MN21] Pedro Henrique Dias Marques and Sebastião Cerqueira Neto. Environmental conflicts and management of marine resources in the Cassurubá Complex, Bahia, Brazil. Marine Policy, 127(??):Article 104435, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000452.

Mazloumi:2023:PCR

[MN23] Nastaran Mazloumi and Simon Nicol. Preparing for climate related impacts: Knowledge assessment of the life history of key exploited fish species managed in a climate hotspot. Marine Policy, 148(??):Article 105381, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004286.

Macusi:2024:FWB

[MNGM24] Edison D. Macusi, Ivy M. Nallos, Maria C. Galveia, and Erna S. Macusi. Fishers’ well-being as affected by anthropogenic stressors in Surigao del Sur, Philippines. Marine Policy, 162(??):Article 106049, April 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000472.

Mohan:2023:IND

[Moh23] Preeya S. Mohan. Implementing nationally determined contributions under the Paris Agreement: an assessment of ocean-based climate action in Caribbean Small Island Developing States. Marine Policy, 155(??):Article 105787, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003202.

Morseletto:2020:NFP

[Mor20] Piero Morseletto. A new framework for policy evaluation: Targets, marine litter, Italy and the Marine Strategy Framework Directive. Marine Policy, 117(??):Article 103956, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19304002.

Mota:2020:FHC

[Mot20] Rui Pedro Mota. Fishery harvest control rules from bioeconomic optimization. Marine Policy, 115(??):Article 103865, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19308140. See corrigen-

dum [Mot22].

Mota:2022:CLF

[Mot22] Rui Pedro Mota. Corrigendum to ‘Fishery harvest control rules from bioeconomic optimization’ [Mar. Policy Vol. 115, May 2020, 103865]. Marine Policy, 137(??):Article 104838, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004498. See [Mot20].

Magliozzi:2023:SPH

[MPD^{+}23] Chiara Magliozzi, Marco Palma, Jean-Noël Druon, Andreas Palialexis, McQuatters-Gollop Abigail, Varkitzi Ioanna, González-Quirós Rafael, Gorokhova Elena, Heyden Birgit, Boicenco Laura, and Artigas Luis Felipe. Status of pelagic habitats within the EU-Marine Strategy Framework Directive: Proposals for improving consistency and representativeness of the assessment. Marine Policy, 148(??):Article 105467, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22005188.

McLennan:2023:NMS

[MPH^{+}23] Shaun McLennan, Beth Polidoro, Tim Huntington, Rod Cappell, Victor Restrepo, Karin Mundnich, and Ernesto Jardim. The new Marine Stewardship Council requirements to improve ghost gear management: Insights from the policy development process. Marine Policy, 157(??):Article 105840, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003731.

Marques:2023:CEI

[MPL^{+}23] Crístofer H. Marques, Paula C. Pereda, Andrea Lucchesi, Ramiro F. Ramos, Olav Fiksdahl, Luiz F. Assis, Newton N. Pereira, and Jean-David Caprace. Cost and environmental impact assessment of mandatory speed reduction of maritime fleets. Marine Policy, 147(??):Article 105334, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003815.

Miret-Pastor:2020:TTD

[MPSF20] Lluís Miret-Pastor, Kristina Svels, and Richard Freeman. Towards territorial development in fisheries areas: a typology of projects funded by Fisheries Local Action Groups. Marine Policy, 119(??):Article 104111, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20304759.

Maxwell:2020:NTM

[MRD^{+}20] Kimberley H. Maxwell, Kelly Ratana, Kathryn K. Davies, Caine Taiapa, and Shaun Awatere. Navigating towards marine co-management with Indigenous communities on-board the Waka-Taurua. Marine Policy, 111(??):Article 103722, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19301605.

Mrdalo:2023:URF

[Mrd23] Zvonimir P. D. Mrdalo. User rights in the fisheries and salmon-farming of the Faroe Islands. Marine Policy, 157(??):Article 105841, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003743.

Ma:2023:RFB

[MRKM23] Deqiang Ma, Jonathan Rhodes, Carissa J. Klein, and Martine Maron. Redistribution of fishery benefits among commercial and recreational fishers caused by offsetting. Marine Policy, 158(??):Article 105881, December 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004141.

Mossop:2020:ADR

[MS20] Joanna Mossop and Clive Schofield. Adjacency and due regard: the role of coastal States in the BBNJ treaty. Marine Policy, 122(??):Article 103877, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307511.

Miranda:2021:WDS

[MS21] Fabiola V. Miranda and Wolfgang B. Stotz. Why do state interventions in artisanal fisheries often fail in Chile? Misalignment to the way of life of fishers. Marine Policy, 132(??):Article 104693, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003043.

Masud:2022:CMA

[MSBA22] Muhammad Mehedi Masud, Sharifah Muhairah Shahabudin, Angathevar Baskaran, and Rulia Akhtar. Comanagement approach to sustainable management of marine protected areas: the case of Malaysia. Marine Policy, 138(??):Article 105010, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000574.

Munoz:2020:CWC

[MSGAR20] Pablo Durán Muñoz, Mar Sacau, Ana García-Alegre, and Esther Román. Cold-water corals and deep-sea sponges bycatch mitigation: Dealing with groundfish survey data in the management of the northwest Atlantic Ocean high seas fisheries. Marine Policy, 116(??):Article 103712, June 2020.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308066.

Martin:2023:CNP

[MSO23] Angela Helen Martin, Maike Iris Esther Scheffold, and Bethan Christine O’Leary. Changing the narrative and perspective surrounding marine fish. Marine Policy, 156(??):Article 105806, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003391.

Markiewicz-Stanny:2023:PGO

[MSP23] Joanna Markiewicz-Stanny and Dorota Pyć. Protection of the global ocean rights. Marine Policy, 148 (??):Article 105468, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200519X.

Markiewicz-Stanny:2022:IMI

[MSSS22] Joanna Markiewicz-Stanny and Marta Szuniewicz-Stepień. Irregular migration and its impact on the marine environment: Legal aspects. Marine Policy, 144(??):Article 105216, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002639.

Moreno:2023:JFP

[MSZ^{+}23] Gala Moreno, Joaquín Salvador, Iker Zudaire, Jefferson Murua, Josep Lluís Pelegrí, Jon Uranga, Hilario Murua, Maitane Grande, Josu Santiago, and Victor Restrepo. The Jelly-FAD: a paradigm shift in the design of biodegradable Fish Aggregating Devices. Marine Policy, 147(??):Article 105352, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003992.

Molony:2020:MSS

[MT20] Brett W. Molony and Damian P. Thomson. Managing the super-shark: Recommendations to improve communication and debate about chondrichthyans. Marine Policy, 118(??):Article 103983, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301135.

McMullen:2023:MLS

[MTC^{+}23] Karly McMullen, Ana Tirapé, Paola Calle, Jessica Vandenberg, Omar Alvarado-Cadena, Yoshitaka Ota, Gustavo A. Domínguez, and Juan José Alava. Marine litter and social inequities entangle Ecuadorian mangrove communities: Perceptions of plastic pollution and well-being concerns in Puerto Hondo and Isla Santay, Ecuador. Marine Policy, 157(??):Article 105857, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003901.

Mendenhall:2023:SRS

[MTN23] Elizabeth Mendenhall, Rachel Tiller, and Elizabeth Nyman. The ship has reached the shore: the final session of the ‘Biodiversity Beyond National Jurisdiction’ negotiations. Marine Policy, 155(??):Article 105686, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002130.

Mukherjee:2021:BRF

[Muk21] Zinnia Mukherjee. Behavioral response of fishers to hypoxia and the distributional impact on harvest. Marine Policy, 133(??):Article 104732, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003432.

Macpherson:2021:HAR

[MUR^{+}21a] Elizabeth Macpherson, Stephen C. Urlich, Hamish G. Rennie, Adrienne Paul, Karen Fisher, Laura Braid, Jill Banwell, Julia Torres Ventura, and Eric Jorgensen. ‘Hooks’ and ‘Anchors’ for relational ecosystem-based marine management. Marine Policy, 130(??):Article 104561, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100172X.

Murray:2021:GAT

[Mur21b] Ruth Murray. A governance analysis of three MPAs in Belize: Conservation objectives compromised by tourism development priorities? Marine Policy, 127(??):Article 104243, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20308897.

Marco:2021:EMR

[MVR21a] Jorge Marco, Diego Valderrama, and Mario Rueda. Evaluating management reforms in a Colombian shrimp fishery using fisheries performance indicators. Marine Policy, 125(??):Article 104258, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309040.

Marco:2021:TBL

[MVR21b] Jorge Marco, Diego Valderrama, and Mario Rueda. Triple bottom line assessment for the historically underperforming Colombian queen conch fishery. Marine Policy, 125(??):Article 104257, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309039.

Martinell:2021:SEE

[MVSP^{+}21] Daniel Peñalosa Martinell, Francisco Javier Vergara-Solana, Marcelo Araneda Padilla, Germán Ponce Díaz, Annie Mejaes, Manuel María Varela Lafuente, and U. Rashid Sumaila. Social effects of energy subsidies and taxes on CO_{2} emissions: the case of Mexican aquaculture public policies. Marine Policy, 128(??):Article 104481, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000920.

Michael:2021:DEB

[MWB^{+}21] Pamela E. Michael, Chris Wilcox, Christophe Barbraud, Karine Delord, Michael Sumner, and Henri Weimerskirch. Dynamic enforcement of bycatch via reproductive value can increase theoretical efficiency. Marine Policy, 132(??):Article 104684, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002955.

Mason:2023:LKA

[MWM^{+}23a] Julia G. Mason, Sarah J. Weisberg, Janelle L. Morano, Richard J. Bell, Mark Fitchett, Roger B. Griffis, Elliott L. Hazen, William D. Heyman, Kirstin Holsman, Kristin M. Kleisner, Katie Westfall, Michele K. Conrad, Margaret Daly, Abigail S. Golden, Chris J. Harvey, Lisa A. Kerr, Gway Kirchner, Arielle Levine, Rebecca L. Lewison, Sean M. Lucey, Wendy Morrison, Brandon Muffley, Jameal F. Samhouri, Matthew Seeley, S. Kalei Shotwell, and Diana L. Stram. Linking knowledge and action for climate-ready fisheries: Emerging best practices across the US. Marine Policy, 155(??):Article 105758, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002919.

McLean:2023:CCC

[MWM^{+}23b] Mercedes McLean, Brian Warner, Robert Markham, Mibu Fischer, Jim Walker, Carissa Klein, Maia Hoeberechts, and Daniel C. Dunn. Connecting conservation culture: the importance of indigenous knowledge in conservation decisionmaking and resource management of migratory marine species. Marine Policy, 155(??):Article 105582, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001094.

Michaelis:2021:CES

[MWWS21] Adriane K. Michaelis, William C. Walton, Donald W. Webster, and L. Jen Shaffer. Cultural ecosystem services enabled through work with shellfish. Marine Policy, 132(??):Article 104689, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003006.

Miao:2021:CDI

[MX21] Dandan Miao and Zhihua Xue. The current developments and impact of land reclamation control in China. Marine Policy, 134(??):Article 104782, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003936.

Ma:2023:APP

[MXZ23] Deyi Ma, Qi Xu, and Jietao Zhou. The application of the precautionary principle in the decision of releasing Fukushima radioactive contaminated water into the sea. Marine Policy, 156(??):Article 105803, October 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003366.

Marentette:2022:MEI

[MZ22] Julie R. Marentette and Fan Zhang. From means to ends: Insights into the operationalization and evaluation of sustainable fisheries. Marine Policy, 141(??):Article 105087, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001348.

Mao:2024:TDO

[MZ24] Zhengkai Mao and Zhijun Zhang. Taking the “UN Decade of Ocean Science for Sustainable Development” as an opportunity to help build a “Community with a Shared Future between China and Pacific Island Countries”. Marine Policy, 159(??):Article 105943, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004761.

Murua:2023:LLF

[MZT^{+}23] Hilario Murua, Iker Zudaire, Mariana Tolotti, Jefferson Murua, Manuela Capello, Oihane C. Basurko, Iñigo Krug, Maitane Grande, Igor Arregui, Jon Uranga, Jose Maria Ferarios, Philippe Sabarros, Jon Ruiz, Yannick Baidai, Maria Lourdes Ramos, Jose Carlos Báez, Francisco Abascal, Haritz Arrizabalaga, Gala Moreno, Laurent Dagorn, and Josu Santiago. Lessons learnt from the first largescale biodegradable FAD research experiment to mitigate drifting FADs impacts on the ecosystem. Marine Policy, 148(??):Article 105394, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004419.

Nalley:2024:HIC

[Nal24] Eileen M. Nalley. Highlighting the impacts of ciguatera poisoning on fishers and fish consumers in Hawai’i to inform future management strategies. Marine Policy, 159(??):Article 105875, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004086.

Nielsen:2024:DFA

[NAN^{+}24] Max Nielsen, Anna Andersson, Rasmus Nielsen, Johan Blomquist, and Staffan Waldo. Does fishery activity affect local tourism? Evidence from Denmark. Marine Policy, 161(??):Article 106051, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000496.

Nicolodi:2021:CGI

[NAP^{+}21] João Luiz Nicolodi, Milton Lafourcade Asmus, Marcus Polette, Alexander Turra, Carlos Alberto Seifert, Fernanda Terra Stori, Deborah Campos Shinoda, Alexandre Mazzer, Vitor Alberto de Souza, and Rafael Kuster Gonçalves. Critical gaps in the implementation of Coastal Ecological and Economic Zoning persist after 30 years of the Brazilian coastal management policy. Marine Policy, 128(??):Article 104470, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21000816.

Nyiawung:2022:CSS

[NAS^{+}22] Richard A. Nyiawung, Raymond K. Ayilu, Neville N. Suh, Norbert N. Ngwang, Frederick Varnie, and Philip A. Loring. COVID-19 and small-scale fisheries in Africa: Impacts on livelihoods and the fish value chain in Cameroon and Liberia. Marine Policy, 141(??):Article 105104, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001518.

Nunes:2023:EIF

[NBB^{+}23] Diogo M. Nunes, Alan C. Bezerra, Wik M. S. Barros, Paulo V. N. Araújo, Ilka S. L. Branco-Nunes, Rafael A. Magris, Pedro H. C. Pereira, Iran C. Normande, Rafael S. L. Barboza, and Andrei T. C. Cardoso. Evidence of illegal fishing within the largest Brazilian coastal MPA: Turning a blind eye to the obvious. Marine Policy, 147(??):Article 105324, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003712.

Nguyen:2023:BAC

[NCD23] Son Nguyen, Peggy Shu-Ling Chen, and Yuquan Du. Blockchain adoption in container shipping: an empirical study on barriers, approaches, and recommendations. Marine Policy, 155(??):Article 105724, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002579.

Nakamura:2021:ULP

[NCE21] Julia Nakamura, Ratana Chuenpagdee, and Mostafa El Halimi. Unpacking legal and policy frameworks: a step ahead for implementing the Small-Scale Fisheries Guidelines. Marine Policy, 129(??):Article 104568, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001792.

Narayanan:2023:KFI

[NEF23] Saratkumar C. Narayanan, Gholam Reza Emad, and Jiangang Fei. Key factors impacting women seafarers’ participation in the evolving workplace: a qualitative exploration. Marine Policy, 148(??):Article 105407, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004547.

Nijamdeen:2023:UEI

[NEH^{+}23] T. W. G. F. Mafaziya Nijamdeen, Ngendahimana Ephrem, Jean Hugé, Kodikara Arachchilage Sunanda Kodikara, and Farid Dahdouh-Guebas. Understanding the ethnobiological importance of mangroves to coastal communities: a case study from Southern and North-western Sri Lanka. Marine Policy, 147(??):Article 105391, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004389.

Nguyen:2023:PSS

[NGA23] Ly Nguyen, Zhifeng Gao, and James L. Anderson. Perception shifts in seafood consumption in the United States. Marine Policy, 148(??):Article 105438, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004857.

Neves:2021:SRS

[NGP^{+}21] João Neves, Jean-Christophe Giger, Nuno Piçarra, Vasco Alves, and Joana Almeida. Social representations of sharks, perceived communality, and attitudinal and behavioral tendencies towards their conservation: an exploratory sequential mixed approach. Marine Policy, 132(??):Article 104660, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002712.

Neis:2022:FSD

[NGRM^{+}22] Barbara Neis, María Andrée López Gómez, Emily Reid-Musson, Brenda Greenslade, David Decker, Joel Finnis, and Christine Knott. Fishing safely during COVID-19 in Newfoundland and Labrador, Canada: Making it happen. Marine Policy, 145(??):Article 105281, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003281.

Nuez:2023:ACS

[NGT^{+}23] Ignasi Nuez, Ioannis Giovos, Francesco Tiralongo, Jaime Penadés-Suay, Ilija Cetkovic, Manfredi Di Lorenzo, Periklis Kleitou, Rigers Bakiu, Mohamed Nejmeddine Bradai, Sara A. A. Almabruk, Roxani Naasan Aga Spyridopoulou, Andréa Sabbio, and Manel Gazo. Assessing the current status of Hexanchus griseus in the Mediterranean Sea using local ecological knowledge. Marine Policy, 147(??):Article 105378, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004250.

Nguyen:2023:BRF

[Ngu23] Hong-Kong To Nguyen. Book review: Freedom of Navigation and the Law of the Sea: Warships, States and the Use of Force, Cameron Moore, in: Routledge Research on the Law of the Sea. Routledge Taylor & Francis, London and New York (2021), ISBN 978-1-138-38814-7. Marine Policy, 152(??):Article 105609, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001367.

Nakamura:2020:ABF

[NH20] Julia Nakamura and Fábio Hazin. Assessing the Brazilian federal fisheries law and policy in light of the Voluntary Guidelines for Securing Sustainable Small-scale fisheries. Marine Policy, 113(??):Article 103798, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306359.

Nielsen:2023:SPC

[NHA23] Rasmus Nielsen, Ayoe Hoff, and Jesper L. Andersen. Structural and productivity changes from introducing strong user rights in the Danish demersal fisheries. Marine Policy, 147(??):Article 105385, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004328.

Nguyen:2023:ELC

[NHD23] Hai Nguyen, Richard. J. Harper, and Bernard Dell. Examining local community understanding of mangrove carbon mitigation: a case study from Ca Mau province, Mekong River Delta, Vietnam. Marine Policy, 148 (??):Article 105398, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004456.

Niu:2022:BRL

[Niu22] Bingru Niu. Book review: The Law of the Sea and Climate Change — Solutions and Constraints, Elise Johansen, Signe Veierud Busch, Ingvild Ulrikke Jakobsen. Cambridge University Press (2021), 435 pp., ISBN 978-1-108-84226-6. Marine Policy, 137(??):Article 104941, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005522.

Noble-James:2023:MBH

[NJBM^{+}23] Tamsyn Noble-James, Ross Bullimore, Fionnuala McBreen, Joey O’Connor, James Highfield, Chris McCabe, Simeon Archer-Rand, Anna-Leena Downie, Jon Hawes, and Peter Mitchell. Monitoring benthic habitats in English Marine Protected Areas: Lessons learned, challenges and future directions. Marine Policy, 157(??):Article 105852, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003858.

Nam:2021:ICO

[NK21] Dong Nam and Mingyu Kim. Implication of COVID-19 outbreak on ship survey and certification. Marine Policy, 131(??):Article 104615, September 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002268.

Neimane:2023:MSP

[NM23] Leila Neimane and Séverine Michalak. Maritime spatial plans as an object of the right of access to information. Marine Policy, 158(??):Article 105870, December 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004037.

Navarro:2024:EUO

[NMCL24] Matt L. Navarro, Jacquomo Monk, Gabrielle H. Cummins, and Tim J. Langlois. Embracing uncertainty in ocean accounts. Marine Policy, 162(??):Article 106040, April 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000381.

Nielsen:2021:MBF

[NNM^{+}21] Pernille Nielsen, Mette Møller Nielsen, Ciaran McLaverty, Kasper Kristensen, Kerstin Geitner, Jeppe Olsen, Camille Saurel, and Jens Kjerulf Petersen. Management of bivalve fisheries in marine protected areas. Marine Policy, 124(??):Article 104357, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310083.

Nuttall:2021:PID

[NNR^{+}21] P. Nuttall, A. Newell, I. Rojon, B. Milligan, and A. Irvin. Pacific island domestic shipping emissions abatement measures and technology transition pathways for selected ship types. Marine Policy, 132(??):Article 104704, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003158.

Nguyen:2024:CSR

[NNRS24] Lan Anh Nguyen, Manh-Hung Nguyen, Arnaud Reynaud, and Michel Simioni. A comparative study of residents and tourists’ valuation for a heterogeneous environmental good: the case of coastal erosion. Marine Policy, 161(??):Article 106038, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000368.

Nakamura:2022:PTD

[NOB22] Katrina Nakamura, Yoshitaka Ota, and Francisco Blaha. A practical take on the duty to uphold human rights in seafood workplaces. Marine Policy, 135(??):Article 104844, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004553.

Nobrega:2023:STS

[NOL^{+}23] Marcelo Francisco Nóbrega, Matheus Assis Oliveira, Marcelo Gomes Lira, Suelen de Souza Rocha, and Jorge Eduardo Lins Oliveira. Sustainability of tunas and swordfish exploitation in the equatorial tropical Atlantic Ocean. Marine Policy, 155(??):Article 105755, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002889.

Nyiawung:2021:MSC

[NRF21] Richard A. Nyiawung, Ajith Raj, and Paul Foley. Marine Stewardship Council sustainability certification in developing countries: Certifiability and beyond in Kerala, India and The Gambia, West Africa. Marine Policy, 129(??):Article 104526, July 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001378.

Nasrudin:2020:CCT

[NRYG20] Rus’an Nasrudin, Budy P. Resosudarmo, Satoshi Yamazaki, and Wardis Girsang. Contribution of cash transfers in moderating household food insecurity in small-island communities: Experimental evidence from Indonesia. Marine Policy, 118(??):Article 104025, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307705.

Noack:2020:SUP

[NSK20] Thomas Noack, Esther Savina, and Junita D. Karlsen. Survival of undersized plaice (Pleuronectes platessa) discarded in the bottom otter trawl and Danish seine mixed fisheries in Skagerrak. Marine Policy, 115(??):Article 103852, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19301599.

Nisa:2022:WBW

[NSN22] Zahidah Afrin Nisa, Clive Schofield, and Francis C. Neat. Work Below Water: the role of scuba industry in realising sustainable development goals in small island developing states. Marine Policy, 136(??):Article 104918, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21005297.

Nunez-Sanchez:2020:GSL

[NSPRSS20] M. J. Nuñez-Sanchez, L. Perez-Rojas, L. Sciberras, and J. Ramos Silva. Grounds for a safety level approach in the development of long-lasting regulations based on costs to reduce fatalities for sustaining industrial fishing vessel fleets. Marine Policy, 113(??):Article 103806, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309977.

NSouvi:2021:FAT

[NSZ^{+}21] Kodjo N’Souvi, Chen Sun, Haiqing Zhang, Debi Ahoefa Broohm, and Mawussé Komlagan Nézan Okey. Fisheries and aquaculture in Togo: Overview, performance, fisheries policy, challenges and comparative study with Ghana, Mali, Niger and Senegal fisheries and aquaculture. Marine Policy, 132(??):Article 104681, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100292X.

Nyman:2020:TCR

[NT20] Elizabeth Nyman and Rachel Tiller. ‘Is there a court that rules them all’? Ocean disputes, forum shopping and the future of Svalbard. Marine Policy, 113(??):Article 103742, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18307401.

Nassiri:2021:HEC

[NTH^{+}21] A. Nassiri, O. Thébaud, S. J. Holbrook, M. Lauer, A. Rassweiler, R. J. Schmitt, and J. Claudet. Hedonic evaluation of coral reef fish prices on a direct sale market.

Marine Policy, 129(??):Article 104525, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001366.

Niella:2023:MYE

[NUD^{+}23] Yuri Niella, Vinay Udyawer, Michael Drew, Brett Simes, Hugh Pederson, and Charlie Huveneers. Multi-year effects of wildlife tourism on shark residency and implications for management. Marine Policy, 147(??):Article 105362, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004092.

Nielsen:2023:CIE

[NVP^{+}23] Rasmus Nielsen, Sebastian Villasante, Jose Manuel Fernandez Polanco, Jordi Guillen, Ignacio Llorente Garcia, and Frank Asche. The Covid-19 impacts on the European Union aquaculture sector. Marine Policy, 147(??):Article 105361, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004080.

Nwafor:2020:PBP

[NW20] Ndubuisi Nwafor and Tony R. Walker. Plastic Bags Prohibition Bill: a developing story of crass legalism aiming to reduce plastic marine pollution in Nigeria. Marine Policy, 120(??):Article 104160, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20303870.

Nie:2021:CER

[NWZ^{+}21] Xin Nie, Jianxian Wu, Wei Zhang, Jun Zhang, Wenhuan Wang, Yinghui Wang, Yaping Luo, and Han Wang. Can environmental regulation promote urban innovation in the underdeveloped coastal regions of western China? Marine Policy, 133(??):Article 104709, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003201.

Owusu:2020:OAR

[OA20] Victor Owusu and Edo Andriesse. From open access regime to closed fishing season: Lessons from small-scale coastal fisheries in the Western Region of Ghana. Marine Policy, 121(??):Article 104162, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309856.

Otumawu-Apreku:2024:FAP

[OAAMS24] Kofi Otumawu-Apreku, Transform Aqorao, Fiona Meke, and Kezyiah Lusa Saepioh. Fishing activities in Pacific Island countries: a human-rights perspective. Marine Policy, 161(??):Article 106008, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400006X.

Otumawu-Apreku:2020:PES

[OAM20] Kofi Otumawu-Apreku and Stephanie F. McWhinnie. Profit Efficiency of the South Australian rock lobster fishery: a Nerlovian and directional distance function approach. Marine Policy, 117(??):Article 103962, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305950.

Otumawu-Apreku:2021:RFC

[OAM21] Kofi Otumawu-Apreku and Stephanie F. McWhinnie. The role of fixed cost and non-discretionary variables in fisheries: the case of South Australian rock lobster. Marine Policy, 133(??):Article 104745, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003560.

Oanta:2022:AAS

[Oan22] Gabriela A. Oanta. The application of administrative sanctions in the fight against IUU fishing: an assessment of Spanish practice. Marine Policy, 144(??):Article 105211, October 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002585.

Osterman:2022:WBH

[ÖB22] Cecilia Österman and Magnus Boström. Workplace bullying and harassment at sea: a structured literature review. Marine Policy, 136(??):Article 104910, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005212.

Ojwala:2024:UWR

[OBNK24] Renis Auma Ojwala, Susan Buckingham, Francis Neat, and Momoko Kitada. Understanding women’s roles, experiences and barriers to participation in ocean science education in Kenya: recommendations for better gender equality policy. Marine Policy, 161(??):Article 106000, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300533X.

Obura:2020:GSE

[Obu20] David O. Obura. Getting to 2030 — scaling effort to ambition through a narrative model of the SDGs. Marine Policy, 117(??):Article 103973, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309340.

OHanlon:2020:CII

[OC20] Zoë O’Hanlon and Valerie Cummins. A comparative insight of Irish and Scottish regulatory frameworks for offshore wind energy — an expert perspective. Marine Policy, 117(??):Article 103934, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930418X.

Ortega:2023:IPD

[OCCSC23] Miquel Ortega, María D. Castro-Cadenas, Jeroen Steenbeek, and Marta Coll. Identifying and prioritizing demersal fisheries restricted areas based on combined ecological and fisheries criteria: the western Mediterranean. Marine Policy, 157(??):Article 105850, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003834.

OKeeffe:2020:SAC

[OCD^{+}20] Jane M. O’Keeffe, Valerie Cummins, Robert J. N. Devoy, Donald Lyons, and Jeremy Gault. Stakeholder awareness of climate adaptation in the commercial seaport sector: a case study from Ireland. Marine Policy, 111(??):Article 102404, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16302408.

Okyere:2023:CDM

[OCD^{+}23] Isaac Okyere, Ernest Obeng Chuku, Selorm Awiah Dzantor, Violet Ahenkorah, and Richard Adade. Capacity deficit and marginalisation of artisanal fishers hamper effective fisheries governance in Ghana: Insights and propositions for promoting sustainable small-scale fisheries. Marine Policy, 153(??):Article 105640, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001677.

Orobko:2022:PCC

[OCT^{+}22] Melissa Orobko, Gillian Curren, Thomas W. Therriault, Peter Lawton, Lucia Fanning, and Isabelle M. Côté. Perception and consideration of cumulative effects and multiple stressors by marine managers in Canada. Marine Policy, 143(??):Article 105122, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001695.

Obregon:2022:LFK

[OCZ^{+}22] Clara Obregón, Joseph Christensen, Dirk Zeller, Michael Hughes, James R. Tweedley, Andrea Gaynor, and Neil R. Loneragan. Local fisher knowledge reveals changes in size of blue swimmer crabs in small-scale fisheries. Marine Policy, 143(??):Article 105144, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001919.

ODonnell:2021:CLF

[O’D21] Tayanah O’Donnell. Coastal Lawscape: a framework for understanding the complexities of climate change adaptation. Marine Policy, 129(??):Article 104532, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001433.

OConnor:2023:PCP

[OEV23] Eamonn O’Connor, Natasha Evers, and Amaya Vega. Port capacity planning — a strategic management perspective. Marine Policy, 150(??):Article 105537, April 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000647.

OMahony:2020:IFC

[OGGC20] C. O’Mahony, S. Gray, J. Gault, and V. Cummins. ICZM as a framework for climate change adaptation action — Experience from Cork Harbour, Ireland. Marine Policy, 111(??):Article 102223, January 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002924.

Orchard:2024:RTR

[OGS24] Shane Orchard, Shawn Gerrity, and David R. Schiel. Re-thinking recreational fishing — how a natural disaster presents insights and opportunities for achieving sustainability and equity objectives. Marine Policy, 159(??):Article 105916, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004499.

Okubo:2023:PSE

[OI23] Ayako Okubo and Atsushi Ishii. Pursuing sustainability? ecosystem considerations in Japan’s fisheries governance. Marine Policy, 152(??):Article 105603, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001306.

Ozkaynak:2024:BSQ

[Öİ24] Ömer Harun Özkaynak and Gönül Tuğrul İçemer. Based on stakeholder questionnaires of ship-derived waste, waste management re-planning in ports and determining the need for inspection. Marine Policy, 159(??):Article 105912, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004451.

Ontomwa:2024:LPF

[OKFN24] Mary Binsari Ontomwa, Edward Ndirui Kimani, Bernerd Mulwa Fulanda, and Chrisphine Sangara Nyamweya. Longline pelagic fishery assemblage in Kenya’s exclusive economic zone marine waters. Marine Policy, 160(??):Article 105989, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23005225.

Obane:2024:AAS

[OKH^{+}24] Hideaki Obane, Kentaro Kazama, Hiroshi Hashimoto, Yu Nagai, and Kenji Asano. Assessing areas suitable for offshore wind energy considering potential risk to breeding seabirds in northern Japan. Marine Policy, 160(??):Article 105982, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23005158.

Ojwala:2022:EGP

[OKNB22] Renis Auma Ojwala, Momoko Kitada, Francis Neat, and Susan Buckingham. Effectiveness of gender policies in achieving gender equality in ocean science programmes in public universities in Kenya. Marine Policy, 144(??):Article 105237, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002846.

Orejas:2020:TCA

[OKR^{+}20] Covadonga Orejas, Ellen Kenchington, Jake Rice, Georgios Kazanidis, Andreas Palialexis, David Johnson, Matthew Gianni, Roberto Danovaro, and J. Murray Roberts. Towards a common approach to the assessment of the environmental status of deep-sea ecosystems in areas beyond national jurisdiction. Marine Policy, 121(??):Article 104182, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19309236.

Oldach:2022:MUW

[OKS^{+}22] Eliza Oldach, Helen Killeen, Priya Shukla, Ellie Brauer, Nicholas Carter, Jennifer Fields, Alexandra Thomsen, Cassidy Cooper, Leah Mellinger, Kaiwen Wang, Carl Hendrickson, Anna Neumann, Pernille Sporon Bøving, and Nann Fangue. Managed and unmanaged whale mortality in the California Current Ecosystem. Marine Policy, 140(??):Article 105039, June 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000860.

Oksas:2023:CES

[Okş23] Olgay Okşaş. Carbon emission strategies for container handling equipment using the activity-based method: a case study of Ambarlı container port in Turkiye. Marine Policy, 149(??):Article 105480, March 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000076.

Ourens:2022:LSS

[OMC^{+}22] Rosana Ouréns, Michael C. Melnychuk, Larry B. Crowder, Nicolas L. Gutierrez, Ray Hilborn, Cristina Pita, and Omar Defeo. Linking small-scale fisheries performance to governance attributes: a quantitative assessment from stakeholders’ perceptions in the Americas and Europe. Marine Policy, 136(??):Article 104876, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004875.

Obane:2021:APA

[ONA21] Hideaki Obane, Yu Nagai, and Kenji Asano. Assessing the potential areas for developing offshore wind energy in Japanese territorial waters considering national zoning and possible social conflicts. Marine Policy, 129(??):Article 104514, July 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001251.

Olsen:2020:MLI

[ONN^{+}20] Julia Olsen, Leticia Antunes Nogueira, Anne Katrine Normann, Bjørn Vidar Vangelsten, and Ingrid Bay-Larsen. Marine litter: Institutionalization of attitudes and practices among fishers in Northern Norway. Marine Policy, 121(??):Article 104211, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20308575.

Osmundsen:2022:APD

[OOGA22] Tonje C. Osmundsen, Marit Schei Olsen, Asle Gauteplass, and Frank Asche. Aquaculture policy: Designing licenses for environmental regulation. Marine Policy, 138(??):Article 104978, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000252.

Okeke-Ogbuafor:2024:CSF

[OOGK^{+}24] Nwamaka Okeke-Ogbuafor, Tim Gray, Sheku Kamara, Edward Sesay, Abdulai Dauda, Selina M Stead, Danielle Robinson, and Kelechi Johnmary Ani. Climate-smart fisheries: CO_{2} emissions reduction and food security are complementary. Marine Policy, 159(??):Article 105926, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004591.

Okeke-Ogbuafor:2020:CRF

[OOGS20a] Nwamaka Okeke-Ogbuafor, Tim Gray, and Selina Marguerite Stead. The controversial role of foreign fisheries consultants in Sierra Leone’s coastal waters. Marine Policy, 118(??):Article 103399, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307036.

Okeke-Ogbuafor:2020:TWP

[OOGS20b] Nwamaka Okeke-Ogbuafor, Tim Gray, and Selina Marguerite Stead. Is there a ‘wicked problem’ of smallscale coastal fisheries in Sierra Leone? Marine Policy, 118(??):Article 103471, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18306341.

Okeke-Ogbuafor:2021:IAP

[OOSG21] Nwamaka Okeke-Ogbuafor, Selina Stead, and Tim Gray. Is inland aquaculture the panacea for Sierra Leone’s decline in marine fish stocks? Marine Policy, 132(??):Article 104663, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002748.

OHagan:2020:ATW

[OPL20] Anne Marie O’Hagan, Shona Paterson, and Martin Le Tissier. Addressing the tangled web of governance mechanisms for land-sea interactions: Assessing implementation challenges across scales. Marine Policy, 112(??):Article 103715, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18309163.

Ofori:2021:MIF

[OR21] Roland O. Ofori and Mark D. Rouleau. Modeling the impacts of floating seaweeds on fisheries sustainability in Ghana. Marine Policy, 127(??):Article 104427, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000373.

Oral:2023:USL

[Ora23] Nilufer Oral. UNCLOS and sea level rise. Marine Policy, 149(??):Article 105454, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22005012.

Olaya-Restrepo:2022:MNA

[ORSB22] Julián Olaya-Restrepo, Alexandre Schiavetti, and Marcos S. Barbeitos. A multilayered network analysis of social participation in the management of Marine Protected Areas in Brazil. Marine Policy, 146(??):Article 105329, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003761.

Osmundsen:2023:PRF

[Osm23] Lori Osmundsen. Port reception facilities and a regional approach: a bridge for abating plastic pollution in the Arctic? Marine Policy, 148(??):Article 105436, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004833.

Osthagen:2020:MBD

[Øst20] Andreas Østhagen. Maritime boundary disputes: What are they and why do they matter? Marine Policy, 120(??):Article 104118, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20302426.

Ovegaard:2021:MAA

[OSV21] Mikael Ovegård, Andreas Sundelöf, and Daniel Valentinsson. Much ado about nothing: an example of how failed incentives thwarted the implementation of the EU landing obligation. Marine Policy, 123(??):Article 104305, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309507.

Oposa:2023:RSC

[OT23] Anna R. Oposa and Erika J. Techera. A review of shark conservation and management legal frameworks in the Philippines. Marine Policy, 155(??):Article 105713, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002464.

Ostwald:2021:MTS

[OTK^{+}21] Ashlea Ostwald, Vivitskaia J. D. Tulloch, Peter M. Kyne, Nicholas J. Bax, Piers K. Dunstan, Luciana C. Ferreira, Michele Thums, Judy Upston, and Vanessa M. Adams. Mapping threats to species: Method matters. Marine Policy, 131(??):Article 104614, September 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002256.

Ovalle:2022:UDL

[OVA22] Juan Carlos Ovalle, Carlos Vilas, and Luís T. Antelo. On the use of deep learning for fish species recognition and quantification on board fishing vessels. Marine Policy, 139(??):Article 105015, May 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000628.

Ostale-Valriberas:2022:AMM

[OVSVPP^{+}22] E. Ostalé-Valriberas, J. Sempere-Valverde, A. Pavón-Paneque, S. Coppa, F. Espinosa, and J. C. García-Gómez. Artificial marine micro-reserves as a new ecosystem-based management tool for marine conservation: the case of Patella ferruginea (Gastropoda, Patellidae), one of the most endangered marine invertebrates of the Mediterranean. Marine Policy, 136(??):Article 104917, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21005285.

Ortiz-Wolford:2021:CSR

[OWCFDC21] Jenniffer Ortiz-Wolford, M. Fabiola Corona-Figueroa, Vanessa Dávila, and Andrea A. Cabrera. Cetacean stranding records along the Pacific coastline of Guatemala, 2007– 2021: Implications for management, conservation and research. Marine Policy, 134(??):Article 104827, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004383.

Odhiambo:2020:EME

[OWS20] Jacob Ochiewo Odhiambo, Joseph Wakibia, and Maurice M. Sakwa. Effects of monitoring and evaluation planning on implementation of poverty alleviation mariculture projects in the coast of Kenya. Marine Policy, 119(??):Article 104050, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19309595.

Okafor-Yarwood:2024:TMS

[OYECdOP24] Ifesinachi Okafor-Yarwood, Oliver Eastwood, Noleen Chikowore, and Lucas de Oliveira Paes. Technology and maritime security in Africa: Opportunities and challenges in Gulf of Guinea. Marine Policy, 160(??):Article 105976, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23005092.

Okafor-Yarwood:2022:SRF

[OYKBA22] Ifesinachi Okafor-Yarwood, Nelly I. Kadagi, Dyhia Belhabib, and Edward H. Allison. Survival of the Richest, not the Fittest: How attempts to improve governance impact African small-scale marine fisheries. Marine Policy, 135(??):Article 104847, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004589.

Pazir:2022:ECP

[PAK22] Mohammad Khalil Pazir, Amirhossein Ahmadi, and Parisa Hossein Khezri. The effect of COVID-19 pandemic on the shrimp industry of Iran. Marine Policy, 136(??):Article 104900, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2100511X.

Pomeroy:2020:FIB

[PALB20] Robert Pomeroy, Carlos Arango, Cristopher G. Lomboy, and Steve Box. Financial inclusion to build economic resilience in small-scale fisheries. Marine Policy, 118(??):Article 103982, August 2020. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19306645.

Patron:2022:CPI

[Pat22] Jose Manuel Sánchez Patrón. The criminal prosecution of illegal fishing and the jurisdiction of Spanish courts. Marine Policy, 143(??):Article 105209, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002561.

Pauksztat:2021:IRC

[Pau21] Birgit Pauksztat. Informal relations and communication about work-related problems in two multilingual crews. Marine Policy, 130(??):Article 103767, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305639.

Patterson:2020:ARM

[PB20] Douglas A. Patterson and Raj Bridgelall. Attack risk modelling for the San Diego maritime facilities. Marine Policy, 121(??):Article 104210, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308563.

Pecarevic:2023:ICC

[PBD^{+}23] Marijana Pećarević, Kruno Bonačić, Tatjana Dobroslavić, Leon Grubišić, Igor Čelić, Branko Glamuzina, Josip Maleš, Gianluca Sarà, and Maria Cristina Mangano. Impact of COVID-19 on Croatian mariculture: Findings from the first national surveys. Marine Policy, 148(??):Article 105395, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22004420.

Poeta:2022:OYA

[PBG^{+}22] G. Poeta, M. Bazzichetto, L. Gallitelli, M. Garzia, F. Aprea, F. Bartoli, C. Battisti, S. Cascone, A. Corradi, D. D’Amelia, E. D’Amico, J. De Luca, F. Del Grosso, L. Iacobelli, S. Langone, C. Lembo Fazio, G. Locchi, M. Perrone, F. Petroni, D. Raimondi, F. Romiti, S. Secco, L. Sonet, A. Spinelli, S. Toscano, S. Vanadia, S. Vecchi, F. Zanon, and M. Malavasi. One year after on Tyrrhenian coasts: the ban of cotton buds does not reduce their dominance in beach litter composition. Marine Policy, 143(??):Article 105195, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002421.

Patrizzi:2023:MBC

[PBG23] Nayla S. Patrizzi, Francisco Barros, and Vinicius J. Giglio. Media bias on communicating conservation interventions: the case of large-scale marine protected areas in the South Atlantic. Marine Policy, 149(??):Article 105475, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000027.

Procksch:2023:HUS

[PBH^{+}23] Natália Procksch, Natalia Bragiola Berchieri, Rafael Kenji Horota, Vinícius Sales, Paulo Henrique Ott, Daniel Danilewicz, Taina Thomassim Guimaraes, Murilo Guimarães, Maurício Roberto Veronez, and Larissa Rosa de Oliveira. Habitat use by South American fur seals (Arctocephalus australis) and sea lions (Otaria flavescens) in a marine protected area in southern Brazil. Marine Policy, 155(??):Article 105693, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002208.

Pelage:2023:FPG

[PBS^{+}23] L. Pelage, A. Bertrand, S. C. W. Siqueira, A. C. A. P. Araújo, K. V. A. Avelino, C. Lopes da Silva, and T. Frédou. Fishers’ perceptions of global change to inform coastal planning in a data-poor socio-ecological system. Marine Policy, 155(??):Article 105784, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003172.

Pennino:2020:DBS

[PBT^{+}20] Maria Grazia Pennino, Ana Helena Bevilacqua, M. Angeles Torres, Jose M. Bellido, Jordi Sole, Jeroen Steenbeek, and Marta Coll. Discard ban: a simulation-based approach combining hierarchical Bayesian and food web spatial models. Marine Policy, 116(??):Article 103703, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307954.

Portman:2020:ABM

[PC20] Michelle E. Portman and André Zhulpa Camporesi. Attitudes and behaviours of marine recreationists towards conservation and environmental protection: a case study of Tel Aviv, Israel. Marine Policy, 122(??):Article 104133, December 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X1930884X.

Palm:2021:MLF

[PCA21] Karl E. Palm, Gary A. Campbell, and Jenny L. Apriesnig. Management of local fisheries: a case study of Laoang, Northern Samar, Philippines. Marine Policy, 132(??):Article 104657, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002682.

Pereira:2023:LIE

[PCA^{+}23] Laura M. Pereira, Guillermo Ortuño Crespo, Diva J. Amon, Renuka Badhe, Salomão Bandeira, Frida Bengtsson, Miranda Boettcher, Gabrielle Carmine, William W. L. Cheung, Bwalya Chibwe, Daniel Dunn, Maria A. Gasalla, Ghassen Halouani, David E. Johnson, Jean-Baptiste Jouffray, Silvana Juri, Patrick W. Keys, Hannah M. Lübker, Andrew S. Merrie, Farah Obaidullah, Juliano Palacios-Abrantes, Lynne J. Shannon, U. Rashid Sumaila, Edoardo Superchi, Naomi Terry, Colette C. C. Wabnitz, Moriaki Yasuhara, and Wei Zhou. The living infinite: Envisioning futures for transformed human-nature relationships on the high seas. Marine Policy, 153(??):Article 105644, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001719.

Portilla-Cabrera:2024:POA

[PCCOS24] Cristiam Victoriano Portilla-Cabrera, Maria Alejandra Cifuentes-Ossa, and John Josephraj Selvaraj. Policyoriented adaptation strategies for artisanal marine fisheries under climate and non-climate stressors in San Andres de Tumaco, Nariño, Colombia. Marine Policy, 161(??):Article 105993, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23005262.

Pascoe:2023:UDE

[PCD^{+}23] Sean Pascoe, Toni Cannard, Natalie A. Dowling, Catherine M. Dichmont, Frank Asche, and L. Richard Little. Use of Data Envelopment Analysis (DEA) to assess management alternatives in the presence of multiple objectives. Marine Policy, 148(??):Article 105444, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004912.

Paiva:2023:MCM

[PCG^{+}23] Sandra Vieira Paiva, Pedro Bastos Macedo Carneiro, Tatiane Martins Garcia, Tallita Cruz Lopes Tavares, Lidriana de Souza Pinheiro, Antonio Rodrigues Ximenes Neto, Tarin Cristino Montalverne, and Marcelo O. Soares. Marine carbonate mining in the Southwestern Atlantic: current status, potential impacts, and conservation actions. Marine Policy, 148(??):Article 105435, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004821.

Pierucci:2022:GRM

[PCK22] Andrea Pierucci, Silvia Columbu, and Laurence T. Kell. A global review of MSC certification: Why fisheries withdraw? Marine Policy, 143(??):Article 105124, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001713.

Prieto-Carolino:2021:GAP

[PCSS^{+}21] Alice Prieto-Carolino, Ida M. Siason, Marieta B. Sumagaysay, Rowena Paz L. Gelvezon, Harold M. Monteclaro, and Rosario H. Asong. A gender analysis of the processing sector of the tuna value chain in General Santos City, Philippines. Marine Policy, 128(??):Article 104477, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000889.

Paramana:2023:ACB

[PDB^{+}23] Th. Paramana, M. Dassenakis, N. Bassan, C. Dallangelo, P. Campostrini, S. Raicevich, F. Ronchi, G. Giorgi, A. Murillas-Maza, M. C. Uyarra, N. Papadopoulou, C. Smith, K. Jarni, Š. Koren Bačovnik, K. Klančnik, M. Pavičić, S. Skejić, O. Vidjak, J. F. Cadiou, L. López-López, I. Alvarez, L. Giannoudi, N. Streftaris, and P. Pagkou. Achieving coherence between the marine strategy framework directive and the maritime spatial planning directive. Marine Policy, 155(??):Article 105733, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2300266X.

Priyashadi:2022:SED

[PDJ22] M. S. V. H. Priyashadi, K. H. M. Ashoka Deepananda, and Asanka Jayasinghe. Socio-economic development of marine ornamental reef fish fishers in eastern Sri Lanka through the lenses of Human Development Index. Marine Policy, 143(??):Article 105136, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200183X.

Penca:2022:SPI

[Pen22] Jerneja Penca. Science, precaution and innovation for sustainable fisheries: the judgement by the Court of Justice of the EU regarding the electric pulse fishing ban. Marine Policy, 135(??):Article 104864, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004759.

Petersson:2022:TGF

[Pet22] Matilda T. Petersson. Transparency in global fisheries governance: the role of non-governmental organizations. Marine Policy, 136(??):Article 104128, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306682.

Parlee:2021:FSS

[PFG^{+}21] Courtenay E. Parlee, Paul Foley, María Andrée López Gómez, Md. Ruyel Miah, Charles Mather, and Robert L. Stephenson. Full spectrum sustainability and a theory of access: Integrating social benefits into fisheries governance. Marine Policy, 134(??):Article 104764, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003754.

Pafi:2020:CTM

[PFM20] Maria Pafi, Wesley Flannery, and Brendan Murtagh. Coastal tourism, market segmentation and contested landscapes. Marine Policy, 121(??):Article 104189, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20300774.

Purwanto:2023:MGB

[PFWM23] Purwanto, Erik C. Franklin, Alan White, and Ses Rini Mardiani. Multiple-goal bioeconomic programming to address conflicting management objectives in Indonesian small pelagic fisheries. Marine Policy, 150(??):Article 105519, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000465.

Pinarbasi:2020:KIT

[PGA^{+}20] Kemal Pınarbaşı, Ibon Galparsoro, Neil Alloncle, Frédéric Quemmerais, and Ángel Borja. Key issues for a transboundary and ecosystem-based maritime spatial planning in the Bay of Biscay. Marine Policy, 120(??):Article 104131, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20303298.

Pincinato:2022:MIS

[PGGA22] Ruth Beatriz Mezzalira Pincinato, Maria A. Gasalla, Taryn Garlock, and James L. Anderson. Market incentives for shark fisheries. Marine Policy, 139(??):Article 105031, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000781.

Pauksztat:2022:ICP

[PGK22] Birgit Pauksztat, Michelle R. Grech, and Momoko Kitada. The impact of the COVID-19 pandemic on seafarers’ mental health and chronic fatigue: Beneficial effects of onboard peer support, external support and Internet access. Marine Policy, 137(??):Article 104942, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005534.

Phayal:2022:IHM

[PGP22] Anup Phayal, Aaron Gold, and Brandon Prins. Interstate hostility and maritime crime: Evidence from South East Asia. Marine Policy, 143(??):Article 105134, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001816.

Pedroza-Gutierrez:2024:WFE

[PGSFRMHH24] Carmen Pedroza-Gutiérrez, Neyra Solano, Francisco Fernández-Rivera-Melo, and Isis Hernández-Herrera. What are the factors that enable or hinder gender equality? Lessons and experiences on gender dynamics in Mexican small-scale fisheries. Marine Policy, 159(??):Article 105944, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004773.

Phelan:2022:TMP

[PGSS22] Kieran Phelan, Alison Gardner, Elizabeth R. Selig, and Jessica L. Decker Sparks. Towards a model of portbased resilience against fisher labour exploitation. Marine Policy, 142(??):Article 105108, August 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001555.

Pham:2020:TMP

[Pha20] Thuy Thi Thanh Pham. Tourism in marine protected areas: Can it be considered as an alternative livelihood for local communities? Marine Policy, 115(??):Article 103891, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19303690.

Palomo:2020:ISM

[PHF20] Leopoldo E. Palomo and Alvaro Hernández-Flores. Integrating a spatial model and decision theory towards optimal boating density and carrying capacity in a recreational fishery. Marine Policy, 112(??):Article 103740, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305967.

Pietkiewicz:2021:LSC

[Pie21] Michał Pietkiewicz. Legal status of Caspian Sea — problem solved? Marine Policy, 123(??):Article 104321, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309684.

Pike:2024:IGF

[PJN24] Felicity Pike, Narriman S. Jiddawi, and Lina Mtwana Nordlund. Intertidal gleaning fisheries: Recognising localscale contributions and management scenarios. Marine Policy, 162(??):Article 106059, April 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000575.

Pineiro:2020:SHW

[PK20a] Laura Carballo Piñeiro and Momoko Kitada. Sexual harassment and women seafarers: the role of laws and policies to ensure occupational safety & health. Marine Policy, 117(??):Article 103938, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930106X.

Pratomo:2020:GAM

[PK20b] Eddy Pratomo and Jonathan Kwik. Good agreements make good neighbours: Settlements on maritime boundary disputes in South East Asia. Marine Policy, 117(??):Article 103943, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19308929.

Pentz:2023:WCC

[PK23] Brian Pentz and Nicole Klenk. Will climate change degrade the efficacy of marine resource management policies? Marine Policy, 148(??):Article 105462, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22005097.

Palialexis:2021:MBE

[PKB^{+}21] Andreas Palialexis, Vasiliki Kousteni, Laura Boicenco, Lisette Enserink, Kalliopi Pagou, Ulla Li Zweifel, Francesca Somma, Anna Cheilari, and David Connor. Monitoring biodiversity for the EU Marine Strategy Framework Directive: Lessons learnt from evaluating the official reports. Marine Policy, 128(??):Article 104473, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000841.

Patra:2023:MLM

[PKB^{+}23] Sivaji Patra, Muhammad Khurshid, Abas Basir, Pravakar Mishra, and M. V. Ramanamurthy. Marine litter management: a sustainable action plan and recommendations for the South Asian Seas region. Marine Policy, 157(??):Article 105854, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003871.

Park:2023:AMI

[PKC23] Sangwon Park, Chang Hoon Kim, and Minsoo Cheon. Analysis of marine incidents and policy implications: Case study of the South Korea. Marine Policy, 155(??):Article 105736, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002695.

Parviainen:2022:ESP

[PKH22] T. Parviainen, S. Kuikka, and P. Haapasaari. Enhancing science-policy interface in marine environmental governance: Oil spill response models as boundary objects in the Gulf of Finland, Baltic Sea. Marine Policy, 135(??):Article 104863, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004747.

Puts:2023:TOB

[PKMT23] Miriam Püts, Alexander Kempf, Christian Möllmann, and Marc Taylor. Trade-offs between fisheries, offshore wind farms and marine protected areas in the southern North Sea — winners, losers and effective spatial management. Marine Policy, 152(??):Article 105574, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300101X.

Park:2022:EEP

[PKR^{+}22] Sunghwa Park, Taeil Kim, Heeyoung Ryu, Hanna Kim, and Janghan Kwon. The economic effect and policy performance of ship registration — a case of Korea. Marine Policy, 144(??):Article 105234, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002810.

Psuty:2020:ISS

[PKS20] Iwona Psuty, Tomasz Kulikowski, and Lena Szymanek. Integrating small-scale fisheries into Polish maritime spatial planning. Marine Policy, 120(??):Article 104116, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19306979.

Pauly:2020:FSC

[PL20a] Daniel Pauly and Cui Liang. The fisheries of the South China Sea: Major trends since 1950. Marine Policy, 121(??):Article 103584, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302015.

Pellowe:2020:IBF

[PL20b] Kara E. Pellowe and Heather M. Leslie. The interplay between formal and informal institutions and the potential for co-management in a Mexican small-scale fishery. Marine Policy, 121(??):Article 104179, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930781X.

Pinn:2023:IEL

[PL23] Eunice H. Pinn and Lewis Le Vay. Interpretation of the European legal framework for the microbiological classification of bivalve mollusc production areas. Marine Policy, 148(??):Article 105479, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000064.

PARK:2024:DMF

[PL24a] Eon Kyung PARK and Seokwoo LEE. Discharge measures for Fukushima wastewater and South Korea’s response and implications of the WTO Korea–Japan fishery dispute (Korea-radionuclides) case. Marine Policy, 159(??):Article 105930, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004633.

Petrossian:2024:PTE

[PL24b] Gohar A. Petrossian and Jess Lettieri. A precautionary tale: Exploring the risks of deep-sea mining. Marine Policy, 162(??):Article 106073, April 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400071X.

Powell:2024:FPM

[PLOG24] Farrah Powell, Arielle Levine, and Lucia Ordonez-Gauger. Fishermen’s perceptions of management in the California spiny lobster and California market squid fisheries. Marine Policy, 161(??):Article 106015, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000137.

Perez-Martin:2020:LMV

[PMAVM20] Ricardo I. Pérez-Martín, Luis T. Antelo, J. A. Vázquez, and Jesús Mirón. An on-land management and valorisation approach for biomass associated with landing obligation compliance. Marine Policy, 116(??):Article 103506, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308443.

Prado:2024:PSP

[PMC24] Deborah Santos Prado, Ivan Machado Martins, and Ronaldo Adriano Christofoletti. Participatory scenario planning as a useful method for transforming ocean conflicts: Insights from a small-scale fishing conflict in São Paulo coast, Brazil. Marine Policy, 160(??):Article 105922, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004554.

Puentes:2022:SMB

[PMFR^{+}22] Vladimir Puentes, Paola A. Mejía-Falla, John G. Ramirez, Luis M. Manjarrés-Martínez, Juan M. Rguez-Baron, Luis A. Zapata, José Tavera, Fabio Gómez-Delgado, Carlos G. Barreto, Emiliano Zambrano, Ángel A. Villa, and Andrés F. Navia. Sharks and marine batoids management in Colombia: Policy instruments, management duty and implications for their populations and stakeholders. Marine Policy, 145(??):Article 105264, November 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003116.

Paolacci:2021:LSP

[PMK^{+}21] Simona Paolacci, Rogério Mendes, Regina Klapper, Amaya Velasco, Graciela Ramilo-Fernandez, Marta Muñoz-Colmenero, Tavis Potts, Sandra Martins, Solene Avignon, Julie Maguire, Enrique De Paz, Martin Johnson, Françoise Denis, Miguel A. Pardo, Dee McElligott, and Carmen Gonzalez Sotelo. Labels on seafood products in different European countries and their compliance to EU legislation. Marine Policy, 134(??):Article 104810, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004218.

Prasetyo:2021:SRT

[PMM^{+}21] Andhika P. Prasetyo, Allan D. McDevitt, Joanna M. Murray, Jon Barry, Firdaus Agung, Efin Muttaqin, and Stefano Mariani. Shark and ray trade in and out of Indonesia: Addressing knowledge gaps on the path to sustainability. Marine Policy, 133(??):Article 104714, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003250.

Perkiss:2022:CCA

[PMN^{+}22] Stephanie Perkiss, Alistair McIlgorm, Rachel Nichols, Anna R. Lewis, Kirti K. Lal, and Michelle Voyer. Can critical accounting perspectives contribute to the development of ocean accounting and ocean governance? Marine Policy, 136(??):Article 104901, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005121.

Pinillos:2023:DCZ

[PMR23] Francisco José Vázquez Pinillos, Juan Manuel Barragán Muñoz, and Francisco Ther Ríos. Delimiting coastal zones for integrated management: the case of the island and the sea of Chiloé (Chile). Marine Policy, 150(??):Article 105535, April 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000623.

Provot:2020:UQM

[PMT^{+}20] Z. Provot, S. Mahévas, L. Tissière, C. Michel, S. Lehuta, and B. Trouillet. Using a quantitative model for participatory geo-foresight: ISIS-Fish and fishing governance in the Bay of Biscay. Marine Policy, 117(??):Article 103231, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1630714X.

Partelow:2020:SNC

[PN20] Stefan Partelow and Katherine Nelson. Social networks, collective action and the evolution of governance for sustainable tourism on the Gili Islands, Indonesia. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X17308734.

Phillips:2023:OSF

[PNS^{+}23] Natasha Phillips, Marianne Nyegaard, Etsuro Sawai, Ching-Tsun Chang, Miguel Baptista, and Tierney Thys. The ocean sunfishes (family Molidea): Recommendations from the IUCN molidae review panel. Marine Policy, 155(??):Article 105760, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002932.

Polejack:2023:DOS

[Pol23] A. Polejack. The UN Decade of Ocean Science stages of grief — skepticism, frustration, fear of failure, and hope. Marine Policy, 152(??):Article 105597, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001240.

Pethiyagoda:2024:MSP

[POM24] N. A. Pethiyagoda, E. M. Ogier, and D. Hatton Mac-Donald. Measuring stakeholder preferences for ecosystem services interacting with salmon aquaculture in Tasmania under expansion: Application of Q methodology. Marine Policy, 162(??):Article 106053, April 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000514.

POONNAWATT:2023:MCA

[POO23] Kornwika POONNAWATT. Multilateral cooperation against maritime piracy in the Straits of Malacca: From the RMSI to ReCAAP. Marine Policy, 152(??):Article 105628, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001550.

Prodhan:2023:FIF

[PPK23] Md. Masudul Haque Prodhan, Md. Salauddin Palash, and Md. Akhtaruzzaman Khan. Factors influencing the financial and relational governance in Bangladesh’s inland fishing industry: an empirical study. Marine Policy, 148(??):Article 105446, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004936.

Prasada:2020:APC

[Pra20] D. V. P. Prasada. Assessing the potential for closure rules in the Sri Lankan sea cucumber fishery: Empirical models of practices and preferences of fishers. Marine Policy, 120(??):Article 104130, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20303006.

Praveena:2024:EPA

[Pra24] Sarva Mangala Praveena. Exploring public awareness, influencing factors and policy implications towards microplastic pollution: Perspectives from Malaysia. Marine Policy, 161(??):Article 106042, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400040X.

Prehn:2021:CSB

[Pre21] Michael Prehn. Climate strategy in the balance who decides? Marine Policy, 131(??):Article 104621, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002323.

Prip:2022:AOG

[Pri22] Christian Prip. Arctic Ocean governance in light of an of an international legally binding instrument on the conservation and sustainable use of marine biodiversity of areas beyond national jurisdiction. Marine Policy, 142(??):Article 103768, August 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19304762.

Parappurathu:2020:HAP

[PRM^{+}20] Shinoj Parappurathu, C. Ramachandran, Muktha Menon, K. K. Baiju, Prathibha Rohit, R. Narayana Kumar, Shelton Padua, and Shiv Kumar. Harnessing artisanal prowess in offshore fisheries: the case of Thoothoor fishers from India. Marine Policy, 121(??):Article 104174, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030049X.

Pinkau:2021:SEA

[PS21] Anne Pinkau and Kerstin S. Schiele. Strategic Environmental Assessment in marine spatial planning of the North Sea and the Baltic Sea — an implementation tool for an ecosystem-based approach? Marine Policy, 130(??):Article 104547, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001585.

Puszkarski:2022:IIS

[PŚ22] Jakub Puszkarski and Olga Śniadach. Instruments to implement sustainable aquaculture in the European Union. Marine Policy, 144(??):Article 105215, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002627.

Pace:2023:EFR

[PSD23] Lisa A. Pace, Ozcan Saritas, and Alan Deidun. Exploring future research and innovation directions for a sustainable blue economy. Marine Policy, 148(??):Article 105433, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22004808.

Petrossian:2020:FSE

[PSMR20] Gohar A. Petrossian, Monique Sosnowski, Dana Miller, and Diba Rouzbahani. Flags for sale: an empirical assessment of flag of convenience desirability to foreign vessels. Marine Policy, 116(??):Article 103937, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306372.

Partelow:2020:JSS

[PSPR20] Stefan Partelow, Tarsila Seara, Richard B. Pollnac, and Victor Ruiz. Job satisfaction in small-scale fisheries: Comparing differences between Costa Rica, Puerto Rico and the Dominican Republic. Marine Policy, 117(??):Article 103949, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1930898X.

Pikesley:2021:BBF

[PSTW21] Stephen K. Pikesley, Jean-Luc Solandt, Colin Trundle, and Matthew J. Witt. Benefits beyond ‘features’: Cooperative monitoring highlights MPA value for enhanced seabed integrity. Marine Policy, 134(??):Article 104801, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004127.

Pearson:2023:STF

[PT23] Natali Pearson and Benjamin S. Thompson. Saving two fish with one wreck: Maximizing synergies in marine biodiversity conservation and underwater cultural heritage protection. Marine Policy, 152(??):Article 105613, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001409.

Puharinen:2023:AGM

[Puh23] Suvi-Tuuli Puharinen. Achieving good marine environmental status in the EU — implications of the marine strategy framework directive for member states and blue economic activities. Marine Policy, 155(??):Article 105712, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002452.

Prellezo:2023:ESI

[PV23] Raúl Prellezo and Sebastian Villasante. Economic and social impacts of the landing obligation of the European Common Fisheries Policy: a review. Marine Policy, 148(??):Article 105437, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004845.

Pedreschi:2021:ITC

[PVH^{+}21] Debbi Pedreschi, Audric Vigier, Hannes Höffle, Sarah B. M. Kraak, and Dave G. Reid. Innovation through consultation: Stakeholder perceptions of a novel fisheries management system reveal flexible approach to solving fisheries challenges. Marine Policy, 124(??):Article 104337, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309878.

Perng:2023:IST

[PWL^{+}23a] Lansing Y. Perng, John Walden, Kirsten M. Leong, Geret S. DePiper, Cameron Speir, Suzana Blake, Karma Norman, Stephen Kasperski, Mariska Weijerman, and Kirsten L. L. Oleson. Identifying social thresholds and measuring social achievement in social-ecological systems: a crossregional comparison of fisheries in the United States. Marine Policy, 152(??):Article 105595, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001227.

Prince:2023:SRC

[PWL^{+}23b] Jeremy Prince, Xuefeng Wang, Kun Lin, Ani Suryanti, Sallehudin Jamon, Mudjekeewis D. Santos, Francisco SB.

Torres, Nicko Amor L. Flores, Vu Viet Ha, and Tran Thi Ngoc Anh. Supporting regional co-operation in the South China Sea with an initial assessment of the data-limited skipjack tuna fishery. Marine Policy, 155 (??):Article 105711, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002440.

Partain:2020:HOE

[PY20] Roy Andrew Partain and Constantinos Yiallourides. Hydrate occurrence in Europe: Risks, rewards, and legal frameworks. Marine Policy, 121(??):Article 104122, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20302062.

Pyc:2023:GOG

[Pyć23] Dorota Pyć. Global ocean governance: Towards protecting the ocean’s rights to health and resilience. Marine Policy, 147(??):Article 105328, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200375X.

Parsmo:2024:EDS

[PYVF24] Rasmus Parsmo, Erik Ytreberg, Maarten Verdaasdonk, and Erik Fridell. Environmental discounts for Swedish ports and fairways: a ship owner perspective. Marine Policy, 159(??):Article 105950, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004839.

Pauksztat:2021:ISI

[PZPB21] Birgit Pauksztat, Minghua Zhao, Claire Pekcan, and Mike Barnett. Introduction to the special issue on “International merchant shipping in the 21st century: Social science perspectives”. Marine Policy, 130(??):Article 104373, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20310241.

Peng:2021:OFC

[PZZ^{+}21] Daomin Peng, Shucheng Zhang, Hongzhi Zhang, Dazhong Pang, Qian Yang, Ruhao Jiang, Yuting Lin, Yongtong Mu, and Yugui Zhu. The oyster fishery in China: Trend, concerns and solutions. Marine Policy, 129(??):Article 104524, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001354.

Qafisheh:2024:GCG

[Qaf24] Mutaz M. Qafisheh. From Constantinople to Oslo to Gaza: Developments of Palestine’s maritime legislation under colonization. Marine Policy, 160(??):Article 105954, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004876.

Quimby:2020:IDE

[QCM^{+}20] Barbara Quimby, Stephen E. S. Crook, Karly Marie Miller, Jorge Ruiz, and David Lopez-Carr. Identifying, defining and exploring angling as urban subsistence: Pier fishing in Santa Barbara, California. Marine Policy, 121(??):Article 104197, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20304863.

Qiu:2020:NNI

[QF20] Wenxian Qiu and Jeremy Firestone. The non-negligible influence of global sea level change on the distribution of maritime zones. Marine Policy, 122(??):Article 104267, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309131.

Qi:2022:CFN

[Qi22] Xiaofei Qi. The conceptual framework of the national blue economic system: a multiagent perspective. Marine Policy, 145(??):Article 105287, November 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003347.

Qi:2024:MNB

[Qi24] Xiaofei Qi. Mapping the national blue economic system space: a case study of China. Marine Policy, 160(??):Article 105959, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2300492X.

Quevedo:2024:SRC

[QK24] Jay Mar D. Quevedo and Ryo Kohsaka. A systematic review of cultural ecosystem services of blue carbon ecosystems: Trends, gaps, and challenges in Asia and beyond. Marine Policy, 159(??):Article 105898, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004311.

Quevedo:2023:ADF

[QLU^{+}23] Jay Mar D. Quevedo, Kevin Muhamad Lukman, Yaya Ihya Ulumuddin, Yuta Uchiyama, and Ryo Kohsaka. Applying the DPSIR framework to qualitatively assess the globally important mangrove ecosystems of Indonesia: a review towards evidence-based policymaking approaches. Marine Policy, 147(??):Article 105354, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004018.

Qi:2024:IGA

[QLZ^{+}24] Xinli Qi, Zhenfu Li, Changping Zhao, Jiaqi Li, and Qiqi Zhang. International governance of Arctic shipping black carbon emissions: Current situation and impediments to progress. Marine Policy, 161(??):Article 106043, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24000411.

Quinio:2023:EEB

[QRK^{+}23] Louise Quinio, Malena Ripken, Thomas Klenke, Brice Trouillet, Henning Sten Hansen, and Lise Schrøder. Exploring ecosystem-based approaches in MSP through actordriven perceptual mapping. Marine Policy, 152(??):Article 105604, June 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001318.

Qin:2021:EMR

[QS21] Man Qin and Mingxue Sun. Effects of marine ranching policies on the ecological efficiency of marine ranching — Based on 25 marine ranching in Shandong Province. Marine Policy, 134(??):Article 104788, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003997.

Qu:2023:ACS

[QTBL23] Zoe Qu, Simon Thrush, Caitlin Blain, and Nicolas Lewis. Assessing the carbon storage value of kelp forest restoration in the Hauraki Gulf Marine Park, New Zealand: Lessons from no-take Marine Protected Areas. Marine Policy, 154(??):Article 105682, August 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002099.

Qin:2021:PIP

[QTD21] Man Qin, Qiaoan Tao, and Yuanwei Du. Policy implementation and project performance: a qualitative comparative analysis based on 29 National Marine ranchings in China. Marine Policy, 129(??):Article 104527, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100138X.

Qu:2021:EVS

[QTPL21] Zoe Qu, Simon Thrush, Darren Parsons, and Nicolas Lewis. Economic valuation of the snapper recruitment effect from a well-established temperate no-take marine reserve on adjacent fisheries. Marine Policy, 134(??):Article 104792, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004036.

Quevedo:2021:BCE

[QUK21] Jay Mar D. Quevedo, Yuta Uchiyama, and Ryo Kohsaka. A blue carbon ecosystems qualitative assessment applying the DPSIR framework: Local perspective of global benefits and contributions. Marine Policy, 128(??):Article 104462, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21000737.

Qiu:2021:VRF

[QWBY21] Yingying Qiu, Erda Wang, Yan Bu, and Yang Yu. Valuing recreational fishery attributes, opportunities and associated activities in China from the tourists’ satisfaction perspectives. Marine Policy, 131(??):Article 104616, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2100227X.

Qin:2020:ECM

[QYD20] Man Qin, Caixuan Yue, and Yuanwei Du. Evolution of China’s marine ranching policy based on the perspective of policy tools. Marine Policy, 117(??):Article 103941, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19308048.

Quan:2021:YRD

[QYX21] Yongbo Quan, Xiao Yu, and Weiwei Xu. The Yangtze River Delta integration and regional development of marine economy: Conference report. Marine Policy, 127(??):Article 104420, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000300.

Qi:2020:WDP

[QZZ20] Xiaofei Qi, Bingxin Zhao, and Jiang-Hua Zhang. What does the product space mean for firms in the development of marine-related products? Marine Policy, 115(??):Article 103822, May 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19306864.

Qi:2020:DNB

[QZZX20] Xiaofei Qi, Bingxin Zhao, Jianghua Zhang, and Wenwen Xiao. The drawing of a national blue product space and its evolution. Marine Policy, 112(??):Article 103773, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19304142.

Rasheed:2020:ESP

[RA20] A. Rifaee Rasheed and Ameer Abdulla. Evaluating stakeholder participatory processes in policy development for Marine Protected Areas. Marine Policy, 112(??):Article 103737, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19301496.

Roll:2022:EIF

[RAB22] Kristin Helen Roll, Frank Asche, and Trond Bjørndal. The effect of introducing fuel tax to the Norwegian fishery industry. Marine Policy, 135(??):Article 104829, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004401.

Rey-Ares:2021:IWC

[RAFLRP21] Lucía Rey-Ares, Sara Fernández-López, and David Rodeiro-Pazos. Impact of working capital management on profitability for Spanish fish canning companies. Marine Policy, 130(??):Article 104583, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001949.

Reis:2020:CCL

[RB20] Jeanette Reis and Rhoda C. Ballinger. Creating a climate for learning-experiences of educating existing and future decision-makers about climate change. Marine Policy, 111(??):??, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18305189.

Rozynski:2022:SQP

[RB22] Grzegorz Różyński and Małgorzata Bielecka. Sediment quality in the Polish part of the trans-boundary Vistula Lagoon: implications for deposition management. Marine Policy, 146(??):Article 105288, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003359.

R:2023:SLS

[RBA23] Sheeja S. R., Rahi. T. B., and Athira Ajay. State led social security and inclusion of marine fisherfolk: Analyzing the case of Kerala, India. Marine Policy, 147(??):Article 105392, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004390.

Ramieri:2024:DIM

[RBB^{+}24] Emiliano Ramieri, Martina Bocci, Daniele Brigolin, Pierpaolo Campostrini, Fabio Carella, Amedeo Fadini, Giulio Farella, Elena Gissi, Fabrizio Madeddu, Stefano Menegon, Micol Roversi Monaco, Francesco Musco, Folco Soffietti, Laura Barberi, and Andrea Barbanti. Designing and implementing a multi-scalar approach to Maritime Spatial Planning: the case study of Italy. Marine Policy, 159(??):Article 105911, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300444X.

Razak:2022:CRR

[RBEA^{+}22] Tries B. Razak, Lisa Boström-Einarsson, Cut Aja Gita Alisa, Rindah Talitha Vida, and Timothy A. C. Lamont. Coral reef restoration in Indonesia: a review of policies and projects. Marine Policy, 137(??):Article 104940, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21005510.

Rickard:2020:FMF

[RBNE20] Laura N. Rickard, Kofi Britwum, Caroline L. Noblet, and Keith S. Evans. Factory-made or farm fresh? Measuring U.S. support for aquaculture as a food technology. Marine Policy, 115(??):Article 103858, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302209.

Richmond:2022:RSC

[RC22] Laurie Richmond and Laura Casali. The role of social capital in fishing community sustainability: Spiraling down and up in a rural California port. Marine Policy, 137(??):Article 104934, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005455.

Roessger:2022:TTP

[RCH22] Julia Roessger, Joachim Claudet, and Barbara Horta e Costa. Turning the tide on protection illusions: the underprotected MPAs of the ‘OSPAR Regional Sea Convention’. Marine Policy, 142(??):Article 105109, August 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001567.

RodriguesdeFreitas:2022:CFE

[RCM^{+}22] Rodrigo Rodrigues de Freitas, Ronaldo Cataldo Costa, Carolina Mattosinho De Carvalho Alvite, Deisi Cristiane Balensiefer, Cecil Roberto de Maya Brotherhood de Barros, and Jonatas Henrique Prado. Challenges for fishers’ engagement in Marine Protected Areas: Lessons from Right Whale Environmental Protection Area, Southern Brazil. Marine Policy, 143(??):Article 105135, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001828.

Rahbek-Clemmensen:2020:HAC

[RCT20] Jon Rahbek-Clemmensen and Gry Thomasen. How has Arctic coastal state cooperation affected the Arctic Council? Marine Policy, 122(??):Article 104239, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030885X.

Ryan:2021:III

[RDD^{+}21] Karina L. Ryan, Cameron J. Desfosses, Ainslie M. Denham, Stephen M. Taylor, and Gary Jackson. Initial insights on the impact of COVID-19 on boat-based recreational fishing in Western Australia. Marine Policy, 132(??):Article 104646, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002578.

Ruiz-Diaz:2020:SEV

[RDLA^{+}20] Raquel Ruiz-Díaz, Xiaozi Liu, Alba Aguión, Gonzalo Macho, Maite deCastro, Moncho Gómez-Gesteira, and Elena Ojea. Social-ecological vulnerability to climate change in small-scale fisheries managed under spatial property rights systems. Marine Policy, 121(??):Article 104192, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20303766.

Rocha:2020:MTS

[RDM^{+}20] Diana Rocha, Benjamin Drakeford, Sarah A. Marley, Jonathan Potts, Michelle Hale, and Angie Gullan. Moving towards a sustainable cetacean-based tourism industry — a case study from Mozambique. Marine Policy, 120(??):Article 104048, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19305767.

Reimer:2023:CRM

[RDT^{+}23] Julie M. Reimer, Rodolphe Devillers, Brice Trouillet, Natalie C. Ban, Tundi Agardy, and Joachim Claudet. Conservation ready marine spatial planning. Marine Policy, 153(??):Article 105655, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23001823.

Rajapakse:2021:UFW

[RE21] Asanka Rajapakse and Gholam Reza Emad. Underlying factors which cause task deviation leading to dangerous situations at sea. Marine Policy, 130(??):Article 104548, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001597.

Rajapakse:2023:FUP

[RE23] Asanka Rajapakse and Gholam Reza Emad. Fatigue, an unsolved puzzle that continues contributing to accidents at sea. Marine Policy, 155(??):Article 105745, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002786.

Reedy:2020:NAS

[Ree20] Katherine Reedy. Neoliberal Aleutians: Seeing Like a fishing company, Seeing Like a coastal community. Marine Policy, 118(??):Article 103981, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301020.

Ren:2021:HRS

[Ren21a] Wenhan Ren. How to realize the sustainable development of mariculture industry — re-examine from the perspective of biased technological progress in China. Marine Policy, 134(??):Article 104791, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004024.

Ren:2021:RDC

[Ren21b] Wenhan Ren. Research on dynamic comprehensive evaluation of allocation efficiency of green science and technology resources in China’s marine industry. Marine Policy, 131(??):Article 104637, September 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002487.

Rector:2024:RSA

[RFG24] M. E. Rector, R. Filgueira, and J. Grant. The role of salmon aquaculture eco-certification in corporate social responsibility and the delivery of ecosystem services and disservices. Marine Policy, 160(??):Article 105948, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004815.

Ribeiro:2021:STL

[RFPS21] Marta Chantal Ribeiro, Rui Ferreira, Eliana Pereira, and Joana Soares. Scientific, technical and legal challenges of deep sea mining. A vision for Portugal — Conference report. Marine Policy, ??(??):Article 103338, ???? 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307346.

Ruslan:2022:MES

[RGH^{+}22] Nur Fatin Nabilah Ruslan, Hong Ching Goh, Caroline Hattam, Andrew Edwards-Jones, and Heng Hing Moh. Mangrove ecosystem services: Contribution to the wellbeing of the coastal communities in Klang Islands. Marine Policy, 144(??):Article 105222, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200269X.

Robertson:2020:LMM

[RGK^{+}20] Thomas Robertson, Suzie Greenhalgh, Isoa Korovulavula, Tomasi Tikoibua, Pio Radikedike, and Philip Stahlmann-Brown. Locally managed marine areas: Implications for socio-economic impacts in Kadavu, Fiji. Marine Policy, 117(??):Article 103950, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19309601.

Roberts:2020:EPM

[RHC20] Kelsey E. Roberts, Olivia Hill, and Carly N. Cook. Evaluating perceptions of marine protection in Australia: Does policy match public expectation? Marine Policy, 112(??):Article 103766, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19300375.

Rutherford:2020:CAA

[RHL20] V. E. Rutherford, J. M. Hills, and M. D. A. Le Tissier. Comparative analysis of adaptation strategies for coastal climate change in North West Europe. Marine Policy, 111(??):Article 102478, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X16304262.

Rubio:2024:APC

[RHO24] Iratxe Rubio, Alistair J. Hobday, and Elena Ojea. Adaption planning to climate change in industrial fisheries: Progress in the Basque tropical tuna fishery. Marine Policy, 161(??):Article 106001, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23005341.

Rybicki:2021:STI

[RHST21] Sandra Rybicki, Katell G. Hamon, Sarah Simons, and Axel Temming. Spatio-temporal interactions between Northeast Atlantic Mackerel and its fishery — Simulating different futures. Marine Policy, 133(??):Article 104740, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003511.

Ren:2021:HDE

[RJ21] Wenhan Ren and Jianyue Ji. How do environmental regulation and technological innovation affect the sustainable development of marine economy: New evidence from China’s coastal provinces and cities. Marine Policy, 128(??):Article 104468, June 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000798.

Rodriguez:2023:APO

[RJAT23] Richard Castillo Rodríguez, Jorge Jiménez, Katherine Arroyo Arce, and Katie R. Thompson. An analysis of policy options available to the International Maritime Organization to protect the Costa Rica Thermal Dome: Building the case for a particularly sensitive sea area. Marine Policy, 148(??):Article 105375, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004225.

Rahman:2020:ITM

[RKR20] Md Sadique Rahman, Mohammad Mizanul Haque Kazal, and Shah Johir Rayhan. Impacts of the training of mud crab farmers: an adaptation strategy to cope with salinity intrusion in Bangladesh. Marine Policy, 120(??):Article 104159, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20305856.

Ruskule:2023:TCG

[RKS^{+}23] Anda Ruskule, Jonne Kotta, Champa Rani Saha, Philipp Arndt, Didzis Ustups, Solvita Strāķe, and Lena Bergström. Testing the concept of green infrastructure at the Baltic Sea scale to support an ecosystem-based approach to management of marine areas. Marine Policy, 147(??): Article 105374, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X22004213.

Roozee:2024:MIO

[RKS^{+}24] Evelyn Roozee, Dongkyu Kim, Antonia Sohns, Jasper R. de Vries, Owen F. Temby, and Gordon M. Hickey. Managing inter-organizational trust and risk perceptions in transboundary fisheries governance networks. Marine Policy, 159(??):Article 105927, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004608.

Rojon:2021:ICP

[RLRS21] Isabelle Rojon, Nicholas-Joseph Lazarou, Nishatabbas Rehmatulla, and Tristan Smith. The impacts of carbon pricing on maritime transport costs and their implications for developing economies. Marine Policy, 132(??):Article 104653, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002645.

Reed:2020:DSM

[RLS20] Jodie R. Reed, Amanda T. Lombard, and Kerry J. Sink. A diversity of spatial management instruments can support integration of fisheries management and marine spatial planning. Marine Policy, 119(??):Article 104089, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19308292.

Reid:2021:SNZ

[RM21a] Oscar Schouten Reid and Fabien Medvecky. Sharks and New Zealand news media. Marine Policy, 134 (??):Article 104751, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003626.

Romero:2021:CTU

[RM21b] Pedro Romero and Oscar Melo. Can a Territorial Use Right for Fisheries management make a difference for fishing communities? Marine Policy, 124(??):Article 104359, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20310101.

Rosas-Munoz:2022:WCW

[RMBR22] Juan Rosas-Muñoz and Marjorie Baquedano-Rodríguez. What can we learn about the natural disasters and multidimensional poverty link for artisanal fishery in Chile? Marine Policy, 144(??):Article 105225, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200272X.

Racine:2021:CSA

[RMF^{+}21] Phoebe Racine, AnnaClaire Marley, Halley E. Froehlich, Steven D. Gaines, Ian Ladner, Ilan MacAdam-Somer, and Darcy Bradley. A case for seaweed aquaculture inclusion in U.S. nutrient pollution management. Marine Policy, 129(??):Article 104506, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001172.

Roberts:2022:PCR

[RMU^{+}22] Nicky Roberts, Buchari Mengge, Muh. Rifadly Utina, Farhan Muhatar, Anugerah, Arham Icwardanhi, R. Muhammad Zulkifli, and Austin Humphries. Patron-client relationships shape value chains in an Indonesian island-based fisheries system. Marine Policy, 143(??):Article 105142, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001890.

Rojas-Nazar:2022:VMR

[RNGRC22] Ursula A. Rojas-Nazar, Jonathan P. A. Gardner, Agnes Rouchon, and Ross Cullen. Valuing marine reserves: a case study from two locations in central New Zealand. Marine Policy, 140(??):Article 105051, June 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000987.

Rosello:2022:RFM

[Ros22] Mercedes Rosello. Regional fishery management organisation measures and the imposition of criminal and administrative sanctions in respect of high seas fishing. Marine Policy, 144(??):Article 105213, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002603.

Relano:2023:CLP

[RP23a] Veronica Relano and Daniel Pauly. Corrigendum to “The ‘Paper Park Index’: Evaluating marine protected area effectiveness through a global study of stakeholder perceptions” [Mar. Policy 151 (2023) 105571]. Marine Policy, 155(??):Article 105702, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300235X. See [RP23b].

Relano:2023:LPP

[RP23b] Veronica Relano and Daniel Pauly. The ‘Paper Park Index’: Evaluating Marine Protected Area effectiveness through a global study of stakeholder perceptions. Marine Policy, 151(??):Article 105571, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S0308597X23000982. See corrigendum [RP23a].

Raut:2023:AAF

[RPK23] Shrutika Raut, S. Prakash, and Amit Kumar. Assessing artisanal fishers’ attitude and perception towards electric rays conservation along Indian coast. Marine Policy, 157(??):Article 105826, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003597.

Rossi:2024:POI

[RPY^{+}24] Valentina Rossi, Carlo Pipitone, Katherine L. Yates, Fabio Badalamenti, Giovanni D’Anna, Cristina Pita, Fátima L. Alves, Jesús E. Argente-García, Jelena Basta, Joachim Claudet, Karsten Dahl, Simonetta Fraschetti, Ioannis Giovos, Peter Mackelworth, Mairi Maniopoulou, Vasiliki Markantonatou, Márcia Marques, Pedro Noguera-Méndez, Joanna Piwowarczyk, Violin Raykov, Gil Rilov, Bob Rumes, Alicia Said, María Semitiel-García, Yael Teff-Seker, Tomás Vega Fernández, and David Goldsborough. Poor online information on European marine protected areas impairs public participation under the Aarhus Convention. Marine Policy, 161(??):Article 106012, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000101.

Roman:2020:BRD

[RQB^{+}20] Catalina Román, Dante Queirolo, Claudio Bernal, Marcelo A. San Martín, and Victoria Escobar. Bycatch reduction device in the Chilean hake (Merluccius gayi) bottom trawl fishery, using a rigid separation grid. Marine Policy, 116(??):Article 103624, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X1830873X.

Reid:2020:IEB

[RR20] John Reid and Matthew Rout. The implementation of ecosystem-based management in New Zealand — A Māori perspective. Marine Policy, 117(??):Article 103889, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19305688.

Rout:2024:IBE

[RRM^{+}24] Matthew Rout, John Reid, Jason Paul Mika, Jay Whitehead, Annemarie Gillies, Fiona Wiremu, Georgia McLellan, and Corey Ruha. Indigenising the blue economy in Aotearoa New Zealand. Marine Policy, 161(??):Article 105987, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23005201.

Robertson:2023:PSN

[RROP23] Kerrie Robertson, Georgina Rood, Metua Okotai, and Mariata Pittman. Purse seiner non grata: Privy Council wades into Pacific fisheries. Marine Policy, 156(??):Article 105801, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003342.

Rakotomahazo:2023:EPI

[RRR^{+}23] Cicelin Rakotomahazo, Nirinarisoa Lantoasinoro Ranivoarivelo, Jacqueline Razanoelisoa, Gildas Georges Boleslas Todinanahary, Eulalie Ranaivoson, Mara Edouard Remanevy, Lalao Aigrette Ravaoarinorotsihoarana, and Thierry Lavitra. Exploring the policy and institutional context of a Payment for Ecosystem Services (PES) scheme for mangroves in southwestern Madagascar. Marine Policy, 148(??):Article 105450, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004973.

RoseteVerges:2024:CLE

[RS24] Fernando A. Rosete Verges and F. Javier Sanz Larruga. Coastal laws evolution in Spain: an analysis from the public policy cycle. Marine Policy, 162(??):Article 106052, April 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24000502.

Ramadhan:2022:HDD

[RSAP22] Andrian Ramadhan, Wilmar A. Salim, Teti Armiati Argo, and Puji Prihatiningsih. The human dimension dilemma in marine spatial planning. Marine Policy, 146(??):Article 105297, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200344X.

Riisager-Simonsen:2022:MNB

[RSFvH^{+}22] Christian Riisager-Simonsen, Gianna Fabi, Luc van Hoof, Noél Holmgren, Giovanna Marino, and Dennis Lisbjerg. Marine nature-based solutions: Where societal challenges and ecosystem requirements meet the potential of our oceans. Marine Policy, 144(??):Article 105198, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002457.

Rees:2020:ETS

[RSS^{+}20] Siân E. Rees, Emma V. Sheehan, Bryce D. Stewart, Robert Clark, Thomas Appleby, Martin J. Attrill, Peter J. S. Jones, David Johnson, Natasha Bradshaw, Simon Pittman, Jenny Oates, and Jean-Luc Solandt. Emerging themes to support ambitious UK marine biodiversity conservation. Marine Policy, 117(??):Article 103864, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930199X.

Romero-Tenorio:2022:IBS

[RTMCVMDE22] Arturo Romero-Tenorio, Manuel Mendoza-Carranza, Javier Francisco Valle-Mora, and Alberto Delgado-Estrella. Interactions between small-scale fisheries and marine mammals from the perspective of fishers in the Mexican tropical Pacific Coast. Marine Policy, 138(??):Article 104983, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000306.

Raubenheimer:2020:RGG

[RU20] Karen Raubenheimer and Niko Urho. Rethinking global governance of plastics — The role of industry. Marine Policy, 113(??):Article 103802, March 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309698.

Ruban:2024:BRA

[Rub24] Dmitry A. Ruban. Book review: Amanda Reichelt-Brushett (Ed.). Marine Pollution — Monitoring, Management and Mitigation, Springer, Cham (2023), xv + 369 pp., ISBN 978-3-031-10126-7, Hardback, EUR 49.99. Marine Policy, 160(??):Article 105990, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23005237.

Runolfsson:2024:MQP

[Run24] Birgir Th. Runolfsson. Measuring quality of property rights: Development of user rights quality in the Icelandic fisheries. Marine Policy, 160(??):Article 105962, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004955.

Russi:2020:GSS

[Rus20] Daniela Russi. Governance strategies for a successful marine protected area — The case of Torre Guaceto. Marine Policy, 115(??):Article 103849, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301770.

Richardson:2021:CMA

[RWVH21] Kelsey Richardson, Chris Wilcox, Joanna Vince, and Britta Denise Hardesty. Challenges and misperceptions around global fishing gear loss estimates. Marine Policy, 129(??):Article 104522, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001330.

Ren:2023:DHT

[RXN23] Wenhan Ren, Yuhan Xu, and Jing Ni. Development, hotspots and trend directions of marine ecological compensation in China: a bibliometric-based assessment. Marine Policy, 156(??):Article 105815, October 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003482.

Ren:2024:RIM

[RXX24] Wenhan Ren, Yuhan Xu, and Han Xiao. Research on the impact of marine ecological civilization demonstration zone policies on the green development level of China’s marine economy: a quasi natural experiment based on coastal cities. Marine Policy, 161(??):Article 106048, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000460.

Retnoningtyas:2021:SPM

[RYS^{+}21] Heidi Retnoningtyas, Irfan Yulianto, Arisetiarso Soemodinoto, Yudi Herdiana, Tasrif Kartawijaya, Mohamad Natsir, and Joko Tri Haryanto. Stakeholder participation in management planning for grouper and snapper fisheries in West Nusa Tenggara Province, Indonesia. Marine Policy, 128(??):Article 104452, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21000634.

Rajagopal:2021:HWU

[RZ21a] Sriram Rajagopal and Pengfei Zhang. How widespread is the usage of the Northern Sea Route as a commercially viable shipping route? A statistical analysis of ship transits from 2011 to 2018 based on empirical data. Marine Policy, 125(??):Article 104300, March 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309453.

Ren:2021:GTP

[RZ21b] Wenhan Ren and Qi Zeng. Is the green technological progress bias of mariculture suitable for its factor endowment? — — Empirical results from 10 coastal provinces and cities in China. Marine Policy, 124(??):Article 104338, February 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2030988X.

Scherer:2021:MEC

[SA21] Marinez E. G. Scherer and Milton L. Asmus. Modeling to evaluate coastal governance in Brazil. Marine Policy, 129(??):Article 104501, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001123.

Santos:2021:MCD

[SAA^{+}21] Catarina Frazão Santos, Tundi Agardy, Francisco Andrade, Larry B. Crowder, Charles N. Ehler, and Michael K. Orbach. Major challenges in developing marine spatial planning. Marine Policy, 132(??):Article 103248, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306213.

Stokes:2020:BUP

[SAB^{+}20] Debra Stokes, Kirin Apps, Paul A. Butcher, Betty Weiler, Hanabeth Luke, and Andrew P. Colefax. Beach-user perceptions and attitudes towards drone surveillance as a sharkbite mitigation tool. Marine Policy, 120(??):Article 104127, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2030347X.

Saha:2022:EEA

[Sah22] Choyon Kumar Saha. Emergence and evolution of aquaculture sustainability certification schemes. Marine Policy, 143(??):Article 105196, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002433.

Sakita:2021:CUD

[Sak21] Seishiro Sakita. Centralization under decentralization: the development of fishery clubs in Lesvos under the administrative reforms of Greece. Marine Policy, 132(??): Article 104655, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X21002669.

Sampson:2022:LBS

[Sam22] Helen Sampson. ‘Beyond the state’: the limits of international regulation and the example of abandoned seafarers. Marine Policy, 140(??):Article 105046, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000938.

Sanlier:2020:APS

[Şan20]  Şengül Şanlıer. Analysis of port state control inspec-

tion data: the Black Sea Region. Marine Policy, 112(??):Article 103757, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19305469.

Sagoe:2021:CPA

[SAO^{+}21] Alberta Ama Sagoe, Denis Worlanyo Aheto, Isaac Okyere, Richard Adade, and Justice Odoi. Community participation in assessment of fisheries related ecosystem services towards the establishment of marine protected area in the Greater Cape Three Points area in Ghana. Marine Policy, 124(??):Article 104336, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309854.

Straume:2024:SNS

[SAO^{+}24] Hans-Martin Straume, Frank Asche, Atle Oglend, Ivar Gaasland, Ingrid K. Pettersen, and Geir Sogn-Grundvåg. The structure of Norwegian seafood trade. Marine Policy, 159(??):Article 105921, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004542.

Sar:2023:CAR

[Sar23] Aysegul Bugra Sar. Considerations on assistance and rescue at sea in the light of the increasing autonomy in shipping. Marine Policy, 153(??):Article 105639, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001665.

Sylvester:2020:PAT

[SB20] Zephyr T. Sylvester and Cassandra M. Brooks. Protecting Antarctica through co-production of actionable science: Lessons from the CCAMLR marine protected area process. Marine Policy, 111(??):Article 103720, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305354.

Schwoerer:2023:FFF

[SBD^{+}23] Tobias Schwoerer, Kevin Berry, Darcy G. Dugan, David C. Finnoff, Molly Mayo, Jan Ohlberger, and Eric J. Ward. Fish or not fish — fisheries participation and harvest diversification under economic and ecological change. Marine Policy, 157(??):Article 105833, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003664.

Siddiqi:2024:UPC

[SBDC24] Aliya C. Siddiqi, Brock J. Bergseth, Amy Diedrich, and Andrew Chin. Understanding the perceived conservation benefits of shark-marine tourism in the Global South. Marine Policy, 161(??):Article 105995, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23005286.

Seminara:2023:AFK

[SBFHD^{+}23] Cecilia Inés Seminara, Márcio Luiz Vargas Barbosa-Filho, Rachel Ann Hauser-Davis, Gabriel Barros Gonçalves Souza, Marcela Alvarenga, Christiana Cabicieri Profice, and Salvatore Siciliano. Artisanal fisher knowledge on the impacts of destructive and illegal practices on the Southern Coast of Bahia, Brazil. Marine Policy, 158(??):Article 105862, December 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003950.

Shumway:2021:PSF

[SBJF^{+}21] Nicole Shumway, Justine Bell-James, James A. Fitzsimons, Rose Foster, Chris Gillies, and Catherine E. Lovelock. Policy solutions to facilitate restoration in coastal marine environments. Marine Policy, 134(??):Article 104789, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004000.

Schwartz-Belkin:2024:EBI

[SBP24] Inbar Schwartz-Belkin and Michelle E. Portman. Exploring barriers to the implementation of geospatial technologies in marine spatial planning: Reports from practitioners. Marine Policy, 159(??):Article 105891, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004244.

Sinan:2021:DPI

[SBS21] Hussain Sinan, Megan Bailey, and Wilf Swartz. Disentangling politics in the Indian Ocean Tuna Commission. Marine Policy, 133(??):Article 104781, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003924.

Smythe:2021:ORI

[SBT21] Tiffany Smythe, David Bidwell, and Grant Tyler. Optimistic with reservations: the impacts of the United States’ first offshore wind farm on the recreational fishing experience. Marine Policy, 127(??):Article 104440, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000506.

Soares:2023:DVN

[SBT^{+}23] Marcelo O. Soares, Daysiane Barbosa Brandão, Carlos Eduardo Peres Teixeira, Rivelino Martins Cavalcante, André Henrique Barbosa de Oliveira, Luis Ernesto Arruda Bezerra, and Eduardo Lacerda Barros. Déjà vu: New oil spill poses cumulative risks to protected coastal environments in the South Atlantic. Marine Policy, 155(??):Article 105764, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2300297X.

Solis:2021:MEA

[SBW^{+}21] Daniel Solís, Kathryn Bisack, John Walden, Paul Richards, and Juan Agar. Measuring the economic abatement cost of sea turtle bycatch in the Northwest Atlantic commercial pelagic longline fishery. Marine Policy, 133 (??):Article 104734, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003456.

Schaar:2021:FHM

[SC21] Siena I. Schaar and Linda J. Cox. The future for Hawai’i’s marine aquarium fishery: a cost benefit analysis compared to an environmental impact assessment. Marine Policy, 127(??):Article 104429, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000397.

Shen:2022:IFS

[SC22] Chuyi Shen and Tinggui Chen. Impact of fuel subsidies on bottom trawl fishery operation in China. Marine Policy, 138(??):Article 104977, April 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000240.

Siano:2020:CPM

[SCA^{+}20] R. Siano, A. Chapelle, V. Antoine, E. Michel-Guillou, F. Rigaut-Jalabert, L. Guillou, H. Hégaret, A. Leynaert, and A. Curd. Citizen participation in monitoring phytoplankton seawater discolorations. Marine Policy, 117(??):Article 103039, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X16306169.

Stratoudakis:2022:OSD

[SCA^{+}22] Yorgos Stratoudakis, Miguel Caetano, Cláudia Afonso, Cátia Bartilloti, Antonina dos Santos, Pedro Terrinha, Ivone Figueiredo, Miguel Miranda, José Lino Costa, Isabel Caçador, Vanessa Fonseca, Paula Chainho, Ricardo Melo, Bernardo Quintella, Maria José Costa, André B. Fortunato, Paula Freire, Marta Rodrigues, Luís Portela, Miguel Bessa Pacheco, Joaquim Pombo, Carla Palma, Anabela Oliveira, António Gonçalves Tavares, Filipe Porteiro, Inês Metelo, José Sá Fernandes, Maria da Conceição Freitas, Henrique Cabral, Marina Cunha, Christopher Pham, and Inês Carvalho. One step from the deep ocean: Linking research and governance for the Tagus river estuary and the Lisbon canyon system — a conference report. Marine Policy, 144(??):Article 105224, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002718.

Seidu:2022:EFS

[SCBB^{+}22] Issah Seidu, Francoise Cabada-Blanco, Lawrence K. Brobbey, Berchie Asiedu, Paul Barnes, Moro Seidu, and Nicholas K. Dulvy. “Every fish in the sea is meat and so are guitarfishes”: Socio-economic drivers of a guitarfish fishery in Ghana. Marine Policy, 143(??):Article 105159, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002068.

Stratoudakis:2020:IPM

[SCBdA20] Yorgos Stratoudakis, C. Correia, A. F. Belo, and P. R. de Almeida. Improving participated management under poor fishers’ organization: Anadromous fishing in the estuary of Mondego River, Portugal. Marine Policy, 119(??):Article 104049, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20300427.

Sjostrom:2021:GLE

[SCCW21] Anja J. C. Sjostrom, Lorenzo Ciannelli, Flaxen Conway, and W. Waldo Wakefield. Gathering local ecological knowledge to augment scientific and management understanding of a living coastal resource: the case of Oregon’s nearshore groundfish trawl fishery. Marine Policy, 131(??):Article 104617, September 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002281.

Sbragaglia:2021:RUS

[SCD21] Valerio Sbragaglia, Ricardo A. Correia, and Enrico Di Minin. Responsible use of social media data is needed: a reply to Maya-Jariego et al. “Plenty of black money: Netnography of illegal recreational underwater fishing in southern Spain”. Marine Policy, 134(??):Article 104780, December 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003912.  See

[MJMAA21, MJMAA22].

Snook:2022:JBY

[SCF^{+}22] Jamie Snook, Ashlee Cunsolo, James Ford, Chris Furgal, Andria Jones-Bitton, and Sherilee Harper. “Just because you have a land claim, that doesn’t mean everything’s going to fall in place”: an Inuit social struggle for fishery access and well-being. Marine Policy, 140(??):Article 105071, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2200118X.

Shih:2020:COF

[SCGC20] Yi-Che Shih, Yen-Chiang Chang, Warwick Gullett, and Wen-Yan Chiau. Challenges and opportunities for fishery rights negotiations in disputed waters — a Taiwanese perspective regarding a fishing boat case incident. Marine Policy, 121(??):Article 103755, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301289.

Smith:2024:DPC

[SCH^{+}24] M. Luke Smith, Guangqing Chi, Hannah Hennighausen, Davin Holen, and E. Lance Howe. Differences in perceptions of COVID-19 risks in a fishing community in Alaska, 2020– 2021. Marine Policy, 161(??):Article 106045, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000435.

Schreiber:2022:BJC

[SCJ22] Milena Arias Schreiber, Ratana Chuenpagdee, and Svein Jentoft. Blue Justice and the co-production of hermeneutical resources for small-scale fisheries. Marine Policy, 137(??):Article 104959, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000069.

Soto:2021:CSF

[SCLM^{+}21] Doris Soto, Carlos Chávez, Jorge León-Muñoz, Carol Luengo, and Yuri Soria-Galvarro. Chilean salmon farming vulnerability to external stressors: the COVID 19 as a case to test and build resilience. Marine Policy, 128(??):Article 104486, June 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2100097X.

Scott:2022:WCE

[SCSR^{+}22] Molly Scott, Edward Cardona, Kaylee Scidmore-Rossing, Mark Royer, Jennifer Stahl, and Melanie Hutchinson. What’s the catch? Examining optimal longline fishing gear configurations to minimize negative impacts on non-target species. Marine Policy, 143(??):Article 105186, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002330.

Scott:2023:CWC

[SCSR^{+}23] Molly Scott, Edward Cardona, Kaylee Scidmore-Rossing, Mark Royer, Jennifer Stahl, and Melanie Hutchinson. Corrigendum to “What’s the catch? Examining optimal longline fishing gear configurations to minimize negative impacts on non-target species” [Mar. Policy 143 (2022) 105186]. Marine Policy, 148(??):Article 105465, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22005164.

Sun:2024:CCT

[SCY24] Ruihong Sun, Yue Chu, and Xinliang Ye. China’s cruise tourism policy evolution and effectiveness analysis. Marine Policy, 159(??):Article 105893, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004268.

Santos:2024:RJO

[SdACdSLdAS24] Thauan Santos, Joilson de Assis Cabral, Paulo Vitor dos Santos Lima, and Matheus de Andrade Santos. Rio de Janeiro’s ocean economy as a key vector for sustainable development in Brazil. Marine Policy, 159(??):Article 105876, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004098.

Soares:2022:ICV

[SdCMN^{+}22] Júlia Benevenuti Soares, Marcus Rodrigues da Costa, Cassiano Monteiro-Neto, Luciana Loto, Maurício Düppré de Abreu, and Rafael de Almeida Tubino. Impacts of COVID-19 on the value chain of a small-scale fishery system in a tropical metropolitan city. Marine Policy, 140(??):Article 105068, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001154.

Sletten:2021:BBH

[SDG^{+}21] Jennifer Sletten, Mimi D’Iorio, Mary G. Gleason, Alex Driedger, Timothé Vincent, Claire Colegrove, Dawn Wright, and Virgil Zetterlind. Beyond the boundaries: How regulation-centered marine protected area information improves ocean protection assessments. Marine Policy, 124(??):Article 104340, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309908.

Sletten:2022:LCB

[SDG^{+}22] Jennifer Sletten, Mimi D'Iorio, Mary G. Gleason, Alex Driedger, Timothé Vincent, Claire Colegrove, Dawn Wright, and Virgil Zetterlind. Corrigendum to “Beyond the boundaries: How regulation-centered marine protected area information improves ocean protection assessments” [124, 2021, 104340]. Marine Policy, 146(??):Article 105203, December 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002500.

Sala:2023:TIR

[SDG^{+}23] Antonello Sala, Jochen Depestele, Aysun Gümüş, Pascal Laffargue, J. Rasmus Nielsen, Hans Polet, Chris J. Smith, Mustafa Zengin, Francois Bastardie, Ole R. Eigaard, Katell G. Hamon, Frank Jensen, Alessandro Lucchetti, Sonia Méhault, Emilio Notti, Nadia Papadopoulou, Andrea Petetta, Mattias Sköld, Benoit Vincent, and Adriaan D. Rijnsdorp. Technological innovations to reduce the impact of bottom gears on the seabed. Marine Policy, 157(??):Article 105861, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003949.

Saviolidis:2020:RBG

[SDI^{+}20] Nína M. Saviolidis, Brynhildur Davídsdóttir, Leena Ilmola, Anastasia Stepanova, Matilda Valman, and Elena Rovenskaya. Realising blue growth in the fishing industry in Iceland and Norway: Industry perceptions on drivers and barriers to blue growth investments and policy implications. Marine Policy, 117(??):Article 103967, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308760.

Sgardeli:2022:ASD

[SDL^{+}22] Vasiliki Sgardeli, Dimitrios Damalas, Angelos Liontakis, Christos D. Maravelias, Danai Mantopoulou-Palouka, and George Tserpes. The Aegean Sea demersal fishery under four climatic and socio-political futures. Marine Policy, 144(??):Article 105194, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200241X.

Setyawan:2022:HAM

[SEG^{+}22] Edy Setyawan, Mark Erdmann, Nikka Gunadharma, Tiene Gunawan, Abdi Hasan, Muhamad Izuan, Meidiarti Kasmidi, Yusdi Lamatenggo, Sarah Lewis, Nugraha Maulana, Ronald Mambrasar, Meity Mongdong, Alberth Nebore, Mochamad Iqbal Herwata Putra, Abraham Sianipar, Kristian Thebu, Syafri Tuharea, and Rochelle Constantine. A holistic approach to manta ray conservation in the Papuan Bird’s Head Seascape: Resounding success, ongoing challenges. Marine Policy, 137(??):Article 104953, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005649.

Sellheim:2020:CAI

[Sel20] Nikolas Sellheim. The CITES appendix II-Listing of mako sharks — Revisiting counter arguments. Marine Policy, 115(??):Article 103887, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307742.

Seta:2023:CIC

[Set23] Makoto Seta. Compulsory insurance for cruise vessels as a preparation for the next pandemic: Law of the sea perspective. Marine Policy, 152(??):Article 105586, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001136.

Seung:2022:DGV

[Seu22] Chang K. Seung. Decomposing global value chain (GVC) income for world fisheries. Marine Policy, 137(??):Article 104950, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21005613.

Song:2022:CCM

[SF22] Annie Young Song and Michael Fabinyi. China’s 21st century maritime silk road: Challenges and opportunities to coastal livelihoods in ASEAN countries. Marine Policy, 136(??):Article 104923, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005340.

Song:2024:GCS

[SFB24] Annie Young Song, Michael Fabinyi, and Kate Barclay. China and seafood trade after coronavirus: the role of state actors. Marine Policy, 160(??):Article 105992, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23005250.

Sanz-Fernandez:2022:PGC

[SFGEPC22] Víctor Sanz-Fernández, Juan Carlos Gutiérrez-Estrada, and Inmaculada Pulido-Calvo. Pagellus genus catches time series in the FAO Major Fishing Areas 27 and 34: Analysis of fishery behaviour. Marine Policy, 136(??):Article 104912, February 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21005236.

Stein:2021:CEP

[SFN^{+}21] Florian Martin Stein, Jens Frankowski, Vincent Nijman, Christine Absil, Irene Kranendonk, and Willem Dekker. Chinese eel products in EU markets imply the effectiveness of trade regulations but expose fraudulent labelling. Marine Policy, 132(??):Article 104651, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002621.

Standal:2020:INB

[SG20a] Dag Standal and Eduardo Grimaldo. Institutional nuts and bolts for a mesopelagic fishery in Norway. Marine Policy, 119(??):Article 104043, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309649.

Stojanovic:2020:GFT

[SG20b] Tim Stojanovic and Kira Gee. Governance as a framework to theorise and evaluate marine planning. Marine Policy, 120(??):Article 104115, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307560.

Suuronen:2020:MMF

[SG20c] Petri Suuronen and Eric Gilman. Monitoring and managing fisheries discards: New technologies and approaches. Marine Policy, 116(??):Article 103554, June 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308716.

Standal:2021:LTP

[SG21] Dag Standal and Eduardo Grimaldo. Lost in translation? Practical- and scientific input to the mesopelagic fisheries discourse. Marine Policy, 134(??):Article 104785, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003961.

Shilland:2021:QSA

[SGA^{+}21] Robyn Shilland, Gabriel Grimsditch, Mohamed Ahmed, Salomão Bandeira, Hilary Kennedy, Maria Potouroglou, and Mark Huxham. A question of standards: Adapting carbon and other PES markets to work for community seagrass conservation. Marine Policy, 129(??):Article 104574, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001858.

Sakib:2021:APC

[SGB21] Nazmus Sakib, Elena Gissi, and Hermanni Backer. Addressing the pollution control potential of marine spatial planning for shipping activity. Marine Policy, 132(??):Article 104648, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002591.

Solo-Gabriele:2021:TIM

[SGFM^{+}21] Helena M. Solo-Gabriele, Tom Fiddaman, Cecilie Mauritzen, Cameron Ainsworth, David M. Abramson, Igal Berenshtein, Eric P. Chassignet, Shuyi S. Chen, Robyn N. Conmy, Christa D. Court, William K. Dewar, John W. Farrington, Michael G. Feldman, Alesia C. Ferguson, Elizabeth Fetherston-Resch, Deborah French-McCay, Christine Hale, Ruoying He, Vassiliki H. Kourafalou, Kenneth Lee, Yonggang Liu, Michelle Masi, Emily S. Maung-Douglass, Steven L. Morey, Steven A. Murawski, Claire B. Paris, Natalie Perlin, Erin L. Pulster, Antonietta Quigg, Denise J. Reed, James J. Ruzicka, Paul A. Sandifer, John G. Shepherd, Burton H. Singer, Michael R. Stukel, Tracey T. Sutton, Robert H. Weisberg, Denis Wiesenburg, Charles A. Wilson, Monica Wilson, Kateryna M. Wowk, Callan Yanoff, and David Yoskowitz. Towards integrated modeling of the long-term impacts of oil spills. Marine Policy, 131(??):Article 104554, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001652.

Sogn-Grundvaag:2022:QEF

[SGH22] Geir Sogn-Grundvåg and Øystein Hermansen. Qualityenhancing fishing in the coastal fishery for Atlantic cod in Norway. Marine Policy, 143(??):Article 105191, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2200238X.

Shao:2021:ERG

[SGK21] Qinglong Shao, Junjie Guo, and Peng Kang. Environmental response to growth in the marine economy and urbanization: a heterogeneity analysis of 11 Chinese coastal regions using a panel vector autoregressive model. Marine Policy, 124(??):Article 104350, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310009.

Standal:2020:GII

[SGL20] Dag Standal, Eduardo Grimaldo, and Roger B. Larsen. Governance implications for the implementation of biodegradable gillnets in Norway. Marine Policy, 122(??):Article 104238, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X20308848.

Stacey:2021:DSS

[SGL^{+}21] Natasha Stacey, Emily Gibson, Neil R. Loneragan, Carol Warren, Budy Wiryawan, Dedi S. Adhuri, Dirk J. Steenbergen, and Ria Fitriana. Developing sustainable smallscale fisheries livelihoods in Indonesia: Trends, enabling and constraining factors, and future opportunities. Marine Policy, 132(??):Article 104654, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002657.

Sguotti:2023:TAC

[SGL^{+}23] Camilla Sguotti, Sanmitra Gokhale, Tin-Yu Lai, Esther Schuch, Christian Möllmann, and Andries Richter. Telecouplings in Atlantic cod — the role of global trade and climate change. Marine Policy, 157(??):Article 105818, November 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003512.

Serra-Goncalves:2023:SMD

[SGLF^{+}23] Catarina Serra-Gonçalves, Jennifer L. Lavers, Andrew M. Fischer, Heidi L. Tait, and Alexander L. Bond. Stopping marine debris at the source: Effectiveness of source reduction plans in Australia. Marine Policy, 155(??):Article 105776, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003093.

Shadrack:2021:SAS

[SGPF21] Ronick Spenly Shadrack, Sompert Gereva, Tim Pickering, and Marta Ferreira. Seasonality, abundance and spawning season of milkfish Chanos chanos (Forsskål, 1775) at Teouma Bay, Vanuatu. Marine Policy, 130(??): Article 104587, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X21001986.

Schuch:2021:IIE

[SGR21] Esther Schuch, Silke Gabbert, and Andries P. Richter. Institutional inertia in European fisheries — Insights from the Atlantic horse mackerel case. Marine Policy, 128(??):Article 104464, June 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000750.

Strating:2022:CCM

[SGRW22] Rebecca Strating, Douglas Guilfoyle, Steven Ratuva, and Joanne Wallis. Commentary: China in the Maritime Pacific. Marine Policy, 141(??):Article 105092, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001397.

Song:2021:EDO

[SGWL21] Weiling Song, Jing Guo, Yue Wang, and Linlin Li. An evaluation of the development of the Ocean Economy in China using an Ocean Economic Development Index. Marine Policy, 132(??):Article 104691, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100302X.

Sogn-Grundvaag:2021:FQM

[SGZH^{+}21] Geir Sogn-Grundvåg, Dengjun Zhang, Edgar Henriksen, Sjurdur Joensen, Bjørn-Inge Bendiksen, and Øystein Hermansen. Fish quality and market performance: the case of the coastal fishery for Atlantic cod in Norway. Marine Policy, 127(??):Article 104449, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000592.

Standal:2022:CED

[SH22] Dag Standal and Bjørn Hersoug. Caught in the environmental discourse — coastal fisheries under pressure. Marine Policy, 143(??):Article 105183, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002305.

Standal:2023:IFC

[SH23] Dag Standal and Bjørn Hersoug. Illegal fishing: a challenge to fisheries management in Norway. Marine Policy, 155(??):Article 105750, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300283X.

Shao:2020:NEM

[Sha20] Qinglong Shao. Nonlinear effects of marine economic growth and technological innovation on marine pollution: Panel threshold analysis for China’s 11 coastal regions. Marine Policy, 121(??):Article 104110, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308371.

Shan:2022:EFO

[Sha22] Desai Shan. Enforcement of fishing Occupational Health and Safety (OHS) standards: Challenges in Atlantic Canada. Marine Policy, 145(??):Article 105282, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003293.

Sprogis:2023:EWW

[SHAC23] Kate R. Sprogis, Dirk Holman, Patricia Arranz, and Fredrik Christiansen. Effects of whale-watching activities on southern right whales in Encounter Bay, South Australia. Marine Policy, 150(??):Article 105525, April 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000520.

Sumby:2021:HFR

[SHFP21] Jonathan Sumby, Marcus Haward, Elizabeth A. Fulton, and Gretta T. Pecl. Hot fish: the response to climate change by regional fisheries bodies. Marine Policy, 123(??):Article 104284, January 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309301.

Shi:2020:SDB

[Shi20a] Yubing Shi. Settlement of disputes in a BBNJ agreement: Options and analysis. Marine Policy, 122 (??):Article 104156, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20302621.

Shih:2020:TPT

[Shi20b] Yi-Che Shih. Taiwan’s progress towards becoming an ocean country. Marine Policy, 111(??):Article 103725, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19302180.

Shields:2022:CWW

[Shi22] Monika W. Shields. Commercial whale-watching reduces vessel incidents in the vicinity of killer whales in Washington State. Marine Policy, 145(??):Article 105290, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003372.

Shi:2023:RBG

[Shi23a] Xinxiang Shi. The relationship between general principles of international law and Article 38(1) of the ICJ statute: a law of the sea perspective. Marine Policy, 148(??):Article 105427, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004742.

Shi:2023:IAC

[Shi23b] Xuefei Shi. International advocacy and China’s distant water fisheries policies. Marine Policy, 152(??):Article 105635, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001628.

Smith:2021:EPI

[SHP^{+}21] David C. Smith, Malcolm Haddon, André E. Punt, Caleb Gardner, L. Richard Little, Stephen Mayfield, Michael F. O’Neill, Thor Saunders, John Stewart, Brent Wise, Elizabeth A. Fulton, and Simon Conron. Evaluating the potential for an increased and sustainable commercial fisheries production across multiple jurisdictions and diverse fisheries. Marine Policy, 124(??):Article 104353, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20310034.

Schutter:2021:BEB

[SHPW21] Marleen S. Schutter, Christina C. Hicks, Jacob Phelps, and Claire Waterton. The blue economy as a boundary object for hegemony across scales. Marine Policy, 132(??):Article 104673, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002840.

Song:2021:UAO

[SHZ^{+}21] Lan Song, Zhiqiu Huang, Huahan Zhang, Kaiwei Tian, Ning Yin, Yanmin Xu, Langxiong Gan, Jiangchao Wang, Jihong Chen, Yaqing Shu, and Chunmiao Zheng. The urgency to address the occupational health of Chinese seafarers for sustainable development. Marine Policy, 129(??):Article 104518, July 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001299.

Sun:2021:CGR

[SHZC21] Yue Sun, Yang Han, Liangfu Zhang, and Yen-Chiang Chang. The construction and governance of regional maritime security order in the South China Sea: Conference report. Marine Policy, 126(??):Article 104438, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000488.

Sakai:2022:SCE

[SII22] Yutaro Sakai, Hiroe Ishihara, and Mitsuhiro Ishino. Sharing in the commons: Evaluating the pooling system in a Danish seine fishery in Japan. Marine Policy, 139(??):Article 105017, May 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000641.

Singh:2021:TYD

[Sin21] Pradeep A. Singh. The two-year deadline to complete the International Seabed Authority’s Mining Code: Key outstanding matters that still need to be resolved. Marine Policy, 134(??):Article 104804, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004152.

Sari:2021:MSS

[SIW^{+}21] Irna Sari, Muhammad Ichsan, Alan White, Syahril Abdul Raup, and Sugeng Hari Wisudo. Monitoring smallscale fisheries catches in Indonesia through a fishing logbook system: Challenges and strategies. Marine Policy, 134(??):Article 104770, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100381X.

Singer:2021:LBM

[SJ21] Rebecca Singer and Peter J. S. Jones. Lyme Bay marine protected area: a governance analysis. Marine Policy, 127(??):Article 103201, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304317.

Samy-Kamal:2020:EFM

[SK20a] Mohamed Samy-Kamal. Estonian fisheries management system under the lens of the Marine Stewardship Council (MSC) standard. Marine Policy, 117(??):Article 103885, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19308826.

Samy-Kamal:2020:OFP

[SK20b] Mohamed Samy-Kamal. Outlook on the fisheries policy reform in Egypt and the draft of the new fisheries law. Marine Policy, 120(??):Article 104136, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302700.

Shiraishi:2024:EWU

[SK24] Hiromi Shiraishi and Kenzo Kaifu. Early warning of an upsurge in international trade in the American eel. Marine Policy, 159(??):Article 105938, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004712.

Szymanowska:2023:SIT

[SKDK23] Beata Barbara Szymanowska, Arkadiusz Kozłowski, Janusz Dabrowski, and Hanna Klimek. Seaport innovation trends: Global insights. Marine Policy, 152(??):Article 105585, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001124.

Shimose:2022:EOR

[SKE22] Tamaki Shimose, Minoru Kanaiwa, and Akihiko Ebisawa. Estimation of the optimal regulation size considering bid price increase with growth in size selectable coral reef fisheries. Marine Policy, 137(??):Article 104961, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000082.

Simon:2020:UAV

[SKL^{+}20] Julien Simon, Dorothée Kopp, Pascal Larnaud, Jean-Philippe Vacherot, Fabien Morandeau, Gaël Lavialle, and Marie Morfin. Using automated video analysis to study fish escapement through escape panels in active fishing gears: Application to the effect of net colour. Marine Policy, 116(??):Article 103785, June 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308078.

Svorken:2023:UOS

[SKL23] Marianne Svorken, Ingrid Kvalvik, and Nicholas Lord. Understanding the organisational structure of fisheries crime in well-regulated fisheries. Marine Policy, 157 (??):Article 105860, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003937.

Smith:2021:DAS

[SKS21] William E. Smith, Gerard T. Kyle, and Steven Sutton. Displacement and associated substitution behavior among Texas inshore fishing guides due to perceived spotted seatrout declines. Marine Policy, 131(??):Article 104624, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002359.

Sorensen:2022:FGA

[SKW^{+}22] JA Sorensen, L. Kincl, R. Weil, J. Dzugan, and D. Christel. Fisheries governance and associated health implications: Current perspectives from US commercial fishermen. Marine Policy, 141(??):Article 105119, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200166X.

Shan:2020:VME

[SL20] Jingzhu Shan and Jingmei Li. Valuing marine ecosystem service damage caused by land reclamation: Insights from a deliberative choice experiment in Jiaozhou Bay. Marine Policy, 122(??):Article 104249, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308952.

Santos:2023:CNA

[SL23] Mateus Rennó Santos and Michael J. Lynch. The cross-national association between institutional imbalance, national culture, and illegal fishing. Marine Policy, 151(??):Article 105601, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23001288.

Schluter:2021:IBT

[SLB21] Maja Schlüter, Emilie Lindkvist, and Xavier Basurto. The interplay between top-down interventions and bottomup self-organization shapes opportunities for transforming self-governance in small-scale fisheries. Marine Policy, 128(??):Article 104485, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21000968.

Strachan:2022:RIF

[SLH22] Laura L. Strachan, Richard J. Lilley, and Sebastian J. Hennige. A regional and international framework for evaluating seagrass management and conservation. Marine Policy, 146(??):Article 105306, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003530.

Suuronen:2023:RMC

[SLK^{+}23] Petri Suuronen, Sven-Gunnar Lunneryd, Sara Königson, Nelson F. Coelho, Åsa Waldo, Viktor Eriksson, Kristina Svels, Esa Lehtonen, Iwona Psuty, and Markus Vetemaa. Reassessing the management criteria of growing seal populations: the case of Baltic grey seal and coastal fishery. Marine Policy, 155(??):Article 105684, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002117.

Scherrer:2024:SRI

[SLL^{+}24] Kim J. N. Scherrer, Tom J. Langbehn, Gabriella Ljungström, Katja Enberg, Sara Hornborg, Gjert Dingsør, and Christian Jørgensen. Spatial restrictions inadvertently doubled the carbon footprint of Norway’s mackerel fishing fleet. Marine Policy, 161(??):Article 106014, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000125.

Soto:2021:SWC

[SLMG^{+}21] Doris Soto, Jorge León-Muñoz, René Garreaud, Renato A. Quiñones, and Francisco Morey. Scientific warnings could help to reduce farmed salmon mortality due to harmful algal blooms. Marine Policy, 132(??):Article 104705, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2100316X.

Strongin:2022:PFT

[SLP^{+}22] Kyle Strongin, Alana Malinde S. N. Lancaster, Beth Polidoro, Alfonso Aguilar-Perera, Leah Gerber, Patricia González Díaz, Juliett González-Méndez, Larry McKinney, Héctor Espinosa Pérez, Daniel Pech, Dorka Cobián Rojas, Steven Saul, and Susana Perera Valderrama. A proposal framework for a Tri-National Agreement on Biological Conservation in the Gulf of Mexico Large Marine Ecosystem. Marine Policy, 139(??):Article 105041, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000884.

Suyo:2020:SNA

[SLS^{+}20] Jee Grace B. Suyo, Virginie Le Masson, Louise Shaxson, Maria Rovilla J. Luhan, and Anicia Q. Hurtado. A social network analysis of the Philippine seaweed farming industry: Unravelling the web. Marine Policy, 118(??):Article 104007, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19309571.

Suyo:2021:NRU

[SLS^{+}21] Jee Grace B. Suyo, Virginie Le Masson, Louise Shaxson, Maria Rovilla J. Luhan, and Anicia Q. Hurtado. Navigating risks and uncertainties: Risk perceptions and risk management strategies in the Philippine seaweed industry. Marine Policy, 126(??):Article 104408, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100018X.

Sidik:2023:BCN

[SLW^{+}23] Frida Sidik, Anissa Lawrence, Tonny Wagey, Franky Zamzani, and Catherine E. Lovelock. Blue carbon: a new paradigm of mangrove conservation and management in Indonesia. Marine Policy, 147(??):Article 105388, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004353.

Sheng:2021:IST

[SLY21] X. U. Sheng, Binbin Lu, and Qingde Yue. Impact of scitech finance on the innovation efficiency of China’s marine industry. Marine Policy, 133(??):Article 104708, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003195.

Schinas:2023:CSC

[SM23] Orestis Schinas and Daniel Metzger. Cyber-seaworthiness: a critical review of the literature. Marine Policy, 151(??): Article 105592, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X23001197.

Simmons:2021:SSA

[SMC^{+}21] Peter Simmons, Michael Mehmet, Belinda Curley, Nicola Ivory, Kane Callaghan, Kim Wolfenden, and Gang Xie. A scenario study of the acceptability to ocean users of more and less invasive management after shark-human interactions. Marine Policy, 129(??):Article 104558, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100169X.

Sparks:2022:WLS

[SMCW22] Jessica L. Decker Sparks, Lee Matthews, Daniel Cárdenas, and Chris Williams. Worker-less social responsibility: How the proliferation of voluntary labour governance tools in seafood marginalise the workers they claim to protect. Marine Policy, 139(??):Article 105044, May 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000914.

Sahri:2020:CRM

[SMDM20] Achmad Sahri, Putu Liza Kusuma Mustika, Hadi Yoga Dewanto, and Albertinka J. Murk. A critical review of marine mammal governance and protection in Indonesia. Marine Policy, 117(??):Article 103893, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304397.

Schwoerer:2023:BHO

[SMF^{+}23] Tobias Schwoerer, Aaron Martin, Ginny Fay, Erik R. Schoen, and Michael Buntjer. Batten down the hatches: Opportunities to protect Alaska from biological invasions through watercraft trade and traffic. Marine Policy, 148(??):Article 105448, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200495X.

Santana-Morales:2021:HMT

[SMHPMO^{+}21] Omar Santana-Morales, E. Mauricio Hoyos-Padilla, Alfonso Medellín-Ortíz, Chugey Sepulveda, Rodrigo Beas-Luna, Marc Aquino-Baleytó, Edgar E. Becerril-García, Daniel Arellano-Millán, Luis Malpica-Cruz, Julio Lorda, and José Leonardo Castillo-Géniz. How much is too much? A carrying capacity study of white shark cage diving in Guadalupe Island, Mexico. Marine Policy, 131(??):Article 104588, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001998.

Smith:2020:BRC

[Smi20] Hance Smith. Book review: Climate Change and Ocean Governance: politics and policy for threatened seas, Paul G. Harris (Ed.). Cambridge University Press, Cambridge (2019), xiii + 450. Marine Policy, 112(??):Article 103504, February 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19302441.

Smith:2022:BRP

[Smi22] Hance Smith. Book review: The Poseidon Project: the struggle to govern the world’s oceans, David Bosco. Oxford University Press (2022), Notes, Selected Bibliography, Index, pp. x + 305. Marine Policy, 143(??):Article 105157, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002044.

Smith:2023:BRF

[Smi23a] Hance Smith. Book review: Fishing Europe’s Troubled Waters: fifty years of fisheries policy, David Symes. Earthscan from Routledge, New York (2023). Marine Policy, 155(??):Article 105744, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002774.

Smith:2023:BRO

[Smi23b] Hance Smith. Book review: Ocean Geopolitics: Marine Resources, Maritime Boundary Disputes and the Law of the Sea, Andreas Osthagen. Edward Elgar, Cheltenham (2022), Bibliography,Index. Pp. ix + 160. Marine Policy, 148(??):Article 105471, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200522X.

Simard:2022:SWP

[SMKS22] Nittya S. M. Simard, Thane A. Militz, Jeff Kinch, and Paul C. Southgate. Shocks within a Pacific island fishery: a historic study of events impacting the mother-of-pearl fishery in Papua New Guinea. Marine Policy, 143(??):Article 105173, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002202.

Salvadeo:2021:FPI

[SMLRB^{+}21] Christian Salvadeo, Hem Nalini Morzaria-Luna, Héctor Reyes-Bonilla, Antonina Ivanova-Bonchera, David Petatán Ramírez, and Eduardo Juárez-León. Fisher’s perceptions inform adaptation measures to reduce vulnerability to climate change in a Mexican natural protected area. Marine Policy, 134(??):Article 104793, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004048.

Shiffman:2021:MSC

[SMW^{+}21] D. S. Shiffman, C. C. Macdonald, J. N. Wester, M. B. Walsh, A. Chevalier, D. Kachelriess, and K. J. Friedman. Marine species conservation at CITES: How does media coverage inform or misinform? Marine Policy, 134(??):Article 104813, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004243.

Seto:2022:TTG

[SMYH22] Katherine Seto, Nathan Miller, Mark Young, and Quentin Hanich. Toward transparent governance of transboundary fisheries: the case of Pacific tuna transshipment. Marine Policy, 136(??):Article 104200, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309133.

Suh:2023:CCD

[SN23] Neville N. Suh and Richard A. Nyiawung. Climate change dynamics and youth participation decisions in aquatic food systems: Case of the oyster sector in The Gambia, West Africa. Marine Policy, 156(??):Article 105804, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003378.

Steenbergen:2020:CRA

[SNK^{+}20] D. J. Steenbergen, P. T. Neihapi, D. Koran, A. Sami, V. Malverus, R. Ephraim, and N. Andrew. COVID-19 restrictions amidst cyclones and volcanoes: a rapid assessment of early impacts on livelihoods and food security in coastal communities in Vanuatu. Marine Policy, 121(??):Article 104199, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20306898.

Sebe:2021:UCE

[SNP21] Maxime Sèbe, Abdelhak Nassiri, and Linwood Pendleton. Using choice experiment designs to evaluate mitigation solutions to reduce whale-ship collisions. Marine Policy, 124(??):Article 104368, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310198.

Shan:2023:ISI

[SNS23] Desai Shan, Barb Neis, and Julie Sorensen. Introduction to special issue on governance and fishing health and safety in the North Atlantic. Marine Policy, 155(??):Article 105729, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002622.

Stoner-Osborne:2020:EMP

[SO20] Blake Stoner-Osborne. The effects of marine protected areas on populations of commercial reef fishes in Moorea, French Polynesia. Marine Policy, 121(??):Article 104177, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19309625.

Sugimura:2022:NPC

[SOK^{+}22] Yoshihisa Sugimura, Tomonari Okada, Tomohiro Kuwae, Yugo Mito, Ryoji Naito, and Yasuyuki Nakagawa. New possibilities for climate change countermeasures in ports: Organic carbon containment and creation of blue carbon ecosystems through beneficial utilization of dredged soil. Marine Policy, 141(??):Article 105072, July 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001191.

Solino:2023:SPL

[Sol23a] Mario Soliño. Social preferences for large marine protected areas in NW Spain. Marine Policy, 155(??):Article 105781, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003147.

Solomon:2023:CCT

[Sol23b] Preethi Solomon. Can cooperatives be a tool for poverty reduction? Social capital perspectives of fisher’s cooperatives in India. Marine Policy, 147(??):Article 105373, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004201.

Soto-Onate:2020:ULA

[SOLN20] David Soto-Oñate and Ana C. Lemos-Nobre. Unreported landings across the EU Exclusive Economic Zones: the role of regional governance and social capital. Marine Policy, 111(??):Article 103723, January 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303434.

Soto-Onate:2021:EUL

[SOLN21] David Soto-Oñate and Ana C. Lemos-Nobre. The European Union landing obligation: the compliance problems derived from its multilevel approach. Marine Policy, 132(??):Article 104666, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002773.

Sheridan:2021:RPL

[SON21] Kate Sheridan, M. Justin O’Riain, and Mark D. Needham. Recreationist perceptions of lethal and non-lethal management of sharks in two of South Africa’s marine areas. Marine Policy, 132(??):Article 104633, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100244X.

Brien:2022:MPD

[SONB^{+}22] Aleke Stöfen-O’Brien, Abolfazl Naji, Amy L. Brooks, Jenna R. Jambeck, and Farhan R. Khan. Marine plastic debris in the Arabian/Persian Gulf: Challenges, opportunities and recommendations from a transdisciplinary perspective. Marine Policy, 136(??):Article 104909, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005200.

Salgueiro-Otero:2020:BUS

[SOO20] Diego Salgueiro-Otero and Elena Ojea. A better understanding of social-ecological systems is needed for adapting fisheries to climate change. Marine Policy, 122(??):Article 104123, December 2020. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20300440.

Seara:2022:LLN

[SOP^{+}22] Tarsila Seara, Adrien Owens, Richard Pollnac, Robert Pomeroy, and Christopher Dyer. Lessons learned from a natural resource disaster: the long-term impacts of the Long Island Sound lobster die-off on individuals and communities. Marine Policy, 136(??):Article 104943, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005546.

Snyder:2020:HCE

[SOS20] Hunter T. Snyder, Natuk Lund Olsen, and Andrew M. Song. How cybernetics explains institutional failure: a case of Greenland’s open-air fish markets. Marine Policy, 117(??):Article 103882, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19301058.

Sosin:2021:BRG

[Sos21a] Claudia Sosin. Book review: Governing Marine Living Resources in the Polar Regions, Nengye Liu, Cassandra M. Brooks, Tianbao Qin (Eds.). Edward Elgar Publishing (2019). Marine Policy, 129(??):Article 104552, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001639.

Storkersen:2021:FPD

[SOS^{+}21b] Kristine Vedal Størkersen, Tonje C. Osmundsen, Lars Helge Stien, Christian Medaas, Marianne Elisabeth Lien, Brit Tørud, Tore S. Kristiansen, and Kristine Gismervik. Fish protection during fish production. Organizational conditions for fish welfare. Marine Policy, 129(??):Article 104530, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2100141X.

Saito:2022:TAN

[SOT^{+}22] Rui Saito, Yoshioki Oozeki, Kenji Takasaki, Tsutomu Saito, Shinji Uehara, and Masanori Miyahara. Transshipment activities in the North Pacific Fisheries Commission Convention Area. Marine Policy, 146(??):Article 105299, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003463.

Sreya:2021:EVS

[SPA^{+}21] P. S. Sreya, Chitra Parayil, N. Aswathy, Binoo P. Bonny, T. P. Aiswarya, and P. O. Nameer. Economic vulnerability of small-scale coastal households to extreme weather events in Southern India. Marine Policy, 131(??):Article 104608, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002190.

Smith:2021:MRM

[SPCCvT21] Christopher J. Smith, K. Nadia Papadopoulou, Eira Carballo-Cárdenas, and Jan P. M. van Tatenhove. Marine restoration in the Mediterranean: red coral and fan mussel discourses, uncertainty and reaching restoration targets. Marine Policy, 128(??):Article 104488, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000993.

Said:2020:SSF

[SPFA^{+}20] Alicia Said, Jose Pascual-Fernández, Vanessa Iglésias Amorim, Mathilde Højrup Autzen, Troels Jacob Hegland, Cristina Pita, Johanna Ferretti, and Jerneja Penca. Smallscale fisheries access to fishing opportunities in the European Union: Is the Common Fisheries Policy the right step to SDG14b? Marine Policy, 118(??):Article 104009, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20302475.

Sanchez:2020:EPP

[SPG20] José L. Fernández Sánchez, José M. Fernández Polanco, and Ignacio Llorente García. Evidence of price premium for MSC-certified products at fishers’ level: the case of the artisanal fleet of common octopus from Asturias (Spain). Marine Policy, 119(??):Article 104098, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304592.

Stebbings:2020:MEU

[SPH^{+}20] Emily Stebbings, Eleni Papathanasopoulou, Tara Hooper, Melanie C. Austen, and Xiaoyu Yan. The marine economy of the United Kingdom. Marine Policy, 116(??):Article 103905, June 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19307390.

Senghor:2023:CGR

[SPHO23] Khadidiatou Senghor, Stefan Partelow, Cesar Giraldo Herrera, and Isimemen Osemwegie. Conflicting governance realities: Aligning historical and cultural practices with formal marine protected area co-management in Senegal. Marine Policy, 155(??):Article 105706, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002397.

Stockin:2022:ERH

[SPJB22] Karen A. Stockin, Matthew D. M. Pawley, Rebecca M. Jarvis, and Rebecca M. Boys. Examining the role of human perceptions during cetacean stranding response in New Zealand. Marine Policy, 145(??):Article 105283, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2200330X.

Speir:2023:MPL

[SPMN23] Cameron Speir, Amanda Phillips, Aaron Mamula, and Karma Norman. A measure of port-level resilience to shocks in commercial fisheries. Marine Policy, 151(??):Article 105575, May 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001021.

Sliskovic:2021:NIS

[SPN^{+}21] Merica Slišković, Marina Piria, Vedrana Nerlović, Katarina Pavičić Ivelja, Ana Gavrilović, and Gorana Jelić Mrčelić. Non-indigenous species likely introduced by shipping into the Adriatic Sea. Marine Policy, 129(??):Article 104516, July 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001275.

Stavroulakis:2020:ESA

[SPT^{+}20] Peter J. Stavroulakis, Stratos Papadimitriou, Vangelis Tsioumas, Ioannis G. Koliousis, Elena Riza, and Freideriki Tsirikou. Exploratory spatial analysis of maritime clusters. Marine Policy, 120(??):Article 104125, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309042.

Squalli:2020:EPE

[Squ20] Jay Squalli. Evaluating the potential economic, environmental, and social benefits of orange-spotted grouper aquaculture in the United Arab Emirates. Marine Policy, 118(??):Article 103998, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20301809.

Squillace:2021:BRP

[Squ21] Mark Squillace. Best regulatory practices for deep seabed mining: Lessons learned from the U.S. Surface Mining Control and Reclamation Act. Marine Policy, 125(??):Article 104327, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2030974X.

Spaniol:2022:AIB

[SR22] Matthew J. Spaniol and Nicholas J. Rowland. Anticipated innovations for the blue economy: Crowdsourced predictions for the North Sea Region. Marine Policy, 137(??):Article 104874, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004851.

Sugimoto:2022:HCN

[SRH^{+}22] Aoi Sugimoto, Raphael Roman, Juri Hori, Norie Tamura, Shingo Watari, and Mitsutaku Makino. How has the ‘customary nature’ of Japanese fisheries reacted to Covid-19? An interdisciplinary study examining the impacts of the pandemic in 2020. Marine Policy, 138(??):Article 105005, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000525.

Stoll:2023:RSS

[SRH23] Joshua S. Stoll, Sarah C. Risley, and Paloma R. Henriques. A review of small-scale marine fisheries in the United States: Definitions, scale, drivers of change, and policy gaps. Marine Policy, 148(??):Article 105409, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004560.

Suarez-Rojas:2023:WDY

[SRLLG23] Chaitanya Suárez-Rojas, Carmelo J. León, and Yen E. Lam-González. What drives you to the sea? Animal rights, environmental protection and sensation seeking. Marine Policy, 147(??):Article 105348, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003955.

Sarker:2021:MZI

[SRR^{+}21] Subrata Sarker, Md Jalilur Rahman, Muhammad Mizanur Rahman, Morgina Akter, Md Shajjadur Rahman, and Md Abdul Wahab. MPA zoning integrating socioecological data in the South East coast of Bangladesh. Marine Policy, 133(??):Article 104736, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100347X.

Strand:2023:CEC

[SRS23] M. Strand, N. Rivers, and B. Snow. The complexity of evaluating, categorising and quantifying marine cultural heritage. Marine Policy, 148(??):Article 105449, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004961.

Suris-Regueiro:2021:AFE

[SRSGMGG21a] Juan C. Surís-Regueiro, José L. Santiago, Xosé M. González-Martínez, and M. Dolores Garza-Gil. An applied framework to estimate the direct economic impact of Marine Spatial Planning. Marine Policy, 127(??):Article 104443, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000531.

Suris-Regueiro:2021:EEI

[SRSGMGG21b] Juan C. Surís-Regueiro, José L. Santiago, Xosé M. González-Martínez, and M. Dolores Garza-Gil. Estimating economic impacts linked to Marine Spatial Planning with input-output techniques. Application to three case studies. Marine Policy, 129(??):Article 104541, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001524.

Sabela-Rikhotso:2021:CAL

[SRvNN21] Phindile Tiyiselani Zanele Sabela-Rikhotso, Dewald van Niekerk, and Livhuwani David Nemakonde. A critical analysis of the legal frameworks governing oil spill management in South Africa. Marine Policy, 127(??):Article 104433, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21000439.

Strating:2024:HRS

[SRY24] Rebecca Strating, Sunil Rao, and Sallie Yea. Human rights at sea: the limits of inter-state cooperation in addressing forced labour on fishing vessels. Marine Policy, 159(??):Article 105934, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004670.

Skerritt:2021:BGD

[SS21a] Daniel J. Skerritt and U. Rashid Sumaila. Broadening the global debate on harmful fisheries subsidies through the use of subsidy intensity metrics. Marine Policy, 128(??):Article 104507, June 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001184.

Sowman:2021:JTN

[SS21b] M. Sowman and Jackie Sunde. A just transition? Navigating the process of policy implementation in small-scale fisheries in South Africa. Marine Policy, 132(??):Article 104683, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002943.

Silver:2022:FIC

[SS22] Jennifer J. Silver and Joshua S. Stoll. A framework for investigating commercial license and quota holdings in an era of fisheries consolidation, concentration and financialization. Marine Policy, 143(??):Article 105179, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002263.

Silva:2023:EBL

[SS23] Luana C. C. Silva and Bráulio A. Santos. Evolution of Brazilian legislation regarding the marine fauna: Advances and shortcomings between 1960 and 2020. Marine Policy, 153(??):Article 105638, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001653.

Scotts:2023:SCR

[SSA^{+}23] Gemma L. Scotts, Michael J. Scales, Gonzalo Araujo, Hollie Booth, and Sarah A. Marley. Socio-cultural relationship between recreational sea anglers and blue sharks (Prionace glauca) in the United Kingdom. Marine Policy, 157(??):Article 105831, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003640.

Snyder:2024:GIH

[SSAP24] Hunter T. Snyder, Simon Stone, Mary Albert, and Chris Polashenski. Are Greenland’s inshore halibut fishers ready for individual transferable quotas? Marine Policy, 159(??):Article 105932, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004657.

Silva:2022:EFL

[SSB^{+}22] Monalisa R. O. Silva, André B. Silva, Jaciana C. Barbosa, Cássia Amaral, and Priscila F. M. Lopes. Empowering fisherwomen leaders helped reduce the effects of the COVID-19 pandemic on fishing communities: Insights from Brazil. Marine Policy, 135(??):Article 104842, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100453X.

Short:2023:MZM

[SSB^{+}23] Charlie Short, Joanna L. Smith, John Bones, Steve Diggon, Aaron Heidt, Chris Mcdougall, and Kylee A. Pawluk. Marine zoning for the Marine Plan Partnership (MaPP) in British Columbia, Canada. Marine Policy, 152(??):Article 105524, June 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000519.

Shannon:2020:GBB

[SSDQ20] Caitriona Shannon, Paul D. Stebbing, Alison M. Dunn, and Claire H. Quinn. Getting on board with biosecurity: Evaluating the effectiveness of marine invasive alien species biosecurity policy for England and Wales. Marine Policy, 122(??):Article 104275, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309210.

Szymkowiak:2023:ICA

[SSF23] Marysia Szymkowiak, Andrew Steinkruger, and Kelsi Furman. Integrating climate adaptation into Gulf of Alaska fishing community planning. Marine Policy, 156(??):Article 105802, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003354.

Sherman:2023:GPU

[SSH^{+}23] C. Samantha Sherman, Colin A. Simpfendorfer, Alifa B. Haque, Eric D. Digel, Patrick Zubick, Jonathan Eged, Jay H. Matsushiba, Glenn Sant, and Nicholas K. Dulvy. Guitarfishes are plucked: Undermanaged in global fisheries despite declining populations and high volume of unreported international trade. Marine Policy, 155(??):Article 105753, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002865.

Shimotsuura:2023:FHS

[SSK23] Taiga Shimotsuura, Tomomi Shoda, and Shigemi Kagawa. Firm heterogeneity in sources of changes in CO_{2} emissions from international container shipping. Marine Policy, 157(??):Article 105859, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003925.

Solaun:2021:SOB

[SSMG21] Oihana Solaun, Yolanda Sagarminaga, Iratxe Menchaca, and Ibon Galparsoro. Spatial overlap between human activities and seabed habitats in European Seas: Insights into EMODnet’s data for management purposes. Marine Policy, 130(??):Article 104551, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001627.

Serra-Sogas:2021:UAS

[SSOP^{+}21] Norma Serra-Sogas, Patrick D. O’Hara, Kim Pearce, Leh Smallshaw, and Rosaline Canessa. Using aerial surveys to fill gaps in AIS vessel traffic data to inform threat assessments, vessel management and planning. Marine Policy, 133(??):Article 104765, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003766.

Sowman:2021:UGF

[SSP^{+}21a] M. Sowman, J. Sunde, T. Pereira, B. Snow, P. Mbatha, and A. James. Unmasking governance failures: the impact of COVID-19 on small-scale fishing communities in South Africa. Marine Policy, 133(??):Article 104713, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003249.

Sunny:2021:AIC

[SSP^{+}21b] Atiqur Rahman Sunny, Sharif Ahmed Sazzad, Shamsul Haque Prodhan, Md. Ashrafuzzaman, Gopal Chandra Datta, Ashoke Kumar Sarker, Mizanur Rahman, and Mahmudul Hasan Mithun. Assessing impacts of COVID-19 on aquatic food system and small-scale fisheries in Bangladesh. Marine Policy, 126(??):Article 104422, April 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000324.

Skerritt:2023:MUG

[SSV^{+}23] Daniel J. Skerritt, Anna Schuhbauer, Sebastian Villasante, Andrés M. Cisneros-Montemayor, Nathan J. Bennett, Tabitha G. Mallory, Vicky W. L. Lam, Robert I. Arthur, William W. L. Cheung, Louise S. L. Teh, Katina Roumbedakis, Maria L. D. Palomares, and U. Rashid Sumaila. Mapping the unjust global distribution of harmful fisheries subsidies. Marine Policy, 152(??):Article 105611, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001380.

Stancheva:2022:SMU

[SSZ^{+}22] Margarita Stancheva, Hristo Stanchev, Jacek Zaucha, Emiliano Ramieri, and Timothy Roberts. Supporting multiuse of the sea with maritime spatial planning. The case of a multi-use opportunity development — Bulgaria, Black Sea. Marine Policy, 136(??):Article 104927, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005388.

Seminara:2024:OCS

[SSZ^{+}24] Cecilia Inés Seminara, Gabriel Barros Gonçalves Souza, Cleverson Zapelini, Márcio Luiz Vargas Barbosa-Filho, and Christiana Cabicieri Profice. Opportunities and challenges for small-scale fisheries co-management: a case study of the unofficial landing records system in a Brazilian fishing community. Marine Policy, 160(??):Article 105991, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23005249.

Stadtlander:2023:BRP

[Sta23a] Christian T. K. H. Stadtländer. Book review: S. Partelow, M. Hadjimichael, A.-K. Hornidge, Ocean Governance: knowledge systems, policy foundations and thematic, 2023 Springer Nature Publishing Cham, Switzerland xx + 436 pp., ISBN: 978-3-031-20739-6 (Hardcover), ISBN: 978-3-031-20742-6 (Softcover), ISBN: 978-3-031-20740-2 (eBook). Marine Policy, 153(??):Article 105645, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001720.

Starr:2023:TAC

[Sta23b] Antony Starr. Total allowable catches and fishing opportunity decisions of the Council of the EU — a voting power index approach. Marine Policy, 156(??):Article 105798, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003317.

Stepien:2023:CSO

[Ste23] Barbara Stepień. Can a ship be its own captain? Safe manning of autonomous and uncrewed vessels. Marine Policy, 148(??):Article 105451, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004985.

Syddall:2021:TAP

[STF21] Victoria Syddall, Simon Thrush, and Karen Fisher. Transdisciplinary analysis of Pacific tuna fisheries: a research framework for understanding and governing oceans as social–ecological systems. Marine Policy, 134(??):Article 104783, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003948.

Stien:2020:GWN

[STG^{+}20] Lars Helge Stien, Brit Tørud, Kristine Gismervik, Marianne Elisabeth Lien, Christian Medaas, Tonje Osmundsen, Tore S. Kristiansen, and Kristine Vedal Størkersen. Governing the welfare of Norwegian farmed salmon: Three conflict cases. Marine Policy, 117(??):Article 103969, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309108.

Schug:2020:UOD

[STIS20] Donald M. Schug, Peter H. Taylor, Suzanne Iudicello, and Jill H. Swasey. Using online data visualization and analysis to facilitate public involvement in management of catch share programs. Marine Policy, 122(??):Article 104272, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309180.

Su:2022:UST

[STKC22] Shu Su, Yi Tang, Jacob P. Kritzer, and Yong Chen. Using systems thinking to diagnose science-based fisheries management in China. Marine Policy, 138(??):Article 104974, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000215.

Storkersen:2021:SMR

[Stø21] Kristine Vedal Størkersen. Safety management in remotely controlled vessel operations. Marine Policy, 130(??):Article 104349, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309994.

Scuderi:2023:SCG

[STS^{+}23] Alessia Scuderi, Rocco Tiberti, Javier García Sanabria, Lucia Merino, Cristina Otero-Sabio, André Pedrosa, and Filomena Cardoso Martins. Sustainability as a common goal: Regulatory compliance, stakeholder perspectives, and management implications of whale and dolphin watching in the Strait of Gibraltar. Marine Policy, 154(??):Article 105663, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23001902.

Socrate:2022:AUA

[SV22] Juliana Socrate and Eleonora Verón. Analysis of uses and activities in the Argentine Sea. Bases for a marine spatial planning in the North Argentina Basin. Marine Policy, 139(??):Article 105014, May 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000616.

Schluter:2020:CRS

[SVF20] Achim Schlüter, Colin Vance, and Sebastian Ferse. Coral reefs and the slow emergence of institutional structures for a glocal land- and sea-based collective dilemma. Marine Policy, 112(??):Article 103505, February 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303579.

Schluter:2020:LSI

[SVHV20] Achim Schlüter, Kristof Van Assche, Anna-Katharina Hornidge, and Nataşa Văidianu. Land-sea interactions and coastal development: an evolutionary governance perspective. Marine Policy, 112(??):Article 103801, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309273.

Steins:2021:COW

[SVKP21] Nathalie A. Steins, Jeroen A. Veraart, Judith E. M. Klostermann, and Marnix Poelman. Combining offshore wind farms, nature conservation and seafood: Lessons from a Dutch community of practice. Marine Policy, 126(??):Article 104371, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20310228.

Sordo:2023:RHW

[SVP^{+}23] Laura Sordo, Paulo Vasconcelos, David Piló, André N. Carvalho, Fábio Pereira, and Miguel B. Gaspar. Recreational harvesting of the wedge clam (Donax trunculus) in southern Portugal: Characterization of the activity based on harvesters’ perception and local ecological knowledge. Marine Policy, 155(??):Article 105694, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300221X.

Seixas:2024:HCP

[SVV24] Sónia Seixas, Tiago Verdelhos, and Helena Veríssimo. How COVID-19 pandemic affected fisheries (catch volume and price): a case study in Europe. Marine Policy, 159(??):Article 105896, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004293.

Strating:2022:MST

[SW22] Rebecca Strating and Joanne Wallis. Maritime sovereignty and territorialisation: Comparing the Pacific Islands and South China Sea. Marine Policy, 141(??):Article 105110, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001579.

Shelton:2024:DRG

[SWF^{+}24] Clare Shelton, Carole S. White, Johanna Forster, Susan Conlon, Georg H. Engelhard, and John K. Pinnegar. Disaster risk in Caribbean fisheries: How vulnerability is shaped and how it can be reduced in Dominica and Antigua and Barbuda. Marine Policy, 160(??):Article 105951, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004840.

Song:2022:MCO

[SWGZ22] Runxi Song, Lei Wu, Matt Geraci, and Hui Zhong. Maritime Cooperation and Ocean Governance 2021: Symposium report. Marine Policy, 146(??):Article 105302, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003499.

Sjafrie:2023:NAR

[SWH^{+}23] Nurul Dhewani Mirah Sjafrie, Pramaditya Wicaksono, Udhi E. Hernawan, Triyono, Doddy M. Yuwono, Muhammad Hafizt, Novi Susetyo Adi, Rohani Ambo-Rappe, Bayu Prayudha, Muhammad Banda Selamat, Sofia Yunia Sani, Setiawan Djody Harahap, Huwaida Nur Salsabila, Jennifer Wijaya, and Chris Roelfsema. Network analysis reveals overlapping roles of stakeholders related to seagrass-data provisioning in Indonesia. Marine Policy, 157(??):Article 105837, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003706.

Shiiba:2022:HBF

[SWHT22] Nagisa Shiiba, Hsing Hao Wu, Michael C. Huang, and Hajime Tanaka. How blue financing can sustain ocean conservation and development: a proposed conceptual framework for blue financing mechanism. Marine Policy, 139(??):Article 104575, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2100186X.

Sari:2022:TEA

[SWI^{+}22] Irna Sari, Alan White, Muhammad Ichsan, Jason Cope, Josh Nowlis, Christovel Rotinsulu, Stephanus Mandagi, Enos Menai, Zulkifli Henan, Rishi Sharma, Syafri Tuharea, Roger Tabalessy, and Melisa Masengi. Translating the ecosystem approach to fisheries management into practice:

Case of anchovy management, Raja Ampat, West Papua, Indonesia. Marine Policy, 143(??):Article 105162, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002093.

Su:2021:TPM

[SWW^{+}21] Ming Ming Su, Geoffrey Wall, Bangle Wu, Huangxuan Xu, Xiaohan Fu, and Yuncheng Deng. Tourism place making through the bioluminescent “Blue Tears” of Pingtan Islands, China. Marine Policy, 133(??):Article 104744, November 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003559.

Su:2021:ATC

[SWXM21] Meng Su, Lingling Wang, Jiahao Xiang, and Yuxia Ma. Adjustment trend of China’s marine fishery policy since 2011. Marine Policy, 124(??):Article 104322, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309696.

Sikveland:2020:DCS

[SZ20] Marius Sikveland and Dengjun Zhang. Determinants of capital structure in the Norwegian salmon aquaculture industry. Marine Policy, 119(??):Article 104061, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930764X.

Spijkerboer:2020:PMS

[SZBA20] R. C. Spijkerboer, C. Zuidema, T. Busscher, and J. Arts. The performance of marine spatial planning in coordinating offshore wind energy with other sea-uses: the case of the Dutch North Sea. Marine Policy, 115(??):Article 103860, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19300703.

Szymkowiak:2020:GFA

[Szy20] Marysia Szymkowiak. Genderizing fisheries: Assessing over thirty years of women’s participation in Alaska fisheries.

Marine Policy, 115(??):Article 103846, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307936.

Tang:2022:SMH

[TAK^{+}22] Lijun Tang, Sanley Abila, Momoko Kitada, Serafin Malecosio, and Karima Krista Montes. Seafarers’ mental health during the COVID-19 pandemic: an examination of current supportive measures and their perceived effectiveness. Marine Policy, 145(??):Article 105276, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003232.

Tang:2022:IFC

[Tan22] Lijun Tang. The impact of 2008 financial crisis and COVID-19 pandemic on the demand and supply of seafarer officers: Evidence from China. Marine Policy, 145(??):Article 105263, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003104.

Tan:2023:BRE

[Tan23a] Ying Tan. Book review: Emerging Technology and the Law of the Sea, James Kraska, Young Kil Park (Eds.). Cambridge University Press (2022), 369, ISBN 978-1-316-51742-0. Marine Policy, 152(??):Article 105622, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001495.

Tan:2023:RHI

[Tan23b] Ying Tan. Book review: Research Handbook on International Marine Environmental Law, 2nd Edition, Rosemary Rayfuse, Aline Jaeckel, Natalie Klein (Eds.). Edward Elgar Publishing, Inc. (2023), 492, ISBN 978-1-78990-907-4. Marine Policy, 157(??):Article 105853, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300386X.

Tang:2024:FFF

[Tan24] Jianye Tang. Form follows function: an initial evaluation of the BBNJ Agreement’s achievements regarding the “not undermining” proviso. Marine Policy, 159(??):Article 105952, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004852.

Turrell:2023:CRI

[TAP^{+}23] W. R. Turrell, W. E. N. Austin, S. P. Philbrick, C. Tilbrook, and H. Kennedy. Clarifying the role of inorganic carbon in blue carbon policy and practice. Marine Policy, 157(??):Article 105873, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004062.

Teschke:2021:PMP

[TBH^{+}21] Katharina Teschke, Patricia Brtnik, Stefan Hain, Heike Herata, Alexander Liebschner, Hendrik Pehlke, and Thomas Brey. Planning marine protected areas under the CCAMLR regime — The case of the Weddell Sea (Antarctica). Marine Policy, 124(??):Article 104370, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20310216.

Turschwell:2020:CBR

[TBPC20] Mischa P. Turschwell, Christopher J. Brown, Ryan M. Pearson, and Rod M. Connolly. China’s Belt and Road Initiative: Conservation opportunities for threatened marine species and habitats. Marine Policy, 112(??):Article 103791, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19307675.

Tran:2023:ROV

[TBV^{+}23] Trung Tien Tran, Thomas Browne, Brian Veitch, Mashrura Musharraf, and Dennis Peters. Route optimization for vessels in ice: Investigating operational implications of the carbon intensity indicator regulation. Marine Policy, 158(??):Article 105858, December 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003913.

Tong:2023:NPG

[TC23] Helong Tong and Shishi Chen. A new pattern of green, intelligent and resilient international shipping: Conference report. Marine Policy, 152(??):Article 105587, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001148.

Tran:2022:PFS

[TCC^{+}22] Nhuong Tran, Long Chu, Chin Yee Chan, Jeffrey Peart, Ahmed M. Nasr-Allah, and Harrison Charo-Karisa. Prospects of fish supply–demand and its implications for food and nutrition security in Egypt. Marine Policy, 146(??):Article 105333, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003803.

Turner:2020:MMP

[TCD^{+}20] Phillip J. Turner, Sophie Cannon, Sarah DeLand, James P. Delgado, David Eltis, Patrick N. Halpin, Michael I. Kanu, Charlotte S. Sussman, Ole Varmer, and Cindy L. Van Dover. Memorializing the Middle Passage on the Atlantic seabed in Areas Beyond National Jurisdiction. Marine Policy, 122(??):Article 104254, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309003.

Tran:2020:PPR

[TCS^{+}20] Nhuong Tran, Quyen Le Cao, Kelvin Mashisia Shikuku, Thanh Phuong Phan, and Lauren K. Banks. Profitability and perceived resilience benefits of integrated shrimptilapia-seaweed aquaculture in North Central Coast, Vietnam. Marine Policy, 120(??):Article 104153, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20300725.

Ting:2021:IPP

[TCT21] Kuo-Huan Ting, Chung-Ting Cheng, and Hou-Yi Ting. Introducing the problem/project based learning as a learning strategy in University Social Responsibility Program — a study of local revitalization of Coastal Area, Yong-An District of Kaohsiung City. Marine Policy, 131(??):Article 104546, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001573.

Tang:2021:ICC

[TCZ21] Yi Tang, Wenjin Chen, and Yanxuedan Zhang. International cooperation and coordination in the Global Legislation of High Seas ABMTs including MPAs: Taking OSPAR practice as reference. Marine Policy, 133 (??):Article 104767, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2100378X.

Thiebot:2021:PMB

[TD21] Jean-Baptiste Thiebot and Magali Dreyfus. Protecting marine biodiversity beyond national jurisdiction: a penguins’ perspective. Marine Policy, 131(??):Article 104640, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002517.

Tortolini:2021:FCS

[TDC21] Virginia Mailen Tortolini, Mariana Degrati, and Mariano Alberto Coscarella. Framing and communicating southern right whale — kelp gull biological interaction in Peninsula Valdés, Argentina: the effects of attribute frames on human’s perceptions and decision-making policies. Marine Policy, 124(??):Article 104314, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309611.

Tsikliras:2020:AUT

[TDP20] Athanassios C. Tsikliras, Donna Dimarchopoulou, and Androniki Pardalou. Artificial upward trends in Greek marine landings: a case of presentist bias in European fisheries. Marine Policy, 117(??):Article 103886, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307377.

Teampau:2020:TPC

[Tea20] Petruţa Teampău. Trouble in paradise: Competing discourses and complex governance in the Romanian Danube delta. Marine Policy, 112(??):Article 103522, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18303737.

Tranter:2022:IFT

[TEA^{+}22] Sylvie N. Tranter, Estradivari, Gabby N. Ahmadia, Dominic A. Andradi-Brown, Dominic Muenzel, Firdaus Agung, Amkieltiela, Amanda K. Ford, Abdullah Habibi, Christian N. Handayani, Mohamad Iqbal, Nils C. Krueck, Muhammad E. Lazuardi, Umi Muawanah, Renoldy L. Papilaya, Tries B. Razak, Agus Sapari, Fikri F. Sjahruddin, Laura Veverka, Safran Yusri, and Maria Beger. The inclusion of fisheries and tourism in marine protected areas to support conservation in Indonesia. Marine Policy, 146(??):Article 105301, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003487.

Teijo:2023:CSI

[Tei23] Carlos Teijo. Criminal sanctions and IUU fishing: the case of Spain. Marine Policy, 147(??):Article 105341, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003888.

Telesetsky:2023:CMC

[Tel23a] Anastasia Telesetsky. Cellular mariculture: Challenges of delivering sustainable protein security. Marine Policy, 147(??):Article 105400, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200447X.

Telesetsky:2023:CRU

[Tel23b] Anastasia Telesetsky. Cross regime UNCLOS and UN-FCCC cooperation to address loss and damage from climate-shifted transboundary fisheries. Marine Policy, 148(??):Article 105426, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004730.

Teo:2023:OMO

[Teo23] Jotham Teo. Ocean materialities and the ontologies of ocean conservation in Areas Beyond National Jurisdiction: the Costa Rica Thermal Dome. Marine Policy, 153(??):Article 105646, July 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001732.

Testa:2020:NPD

[Tes20] David Testa. A note on the potential designation of the Mediterranean Sea as a sulphur emission control area. Marine Policy, 121(??):Article 104145, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20305765.

Trueman:2022:TCN

[TFF22] Justin D. Trueman, Ramón Filgueira, and Lucia Fanning. Transparency and communication in Norwegian and Nova Scotian Atlantic salmon aquaculture industries. Marine Policy, 138(??):Article 104958, April 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000057.

Teneva:2023:SIN

[TFH^{+}23] Lida Teneva, Christopher M. Free, Andrew Hume, Vera N. Agostini, Carissa J. Klein, Reg A. Watson, and Steven D. Gaines. Small island nations can achieve food security benefits through climate-adaptive blue food governance by 2050. Marine Policy, 151(??):Article 105577, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001045.

Turschwell:2023:SAS

[TFMT^{+}23] M. P. Turschwell, E. A. Fulton, J. Melbourne-Thomas, M. Lacharité, K. R. Hayes, G. Wood, K. Evans, D. Hatton MacDonald, J. Dambacher, R. H. Bustamante, R. Abbassi, P. Fidelman, and C. J. Brown. A synthesis of approaches to support integrated assessments of hazards for the emerging blue economy. Marine Policy, 155(??):Article 105696, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002233.

Touron-Gardic:2022:SPV

[TGHF^{+}22] Gregoire Touron-Gardic, Øystein Hermansen, Pierre Failler, Abdou Daïm Dia, Mohamed Ould Lemine Tarbia, Khallahi Brahim, Andy Thorpe, El Hadj Bara Dème, Eli Beibou, Elimane Abou Kane, Moustapha Bouzouma, and Juliana Arias-Hansen. The small pelagics value chain in Mauritania — recent changes and food security impacts. Marine Policy, 143(??):Article 105190, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002378.

Tudela:2022:RPB

[TGM22] Alberto Tudela, Vincent Giroud, and Llibori Martinez. Rebuttal to “Plenty of black money: Netnography of illegal recreational underwater fishing in southern Spain” Maya-Jariego et al., Marine Policy 126 (2021). Marine Policy, 144(??):Article 105235, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002822. See response [MJ23].

Titterton:2022:KPI

[TH22] Haydn Titterton and Marcus Haward. The Kerguelen Plateau: Interactions between the Law of the Sea and the Antarctic Treaty. Marine Policy, 138(??):Article 104993, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000409.

Thayer:2020:IEC

[THGR^{+}20] J. A. Thayer, E. L. Hazen, M. García-Reyes, A. Szoboszlai, and W. J. Sydeman. Implementing ecosystem considerations in forage fisheries: San Francisco Bay herring case study. Marine Policy, 118(??):Article 103884, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304051.

Thompson:2022:BBM

[Tho22] Benjamin S. Thompson. Blue bonds for marine conservation and a sustainable ocean economy: Status, trends, and insights from green bonds. Marine Policy, 144(??):Article 105219, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002664.

Teniwut:2022:TSG

[THP22a] Wellem Anselmus Teniwut, Cawalinya Livsanthi Hasyim, and Frischilla Pentury. Towards smart government for sustainable fisheries and marine development: an intelligent web-based support system approach in small islands. Marine Policy, 143(??):Article 105158, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002056.

Thorpe:2022:UTT

[THP^{+}22b] Andy Thorpe, Oystein Hermansen, Iain Pollard, John Isaksen, Pierre Failler, and Grégoire Touron-Gardic. Unpacking the tuna traceability mosaic — EU SFPAs and the tuna value chain. Marine Policy, 139(??):Article 105037, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000847.

Tsvetkova:2024:TTC

[THS^{+}24] Anastasia Tsvetkova, Magnus Hellström, Henry Schwartz, Rodrigo Rabetino, and Hammad Syed. A transition towards clean propulsion in shipping: the role of PES-TLE drivers and implications for policy. Marine Policy, 161(??):Article 106002, March 2024. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23005353.

Teniwut:2023:DCC

[THT^{+}23] Wellem Anselmus Teniwut, Syahibul Kahfi Hamid, Roberto Mario Kabi Teniwut, Maimuna Renhoran, and Christy Desta Pratama. Do coastal communities in small islands value marine resources through marine protected areas?: Evidence from Kei Islands Indonesia with choice modelling. Marine Policy, 157(??):Article 105838, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003718.

Thurlow:2021:GAS

[TJ21a] Grace Thurlow and Peter J. S. Jones. A governance analysis of Soufriere Marine Management Area, Saint Lucia: Previously effective but increasingly challenged by driving forces. Marine Policy, 127(??):Article 104220, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308666.

Trouillet:2021:CRB

[TJ21b] Brice Trouillet and Stephen Jay. The complex relationships between marine protected areas and marine spatial planning: Towards an analytical framework. Marine Policy, 127(??):Article 104441, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000518.

Tyabji:2022:UTS

[TJS22] Zoya Tyabji, Rima W. Jabado, and Dipani Sutaria. Utilization and trade of sharks and rays in the Andaman Islands, India. Marine Policy, 146(??):Article 105295, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003426.

Twumasi:2021:IGF

[TJZ^{+}21] Martinson Ankrah Twumasi, Yuansheng Jiang, Xiaoshi Zhou, Bismark Addai, Kwabena Nkansah Darfor, Selorm Akaba, and Prince Fosu. Increasing Ghanaian fish farms’ productivity: Does the use of the Internet matter? Marine Policy, 125(??):Article 104385, March 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310368.

Tseng:2022:TRF

[TK22] Huan-Sheng Tseng and Shih-Ming Kao. Towards responsible fisheries: Management of gillnet fishing in Taiwan. Marine Policy, 141(??):Article 105098, July 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001452.

Tseng:2023:ODM

[TK23] Hui-Yi Katherine Tseng and Shih-Ming Kao. Oceanographic data management in Taiwan: an appraisal under the transnational legal ordering perspective. Marine Policy, 155(??):Article 105782, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003159.

Tian:2023:MTL

[TKC23] Huimin Tian, Mehran Idris Khan, and Yen-Chiang Chang. Maritime traffic law of the People’s Republic of China and its implications in international law. Marine Policy, 151(??):Article 105570, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000970.

Tokunaga:2023:IFA

[TKP23] Kanae Tokunaga, Lisa A. Kerr, and Andrew J. Pershing. Implications of fisheries allocation policy on anticipated climate change impacts. Marine Policy, 148(??):Article 105402, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004493.

Tian:2022:AUN

[TLT^{+}22] Hao Tian, Yang Liu, Yongjun Tian, Youfu Jing, Shuhao Liu, Xudong Liu, and Yong Zhang. Advances in the use of nighttime light data to monitor and assess coastal fisheries under the impacts of human activities and climate and environmental changes: a case study in the Beibu Gulf. Marine Policy, 144(??):Article 105227, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002743.

Tang:2020:CPD

[TLYC20] Jianye Tang, Yunzhou Li, Xiaoming Yang, and Yong Chen. Can the presidential diplomacy of China and France promote the East Antarctic marine protected area? Marine Policy, 118(??):Article 104002, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301299.

Tanaka:2023:EEM

[TM23a] Hajime Tanaka and Mitsutaku Makino. Economic evaluation of MSY-based fishery policy using input–output table: a case study of squid-related industries in Hakodate City, Japan. Marine Policy, 157(??):Article 105843, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003767.

Tiller:2023:IBA

[TM23b] Rachel Tiller and Elizabeth Mendenhall. And so it begins — the adoption of the ‘Biodiversity Beyond National Jurisdiction’ treaty. Marine Policy, 157(??):Article 105836, November 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2300369X.

Tiller:2023:SIN

[TMDN23] Rachel Tiller, Elizabeth Mendenhall, Elizabeth De Santo, and Elizabeth Nyman. Shake it off: Negotiations suspended, but hope simmering, after a lack of consensus at the fifth intergovernmental conference on biodiversity beyond national jurisdiction. Marine Policy, 148(??):Article 105457, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22005048.

Tunnicliffe:2021:SEG

[TML^{+}21] Verena Tunnicliffe, Anna Metaxas, Jennifer Le, Eva Ramirez-Llodra, and Lisa A. Levin. Strategic Environmental Goals and Objectives: Setting the basis for environmental regulation of deep seabed mining. Marine Policy, ??(??):Article 103347, ???? 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830321X.

Toupin:2023:MES

[TMM^{+}23] Rémi Toupin, Kory Melnick, Tamanna Moharana, Blake Curry, Catherine Gracey, Naomi Richards, Bertrum H. MacDonald, and Philippe Mongeon. Measuring engagement with shark research in science, news and social media, and policy: a bibliometric analysis. Marine Policy, 157(??):Article 105829, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003627.

Trialfhianty:2022:SMC

[TMS^{+}22] Tyas Ismi Trialfhianty, Fajrun Wahidil Muharram, Suadi, Claire Helen Quinn, and Maria Beger. Spatial multi-criteria analysis to capture socio-economic factors in mangrove conservation. Marine Policy, 141(??):Article 105094, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001415.

Tissiere:2020:FES

[TMT20] Laurie Tissière, Stéphanie Mahévas, and Brice Trouillet. Findings from an exploratory study on the governance of a French fishery. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16307072.

Trebilco:2020:PRS

[TMTC20] Rowan Trebilco, Jess Melbourne-Thomas, and Andrew John Constable. The policy relevance of Southern Ocean food web structure: Implications of food web change for fisheries, conservation and carbon sequestration. Marine Policy, 115(??):Article 103832, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X1930212X.

Torabi:2021:LRR

[TN21] Foroogh Torabi and Seyed Mohammad Tababaye Nejad. Legal regime of residual liability in decommissioning: the importance of role of states. Marine Policy, 133(??):Article 104727, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003389.

Toygar:2023:WFC

[TNY23] Arda Toygar, Senem Nart, and Umut Yıldırım. Workfamily conflict and work alienation among seafarers: the mediating role of emotional exhaustion. Marine Policy, 155(??):Article 105759, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002920.

Tanner:2021:SDS

[TOAPJ21] Michael K. Tanner, Marcelo Olivares-Arenas, Luciana Puebla, and Jose R. Marin Jarrin. Shifting demand to sustainable fishing practices in Darwin’s Archipelago: a discrete choice experiment application for Galapagos’ certified yellow-fin tuna. Marine Policy, 132(??):Article 104665, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002761.

Todorov:2020:RMA

[Tod20] Andrey Todorov. Russia in maritime areas off Spitsbergen (Svalbard): Is it worth opening the Pandora’s Box? Marine Policy, 122(??):Article 104264, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309106.

Todorov:2022:DSR

[Tod22a] Andrey Todorov. Dire straits of the Russian Arctic: Options and challenges for a potential US FONOP in the Northern Sea Route. Marine Policy, 139(??):Article 105020, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000677.

Todorov:2022:SGB

[Tod22b] Andrey Todorov. Shipping governance in the Bering Strait Region: Protecting the Diomede Islands and adjacent waters. Marine Policy, 146(??):Article 105289, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003360.

Tsakiridis:2020:CEE

[TOHK20] Andreas Tsakiridis, Cathal O’Donoghue, Stephen Hynes, and Kevin Kilcline. A comparison of environmental and economic sustainability across seafood and livestock product value chains. Marine Policy, 117(??):Article 103968, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19305561.

Tindall:2022:IMM

[TOL^{+}22] Charlotte Tindall, Oluyemisi Oloruntuyi, Samantha Lees, Catherine S. Longo, David Schley, and Rohan J. C. Currey. Illuminating the mechanisms to mitigate forced and child labour risks within Marine Stewardship Council certified fisheries. Marine Policy, 143(??):Article 105140, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001877.

Tong:2022:FDR

[Ton22a] Helong Tong. The future development and restructuring of the international shipping industry: Conference report. Marine Policy, 137(??):Article 104956, March 2022.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000033.

Tong:2022:SDW

[Ton22b] Helong Tong. Sustainable development and win-win cooperation in Asian marine tourism: Conference report. Marine Policy, 146(??):Article 105304, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003517.

Tukana:2023:BSF

[TPG^{+}23] Max Tukana, Jeremy Prince, Kerstin B. J. Glaus, Kalisiana Marama, and Cherie Whippy-Morris. A baseline study of Fiji’s small-scale lobster fishery using value chain analysis and size at maturity thresholds. Marine Policy, 149(??):Article 105513, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000404.

Trathan:2023:FSG

[Tra23a] Philip N. Trathan. The future of the South Georgia and South Sandwich Islands marine protected area in a changing environment: the choice between industrial fisheries, or ecosystem protection. Marine Policy, 155(??):Article 105773, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003068.

Trathan:2023:WNI

[Tra23b] Philip N. Trathan. What is needed to implement a sustainable expansion of the Antarctic krill fishery in the Southern Ocean? Marine Policy, 155(??):Article 105770, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003032.

Trudo:2023:PMP

[Tru23] Karin Otsuka Trudo. Perceptions of marine pollution when intertwined with displacement: the case of Miyakojima, Okinawa Prefecture, Japan. Marine Policy, 155 (??):Article 105749, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002828.

Teneva:2022:EPO

[TSA^{+}22] Lida Teneva, Aaron L. Strong, Vera Agostini, Kenneth J. Bagstad, Evangelia G. Drakou, Zachary Ancona, Kristina Gjerde, Andrew C. Hume, and Nicholas Jickling. Estimating the pelagic ocean’s benefits to humanity can enhance ocean governance. Marine Policy, 136(??):Article 104906, February 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21005170.

Tseng:2021:MLE

[Tse21] Huan-Sheng Tseng. Maritime law enforcement on bordercrossing fishing vessels: Experience of Taiwan’s practices under its unique political relationship. Marine Policy, 133(??):Article 104749, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003602.

Temple:2022:IUU

[TSH^{+}22a] Andrew J. Temple, Daniel J. Skerritt, Philippa E. C. Howarth, John Pearce, and Stephen C. Mangi. Illegal, unregulated and unreported fishing impacts: a systematic review of evidence and proposed future agenda. Marine Policy, 139(??):Article 105033, May 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200080X.

Thorvaldsen:2022:IFM

[TSH22b] Trine Thorvaldsen, Signe A. Sønvisen, and Ingunn M. Holmen. The impact of fisheries management on fishers’ health and safety: a case study from Norway. Marine Policy, 140(??):Article 105066, June 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001130.

Thorbjornsen:2023:HCC

[TSLK23] Susanna Huneide Thorbjørnsen, Ann-Elin Wårøy Synnes, Ingrid Disch Løset, and Alf Ring Kleiven. Hazard and catch composition of ghost fishing gear revealed by a citizen science clean-up initiative. Marine Policy, 148(??):Article 105431, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200478X.

Taniu:2024:IEC

[TSS^{+}24] Silviana Taniu, Dyah Wulan Sari, Dias Satria, Tri Haryanto, and Wahyu Wisnu Wardana. Impact evaluation of cooperative membership on welfare: Evidence from captured fishery households in Indonesia. Marine Policy, 159(??):Article 105923, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004566.

Tsai:2023:OLA

[TSWC23] Liang-Ting Tsai, Tsuyoshi Sasaki, Chin-Kuo Wu, and Cheng-Chieh Chang. Ocean literacy among Taiwanese and Japanese high school students. Marine Policy, 150(??): Article 105555, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X23000829.

Tirumala:2022:IFM

[TT22] Raghu Dharmapuri Tirumala and Piyush Tiwari. Innovative financing mechanism for blue economy projects. Marine Policy, 139(??):Article 104194, May 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20303262.

Teh:2020:SSF

[TTA^{+}20] Lydia C. L. Teh, Louise S. L. Teh, Keita Abe, Gakushi Ishimura, and Raphael Roman. Small-scale fisheries in developed countries: Looking beyond developing country narratives through Japan’s perspective. Marine Policy, 122(??):Article 104274, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309209.

Talib:2022:TCM

[TUBA22] Naimah Lutfi Talib, Ariane Utomo, Jon Barnett, and Dedi Supriadi Adhuri. Three centuries of marine governance in Indonesia: Path dependence impedes sustainability. Marine Policy, 143(??):Article 105171, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002184.

Tolochko:2021:USD

[TV21] Petro Tolochko and Alice B. M. Vadrot. The usual suspects? Distribution of collaboration capital in marine biodiversity research. Marine Policy, 124(??):Article 104318, February 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20309659.

Tran:2023:AHI

[TVN23] Tuyen Quang Tran, Huong Van Vu, and Thanh Viet Nguyen. Aquaculture, household income and inequality in Vietnam’s coastal region. Marine Policy, 153(??):Article 105634, July 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001616.

Townsend:2022:EDV

[TW22] Ralph E. Townsend and Scott Walker. The economics of deemed values. Marine Policy, 142(??):Article 105105, August 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2200152X.

Tsujita:2022:IIC

[TYS22] Takahiro Tsujita, Nobuyuki Yagi, and Yutaro Sakai. Industry-initiated catch limit management: the case of purse seine fishery in Japan’s EEZ waters of the North Pacific Ocean. Marine Policy, 140(??):Article 105053, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001002.

Tajima:2024:CBF

[TYS24] Satoshi Tajima, Kensuke Yamaguchi, and Hideaki Shiroyama. Consensus building with fishermen on offshore wind farms in Japan: Current status and policy recommendations. Marine Policy, 160(??):Article 105975, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23005080.

Tang:2023:ECB

[TZC23] Hao Tang, Khadija Zulfiqar, and Yen-Chiang Chang. Ecological civilization building and the law of the sea: Conference report. Marine Policy, 157(??):Article 105871, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004049.

Tanaka:2022:SDM

[TZM^{+}22] Kotaro Tanaka, Mengyao Zhu, Kohei Miyaji, Tadayuki Kurokawa, and Tomonari Akamatsu. Spatial distribution maps of real-time ocean observation platforms and sensors in Japanese waters. Marine Policy, 141(??):Article 105102, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2200149X.

Teng:2021:IMF

[TZZ^{+}21] Xin Teng, Qiwei Zhao, Panpan Zhang, Liang Liu, Yue’e Dong, Heng Hu, Qi Yue, Ling Ou, and Wei Xu. Implementing marine functional zoning in China. Marine Policy, 132(??):Article 103484, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304093.

Uriarte:2023:LLM

[UISM^{+}23] Andrés Uriarte, Leire Ibaibarriaga, Sonia Sánchez-Maroño, Pablo Abaunza, Marga Andrés, Erwan Duhamel, Ernesto Jardim, Lionel Pawlowski, Raúl Prellezo, and Beatriz A. Roel. Lessons learnt on the management of short-lived fish from the Bay of Biscay anchovy case study: Satisfying fishery needs and sustainability under recruitment uncertainty. Marine Policy, 150(??):Article 105512, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000398.

Uduji:2020:DCS

[UOO20] Joseph Ikechukwu Uduji and Elda Nduka Okolo-Obasi. Does corporate social responsibility (CSR) impact on development of women in small-scale fisheries of sub-Saharan Africa? Evidence from coastal communities of Niger Delta in Nigeria. Marine Policy, 118(??):??, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306195.

Urbina:2022:TIL

[Urb22] Julio Jorge Urbina. Towards an international legal definition of the notion of fisheries crime. Marine Policy, 144(??):Article 105214, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002615.

Vilas:2020:UCV

[VAMR^{+}20] C. Vilas, L. T. Antelo, F. Martin-Rodriguez, X. Morales, R. I. Perez-Martin, A. A. Alonso, J. Valeiras, E. Abad, M. Quinzan, and M. Barral-Martinez. Use of computer vision onboard fishing vessels to quantify catches: the iObserver. Marine Policy, 116(??):Article 103714, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308340.

Visalli:2020:DDA

[VBC^{+}20] Morgan E. Visalli, Benjamin D. Best, Reniel B. Cabral, William W. L. Cheung, Nichola A. Clark, Cristina Garilao, Kristin Kaschner, Kathleen Kesner-Reyes, Vicky W. Y. Lam, Sara M. Maxwell, Juan Mayorga, Holly V. Moeller, Lance Morgan, Guillermo Ortuño Crespo, Malin L. Pinsky, Timothy D. White, and Douglas J. McCauley. Datadriven approach for highlighting priority areas for protection in marine areas beyond national jurisdiction. Marine Policy, 122(??):Article 103927, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309194.

Filho:2021:RFP

[VBd^{+}21] Márcio Luiz Vargas Barbosa Filho, Gabriel Barros Gonçalves de Souza, Sérgio de Faria Lopes, Rachel Ann Hauser-Davis, Salvatore Siciliano, and José da Silva Mourão. Reef fisher perceptions acknowledge the socioenvironmental effectiveness of a 20-year old Brazilian Marine Protected Area. Marine Policy, 134(??):Article 104797, December 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004085.

Verling:2023:ARB

[VBH^{+}23] Emma Verling, Cátia Bartilotti, Claudia Hollatz, Miriam Tuaty-Guerra, Jorge Lobo-Arteaga, and Tim O’Higgins. Applying risk-based approaches to implementation of the Marine Strategy Framework Directive in the North-East Atlantic: Learning lessons and moving forward. Marine Policy, 153(??):Article 105667, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300194X.

VanDover:2020:RNI

[VCC^{+}20] C. L. Van Dover, A. Colaço, P. C. Collins, P. Croot, A. Metaxas, B. J. Murton, A. Swaddling, R. E. Boschen-Rose, J. Carlsson, L. Cuyvers, T. Fukushima, A. Gartman, R. Kennedy, C. Kriete, N. C. Mestre, T. Molodtsova, A. Myhrvold, E. Pelleter, S. O. Popoola, P.-Y. Qian, J. Sarrazin, R. Sharma, Y. J. Suh, J. B. Sylvan, C. Tao, M. Tomczak, and J. Vermilye. Research is needed to inform environmental management of hydrothermally inactive and extinct polymetallic sulfide (PMS) deposits. Marine Policy, 121(??):Article 104183, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302943.

Vilas:2021:CPC

[VCC^{+}21] Daniel Vilas, Marta Coll, Xavier Corrales, Jeroen Steenbeek, Chiara Piroddi, Diego Macias, Alessandro Ligas, Paolo Sartor, and Joachim Claudet. Current and potential contributions of the Gulf of Lion Fisheries Restricted Area to fisheries sustainability in the NW Mediterranean Sea. Marine Policy, 123(??):Article 104296, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309416.

vandenBurg:2023:MEM

[vdBSvdV^{+}23] S. W. K. van den Burg, M. Skirtun, O. van der Valk, W. Rossi Cervi, T. Selnes, T. Neumann, J. Steinmann, G. Arora, and P. Roebeling. Monitoring and evaluation of maritime spatial planning — a review of accumulated practices and guidance for future action. Marine Policy, 150(??):Article 105529, April 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000568.

vanderGrient:2021:PSI

[vdGD21] J. M. A. van der Grient and J. C. Drazen. Potential spatial intersection between high-seas fisheries and deep-sea mining in international waters. Marine Policy, 129(??):Article 104564, July 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001755.

Vecchio:2022:PHR

[VDP^{+}22] Y. Vecchio, J. Di Pasquale, G. Pauselli, M. Masi, and F. Adinolfi. Public health risk management during the Covid-19 pandemic, new amendments in the European Maritime and Fisheries Fund to meet fishers’ needs. Marine Policy, 135(??):Article 104873, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100484X.

Villegas:2021:CCS

[VGAH^{+}21] Ciera Villegas, Nicolás X. Gómez-Andújar, Michael Harte, Sarah M. Glaser, and James R. Watson. Cooperation and conflict in the small-scale fisheries of Puerto Rico. Marine Policy, 134(??):Article 104809, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004206.

Villasante:2022:SES

[VGC^{+}22] Sebastian Villasante, Ignacio Gianelli, Mauricio Castrejón, Laura Nahuelhual, Leonardo Ortega, U. Rashid Sumaila, and Omar Defeo. Social-ecological shifts, traps and collapses in small-scale fisheries: Envisioning a way forward to transformative changes. Marine Policy, 136(??):Article 104933, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21005443.

Vukic:2023:APE

[VGK23] Luka Vukić, Giambattista Guidi, and Iven Krämer. Air pollutant emission calculation and shipping costs: the case of yacht transport between the North Sea and Mediterranean ports. Marine Policy, 148(??):Article 105447, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004948.

Valdor:2020:GAM

[VGS^{+}20] Paloma F. Valdor, Aina G. Gómez, Peter Steinberg, Edwina Tanner, Antony M. Knights, Rochelle D. Seitz, Laura Airoldi, Louise B. Firth, Christos Arvanitidis, Massimo Ponti, Eva Chatzinikolaou, Paul R. Brooks, Tasman P. Crowe, Alison Smith, Gonzalo Méndez, Aida Ovejero, Abilio Soares-Gomes, John A. Burt, Catriona MacLeod, and José A. Juanes. A global approach to mapping the environmental risk of harbours on aquatic systems. Marine Policy, 119(??):Article 104051, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302684.

Vierros:2020:CBT

[VHD20] Marjo K. Vierros and Harriet Harden-Davies. Capacity building and technology transfer for improving governance of marine areas both beyond and within national jurisdiction. Marine Policy, 122(??):Article 104158, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20301469.

Vierros:2020:CIP

[VHS^{+}20] Marjo K. Vierros, Autumn-Lynn Harrison, Matthew R. Sloat, Guillermo Ortuño Crespo, Jonathan W. Moore, Daniel C. Dunn, Yoshitaka Ota, Andrés M. Cisneros-Montemayor, George L. Shillinger, Trisha Kehaulani Watson, and Hugh Govan. Considering Indigenous Peoples and local communities in governance of the global ocean commons. Marine Policy, 119(??):Article 104039, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19309212.

vanHoof:2020:CPC

[vHSSK20] L. van Hoof, N. A. Steins, S. Smith, and M. Kraan. Change as a permanent condition: a history of transition processes in Dutch North Sea fisheries. Marine Policy, 122(??):Article 104245, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308915.

VanAssche:2020:GCC

[VHSV20] Kristof Van Assche, Anna-Katharina Hornidge, Achim Schlüter, and Nataşa Vaidianu. Governance and the coastal condition: Towards new modes of observation, adaptation and integration. Marine Policy, 112(??):Article 103413, February 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X18303865.

Vescovo:2021:SCD

[VJL^{+}21] Victor L. Vescovo, Alan J. Jamieson, Patrick Lahey, Rob McCallum, Heather A. Stewart, and Casey Machado.

Safety and conservation at the deepest place on Earth: a call for prohibiting the deliberate discarding of nondegradable umbilicals from deep-sea exploration vehicles. Marine Policy, 128(??):Article 104463, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000749.

Vigfusson:2023:SIO

[VJO23] Kristjan Reykjalin Vigfusson, Lara Johannsdottir, and Snjolfur Olafsson. Strategy implementation obstacles: Iceland fishery CEO perspectives. Marine Policy, 147(??): Article 105353, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X22004006.

Voldnes:2020:TCH

[VKN20] Gøril Voldnes, Ingrid Kvalvik, and Bjørg Nøstvold. Taking care of a highly valuable resource throughout the value chain — Lack of market orientation in red king crab export? Marine Policy, 117(??):Article 103965, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303914.

Vinata:2023:CCR

[VKS23] Ria Tri Vinata, Masitha Tismananda Kumala, and Cita Yustisia Serfiyani. Climate change and reconstruction of Indonesia’s geographic basepoints: Reconfiguration of baselines and Indonesian Archipelagic Sea lanes. Marine Policy, 148(??):Article 105443, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004900.

vanLeeuwen:2022:DSS

[vLM22] Judith van Leeuwen and Jason Monios. Decarbonisation of the shipping sector — time to ban fossil fuels? Marine Policy, 146(??):Article 105310, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003578.

vanLeeuwen:2021:MMS

[vLPBK21] Wessel M. A. van Leeuwen, Claire Pekcan, Mike Barnett, and Göran Kecklund. Mathematical modelling of sleep and sleepiness under various watch keeping schedules in the maritime industry. Marine Policy, 130(??):Article 104277, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309234.

Vogel:2023:DCR

[VLS^{+}23] Jacqueline M. Vogel, Catherine Longo, Jessica Spijkers, Juliano Palacios-Abrantes, Julia Mason, Colette C. C. Wabnitz, William Cheung, U. Rashid Sumaila, Gordon Munro, Sarah Glaser, Johann Bell, Yongjun Tian, Nancy L. Shackell, Elizabeth R. Selig, Philippe Le Billon, James R. Watson, Cullen Hendrix, Malin L. Pinsky, Ingrid van Putten, Kendra Karr, Eva A. Papaioannou, and Rod Fujita. Drivers of conflict and resilience in shifting transboundary fisheries. Marine Policy, 155(??):Article 105740, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002737.

Vandergeest:2021:BSS

[VM21] Peter Vandergeest and Melissa Marschke. Beyond slavery scandals: Explaining working conditions among fish workers in Taiwan and Thailand. Marine Policy, 132(??):Article 104685, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002967.

Venturini:2021:RBP

[VMC^{+}21] S. Venturini, F. Massa, M. Castellano, G. Fanciulli, and P. Povero. Recreational boating in the Portofino Marine Protected Area (MPA), Italy: Characterization and analysis in the last decade (2006–2016) and some considerations on management. Marine Policy, 127(??):Article 103178, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17306942.

Vecchio:2024:TIF

[VMD^{+}24] Yari Vecchio, Margherita Masi, Teresa Del Giudice, Marcello De Rosa, and Felice Adinolfi. Technological innovation in fisheries and aquaculture: What are the “discourses” of the Italian policy network? Marine Policy, 159(??):Article 105947, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004803.

Vandergeest:2021:SFY

[VMM21] Peter Vandergeest, Melissa Marschke, and Mallory Mac-Donnell. Seafarers in fishing: a year into the COVID-19 pandemic. Marine Policy, 134(??):Article 104796, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004073.

VanRoy:2024:PRI

[VMV^{+}24] Ward Van Roy, Jean-Baptiste Merveille, Annelore Van Nieuwenhove, Kobe Scheldeman, and Frank Maes. Policy recommendations for international regulations addressing air pollution from ships. Marine Policy, 159(??):Article 105913, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004463.

Vakili:2020:DPFa

[VÖB20a] S. V. Vakili, A. I. Ölcer, and F. Ballini. The development of a policy framework to mitigate underwater noise pollution from commercial vessels. Marine Policy, 118(??):Article 104004, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308942.

Vakili:2020:DPFb

[VÖB20b] S. V. Vakili, A. I. Ölçer, and F. Ballini. The development of a policy framework to mitigate underwater noise pollution from commercial vessels: the role of ports. Marine Policy, 120(??):Article 104132, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309480.

Virgili:2024:EFS

[VPB^{+}24] M. Virgili, A. Petetta, G. Barone, D. Li Veli, G. Bargione, and A. Lucchetti. Engaging fishers in sea turtle conservation in the Mediterranean Sea. Marine Policy, 160(??):Article 105981, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23005146.

Vilela:2021:USM

[VPRR^{+}21] Raul Vilela, Maria Grazia Pennino, Gonzalo Rodriguez-Rodriguez, Hugo M. Ballesteros, and Jose Maria Bellido. The use of a spatial model of economic efficiency to predict the most likely outcomes under different fishing strategy scenarios. Marine Policy, 129(??):Article 104499, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100110X.

vanRensburg:2024:UPP

[vRB24] Thomas M. van Rensburg and Noreen Brennan. Understanding public preferences towards Ireland’s offshore wind sector: a study on renewable energy trade, public involvement, and setback distance. Marine Policy, 160(??):Article 105988, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23005213.

Verling:2021:ARB

[VRBC^{+}21] Emma Verling, Ramon Miralles Ricós, Manuel Bou-Cabo, Guillermo Lara, Maria Garagouni, Jean-Marc Brignon, and Tim O’Higgins. Application of a risk-based approach to continuous underwater noise at local and subregional scales for the Marine Strategy Framework Directive. Marine Policy, 134(??):Article 104786, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003973.

Vaughan:2023:RFT

[VSD^{+}23] Duncan Vaughan, Daniel J. Skerritt, James Duckworth, U. Rashid Sumaila, and Mark Duffy. Revisiting fuel tax concessions (FTCs): the economic implications of fuel subsidies for the commercial fishing fleet of the United Kingdom. Marine Policy, 155(??):Article 105763, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002968.

Villasante:2022:SEU

[VSDR^{+}22] Sebastián Villasante, U. Rashid Sumaila, Jose María Da-Rocha, Natacha Carvalho, Daniel J. Skerritt, Anna Schuhbauer, Andrés M. Cisneros-Montemayor, Nathan J. Bennett, Quentin Hanich, and Raúl Prellezo. Strengthening European Union fisheries by removing harmful subsidies. Marine Policy, 136(??):Article 104884, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004954.

vonSchuckmann:2020:OSD

[vSHHT20] Karina von Schuckmann, Elisabeth Holland, Peter Haugan, and Peter Thomson. Ocean science, data, and services for the UN 2030 Sustainable Development Goals. Marine Policy, 121(??):Article 104154, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302050.

Vergara-Solana:2022:VVM

[VSPMS^{+}22] Francisco Vergara-Solana, Daniel Peñalosa-Martinell, Daniel Skerritt, Annie Mejaes, German Ponce-Diaz, Fernando Aranceta-Garza, Fernando González-Laxe, Juan Carlos Seijo, and U. Rashid Sumaila. Volatility and vulnerability in Mexican fisheries and aquaculture: Enhancing resilience via public policy. Marine Policy, 136(??):Article 104888, February 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004991.

VanRoy:2023:CPD

[VSV^{+}23] Ward Van Roy, Kobe Scheldeman, Annelore Van Nieuwenhove, Jean-Baptiste Merveille, Ronny Schallier, and Frank Maes. Current progress in developing a MARPOL Annex VI enforcement strategy in the Bonn Agreement through remote measurements. Marine Policy, 158(??):Article 105882, December 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004153.

vonThenen:2021:GMP

[vTHS21] Miriam von Thenen, Henning Sten Hansen, and Kerstin S. Schiele. A generalised marine planning framework for site selection based on ecosystem services. Marine Policy, 124(??):Article 104326, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309738.

Vaidianu:2020:MCP

[VTRS20] Nataşa Văidianu, Florin Tătui, Mădălina Ristea, and Adrian Stănică. Managing coastal protection through multi-scale governance structures in Romania. Marine Policy, 112(??):Article 103567, February 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304123.

Virdin:2022:CIF

[VVR^{+}22] John Virdin, Tibor Vegh, Blake Ratcliff, Elizabeth Havice, Jack Daly, and Jack Stuart. Combatting illegal fishing through transparency initiatives: Lessons learned from comparative analysis of transparency initiatives in seafood, apparel, extractive, and timber supply chains. Marine Policy, 138(??):Article 104984, April 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000318.

Vrooman:2022:UDR

[VvSvH^{+}22] Jip Vrooman, Christiaan van Sluis, Floris van Hest, Han Lindeboom, and Albertinka Murk. Unambiguously defined and recognized seabed protection targets are necessary for successful implementation of MPAs. Marine Policy, 140(??):Article 105056, June 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001038.

VanDiggelen:2022:CLL

[VWFW22] Amanda D. Van Diggelen, Sara E. Worden, Adam J. Frimodig, and Stephen P. Wertz. California’s lessons learned and recommendations for effective marine protected area network management. Marine Policy, 137(??): Article 104928, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X2100539X.

Vye:2020:EPM

[VWJMAJ20] Siobhan R. Vye, Sophie Wynne-Jones, Patricia Masterson-Algar, and Stuart R. Jenkins. Exploring perceptions of marine biosecurity interventions: insights from the commercial marina sector. Marine Policy, 118(??):Article 104027, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19302222.

Vylegzhanin:2020:CAO

[VYB20] Alexander N. Vylegzhanin, Oran R. Young, and Paul Arthur Berkman. The Central Arctic Ocean Fisheries Agreement as an element in the evolving Arctic Ocean governance complex. Marine Policy, 118(??):Article 104001, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301780.

Wang:2024:GAM

[WA24] Weishan Wang and Claudio Aporta. Arctic marine shipping development and governance in Canada: a historical overview. Marine Policy, 160(??):Article 105958, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004918.

Weaver:2022:API

[WABR^{+}22] P. P. E. Weaver, J. Aguzzi, R. E. Boschen-Rose, A. Colaço, H. de Stigter, S. Gollner, M. Haeckel, C. Hauton, R. Helmons, D. O. B. Jones, H. Lily, N. C. Mestre, C. Mohn, and L. Thomsen. Assessing plume impacts caused by polymetallic nodule mining vehicles. Marine Policy, 139(??):Article 105011, May 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000586.

Wang:2023:SCM

[Wan23a] Chuanliang Wang. The significance of the concept of a maritime community with a shared future in the international law-making process: Taking the BBNJ Agreement as an example. Marine Policy, 149(??):Article 105509, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000362.

Wang:2023:ILM

[Wan23b] Sen Wang. International law-making process of combating plastic pollution: Status Quo, debates and prospects. Marine Policy, 147(??):Article 105376, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004237.

Wang:2023:URG

[Wan23c] Weishan Wang. UNDRIP rights to guide the governance of the Northern Low-impact Shipping Corridors initiative. Marine Policy, 155(??):Article 105737, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002701.

Warner:2021:IEL

[War21] Robin Warner. International environmental law principles relevant to exploitation activity in the Area. Marine Policy, ??(??):Article 103503, ???? 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303609.

Ward:2023:FAP

[War23] Sarah Ward. The Fourth Asia Pacific Regional Conference on Underwater Cultural Heritage: a marine policy perspective. Marine Policy, 148(??):Article 105358, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004055.

Waters:2022:EBD

[Wat22] Yolanda L. Waters. Establishing a baseline decision context for restoring the Great Barrier Reef — implications for longterm management. Marine Policy, 141(??):Article 105106, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001531.

Willis:2023:HCG

[WBH^{+}23] Sam Willis, Despoina Andrioti Bygvraa, Md. Sazedul Hoque, Emily S. Klein, Cagri Kucukyildiz, Joseph Westwood-Booth, and Eric Holliday. The human cost of global fishing. Marine Policy, 148(??):Article 105440, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22004870.

Wilson:2020:BPF

[WBPG20] Jono R. Wilson, Darcy Bradley, Kristina Phipps, and Mary G. Gleason. Beyond protection: Fisheries cobenefits of no-take marine reserves. Marine Policy, 122(??):Article 104224, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20308708.

Wang:2020:IEI

[WCC^{+}20] Grace Wang, Wen-Huei Chang, Yue Cui, Guanqiu Qi, and Kevin X. Li. Introducing an economic impact platform to navigate cruise value-added chain with environmental considerations. Marine Policy, 112(??):Article 103713, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19303331.

Weiss:2021:MAE

[WCF^{+}21] Sarah G. Weiss, Danielle Cholewiak, Kaitlin E. Frasier, Jennifer S. Trickey, Simone Baumann-Pickering, John A. Hildebrand, and Sofie M. Van Parijs. Monitoring the acoustic ecology of the shelf break of Georges Bank, Northwestern Atlantic Ocean: New approaches to visualizing complex acoustic data. Marine Policy, 130(??):Article 104570, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001810.

Wood:2024:MNI

[WCM^{+}24] Louisa E. Wood, Stacey A. Clarke, Debbie Murphy, Phil I. Davison, Mark A. Thrush, David Bass, Silvana N. R. Birchenough, Edmund J. Peeler, and Hannah J. Tidbury. Monitoring of non-indigenous marine species for legislative and policy goals in the UK. Marine Policy, 162(??):Article 106027, April 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24000253.

Wilson:2020:ERI

[WCT^{+}20] Susan C. Wilson, Imogen Crawford, Irina Trukhanova, Lilia Dmitrieva, and Simon J. Goodman. Estimating risk to icebreeding pinnipeds from shipping in Arctic and sub-Arctic seas. Marine Policy, 111(??):Article 103694, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302039.

Wang:2021:MEW

[WCZS21] Shuhong Wang, Suisui Chen, Hongyan Zhang, and Malin Song. The model of early warning for China’s marine ecology-economy symbiosis security. Marine Policy, 128(??):Article 104476, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21000877.

Ware:2020:CHM

[WD20] Suzanne Ware and Anna-Leena Downie. Challenges of habitat mapping to inform marine protected area (MPA) designation and monitoring: an operational perspective. Marine Policy, 111(??):Article 103717, January 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301022.

Wood-Donnelly:2022:IS

[WD22] Corine Wood-Donnelly. Iceberg sovereignty. Marine Policy, 143(??):Article 105139, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001865.

Walden:2023:BEP

[WD23] John B. Walden and Geret DePiper. Benchmarking ecosystem performance using index numbers: a case study of the northeast large marine ecosystem. Marine Policy, 155(??):Article 105708, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002415.

Williams:2020:ECS

[WDC^{+}20] Chris Williams, William Davies, Robert E. Clark, Angela Muench, and Kieran Hyder. The economic contribution of sea angling from charter boats: a case study from the south coast of England. Marine Policy, 119(??):Article 104066, September 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X1930795X.

Wang:2020:FNS

[WDG^{+}20] Dan Wang, Renke Ding, Yu Gong, Rui Wang, Jie Wang, and Xiaoling Huang. Feasibility of the Northern Sea Route for oil shipping from the economic and environmental perspective and its influence on China’s oil imports. Marine Policy, 118(??):Article 104006, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301792.

Watterson:2023:CSL

[WdSRY23] Christopher J. Watterson, Jean-Annet de Saint Rapt, and Eunha Yun. China’s state-led international port development: Challenges for port states. Marine Policy, 155(??):Article 105697, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002245.

Willette:2021:LMC

[WEF^{+}21] Demian A. Willette, Sofia C. Esteves, Ben Fitzpatrick, Marie L. Smith, Kesterlyn Wilson, and Xiaoya Yuan. The last mile challenge: Certified seafood and federal labeling laws out of sync at the end of the supply chain in Los Angeles, California. Marine Policy, 125(??):Article 104380, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20310319.

Wendebourg:2020:SOF

[Wen20] Mara R. Wendebourg. Southern Ocean fishery management — is CCAMLR addressing the challenges posed by a changing climate? Marine Policy, 118(??):Article 103847, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18308121.

Weitzman:2022:IKF

[WFG22a] Jenny Weitzman, Ramón Filgueira, and Jon Grant. Identifying key factors driving public opinion of salmon aquaculture. Marine Policy, 143(??):Article 105175, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002226.

Wichmann:2022:PPE

[WFG^{+}22b] Carla-Sophie Wichmann, Daniel Fischer, Sonja Maria Geiger, Daniela Honorato-Zimmer, Katrin Knickmeier, Katrin Kruse, Anna Sundermann, and Martin Thiel. Promoting pro-environmental behavior through citizen science? A case study with Chilean schoolchildren on marine plastic pollution. Marine Policy, 141(??):Article 105035, July 2022.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000823.

Wrobel:2022:TME

[WFG22c] Krzysztof Wróbel, Kamil Formela, and Mateusz Gil. Towards maritime employment and retention policy — a longitudinal study on entrants’ perspectives. Marine Policy, 144(??):Article 105232, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002792.

Wahler:2023:VBB

[WFR23] Hannah Wahler, Lucia Fanini, and Maraja Riechers. Valuing beaches for beauty and recreation only? Uncovering perception bias through a hashtag analysis. Marine Policy, 155(??):Article 105707, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002403.

Wright:2021:MSP

[WGJ^{+}21] Glen Wright, Kristina M. Gjerde, David E. Johnson, Aria Finkelstein, Maria Adelaide Ferreira, Daniel C. Dunn, Mariamalia Rodriguez Chaves, and Anthony Grehan. Marine spatial planning in areas beyond national jurisdiction. Marine Policy, 132(??):Article 103384, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304408.

Wongbusarakum:2021:MCA

[WGP^{+}21] Supin Wongbusarakum, Matt Gorstein, Robert Pomeroy, Cheryl L. Anderson, and Alexander Mawyer. Mobilizing for change: Assessing Social adaptive capacity in Micronesian fishing communities. Marine Policy, 129(??):Article 104508, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001196.

Wedding:2022:LMS

[WGR^{+}22] L. M. Wedding, S. J. Green, S. Reiter, K. R. Arrigo, L. Hazen, M. Ruckelshaus, J. M. A. van der Grient, R. M. Bailey, M. A. Cameron, J. Leape, M. Levi, A. Merkl, M. M. Mills, S. Monismith, N. T. Ouellette, G. van Dijken, and F. Micheli. Linking multiple stressor science to policy opportunities through network modeling. Marine Policy, 146(??):Article 105307, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003542.

Winter:2020:IFR

[WH20] Anna-Marie Winter and Jeffrey A. Hutchings. Impediments to fisheries recovery in Canada: Policy and institutional constraints on developing management practices compliant with the precautionary approach. Marine Policy, 121(??):Article 104161, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308358.

Wang:2022:ICP

[WHL22] Qiang Wang, Rui Huang, and Rongrong Li. Impact of the COVID-19 pandemic on research on marine plastic pollution — a bibliometric-based assessment. Marine Policy, 146(??):Article 105285, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003323.

Whomersley:2023:UWA

[Who23] Chris Whomersley. UNCLOS at 40: What about human rights? Marine Policy, 148(??):Article 105424, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004717.

Wang:2020:EBF

[WHPF20] Ying Wang, Jianfeng Hu, Haoran Pan, and Pierre Failler. Ecosystem-based fisheries management in the Pearl River Delta: Applying a computable general equilibrium model.

Marine Policy, 112(??):Article 103784, February 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306001.

Willaert:2020:CPD

[Wil20] Klaas Willaert. Crafting the perfect deep sea mining legislation: a patchwork of national laws. Marine Policy, 119(??):Article 104055, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308954.

Willaert:2021:DSM

[Wil21a] Klaas Willaert. Deep sea mining and the United States: Unbound powerhouse or odd man out? Marine Policy, 124(??):Article 104339, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309891.

Willaert:2021:ESA

[Wil21b] Klaas Willaert. The Enterprise: State of affairs, challenges and way forward. Marine Policy, 131(??):Article 104590, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002013.

Willaert:2022:TFD

[Wil22] Klaas Willaert. Transparency in the field of deep sea mining: Filtering the murky waters. Marine Policy, 135(??):Article 104840, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004516.

Waldo:2020:LAT

[WJB^{+}20] Åsa Waldo, Maria Johansson, Johan Blomquist, Torbjörn Jansson, Sara Königson, Sven-Gunnar Lunneryd, Anders Persson, and Staffan Waldo. Local attitudes towards management measures for the co-existence of seals and coastal fishery — a Swedish case study. Marine Policy, 118(??):Article 104018, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306098.

Watson:2021:CMP

[WJLSH21] Maryann S. Watson, Anne-Marie Jackson, Georgia Lloyd-Smith, and Christopher D. Hepburn. Comparing the Marine Protected Area Network Planning Process in British Columbia, Canada and New Zealand — Planning for cooperative partnerships with Indigenous communities. Marine Policy, 125(??):Article 104386, March 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2031037X.

Wilhelm:2020:PGH

[WKBS20] Miriam Wilhelm, Alin Kadfak, Vikram Bhakoo, and Kate Skattang. Private governance of human and labor rights in seafood supply chains — The case of the modern slavery crisis in Thailand. Marine Policy, 115(??):Article 103833, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19304403.

Walton:2022:CGT

[WKH22] Grant W. Walton, Meg Keen, and Quentin Hanich. Can greater transparency improve the sustainability of Pacific fisheries? Marine Policy, 136(??):Article 104251, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20308976.

Wakita:2022:PSP

[WKO^{+}22] Kazumi Wakita, Hisashi Kurokura, Zaida A. Ochavo, Reyda I. Inolino, Hiroshi Fushimi, and Satoshi Ishikawa. Potential signals promoting behavior for coastal conservation: Conformity in small-scale fishing communities in the Philippines. Marine Policy, 146(??):Article 105292, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003396.

Wang:2020:AIA

[WL20] Leifan Wang and Fenghua Li. Alternative interpretation to Article VIII of the International Convention for the Regulation of Whaling: Implications from Article XX of the GATT. Marine Policy, 117(??):Article 103944, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306463.

Wang:2022:SDC

[WL22a] Leifan Wang and Fenghua Li. State duty to cooperate on the Fukushima nuclear contaminated water release. Marine Policy, 136(??):Article 104878, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004899.

Wu:2022:STE

[WL22b] Jinghui Wu and Bo Li. Spatio-temporal evolutionary characteristics of carbon emissions and carbon sinks of marine industry in China and their time-dependent models. Marine Policy, 135(??):Article 104879, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004905.

Wilde:2023:ESD

[WLCC23] Daniel Wilde, Hannah Lily, Neil Craik, and Anindita Chakraborty. Equitable sharing of deep-sea mining benefits: More questions than answers. Marine Policy, 151(??):Article 105572, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000994.

Williams:2023:SPC

[WLCG23] Lacey Williams, Serena Lucrezi, Paul Cowley, and Enrico Gennari. Stakeholders’ perceptions of the conservation and management of elasmobranchs in South Africa. Marine Policy, 157(??):Article 105847, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003809.

Wan:2021:HDC

[WLQD21] Xiaole Wan, Qianqian Li, Lulian Qiu, and Yuanwei Du. How do carbon trading platform participation and government subsidy motivate blue carbon trading of marine ranching? A study based on evolutionary equilibrium strategy method. Marine Policy, 130(??):Article 104567, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001780.

Wilson:2020:STM

[WLRH^{+}20] Margaret W. Wilson, Julia M. Lawson, Maria Ignacia Rivera-Hechem, Juan Carlos Villaseñor-Derbez, and Steven D. Gaines. Status and trends of moored fish aggregating device (MFAD) fisheries in the Caribbean and Bermuda. Marine Policy, 121(??):Article 104148, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20304796.

Wu:2022:BMR

[WLW^{+}22] Haiping Wu, Qiuhui Li, Chunyan Wang, Qianwen Wu, Chongwei Peng, Thomas A. Jefferson, Zhizhen Long, Fangqiang Luo, Youhou Xu, and Shiang-Lin Huang. Bycatch mitigation requires livelihood solutions, not just fishing bans: a case study of the trammel-net fishery in the northern Beibu Gulf, China. Marine Policy, 139(??):Article 105018, May 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000653.

Wang:2021:PPE

[WLX21] Na Wang, Jing-Mei Li, and Zhi-Hua Xu. Public preference for the ecological restoration of coastal wetlands in Jiaozhou Bay in China based on a choice experiment. Marine Policy, 128(??):Article 104487, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000981.

Wu:2022:CET

[WLXY22] Min Wu, Kevin X. Li, Yi Xiao, and Kum Fai Yuen. Carbon Emission Trading Scheme in the shipping sector: Drivers, challenges, and impacts. Marine Policy, 138(??):Article 104989, April 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000367.

Wang:2021:FDP

[WLY21] Shuhong Wang, Binbin Lu, and Kedong Yin. Financial development, productivity, and high-quality development of the marine economy. Marine Policy, 130(??):Article 104553, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001640.

Wakamatsu:2021:ERS

[WM21] Hiroki Wakamatsu and Tsutom Miyata. Effects of radioactive safety information on consumer fears of radioactive contamination from oyster products in Japan. Marine Policy, 126(??):Article 104401, April 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000117.

Ward:2022:PUC

[WM22a] Sarah Ward and Mingfei Ma. Protection of the Underwater Cultural Heritage in Asia: Policy, law and practice. Marine Policy, 143(??):Article 105193, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002408.

Wuor:2022:DLF

[WM22b] Mehnwon Wuor and Leslie Mabon. Development of Liberia’s fisheries sectors: Current status and future needs. Marine Policy, 146(??):Article 105325, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003724.

Wiber:2021:RBE

[WMKG21] Melanie G. Wiber, Charles Mather, Christine Knott, and María Andrée López Gómez. Regulating the Blue Economy? Challenges to an effective Canadian aquaculture act. Marine Policy, 131(??):Article 104700, September 2021.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003110.

White:2022:BRN

[WMRB22] Casey M. White, Sangeeta Mangubhai, Lukas Rumetna, and Cassandra M. Brooks. The bridging role of nongovernmental organizations in the planning, adoption, and management of the marine protected area network in Raja Ampat, Indonesia. Marine Policy, 141(??):Article 105095, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001427.

Witter:2021:CSP

[WMS21] Allison Witter, Grant Murray, and U. Rashid Sumaila. Consumer seafood preferences related to alternative food networks and their value chains. Marine Policy, 131 (??):Article 104694, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003055.

Wootton:2022:PPP

[WNBRSG22] Nina Wootton, Melissa Nursey-Bray, Patrick Reis-Santos, and Bronwyn M. Gillanders. Perceptions of plastic pollution in a prominent fishery: Building strategies to inform management. Marine Policy, 135(??):Article 104846, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004577.

Wang:2023:BRM

[WNDZ23] Lei Wang, Wenzheng Niu, Lili Ding, and Zhongchao Zhao. A bibliometric review on marine economy: Current status, development and future directions. Marine Policy, 155(??):Article 105705, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002385.

Wakamatsu:2022:VWS

[WNM22] Mihoko Wakamatsu, Hiroki Nakamura, and Shunsuke Managi. The value of whaling and its spatial heterogeneity in Japan. Marine Policy, 135(??):Article 104852, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004632.

Willis-Norton:2024:SSS

[WNMS^{+}24] Ellen Willis-Norton, Tracey Mangin, Donna M. Schroeder, Reniel B. Cabral, and Steven D. Gaines. A synthesis of socioeconomic and sociocultural indicators for assessing the impacts of offshore renewable energy on fishery participants and fishing communities. Marine Policy, 161(??):Article 106013, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000113.

Watterson:2020:ORE

[WOG20] Christopher J. Watterson, Stephen Osborne, and Samuel Grant. Open registries as an enabler of maritime sanctions evasion. Marine Policy, 119(??):Article 104090, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20300385.

Wang:2023:CCS

[WP23] Liang Wang and Haoxin Peng. Collaboration or competition? A science map for promoting the sustainable transition of port relationships. Marine Policy, 155(??):Article 105795, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003287.

Ward-Paige:2022:FMM

[WPWM^{+}22] C. A. Ward-Paige, E. R. White, E. M. P. Madin, G. J. Osgood, L. K. Bailes, R. L. Bateman, E. Belonje, K. V. Burns, N. Cullain, P. Darbyshire-Jenkins, R. S. de Waegh, A. M. Eger, L. Fola-Matthews, B. M. Ford, C. Gonson, C. J. Honeyman, J. E. House, E. Jacobs, L. K. Jordan, J. J. Levenson, K. Lucchini, Marti-PuigP Martí-Puig, L. A. H. McGuire, C. Meneses, P. H. Montoya-Maya, R. A. Noonan, P. A. Ruiz-Ruiz, P. E. Ruy, R. A. Saputra, G. Shedrawi, B. Sing, M. D. Tietbohl, A. Twomey, D. C. Vergara Florez, and L. Yamb. A framework for mapping and monitoring human-ocean interactions in near real-time during COVID-19 and beyond. Marine Policy, 140(??):Article 105054, June 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001014.

Wang:2023:AIG

[WRL23] Qiang Wang, Feng Ren, and Rongrong Li. Assessing the impact of geopolitics on international scientific cooperation — the case of us–china marine pollution research. Marine Policy, 155(??):Article 105723, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002567.

Watzel:2021:MMR

[WRV21] Ralph Watzel, Carsten Rühlemann, and Annemiek Vink. Mining mineral resources from the seabed: Opportunities and challenges. Marine Policy, ??(??):Article 103828, ???? 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20300622.

Wakamatsu:2021:CJF

[WS21] Hiroki Wakamatsu and Yutaro Sakai. Can the Japanese fisheries qualify for MSC certification? Marine Policy, 134(??):Article 104750, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003614.

Williams:2023:FIP

[WS23] Chris Williams and Jessica L. Decker Sparks. Fishery improvement projects: a voluntary, corporate “tool” not fit for the purpose of mitigating labour abuses and guaranteeing labour rights for workers. Marine Policy, 147(??):Article 105340, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003876.

Wang:2021:ERA

[WSD21] Juying Wang, Xiao Shi, and Yuanwei Du. Exploring the relationship among marine science and technology innovation, marine finance, and marine higher education using a coupling analysis: a case study of China’s coastal areas. Marine Policy, 132(??):Article 104706, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003171.

Woker:2020:LSC

[WSDB20] Hilde Woker, Bernhard Schartmüller, Knut Ola Dølven, and Katalin Blix. The law of the sea and current practices of marine scientific research in the Arctic. Marine Policy, 115(??):Article 103850, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304130.

Williamson:2021:CSI

[WtDD^{+}21] Michael J. Williamson, Mariel T. I. ten Doeschate, Rob Deaville, Andrew C. Brownlow, and Nicola L. Taylor. Cetaceans as sentinels for informing climate change policy in UK waters. Marine Policy, 131(??):Article 104634, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002451.

Wu:2022:STR

[Wu22] Hsing-Hao Wu. A study on transnational regulatory governance for marine plastic debris: Trends, challenges, and prospect. Marine Policy, 136(??):Article 103988, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20301822.

Wagner:2021:SGN

[WvdMG^{+}21] Daniel Wagner, Liesbeth van der Meer, Matthias Gorny, Javier Sellanes, Carlos F. Gaymer, Eulogio H. Soto, Erin E. Easton, Alan M. Friedlander, Dhugal J. Lindsay, Tina N. Molodtsova, Ben Boteler, Carole Durussel, Kristina M. Gjerde, Duncan Currie, Matthew Gianni, Cassandra M. Brooks, Marianne J. Shiple, T. ‘Aulani Wilhelm, Marco Quesada, Tamara Thomas, Piers K. Dunstan, Nichola A. Clark, Luis A. Villanueva, Richard L. Pyle, Malcolm R. Clark, Samuel E. Georgian, and Lance E.

Morgan. The Salas y Gómez and Nazca ridges: a review of the importance, opportunities and challenges for protecting a global diversity hotspot on the high seas. Marine Policy, 126(??):Article 104377, April 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310289.

Walters:2021:DEA

[WW21a] David Walters and Emma Wadsworth. Determinants of effective action on workplace safety and health in global companies — The case of global network container terminal operators. Marine Policy, 124(??):Article 104374, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20310253.

Wang:2021:MFP

[WW21b] Sun Ling Wang and John B. Walden. Measuring fishery productivity growth in the Northeastern United States 2007– 2018. Marine Policy, 128(??):Article 104467, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000786.

Wang:2022:CEE

[WW22] Qiang Wang and Shasha Wang. Carbon emission and economic output of China’s marine fishery — a decoupling efforts analysis. Marine Policy, 135(??):Article 104831, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004425.

Wu:2023:ALH

[WWC23] Ling-I Wu, Chia-Hui Wang, and Chih-Shin Chen. Applying life history traits to eco-labelling schemes for fisheries management: a case study in Taiwan. Marine Policy, 148(??):Article 105412, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004596.

Wan:2021:SSR

[WWCS21] Zheng Wan, Likun Wang, Jihong Chen, and Daniel Sperling. Ship scrappage records reveal disturbing environmental injustice. Marine Policy, 130(??):Article 104542, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001536.

Waters:2023:PAD

[WWD23] Yolanda L. Waters, Kerrie A. Wilson, and Angela J. Dean. Plastic action or distraction? Marine plastic campaigns influence public engagement with climate change in both general and engaged audiences. Marine Policy, 152(??):Article 105580, June 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001070.

Ward:2023:LSR

[WWG^{+}23] Tim M. Ward, Barrett W. Wolfe, Gretchen L. Grammer, Alex R. Ivey, Edward King, Andreas Schiller, Karlie S. McDonald, and Jeffrey M. Dambacher. Large sardine resource discovered off south-eastern Australia: Potential risks, challenges and benefits of establishing a new fishery. Marine Policy, 155(??):Article 105739, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002725.

Wan:2021:EPT

[WXLD21] Xiaole Wan, Shuwen Xiao, Qianqian Li, and Yuanwei Du. Evolutionary policy of trading of blue carbon produced by marine ranching with media participation and government supervision. Marine Policy, 124(??):Article 104302, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309477.

Wang:2022:HDE

[WYD^{+}22] Han Wang, YanChun Yao, XiaoLing Dai, ZhouPeng Chen, Jianxian Wu, GuangLong Qiu, and Ting Feng. How do ecological protection policies affect the restriction of coastal development rights? Analysis of choice preference based on choice experiment. Marine Policy, 136(??):Article 104905, February 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21005169.

Walther:2021:SAP

[WYH21] Bruno Andreas Walther, Ning Yen, and Chieh-Shen Hu. Strategies, actions, and policies by Taiwan’s ENGOs, media, and government to reduce plastic use and marine plastic pollution. Marine Policy, 126(??):Article 104391, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21000014.

Wan:2021:ISE

[WZZC21] Zheng Wan, Xujun Zhou, Qiang Zhang, and Jihong Chen. Influence of sulfur emission control areas on particulate matter emission: a difference-in-differences analysis. Marine Policy, 130(??):Article 104584, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001950.

Xu:2022:REE

[XHY22] Jie Xu, Limin Han, and Wei Yin. Research on the ecologicalization efficiency of mariculture industry in China and its influencing factors. Marine Policy, 137(??):Article 104935, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21005467.

Xu:2021:RCM

[XHZ^{+}21] Xin Xu, Yingying Hou, Changping Zhao, Lei Shi, and Yu Gong. Research on cooperation mechanism of marine plastic waste management based on complex network evolutionary game. Marine Policy, 134(??):Article 104774, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003857.

Xu:2021:ECR

[XLLS21] Zhihua Xu, Shuqin Li, Jingmei Li, and Jingzhu Shan. The effect of consequentiality on the residents’ willingness to pay for the governance of Ulva prolifera bloom in Qingdao, China. Marine Policy, 126(??):Article 104404, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000142.

Xu:2021:EEC

[XSCL21] Lang Xu, Jia Shi, Jihong Chen, and Li Li. Estimating the effect of COVID-19 epidemic on shipping trade: an empirical analysis using panel data. Marine Policy, 133(??):Article 104768, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003791.

Xu:2023:BMC

[XT23] Qi Xu and Ziyue Tan. Building a maritime community with a shared future: Scholarly reflections on China’s new ocean vision. Marine Policy, 149(??):Article 105508, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000350.

Xixi:2022:OSS

[XTY22] Zhang Xixi, Qu Tongkun, and Wang Yecheng. Optimal strategies for stakeholders of Fukushima nuclear waste water discharge in Japan. Marine Policy, 135(??):Article 104881, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X21004929.

Xu:2022:MCR

[XWH^{+}22] Nan Xu, Yuqing Wang, Conghong Huang, Shuai Jiang, Mingming Jia, and Yue Ma. Monitoring coastal reclamation changes across Jiangsu Province during 1984–2019 using Landsat data. Marine Policy, 136(??):Article 104887, February 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2100498X.

Xiao:2020:ENI

[XWL^{+}20] Yi Xiao, Grace Wang, Kun-Chin Lin, Guanqiu Qi, and Kevin X. Li. The effectiveness of the New Inspection Regime for Port State Control: Application of the Tokyo MoU. Marine Policy, 115(??):Article 103857, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302659.

Xu:2024:IFR

[XXLW24] Zhihua Xu, Juan Xu, Shuqin Li, and Chuwei Wang. The influencing factors of residents’ willingness to pay in marine ecological restoration: the integration of the theory of planned behavior and social capital theory. Marine Policy, 161(??):Article 106031, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000290.

Xu:2023:URP

[XXMZ23] Zhihua Xu, Juan Xu, Wenchao Meng, and Gang Zhou. Understanding residents’ policy preferences to construct effective marine debris governance policies. Marine Policy, 158(??):Article 105872, December 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004050.

Xing:2023:ELG

[XZ23] Wangwang Xing and Ling Zhu. Exploring legal gaps and barriers to the use of unmanned merchant ships in China. Marine Policy, 153(??):Article 105662, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001896.

Yuen:2020:SBS

[YBW20] Kum Fai Yuen, Xiwen Bai, and Xueqin Wang. Safety behaviour at sea: Policy implications for managing seafarers through positive psychology. Marine Policy, 121 (??):Article 104163, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20306953.

Yuan:2021:MCO

[YC21a] Wei Yuan and Yen-Chiang Chang. Maritime cooperation and ocean governance: Conference report. Marine Policy, 125(??):Article 104388, March 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310393.

Yunitawati:2021:GES

[YC21b] Dinah Yunitawati and Julian Clifton. Governance in the early stages of marine protected area development: a case study of Nusa Penida District Marine Conservation Area, Indonesia. Marine Policy, 127(??):Article 103653, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19305172.

Yuen:2021:PCS

[YCBW21] Kum Fai Yuen, Yangyi Cao, Xiwen Bai, and Xueqin Wang. The psychology of cruise service usage post COVID-19: Health management and policy implications. Marine Policy, 130(??):Article 104586, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001974.

Yang:2022:SSF

[YCL22a] Tsung-Yu Yang, Tsun-Feng Chiang, and Wen-Hong Liu. Small-scale fishers’ catch production in Taiwanese coastal areas. Marine Policy, 143(??):Article 105182, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002299.

Yu:2022:EGR

[YCL22b] Derek Chin-Yun Yu, Wen-Yan Chiau, and Hsueh-Jung Lu. Effective governance of a remote marine protected area: the case of Dongsha Atoll National Park. Marine Policy, 139(??):Article 105013, May 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000604.

Yu:2022:LPM

[YD22] Xiao Yu and Yichong Dong. Local practice of marine protected areas legislation in China: the case of Zhoushan. Marine Policy, 141(??):Article 105084, July 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001312.

Yadav:2021:SFC

[YFD^{+}21] Shreya Yadav, Abdulla Fisam, Rachel Dacks, Joshua S. Madin, and Alexander Mawyer. Shifting fish consumption preferences can impact coral reef resilience in the Maldives: a case study. Marine Policy, 134(??):Article 104773, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003845.

Yadav:2020:OCC

[YG20] Siddharth Shekhar Yadav and Kristina Maria Gjerde. The ocean, climate change and resilience: Making ocean areas beyond national jurisdiction more resilient to climate change and other anthropogenic activities. Marine Policy, 122(??):Article 104184, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302074.

Yu:2020:FSM

[YH20] Jinkai Yu and Qingchao Han. Food security of mariculture in China: Evolution, future potential and policy. Marine Policy, 115(??):Article 103892, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930675X.

Yang:2022:SAF

[YHU22] Xue Yang, Ingunn Marie Holmen, and Ingrid Bouwer Utne. Scenario analysis of fish escapes in Norwegian aquaculture for implementation of barrier management and improved learning from accidents. Marine Policy, 143(??):Article 105208, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200255X.

Yi:2020:POA

[YK20] Sangchoul Yi and Wonkook Kim. Potential for offshore aquaculture development in North Korea: Focusing on Atlantic salmon farming. Marine Policy, 119 (??):Article 104092, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20302323.

Young:2021:NSA

[YK21] Oran R. Young and Jong-Deog Kim. Next steps in Arctic Ocean Governance Meeting the challenge of coordinating a dynamic regime complex. Marine Policy, 133(??):Article 104726, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003377.

Yen:2021:RML

[YL21] Kuo-Wei Yen and Li-Chuan Liuhuang. A review of migrant labour rights protection in distant water fishing in Taiwan: From laissez-faire to regulation and challenges behind. Marine Policy, 134(??):Article 104805, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004164.

Yu:2023:TTS

[YLSC23] Xuan Yu, Sen Lin, Weiteng Shen, and Qi Chen. The tiered transaction system for sea use rights in China: Theoretical basis and market construction. Marine Policy, 147(??):Article 105403, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200450X.

Yanda:2023:CCL

[YMG23] Pius Z. Yanda, Anselm R. Mwajombe, and Brown Gwambene. Coastal communities’ livelihoods conditioned by resource use dynamics and changing environments in Lindi Region, South-Eastern Tanzania. Marine Policy, 150(??):Article 105532, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000593.

Yu:2021:LMP

[YML^{+}21] Lance Yu, Linna Ma, Vicky Lam, Xiaomu Guan, Yujie Zhao, Sheng Wang, Yongtong Mu, and Rashid Sumaila. Local marine policy whacking the national Zhikong scallop fishery. Marine Policy, 124(??):Article 104352, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20310022.

Yosef:2022:API

[YNZ22] Reuven Yosef, Tal Nachshonov, and Piotr Zduniak. Anthropopause positively influenced Red Sea Clownfish (Amphiprion bicinctus) populations but not the host sea anemone (Actiniaria spp.) in Eilat, Israel. Marine Policy, 145(??):Article 105280, November 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200327X.

Yea:2021:CVC

[YS21] Sallie Yea and Christina Stringer. Caught in a vicious cycle: Connecting forced labour and environmental exploitation through a case study of Asia–Pacific. Marine Policy, 134(??):Article 104825, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100436X.

Yea:2023:IPW

[YS23] Sallie Yea and Christina Stringer. The informalisation of precarious work in fishing crew: Experiences of Fijian fishers on distant water vessels. Marine Policy, 155(??):Article 105709, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002427.

Yosef:2022:ABA

[YSEK22] Reuven Yosef, Veronica Spivak, Shalev Edelman, and Jakub Z. Kosicki. Anthropogenic beach-activities influences the endemic Red Sea Ghost Crab (Ocypode saratan). Marine Policy, 138(??):Article 104986, April 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000331.

Yi:2021:CED

[YSWW21] Xiaojing Yi, Kun Sheng, Yuanyue Wang, and Shuhong Wang. Can economic development alleviate storm surge disaster losses in coastal areas of China? Marine Policy, 129(??):Article 104531, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001421.

Yin:2020:CAM

[YT20] Mengmeng Yin and Erika J. Techera. A critical analysis of marine protected area legislation across state and territory jurisdictions in Australia. Marine Policy, 118(??):Article 104019, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19302581.

Yandle:2022:AVC

[YTBM^{+}22] Tracy Yandle, Jennifer Sweeney Tookes, Paulita Bennett-Martin, Sherry Larkin, and Michael Page. An assessment of the viability of a commercial lionfish fishery in the US Virgin Islands. Marine Policy, 143(??):Article 105103, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001506.

Yildirim:2022:CEW

[YTÇ22] Umut Yildirim, Arda Toygar, and Cavit Çolakoğlu. Compensation effect of wages on decent work: a study on seafarers attitudes. Marine Policy, 143(??):Article 105155, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002020.

Yu:2023:VCD

[YTG^{+}23] Lance Yu, Xiaolan Tan, Xiaomu Guan, Yongtong Mu, Vicky Lam, and Rashid Sumaila. Value chain of the data-poor Chinese bay scallop aquaculture. Marine Policy, 150(??):Article 105556, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000830.

Yildiz:2020:AGP

[YU20] Taner Yildiz and Aylin Ulman. Analyzing gaps in policy: Evaluation of the effectiveness of minimum landing size (MLS) regulations in Turkey. Marine Policy, 115(??):Article 103829, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19301587.

Yu:2022:OOD

[Yu22] Chuxiao Yu. Operational oceanography as a distinct activity from marine scientific research under UNCLOS?

— An analysis of WMO Resolution 45 (Cg-18). Marine Policy, 143(??):Article 105131, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001786.

Yu:2020:EGO

[YW20] Jin-Kai Yu and Yan-Ling Wang. Exploring the goals and objectives of policies for marine ranching management: Performance and prospects for China. Marine Policy, 122(??):Article 104255, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309015.

Yang:2020:MBS

[YWY20] Lijing Yang, LingLing Wu, and Shuqing Yan. Market-based sea use management in China: Features and countermeasures. Marine Policy, 120(??):Article 104091, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309182.

Yu:2022:PNA

[YWYL22] Huijuan Yu, Yezhi Wang, Wei Yin, and Dahai Li. Policy network analysis of China’s ocean ranching policy: Network structure, actors and interaction. Marine Policy, 140(??):Article 105070, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001178.

Yu:2023:CEH

[YYD23] Wencheng Yu, Miao Yu, and Lili Ding. Characteristics and evolution of hierarchical fishery policies in China — a textual analysis based on 5311 policies from 2003 to 2022. Marine Policy, 155(??):Article 105699, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002269.

Yu:2020:EMP

[YYL20] Jinkai Yu, Wang Yin, and Dahai Liu. Evolution of mariculture policies in China: Experience and challenge. Marine Policy, 119(??):Article 104062, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304798.

Yu:2023:GER

[YZG23] Lianghong Yu, Shan Zheng, and Qiang Gao. Government environmental regulation strategy for new pollutants control in mariculture. Marine Policy, 150(??):Article 105545, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000726.

Yin:2022:ISM

[YZH22] Kedong Yin, Kai Zhang, and Chong Huang. Institutional supply, market cultivation, and the development of marine strategic emerging industries. Marine Policy, 139(??):Article 105045, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000926.

Zuercher:2022:ECE

[ZBF^{+}22] Rachel Zuercher, Natalie C. Ban, Wesley Flannery, Anne D. Guerry, Benjamin S. Halpern, Rafael Almeida Magris, Shauna L. Mahajan, Nicole Motzer, Ana K. Spalding, Vanessa Stelzenmüller, and Jonathan G. Kramer. Enabling conditions for effective marine spatial planning. Marine Policy, 143(??):Article 105141, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001889.

Zimmerhackel:2023:FIA

[ZCA^{+}23] J. S. Zimmerhackel, J. Clifton, F. Ackermann, M. P. Burton, C. E. Elrick-Barr, G. Hill, and E. S. Harvey. A framework for the integrated assessment of social and economic values associated with man-made marine structures. Marine Policy, 152(??):Article 105612, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001392.

Zhao:2022:SCR

[ZCCQ22a] Qilei Zhao, Xinjun Chen, Guiying Chen, and Siqiong Qi. The socioeconomic contribution of the recreational fishery based on input–output analysis: the case of China. Marine Policy, 143(??):Article 105177, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200224X.

Zhao:2022:UDR

[ZCCQ22b] Qilei Zhao, Xinjun Chen, Guiying Chen, and Siqiong Qi. Understanding the development of recreational fisheries with an industry competitiveness framework: Evaluation of a case study from China. Marine Policy, 143(??):Article 105121, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001683.

Zhao:2023:MCE

[ZCG^{+}23] Qiqi Zhao, Yan Chen, Keshava Pallavi Gone, Emily Wells, Keahna Margeson, and Kate Sherren. Modelling cultural ecosystem services in agricultural dykelands and tidal wetlands to inform coastal infrastructure decisions: a social media data approach. Marine Policy, 150(??):Article 105533, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2300060X.

Zhou:2023:ICC

[ZCS^{+}23] Jingen Zhou, Shu-Ling Peggy Chen, Wenming Wendy Shi, Maneerat Kanrak, and Jiawei Ge. The impacts of COVID-19 on the cruise industry based on an empirical study in China. Marine Policy, 153(??):Article 105631, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001586.

Zhou:2023:CSC

[ZCSK23] Jingen Zhou, Shu-Ling (Peggy) Chen, Wenming (Wendy) Shi, and Maneerat Kanrak. Cruise supply chain risk mitigation strategies: an empirical study in Shanghai, China. Marine Policy, 153(??):Article 105600, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001276.

Zhang:2020:OCS

[ZCZ20] Weibin Zhang, Yen-Chiang Chang, and Liangfu Zhang. An ocean community with a shared future: Conference report. Marine Policy, 116(??):Article 103888, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301081.

Zhang:2020:FSE

[ZD20a] Pengfei Zhang and Lukas Drumm. The flagging-out strategy: an examination of the impacts on the decreasing German national fleet. Marine Policy, 115(??):Article 103872, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19309832.

Zhang:2020:GSF

[ZD20b] Pengfei Zhang and Lukas Drumm. The German Shipping Foundation: Has it been effective in maintaining maritime expertise in Germany? Marine Policy, 115(??):Article 103871, May 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20300117.

Zhao:2020:EFB

[ZFGL20] Changping Zhao, Chao Fang, Yu Gong, and Zudi Lu. The economic feasibility of Blue Carbon cooperation in the South China Sea region. Marine Policy, 113(??):Article 103788, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19304725.

Zhou:2021:MFR

[ZGY21] Qin Zhou, Ruobin Gao, and Kum Fai Yuen. Marine fuel refining technology improvement trade-offs: a game theoretic approach. Marine Policy, 132(??):Article 104677, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002888.

Zhang:2021:UTC

[Zha21] Hongzhou Zhang. Understanding the tragedy of the commons in the South China Sea fisheries: the prisoner’s dilemma model revisited. Marine Policy, 125(??): Article 104376, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X20310277.

Zhao:2022:ICN

[Zha22] Na Zhao. The impact of China’s new-type urbanization on individual sustainable development: a comparison between coastal and inland areas. Marine Policy, 136(??):Article 104938, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21005492.

Zhao:2023:RCN

[ZHL^{+}23] Changping Zhao, Yingying Hou, Mengru Liu, Yu Gong, and Junbin Wang. Research on the cooperative network game model of marine plastic waste management. Marine Policy, 149(??):Article 105504, March 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000313.

Zhou:2023:PVE

[Zho23] Hang Zhou. Provincial variations and entrepreneurialism in the development of China’s Distant Water Fisheries (2011–2020). Marine Policy, 147(??):Article 105344, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003918.

Zhu:2022:QBS

[Zhu22] Ling Zhu. A questionnaire based study on the status quo and future action over the regulation of marine greenhouse gas emissions in Hong Kong. Marine Policy, 145(??):Article 105278, November 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003256.

Zhu:2023:STA

[Zhu23] Ling Zhu. Some thoughts on application of the polluter pays principle for controlling marine greenhouse gas emissions. Marine Policy, 158(??):Article 105877, December 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004104.

Zhao:2020:TSS

[ZJ20] Xin Zhao and Peihong Jia. Towards sustainable smallscale fisheries in China: a case study of Hainan. Marine Policy, 121(??):Article 103935, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305998.

Zaucha:2022:EMS

[ZJ22] Jacek Zaucha and Stephen Jay. The extension of marine spatial planning to the management of the world ocean, especially areas beyond national jurisdiction. Marine Policy, 144(??):Article 105218, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002652.

Zaucha:2021:ESM

[ZK21] Jacek Zaucha and Anja Kreiner. Engagement of stakeholders in the marine/maritime spatial planning process. Marine Policy, 132(??):Article 103394, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304081.

Zhang:2020:FGS

[ZL20] Hongzhou Zhang and Mingjiang Li. Fisheries governance in the south China sea: Problems, progress, and prospects. Marine Policy, 121(??):Article 104237, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308836.

Zhu:2021:TTA

[ZLD^{+}21] Wenbin Zhu, Zhanhui Lu, Qian Dai, Kaner Lu, Zhenhua Li, Yongdong Zhou, Yazhou Zhang, Ming Sun, Yunzhou Li, and Wei Li. Transition to timely and accurate reporting: an evaluation of monitoring programs for China’s first Total Allowable Catch (TAC) pilot fishery. Marine Policy, 129(??):Article 104503, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001147.

Zhao:2020:SAQ

[ZLGX20] Changping Zhao, Rui Li, Yu Gong, and Xin Xu. Study on the asymmetric QRE network game simulation of the South China Sea route trade cooperation. Marine Policy, 111(??):Article 103724, January 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19300302.

Zhu:2021:SAR

[ZLH21] Wenchao Zhu, Bo Li, and Zenglin Han. Synergistic analysis of the resilience and efficiency of China’s marine economy and the role of resilience policy. Marine Policy, 132(??):Article 104703, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003146.

Zhang:2023:MSP

[ZLPT23] Yiqi Zhang, Christine Loh, Gerald R. Patchell, and Kellee S. Tsai. Multi-scale policy diffusion of marine emissions governance. Marine Policy, 153(??):Article 105637, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001641.

Zhou:2022:ICS

[ZLWY22] Yusheng Zhou, Xue Li, Xueqin Wang, and Kum Fai Yuen. Intelligent container shipping sustainability disclosure via stakeholder sentiment views on social media. Marine Policy, 135(??):Article 104853, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004644.

Zhou:2023:SSC

[ZLY23] Yusheng Zhou, Xue Li, and Kum Fai Yuen. Sustainable shipping: a critical review for a unified framework and future research agenda. Marine Policy, 148 (??):Article 105478, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000052.

Zheng:2021:HDF

[ZLZ21] Hui Zheng, Jie Li, and Xin Zhao. How does financial policy support the development of China’s fishery? Characteristics, experience and prospects. Marine Policy, 132(??):Article 104678, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100289X.

Zudaire:2023:BDF

[ZMM^{+}23] Iker Zudaire, Gala Moreno, Jefferson Murua, Paul Hamer, Hilario Murua, Mariana T. Tolotti, Marlon Roman, Martin Hall, Jon Lopez, Maitane Grande, Gorka Merino, Lauriane Escalle, Oihane C. Basurko, Manuela Capello, Laurent Dagorn, Maria Lourdes Ramos, Francisco J. Abascal, José Carlos Báez, Pedro J. Pascual-Alayón, Santiago Déniz, and Josu Santiago. Biodegradable drifting fish aggregating devices: Current status and future prospects. Marine Policy, 153(??):Article 105659, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001860.

Zambrano-Monserrate:2020:EDC

[ZMR20] Manuel A. Zambrano-Monserrate and Maria Alejandra Ruano. Estimating the damage cost of plastic waste in Galapagos Islands: a contingent valuation approach. Marine Policy, 117(??):Article 103933, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309820.

Zhang:2023:CMS

[ZMZ23] Shiyang Zhang, Zhengkai Mao, and Zhijun Zhang. China’s maritime strategy in the new vision. Marine Policy, 155(??):Article 105769, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003020.

Zhu:2023:SEI

[ZNA^{+}23] Shengda Zhu, Adolf K. Y. Ng, Mawuli Afenyo, Roozbeh Panahi, and Michael G. H. Bell. Socio-economic impacts of shipping along the Northwest Passage: the cost to locals. Marine Policy, 153(??):Article 105647, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001744.

Zou:2023:CIR

[Zou23] Yizheng Zou. China and Indonesia’s responses to maritime disputes in the South China Sea: forming a tacit understanding on security. Marine Policy, 149(??):Article 105502, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000295.

Zhao:2022:EME

[ZPX^{+}22] Yiyi Zhao, Ellen K. Pikitch, Xiangxin Xu, Timothy Frankstone, John Bohorquez, Xu Fang, Jie Zheng, Yunyi Li, Zhu Chen, Wuying Lin, Wenjia Hu, and Guifang (Julia) Xue. An evaluation of management effectiveness of China’s marine protected areas and implications of the 2018 Reform. Marine Policy, 139(??):Article 105040, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000872.

Zhou:2020:KCC

[ZSLY20] Yusheng Zhou, Ying Shan Soh, Hui Shan Loh, and Kum Fai Yuen. The key challenges and critical success factors of blockchain implementation: Policy implications for Singapore’s maritime industry. Marine Policy, 122(??):Article 104265, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309118.

Zuidema:2023:PBS

[ZSWJ23] Jason Zuidema, Desai Shan, Kevin Walker, and Alireza Jahanbakhsh. Port-based seafarers’ welfare in Canada: resources and gaps. Marine Policy, 147(??):Article 105336, January 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003839.

Zheng:2020:CPF

[ZSZ20] Hui Zheng, Min Shang, and Xin Zhao. Chinese policy on fishery insurance: Evolution, characteristics and challenges. Marine Policy, 119(??):Article 104099, September 2020.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20300026.

Zheng:2021:COT

[ZT21] Li Zheng and Kailan Tian. The contribution of ocean trade to national economic growth: a non-competitive input-output analysis in China. Marine Policy, 130(??): Article 104559, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X21001706.

Zhao:2023:ICM

[ZT23a] Zhiwei Zhao and Lijun Tang. The impact of COVID-19 on maritime pilots: Evidence and lessons. Marine Policy, 153(??):Article 105664, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001914.

Zou:2023:UCD

[ZT23b] Keyuan Zou and Anastasia Telesetsky. UNCLOS and its contributions to the development of international law: an editorial note. Marine Policy, 155(??):Article 105703, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002361.

Zhu:2023:VAT

[ZTA23] Mengyao Zhu, Kotaro Tanaka, and Tomonari Akamatsu. Visualizing the annual transition of ocean policy in Japan using text mining. Marine Policy, 155(??):Article 105754, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002877.

Zhao:2023:FSD

[ZTM^{+}23] Zhiwei Zhao, Lijun Tang, Yunlei Ma, Yueyan Wu, Shiqi Lin, Zefan Wu, Zheng Zhou, Xinyi Wang, and Xinjian Wang. Fatigue at sea during and after the COVID-19 pandemic: a comparative study of two matched samples of seafarers. Marine Policy, 155(??):Article 105730, September 2023.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002634.

Zhao:2023:FDC

[ZTW23] Zhiwei Zhao, Lijun Tang, and Yueyan Wu. Fatigue during the COVID-19 pandemic: the experiences of Chinese seafarers. Marine Policy, 153(??):Article 105643, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001707.

Zhang:2021:ICP

[ZTZ^{+}21] Yanxuedan Zhang, Yi Tang, Yu Zhang, Yuqiong Sun, and Haoran Yang. Impacts of the COVID-19 pandemic on fish trade and the coping strategies: an initial assessment from China’s perspective. Marine Policy, 133(??):Article 104748, November 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003596.

Zhang:2020:CPB

[ZV20] Xiong Zhang and Amanda C. J. Vincent. China’s policies on bottom trawl fisheries over seven decades (1949– 2018). Marine Policy, 122(??):Article 104256, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309027.

Zhang:2021:IMI

[ZW21] Ying Zhang and Shuhong Wang. Influence of marine industrial agglomeration and environmental regulation on marine innovation efficiency — From an innovation value chain perspective. Marine Policy, 134(??):Article 104807, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004188.

Zheng:2023:PEA

[ZWC^{+}23] Hui Zheng, Chong Wang, Jingya Chen, Xuan Liu, and Xin Zhao. Policy evolution and analysis of characteristics and values of China’s marine seed industry development under the food security strategy. Marine Policy, 155(??):Article 105681, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002087.

Zhao:2021:EPN

[ZWG^{+}21] Changping Zhao, Yecheng Wang, Yu Gong, Steve Brown, and Rui Li. The evolution of the port network along the Maritime Silk Road: From a sustainable development perspective. Marine Policy, 126(??):Article 104426, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000361.

Zhao:2020:CPF

[ZWJvL20] Zhiwei Zhao, Emma Wadsworth, Jørgen Riis Jepsen, and Wessel M. A. van Leeuwen. Comparison of perceived fatigue levels of seafarers and management approaches in fatigue mitigation: Case studies from two Chinese and two European shipping companies. Marine Policy, 116(??):Article 103897, June 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19307638.

Zheng:2021:RJT

[ZWXL21] Sining Zheng, Shuqin Wang, Wei Xu, and Qiang Liu. Research on the job transfer pathway of Chinese marine fishermen and its driving factors. Marine Policy, 129(??):Article 104572, July 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001834.

Zhao:2021:PSF

[ZXGL21] Changping Zhao, Xiya Xie, Yu Gong, and Beili Liu. The propagation of sustainable fishery by Arctic shipping route stakeholders. Marine Policy, 131(??):Article 104619, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2100230X.

Zhou:2021:LSF

[ZY21] Qin Zhou and Kum Fai Yuen. Low-sulfur fuel consumption: Marine policy implications based on game theory.

Marine Policy, 124(??):Article 104304, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309490.

Zou:2023:RBU

[ZY23] Keyuan Zou and Qiang Ye. The relationship between UNCLOS and Customary International Law: Some reflections. Marine Policy, 154(??):Article 105691, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300218X.

Zhou:2021:EMK

[ZYTT21a] Yusheng Zhou, Kum Fai Yuen, Beverly Tan, and Vinh V. Thai. The effect of maritime knowledge clusters on maritime firms’ performance: an organizational learning perspective. Marine Policy, 128(??):Article 104472, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100083X.

Zhou:2021:MKC

[ZYTT21b] Yusheng Zhou, Kum Fai Yuen, Beverly Tan, and Vinh V. Thai. Maritime knowledge clusters: a conceptual model and empirical evidence. Marine Policy, 123(??):Article 104299, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309441.

Zheng:2023:PDA

[ZZD^{+}23] Li Zheng, Hongyang Zou, Xiaofeng Duan, Zhongguo Lin, and Huibin Du. Potential determinants affecting the growth of China’s ocean economy: an input-output structural decomposition analysis. Marine Policy, 150(??):Article 105520, April 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000477.

Zhao:2021:PBW

[ZZH21] Minghua Zhao, Pengfei Zhang, and Gaochao He. Portbased welfare services for seafarers in Chinese ports: Their roles, changes and challenges. Marine Policy, 130(??):Article 104190, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20300701.

Zheng:2021:ACP

[ZZW^{+}21] Hui Zheng, Li Zhang, Shuyi Wang, Juan Xu, and Xin Zhao. The affecting channels and performances of financial development and poverty reduction: New evidence from China’s fishery industry. Marine Policy, 123(??):Article 104324, January 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20309714.